A Complete Bibliography of Publications in Journal for the History of Astronomy

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

01 April 2025 Version 1.37

Title word cross-reference

$100 [Apt14]. $120 [Hen15b]. $120.84 [Kra22]. $127 [Llo15]. $135.00 [Smi96]. $14 [Sch15]. $140 [GG14]. $15 [Jar90]. $17.95 [Had84]. $19.95 [Mul83, Nau98]. 20 [Ost07]. 2000 [Eva09b]. $24.95 [Lep14, Bro90]. $27 [Wło15]. $29 [Sha14]. $29.25 [Hea15]. $29.95 [Ger17, Mes15]. 30 [Mau13]. $31.00 [Bra15]. $34 [Sul14]. $35 [Ano15f, Dev14b, Lau14, Mir17, Rap15a]. $38 [Dan19b, Vet19]. $39.95 [Mol14b]. $39.99 [Bec15]. $40 [Dun20, LF15, Rob86]. 40 [Edd71b]. $42 [Nau98]. $45 [Kes15]. $49.95 [Ree20]. $49.99 [Bec15]. $50 [Kru17, Rem15]. $55 [Bon19b, Sil23]. $60 [Cro23a, Mal15]. 600 [GB95]. $72 [Ave18]. $79 [Wil15]. $8.95 [Had84]. $87 [CWW17].

$89.95 [Gan15]. 8 = ψ^{1,2,3} [Cov15]. $90 [Ano15f]. $95 [Swe17c]. ^{c} [Kin87, NRKN16, Rag05]. ^{mul} [Kur19]. ^{muld} [Kur19]. _{3} [Kur19]. · [Ano15f, Ash82, Mal10, Ste12a]. Δ [MS04]. ΔT [Ma23, MHZS23].

-1000 [Hub83]. -4000 [Gin91b]. -601

[Hub83]. -86 [Mar75]. -Ft [Edd71b, Mau13]. -inch [Ost07]. -year [GB95]. -Year-Old [Eva09b].

Żagań [CM10].

/Catalogue [Kun91]. /Charles [Tur07]. /Collected [Gin93]. /Heretic [Tes10]. /the [War08].

0 [Ave18, Hei14a, Oes15, Swe17c, Wło15, Ash82].

1 [Ano15f, BH73a, Ber14, Bru78b, Eva87b, Ito09, JLS^{+}05, Kle14, ML14, Moz13b, Ost84b, Rap15a, Ree20, Sha71b, Smi08, JLGL14]. 10-foot [Kär24]. 10.1484 [Fal23]. 10.1484/M.ALFA [Fal23]. 10.1484/M.ALFA-EB.5.12503 [Fal23]. 1006 [Bot76, Por74]. 102 [Cul07]. 104 [Cul93]. 1048 [Ken78]. 1087 [GS13, Lip15, Lip15]. 11 [Gin02d]. 1133 [MHZS23]. 1170s [HS16]. 12.50/$17.95 [Car86]. 1200 [Pin72]. 1239 [UCI16]. 13 [Gol84]. 1330 [Dob86]. 1354 [UCI16]. 13th [Nor92a]. 14 [Eva07]. 140. [Too95]. 1420 [Mal97]. 1450 [Mol93]. 14th [Not23a]. 1500 [Moz13b, Moz13c, Gra05c]. 1528 [Dob77]. 1550 [Gin13b, Kre13]. 1568 [Oes10]. 1577 [Don73a]. 1589 [Kre07b]. 1593 [CR22]. 15th [Gin90d, Tur70a]. 16 [Jon12]. 1600 [SG09, Gin04a, War08]. 1601 [Wil00]. 1603 [Gin12b]. 1610 [Hei12, Fin99]. 1611 [Dij09]. 1616 [Fin19]. 1630 [Oos15]. 1631 [Gin03f]. 1640 [Vic85]. 1650 [Gin85e, Swe86a]. 1660 [Bre11]. 1676 [Fab19]. 1687 [Eas96]. 16th [CMP24, Gin90d, Gin11b, Sch14a]. 16th-Century [Gin11b, CMP24]. 17 [MSH19a, MSH22a]. 17-36-5 [EP22]. 17-36-5/1741 [EP22]. 1700 [Cam11, Llo89a]. 1703 [Coh99]. 1719 [Apt99, Fei03]. 1730 [Ver87]. 175 [SS06]. 1750 [Hal03]. 1751 [Gin80a]. 1755 [Wil11]. 1777 [Hut12]. 1799 [Gin87c, Gre02]. 17th [Car23, Gin90d, Has11]. 17th-Century [Has11]. 1800 [Gin90b]. 1805 [Coh86a]. 1817 [Sch12b]. 1835= [KKSS21]. 1835=-1887 [KKSS21]. 1838 [Eva86, Hos08c, Wel94]. 1840 [Eis85, Gin85b]. 1845 [Eva87a]. 1850 [Ben91a]. 1860s [Pan94b]. 1870 [Kol01a]. 1870s [Pan94b]. 1877 [Cen14]. 1880S [Sta04]. 1885 [Jon76]. 1889 [Ban13]. 1897 [Sut72]. 18th [Bec15, Cha21, Hos14c]. 18th-Century [Bec15, Hos14c]. 1900 [Wil88]. 1909 [Dic90]. 1918 [Gra12c]. 1920 [Gra12b, Dew00]. 1921 [Rud92]. 1924 [Ber97a]. 1930 [Gol73a, Hos04f]. 1931 [Ber82]. 1935 [Bat87]. 1939 [Bel09a, Ost05a]. 1940 [DeV98, Dea94]. 1945 [Her94]. 1950 [Ham09, Kri94, PdV85]. 1960 [Bru96]. 1963 [KI10, Lan87]. 1968 [Dic97a]. 1970 [Kru08]. 1971 [Wil75]. 1973 [Doe19]. 1975 [Ost80]. 1976 [Gin02d]. 1979 [Ano78d]. 198 [Wes97]. 1980 [How88]. 1981 [Orc84]. 1982 [Ash86]. 1984 [Déb87]. 1987 [Mel89]. 1990 [Jas96b]. 1990s [Hos96]. 1994 [Dic95]. 1998 [Dic00b]. 1999 [Rot01]. 19th [GJ22, Hol10b, Sch14a, Tur70a]. 1st [Wal13].

2 [Ano01c, Ano05a, BH73b, Bru78c, Cen14, Eva87c, GG14, Hei14b, Hos11d, JHLS05, Jus20, Moz13c, Ost84c, SB72b, Smi09a, Wil14b, LGJLF14]. 2.1-9 [Eva07]. 2.10 [Eva07]. 2.10-14 [Eva07]. 2.3 [Fat00]. 200 [Cul02]. 2000 [Ost04]. 2002 [Bel07]. 2004 [Doe08]. 2006 [TF16, Wal13]. 2012 [Ald11, Liv13]. 20th [Str99, de 22]. 20th-Century [Str99]. 21st [Goe09, GHS08]. 23 [Ma23]. 24 [Azz11]. 267A [TF16]. 2800 [Gin91b]. 2A [Wes90]. 2nd [Hen15b].

3

[Apt14, Cro23a, Gan15, Hos96, Joh19, Mes15, Qua17, Rem15, Sul14, Vet19, Sil23, Ste99]. 3404 [Hal03]. 36948 [Ste02a]. 3rd [Cha15b, Mol14a].

4 [Ano94a, Ben20, Dev14b, Kes15, Lip15, Llo15, Wil15, Ste06a]. 400 [Ram07, Pas14]. 400th [Ash86, Pas15]. 40ft [Hos03b, Tur77]. 41 [Ano95g]. 430H [Kin81]. 45 [Bow06]. 4500-2000 [Hos00b]. 47914 [Oss24]. 4A [PdV85]. 4th [Llo15].

5 [Ano96a, Bel14a, Cha15b, Dun20, Gre14, Mir17, Pas14, For73]. 5/1741 [EP22]. 50/$100 [Rad90]. 500 [Ram07]. 527 [Ave95].

6 [Car00, Hen15b, Kra22, Lau14, Löh19, Mar14]. 600 [TQ17, Kre80a]. 601 [Hub83]. 60s [RS22]. 61 [Ser90]. 6231 [Ash81]. 684 [OP89].

7 [Bak15, Bec19, Cro15, Dan19b, Eva17, Lep14, Mol14b, Rot17, Sch15, Sha14, Wło23].

8 [Bel14b, CWW17, Hea15, Mal15, Roc17, Smi18a, Zep19]. 807 [NN24]. 837 [Yaz14]. 84i [Gol84]. 85 [Too95]. 85/140. [Too95].

9 [Bon19b, Bra15, Cha19, Gee18, Ger17, LF15, McC19, War19]. 90 [Hal03]. 90-698-3404 [Hal03]. 96h [Smi96]. 978 [Ano15f, Apt14, Ave18, Bak15, Bec19, Bel14a, Bel14b, Ben20, Ber14, Bon19b, Bra15, Cen14, CWW17, Cha19, Cha15b, Cro15, Cro23a, Dan19b, Dev14b, Dun20, Eva17, Gan15, Gee18, Ger17, GG14, Gre14, Hea15, Hei14a, Hei14b, Hen15b, Joh19, Jus20, Kes15, Kle14, Kra22, Lau14, LF15, Lep14, Lip15, Llo15, Löh19, Mal15, Mar14, McC19, Mes15, Mir17, Mol14b, ML14, Oes15, Pas14, Qua17, Rap15a, Ree20, Rem15, Roc17, Rot17, Sch15, Sha14, Smi18a, Sul14, Swe17c, Vet19, War19, Wil15, Wil14b, Wło15, Wło23, Zep19]. 978-0-19-879955-9 [McC19]. 978-0-19-995978-5 [Gre14]. 978-0-226-08440-4 [Kes15]. 978-0-226-16697-1 [Rap15a]. 978-0-226-30442-7 [Sha14]. 978-0-226-49697-9 [Bon19b]. 978-0-226-52228-9 [Bra15]. 978-0-262-01833-3 [Sul14]. 978-0-262-03691-7 [Dan19b]. 978-0-520-28799-0 [Swe17c]. 978-0-578-49710-5 [Dun20]. 978-0-674-07282-4 [Dev14b]. 978-0-691-14753-6 [Lau14]. 978-0-691-15335-3 [Mes15]. 978-0-691-17489-1 [Ree20].

978-0-7190-9144-5 [Cha15b]. 978-0-7486-7889-1 [ML14]. 978-0-8101-2652-7 [Lep14]. 978-0-8229-4530-7 [Bec19]. 978-0-901180-54-4 [Ben20]. 978-0-9566948-6-7 [Sch15]. 978-1-107-03174-6 [Hen15b]. 978-1-107-04406-7 [Rot17]. 978-1-10703-208-8 [Bel14b]. 978-1-108-47159-6 [Kra22]. 978-1-137-48178-8 [Ano15f]. 978-1-138-10117-3 [Qua17]. 978-1-250-00877-0 [Wło15]. 978-1-4094-4201-1 [Ber14]. 978-1-4094-6643-7 [Bak15]. 978-1-4197-1387-3 [Rem15]. 978-1-4214-1442-3 [Gan15]. 978-1-4331-2860-8 [CWW17]. 978-1-4426-3017-8 [Smi18a]. 978-1-4614-4096-3 [Apt14]. 978-1-4614-4423-7 [Mol14b]. 978-1-4614-5362-8 [Hea15]. 978-1-4724-2238-5 [Bel14a]. 978-1-58093-343-8 [Mal15]. 978-1-60618-005-1 [Ano15f]. 978-1-60732-641-0 [Ave18]. 978-1-63149-016-3 [Vet19]. 978-1-78297-643-1 [Kru17]. 978-1-929154-49-4 [Wil15]. 978-1-930972-83-4 [Llo15]. 978-1-935623-19-9 [Ger17]. 978-1-944466-18-3 [Doe19]. 978-2-503-54129-7 [Eva17]. 978-2-503-56876-8 [Zep19]. 978-2-503-59609-9 [Fal23]. 978-3-030-86599-3 [Cro23a]. 978-3-05-006365-2 [Hei14b]. 978-3-319-09640-7 [Cro15]. 978-3-319-21776-5 [Mir17]. 978-3-319-25863-8 [Roc17]. 978-3-319-31992-6 [Löh19]. 978-3-319-56783-9 [Gee18]. 978-3-319-65598-7 [Mit19]. 978-3-319-91841-9 [Sil23].

978-3-319-91843-3 [Sil23]. 978-3-515-10468-5 [Bur15]. 978-3-8440-2450-0 [Oes15]. 978-3-941563-19-3 [Joh19]. 978-3-944913-39-1 [Kle14]. 978-3-944913-42-1 [Ano15f]. 978-3-944913-43-8 [Ano15f]. 978-5-94052-225-6 [Ano15f]. 978-83-86062-94-2 [Wil14b]. 978-84-475-3588-0 [Hei14a]. 978-84-9168-164-9 [Cha19]. 978-84-9965-031-9 [Ano15f]. 978-87-7533-060-7 [Wło23]. 978-88-222-6254-7 [Oos15]. 978-88-222-6584-5 [Fin19]. 978-88-470-2828-9 [LF15]. 978-88-569-0522-9 [Cen17]. 978-88-7642-485-4 [Lip15]. 978-88-88649-38-2 [Cen14]. 978-90-04-18107-6 [Mar14]. 978-90-04-34227-9 [War19]. 978-90-04-38380-7 [Boh19]. 978-90-04-39234-2 [Jus20]. 978-90-481-2232-5 [Pas14]. 978-94-007-6874-1 [Bec15]. 978-94-007-6875-8 [Bec15]. 978-956-353-131-2 [GG14]. 99h [Nau98]. 99j [Nau98]. 9x [Fat00].

= [Ano15f, Chi16, Jar14]. Īsā [NRKN16].

Âge [Har85].

A. [Smi22, Too84]. A.D [Bak15, Coh86a, Mol93, Ols93, Osl06]. A.D. [Cul02, Cul07, Jon12, Ma23, Moz13b, Moz13c, NKM^{+}16, NN24, OP89, Pin72, Por74, Ram07, SG09, TQ17, UCI16, WS16, Yaz14]. AAAS [Rob86]. Aaboe [Ano07a, Gin88a, Gin88a]. Aachen [Oes15]. År [Hol03]. Aaron [Sch15]. AAS [Hin04]. AAVSO [Smi12]. Abī [KF70]. Abū [VB01]. Abû [Loc95]. Abbot [DeV90]. Abbreviating [Kre07a]. Abbreviation [Sal95]. ‘Abd [GKS72]. Aberdeen [Dob73, Rei82]. Aberration [GB09]. Abi [RA18]. above [LB06]. Abraham [Gin91d, Jus20, Kre02, RA18, Gol98, GC99, Gol11, Gol12, Jus20, Kre02, Gol11, Gol12, Jus20, War19]. Abrams [Rem15]. Abridged [Gin16a, Str99, Car99]. Absolute [Eva93a]. Absolutism [Sha94b]. Absorption [SB72a, SB72b]. Abstracts [Mea70a]. Abu [Ken78, Ada21, NKM^{+}16, Nor92a, Sal95]. Abu-r-Raihan [Ken78]. Abundances [DK83b]. Abyss [Dra80]. Académie [Ler06]. academies [Cha21]. Academy [Bok73, Oes15, Tur79a]. Accademia [Tur84]. Accademici [GM03]. Accelerating [Ell08]. Acceleration [Dog87, Edd82]. Accelerations [MSH20, Ste77]. access [Doe19, Fal23, Sil23]. According [Kre93, SC88]. Account [Gol94, Ree20]. Accounting [CWW17, Ush15]. Accuracy [Cha83, Edm11, FS96, LZ99, MSH21, MSH22b, Moz13b, Moz13c, Moz19, Nev96, SS91a, SSM97, SS98a, SH81, SF95, Wes78, Wün99]. Achievement [Wal77]. Achievements [Ano84b, Hos02b, Woo80]. Achilles [Dan04a]. Achromatic [Bak15, Bre18, Ost94, Kär24]. Achttiende [Van02a]. Ackermann [Boh19]. across [McC07, Nor06, Rug04]. Act [Ste13]. Acta [Ano15f, Kle14, Voe03]. Activities [Tat99]. Activity [CG10, Sak80, UC16]. Actors [Gre23, Tat99]. Actuelles [Cap12]. Actus [Hei14b]. AD [Wal13, Gin97a, GDEK24, CR22, MHZS23, Sch23a]. AD1239 [UC22]. AD1478 [UC23]. Adair [Sch15]. Adams [Dic91a, HB92]. Adaptations [CG18]. Adaptive [Max12]. Addenda [Hos91b, PHP93]. Addendum [EBA94, GV98b, HP99, Le 09, Sch84a, Van71a]. Additional [Zed05, Swe86a]. Addresses [Tur71b]. Adelard [Eas89, Sal95, Eas89]. Adjusting [BG10]. Adler [Ack10, Dun20, Van99, Ack10, Van99]. Ado [Eva82]. Adolf [Hen13, Hen13]. Adriaan [BH73a, BH73b, Het74, HB92]. aduqa [Ste12a]. Advance [Lov77]. Advancing [Smi12]. Adventures [Ber94, Dic00a, Ost02]. Aesthetics [Olo04]. Aether [Gol73a, Gra00]. Aether-Drift [Gol73a]. Aevi [Gin87c]. Affair [Coy07, Dra90, Gin05d, McM09, Déb96, Ped83]. Africa [BEG98, BBMG03, EBB^{+}01, Sne98, War77, War79]. African [KHM^{+}12, Mal15]. After [CJ20, Eas96, Gra11b, Ano80a, Ash79a, Gin11a, Gol87a, Hir85, Hos73, Mea70b, Rap15b, Swe70, Van95b]. Aftermath [HW07]. Again [Bar19, Eis06, Eva93b, Wło15, Gle07]. Against [Lep14, Ald97, Don79, Gin02c, Gra10, Gra12d, GB86, Gue12, Sha07b, Van97]. Age [Ber97b, Bru09, Edd71a, Gin01d, Kra07c, Mac93, Mar97, Mil04, Pan95, Sch03, Ste00a, Tri10, Mer09]. Ageing [Gin88a]. Agency [Omo20]. Agenda [Hos12b]. Ages [Aul16, DeV00a, Ell08, Eva82, McC99, Mer09, Mor08, Obr14, Sch84a, Sch84b, Ano78b, Dal78, Eas96, GC06, Hos97a, Pou75, Sam13, Ste96, UC17, Sam13]. Agial [BEBM02]. Agnes [Eis03, Eis03]. AHU [Lil89]. A’Imma [GKS72]. Air [Lat20]. Aires [Nor91b]. Airy [Bel20, Cha88, Cou18]. Aiton [Ano91b]. Akademie [Hal03, Hei14b, Oes15, RR13]. Akademische [Ano15f, Kle14]. Akkadian [Sol09, Sol09]. Akkadskie [Sol09]. Akten [Swe98]. Al [Ans05, Ken78, Sal09, Ack05, Ave87, Gli88, Gol73b, Hos01b, Ken94, Kin87, Liz17, Rag85, Ros74b, Sal94b, Sch07b, Sha92, Ves73, GKS72, GC06, Kin73, RPPA15, Sam07, Sch13, Yaz14, GC19, Hos01b, Moz22, Naò16, Not24b, Rag16b, RA18, Sal87, Sal94b, SM94, Swe17c, VB01, Ves73, Ken73, Rag83, Sam07, Ack05, Kin95a, NRKN16, Sal79]. Al-A’Imma [GKS72]. al-’Amakin [Sal79].

Al-Andalus [Sal09, Hos01b, Naò16, Swe17c]. Al-Asfi [RPPA15]. Al-Birjandion [Ans05]. Al-Biruni [Ken78, Ros74b, Rag83]. al-Bitrūjī [Sha92]. Al-Bitruji [Ken73]. al-Corsuno [RA18]. al-Dīn [Sal94b, Ves73]. al-Din [Ack05, Ken94]. al-Farabi [Liz17]. Al-Farghani [Sam07]. Al-Farghani’s [Sam07]. Al-Hadib [GC06, GC19]. al-Hay’a [Rag85]. al-Haytham [Gol73b]. al-Haytham’s [Kin95a]. al-Kūhī [VB01]. al-Khafrī [Sal94b]. Al-Khalīlī [Kin73]. Al-Kindi [Yaz14]. al-Magrib [Hos01b]. al-Misri [Ack05]. al-Rūh [NRKN16]. al-Shātir [Rag16b, Sal87]. al-Shatir [Moz22]. al-Shukuk [Gol73b]. al-Tūnisī [Not24b]. al-Tūsī [Ves73]. al-Tusi [Ken94]. al-zījat [Kin87]. al-Zarqālluh [SM94]. Alīqushjī [Rag05]. ala [Gol73b]. Alain [Ler11a]. Alain-Philippe [Ler11a]. Alan [Gin13a]. Albany [Bel07]. Albert [Ano84b, Cro23b, Bar13a, HE82, Mes20, Mes21, Mil81, Woo80, Ano84a, Gin82a]. Alberto [Ped79]. Albertus [Ped79, Pri91, Ver06]. Album [OP92, Dar91]. Alchandreana [Cha09]. Alchemy [Whi00]. Aldana [Kru17]. Alexander [Duk11, For81, Rad90, Roc17, Atk86, Hos04a, Spa03, Smi94b]. Alexandra [GG14]. Alexandria [Eva88, Eva01, Jon12, Loc93, Too96, Too96]. Alfonsii [Not15b]. Alfonsine [Swe04b, Ano13a, Cár88, CG03, Cha07, CG13, Cha17, GC01, GC08, Jus23, KD98, Kre07a, Kun86b, Not15b, Not23b, Pou88, SC88, Sam19, Swe04b]. Alfonso [Nor00b, Nor00b, Pou88]. Alfred [Bur15]. Algeria [SK08]. Algol [MD21]. Alhazen [Gol73b]. Ali [Kin79]. Alice [Wen99]. Alien [Cro12]. Alignment [Wal06, Wal07]. Alignments [FQ77, PC89, Sch86, Spr00, TT72a, TT72b, TT74]. Alimberto [Don79]. Allāh [NRKN16]. alla [Ano06a, Bel11, May11]. Alle [Cen14]. Allées [Hos07c, SK08]. alleged [Wil79].

Allemagne [Kus10]. Almagest [Duk11, Eva92, Gin00b, Too77, Gin16a, MHS17, Ano95b, Ano01b, Duk02, Duk11, Gol76, Kin95a, OT13, Ped87, Sch01a, SL13, Swe21, Thu95, Too77, Van94a, Wri07, Zie23, Gol76]. Almagestum [Rap11a]. Almanac [CG19, Dun92, Rot01]. Almanac-Makers [Dun92]. Almanach [Kus11]. Almanacs [BBW97, Bri01a, Dun92, Gin81a]. Almeria [HAG94b, HAG95b]. Almosnino [She15]. Aloft [DeV19]. along [UAGGB21]. Alpetragius [Ken73]. Alphonsi [Cár88]. alphonsine [PS98, Ler06, Nor85, PS98]. Alphonsines [Nor85]. Alquibla [Hos01b]. altbabylonischen [Ste12a]. alten [Siv77]. Alternate [Mal15]. Alternative [Kra95, Wes09a]. Altertum [Eva06]. Altitude [Arn11, Bog02, CL05]. Altitudes [SS95]. Altnordische [Jor94]. Altobelli [Lat12]. Altos [Had84, Mul83]. Alvan [Dew70, Gin97b, Dew70, Gin97b]. am [Rut12, Sev00]. ’Amakin [Sal79]. Amalgamator [Jar09b]. Amalthea [WDR11]. Amateur [Bec18, Dew00, RC16, Sta00a, War80, Gin00a]. Amateurs [Rot18, Van02a]. Amazing [Lau14, Ost02]. Ambassador [Cam02]. Ambition [O’D06]. Ambitions [Hos02b, Hos04b]. Amendment [How94]. America [Bed10, Bru96, Dun92, Vet12, Ano73a, Bru78a, Car76, Dun92, Had84, HS92, Kee91, Kru22, Nal16, Slu97a, Ave09, Car76, Van95a, Waf02]. American [Ano15f, Ave95, Bru78a, Dic00a, Doe01, Doe08, Gub88, Hev94, Hin04, Hir92, Jar90, Kel83, Kid87, Kru08, Rot03, Smi94c, Smi12, Ano75a, Ano80e, Ave09, BD82, Ber74, Bro91a, DeV98, DOL98, Lan83, Nee91, Num73, Rom06, Rot18, Sch12b, You86, DeV98, Vet19]. Americans [Hir92]. Americas [Bro91a, Mul83]. Amici [Cas21]. Amico [Swe92b, Swe72, Swe92b, dB95]. amid [How05]. Amidst [Don73a, Waf04]. Ammi [Ste12a]. Ammi- [Ste12a]. Among [Sne98, Ave92, Bru78b, Bru78c, Rya09].

Amphiareion [Sch04b]. Ampurias [Bor98]. Amsterdam [Ano78d, Ban13, Hal03, vN12]. analemma [Sid24a]. Analemmatic [Par85]. analemmatiques [Par85]. Analogue [Muc75]. Analogy [Num73]. Analyse [Joh19]. Analysed [Eas89, Gin82a, Gol76, Siv98]. Analysing [Gin89c]. Analysis [Bel14a, BG12, Bon05, CS21, CMP24, Duk06, Eis88a, Ext92, GB10, Gol75, Hen15b, Iwa03, Jon71, Jon76, Kri93, Mac98, Not15b, Oss12, RAGGB18, Ros82a, Top03, Yor01, vdW17, Car24, GH24, Hea14, Moz22, She90]. analyzed [Hei22]. Analyzing [Bel14a]. ANAsh.Mus.D.O.633 [EP22]. Anaxagoras [Gre14, Tho71a]. Anciens [Cha09, Hos04f]. Ancient [Ald09, Arn11, Ave97, Bel14b, Duk03, Duk08, Duk09, Edd82, Gin88a, Gin11c, Gol03a, GG14, Gra14a, Gre02, Han09a, Han09b, Har77, Har78, Haw74, Hiv81, Hos05a, Jon04, Jon15, Ken87, Llo80, Llo18, Mar12b, Mer03, Mit73, Moe83, Nor88, Obr19, Oss18, Pin84, Qua17, Roc08, Roc16, Roc17, Roc19, Rug00, Rug07, Sid04, Sid10, Siv98, Sol09, Sym15, SK16, Tau09, Voe01, Was77, vdW90, Cul17, Han21, Imh22, Kur21, Sid24a, TW25, Win25, ABB03, BS06, BSF08, BFAH^{+}10, CS14, Cul93, Cul05, Dav85, Duk04, Eva08, Han09b, JKKC19, Loc85, Loc90, Loc96, Loc99, LB06, Roc14, Sch00a, SB05, SBF07b, SBF07a, SK97, SC13, Tre90, Tum74, Yor01, Bur82, Eva02, HS81, Llo09, Loc96, Nor96, Ste77]. Andalucia [HAG95c]. Andalus [Hos01b, Sal09, Naò16, Swe17c]. Andean [McC83a, Ald97]. Andes [Ave97, Ave18, Ave18]. André [Mar14]. Andrea [Cro23a]. Andreas [Oes15, RR13, Gre10, Kre10, Kus11, Rem08]. Andrew [Ave18]. Andromeda [Van14b]. Andromedae [Jon76]. Andrzej [Ave95]. Anfängen [Dob99a, Kri94]. Ángel [Cha19]. Angelus [DK96]. angesichts [Sev00]. Angle [Whi75]. Angles [And90, Ben91a, Dic00b]. Anglo [Hea94, Hea94, MSH22b]. Anglo-Australian [Hea94, Hea94]. Anglo-Saxon [MSH22b]. Angular [Ben91a, Cha83]. Angus [Ano76a]. anhand [Row07]. Anima [Bon05]. Animations [Duk04, Van98a]. Anita [Bak15]. Anna [Lip15]. Annals [Bok73, NN24]. Années [HRVM13, Hos04f]. anni [CR22]. Annibale [Nal22]. Anniversaries [Gin90h]. Anniversary [Ash86, Pas15, Tri10]. Anno [Gin01c]. Annotated [Gin83b, Has11, LGJLF14, Mos03, Pan09a, Pin03]. Annotation [Duk11]. annotationes [Mal13]. Annotations [GW81, SS16, Swe86a]. annual [Wil79]. Annulo [Van74a]. Anomalous [Lan84]. Anomaly [Car19, ECT10, She99, MD21]. Anonymous [CS21]. Ans [CL05]. Ansätze [Hen92, vdW78]. Anses [Van74b]. Antão [LN09]. Antebellum [Eis88b]. Antecedents [Har79]. Antediluvian [Spe13]. Anthes [Kle14]. Anthology [Arm74, Hos76a, Nor94, Smi93]. Anthony [Ald09, Ano01c]. Anthropological [Had84]. Anthropology [Ave95, Bra15, Pla91, Mia12]. Anti [Bal85, Fin19, Mil80]. Anti-Copernican [Fin19, Bal85]. Anti-Einstein [Mil80]. Antiche [GS13, Lip15]. Antichrist [Rut21]. Antiguas [Hos00a]. Antiguo [Bel06, Hos12a]. Antike [Bur84]. antiken [Eva08]. Antikythera [ASE^{+}13, ASCE14, CTE12, Car17a, Edm11, Eva09b, ECT10, Mer77, VMV18]. Antiochos [BG10]. Antiquarian [Ben20]. Antique [Bre18, WW82]. Antiquities [dP82]. Antiquity [Aab72, Bon11, Bur84, Cer99, Cou18, Duk12, Gin10c, Har78, Haw74, Hos97a, Llo09, Mer11, Pin73a, Rüp15, ZJJ^{+}99, Eva11]. Antología [Smi93]. Antonella [Cen14]. Antoni [Ano15f]. Antonio [VG01]. anywhere [Cov17]. Apex [Hos80b]. Apian [Mal90, GE24, Kun87]. Apianus [Gin71, Gin91a, Gra11a]. Aplanatic [Ben82b, Bre18]. Apollo [Sch12b, Bro10, Kes10]. Apollonius [Jon99]. Apologetik [Sev00]. Apology [MSH22a, Omo19]. Apotheosis [Cog08]. Apotheosized [Gin88a]. Apparatus [Eva04, Gen87]. Apparent [Gol96b, Sha75]. Appearances [Pet84, Ush15, CWW17]. Appearing [Vai12]. Appended [Kun86b]. Appendix [CR22]. Application [CTE12, Hos16b, Not14b, Pop22]. Applications [LS05, Edd80]. Applied [Hos83a, Not19, Ste04]. Appointment [Ost92b]. Appraisal [Pal08, SS86]. Apprehensions [Waf86b]. Apprenticeship [Vic85]. Approach [Don96, GZB14, HC99, Tur70b, Wei82b]. Approaches [Hen15a, Hen92, Rut13]. Appropriate [Jar90]. Approximative [Van97]. Après [Hen76]. April [Ave95, Gin02d, Bou12]. Apsidal [Moz13a, VWH97]. Aqsà [Hos01b]. Aquarii [Cov15]. Arab [Gin06b, Nor92a, Nor92a, Por74]. Arabe [Cha09]. Araber [Kin83]. Arabia [NN24]. Arabic [Ber97b, Ber12, Gin01a, Mal90, Ras13, Dal01, Duk04, Kin02b, Kun85, Kun86a, Kun87, Mer06, NRKN16, Not14b, SS91a, Sal89, Sal90, Sal94a, Sal94b, Tur23, Ans05, Cha09, Dal01]. Arabic-Islamic [Kun85]. Arabic-Latin [Mal90]. arabisch [Eva92, Gin00b]. arabisch-Mittelalterliche [Gin00b]. Arabische [Mal90]. arabischen [Kin81]. arabischlateinischer [Too77]. Arabist [Eas89]. Arabo [Eva89, Ves98]. Arabo-Byzantine [Eva89]. Arabs [Ber92, Ber92]. Arachne [Sch04b]. Arago [Lev17a]. Aratea [Gee18]. Aratean [Lip15]. Aratus [McM01, Eva12]. Aratus. [Eva12]. Arbours [Hos81]. Arc [Ste10a, Thu03]. Arcetri [Can12, Can12]. Archaeoastronomical [AR86, Iwa03]. Archaeoastronomy [AB14, Ano73a, Ano79a, Ano03a, Ash00, Bel97, BEG98, BEBM02, Bel11, Bel16a, Bro91a, Car76, Ext92, Had84, HAG94b, HAG95b, HAG95c, Hos96, HP98, HMA^{+}98, HS99, HVT^{+}01, Hos02d, Iwa96, Jas96a, Jas96b, Kru83, Kru17, McC07, McC17, Mur98, Pla91, Rug07, RH00, Tir22a, Ano95a, Ano01c, Car76, Kru83, Kru17, Mul83, Rob86, Rug07, Shi11, Wil87a]. Archaeologi [Bel11]. Archaeological [Ell85, Jas96b, Rug82]. Archaeology [Ave95, Edd74, McC98b, Tir22a]. Archaeotopography [Bel97, BEG98, Bel14b]. Archaic [Ram10, Ste08a]. Archduke [Mes20, Mes21]. Archenhold [For73]. Archenhold-Sternwarte [For73]. Archeoastronomia [Bel11, Hos94]. Archeologia [Rug95]. Archetypes [Dav12]. Archimedes [Sha75, Swe04b]. Archipiélagos [Iwa96]. Architect [Hos81, Jar00]. Architectonica [Pou09]. Architectural [Tur07]. Architecture [BM15, Bel16b, Ben75, Her76, Kin95b, Bel14b, Hos04e]. Architektur [Kri94, Her76]. Archiv [Bur15]. Archival [MG83]. Archive [Gin99c, Hat08, Hos05d, AKJ^{+}23, Ano83c]. Archives [Ano73b, Ano82a, Ano90h, BH87, DH73, Haz02, Hos82a, Jar75, Jar77, Mes20, Sha71a, War77, War78, Mor80, Ell90, Haz02, Hos05a]. Arcium [HH07]. Arctic [Slu97a]. Arcus [dJ12]. Ardnacross [MR93]. Arduous [WDR11]. Area [CLL^{+}00, Dav15]. Area-Measure [Dav15]. Argelander [Gef75]. Argentina [Gra12b]. Argentine [Hod77]. Argument [Gra13b, Hus12, Mue00, Pal08, Sal76, Top03, Nau98]. Arguments [Blå14, Nor92b]. Argyll [Rug85]. Aries [Mol03]. Aristarch [Zbi97]. Aristarchus [Tho87, Gin85f, Zbi97]. Aristotelian [Mar14, Sch03, Dob99b, Hoo87, Swe72]. aristotelischen [Ham06]. Aristotilis [Nor99]. Aristotle [Bur10, Eva07, Kre08, Lat00, Nor99, Osl06, Tau11, Cro07, De 14, Eva93a, Eva07, Fek24, GT00, Llo15, Mar14, Ste00b, Llo15, Was72]. Arithmetica [Gin89b]. Arithmetical [Neu82]. Arizona [Gin01e, Hoy84, Liv13, Hoy84]. Armagh [Sch91, BH87, Sch91]. Armenia [Kul75, Tum74, Kul75]. Armenian [Kaz73]. Armenoi [PHP92, PHP93, PH96]. Armies [Gra13b]. Armillary [Gin03d, Wło87]. Armillis [Cha15a]. Armin [Kle14]. Armitage [Ano76a]. Arms [Hos16b]. Arnold [Ave95]. Arpenteur [Ver09]. Arqueoastronomía [Iwa96, Spr93a]. Arqueología [Hos12a]. Arrangement [Hos78a]. Arribas [Boh19]. Art [Azz11, Bol05, Dan17, Dev14b, Don82b, Dup14, GHS08, Gin86a, Gin99a, Goe09, Lüt92, Mon94, Mur98, Ste00a, Yeo99, You86, Hut13, Had84, Gin86a]. Artes [Ste96]. artful [GE24]. Arthur [Bro88b]. Article [Ano96e, Ano97i, Ano98h]. Articles [Ano92h, Ano93i, Ano94h, Ano95i]. Artifact [Gin11c]. Artificial [Ree11]. Artist [Eva87a]. Artists [Dew70, Gin97b]. Arts [Tay10a, KHM^{+}12, Mal15]. as-saniya [Rag85]. as-Sunniya [Rag85]. as-Suyuti [Rag85]. Ascendancy [CT12]. Ascensions [Neu82]. Ascribed [Don79]. Asfi [RPPA15]. Asger [Gin88a, Ano07a]. Ash [She79a]. Ashbrook [Ano80d, Gol87b]. Ashgate [Bak15, Bel14a]. Asia [Shi11, CMP24, Slu97a]. Asian [Ano03a, SG09, Ano03a]. Aspects [Eis88a, Gin99b, Wea90, Hen92, Pou09]. Aspekte [Sev00]. Aspekten [Hen92]. aspera [GDEK24]. Assemble [Bed80]. Assembled [Dal01]. Assembly [Ano89a, Déb95, Dic97b, Dic00c, Nor91b]. Asservantur [Gin87c]. Assessed [Kre05, Lon06]. Assessment [Edm11, SMH20, TTF76, MHS23].

Assistants [Wil00]. Associated [NR14]. Association [Duk03, Smi12]. Assumptions [Sza18]. Assyrian [Iwa03]. Ast [Bow06]. Asteroid [Mir17, Cun16]. Asteroids [Her20, RS18, Wil02, Con24]. astonishment [HDG^{+}13]. astra [GDEK24]. Astral [Jar09a, Jon06b, Oss18, Jon21, Oss24, Bri01b]. Astrarii [Tur03]. Astrarium [Gin89a, Gin89a]. Astro [Edd74, Ost05b, Sta99, DeV07]. Astro-Archaeology [Edd74]. Astro-Optiker [Ost05b]. Astro-Physicists [DeV07]. Astro-Technik [Sta99]. Astroarchaeology [Har73]. Astrobiology [Ram06, Str16b, Ram06]. Astrofisica [Cen14]. Astrogeology [Yeo02]. Astrograph [Ost07]. Astrolabe [Dav17, Dav19, Eva09a, Gib78, GKS72, Gin01a, Hen75, Nor90, Nor02, Plo00, Pou79, Pou09, Sam07, SV16, Tur96, Tur15, Wło87, DL15, Hus13, Tur23, Sha84, Eva09a, Gib78, Hen75, Nor90, Sam07, She15, Tur96]. Astrolabes [Ack10, Ano74a, Boh19, Dek12, Gin88c, Gin03e, Hen75, Nor04, Pou79, Ree11, RA18, Ste06b, Van99, Ack10, Boh19, Dek12, Gin78c, Ste06b]. Astrolábu [Nor02]. Astrologer [Ano18a, Bow83, FJ11, Gin07a, Gre10, Azz11, Eva04]. Astrologer-Friend [Gin07a]. Astrologers [Rut16]. Astrologi [Kre89]. Astrologia [Gin10a]. Astrologica [EP22]. Astrological [Azz11, Bur20, Gol72, Gol11, Gol12, GC13, Jus20, Mol03, RL18, Ada21, Hei22, LS24b, Mon22, War15, Wes01]. Astrologie [Eva06, Gin07a]. Astrologiques [Cha09]. Astrologische [Gau10]. astrologue [Bow83]. Astrology [Bon05, Bon09a, Bon11, Cam11, Cha09, Dal01, Doo18, Dup14, Fek13, Gin81a, Gin10a, Gin14, Gol72, Gol95, Gol12, Han09a, Kre89, Li19, Mal97, Mea81, Mea86, Mer09, Mes21, Nor96, Nor07, Not19, Pin73a, Pin90, RL18, Rut03, Rut13, Sal95, Sal09, Smo09, Tau09, Tho90a, Tho90b, War19, WP22, vB17, Mon21, Rib24, Gin14, Azz06, Bon11, Fek13, Mea81, Nor96, Nor07, Pin73a, Pin90, Rut03, Rut13, Ste01, Tau09]. Astrometric [Wün99]. Astrometry [Dic00b]. Astronauts [Ger17, Neu13, Kes10]. Astronom [Gol84, Nor80, Sha08]. Astronome [LC10, Bow83, Can09]. Astronomen [Gin06a, Kin83, Gin90f]. Astronomer [Apt99, Aub16, Bec18, Coh97, Coh99, Déb13, Dew00, Eva87a, Fei03, Gin81c, Has11, Hos88, How05, Kle14, Kri88, Lev17a, Mir17, Naz16, Per00, Sha07b, Spe13, Sul77, Tho71b, Tri13, Ver17, War93, Wil14b, CJ22, Sel12, Rei18, Siv05, Apt14, Gen90b, Phi91, Sel14]. Astronomers [Ave18, Ban12, Ben84, Ber97b, Bru78b, Bru78c, Cow71, Dew00, Dic97a, Dob99b, FS96, Gin91b, Glu22, Hos89a, Hos08a, Kid90, Kin83, Kre18, Kul80, Kul87, Maf05, McC82, McC94, Mor01a, Nas20, O’D06, Ome79, Pas03a, Rot80, Sch12a, Sha87b, Shi00, SF95, Oes15, SS96, SS97a, Sil23, Ben84, Doe01, Gin85a, Hol10b, Hos08a, Rot80]. Astronomes [Ver87]. Astronomi [Bel11, Gin87c, Hei00]. Astronomia [Cár88, Gin96c, Pan09b, vN12, Ste02b, Bel06, Hos00a, Hos12a, Dav15, Has87, Rug95, Ste02b, Ste12b]. Astronomiae [Voe03, vB10, Ano15f, Kle14, Ped79, Ves98]. Astronomica [Don00, Voe97, Cha19, Don92, Tau13]. Astronomical [ASE^{+}13, Ano70b, Ano77l, Ano79b, Ano03a, Ano06a, Ano15f, AZB02, Ash87, ASS13, Ban13, dB70, Bel14a, BG10, BGP13, Ben87b, Ber94, Bol20, BH12, Bra90, BBW97, Bro88b, Bry81, Car08, Cau17, CK07, CG10, Cha12, CL22, Chi21, CMP24, Con70, Cul93, Cul05, DeV98, DeV00a, Déb96, Dew00, Dob77, Eas00, Eas10, Edd80, Eis88a, Eva93b, FV17, For73, For76, Gee18, Gen90a, Gen92a, Gin72, Gin85a, Gin88b, Gin90d, Gin90e, Gin90h, Gin92c, Gin05b, Gin11b, Gol87b, Gol98, Gol01b, GC06, Grö83, Haa84, Haz02, Hei14b, Her77, Hin04, Hol10b, Hos84a, Hos06a, How80, How86, How94, Hub91, Huf91, JJ10, Jar75, Jar77, Jon91a, Kel83, Kin79, Kin87, Kra17, Kre81b, KD98, Kre12, Kun90, Lan89, Lil99, Lüt10, McC13c, Mea71, MS85, Mon22]. Astronomical [MJT03, Nor76, Nor92a, Not15a, Not23a, O’D06, OP92, Oss12, Ost92b, Ost94, Pan09a, Pan10, Pap91, Pas14, Pau81, Péo08, PC87, Pet04, Pin71a, Pou85, Pow95, PRA87, PC89, Qua17, Rag92, Ram10, Rap11a, RPPA15, Roc19, Ron89, Ros74a, RMH91, RM92, Rut21, Sal90, Sal06, Sam13, Sch04a, SS16, Sch20, Sch84a, Sch84b, Sco90, Siv05, Spe80, Spr00, Sta11, Sta04, Sta10, SS98a, Ste77, SY92, SS06, SMH20, Sym15, TT71, UL95, UCI16, War77, Way96, Wil84, You86, Zru21, vB10, Aug24, CS21, Cha21, Che72, CP23, Cul17, Fal21, GH24, GJ22, Hea14, Hen15b, KG82, Kra24b, Kre80b, Mer98a, MD21, Pin98, Pri91, Sho21, SP22, Smi22, Gin91b, Her71, PP82, Dew00, Gee18, Gin05b, Gol00, Hos73, Hub91]. Astronomical [Mül91, Orc88, Sta10]. Astrónomicas [Iwa96]. astronomiceskie [Ano15f]. Astronomicheskaya [Kul87]. Astronómicos [Gra12b]. Astronomicum [Gau17, Gin71]. astronomiczne [Hil77]. Astronomie [Ash79a, Bat87, Cha89, Dob99a, Dup08, Eva06, For70a, For82, Gin07b, Gol95, Hat88, Her76, Kin81, Kre80a, Kre99, Kre07b, Oes10, Sta99, Gra80, PS98, Van98b, Cha75, Ham11, Her94, Mos09, Ver87, vdW78]. Astronomiegeschichte [Voe03]. Astronomies [Mal15, Rug00, Wal99, Hos96, Wal99]. Astronomii [Prz79, Wer10]. Astronomique [JT92, Mol76, PP82]. Astronomiques [Dob83b, Dob77]. Astronomische [Eas10, Gau10, Kre07b, Ste12a, Loc99]. Astronomischen [Her76, Fei13, Hei14b, Sta99]. Astronomischer [Gau17].

Astronomisches [Swe90]. Astronomiya [McC08]. Astronomy [Ack10, Ald97, Ald09, And90, Ano70b, Ano70g, Ano70h, Ano71c, Ano71d, Ano71k, Ano72h, Ano72i, Ano73b, Ano75a, Ano77k, Ano80a, Ano80e, Ano82l, Ano94c, Ano97c, Ano01a, Ano01c, Ans05, Arm72, Ash72, Ash79a, Atk75, AH94, Ave95, Ave99, Ave11, Ban10, Bar18, Bat87, Bat90, Bat15, Bec19, Bee71, Bel07, Bel06, Bel09b, Bel14b, BM15, Ben75, Ben91a, Ben92, Ber74, Ber81, Ber82, Ber97b, Ber14, Bru96, Bur82, CL05, Cár88, Cas18, Cau08, CG97, Cha98, Cha00, Cha17, Cha83, Cha15b, Cha75, CJ20, Dal01, De 13, DeV98, Dek08b, Dev14a, Dic97a, Dob83b, Dob83a, Dob86, DK96, Dob99a, Dob99b, Doe05, Don02, DH73, Duf17, Duk12, Dun92, Eas97, Eas01, Edd72, Edd80, EB04, Ell85, Ell90, Eva99, Eva01, Eva02, Eva05, Eva07, Far93]. Astronomy [For70a, For80a, For16, GB11, Gau10, Gau18, Gin72, Gin77b, Gin83b, Gin89c, Gin91b, Gin96a, Gin01d, Gin02a, Gin03b, Gin06b, Gin07b, Gin11a, Gin13c, Gol95, Gra12b, Gra12c, Gra05c, Gua09, Gue12, Had84, HH07, Hal03, Ham11, Har72, Har77, Har78, Har14, Hat88, Hea97, Heg85, Hei00, Hen08, Her22, Her94, Hew74, Hig16, Hir85, Hiv81, HH82, HH84, Hos71a, Hos73, HG80, Hos86b, Hos95b, Hos97a, Hos00a, Hos05a, Hos14c, How80, Jar98, JT92, Jon06b, Jon10, Kak95, Kel10, Ken73, Ken87, Kin78, Kin81, Kin87, Kre80a, Kre81b, Kre01, Kre07b, Kre10, Kre11, Kri88, Kun87, Lau14, LF15, Lei17, Lév17b, Li19, Lia19, Lil89, Llo80, Llo94, Loc96, Lon15, Lov77, Mac90, Mac93, Mar14, Mar12b, McC77, McC82, McC89, McC90, McC91, McC98b]. Astronomy [McC04, McC19, McC13b, McC08, McM01, Mea71, Mea85, Mer88, Mer96, Mer03, Mes20, Mit73, Mit86, Mit19, Moe83, Mol76, ML14, Mor01b, Mor01a, Mor08, Moz19, Mul84, Nak71, Nal16, Naò16, Nee91, Nee71, Nic90, Nor86, Nor97, Nor06, Nor08b, Ogj11, Olo04, Orc90, Ost80, Ost92a, Pal77, Pan09a, Pan09b, PP07, Pau95, Ped79, Pin72, Pin73b, Pin74, Pin76, Pin81, Pin84, Pla91, PdV85, Pou85, Prz79, Rag09, Ram16, Reb03, Rec21, Roc08, Roc14, RAGGB18, Rot18, Rug04, RC11, RC16, Sal87, Sal90, Sal97, Sal09, Sch81, Sch82, Sch91, Sch01c, Sha98, Sha87b, Sha70, Shi09, Sil23, Siv77, Smi85, Smi97, Smi08, Smi09a, Smi12, Smi20, Spa04, Spe13, Spe17, Sta72, Sta15, Ste01, Ste03, Ste04, Ste13, Ste16, Ste90a]. Astronomy [Ste94, Str19, Sul99, Sul14, Swe13, Swe17c, Tay10b, Ten00, Thi88, Tih98, Too77, Too95, Tri13, Van99, Van92, VB87, Ver87, Ves98, Voe01, VG01, Voe03, War79, Was72, WW71, Wer10, Wes90, Whi70a, Wil96, Wil04, Wil05a, Wło15, WKH18, Yan86, Yav99, Zso20, dV91, vG05, vdW78, vdW80, Dic13a, Hen97, Kin73, Lin04, Maa16, Mol14a, Mon21, Mor21, Mun13, Not23b, Oos15, Pin96, PS98, Sal94b, Ste23, Str23, Tih21, Urt21, Wło23, de 22, Ave06, Löh19, SGDO02, Bec15, Bok73, Gre14, LF15, Mit19, Sul14, Too95, Arm73, Arm74, Ash72, Ash00, Aub12, Ave97, Bar13b, Bat90, Ber89, Ber01, Bur80, Coh95, DeV00a, Dic97a, Doe00, Doe05, Don88b, Don02, Dra80, Eva89, Gin96b, Gin96a, Gin01a, Gin02a, Gin03b, Gin05c]. Astronomy [Gol87b, Gra76, Har78, Hea97, Heg82, Hew74, Hir09, Hos73, Hos83a, Hos97a, Ito09, Jar00, Ken73, Ken85, Ken87, Ken94, Kra99, Kre02, Liz17, Llo94, Loc96, Mac93, McC91, McC99, McC13b, Mer96, Mer03, Mit86, Nee71, Nor06, Nor08b, Olo04, Orc90, Ost80, Rug04, Sal09, Sha02, Shi09, Sid10, Siv89, Siv98, Smi85, Smo09, Ste05, Ste17, Ste90b, Thi88, Tho88, Tih07, Tur70a, Van98c, Van97, Voe01, Wil05a]. Astronomy/Scripture [Gin07b]. Astronoom [Nor91a]. Astronumerology [Cog08]. Astrophotographers [Bec14]. Astrophysical [Bar94, Doe90, Doe19, Dup90, Edd90, TPYR21]. Astrophysicist [Kid87, Ost95, Ost96, Ost97b, Ost98, Woo87, Kri88]. Astrophysics [Bec01, Cen17, Cro15, DeV05, Eis88a, Gin83b, Kai07, Ost80, Ost01, PdV85, Rot17, Smi18a, WW71, Wes90, Chi16, Nat13, Doe19, Hea15, Dev14a, Eis03, Kid87, SK03]. Astrum [Gin11a]. Athanasius [Row07]. Athens [Loc89]. Athribis [EP22, Win22]. athribitana [EP22]. Atkinson [Ano95h]. Atlante [Mos11]. Atlantique [Hos00b]. Atlas [Ash82, Dre15, Gin12a, Hos89b, KPG03, Rem08, Sch05a, War71, Gin10b, Gin12a, Kul80, War71]. Atmospheres [DK83b]. Atmospheric [Ext92, Lin07, Sch86, Sch13, dJ12]. Atomic [Gin01d, Ste12c]. Atoms [Lay02, Gin01g]. atque [Hos14a]. Att [Sha07a]. Attack [Edd82]. Attempt [McC99]. Attempted [CB05]. Attempts [Cha72]. attributed [Sha92]. Aube [Gra07a]. Audience [Gen90a]. Audomaro [Nor86, Nor86]. Aufsätze [Siv77]. Augsburg [dC01, dC01]. August [Ano78d, Gau10, Oes10]. Augustanus [dC01]. Augustus [McM01]. Aula [LN09]. Aura [Bla98a, Edm91, Bla98a]. Aureae [Kre07a]. Aurifaber [Gre10, Kre10]. Aurora [Far96, Sis85]. Aurorae [Dal80]. Ausbreitung [Wes97]. ausnehmende [Ham09]. Aussagen [Pin92]. Austin [SM04]. Australia [Ano73b, DH73, Mit19]. Australian [Hea94, Hea97, Sil23, Hea94]. Austria [Mes20, Mes21]. Austrian [Grö83]. Authenticity [Bow06]. Author [McK78]. Authors [Ped15, SC88]. Autobiographie [Kre86]. Autobiographies [Cor03]. Autobiography [DeV85a, Hen13, Hir09, Phi91]. Autograph [She87]. Autographensammlung [Oes15, RR13]. Autolycos [Pin81]. Autolykos [Sta72]. Autour [Ste12b]. aux [Cap12, Eva88, Eva01, Gin07b, Hos04f, Pin80]. auxiliary [Kin73]. av [Siv10]. Available [Bar19]. avant [Hos00b, Hen76]. Avebury [Fre77, TT76, TTF76]. Aveni [Ano01c, Bro90, Ald09, Ano95a, Ald09]. aventura [Gli88]. Avenue [TT76]. Avestan [Har79]. Avete [Hos14a]. Aviva [Bon19b]. Awakening [Gin75b]. Awesome [Lad10]. Azophi [Mal90, Kun87].

B [Hea15, Hen15b, Joh19, KF70, Mar23, Pet20, SIN85, NRKN16, Hol10a, Ost98, Kul87]. B. [Her71]. B.C [Cer97, Haw74, Mol93, Ols93, Osl06]. B.C. [Cul93, Ram07, Heg82]. Bīja [Pin96]. Bīja-corrections [Pin96]. BA [HK82]. Baade [SK03, Fea00, Ost95, Ost96, Ost97b, Ost98, SK03]. Babcock [Van14b]. Baby [Bon09a, Gin96d]. Babylon [Bri98, Eva02, Gol95, Ste99, Ste02a, Ste06a, Ste10a, Ste12a, Ste19, Doe05]. Babylonia [Gin92c, Hub91]. Babylonian [Bri98, Gin92c, Mer09, Oss12, Swe13, Bri05, CS02a, Fat98, FSAD99, Gau12, Gau14, Gau18, Gol02b, Har79, Hub91, Iwa03, Jon91a, Kol01a, Moe83, Mon21, Oss24, RC92, Ste01, Ste03, SG07, Ste16, SF93, SF94, SS06, Ste08b, Swe99a, Swe13, dJ12, Bri05, Ste99, Ste06a, Ste13]. Babylonians [SSM97, SS97b, Ste97]. babylonischen [Swe99a]. Back [ASCE14, Eva13]. Background [Gol97, Har72, Hoo87, Ped73, Sha12, SS96, Mer10]. Backsights [Rug82]. Backstaffs [Ree11]. Badajoz [HS99]. Baers [Dob77, Dob77]. Baghdad [Gin99b]. BaHena [Had84]. Bailey [Smi00]. Baja [vH86]. Baldewein [Oes10]. Balear [Iwa96]. Ball [BD82]. Ballena [Had84, Mul83]. Ballooning [Van95a]. Balloons [Van95a]. Balls [BD81]. Baltic [Omo17]. Baltimore [Ano89a, Gan15]. Baltinglass [Sto94]. Balts [SK97]. Band [Kin81, Sha71b, Swe76, Voe97]. Bang [Kra10, Smi98, Kra10]. Bank [Sha78, Sul02, HH84, Sul02]. Banned [Fin19]. Banneker [Kel74, Kel74]. Baptist [JLGL14, LGJLF14]. Baptiste [vH86]. Barcelona [Ano15f, Cha19, Gin99b, Hei14a, RA18]. Bard [CWW17]. Barnard [Bra90, Chr98, Mac10, Ost94, WDR11, Chr98]. Barnes [Joh19]. Barnhart [Kru17]. Baron [Mes11, Vet19, Pet20]. Baroque [Llo89b, Wer13]. Barriers [Het94]. Barringer [Hoy84]. Barrow [Pow95]. Bart [Smi94a, Smi94a]. Bartels [Oes15, RR13]. Based [Mer88, WS19]. Basel [Mos03, Kel78]. Basic [Bro90]. Basis [AH94, Jon71]. Bath [Bon09a, Eas89, Gin96d, Hos12c, Sal95]. Batlamyus [Gol73b]. Battista [Ser88, Swe72, Swe92b]. Battle [Rec20]. battlements [Sha07a]. Bavaria [Bre17, Her94, Not14b]. Bayer [Gin04c, Gin12b, Gin12b, Lat14, Swe86b]. Bayern [Her94]. BC [Bri05, De 14, MSH19a, Llo15, Bri05]. BC188 [MSH22a]. Bd [Her76]. Be [Bec99a, CT12, Rya09, Waf91]. Beach [Wil15]. Bearing [Péo08]. Beatrice [Eis01]. Beautiful [Gan15]. Beck [Wes97]. becoming [Sil23]. Bede [Pal12, SMH20]. bedeutender [Bre13]. Bedeutung [Gre00, Ste12a]. Bedini [Min00]. Bedside [Pas03a]. Before [Gol01a, Gre14, Whi70a, Bru96, Cro00, Fin19, Gra06, Gra76, Hos97a, Hos04e, Hos07a, Jak71, Oss12, Pin72, Sal85, Slu97b, Swe70, Swe86a, Thi88, Wil88]. Beg [Kri93, She90, Kul81]. Began [PP07]. Beginning [Smi81, Tip88b]. Beginnings [Bog02, Han09b, Mol93, Sim09, Tat99, War19]. Begun [Her77]. Behaim [Gin94]. behind [Gle04, Kin87, Sza18, How08]. Beihefte [Bur15]. Beijing [CJ20, CJ22]. Beispiel [Spr92]. Beiträge [Kle14, Voe03]. Beldomandi [Cha07]. Belief [Coy07]. Beliefs [Str16b, Het94]. Believing [Lip00]. Belknap [Dev14b]. Bell [Pow95]. Bell-Barrow [Pow95]. Bellarmine [Car97, She87, She87]. Bellarmino [Wes89]. Belles [Ler06]. belonging [Cha21]. Ben [Gol75, Gol77, Gol96a, Mea71, Nor76, Hos95a, RA18]. Bengt [Reb03]. Benjamin [Kel74, Kel74, McK78, McK87, McK78, McK87]. Benno [Bel14a]. Benson [Rem15]. Beobachtung [Swe99a]. Bequest [Dic91a]. Berechnung [Swe99a]. Berichten [Hor74]. Berkeley [Nau98, Swe17c]. Berlin [Hei14b, Roc17, Her22]. Bernard [Smi04, dB70, Kre80b, Kre81b]. Bernardino [Omo13]. Bernardo [Tur90a]. Bernhard [Eel83, Ost05b, Ano78a, Lep14, Ost05b, Eel74]. Bernoulli [Sha71b, Sha71b]. Berücksichtigung [Sta99]. Beruni [Ros74b]. Beschreibung [Hen92, Jor94]. Bessel [Kre86, Sta06, Wil86a, Wil86a]. Best [Bre17, Sal06]. Bethlehem [Gin81c, GHHB02, Sch15, AB13, Fle92, Gin81c, GHHB02, Nor08a, Sch15]. Betrayal [Sch07b]. Better [Smi07, Van15]. Between [Bon19b, Bon19a, Moz19, Bar18, Car21, Cha83, Chi09, Don88a, Duk03, Gin90c, Gin03a, Gin11c, Gra09, HB92, Hoy84, Kra95, Kre93, Loc90, Pal12, Sch14b, Thu03, Tir22a, Ban13, Bla97, Hos00c, Rem11, Wil11]. Bevis [KPG03]. bewegenden [Gin90f]. Beyond [Doe05, Smi06, Smi08, Smi09a, Cro12, NKM^{+}16, Sad90, Ste13, Ben87a, Edd74, Gol87a, Mor97]. Bhattotpala [Pin71a]. Bi [Bra20, Siv98]. Bi-Daily [Bra20]. Bialas [Mos07, Mos07]. Bianchini [vB10, Swe90, vB10]. Bible [Gin01f, Car97]. Biblical [Gin01f, Mar03, Rib24]. Bibliographia [Gin99e]. Bibliographical [Gin75b, Gin05e]. Bibliographies [Kin83, Waf86a]. Bibliography [Ano72b, Ano80a, Gin83b, Gin99e, Gin10a, Har75, Hos71a, Hos83b, Hos84b, Gin83b, Waf86a]. Biblioteca [Oos15]. Bibliotheca [Gin87c]. Bibliothek [Gau10]. bibliothèque [Feu75]. Biddell [Cha88, Cou18]. Big [Eva88, Hev94, Kra10, Smi98, Smi06, Sul99, Tim20, Yav99, Lan88, Sul77]. Bilans [Hos00b]. Bildern [Hor74, Kri94, Rut12]. Bildgeschichte [Ano06a]. Bilingüe [Omo13]. Bill [Car86]. Billard [Ano01j, Mer01]. Biman [Hea15]. Binding [Vai12]. Bio [Kin83]. Bio-Bibliographies [Kin83]. biografia [Gua09]. Biografie [Ost05b]. Biographers [Ree20]. Biographical [Gin85a, Hen92, Hos08a]. Biographie [Mos86]. Biographien [Gin05a]. Biographies [Ait82, Bec14, Dra84a, Gin85a, Hos82b, Sha87a, Tho22, Ree20]. biographischen [Hen92]. Biography [Big10, Chr98, DeV20, Din73, Don86, Gen90b, Har75, Lat93, Leq18, Orc88, Per00, Rea92, Sha87a, Smi94a, Tir22b, Wil86a, Gaa23, Don86, Gal83, Hos82b, Ken78, Kre09, Ler17, Low09, Tho71b, Tur71a]. Biological [Cor98, Str99]. Bird [Loc90]. Birjandion [Ans05]. Birr [Arm72, Ben14a]. Birth [Bla98b, Dan04a, Gin01c, GB86, Gre99, Hea15, Kra07b, Sch11, Sha14, Str93, Str19, VB21, Way96, Far05, Mel09, Nat13]. Biruni [Ken78, Ros74b, Sal79, Moz13a, Moz14, Rag83, Sal79, Rag83]. bis [Ano06a, Ash79a, Dob99a, Gin05a, Gre00, Har84, Hub83, Jor94, Kin81, Kri94, Mar75, Voe02]. Bisanzio [GS13, Lip15]. Bitrūjī [Sha92]. Bitruji [Ken73]. bizarre [LS24b]. Black [Doe08, Mon82, Sch01b, Sch79, Sch11, Sch07b, Mel09]. Blagrave [Tay10a]. Blazing [Gen90a]. Blick [Ost05b]. Blue [Mor97]. Blunt [Tur79b]. BM [Oss24, Ste02a]. BNP [LN09]. Board [For71b, FJ11, Ler80, For70b, For71c]. Bode [Jak72b, KH73, Ano15f, KH73, Sch14b]. Bodies [Far04, Sol09]. Bodleian [Azz11]. Bok [Smi94a, Smi94a]. Bologna [GB09, Not23a]. Bombay [Sch12a]. Bond [Ben20, Ste90a, POH96]. Bonelli [Ano82d]. Boner [Oos15]. Bonjorn [Hos95a]. Book [Ack05, Ack10, Ait82, Ait86, Ald97, Ald11, All71, And90, And00, Ano82b, Ano84b, Ano84a, Ans05, App02, App05, Apt89, Apt99, Apt14, Arm70, Arm72, Arm73, Arm74, Ash72, Ash79a, Ash79b, Ash82, Ash86, Ash87, Atw79, Aub12, Ave87, Ave94, Ave97, Ave09, Ave11, Axw17, Azz06, Bak15, Ban13, Bar05, Bar13b, Bat87, Bat90, Bec99a, Bec10, Bec15, Bec19, Bed80, Bed10, Bee71, Bel77, Bel07, Bel09a, Bel14a, Bel01a, Bel09b, Bel11, Bel14b, Bel19, Ben76a, Ben82a, Ben84, Ben87a, Ben87b, Ben91a, Ben91b, Ben93a, Ben99, Ben09, Ben13, Ben20, Ber70, BS73, Ber81, Ber89, Ber94, Ber97a, Ber01, Ber92, Ber97b, Ber12, Ber14, Ber19, Bia04, Big10, Bla98a, Bla91b, Bla97, Bla01, Bla79, Blu00, Boh19, Bok73, Bol05, Bol20, Bon76, Bon09a]. Book [Bon11, Bon12, Bon19b, Bon19a, Bow83, Bow10, Bra90, Bra15, Bra05, Bre00, Bre13, Bre11, Bri01b, Bri05, Bro83, Bro10, Bro90, Brü87, Bru78a, Bru82, Bru96, Bru09, Bry81, Ano81a, Bur82, Bur84, Bur15, Bur20, Bur80, Bur10, Cah19, Cam02, Cam11, Can12, Can09, CL05, CL08, Cap12, Cár88, Car76, Car86, Car97, Car99, Cau08, Cau17, Cen14, Cer97, Cer99, Cha00, Cha09, Cha19, Cha85, Cha89, CT12, Cha15b, CM10, Chr98, Cog08, Coh79, Coh80, Coh86a, Coh91, Coh97, Coh99, CS02b, CJ05, Coo75, Coo97, Coo00, Cor98, Cow71, Coy07, Cro15, Cro00, Cro08, Cro12, CMM^{+}15, Cro23a, Dan04b, Dan11, Dan12a, Dan12b, Dan19a, Dan19b, Dar91, Dav06, Dav08, De 13, DeV85b, DeV85a, DeV89, DeV96, DeV00a, DeV02a, DeV03, DeV05, DeV07]. Book [Dea94, Déb87, Déb91, Déb05, Déb13, Dek12, Dev14a, Dev14b, Dew70, Dew87, Dew00, Dic82, Dic90, Dic95, Dic97a, Dic99, Dic00a, Dic00b, Dic06, Dic09, Dic13b, Dij09, Dob77, Dob79, Dob83b, Dob83a, Dob86, Dob92, Dob96, Dob99a, Dob00, Doe94, Doe00, Doe01, Doe05, Doe08, Doe19, Dog87, Don71b, Don71a, Don73a, Don73b, Don82b, Don82a, Don86, Don88a, Don98, Don00, Don02, Don09, Dra80, Dra84a, Dra90, Dru84, Duk10, Duk11, Duk12, Dun92, Dun13, Dun20, Dup08, Dup14, Eag08a, Eag08b, Eas89, Eas96, Eas10, Edd71a, Edd74, Edd80, Edd82, Edd86, Edm84, Eel74, Eel83, Eis85, Eis88a, Eis88b, Eis96, Eis03, Ell08, Eng02, Eva82, Eva86, Eva87a, Eva88, Eva89, Eva93a, Eva00, Eva02, Eva05, Eva06, Eva07, Eva08, Eva09a, Eva09b, Eva11, Eva12, Eva13, Fal23, Fan05]. Book [Fan09, Fat00, Fei84, Fei86, Fei90, Fei95, Fei03, Fek13, Fin90, Fin99, Fin10, Fin19, Fle73, For70a, For73, For81, For82, For83, Gal83, Gan15, Gau10, Gau17, Gen87, Gen90b, Gen92b, Gen96, Ges19, Gib78, Gib82, Gie20, Gin70a, Gin72, Gin75b, Gin75a, Gin77b, Gin77d, Gin77c, Gin78a, Gin78c, Gin78b, Gin80a, Gin80b, Gin80c, Gin81a, Gin81b, Gin81c, Gin82a, Gin82b, Gin82c, Gin83a, Gin83b, Gin84b, Gin84a, Gin85a, Gin85c, Gin85b, Gin85d, Gin85e, Gin86a, Gin86b, Gin87b, Gin87a, Gin87c, Gin88a, Gin88b, Gin88c, Gin89a, Gin89b, Gin90c, Gin90d, Gin90a, Gin90b, Gin90e, Gin90f, Gin91a, Gin91b, Gin91d, Gin91c, Gin92b, Gin92a, Gin92c, Gin93, Gin94, Gin95, Gin96b, Gin96c, Gin96a, Gin96d, Gin97a]. Book [Gin97b, Gin97c, Gin97d, Gin98, Gin99b, Gin99a, Gin99c, Gin99d, Gin99e, Gin00a, Gin00b, Gin00c, Gin00d, Gin00e, Gin01c, Gin01g, Gin01d, Gin01e, Gin01b, Gin01a, Gin01f, Gin02a, Gin02b, Gin02c, Gin03a, Gin03b, Gin03c, Gin03d, Gin03e, Gin03f, Gin03g, GM03, Gin04b, Gin04a, Gin04c, Gin05b, Gin05a, Gin05c, Gin06a, Gin06b, Gin06c, Gin07a, Gin07b, Gin08, Gin10a, Gin10c, Gin10b, Gin11a, Gin11b, Gin11c, Gin12b, Gin12c, Gin12a, Gin13b, Gin14, Gin15, Gle04, Gle07, Gli88, God07, God13, God19, Goe09, Gol87a, Gol72, Gol73b, Gol87b, Gol95, Gol00, Gol11, Gol12, Gol03b, GG14, Goo09, Gra90a, Gra14a, Gra07a, Gra12b, Gra12c, Gra13a, Gra00, Gra01, Gra05c, Gre99, Gre00, Gre02, Gre04, Gre09, Gre14, Gua09]. Book [Gub88, Gur06, Had84, Hal88, Hal93, Hal03, Ham06, Ham09, Ham11, Han12, Han09a, Han09b, Har73, Har94, Har92, Har84, Har85, Has11, Has19, Hat83, Hat88, Hat04, Haw74, Haz02, Hea94, Hea97, Hea99, Hea15, Heg78, Heg82, Heg85, Hei00, Hei14a, Hei99, Hei08, Hei12, Hei14b, Hen75, Hen76, Hen08, Hen09, Hen13, Hen15b, Her72, Her76, Her80, Her90, Her94, Het76a, Het78, Het94, Hev94, Hew74, Hig12b, Hig12a, Hig17, Hil77, Hin04, Hir85, Hir92, Hir09, Hol03, Hol10b, Hol95, Hor74, Hor82, Hos71b, Hos72, Hos73, Hos75, Hos76a, Hos81, Hos82b, Hos83a, Hos83b, Hos84b, Hos86a, Hos88, Hos89b, Hos92a, Hos94, Hos95a, Hos95b, Hos96, Hos97a, Hos97b, Hos00a, Hos00b, Hos00c, Hos01a, Hos01b, Hos02a, Hos04b, Hos04c, Hos04d, Hos04f]. Book [Hos04e, Hos05a, Hos06a, Hos06b, Hos08b, Hos08a, Hos08c, Hos11a, Hos11b, Hos12a, Hos13, How02, How05, How08, How80, How88, Hoy84, Hub83, Hub91, Huf06a, Hug10, Hus13, Hut12, Jar00, Jar73, Jas96a, Jas96b, Jhi74, Joh19, Jon91a, JT92, Jon97, Jon99, Jon10, Jor94, Jus20, Kai07, Kak95, Kaz73, Kel10, Kel78, Kel83, Kel99, Kel74, Ken78, Ken82, Ken85, Ken87, Kes10, Kes15, Kid87, KH73, Kin79, Kin83, Kin87, Kin95a, Kle14, Kni70, Kra07a, Kra99, Kra07b, Kra10, Kra11, Kra13, Kra22, Kre80a, Kre81a, Kre82, Kre86, Kre89, Kre93, Kre99, Kre02, Kre05, Kre08, Kre13, Kre19, Kre20, Kri88, Kri94, Kri97, KI10, Kru83, Kru97, Kru08, Kul80, Kul81, Kul75, Kul87, Kun85, Kun88, Kun90, Kun91, Kus10, Kwa12, Lad10, Lan82, Lan84, Lan87]. Book [Lan88, LN09, Lás19, Lat93, Lat00, Lat12, Lat19, LF15, LC10, Lep14, Ler06, Ler11b, Ler17, Lia19, Lil99, Lin79, Lin93, Lin07, Lip15, Lip19, Lip00, Liv13, Liz17, Llo80, Llo89a, Llo90, Llo94, Llo09, Llo11, Llo15, Loc96, Loc99, Loc00, Löh19, Lov87, Lut91, Lüt92, Lüt00, Mac76, Mac93, Maf05, Mal90, Mal97, Mal13, Mal15, Mar75, Mar87, Mar97, Mar99, Mar03, Mar12a, Mar14, Mar12b, Max12, May11, McC82, McC83a, McC85, McC91, McC98b, McC99, McC11, McC19, McC13b, McC13c, McC08, McK78, McK87, McM01, McM03, Mea70b, Mea71, Mea81, Mea85, Mea86, Mea02, Mel89, Mer77, Mer88, Mer92, Mer96, Mer03, Mer06, Mer09, Mer10, Mer11, Mes11, Mes15, Mil80, Mil04, Min00, Mit86, Mit13, Mit19, Moe77, Moe80]. Book [Mol14b, Mol70, Mol76, Mol93, ML07, ML14, Mor97, Mor01b, Mor08, Mos07, Mos09, Mos10, Mos86, Mul82, Mul84, Mul80, Mul83, Mul81, Mül91, Mül10, Nak71, Nee71, Nee74, Nee08, Neu81, Nor76, Nor78, Nor80, Nor83, Nor85, Nor86, Nor88, Nor90, Nor91a, Nor92b, Nor92a, Nor96, Nor97, Nor99, Nor00a, Nor00b, Nor02, Nor04, Nor06, Nor07, Nor08b, Nor08a, Not19, O’B88, O’D06, Obr19, Oes10, Oes15, Ogj11, Olo04, Ols93, Ome79, Omo11a, Omo13, OT13, Omo19, Oos15, Orc84, Orc88, Orc90, Osl06, Oss12, Ost80, Ost92a, Ost97a, Ost04, Ost05b, Ost05a, Pal12, Pal77, Pal84, Pan09a, Pan09b, Pan15, Par86, Par90, Par92, Pas03a, Pas03b, PP07, Pas14, Pau78, Pau82, Pau88, Pau95, PG79, Ped73, Ped79, Péo08, Per00, Pet80]. Book [Phi91, Pin71a, Pin73a, Pin74, Pin80, Pin81, Pin84, Pin88, Pin90, Pin92, Pou75, Pou09, dSP78, Prz79, Rad90, Rag83, Rag85, Rag01, Rag09, Ram10, Ram06, Rap11b, Rap13, Rap15a, Rap11c, Rap20, Ras13, Rav72, Rav73, Rea81, Rea84, Rea92, Ree10, Ree11, Ree20, Rem08, Rem11, Rem15, Ren89, Ric07, Rob86, Roc08, Roc19, Rom06, Ron89, Ros79, Ros73, Ros81, Ros74b, Ros19, Rot80, Rot01, Rot03, Row07, Rud92, Rug07, RC11, Rut03, Rut12, Rut13, Rya71, Rya09, Sal79, Sal97, Sal09, Sam07, San99, Sar78, SS08, Sav12, Saw80, Sch07a, Sch15, Sch82, Sch91, Sch99, Sch05b, Sch00b, Sch03, Sch12b, Sch14a, Sch70, Sch11, Sch07b, Sel14, Ser88, Sev00, Sha07a, Sha78, Sha82, Sha85, Sha89, Sha99a, Sha08, Sha71b, Sha02]. Book [Sha14, Sha87b, Sha96, Sha70, Sha07c, Sha87c, She79b, She86, She87, Shi09, Shi11, Shu82, Sid10, Sil23, Sis85, Sis87, Siv77, Siv89, Siv98, Siv05, Siv10, Smi81, Smi85, Smi89, Smi92, Smi93, Smi94a, Smi94b, Smi94c, Smi98, Smi02, Smi07, Smi09b, Smi12, Smo09, Sol09, Sta72, Sta95, Sta07, Sta13, Sta99, Sta00a, Sta10, Ste99, Ste05, Ste06a, Ste12a, Ste13, Ste17, Ste77, Ste90b, Ste98b, Ste00a, Ste02b, Ste06b, Ste12b, Ste12c, Ste88, Ste96, Sti07, Str10, Str99, Str19, Str81, Sul77, Sul99, Sul02, Sul12, Sul14, Sut72, Swe73, Swe90, Syl07, Tat99, Tat09, Tau94, Tau96, Tau09, Tau11, Tau13, Tay10b, Ten00, Tes10, Thi88, Tho71b, Tho71c, Tho84, Tho86, Tho87, Tho88, Tho90a, Tho90b, Tih07, Tih17, Tip88a, Too96, Tri97, Tri03]. Book [Tri13, Tri10, Tuc00, Tur70b, Tur70a, Tur71b, Tur71a, Tur72, Tur79a, Tur84, Tur90a, Túr90b, Tur96, Tur03, Tur07, UL95, Vai12, Van95a, Van95b, Van99, Van02a, Van85, Van89, Van90, Van92, Van96, Van98b, Van98c, Ver13, Ver87, Ver09, Ver12, Ver17, Vet12, Vet19, Vic85, Vil90, Voe94, Voe01, Voe02, Voe03, Voe07, Waf86a, Waf91, Waf02, Waf04, Wag88, Wal77, Wal99, Wal13, War71, War79, War80, War93, War98, War08, War13, War19, Was77, Way96, Wea90, WW74, WW82, Wel84, Wel94, Wel05, Wen99, Wer10, Wer13, Wes77a, Wes84, Wes90, Wes96, Wes77b, Whi73a, Whi76a, Whi02, Whi72, Whi73b, Whi92, Whi00, Whi71, Wil15, Wil86a, Wil97, Wil02, Wil07, Wil87a, Wil14b, Wil86b, Wil87b, Wil95, Wil00, Wil01, Wil03, Wil04]. Book [Wil05b, Wil05a, Wis06, Wło10, Wło15, WW19, Wło23, Woo88, Woo87, Yan86, Yav99, Yeo88, Yeo99, Yeo00, Yeo02, Yeo07, Zbi97, Zep19, Zui19, Zuj11, dC01, dV91, vB10, vG05, vH86, vH87, vB17, vdW78, vdW90, Gra21a, SL13, Omo11a, Pin74, Rec22]. Bookish [Doo18]. Booklet [Gib78]. Books [Ano70e, Ano70f, Ano71i, Ano71j, Ano72f, Ano72g, Ano73j, Ano73k, Ano74h, Ano75g, Ano75h, Ano75i, Ano76h, Ano76i, Ano77g, Ano77h, Ano78h, Ano78i, Ano78j, Ano79h, Ano79i, Ano79j, Ano80b, Ano80i, Ano81b, Ano81f, Ano81g, Ano81h, Ano82c, Ano82i, Ano82j, Ano82k, Ano83a, Ano84d, Ano84e, Ano84f, Ano84g, Ano85a, Ano85b, Ano87e, Ano87f, Ano87g, Ano88e, Ano89i, Ano89j, Ano90f, Ano90g, Ano91h, Ano92f, Ano92g, Ano93f, Ano93g, Ano93h, Ano94g, Ano95f, Ano96d, Ano97g, Ano97h, Ano98g, Ano99f, Ano99g, Ano99h, Ano00f, Ano00g, Ano00h, Ano01i, Ano02g, Ano02h, Ano02i, Ano02j, Ano03f, Ano03g, Ano03h, Ano04h, Ano04i, Ano04j, Ano05c, Ano05d, Ano05e, Ano06b, Ano06c, Ano07f, Ano07g, Ano08e, Ano08f, Ano08g, Ano08h, Ano09f]. Books [Ano09g, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano12f, Ano12g, Ano12h, Ano13f, Ano13g, Ano13h, Ano15e, Ano15f, Ano16b, Ano16c, Bro90, Con70, Gin87a, Gin90d, Hen15a, Kru17, Mar03, Sch15, Sta72]. Borana [Rug87]. Borders [FV17]. Borelli [Bos09, Ros74a]. Borgia [Ano01c, Ave99, Ave06, Bri01a]. Borri [Car08]. Borrowed [Blo78]. Boscovich [Tuc00, Tuc00]. Boston [Ben20, DeV02a]. Bother [Ber81]. Boulder [Ave18, Rob86]. Boulliau [AH83]. Bound [Kwa12]. Boundaries [Ste05]. Boundless [Gra00]. Bowditch [Mon82]. Bracewell [Mit19]. Brackenridge [Nau98]. Bradlee [Che73]. Brahe [Don88a, Gin89b, Gle04, Kre05, Lat93, Péo08, Sha07a, Sha99b, Tho71b, Wes90, Wil00, CR22, Car08, GV98a, GV98b, Gin07a, Gra13b, Mes21, NKM^{+}16, Sch23b, Sch25, Sha07a, Tho71b, Van06, Wes78, Wes79, Lat93]. Brahmagupta [Pin71a, Pin71a]. Brahmin [DeV02a]. Brass [Gin91c]. Brazil [de 22]. Breakdowns [SO00]. Breaking [Dic06]. Bredekamp [Hei14b]. Bregon [Rod73]. Breifwechsel [Swe90]. Brengger [Mes20]. Brepols [Eva17, Fal23, Zep19]. Brera [Gin87c]. Brian [Bak15]. Brides [Sch12b]. Bridging [Tir22a]. Brief [Ano76b, Ano76c, Edm84, Gin05c, Gub88, Hat88, Hos82b, Rag09, Tau09, Wel84, Wil86a]. Briefe [Oes15, RR13, Ren89, Swe98, Brü87]. Briefen [Kre86]. Briefing [Yeo00]. Briefwechsel [Ano15f, Eel83, Sch14b, For81]. Bright [Lan88, Moz16, Mar75]. Brightest [Sch07a].

Brightly [Cha72]. Brightness [Gol96a, MD21]. Brill [Boh19, Jus20, Mar14, War19]. Brilliant [Gin05c]. Britain [Ash00, Dew00, Heg82, Hos01a, Bry72, TT78, Dic09, Heg78]. Britannica [Tho84, KPG03]. British [Gin05c, Hos04d, Lov87, dP82, Sul02, Tho84, Yeo99, Ash00, Bel09a, Dav19, Hos04d, LH05, Lov87, McC93, McC89, Mea81, Ogj11, dP82, Sym02, Win06, Bel09a]. Brittany [Hos03c, Hos07b, Hos07c, Hos01a]. Broadsides [Gin76]. Brogar [TT73b, TT75, TT77]. bronze [Jon21]. Broom [Hug10]. Broughton [Smi18a]. Brown [CT12, CT12]. Bruce [Nau98]. Brudzewo [Bar13a]. Bruin [Kin02a]. Bruno [Ano22b, Blu00, Pan09b, Sha99b, Tes10, Blu00, Blu21, Don71a, Gra08b, Lüt10, Omo19, Pan09b, Tes10, Don71a]. Buch [Rut12, Gin91a]. Buenos [Nor91b]. Bugge [Hut12]. Build [Ano05a, Cha72, JLS^{+}05, JHLS05]. Build-up [Ano05a, JLS^{+}05, JHLS05]. Builder [Gin02d]. Building [Bre17, Doe19, Gra12e, Hev94, Kra20, Pau88, War98, Yav99, Mul82]. Buildings [RPPA15]. Built [Tat09]. Bull [vdW90]. Bunge [Ano81a]. Bureau [CT86, Gre23]. Bürgi [Gin80b]. Burial [BEC^{+}99b, HN91]. Burnett [Boh19]. Bute [Ano74d]. Butte [SS86, Zei85a]. Byzantine [Cau17, Eva89, Tih17, Cha00, Dob86, Eva89, JT92, MHS23, Pin98, Tih21, Eva01]. Byzantium [Eva01, Sal90].

C [Ber01, Bre18, For81, Hos00b, Kin07, McC19, Nau08, Pau86, Rad90, Rob86, Siv10, Wes97, Zep19, Ano15f, BH12, Cul02, Eas97, HK82, KD98, Not24b, SB72a, SB72b, Sha12]. C. [Ano73e, Dew87, Eva87a, SCV20]. C.E. [TQ18]. c/5 [Ano15f]. CA [Swe17c, Gra05c, Kin81, Kri94, Sha08]. Caballeros [Gli88]. Cacciatore [Rap11c]. CAD [Smi18a]. cadrans [Par85]. Caelo [Tau11, Tau11]. Caesar [Cer97, Cer97, Han09b]. Caesareum [Gin71]. Cahokia [Rom06]. Caille [Déb13]. Cairns [TT73b]. Cairo [Vil90]. Caius [DL15]. Calcul [Dob86]. Calculated [Ste02a]. Calculating [Car21, Gin91b, Gin01c, Kre10, CML23]. Calculation [Ano24a, Mon82, MS04, Wil93, ZH12, Sho21, Sal06]. Calculations [BSL86, Jon16, Kel99, LS05, LZ99, Gin91b]. Calculator [Gin91b, Gin91b, McC82, McC82]. Calculators [McC82, Gin91b]. Calculs [Siv10]. Calendar [Ash86, BMA90, Far93, Fat00, Gin99c, Gra11a, Han09b, Hen15a, HS16, HS18, Joh19, Kel99, LZ99, Loc00, McC19, Mer98b, Nor83, Omo17, Ped15, Rug87, Rüp15, Swe74a, Swe86c, War13, Bir74, CML23, Dic21, Gin99d, McC13a]. Calendar-Maker [Omo17]. Calendar-Making [Rüp15]. Calendaria [Voe97]. Calendars [Ave92, Bie21, Bro90, Gin78a, Gra90a, Han09b, Har79, Kre20, Lév17b, Loc96, Mar12b, McC90, McC13a, Pet80, Ric07, Sid10, Gin78a, Mar12b, Roc08]. Calendrical [Kel99, Mal78, McC19, Nor92a, CML23, Kel99]. Calendrics [Roc08]. Calendrier [Siv10, Bir74]. California [Had84, Mul83, Nau98, Swe17c, vH86]. Call [Sal06]. Callahan [TH79]. Callanish [CFMR77]. Called [Wes84]. Calvinist [Hal03]. Cambridge [Bel14b, Dan19b, DeV00a, Dev14b, Dob00, Gin00e, Gin12a, Hen15b, Kra22, Rad90, Rot17, Sul14, Too95, Van98c, DL15, Dup90, Gen90a, Vic85, Sha78]. Came [Gin88d, Had84, Waf91]. Camera [Edd71b]. Camille [She14, Cro15]. Campaign [GV98a, GV98b]. Campaigns [Goo09]. Campani [Tur84]. Campanus [Swe73, Swe73]. Campbell [Ost89]. Canada [Smi18a, Bat90, Jar75, Jar77, Smi81, Smi18a, Smi81, UL95]. Canadian [Bat90]. Canaria [ESBG96, ESBG97]. Canario [Iwa96].

Cancri [Sch23b]. Candrārkī [KMD^{+}18]. Canon [Azz11, Ols93]. Canons [Hus12, Nor85]. Canvas [Ste00a]. Cape [Eva86, Eva87a, Hea99, Hos04b, Rot80, Wel94, Eva86, Eva87a, Hea99, Hos04b, Hos08c, Rot80, War78, Wel94, Hos08c]. Capella [Eas00]. Capital [FV17]. Capponi [GM03]. Capri [Fal21, Nor78]. Capricorn [Orc90]. Captain [Pau88]. Cardano [Azz11]. Cárdenas [vH86]. Cardinal [Das15]. Career [Edd86, She99, Smi22]. Careers [DeV98]. careful [MD21]. Carl [Ano15f, Sch14b, Sch14b]. Carnac [Fre75, TT71, TT72a, TT72b, TTG76]. Carnegie [Dic06, Pau95, Nee91, Pau95]. Caroline [Bec15, Hos14c, Kra11, Cor03, Hig14, HW81, Hos05b, Hos05c, Hos06c, Hos06d, Hos14b, Tur77, Cor03]. Carolingian [Cau08, McC04, Pal12, Cau08, Dek08a, Eas00, Eas01, Obr19]. Carried [Slu97a]. Carrington [PP07]. Carteggio [Cen14]. Cartesian [Whi73b]. cartésienne [Lüt00]. Cartographer [Kle14, Wil14b]. Cartography [Van02b]. Carved [Ave05]. Case [Bec19, Cle71, Dek08a, GC08, Gra12d, Lan83, Rav73, Sch70, SB84, Tay10b, TW25, GB09, Sch12a]. Caspar [Gau17, Gin77c, Hol95]. Cassini [Van96, BS92, Fab19, Gra21b, Leq18, Van96, VvdS19]. Cassiopeia [Ash80]. Castelli [Sie05]. Castigator [Jar06]. Castilian [Cha19]. Castille [Pou88]. Casting [Fin08]. Castle [Cha15b, Mol14a, Ben14a, SW18, Arm72]. Castlehill [Rei82]. Catalan [Hos08b]. Catalans [Ano15f, Hos08b, Hos06b]. Catalog [Gin75a, Gre02, Gre04, Gub88]. Catalogo [Gin87c]. Catalogue [Ano76c, Ano01b, Con70, DE00, Dob02, Duk02, Dun20, Eva87b, Eva87c, Eva92, Gen87, Gin70a, Gin75a, Gin77d, Gin81b, Gin87a, Gin90d, Gin03d, Gin04a, Gin10c, Hos06c, Kri93, Kun86b, Kun91, dP82, Ram07, Ree11, SC88, Saw80, Sch01a, Sch05a, Sch14a, SG09, Swe86b, Swe92a, SK16, Tau94, Tho84, Tur70a, Wło90, Wri07, Eva92, Gin04b, GWZ22, Hos11b, Kun91]. Catalogues [Duk03, Gin87c, Gin04c, Sch13, Zso94, Zso20, She90]. Cataloguing [Hos11b]. Catalogus [Gin87c]. Catalonia [Hos06b]. Catalunya [HP98, HP99]. Catastrophist [Car86]. Catch [Vet19, Whi08]. Cathedral [Bre00, Pou83]. cathédrale [Ano83b, Pou83]. Cathedrals [Bre00]. catoptric [Nal22]. Causality [Bon07]. Cause [Blu00]. Causes [Gre99, Kre08, Cro23b]. Cautious [Eis01]. Cave [FJ11]. Cavellat [Gin88b]. CD [Bur15, Eva00]. CD-ROM [Bur15, Eva00]. cdix [Cro15]. CE [MHS23]. Cecco [Don79]. Cecilia [DeV03, DeV85a]. Cedillo [Cha19]. Celebrate [Ano84b, Woo80]. Celebrating [Kra22]. Celebration [Ano84b]. Celebrations [Gin97a, How88, Sne98]. Celeste [Ste12b, Can09, Hos89b, Wil03, Wil95, Van94b]. Célestes [Cap12]. Celesti [Omo11a]. Celestial [Ack18, Ave11, Bec19, Ben20, Bol05, Bon19b, Bon19a, Bra20, Bur80, Cam11, Dup03, Eva02, Gin85e, GH07, Gra06, Gra11b, Hug10, Jar00, JR10, Kun88, LG12, Mor97, Qua17, Sal94a, Sch03, Ste00a, Swe13, TB13, Cul17, Fek24, HGGVE^{+}23, Hos02a, Oos15]. celestiales [Cha19]. celestis [Ste88]. Cellarius [Rem08]. Cemetery [PHP92, PH96, PP82]. Census [Dan04b, Gin06e, Mos03, Pou85, All71, Sar78, Sha96, Shu82]. Cent [CL05]. Centenary [Hos92a, Kra22, DeV03]. Centennial [Ano84b, Ano84a, Dic95, Dic06, Gin84b, Smi12, HE82, Woo80]. Center [Doe19, Had84, Ost92b]. Centers [Mül91]. Centiloquium [Bou23]. Central [Nau98, BG12, Hos02c]. Centrality [Fei86, Sza18]. Centre [Ave95, Lüt10, Moz14]. Centred [Eas01]. Centuries [Bro88a, CG97, Dic00b, Gin90d, Hen15b, Kre20, Lil99, Mul84, Sch14a, Sta95, TQ18, Tur70a, UC16, WW74, Yeo99, Hea14, MHS23, Bro91b, Sha87b, Sta95]. Century [Ack05, Apt14, Arn11, Bec10, Bec15, Ber97a, Bla98b, Cha15b, Cor98, Dun92, Eis85, Eva89, Eva09b, Gin75b, GM03, Gin11b, Goe09, Gol03b, Gra21b, Has11, Hin04, HS18, Hol10b, Hos04d, Jar90, Kai07, Llo15, Min00, Nal16, Nor92a, Ost80, PdV85, Pou09, RC11, Str99, Sut72, Ten00, Waf02, Way96, Whi73a, vH87, vdW17, Car23, Cha21, CMP24, De 14, For16, GJ22, Maa16, Mol14a, NKM^{+}16, Not23a, Sel12, Sho21, de 22, Aub16, Bau87, Ben84, Bre13, Cha98, CG10, Che72, Chi15, Con70, CMM^{+}15, CO13, Dan12b, DOL98, DS89, Far04, FV17, Fle92, GHS08, Gen87, Gin76, Gin83a, Gin97c, GM03, Gin13c, Gre00, Grö83, HH07, Has19, Her73, Hom72, Hos95a, Hos04d, Hos14c, Hus12, JT92, Kai07, Ker19, KG82, Kin87, Kin95a, Kok98]. Century [Kre89, Kre11, Ler80, LG12, Liv16, McC93, Mea85, Mol76, Nas13, Nor80, Nor83, Not14b, Not15a, Not15b, Pau86, Pin98, Rav72, RS18, RC11, Rus74a, Rus74b, Sal94a, Sal94b, Sam98, Sha99a, Smi91, Smi97, Spe13, Ste90b, Str10, TG87, Tur02, VG01, War98, Wer13, Whi72, Wil93, Ano81a, Kes15, Mer11, Ben87b, Dan12b, Dar91, Eas89, Fei86, Gin77c, Gin87a, Hin04, Péo08, Rot80, Smi89, Tur72, Ver17]. Century-Long [Eva09b]. Century-Old [Ber97a, Gra21b]. Cepheid [Wil07]. Ceragioli [Bre18]. Ceres [CMO11, Cun16, For71a, Mir17, Mir17]. CESLA [Ave95]. Châtelet [Nau16]. Chabás [Fal23, Gra24]. Chacona [BAE93]. chair [Glu22]. Cham [Cro23a, Gee18, Löh19, Mit19, Sil23]. Chamber [Tur79a]. Chamberlain [Had84]. Chamberlin [Bru78b, Bru78c]. Champion [Wag92]. Chandra [Sch07b]. Chandrasekhar [Smi96, Smi96]. Change [Cha90, Cul93, Kre13, Sch01c, Kra24b, Mar12a]. Changed [Bia04, Wil05b, Zuj11]. Changes [SS90]. Changing [De 06, DeV07, McC03]. Channel [Le 08, Le 09]. Chaos [Hol03]. Chappe [vH86]. Chapter [Ans05, Tur84]. chapters [Swe86a, Pin03]. Characteristics [Cha12]. Characters [DeV89]. Charge [Van97]. Charles [Boh19, Cha15b, Hos88, Plo90, Tur07, DeV90, Hod77, Hos88, Eva87a]. Chart [Sta06]. Charting [Sch00b]. Charts [Ash87, GH07]. Chase [Ber94]. Chasing [Dan04b, Duk12, Dun13]. Chaucer [Gra90b, Mea71, Gra90b, Mea71]. Chaucerian [Dav19]. Cheaper [Smi07]. Checklist [Ano74a, WW71]. Chéseaux [Jak70]. Chet [Ano15f]. chetverti [Rya71]. Chicago [Bon19b, Bra15, Dun20, Kes15, Rap15a, Sha14]. Children [Hig17]. Chimney [MEA91]. China [Ano03a, Gin98, Gin03a, Has11, Shi11, Siv77, Siv98, Tur71b, CL22, Cul93, FS96, FV17, Gin96c, Has87, Has11, Li19, Lia19, Nee74, Nor91a, Shi00, Siv05, Siv10, SS98a, SF95, Tur71b, Yor01, Lan89]. Chinese [Har72, Nee71, Pan09a, Qua17, Siv05, Cul02, Cul05, Cul17, Lan89, LZ99, MSH21, Nak71, Nee71, Pan09a, Qua17, Siv77, Siv98, Ste98a, SB12, SMH18, TQ18, TW25, WS16, WS19, Yan86, ZJJ^{+}99]. Chinnici [Cen14, Cen17]. Chinois [Siv10]. Chinoises [Yan86]. Chioniades [Sal90]. Chios [Wil08b]. Ch’iu} [SY92]. Chr [Wło23]. Chris [Lau14]. Christ [Gin01c, Sha11, Nor08a]. Christiaan [Ano78d, Ben82a, Kra22]. Christian [For81, Gin01c, Gra00, Mal97, Nor92a, Eas97, McC11, Nor92a, Obr19, RS18]. Christianity [Gin90c]. Christiansen [Mit19]. Christine [Mal15, Roc17]. christlichen [War13]. Christmas [CJ20]. Christoph [Bla97, Kre07b, EZ16]. Christopher [Fei84, Qua17, Ben75]. Chronicle [SY92, Car25, Bot76, MSH22b, OP89]. Chronicles [SS91a, Ste77]. Chronicling [Ost04, Way96]. Chronography [Nor92a]. Chronological [Mar12b, Nor92a, Ost92a].

Chronology [Kaz73, Mal78, Mer03, Nor92a, Sch00a, Ste12a, Swe74b, Ver12, CS21, Rib24, Kaz73]. Chronometern [Bre13]. Chugoku [Nak71]. Chumash [Had84]. Ch’un [SY92]. Ch’un-Ch’iu} [SY92]. Church [Bre00, Bro96, Gin05d, Hos04c, Mul82, Wes96, Gol01a, Hos04c, Sch91]. churches [UAGGB21]. Ciência [LN09]. Ciel [Bla01, Kus10, Mol70, Ver09]. Cielo [Hos00a, LS24b]. Cielos [Gra12b]. Ciencia [Gli88, Hos06b]. científicos [LN09]. Cimento [Tur84]. cimétiére [PP82]. Cingitur [Van74a]. Cinquecento [Swe92b, Pan09b]. circa [Her22]. Circle [And90, Bel20, Ben91a, Lüt10, Sta99, Wło23, Meu10]. Circles [Chi09, Hos01a, PF83, Ree11, Gin87b]. Circular [Gin87b, Llo15]. Circulation [CK07]. Circumference [Lüt10]. Circumpolar [Loc85]. Circumsolarity [Bra20]. Cities [Kin95b]. Citizen [Smi12, Tur75]. Citizen-Scientists [Smi12]. City [Spr93a]. Civic [Bel07]. Civil [Bro10, Dew87, Swe74a]. Civilizaciones [Hos00a]. Civilization [Ave11, Sal06, BF13, Ave11, Pan15]. Civilizations [Hos83b]. Civilizing [Dea94]. Clairaut [Wil93]. Clarendon [Smi96]. Clark [Dew70, Gin97b, Dew70, Gin97b]. Clarke [ML14]. Clash [Eis88b, Wes89]. Classic [Ald02, Cog08, Smo09, Vai12, Ada21, Ler25]. Classical [Eva11, Han09a, Mon82, Plo16a, Sid10, Mol14b]. classics [War15, Mit86]. Classification [HB71b, HB71a, Har14, Pin70, Sym15, Tee89, Zso10]. Classique [Gra07a, Llo80, Wil87b]. Classroom [Rap15b]. Claudius [Moe80, Too77, She15, Eva92]. Claus [Wło23]. Clavius [Bla97, Bla97]. Cleomede [Eva05, Neu81]. Cleomedes [Neu81]. Clerk [Wil05b]. Clerke [Eis03, Eis03, Gin01b]. Clerks [Tur71b]. Clifford [Mir17]. Climaci [GH24]. Climb [Ost89]. Clive [Dun23, Mag23]. Clock [Loc83, MS04, MS05, Oes10, Pou83]. Clockmaker [Bre13, Syl07]. Clockmaking [Ben20, Hai21]. Clocks [Ano77l, Bro90, Gin85c, Gin86b, How86, How94, Lan89, Loc92, Loc96, SS98a, Str81, Sym02, Gin86b]. Clockwork [Bry81, Dic99, Lan89, Str81, Tur90a]. Close [Gin87a, Li19, Yor01]. Clouds [Ash79b]. Clube [Car86]. Clues [Ano04b, BGP13, ZB04]. Clusters [Hos08g, Hos11b, Hos13, Smi00, Pop22, Gin70a]. Clyde [Rea92, Rea92]. CO [Ave18, Rug95]. Co-Operation [Rug95]. Coast [HGGVE^{+}23]. Coat [Hos16b]. Cocktail [Mes15, Fre14]. Code [Dic06, Sch11, Tur79b, Mel09]. Codex [Ano01c, Ave99, Ave06, Pin74, Vai12, BBW97, Bri01a, Kel83, Kno03, Mar95, GH24]. codice [GS13, Lip15]. Codices [AMP95, AMP96]. Codicum [Gin87c]. Codified [Hos01a]. Coelestis [Tho84]. coelestium [Eva00, Cha19, Wil86b]. coelum [Ham06]. Coffee [Gin16b, Rem15]. Coffin [Loc83, Loc92]. Coffin-Lid [Loc83]. Cognition [Van89]. Cognitive [How02, Wes09b]. Cognizing [Wes09b]. Cohen [Smi04]. Coherent [Ber19]. Coignet [Mes20]. Coin [SM05]. Coincidence [Bon07]. Cold [Bat90]. coleccões [LN09]. Colégio [LN09]. Collaboration [Cer97]. Collapse [Bla97]. Collectanea [Gin78b]. Collected [Bel14a, DeV85b, Gin93, Mes11, Min00, DeV85b]. Collecting [Mol14b]. Collection [Gin87c, Gin99a, Lat20, Sch14a, Sta04, Tau94, Tur70a, SG16]. Collections [Gin77d, Gub88, Sch14a, Tur70a, Tur79a, WW71]. Collectives [Hos00b]. Collector [WW82]. College [Bel77, GDEK24, Had84, Mac90, Rob90, Rot90, Ste90a, DL15, Mor80, Spe80]. Collide [Ave95]. Collimitius [Sha99a]. Collinder [Ano76j]. Colloquia [Jar73, Nee74]. Colloquio [Rug95]. Colloquium [Llo80, Siv89]. Colluding [Top03]. Colonial [Hea99, dA17]. Colorado [Ave18]. Colour [Gin10b, Hoc92]. Columbian [Ave18, Car76, Ano73a, Ave18, Hiv81, Kru22, Spr92]. Comet [Bro85, Car17b, Cer97, Don73a, FF20, Fle92, Gin86a, HW81, Kok81, Mar87, OP89, POH96, Sei17, Waf86a, Car25, NKM^{+}16, HD84, JKKC19, OP92, Waf86b, Wil93, Yaz14, Rec22]. Cometa [Gin82c]. Cometarum [Ruf10]. Cometary [Hen09, Jas20, Ost92a, Ruf71, vH87]. Cometas [Hei14a, Omo13, Gra12a]. Comete [Str10]. Cometography [Gre02, Gre04, Gre02, Gre04]. Comets [Bos09, Bur10, Chi15, Coo99, Dal80, Gen92b, Gin20, Gin86a, Gra20, Gre99, Gre00, Gre02, Hei14a, Hug10, KKSS21, Mar75, Omo13, Ram07, Sch03, Shi11, Str10, Wil08b, Yeo99, Zso10, vH87, CL22, Cro23b, Gen92b, Gre99, Gre04, Ost92a]. Coming [CT12, McM89, Nor08a]. Coming-to-Be [CT12]. Commemorate [Ash86]. Commemorated [Cha89, Dic95, Gin03f, Sha08, Woo87]. Commemorating [Sha89]. Commemorative [Rag83]. Commensuration [Hen15a]. Comment [Duk05]. Commentaire [Eva88, Eva01, Pin80]. Commentaria [Don00]. Commentaries [Byr11, Gin96c]. commentario [Omo11a]. Commentariolus [Gra80, Dob73, DS89]. Commentary [Bar13a, CS21, Duk11, Eva88, Hos97b, Lat00, Lin07, Pin71a, Pin74, Pin80, RA18, Sal79, Don93, Knu14, Swe86a, Swe17c, Fal23, Gan15, Zep19, Bel09b, dSP78]. Commentators [Lat00]. Commented [OT13]. Commenting [Lat00]. Comments [Atk75, Tur77, TH79]. Commercial [Sim09]. Commission [Ano70b, Ano95g, Gin82d, SGDO02, Ano71e, Ano76d, Ano79b, Ano89a, Déb95, Dic97b, Dic00c, Nor91b].

Committee [Ano95g]. Common [BS92, Hen15a, Smi96]. Commonly [Kun86b]. Communities [Bru96]. Community [DeV98, Pau81, Péo08, Smi18b, Sul14, Mun13, Bon19b]. Companion [Gin00e, HW07, WR08]. Company [Sch12a]. Comparative [GZB14]. Comparison [Hei12, TQ18]. Compass [Gin03e]. Compatibles [Ols93]. Compendium [Jon91a, Oss24]. Compilation [Ste19, Wło90]. compiler [Fal23]. Complete [Cul02, Hos13, Kno90a, Kri93, Pou96, Smi09b, Low09]. Completed [Fei03, Gin08]. Complètes [Wil03, Pou96]. Completing [Gat20, Kre19, Sha22]. Complex [Spr93b, Spr93c, Swe17c]. Composing [Ste02b]. Composita [Dav09]. Composition [Siv05, Ste02b, Swe74b, Whi78]. Compound [Van77]. comprehensive [Kra24a]. Compromises [Mac90]. Computation [Kre12, Nee71, Whi75]. Computational [CG97, GC08]. Computations [Jon91b, Ter84]. Computer [CS02a, Duk04, Eva09b, Fat98, Muc75, Ols93, Van98a, Gin91b]. Computerized [Ano74a]. Computers [Mul20]. Computing [CG13, Kre11, Ols93, Siv10, Ano13a]. Computistical [McC03]. Computistics [Not14b, Pal12]. Computus [Gra05a, Loh13, McC11, Pal12, Wal13, Bur15, Wal13]. Comune [Bel11]. Con [Swe92b, Omo11a]. Concave [Ost86]. Concentric [Swe99b]. Concentric-Sphere [Swe99b]. Concept [Cle71, Jhi74, Kru97, Smi92]. Conceptions [Hos04f, Spe13, Ell08]. Conceptual [McC91, Wil88]. Concerning [Tur15, Cer95, Gin85g, Gin97c, Moo94, Rus02]. Conches [Bra20]. Concise [Ano75c, DeV00a, DeV05, Bec15, Hos82b]. Conclusions [Hos08h, VMV18]. Concordance [Gin00b]. Condamnation [Fel02]. condensation [TH79]. Conducting [Dan04b]. Conference [Ano94a, Ano96a, Ash86, Car00, Jas96b, Kre01, Liv13, Mel89, Mit73, Tho90b, Wal13, Wil87a, Tri10]. Conferences [Sal97, Whi76b]. Confession [Bon19b]. Confessional [Joh19]. Configuration [Dek08a]. Confirmation [Gin81c]. Conflict [Eis88b, Jar90, Lon06]. Conflit [Gin07b]. Confounded [Cas19]. Congregation [Fel02]. Congress [Ano79b]. Conical [Loc89, Loc95]. Conjectures [Pal77]. Conjunctions [Li19, Yor01]. Connection [Jar90, Rec20]. conoscere [Ano06a]. Conquest [Hos02a]. Consensus [Ben92]. conséquences [Hen76]. Consideration [PC89]. Considerations [Don79]. Considered [Blu00, Rap11b]. Constantinople [Has11]. Constellation [Ave09, Loc90, Kur21]. Constellations [Dav85, Dek20, Dob00, Gin72, Kun87, Kur19, Loc85, LB06, Mal90, Obr14, Obr19, Pal77, Sch02, Sch05a, Sol09, Dob00]. constitutione [Ste88]. Constructed [GvH11]. Constructing [Lat20]. Construction [ASCE14, Ben13, Gin01a, Kra99, Kun88, RA18, SW18, Tur70b, Hen97]. Constructions [Ave18]. Contacts [Gin03a]. Contained [Con70]. Contemplation [Cap12]. contemporaine [Har85]. Contemporaries [Ben87b, Mil04]. Contemporary [Bar05, Jon76, Lat12, UCI16, UC22, UC23]. Context [Ash86, CM10, Dob96, Eis03, Gin89c, Hal93, Hay81, Hos03c, Hos04b, Hos08b, How02, JLGL14, Kra07b, Lat20, Llo94, Loc00, Mer10, Mol93, Nee08, Ree11, Row07, RC11, Sha07a, Tho86, Wal13, Mit13, McM03]. Contexts [Gre09, How02, McC03]. Conti [Tuc00]. Continuation [SMH20]. Continue [Car17b]. Continuity [CM17, Gin96d, Mar12a]. Contra [Jar06]. Contraction [TW25]. Contrasts [Tur07]. Contribution [Bar88, CS02b, Hoy82, Kri88, Nau01].

Contributions [For70c, For80a, TG87]. Contributor [Ano88a]. Contributors [Ano70c, Ano70d, Ano71f, Ano71g, Ano71h, Ano72c, Ano72d, Ano72e, Ano73g, Ano73h, Ano73i, Ano74e, Ano74f, Ano74g, Ano75d, Ano75e, Ano75f, Ano76e, Ano76f, Ano76g, Ano77c, Ano77d, Ano77e, Ano77f, Ano78e, Ano78f, Ano78g, Ano79e, Ano79f, Ano79g, Ano80c, Ano80f, Ano80g, Ano80h, Ano81c, Ano81d, Ano81e, Ano82f, Ano82g, Ano82h, Ano83d, Ano83e, Ano84c, Ano86a, Ano86b, Ano87a, Ano87b, Ano87c, Ano87d, Ano88b, Ano88c, Ano88d, Ano89e, Ano89f, Ano89g, Ano89h, Ano90b, Ano90c, Ano90d, Ano90e, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano92b, Ano92c, Ano92d, Ano92e, Ano93b, Ano93c, Ano93d, Ano93e, Ano94d, Ano94e, Ano94f, Ano95e, Ano96c, Ano97d, Ano97e, Ano97f, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b]. Contributors [Ano99c, Ano99d, Ano99e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03b, Ano03c, Ano03d, Ano03e, Ano04d, Ano04e, Ano04f, Ano04g, Ano05b, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12b, Ano12c, Ano12d, Ano12e, Ano13b, Ano13c, Ano13d, Ano13e, Ano14c, Ano14d]. Controversia [Nor92b]. Controversial [CMO11]. Controversies [Mar89]. Controversy [Bak15, Bau87, Big10, Bon12, Din73, GMML14, Gin03h, Gin11c, She99, Smi98, Bak15, Gol87a, Voe94]. Conversations [Liv16, McC07]. Conversion [Fat98, Fat00, Mer98b]. Convinced [Car17b]. coolest [Del23]. Coon [Mar89]. Copán [Ald02]. Copenhagen [Wło23]. Copernic [Gra80, Hen76, Pou96]. Copernican [Bar05, Fin19, Hal88, Har75, Kra07a, Mar03, Par92, Ped73, Voe94, Wal77, Ano83b, Bal85, Car24, GV98a, GV98b, Gin06e, Gra12d, Gra13b, Gre10, Hal88, Har75, Het76a, Jar82, Kra07a, Kre10, LG12, Ler25, OT13, Rag05, Tim25, Wes09b, Wes18, WKH18]. Copernicana [Ait82, Jar73, Nee74, Swe98, Ano74b, Nee74]. Copernicanesimo [Fin99]. copernicanischen [Wes97]. Copernicanism [Bro96, Fei86, Hos04c, Bau86, Dan04a, Fin99, Kus10, McM89, Rap15b, Rav72, Rot09, dA17, Van89]. copernicanos [dA17]. Copernicans [Hal03]. Copernicienne [Kus10]. Coperniciens [Fel02]. Copernico [Omo11a, Cha19]. Copernicu [Mos03, Pou96]. Copernicus [Ait82, App05, Bar19, Ber94, Bla97, Dan12a, Dob96, Gin89c, Gin03f, Gin05a, Het76a, Kno90a, Kre93, Kre07b, Mar14, Moe77, Nor94, Ped73, Pou96, Ros74a, Swe88, Swe90, Swe98, Too81, Wal77, Wes93, Ait82, Ano97b, App05, Bar19, Bar99, Bar05, Blå14, Cha19, CB05, Dan12a, Dob73, Dob79, DS89, Dob96, Dob01, Eva00, Fei86, Fel02, Gin81d, GW81, Gin85f, Gin89c, Gin05a, Gin08, God06, God07, God15, God19, God24, Gol94, Gol02a, Gra80, Gra10, Hal03, Has19, Hen76, Hil77, HG04, Hoo87, Jar73, Juz06, Kno90a, Ler00, Mar14, Mor14, Mos03, Omo11a, Ped73, Pou96, Rag16b, Rav72, Ros79, Ros73, Ros86, SS08, SL13, Sev00, Sha17, Sha22, She79b, She79a, Shr00, Str16a, Swe88, Swe17a, Swe17b, SW18, Tho86, Too81]. Copernicus [Ver06, Ves73, Wes97, Wło07, dB95, vN12, Bar05, Dan04b, Fin19, Gin87b, Mar14, Moe77, Osl06, Ros79, Ros73, Tho86, Wes77b]. Copernicus-Biographien [Gin05a]. Copies [GW81, Hei12]. Copy [Hei12, Mor14, She87, Str16a, Dob73]. Córdoba [Gra12b, Rie13]. Corfu [PH94]. Cork [PF83]. Cornell [Tay10b, Tay10b]. Corona [Bec00, Pan94b, Pan95, UC22]. coronal [RS22]. Coronelli [Ben18, Mos11]. Corporation [Vet19, Bra15]. Corpus [GC08, Siv05, Lip15, MD21, Ano13a, CG13]. Correcte [Cha15a]. Corrected [Fre75, Sch13]. Correction [Hen15a]. Corrections [Gol84, Pin96]. Correlations [Sch84a, Sch84b]. Correspondance [HRVM13, Eva17]. Correspondence [Cen14, Coh80, Coh97, Coh99, Fea00, Fei03, For81, Gin99d, Gin06a, Hat08, Hos84a, Hoy84, Jas20, Kre86, MLH84, Mes11, Péo08, SCV20, Swe90, Sch14b, TH79, Fle73]. Correspondent [Wil14b]. Corrigenda [Ano83b, Ano95a, Ano13a]. Corrigendum [Ano77a, Ano95b, Ano01b, Ano14a, Ano15a, Ano15b, Ano17a, Ano22a, Wil87c, Wil09]. Corrispondente [Lat12]. Corrispondenza [Gin87c]. Corsican [HAG94a, HAG95a, TBSKO06]. Corsuno [RA18]. Cosmic [Car86, Dic06, Dic13b, How05, Kai07, Kra07c, Mes15, Mit13, Tho87, Fre14, Hos83a, McC13b, Woo88]. Cosmigraphics [Ben14b, Rem15]. Cosmique [Gra07a]. Cosmo [Omo11a]. Cosmogony [Bru78b, Bru78c, Jon71, Llo09, Dra73, Hos72, Llo09]. Cosmographia [dA74]. Cosmographical [Hor82]. Cosmographicum [Ait86, Wil82]. Cosmography [Dun24, Jor94, Mos10, Mos11]. Cosmológica [Hei14a, Gra12a]. Cosmological [Axw17, Gin14, Hor82, Hos06a, Kra20, Rot17, Rus74a, Rus74b, Rya09, Har13, Hos78a]. Cosmológico [Sha99b]. cosmologie [Har85]. Cosmologies [Kra12, CM23]. cosmologiques [Hos78a]. Cosmologists [Liv16]. Cosmology [Apt89, Azh20, Bec10, Ben81a, Ben13, Ber97a, Car08, Cau08, CB05, Don09, Eis96, Eis01, Ell08, Fek13, God06, Gra14a, Had84, Har92, Har84, Har85, Hay81, Hen08, Hol03, Hoo87, Hos70, Hos78a, Hos85c, Hos87, Hos97b, Ken94, Kra95, Kra10, Mar03, McM05, O’B88, Pal84, Rag85, Rya09, Sal94a, Sch78, SIN85, Sha94a, Sha99b, Ste88, SK97, Tri97, WKH18, vdW90, And21, Far05, Jar90, Apt89, Ben13, Bla97, Eis96, Ell08, Gre99, Gre09, Har92, Hay81, Hos95b, Kai07, Kra07b, Mar12a, McC83a, McM05, Smi98]. Cosmonauts [Ano15f, Ger17, Ger14, Neu13]. Cosmos [Bon19b, Dan12b, DeV07, Eas96, Ell08, GG14, Gra00, Har92, Hol03, Lon06, Mal15, McC13c, fH21, KHM^{+}12, Dan22, DeV07, Eas96, Gra97, Gra00, Lip19, Tur90a, Wes93, Rug99, Dan12a, Mit13, Nee71, Rom06, Sch82]. Cosmovisions [Hos06a, McC98b]. Cost [Rot17, Har13]. coucher [BN81]. couchers [Pin81]. coude [Wei82a]. Council [Ped83]. Count [Gen96]. Counter [Mil80, Pal08, TH79]. Counter-Argument [Pal08]. Counter-Revolution [Mil80, TH79]. Countries [Azz06]. Country [Ben84, McK78]. Couple [Gin84d]. Court [Sha94b]. Courtier [Sha94b]. Couvertes [Hos07c, SK08]. cover [Del23]. Covering [Don88b]. Cracking [Mel09, Sch11]. Cracovie [Gin87c]. Cracow [Mel89, Pou85, Dob99b, Hil77, Kre07a, Pou85]. Craft [Mül10]. Craftsmanship [Tur03]. Craftsmen [Don98, Tur71b]. Crater [Hoy84, Mar89]. Craters [Hom72, Hoy81, Lan80, Mar89]. Crätschmair [Wło25]. CrB [Sch23a]. Created [Van95b]. Creating [And00, Ben14a, Lon15]. Creation [Gin00c, Gre09, Hea94, Lie02, Moe75, Tip88b, Bru78a]. Creator [Sha87a, Yeo02]. Credentialling [Gin13c]. Credit [Sha07c]. Cremona [Eva92, Zie23]. Crescent [FSAD99, Gau14, Hog88, Kin88]. Crespo [Cha19]. Crete [PHP92, PHP93]. Crime [Moe80]. Crimean [TPYR21]. Crisis [Azz06, CL22, CJ22]. Cristina [May11]. Cristoforo [Car08]. Criteria [McC03]. Criterion [FSAD99, Gau14]. Critical [Ben91a, Gol11, Gol12, JJ10, Jon71, Jus20, Kel10, Lin07, Not15b, War19, Wes96, vdW17, MHS23]. Criticism [Mer11, Rib24]. Critics [Bar99]. Critique [Blå14, Eva88, Sal94a, Top03, Sal94b]. Croatia [Mul84]. Croatian [Mul82]. Croats [Mul84]. Crocodiles [Loc90]. Cromlechs [TT73a]. Crommelin [JKKC19]. Cross [Bre11, Mul82, Ree11, RS84, Rug04, Wes96, Bre11]. Cross-Cultural [Bre11, Rug04, Bre11]. Cross-Fire [Wes96]. Cross-Staffs [Ree11]. Crossing [Ste05]. Crossover [SP12]. Crossroads [Kre05, McC83a, McC07, Rav73]. Croton [Tau96]. Crystals [Had84]. CS1040 [Bel14a]. cuerpos [Cha19]. Cullen [Qua17]. Culminating [Gau18]. cultivated [Rem15]. Cultivation [Eis85]. Cultural [Ald09, Ano84a, Ave97, Bel09b, Bel19, Bre11, Ger17, GG14, Han09a, Mal15, Mit13, Rug04, Ste05, Str16b, Wal13, HE82, Bre11]. Culture [Ald97, Ano94c, Bee71, DeV02a, Eva99, GHS08, God13, Gre99, Hos85d, Jar98, Mit13, MJT03, Omo17, Rem11, Sal97, Sch01c, Sha94b, Smi94c, GJ22, Rav72, Goe09, Jon06b, RC11, Sul02]. Cultures [Gra01, Hos96, Hos00a, Nor06, Rug04, Wal99, Imh22, Boh19, Bro90, McC07]. Cum [Cha15a, She74]. Cummian [Nor92b]. Cumulative [Ano72b, Ano79c, Ano89b, Ano20a, Hos83b]. Cuneiform [Hun18, Mon22, Jon91a]. Cunningham [Mir17]. Curate [Gin05e]. Curiosities [Ben93a, Rag16a]. Curiosity [Bra90]. Curious [Eis06, Gle07]. Current [Jas96b, Fat00, McC07]. cursuum [Hus13]. Curtis [Gin13a, Hos76c, Ste90b]. curvature [Nal22]. Curved [Ruf71]. Customs [HAG95c]. Cuzco [Fin08]. cxvi [Cen14]. cxxxv [Ano15f]. Cycle [Cha02, Gin92c, Hos95a, McC93, GB95, Gin92c]. Cycles [Sch84a, Sch84b]. Cygni [Ser90]. Cylinder [AS97]. Cyriacus [Sal76].

D [BS92, CWW17, Gin99a, Hos00c, Sha14, Sul14, Ano15f]. Dīn [Sal94b, Ves73]. Dacia [Nor86]. Dag [BG10]. Dahshur [BM15]. d’Ailly [Mal97]. Daily [Bra20, Sch01c]. Da’ire [Kin79]. Da’ire-yi [Kin79]. Daivajña [CMP24]. Dal [Ano06a]. D’Alembert [Wil03, Wil03]. Dalen [Bel14a]. d’Alexandrie [Eva88, Eva01, Pin80]. Dallas [DeV18]. Dall’Olmo [Ano80j, UC17]. Damascus [Sal87]. Dan [Wło15]. Dana [Nau98]. Danica [Has87]. Daniel [Hoy84, Num73, Wil15]. Danish [Wło23, CKE11]. dans [Bir74, Bla01, Kus10, Llo80, Par85]. Daramona [MH88]. Dargestellt [Voe02]. Dario [Oos15]. Dark [Dor20, Sco16, Van14b, Fre14, Mes15]. Darkness [Kra13, Lan82, Tip88a]. Darstellung [Kre82]. Darwinian [Hen08]. Dash [Smi81]. d’Astronomie [Eva01, Rag92, Cha00]. Data [Ext92, FSAD99, Gau14, KY95, Not15a, GH24]. Datchet [Ash03]. Date [BG10, Car17a, Car19, Moe75, Ver12, CR22, Kno97]. datées [HRVM13]. Dates [Fat98, Gin92c, McC98a, Ros82a]. Datierung [Ste12a]. Dating [Bri01a, DE00, Dav19, Dob02, Duk02, SS06, Whi78, Ver12]. d’Auteroche [vH86]. David [Apt14, Doe19, Gin09, Hos95a, Kra22, ML14, Sul14, Vet19, Jon06a, Omo11b, Rei18, SM04]. Dawn [Bat90, Han09a]. Day [Ald11, Bec10, Kra07b, Sha94b, Far05]. Days [McC89]. DC [Doe19, Ger17]. Dead [Ste17]. Dean [Doe01, TH79]. Death [Bla98b, Pau81, HGGVE^{+}23]. Debate [Ber82, Bro83, Cer95, Cor98, Far04, Fei90, Gin88d, Gin11c, Hos76b, Sha99b, Str99]. Debated [Bel11, Hei12]. Debates [Axw17]. Debating [God15]. Debris [Woo88]. Debt [Gin85f, Gin95]. Débuts [Cha89]. Decade [Cha90, Ost05a]. Decans [Gin72]. December [Gin02d, How88, Wis06, Sch23a].

Decepit [Pal08]. Deceptus [Pal08].

Decisions [McM09]. Decisive

[Bro96, Yeo99]. Declaration [She87].

Declinations [CFMR77]. Decline [Sch03]. Decoding [Eva09b, Kel83]. Découverte [CL08]. Decree [Fin19]. Deduced [Dog87]. Dee [Bur80, Bur80]. Deep [Bed10, Gin10b, Hir09, Ste11]. Deep-Sky [Ste11]. Defence [GB86, Lov77, Sha07b, Van97]. Defending [DeV90]. Defense [Fei86]. Deferent [Nor88]. definitive [Mor21]. degli [Gin87c]. Degree [TW25]. Del [Had84, Ave87, Bel06, Cha19, Fin99, Gin82c, Gin02c, Gli88, Gra12b, Hos00a, Omo11a, Pan09b, Rap11c, She86, Swe92b, Tur84, Van96, Swe92b]. Delambre [CT86]. Delft [Mil04]. Delhi [Siv89]. dell [Cen14]. della [Gin87c, Lip15]. dell’Astrofisica [Cen17, Chi16]. dell’Infinito [Pan09b]. Delray [Wil15]. Demarcation [Waf01]. Democritus [Bro83]. Demokratischen [For70a]. Demonstration [Hos00c]. demotic [EP22, Win22]. demotic-hieratic [EP22, Win22]. Dendera [Gin11c]. Dendrochronology [Edd86]. Denmark [Nor86]. Dense [Hen92]. Densmore [Nau98]. Department [Ave95, dP82]. Dependence [Jak70]. Depictions [Rya09]. Depth [Duk03, Gol03b]. Derek [Ano98a, Hos08d]. Derivation [Swe17a]. Derivative [LH05]. Derivatives [Pin88]. Derived [New77]. Derwentwater [Far96]. Descartes [McC91]. Describing [SS06]. Description [RPPA15, Hen92]. Descriptive [Ano76c]. Desert [BS06]. Deserter [Hos04g]. Design [Ost07]. Designs [Ave05]. detail [Kra15]. detailed [Sch25]. Detectors [Lat20]. Determination [All92, FF20, Hos80b, Pet20, SMH18]. Determined [Sti07]. Determining [BBW97, Bru09, GJ22, Moz13a]. Detestable [Sel14]. Detroit [Gin00c, Gin00c]. d’Europe [Hos00b]. deutsche [Gra08a]. Deutschen [For70a, Joh19, Kre07b]. deutscher [Oes15, RR13]. Development [Ano97c, Ben91a, Ber01, Bro91b, Car19, Car08, For70c, Kid87, Mar89, Ram06, SB72a, SB72b, Smi98, Smi08, Smi09a, Sut72, Tat99, Van77, Wil96, Swe86a]. Développement [Bat87]. Device [CTE12, dB95]. Devices [War08, Str23]. DeVorkin [Doe19, Doe01]. Dia [LS24a]. Diagonal [CS14, Loc83, Loc92, Sym02, SC13]. Diagram [LD12, Sym15]. Diagrammatic [Mos11]. Diagramme [Eas10]. Diagrams [Azz11, Eas92, Eas10, Gra05c, Gra12e, Hor82, JJ10, McC91, Pan10, Sha12, Mon22, Sid24b, Sid25]. Dial [AS97, SW18, Tur70b]. Dialogo [Car99, Gin02c, Pal08]. Dialogue [Don82b, Wes96, Don79, Mue00]. dialogues [Chi21]. Diameter [Bro88a, Sha75]. Diameters [DeV75]. Diamond [Sch07a]. Diaries [Eva86, Hub91]. Diario [McM89]. Dibattito [Bel11]. Dibner [Gin87c]. Dicati [LF15]. dichten [Hen92]. Dick [Ano15f]. Dickens [Lep14]. Dickinson [SM04]. Dictatorship [RC16]. Dictionary [Gin85a, Hos76a, Hos82b, Hos04d, Kre09, Low09, Hos76a, Kun90, Wel84, Wil02]. Did [Gap15, Gin85f, Gra06, VvdS19, Cov17]. Diesmal [Hor74]. Dietmar [Ano15f]. Dieu [Gra07a]. Difference [TW25]. Difficulties [Coo75]. Diffusion [GC06, Sha12, Siv05]. Diffusione [Fin99]. Digby [Sel06]. Digging [CWW17]. Dilemma [Plo77]. Dillingen [Gin96c]. Dillon [Hod77]. Dimensions [Pan09a, Tho87, How02]. Din [Ack05, Ken94]. Dinakara [KMD^{+}18]. Dintorni [Pan09b]. Dion [Loc93]. Dionysius [Ter84]. Dionysos [Loc89]. Dioptrique [Lüt00]. Diplomacy [Blu13]. Diptych [Tau94]. Direction [Mel89, Moz13a, Rag01]. Directions [Bro91a, Das15, Gau12]. Director [Hos16a, Ost92b]. Disappointments [Hos14b]. Disciple [CG19]. Discipline [Tay10a]. Disciplines [Ave95, Jar09b]. Discontents [dA17]. Discourse [Olo04]. Discover [Eva09b, VvdS19]. Discovered [Moo94, Rus02, Str16a, Wil07]. Discoverer [Rea92]. Discoverers [Ben76a, Kra11]. Discoveries [Gle04, Pas03a]. Discovering [Rea84, Smi15, Arm73]. Discovers [Pau78]. Discovery [Ash81, Bee06, Bou12, CMO11, Cun16, Dic82, Eva93a, Fab19, For71a, Gin70b, Gol02b, GB09, Het75, Hoc92, HW07, Hol09, Hos75, Hos76c, Hos83a, Hos03a, How05, HS92, Jon04, Lan82, Mea70b, Mil08, Mir17, Nau08, Pal01, Pas03b, Plo77, Pum11, Rea81, Ser90, SL83, Smi83, Smi97, Wes94, Con24, Bra90, Dob79, Hir85, Hos05a, Mil04, Mir17, Wil97]. Discuss [Ped15]. Discussion [Gra11b, Wil95]. Dishes [Kre11]. disks [Jon21]. disparition [Wei82a]. Displays [ECT10]. Disputationes [Sha12, Van23]. Dispute [Ser13]. Dissemination [Rot09]. Dissertatio [She74]. Dissertation [Kre81a]. Dissertations [Ano70g, Ano70h, Ano71k, Ano72h, Ano72i, Ano75a, Ano77k, Ano80a, Ano80e, Ano82l]. Dissolution [Gin85e, Gra00]. Distance [Fea00, Rag01, VB01, Smi79]. Distances [Byr11, KH73, Kre93, Tho87]. Distant [Ome79, Spa04]. Distinguished [War80, Woo87]. Disziplingeschichte [Kre82]. Diversity [Dav09]. Divided [And90]. Dividing [Ben91a]. Divination [Jon06b, Swe13, Eva02]. Divinatory [Hen15a]. Divini [Tur84]. Divinity [Coh95]. Djebel [SK08]. DKK [Wło23]. DM [Wes97]. do [LN09, Rav72]. Dobrzycki [Ano04c, Bar13b, Gin04d]. Dock [Gle04]. Doctor [Ros79]. Doctoral [Ano73c]. Documenta [Swe98]. Documentaires [Hos00b]. Documentary [Dra90]. Documentazione [Jon97]. Documente [Hol77]. Documenti [She86]. Documenting [Bro10, Tim25]. Documents [Bro10, Dra90, Fel02, She86, UC23]. doi [Fal23]. Dokumente [Hor74, Ren89]. Dolan [Gee18]. Dollond [Bak15, GMML14, Kär24]. Dollondiana’ [Kär24]. Dolmenic [Bel97]. Dolmens [BBMG03, BG12, Che99, HAG94a, HAG95a, HZ97, Hos00b, Hos02c, Hos07b, Hos08b, Hos08h, Iun02, Sau00, TBSKO06, Zed05, Hos08b]. Domenico [Van96]. Domini [Gin01c]. Dominion [Jar91b]. Don [Ost02]. Donahue [Too95]. Donald [Ben20, Bog02, Gin02d, Lay02, Lie02, Ost01]. Donati [POH96]. Dondi [Tur75, Gin89a, Tur75]. Dondis [Gin89a]. Dongjuma [Van98b]. Donning [Gin02c]. Doom [Ald11]. d’optique [Lüt00]. Dordrecht [Bec15]. d’Origine [Cha09]. Doris [Ano73e]. Double [Hus12, Pin71b, Sal76, SM05, WS19, Cas21]. Double-Argument [Hus12, Sal76]. Double-Epicycle [WS19]. doubt [HDG^{+}13]. Douglass [Edd86]. Down [Had84, DeV00a, Ell08, Hea97, Mer09, Sta95]. Downshift [Par85]. Doz [vH86]. dracone [Zep19]. Drake [Gin92b, Mac94]. Draper [Plo77]. Drawing [Hos82c, Ree22]. Drawings [Mon94, RP87]. Draws [Gra11a]. Dream [DeV90, Tur07]. Dreams [Lau14, Ost97a]. Dresden [Pin74, Vai12, War08, Kno03, Mar95, Vai12]. drevnei [McC08, Sol09]. Dreyer [Hos11b, Hos11b]. Dreyfus [Déb96]. Drift [Gol73a]. Drop [Sch01b]. drugikh [Sol09]. DSB [Hei99, Hos08a, Kre09, Low09]. Dubitabilium [Ves98]. Dubitationes [Gol73b]. Duchess [May11]. Dudley [Bel07, Eis88b, Bel07, Eis88b]. Due [Ano95b, Gin02c, Pou84, Thu95]. Duhem [O’B88]. Dumée [Aub16]. Duncan [Dob73]. Dunsink [Chi09, Hos82a, Meu10]. durch [Gau17]. Durchmusterung [Her77]. Durées [BN81]. Durham [Hos70]. during [Ber97b, Dun92, Gin75b, Heg82, Hei00, Hol03, Mil04, RC16, Sha87b, UC16, Zru21].

Dutch [Hal03, Ban10, GGCF03, Maa16, Pau81, Van02a, Van98b, Zui19]. Duty [Cha88]. Duzer [Ano15f]. Dwarf [Hol09]. Dwarfs [Nau08]. Dynamic [Kra99, Hen97]. Dynamical [Eva93a, Whi70a]. Dynamics [Eva93a, Nau98, Eva93a]. Dziela [Gin08].

E. [Bra90, Hos85b, Mac10, WDR11]. Ear [Sul77]. Earl [Cha15b, Mol14a]. Earliest [BH12, Gin87a, Kee91, POH96, SV16, SG09, vdW74]. Early [Ait82, Ano06a, Ash81, Bau87, Bec10, Bla91a, Bon16b, Cen17, Cha17, Cro00, Cro23a, CO13, Dew79, Dob99b, Dob01, Don98, Don09, Duk10, Duk12, Eas89, Edd80, Fle92, For76, Gen90a, Ges19, Gin20, Gin81d, Gin82a, Gin88b, Gin90d, Gin90b, Gin96a, Gin96d, Gin98, Gin02a, Gin12c, Gra08a, Har85, Hir85, Hos79b, Hos07b, Hus12, Jar98, JJ10, JR10, Ken87, Kid87, Kin81, Kre81b, Kur19, Lei17, Lév17b, Mac15, Mar03, Mar12a, Mar14, McC04, McC13b, McC08, Mea02, Mor01a, Nee71, Nic90, Nor08b, Not20, Ogj11, Ost86, Pal77, Pau86, Pin73b, Pin76, Pop22, Ram10, Rec21, Ree20, Rem11, Rya09, Sal94a, Sha71b, Sie05, Smi92, Smi18a, Spe17, Sta11, Sta06, Ste19, SB12, SMH18, Ste96, Str81, Sut72, Swe17b, Tho90a, Thu01, Tur84]. Early [Van71a, Van71b, Wal99, Was72, Wes84, Whi73a, Wło11, Yan86, Yeo88, Zso94, CM23, DL15, GE24, Mil81, Not23a, Not24a, Oos15, Smi22, Wes01, WZR23, dH21, Lia19, Omo20, TQ18, Axw17, Edd71a, Gin96a, Gra76, PG79, Was72]. Early-Modern [Har85, Jar98, JJ10, Rem11]. Earth [Car86, Déb05, Duk10, Fei86, Gra14a, Had84, Kwa12, Mor01a, Pau95, RC11, Sch82, Ste77, Ste12c, Swe86a, Waf04, Bru09, Déb91, Gra19a, GB09, Hoo84, Hos03a, Kra09, Kre93, McC83a, McC17, Mol70, MS04, MS05, Nal22, Pal08, PP07, Ste77, SMH18, Sza18, Ver06, Ver09, Voe02, Wil87a]. Earth-Bound [Kwa12]. Earthly [Gin03c]. Earth’s [Moz13b, Moz13c]. Earth’s/Sun [Moz13b, Moz13c]. East [For70a, Haw74, Nor88, Shi11, Ano03a, Bri01b, Gin03a, Moz13b, Moz13c, Rag09, Roc08, Roc14, Sch12a, SG09, UAGGB21, Ano03a, Rag09, Roc08]. East-Asian [Ano03a]. Easter [GG14, EB04, GG14, LL87, Lil89, McC93, McC11, Nor92b, UAGGB21, Ver12]. Eastern [Ack10, Ave09, Bre11, Sid10, Ack10, Bar18, HP98, HP99, Iun02, MSH21, Shi11]. EB.5.12503 [Fal23]. Eberhard [Oes10]. eBook [Bec15]. Eccentric [Rag05, Swe17a, VWH97]. Eccentricities [Car21]. Eccentricity [Moz13a, ZH12, Not24a]. Eccentrics [Mor08]. Eccho [Whi08]. Ecclesiastical [SMH20]. Echoes [Gra21b, Gue12]. Eclipse [Bec00, Bri05, Duk12, Edd71b, FS96, Gau12, Jon12, Kno03, LZ99, Mar95, Mor01a, MSH19b, MSH19a, New77, OP92, Pan94a, Rap11c, SS91a, SM05, Sch05b, Shi00, SSM97, SS97b, Ste97, SS98b, Ste98a, SS91b, SF93, SF95, SF97, SS06, SMH20, Vet19, Wło07, Dob86, KF70, Ma23, MSH22b, MSH22a, MHZS23, MHS23, SS96, SS97a, UC22, UC23, Arm70]. Eclipses [Aab72, Eng02, FA97, GT00, Haw74, Mor01a, MS04, Ols93, Pas02, Por09, Ste08a, Ste10b, SMH18, Thu94, UCI16, MSH21]. Ecliptic [Jon02, Loc99, Nev96]. écrits [Pou96]. Écriture [Gin07b]. ed [Bel11, Hei14b, Wes97]. Eday [TF77]. Eddington [Gin00d, Dav08, Gin00d, Sch07b, Smi22, Woo87, Dav08, Woo87]. edición [Omo13]. Edicions [Cha19, Hei14a]. Edifice [Hos78a]. Edinburgh [ML14, Heg85, Heg85]. Edison [Edd72]. Edited [Ano15f, Ave95, Cha15b, Cro15, Cro23a, Eva17, Ger17, Hei14a, Kle14, Lip15, Mal15, Mul83, Oos15, Rad90, Rob86, Roc17, She86, Swe17c, Bak15, Bre18, Wil15]. Edith [Wen99]. Editing [Don92, JJ10, Swe88]. Edition [Bur15, Eva88, Fal23, Gat20, Gol11, Gol12, Has11, Hei12, Hos85a, Jus20, Ler00, Lin07, Loh13, OT13, Pin80, dSP78, RA18, Ros86, Sha12, Sha17, Str16a, Swe17b, War19, Zep19, vN12, Van23, Hen15b, Wes97]. Editions [SV16, Bou23]. Editor [Big10, Ano07h, Ano24a, Eva15]. Editore [Fin19, Oos15]. Editorial [Ano70a, Ano71a, Ano72a, Ano73d, Ano74c, Ano75b, Ano77b, Ano89c, Ano90a, Ano91a, Ano92a, Ano93a, Ano94b, Ano95c, Ano96b, Ano97a, MJT03]. Editors [Ano09a]. Editrice [Cen17]. Edizioni [Cen14, Lip15]. Edmond [Arm70, Coo84, HD84, Kol92, Arm70, Sch00b]. Edmund [Nau08]. Edmundo [GG14]. Educating [Mar99]. Education [Bok73, Mar14, Pet04, Bok73]. Educator [Gin02d]. Edward [Ave18, Chr98, Dev14b, Coh86b, Kin10, Ost84b, Ost84c, Plo90, Str93, WKH18]. Edwards [GG14]. Edwin [DeV96, Gin92a, Kru17, HB92]. Eeuw [Van02a]. Effect [Bee06, Pal08, Sch01b]. effecte [Hos89b]. Effects [Lov77, Sch86]. Effort [Bla91a]. Efforts [Bec00, Hoy84]. Eggs [Whi74]. Egipto [Hos12a, Bel06]. Egypt [Ack05, BFAH^{+}10, Bel14b, Eva04, EM08, Gin72, Hos12a, KG82, Loc99, SB05, SBF07b, SBF07a, SC13, WZR23, Bel14b]. Egyptian [Gin11c, Loc96, Loc00, Vil90, Bel01b, Bel06, BS06, BSF08, BFAH^{+}10, CS14, Dav85, Gau14, Haa84, Loc83, Loc85, Loc90, Loc92, Loc96, Loc00, Rag16a, Sch00a, SB05, SBF07b, SBF07a, SC13, Sym15, SK16, Wal06, Win25, Gin72]. Egyptians [Das15, LB06]. egzaktnih [Mul84]. Eichstadt [Omo17]. Eiffel [Tur07]. eigenen [Hor74]. Eigenschaften [Bur84]. Eighteenth [WW74, Bro88a, Bro91b, Che72, Far04, Hom72, Rav72, Tur02, Wil93]. Eighteenth-Century [Che72, Hom72, Tur02, Wil93]. Eighth [TQ18, UC16]. Einblattdrucken [Gin76]. Einstein [Ano84b, Ano84a, Gin82a, Gle07, Hen92, HE82, Kra99, Kra07b, Mil80, Sch11, TH79, Wil05a, Ano84b, Ano84a, Cre06, Cro07, Eis06, Far05, GHS08, Gim12, Gin82a, Gin84b, Goe09, GMW80, Hen97, Kra99, Lin04, Mel09, Mil80, Mil81, Nor00a, Ros82b, Sta13, TH79, Woo80, Gin84b, Goe09, Kre08, Lan84, Sta07, Sta13]. Einstein-Turm [Hen92]. Ejnar [Ano83c]. ekoprussia.home.pl [Bar19]. elbow [Wei82a]. Election [Coo84]. Elections [Gol12]. Elector [Oes10]. Electric [Lat20]. Electronic [Hos05d]. Electrons [Ost97a]. Elektrizität [Hen09]. élémentaire [Neu81]. Elements [DK83b, Gin01g, Hos73, Moz13b, Moz13c, Moz19]. Elena [Oes15]. Elert [Ano15f, Sch14b]. Eleventh [Eva89, Pou09, Kin95a]. Eleventh-Century [Eva89, Pou09, Kin95a]. Elf [SO00]. Elias [Wło25]. Elihu [Hoy84]. Eliminate [Gra06]. Eliocentrico [Omo11a]. Elisabeth [Bre18]. Elites [Eis88b]. Elizabethan [Gin02b, Gin02b]. Ellery [Ber74, SO00]. Ellipse [Mil08]. ellipses [Mar23]. Elliptical [Ruf10]. Els [Ano15f, Mal10, Hos08b]. Elucidated [Sha11, Tau13, Tur03]. Elucidating [Tho88]. Elusive [Don73b]. Elveden [DL15]. Emergence [Gin82a, Smi85, Ver12, Mil81]. Emerson [Chr98]. Émilie [Nau16]. emission [RS22]. Empathy [Jar14]. Emperor [Cul93]. Emphasis [Rot80]. Empire [Ave97, Doe19, Bar18, MHS23, Sch20, Ave92, Hos04b, Sch05b, Sch07b]. Empires [Bro90]. Empirical [Swe99a, Swe04a, Van97]. Empiricism [Llo15, De 14]. Employing [Pin71b]. Empress [Car25]. Encounter [Gin90c]. encountered [Cha21]. Encounters [Cas15]. Encyclopaedia [HH07]. Encyclopaedic [DeV20]. encyclopaedism [For16]. Encyclopedia [Dev14a, Hos08a, Nor97, Par90, Eis96, Dev14a, Nor97]. Encyclopedic [Rug07]. End [Ald11, Hal93, Kre89, Sam98, Rav72]. endangered [Wei82a]. Ende [Mos86, Voe02]. Energy [Dor20, Sco16, Wes09a]. Engineers [Ano15f, Ger14]. Engines [Smi97]. England [Hut12, Vic85, Dun92, Tho90a, Wło11, Dun92, Sav12, Tho90a]. English [Eas89, Gan15, Gol11, Gol12, Jus20, Lin07, Mea86, Mol14b, dSP78, War19, Ada21, Ben84, Ben92, Bie21, Cro15, DL15, Dav17, Knu14, Ler25, Mea86, Sei17, Sha70, Waf86b, Wes77b, Gin81a]. Englished [O’B88, Too95]. Engravings [UCI16]. Enigma [Pas03b, Swe92a, Gin00d, Gin02c]. Enlightened [Gin82b]. enlightenment [Her22, Waf04]. Enlumyned [Mea71]. Entangled [Fin19]. Enterprise [Dic09, Rob90]. Enters [Cha09]. Entladungsröhre [Hen09]. entre [Gin07b, Gra12a, Hei14a]. Entropic [Gre09]. Entstehung [Swe99a, Wes97]. Entwicklung [Gre00, Sta99]. Entwicklungsphasen [Ash79b]. Ephemeriden [Mar75, Hub83]. Ephemerides [CG19, Not21, Wil86b, Coh86a, Siv10, DK96]. Ephemeris [Cle71, Gin91b, HS81, Kel83, Moe83, Sho21, CS02a, Kok98, LS05, Ste02a]. Epic [Vet19]. Epicepicycles [Mar23]. Epicureans [Bon07]. Epicycle [Moz14, Pin71b, WS19, vdW74]. Epicycles [Mor08]. Epilogue [BSF08]. Episode [Dob99b]. Episodes [Duk08, Ber89, Gin02a]. épistémologique [Lüt00]. épistémologiques [Hen76]. Epistolary [MJT03]. Epoch [DE00, SC88, Sch02, Sch04a, Sch05a]. Equant [Nor88, RC18, Ber97b, Duk05]. Equation [Rie13, Sta95, Swe00, GC01]. Équatoires [Nor82]. Equatoria [Nor82]. Equatorial [Ben81b, Hoc92, Low88, Tur70b, Wei82a, Wei82a]. Equinoctial [ESBG96, ESBG97]. Equinox [EM08, Rug97, SS95]. Equinoxes [Thu01]. Equivalence [Ito09, Van97, Gin95]. Era [Ger17, Gin17, McC11, Smi07, Zru21, Neu13, Dew87, Gin01c]. Erasmus [OT13]. Eratosthenes [Jon02]. Erde [Voe02].

Erdgestalt [For83]. Erfinder [Nor80]. Erfindung [Ano06a]. Erforschung [Ash79b]. Ergo [Ham06]. Erhard [Kle14]. Eric [Lep14, Ano91b]. Eriugena [Eas01]. Erkenntnis [Ano06a]. Erlangen [Wes97]. Erläuterungen [Loh13, Bur15]. Ermitazha [Rya71]. Ernest [Had84]. Ernst [Wes97, Ano71b, Naz16]. Eros [Atk82]. Err [Ber89, Gap15]. Errata [Ano97c, Ano97b, Ano01c, Ano05a, Ave06, Gin81d]. Erratum [Ano78b, Ano04a, Ano04b, Ano22b, MS05]. Error [Gin03g, McC19, MS04, MS05]. Errors [Bro88a, Gol77, Kri93, She90]. Erschröckliche [Gin76]. erste [Gin80b, Gra08a]. Erudizione [GM03]. eruptions [Sch23a]. Erwin [Hen92, Hen92]. Esfera [LN09]. ESO’s [Smi92]. Esra [Gol11]. Essai [Lüt00]. Essay [Ald09, Ano73e, Ano75c, Ano03a, Ano06a, Arn95, Ash00, Ave92, Ave95, Bel06, Ber74, Ber82, Bla98b, Bol10, Bri98, Bro91a, Bro96, Cha75, Coh95, Cor03, Dal01, DeV98, Den00, Dew79, Din73, Don79, Don92, Ell85, Eva92, Eva93b, Eva01, Fel02, Gin76, Gin89c, Gin05d, Gol73a, Gol76, Gra80, Gra08a, Gra76, GB86, Gra90b, HS81, Har72, Har75, Har78, Hat08, Hay81, Hei12, Hiv81, Hol77, Hos78a, Iwa96, Jar90, Jar91a, Jar09a, Jon06b, Ken73, Ken94, Kin78, Kin81, Kin87, Kno90a, Kre07b, Kre09, Lan89, LG12, Loh75, Lon06, Low09, Mar89, McC04, McC07, McC13a, McM74, McM89, McM05, Mer88, MG86, Mor01a, Mos03, Nau98, Nau01, Nor82, Nor94, PG72, Ped87, Pin03, Pla91, PdV85, Pou75, Pou79, Pou84, Pou85, Pou96, dP82]. Essay [Rag92, Rei92, Ros74a, Rug95, Rug99, Rug04, Sal90, Sal95, Sal06, Sam13, Sch90, SK03, Sha87a, Sha94a, Sha94b, Sha99b, Sha11, Sha17, She74, She79a, Smi96, Spr92, Spr93a, Ste01, Swe76, Swe86a, Swe88, Swe92b, Swe98, Swe99a, Swe01, Swe04b, Swe13, Too77, Too81, Too95, Tur75, Van94b, Van02b, Voe97, Waf01, Was72, Wes94, Wes93, Wes97, Wes09b, Whi76b, Wil75, Wil82, Wil84, Wil88, Wil11, dA74, Ano77a, Ano97b, Cen17, CWW17, Eva17, Ger17, Kru17, Mir17, Qua17, Roc17, Rot17, Swe17c]. Essays [Ben99, Bra90, Coh91, Gen96, Gin88a, Gin90c, Gin96d, GB86, Hos00c, Hos11a, Ken87, McC91, Nor07, Bla91b, Ken85]. Essential [Ree10, Swe17b]. Established [War93]. Establishing [Swe81]. Establishment [DK83a, Her73, Sch81, GJ22]. Estonian [Lep14]. ’Estrange [Gen96]. Estrellas [Hos12a]. Estudios [Nor00b]. Estudis [Ano15f]. Eternity [Loc00]. Etherial [Gol73a]. Ethical [Llo94]. Ethiopic [Nor92a]. Ethnoastronomical [Sne98]. Ethnoastronomy [Ave09, Zei86]. Etnoastronomía [Spr93a]. étoiles [Kun91, Van94b, dV91]. Étude [Cha09, Har85]. Études [Gol11, Gol12, Eva01]. Etymologies [Pin88]. Euch [Kre86]. Euclid [Gin03b, Swe21, TH79, Gin03b]. Eudoxan [Dek08b]. Eudoxus [Wil05a, Kno90b, Lin04, Sch04a, Tho71a]. euerlasting [Hos89b]. Eugen [Hen09]. Euler [Fle73, Wil08a, Wil09, Fle73]. Euler-Mayer [Fle73]. EUR [Ano15f, Boh19, Bur15, Cen14, Cen17, Cha19, Cro15, Eva17, Fin19, Gee18, Hei14a, Hei14b, Joh19, Jus20, Kle14, Lip15, Löh19, Mar14, Mit19, Oes15, Oos15, Roc17, War19, Zep19]. EUR49 [Hal03]. Europa [Hos08b]. Europaea [Gin96c]. Europäische [Her76]. europäischen [Voe02]. Europas [Gin80b]. Europe [Ano06a, Cro23a, Gin90b, Gra05c, Rut13, Tat99, Wal13, Ben92, CK07, CM23, Eas97, Hos08h, Not21, Obr19, Omo20, Tur02, McC19, Ber92, Bur82, Dav06, Dek12, Don98, McC04, Pou09, Wal99, vH87]. European [Gin90d, Kre05, Mol93, dP82, Sam13, SS08, Siv05, Smi92, Ano78b, Ano94c, Bla91a, Cha21, Con70, Dal78, Edm98, Gin90d, Not24a, Por74, Sha87b, Ste10b, Tat99, UC17, UC19]. Europeanization [Bla91a]. Europese [Nor91a]. Eustachio [GB09]. Evaluating [Ros84]. Evaluation [LL87]. Evans [Duk05]. Evanston [Lep14]. Eve [Hoc92, Lüt92, Oos15, TB13]. Event [Bra20]. Events [ASE^{+}13, Sis87]. ever [Del23]. Everybody [CMM^{+}15, Cro23a, Gra21a]. Everyone [Sco16]. Everything [Wil05b]. Evidence [BG10, Far93, Mol03, Not14b, SS98a, CS21, GWZ22, Not21]. Evolution [Dic06, Dic13b, Edm91, Hew74, Huf06b, Mit13, Pan12, Sul99]. Evolutionary [De 06, Sch78]. Exact [All71, Dea94, Gin88a, Gin99b, Hei99, Kun85, Mul84, Sar78, Sha96, Shu82]. Examination [Moz14]. Examples [Bel97]. Excavations [MR93]. Excel [Cul05]. ‘Excellentissimo [Kär24]. excerpts [TH79]. Exchange [Cro23a]. Exchanges [Bre11]. Exercising [Sta06]. Exhibit [Cen17]. Exhibition [Gin10c, Lau07]. Exiguus [Ter84]. Existence [Hol09]. Exotic [Ram10]. Expand [Smi94b]. Expanding [Ber82, Liv13, Nor00a, G^{+}99]. Expansion [Dea94, TW25]. Expansion-Contraction [TW25]. Expected [Waf86b]. Expecting [Car17b]. expedición [Gli88, Ave87]. Expedition [Ave87, Huf91, Ost89, Rap11c, Hos04b]. Expeditions [DOL98, KM91, Sch05b, Sch05b]. Experience [KD98, Llo15, De 14]. Experiment [Gin92b, Gol73a, Nor78]. Experimenting [Kre11]. Experiments [BS06, Gol73a, Ost86, Swe70]. Explained [Gin02a, SD03]. Explaining [Bra20, Eva09a, Hos75]. Explanation [Pap91]. Exploration [Lau14, Doe94]. Explorations [Wil11]. Explored [Kel99, Pin03, Gin80c]. Explorers [Dic95, Hea97]. Exploring [Doe94, GG14, Haz02, Bro10, Rug07, Whi92]. Explosion [CO13]. Exponential [Her73].

Expositio [Not15b]. Exposition [Pin84, Tih17]. Exstaticum [Row07]. Extended [Gin81d]. Extinction [Sch86, Sch13, dJ12]. Extracts [Hos05a]. Extragalactic [Gin92a, SL83, Smi08, Smi09a]. Extraterrestrial [Bro83, Fei90, Str99]. Extratesrrestrial [Cor98]. Extrusion [Pal08]. Eyck [Mon94]. Eye [Lat20, Nor94, Smi89]. Eyepieces [Van77]. Eyes [Edm84, Kes10, RL18]. Ezra [Gol11, Gol12, Jus20, War19, Ave21, Gol12, Jus20, War19].

F

[Ald09, Bro90, Hen92, Moo94, RS22, Rus02]. F. [Duk16, Gef75, Sta06]. fī [Kin87]. Fabbri [Fin19]. Faber [Car86]. Fabio [Lip15]. fabrica [Cha19]. Fabricated [Don88b]. Fabricius [Wło23]. Face [DeV07]. Faced [Sch03]. Faces [WW19]. Facets [Sch00b]. Faciles [Eva88, Eva01, Pin80]. Facing [Hos01b]. Facini [Tur90a]. Facsimile [Gin71, Gin91a, Gin12b, Hos89b, Pou96, Wil15]. Facsimiles [Gin78b, Pou96, Pou96]. Fact [Bro96, Dan12a]. Facts [Gee18, HH13]. Factum [Mil08]. Failed [Ban13]. Failings [Moe80]. Failure [Ost86, SB84]. Faith [How02, Bro96]. Fajada [SS86, Zei85a]. Faksimiledruck [Gin76]. Falk [Wło15]. Fall [Bru78b, Bru78c, Cha02, Fan05, Ost84b, Ost84c]. Falling [Coo75]. Fallstudien [Kre82]. Family [Bro88b, Cro08, Kre86, Smi02, Ben20, Smi02]. Far [Shi11]. Far-Eastern [Shi11]. Farabi [Liz17]. Fareham [Sch15]. Farghani [Sam07]. Farghani’s [Sam07]. Farmer [Van98b]. Farnese [Tur90a, Duk06, Sch05a, Tur90a]. Farnham [Bak15, Bel14a]. Fascicule [Gin88b]. Faster [Smi07]. Fasti [McM01]. Fate [DeV94, Tre04]. Father [Gin05c]. Faujas [Ash03]. Faujas-De-Saint-Fond [Ash03]. Fauth [Sta00a]. Favino [Fin19]. Fe [Nau98]. features [RP87]. February [Swe93]. Federica [Fin19]. Feet [God19]. Feist [Hei14b]. Felix [Ban13, Wes97, Cha19]. Fenomenen [Van02a]. Ferdinand [Gin96c, Nor91a, Has11, Siv05]. Ferguson [Túr90b, Hos84b, Hos84b]. Fernrohre [Dup08]. fest [Ost05b]. Festland [Eva08]. Festschrift [Ano15f, DF14, Mos07, Dun23, Don98, Duk16, Gen96, Gin92b, Hos00c, Mag23, ML07, Mos07, Nor88, Pau82, Rec21]. Fezzan [BEBM02]. fi [Rag85]. Fiction [Dan12a, HH13, Har84, Par92]. Fictional [Dan12b, Ros79]. Fictions [Dan12b]. Field [Das15, Ost94, Tat99]. Fiery [Cam11]. Fifteen [CLL^{+}00]. Fifteenth [Grö83, HS18, Tur72, UC16, vH87, vdW17, Cha98, HH07, Has19, Kok98, Kre11]. Fifteenth- [Grö83, Kre11]. Fifteenth-Century [vH87, HH07, Has19]. Fifth [McC93]. Fifth-Century [McC93]. Fifty [Eis88a, Hir85]. Figure [Gra05a]. Figures [Ave05, Don88b]. Fijo [Gli88]. Film [Bar19, Sul18]. Filosofa [May11]. Filosofia [GM03]. Final [Car17a, Ros19]. Find [Das15, Coo75]. Finding [CG97, Dev14b, For71d, GC19, Neu82, Rag01, SCV20, Wil07, Hut13, PC01, Kra10, Vet12]. Fine [Ben99, Rap11b, Rap11b]. Finé’s [Pan10]. Finest [Rem08]. Finland [Gra12c, Gef75, Gra12c]. Fire [Chr98, San99, Sch78, Wes96, Gin86a, Gin86b, Yeo99]. Fire-clocks [Gin86b]. Fireflies [Str93]. Firenze [GM03]. Firmament [LB06, Rya09, Sha70]. First [Apt99, BD81, BD82, Ben82b, Bla98b, Bro85, Bry72, Cau17, Cha02, Chr98, Coh97, Coh99, Cul02, Dra76, FSAD99, Fei03, For74a, Gin75b, Gin84b, Gra13a, Gre10, Hei12, Hin04, Hos82c, Hos85a, Hos85d, Kra07a, Mir17, Mon94, Pan12, Per00, Rug94, Slu97a, Smi89, Sne98, Tri13, Waf86b, Waf01, Way96, WDR11, Car23, Cun16, Don93, GMW80, Wes97]. Fission [Ost05a]. Fits [Goo09]. Five [Bec99a, CG97, ML14, PRA87]. Fixed [Ber92, Wag88]. Fixing [HS18]. Fizmatlit [Ano15f]. FL [Wil15]. Flagellation [Sha11]. Flagstaff [Liv13]. Flammarion [Cro15, Cro15, She14]. Flamsteed [Coh97, Coh99, Fei03, Gin04c, Per00, Tho84, Apt99, Ash80, Coh79, Coh80, Coh97, Coh99, Dre15, Fei03, For76, KY95, Lat17, Per00, Ste14, Tho84, Wag87, Wil79, Apt99, Coh79]. Flashing [Str93]. Flat [Ben81b]. Flattened [Waf04]. Flavius [GC06]. Flawed [App02]. Fleckenstein [Hos80a]. Florence [Cen14, Fin19, GM03, Oos15, Gin04b]. Florentine [GM03]. Floyd [Pau88]. Fludd [Hal93, Hal93]. fly [Sha07a]. flyga [Sha07a]. Focus [Bla97, Gin94, Lan83]. Foerster [Hen09]. Folk [McC89, Bir74]. Folklore [Ost92a]. Followed [Pin84]. Fond [Ash03]. Fondazione [Cen14, Omo11a]. Fondements [Wil87b]. Fonni [PRA87]. Fontana [Hos95b, Gin03c]. Foot [Ben81b, Kär24]. Foppes [Van98b]. Forbes [Hos85b]. Force [Tat99, Kut23]. Forces [Ano84b]. Ford [Edm98]. Forecasting [Bur20]. Foresight [TT77]. Foresights [Rug83]. Forever [Bia04]. Foreword [Ano70a, Ano71a, Ano72a, Ano73d, Ano74c, Ano75b, Smi18a]. Forged [Sul14, Mun13]. Forgeries [GKS72]. Forgery [Van14a, Ano14a]. Forgotten [AS97, CS02b, Hos04a, KPG03, Mea86, Nau01, Tur84, Tau13]. Form [Gee18, Gin01a, vdW74]. Format [Kre12]. Formation [Gin01g, Sch02]. Formats [Hos05d]. Formeln [RR13, Oes15]. Forms [JR10, Rot09]. Formulae [Gin91b]. Formulation [Jak72b]. Forschungen [For70a, Kre07b]. Fortin [Dre15]. Fortschritte [Sev00]. Fortune [Hos10a]. Fotheringham [MSH20]. Fouad [Jon16, TF16]. Foucault [Déb05, Déb05]. Found [Bec10, Ben93b, Eis06, Gle07]. Foundation [Bar19, For74a, Edm98]. Foundations [Eis96, Hat04, Llo94, Omo11a, Qua17, Swe99a, Swe04a, Cul17, Bel09b].

Founder [Big10]. Founding [And00, Dic91a, Edd90, Hea99, Pan10, Plo93, Yeo88]. Four [Ben81b, Duk08, EP22, Gin11a, Kre20, Moz13a, Tim20, Whi76b, Cha21, Mit19]. Four-Foot [Ben81b]. Four-Point [Moz13a]. Fourteenth [Ack05, CG10, Hos95a, Hus12, JT92, Mol76, Nor83, Not15a, Not15b, Pin98]. Fourteenth-Century [Ack05, Hos95a, Hus12, Mol76, Nor83, Not15a, Pin98]. Fourth [TT77, De 14, MHS23]. Fra [Hol03]. Fracastoro [dB95]. Fragment [CS14, TF16, Yaz14]. Fragments [FJ11, Sym02]. Frame [Pal84]. francaise [Pou96]. France [Bla01, Hos00b, Ler06, Aub16, Che99, Gin88b, Hos00b, Hos02c, Hat83, Rap11b]. Francesa [Ave87]. Francis [Sha94a, GMML14, Bak15, Dew79]. Franciscan [Fal21]. Franciscus [Sha94a]. Franco [RC16]. Francois [Car23, Car24, Swe75]. Frank [Ost07]. Frankfurt [Gin97a]. Frankia [NN24]. Frankish [NN24]. Frans [Kin02a]. Franz [Brü87, Bur15, Mes11]. Frater [Mit19]. Fraud [Van97]. Fred [Lon06, Doe19, Han12]. Free [Ram16]. Freese [Mes15]. Freestanding [Hos04e]. French [Dea94, Tur07, Ace20, Ait86, Bau86, Bir74, Bla01, BH12, BN81, Cha72, CT86, Dea94, Feu75, Ler80, Nau17, Pan10, PP82, Par85, Pou83, PS98, Sha17, Van94b, Wei82a, dV91]. Fresh [Ext92, Gin05d, Wei82b]. Freundlich [Hen92]. Friedhelm [Ano15f]. Friedmann [Smi94b, Smi94b]. Friedrich [Ano15f, Sch14b, Wil86a]. Friend [Gin07a]. Friendliness [Ano13a, CG13]. friendly [Hen92]. Friendship [Sch07b]. Fringe [Sha14]. Frisius [Ver09, Ver09]. Friuli [AR86]. Frombork [Ave95]. Frontier [Gin01e, Par92, Ros19, Tat99]. Frontispiece [JLGL14, LGJLF14]. FRS [Rei18, Túr90b]. frühen [Mal90, Sha99a, Ste96, War13]. Frühneuhochdeutsche [Gau17]. frühneuzeitlichen [Fei13]. frühzeitlichen [Hei14b]. Fruits [Sne98]. Ft [Edd71b, Mau13]. Fulfill [Ano84b]. Full [Coh86a, Da 04]. Function [Mur95, VB21]. Functional [VMV18]. Functions [Hos03b, JR10, McC90]. Fundamental [Swe75]. Fundamentum [Gau17]. Funding [Hos03b]. Funéraires [Hos00b]. Funeral [Cer97]. Funktionen [Ano06a]. Furnace [Gin01g]. Fürst [Ano15f]. Further [CLL^{+}00, HAG95a, Loc83, Loc89, Loc92, Loc95, Nor00a, PC89, TT75]. Future [Ano71c, Ano76b, Gin00a, Jas96b, Smi20, Wea90, Eng02]. Fysica [Van98b].

G [Hos85b, Swe86a, Swe17c]. G.

[BS92, Bru82, Hoy82, Mar23, SIN85, Wil08a, Wil09]. Gabaud [Kra22]. Gaetano [Rap11c]. Gaia [Pet20]. Galactic [Cas15, HB71b, HB71a, TG87]. Galaxie [Hos04f]. Galaxies [Gin90g, Smi02, Van14b, Kra24b, Hos13, Lad10, Pau78]. Galaxy [Ber97a, Smi06, Smi08, Smi09a, Whi00, Gur06, Hos75]. Galilaeana [Oos15]. Galilean [Dra86, Gin83b, Lut91, Sch90]. Galilée [Lüt00]. Galilei [Mos86, Ren89, Van94b, Wil15, EZ16, Dra84a, Lat12, May11, She86]. Galileian [Dra84a, Dra90, McM03, Ree10, Smi93, Van85, Voe94]. Galileiana [Gin02c, McM89]. Galileiani [Hos04c]. Galileo [Ano14a, Bla97, Bro96, Dra90, GM03, Gin05d, How08, Lat12, May11, Ree20, Rem11, She86, Str10, Van94b, Wil15, Ano04a, Ano18a, Bar16, Bla91b, Car99, Coy07, DN98, Don79, Don82b, Dra73, Dra76, Dra79, Dra84b, Dup03, EZ16, Fan05, Fan09, Fel02, Gal83, Gat20, Gin00a, Gin00e, Gin03h, GV03, Gin05d, Gin11a, GvH11, Gin17, Gue12, Hei12, HG04, Hos04c, Hos08g, How08, Jak71, Jak72a, Lat12, Lew07, Mac76, May11, McM74, McM09, Mos86, Mue00, Ome73, Pal01, Pas15, Ped83, Rap13, Rap15a, Ree16, Ree20, Ren89, Ros81, Rot09, Sak80, Sha94b, Sha07c, She74, She86, Sie05, Slu97b, Swe04c, Top03, Van14a, Van15, Wes84, Wes89, Wes96, Whi78, Wil14a, Wil15, Wil95, Bla91b, Bow10, Bro96, Car97, Car99, Coy07, Cro00, Don79, Don82b]. Galileo [Dra84a, Fan05, Fan09, Fin10, Gal83, Gin00a, Gin00e, Gin05d, Gin10c, Gin11a, Hei08, Lut91, Mar97, McM74, McM03, Mos86, Rap13, Rap15a, Ree10, Ren89, Sch90, Sha94b, Sha07c, She87, Ste00a, Van94b, Wes84, Wes96, Whi92, Zuj11, Mac76, Ros81, Smi93]. Galileus [Pal08]. Gallen [Bot76]. Gallic [Gra90a]. Gallo [Jon21]. Galway [Wal13]. Gama [vdW19]. Game [Gin77b]. Games [Cer97]. Ganeśa [CMP24]. gap [Tir22a]. Gaposchkin [DeV85a, DeV03, DeV85a, DeV03]. Garamantes [BEBM02]. Garnier [Tur07]. Gascoigne [Apt14, Apt14, Sel06, Sel12]. gases [Pop22]. Gasperini [Cen14]. Gassendi [Hat08, Hat08]. Gasser [Dan04a]. Gaszowiec [Kre07a]. Gattei [Ree20]. Gauge [Ano84b]. Gauging [Bec22]. Gauloises [Gra90a]. Gauß [Ano15f, Sch14b, For71a, Sch14b]. Gdańsk [Gre16]. Gear [Edm11]. Geared [Bry81]. Gears [Mer77]. Geburtstag [Ano15f, DF14, For81, Mos07]. Gedruckten [Gin99e]. Gee [Bak15]. Gegenwart [Dob99a]. Geistesgeschichte [Har84]. Gelehrter [Oes15, RR13]. Gemini [Bro10]. Geminos [Tih07, Tih07]. Geminus [Bow06]. Gémiste [Cha00]. Gemistus [Mer98a, Tih98]. Gemma [Ver09, Ver09]. Gender [Hen08]. General [Ano89a, Déb95, Dic97b, Dic00c, Dog87, Fat00, For82, Her90, Hos11b, Mer98b, Nor91b, Nor00a, PdV85, Wes90, G^{+}99]. Generated [CS02a]. Generations [Dic06]. Genesis [Hal88, Her90, Sta99, CO13, Sta99]. Genetic [BG12]. Genius [Arm70, Fan05]. Gennady [Ano15f]. Gennem [Hol03]. Genoa [Not23a]. Genre [Jar14, Pan12]. geo [Mar23]. geo-heliocentric [Mar23]. Geocentric [Gra10, Kul75]. Geodäsie [For83]. geodésica [Gli88, Ave87]. Geodesist [Déb13]. Geodesy [For83]. Geograph [Nor80]. Geographical [Pin03]. Géographie [Ler11b]. Geographies [Sul12]. Geography [Ler11b, Sal79, Pin03, Ber14]. Geologist [Bru78b, Bru78c, Doe94]. Geomagnetic [Gub88]. Geomagnetism [Gub88]. Geometrical [Apt89, Dav15, Hos73, Sha12]. Geometrie [Pou84]. Geometrización [Ave87]. Geometry [HH82, HH84, Lüt92, TTF76, War08, Coo84]. Geophysical [Wea90]. Geophysics [Kre82]. Geophysik [Kre82]. geordnete [Spr92]. Georg [Gin03e, Oes15, RR13, Sha99a, Kra07a]. George [Ber74, Sha07b, Bar13a, Cha88, Cou18, Hos08e, Mer98a, SO00, Tih98]. Georges [Cha00]. Georgii [Mal13, MG86]. Geosciences [Sis87, Sis87]. Gerard [Gen96, Gen96, Zie23]. Gerardo [Kru17]. Gerate [Gin90f]. Gerhards [Eva92]. Gerland [Bur15]. Gerlandi [Bur15, Loh13]. Gerling [RS18, SCV20]. German [Dob99a, Eag08a, Eag08b, Hen92, Oes15, Ram16, Ren89, Sch14b, Van90, Wil01]. Germanisches [Gin94]. Germany [For70a, Hut12, For73, Gin11b, Joh19, Kus10, Llo89b, Plo16b, Sta15, Voe03, Gin82b, Gin11b]. Gerovitch [Ano15f]. Gerson [Gol75, Gol77, Gol96a, Nor76]. Gesammelte [Don92, Don00, Gin14, Hol77, Siv77, Tho71c, Voe97, Voe07, Wil86b, Gin93, Kre19]. Gesamtausgabe [Gin05a, Swe88, Swe98, Sha22]. Gesamtkordanz [Gin00b]. geschaffen [Oes10]. Geschichte [Dru84, For70a, For83, Gin81c, Ham09, Her76, Kri94, Kri97, Pet80, Sha85, Ash79a, Dob99a, Dup08, Kin81]. Gesellschaft [Kel78]. Getting [Ost94]. GHA [Gin06d, Wes90]. Ghost [Dic99]. Gianfrancesco [Rap15a, Wil14a]. Giant [De 06, O’D06]. Giants [Fei95, Mil04]. Giganti [HAG93, ZHGM96]. Gilbert [Bru82, Hoy82, Pum11, Bru82]. Gilding [And21]. Gill [Rei18]. Gilliss [SCV20]. Gingerich [Jar90, KE23, Nor94]. Gio [Van96]. Giordano [Blu00, Don71a, Pan09b, Lüt10, Omo19, Tes10, Blu00, Tes10]. Giorgio [Cen14]. Giotto [Lüt92]. Giovan [Ser88, Swe92b, Tuc00]. Giovanni [Tur75, vB10, Bos09, Swe72, Tur75, VG01, Can12]. Girl [Sch18]. Girl-Hours [Sch18]. Giulio [Bel14b]. Giuseppe [Tuc00, CMO11, Ser90]. Giza [NR14]. Glasgow [ML14, BC16, ML14]. Glass [Bel20, Ben82b, Gin91c, Hor82, Low88, Zru21]. Glasses [Gin86b]. Glassworks [Hei08]. Glimpsing [Smi85]. Global [BM15, Goo09, Rap15a, SCV20]. Globe [Ano15f, Can09, Duk06, Van10]. Globe-Trotter [Can09]. Globemaker [Gin13b]. Globen [Gin90f]. Globes [Ack18, Ben18, Bol05, Gin90f, Gin03d, Gin04b, Kun88, Gin03d, Gin04b]. Globular [Smi00]. Globus [Gin94]. Glorious [Eng02]. Glory [Vet19]. Gmunden [Sha89, Sha08]. Gnomonics [Aul16, Sch04b, Tur70a]. Go [Car97]. God [Gin87b, Gra12d, Gra13b, Hos14d, How05, Gin90c, Mar03, Syl07]. Goddu [Mar14]. Gods [Mer92, Sha87c]. Goeppingensis [Gin03f]. Goes [Bow10]. Gold [Ano17c, Eva13]. Golden [Ber97b, Aul16, Bee70, Edd71a, Sil23]. Goldstein [Hen09]. Gómez [Cha19]. Gonville [DL15]. Good [Eva86, Hea99, Rot80, Spa04, War78]. goode [Hos89b]. Goodricke [Hos79a]. Goods [Gin97d]. Goodwin [Bre18]. Gordin [Sha14]. Gorgon [Mor01b]. Gorm [Wło23]. Goss [Mit19]. Gosudarstvennogo [Rya71]. Gottfried [Gin06a, Oes15, RR13, Gaa23, Gin06a]. Göttingen [For74a, Ost97b, Ost98]. Gould [Her71]. Govern [Het94]. Gown [Gin02c]. Graecia [AR00]. Grah [Rod73]. Graha [CMP24]. Graha-laghava [CMP24]. grammōn [LS24a]. Gran [ESBG96, ESBG97]. Granada [Cha19, Hei14a, HAG94b]. Grand [Cul93, Eva88, Eva01, May11]. Grande [CL08, Eva01]. Granduchessa [May11]. Granting [Pan09a]. Graphical [Joh10]. Gratings [Ost86]. grave [God07]. Graves [Bor98, CKE11]. Gravitation [Gui85, Wil87b]. Gravitational [Hol10a, Smi18b]. gravitatione [Wil87b]. Gravity [Eis06, Hos08f]. Graz [Gin84a, LD12, Gin84a]. Grazer [Gra08a]. Great [Ber82, Ber94, Car17b, Hig12b, Hoc92, Hos76b, Kul80, Lan89, Lau07, Low88, San99, Sch80, Smi07, Cro23b, Hig12b, Kul87, Rus02, Sha84, Wal07]. Greater [Duf17]. Greatest [Gle04, Hug10, Ver13]. Grec [Lip15]. Greci [Jon97]. Greco [Lip15, GS13, CLL^{+}00, Eva04, Loc93, Loc95, Ram07]. Greece [Sch04b, Sev00]. Greek [Gra14a, Llo09, Was72, Duk04, Eva99, Eva01, Eva05, Eva08, Gol00, Gra14a, Han21, Lás19, Mas21, MHS17, Neu82, PH94, Pin71b, Pin76, Saw80, Sch02, SL13, Sta72, Thu01, Was72, vdW78, Saw80]. Greeks [Cer99, Gin87b, Oss12, Mer77]. Greeley [DeV90]. Green [Nau98]. Greenwich [Ano90h, Dew79, For70b, For71b, For71c, For74b, Gin03d, Hig17, Ree11, War78, Ano71d, Ano82a, Déb87, Dic82, Gin03d, How86, How94, Ker19, Low88, Mac20, Mul20, Ree11, Smi91, Ste06b, And00, Dic82, Gin03d, Per00, Ree11, Ste06b]. Grégoras [Dob86]. Gregorian [Ric07, Ash86, Swe74a, Swe86c, War13, Ash86]. Gregorianische [War13]. Gregory [Sal90, Gin85g, Sim92]. Grenoble [Ano76d]. Gresham [Coh79, WKH18]. Grid [PC87]. Griechenland [Sev00]. Griechenlands [Eva08]. griechisch [Sev00]. griechischen [vdW78]. Grigor’evich [Gur04]. grobu [God07]. Groot [Llo15]. Große [Row07]. Grosseteste [Cro23b]. Grottos [Hos81]. Group [Loc90, Ano12a]. Groups [Sco90]. Grove [Bru82]. Growing [Reb03]. Growth [Ben92]. Grubb [War98]. Grund [Gau17]. Gúdhártha [Pin84]. Gúdhártha-praká’saka [Pin84]. Guericke [Jas20]. Guest [IK80, SG09]. Guest-Star [IK80]. Guide [Ano22b, App02, Blu21, Dob83a, Dob00, Hos13, Mea86, Mol14b, Pin88, Pin03, WW82, Car99]. Guided [Hos97b]. Guidetti [Lip15]. Gula [Kur19]. Gunther [Gin78c]. Guo [Ste94]. Gustave [Tur07].

H [Bed10, Doe19, For81, Gin02d, Lay02, Mit19, Ost01, Plo93, Too95, Wes97]. habe [Kre86]. habitationibus [Ras13]. Habitations [Ras13]. Habits [Het94]. Hadib [GC06, GC19]. Hague [Déb95]. Hahn [Eva17]. Hale [Hev94, SO00, Ber74]. Half [Bla98b, Way96]. Half-Century [Bla98b, Way96]. Hall [Don98]. Halley [Arm70, Sch00b, Arm70, Ash81, Bro85, Coo84, Fei95, Gin85g, Gin86a, Hos85a, HD84, Kol92, Mar87, OP89, Sch00b, VvdS19, Waf86a, Waf86b, Wil93, Yaz14, Fei95, Mar87, Tho87, Waf86a]. hallucinati [Kre89]. Hamel [Ano15f, DF14]. Hamilton [Hos84a]. Hamlet [PG72]. Hampshire [Sch15]. Han [Cul07]. Hand [Gin82c, Kes15, Nas13]. Handbook [Jon91a, Neu81, CP23, Hen15b]. handbooks [CMP24]. Handbuch [Kre07b]. Hands [Hol10b]. Hands-on [Hol10b]. Handschrift [Gra08a]. Handschriften [Eas10]. Handwerk [dC01]. Handy [Jon17, CT93]. Hangaroa [GG14]. Hanover [Cro08]. Hanson [Pal77]. Happened [Hos76b]. Hardy [Hen08, Hen08]. Harkányi [Pet20]. Harkins [Wes09a]. Harkness [DOL98]. Harlow [Hos16a]. Harmonia [Hos89b, Rem08]. Harmonic [Coh95]. Harmonie [Dru84]. Harmonies [Dru84]. Harmony [Bon16a, Bon19b, Bon19a, Ste98b, Ste98b]. Harriot [Don86, Loh75, Blo78, Don86, Loh75].

Harris [Ste14]. Harrison [Atw79, Ben93b, Hai21, Sch99, Atw79]. Harry [Mit19]. Hartmann [Gin03e]. Hartner [Bee70, Gin85d]. Hartvig [Ano07a]. Harvard [Bel77, Dev14b, Tau94, Bel77, Doe19, Ell90, Gin88d, GDEK24, Kid90, Mac90, Rob90, Rot90, Sch18, Ste90a, Zru21]. Harvard-Smithsonian [Doe19]. Harwit [Rot17]. Hattusha [GB11]. Hauptinstrumentes [Sta99]. Hauptsternwarte [Kri97]. Hauptwerkes [Kre07b]. Haute [CL05]. Hawanat [BBMG03]. Hawking [Bra15, Bra15, Mia12, Bra15]. Hay’a [Rag85, Boh23]. Haynald [TMG02]. Haytham [Gol73b]. Haytham’s [Kin95a]. Hazards [Ash81]. Headbands [Vai12]. Heads [SM05]. Hearnshaw [Hen15b, Kra22]. Heart [Gra22, Kra13]. Heaven [Dan12a, Nor94, Rya09, Siv05, Sol09, Whi71, Sha07a, Gra14a]. Heavenly [Far04, Gin03c, Jon06b, Lan89, Sol09, Wes77b, Gle04]. Heavens [Ben13, Bla01, Car97, Cau08, Coh95, Duk10, Eva06, Eva07, Eva09b, Gin01d, Gra09, Gra00, Kun90, Lad10, Maf05, Mol70, Pau95, Péo08, RC11, Sal06, San99, Sch00b, She87, Ste00a, Tho87, Túr90b, Voe94, Whi00, Ash87, Bia04, Dun13, Eva05, Rem08, Tho84]. Hebenstreit [JLGL14, LGJLF14]. Hebrew [Gol11, Gol12, War19, Ave21, Gol98, Gol01b, Lan98, War15]. Hebrew-English [War19]. Hegel [CH92, Kre81a]. Heidelberg [Cro15, Mir17]. Heidelberger [Swe76]. Heilbron [Gat24a]. Heinrich [For81, Bre13, Gin07a]. Held [Jas96b, Liv13, Sal97, Whi76b, Ave95]. Helen [Ano95d]. Hélene [Bra15]. Heliacal [Kol01b, Kol07, Sch87]. héliaques [Pin81]. Helical [Sch00a]. Heliocentric [Ano22a, Gol02a, Kul75, Swe17a, Gap21, Mar23]. Heliocentricism [Mar14]. Heliocentrism [Dob01, Eas92, Rav72]. heliocentryzm [Rav72]. Helium [Hea15, Nat13]. Helle [Mar75]. hellen [Hub83]. Hellenistic [Gin03b, Tih07, FJ11, Loc89]. Hellman [Ano73e]. Hemelse [Van02a]. Hemisphere [Orc90]. Hemispheres [Dan19b, Dan19b]. Hendrik [Gin99c]. Henri [Dob77]. Henry [Lau14, Wis06, DK83b, KD90, Kin07, Plo77, Wis06, Doe01]. Her [Aub16]. Heraclides [Eas92, Gra14a]. Heralded [Nor08a]. Heranziehung [Kre82]. Herbert [Bar88]. Hercules [Wag92]. Heresy [How05]. Heretic [Tes10]. Heribert [Wes97]. Heritage [And00, Gin11a, Het76a, Hos83a, Lüt92]. Herman [Ost01]. Hermes [She79b]. Hermeticism [She79b]. Hero [Smi94b]. Heroic [Ger17, Neu13, Hoy84]. Herschel [Ash79a, Bec15, Ben13, Cor03, Eva86, Hos04b, Hos08c, Hos11b, Hos13, Mea85, Tri97, Wel94, Ash79a, Ben76b, Ben13, Bol20, Cas18, Cas19, Cor03, Cro08, Eva86, Gin99d, HH13, Hig14, Hos78b, Hos79b, Hos80b, HW81, Hos87, Hos03b, Hos04a, Hos04b, Hos04g, Hos05b, Hos05c, Hos05d, Hos05e, Hos06c, Hos06d, HD06, Hos08c, Hos08e, Hos08g, Hos10a, Hos10b, Hos11b, Hos11c, Hos11d, Hos12b, Hos12c, Hos13, Hos14b, Hos14c, Hos14d, Hos14e, Hos15a, Hos15c, Hos16b, Hys96, Kra11, LH03, Mau13, MBH16, Moo94, Rus02, Sch80, Sch81, Spa03, Spa04, Ste14, Sul18, Tur77, Wel94, Whi08, Win06, Bec19, Gin85b, Gin99d, Kra11]. Herschels [Cro08, Ash03, Bec15, Bec22]. Hertzsprung [Ano83c, Ash79a, Mea85]. Herzog [Gau10]. Hessen [Mos09]. Hesser [Smi18a]. Hessler [Wil15]. Heures [Not15a]. Hevelius [Gin75a, Wil14b, Gie20, Gin75a, Gre16, Mül10, Ruf71, War71, Wil14b, Wło11, Wün99]. Heybech [GC08]. Hi [Bee06]. Hidden [Duk09, Gin03g]. hieratic [EP22, Win22]. Hieroglyphic [ASS13, Pin74]. Hieronymus [Gin77a]. Hierothesion [BG10]. Hiesca [Nor00b]. High [Bog02, Rug82, Rug83, Sha70, Not21, HH84]. Highest [Ost89]. Highland [Ave94]. Highlights [Gin10b]. Hill [CT12, Dic95, RM92, CT12, Wil08a, Wil09]. himlens [Sha07a]. Himmel [Kre86, Ost05b]. Himmels [Jar09a, Rut12]. Himmelsbahnen [Ano15f, DF14]. Himmelsgloben [Eva06]. Hindu [Pin84, Pin84, Mer03]. Hipparchan [DE00]. Hipparchos [Sid24a]. Hipparchus [New77, Ace20, Ano95b, Ano24a, Bou13, GB95, GH24, GWZ22, Jon91b, Jon02, LS24a, MSH19b, Sch05a, Sid04, Thu95]. Hipparcos [Ver13]. Hippocrates [Wil08b]. Hispanic [ESBG96, ESBG97]. Hispánica [Iwa96]. Hispano [Ave87]. hispanofrancesa [Gli88]. Hist [Gol84]. Histoire [Eva88, Hos04f, Lan87, Van92, Cap12, Hat88, Ler06]. Historia [Ruf10, Tho84, Hos06b, Prz79]. Historian [Jar06, Ber94]. Historians [Ano71c]. Historic [Bec13, Ben87a, Hos05a, Rya71, SG16, Tur79a, Van99, Zei86, Ack10]. Historica [Voe03, Ano15f, Kle14]. Historical [Ano73f, Ano84a, Ano03a, Ano15f, Arm74, Ave95, Bel14a, Ber82, BSL86, CM17, Cou18, Eis96, Gen96, Gin87a, Gin90c, Gol94, Hos86b, How02, LS05, Lat20, McC89, Mir17, Mor01a, MS04, PG79, Ree20, Rya71, Sch14a, SS16, Shi11, Sis87, Smi98, Ste04, Ste90b, SG05, Tau94, Tri03, Yeo07, Cun16, Hen92, HE82, MS05, Tri03]. Historicity [BBW97]. Historické [Sch14a]. Historie [Hol03]. Histories [Kre01, Ano01a, Dun24]. Historiography [Spe17]. historischen [Eag08a, Eag08b, Kre82]. History [And90, Ano70b, Ano70g, Ano70h, Ano71d, Ano71k, Ano72h, Ano72i, Ano75a, Ano75c, Ano77k, Ano80a, Ano80e, Ano82l, Ano09a, App02, Arm73, Arm74, Ave18, Bat90, Ben91b, Ben99, BS73, Ber89, Ber97b, Bok73, Bon11, Bro10, Bru09, Bur10, Cer97, Cha85, Cog08, Coo97, DeV00a, DeV05, Dev14a, Dew79, Dic90, Dic95, Dic06, Dob83a, Dob99a, Doe94, Doe00, Doe05, Duk10, Duk12, Dup08, Edd80, Edm84, Eis06, Ell90, Ell08, Eva92, For82, For83, Gin77d, Gin83b, Gin86a, Gin87c, Gin96d, Gin01d, Gin01e, Gin02a, Gin12c, Gle04, Gle07, Gol73a, Gol87a, Gol72, GB86, Gub88, Han09a, Har92, Har72, Har78, Hay81, Hea97, Hei00, Her90, Hew74, Hir92, Hir09, Hol03, Hol10b, Hos76a, HG80, Hos83a, Hos95b, Hos00c, Hos04e, How80, Hug10, Jar14]. History [Kai07, Ken85, KH73, Kul81, Kun88, Kur19, Lan88, Lia19, Maf05, Max12, McC13b, Mea85, Mer09, ML07, Mor01a, Mos03, Mul84, Mul80, Nor88, Nor07, Not20, Orc90, Ost92a, Pas14, Pin90, PdV85, Prz79, Ram10, Rie13, Ron89, Rot80, Rug04, Sam98, Sch82, Sch70, Sis87, Siv89, Smi81, Smi89, Smi92, Smi97, Smi09b, Smi12, Ste98a, SF93, SGDO02, SG09, SMH20, Sul99, Tau09, Ten00, Tur71b, Tur75, Tur79a, Tur84, UL95, Van92, Van98c, Van02b, Ver13, Voe01, Voe03, WW71, Wel84, Wes90, Whi92, Wil97, Wil88, Wil04, Wil05a, Yeo88, vdW17, Kra24a, Kra24b, LS24b, Lin04, Oos15, Sho21, Ste23, Car86, Sil23, Ber94, Coo00, Dra90, For70a, Gin90a, Gin01d, Gra12c, Gub88, Han09b, Kaz73, Kel99, Lov87, Mal97, Mar87]. History [Max12, Nor97, Par90, Ron89, Rot03, Rut03, Smi09b, Sta10, Ste05, Way96, Woo88]. Hitherto [Don73c]. Hittite [GB11]. HKD [Joh19]. Hoary [Hir09]. Hoax [Rus02]. Hochschulschriften [Her76]. Hodierna [Cov15, Jon86, SIN85, Ser88, Ser88]. Hof [Sha99a]. Holbein [Cam02]. Holden [Ost84b, Ost84c]. Hole [Mon82, Sch11, Mel09]. Holes [Sch79, Sch11, Sch07b]. Holland [Hut12, Hal03]. Hollow [Kol92]. Holly [Lau14]. Holograph [Swe74b]. Holy [Mul82, Swe86a, Hoo84, How08]. Homer [Pow88, Gen92a]. hommes [dV91]. Homocentric [Sha98, Swe72, Swe92b, Swe17c]. Hong [Cul02]. Honor [Ald09, Nor88]. Honour [Gin99b, Whi76b]. Honoured [Ald09]. Hooke [Joh10]. Hope [Eva86, Hea99, Rot80, War78]. Hopewellian [HH84]. Hopi [McC77, McC90]. Hopkins [Gan15]. Horace [Van14b]. Horizon [Ext92, Mor97]. Horizons [Hos71b, Ree11]. horloge [Pou83]. Horlogerie [Nor82]. Horns [Bee70]. Horological [Ben20]. Horology [Ben99]. Horoscope [Kol01a, Mes21, WZR23]. Horoscopes [Ano04a, Gin90a, Jon97, Swe04c, EP22, Win22, Gin90a, Nor07]. Horoscopic [Gin90a, Mer09]. Horoscopy [Jon06b]. Horrocks [Gin05c, Wil87c, AH83, Gin05c, Gin05e, Gra13a, Sta11, Wil87d, Gra13a]. Horský [Gin88f]. Horst [Hei14b]. Hoskin [Bec15, Ano17c, EK22]. Hotaling [Ano95a]. Hours [Sch18, Imh22]. House [MH88]. Houzeau [Dew87]. Howard [War98]. Howse [Ano98a]. Hoyle [Han12, Lon06, Han12, Lon06, Ros73]. Hrvata [Mul84]. HST [Lat23]. Hubble [DeV96, Gin92a, Huf06a, McC13c, Tat09, Gin92a, HB71b, HB71a, HB92, Hos97b, Huf06a, McC13c, Tat09, Hos97b, Tri97]. Hudson [Had84]. Huelva [RH00]. Huggins [Aub12, Aub12, Bec00, Bec01, Bec13, Mor80]. Hugo [Nor99]. Human [Ber89, Bro10, McC99]. Humanismus [Sha99a]. Humanist [Gin91d, Sha85]. Humanistische [Sha85]. Humanists [Don98]. Humans [Ger17]. Humboldt [For81, For81]. Hundred [Eis88a, Gin84b, Gin11a, ML14, Whi76b, Dic21, GMW80]. Hunebedden [GGCF03]. Hungary [Bar94, FZ07, VB87]. Hunting [Haz02]. Huo [Cul07]. Hurbanovo [SP22]. Huth [Dev14b, Oes15, RR13]. Hüttermann [Kle14]. Huygens [Ano78d, Ben82a, Hat83, Ben82a, Hat83]. Hven [Wil00]. Hvezdarstvi [Gin83a]. Hypatia [Too96, Too96]. Hypothesen [Whi72]. Hypotheses [Ito09, Mur95, Smi22].

Hypothesis [Bru78a, CO13, Gra12d, GB86, Num72, SC88, Wal07, Whi71, CR22].

I. [Bee06, Smi04, Ano06a]. IAU [Ano95g, Kra22, Way96, Ano71e, Ano89a, Déb95, Dic97b, Dic00c, Gin82d, Kra22, Nor91b, Way96]. Iberia [HGGVE^{+}23, HMA^{+}98]. Iberian [Kre02, BG12, GB10, GC06, HAG94b, HAG95b, HAG95c, HP98, HMA^{+}98, HS99, HVT^{+}01, Hos02d, Mor21, RAGGB18, RH00]. Ibérica [Iwa96]. IBM [Ols93]. IBM-PC [Ols93]. Ibn [GC06, Gol11, Gol12, Jus20, Rag92, Swe17c, War19, Ave21, GC19, Kin88, Moz22, Not24b, Rag16b, SM94, Sam19, War19, Sal87, Gol73b, Kin95a]. Ice [Gin86a]. Ich [Kre86]. Iconoclast [Mos86]. Iconography [Kno03]. Idea [BG12, Edm91, Fei90, Gat24b]. Ideal [Rug95]. Ideas [Hos76a, Omo20, Gin03a, Hos76a]. Idee [Dru84]. Identical [Rod73]. Identification [Loc85, Wag88, PM89]. Identifying [Dav85]. Identity [Gau87]. Idol [Rap15a, Wil14a]. Idyll [JLGL14, LGJLF14]. if [Cov17]. Ihre [Pin92, Ano06a, Gre00, Kin83, Ste12a]. ihren [Gin80b]. ihrer [Ham09]. II [Ack10, Ano06a, Ans05, Coh99, Dog87, Don00, Eva92, Eva01, Gol11, Loc96, Mar12b, Nor86, Pou96, Sal06, Swe88, Tuc00, Voe03, Wes90, Ano76c, AMP96, DK83b, ESBG97, For71b, HRVM13, Hen99b, Rus74b, SS97a, SBF07a, Spr93c, TT76, Zei86, Gin85d, Van95b]. III [Ano06a, Eva01, Fei03, Gin00b, Gin08, Kno90a, Sil23, Oos15, And17, Cha19, For71c, Hos08e, KD90, Gin87c]. ilal [Kin87]. Ilario [Lat12]. Ileana [Cen14, Cen17]. Ilkhanate [Bar18]. Ilkhani [Moz16]. Illuminated [Gra90b]. Illuminating [Gin00d]. Illusion [SD03]. Illustrated [Ack05, Dar91, Eng02, For82, Gee18, Gen87, Her80, Kin78, Mul81, Sil23, Smi09b, Van98c, Arm74]. Illustrating [Gin01d, Lan87]. illustrations [Lip15]. Illustriert [Hor74]. Illyrian [FJ11]. im [Fei13, Har94, Hen09, Joh19, Kle14, Kre99, Mal90, Ost05b, Sha99a, Siv77, Spr92, Ste96, Too77]. Image [JLGL14, LGJLF14, Rap11a]. Imagery [Gen92a, JR10]. Images [Ano06a, Eas00, Ger17, Gin01d, Gin05a, Gin10c, Lew07, McC13c, Mül10, Obr19, Pan12, Sha07c, Neu13]. Imagination [Bon16b, Dan17, Hos00c, Hos04b, Dic00a, Jar09a, Whi02]. Imagined [Liv16]. Imagineering [Ano06a]. Imaging [Huf06a, Lad10, Lip19, Yeo99, GWZ22]. Imagining [Par92]. imago [Hei14b]. Immagini [Ano06a]. Immanuel [De 13, Sha14]. Immateriata [Rab05]. Immobility [Fei86]. Immortal [Chr98]. Impact [Hos10a, Mar89, Har72, Yeo02]. Imperial [Don88a, Lia19, Sch05b]. Impey [Lau14]. Implications [ASS13, DK96, Ext92, Gau14, Gin06b, Kno03, UCI16, VG01, Oss24]. Importance [Van76]. important [Van24]. Imposed [HB92]. Imprimeurs [Gin88b]. Improbable [Gin11c]. Improving [She79a]. Inadvertent [Cle71]. Inaugural [Rug00]. Inca [Ave92, Ave97, Ave92, Fin08]. Incas [Ald97, Ave97]. Inception [Cul93]. inch [Ost07]. including [Gin95]. Incorporated [Bra15, Mia12]. Incorporation [Gra13b]. Increased [Dav12]. Independent [Dew00]. Index [Ano78c, Ano79d, Ano89d, Ano14b, Ano15c, Ano16a, Ano17b, Ano18b, Ano19, Ano20a, Ano20b, Ano21a, Ano22c, Ano23, Ano24b, Fel02, For70b, For71b, For71c]. India [Sch12a, Sha87b, Ano75c, Cha75, Pin76, Plo00, Sha87b, Ano75c, Bla79, Jon10]. Indian [Ram10, Gin05b, Jon10, Kak95, Mer88, Mer03, Moz21, Pin70, Pin71b, Pin72, Pin73b, PM89, Pin96, Ram10, TQ18, vdW80, Kak95, Mer88]. indianischer [Spr92]. Indians [Had84, McC77]. Indicated [CFMR77]. Indicating [Imh22]. Indications [Pow95]. Indicator [EM08]. Indicators [Rug82]. Indices [Ano79c, Ano89b]. Indicted [Eva93b]. Indienne [Cha75]. Indigenous [Hos06a]. Industry [Hat04]. Inedita [Her72]. inédits [Fel02]. Inequality [Swe17a]. Infancy [McC13b]. Inferior [Swe89, Swe17a]. Infinie [Gra07a]. Infinite [Don71a, Gra07a, Ano22b, Blu21]. Infinity [Gra08b, O’B88]. Influence [Azz06, BH73a, BH73b, Ber92, Blå14, SS08, SK03, Zbi97]. Influences [Bur80, Dup14]. influential [Bar22]. Infra [Edd72]. Infra-Red [Edd72]. Infrared [Lon15]. Ingolstadt [Kre99, Mal90, Kre99]. Inhabitants [Mul80]. Initial [Edm11]. Innovation [Rag01, WS19, Mar22]. Innovations [Cul02]. Innovator [dB70, Bro96]. Inquisition [Fan09, Fel02, Wes96]. ins [Gau17]. Inscriptions [AH94, Gra90a, Ler06, Jon21]. Insider [Lau18]. Insights [Ste12b, Tau11]. Insolita [Car25]. Inspiration [Zui19]. Instaurata [vN12]. Institut [Ano15f]. Institute [Ave95, Sho21]. Institution [Ash72, Cha90, Dic06, Doe19, Ger17, Gin87c, Nee91, Pau95, Ash72, Bla98b]. Institutional [Mol93, Hen92, Kra24a]. institutionellen [Hen92]. Instructions [Ano77c, Ano80c, Ano89e, Ano91c, Ano02a]. Instrument [Ben09, Déb96, Gin91c, Gin02b, Gin04a, Kin79, McK78, Rud92, Sch12a, Sta00b, Tur70a, TA02, WW74, WW82, ZM20, Zuj11, dC01, Wei82a, Wło22, GE24, Gin91a]. Instrument-Maker [McK78]. Instrument-Makers [WW74]. Instrumental [Gol77, SB84]. Instrumentation [Ack05, McC99, Sal06, SP22]. Instrumenten [Gin80b]. Instrumentmaker [Van98b]. Instruments [Ack05, Ack10, And90, Ano76c, Ano77j, Ano77l, Ano06a, Bar94, Bed80, BH81, Ben91b, Ben92, Bon76, Cap12, Cha83, Dal01, Dew00, Gen96, Gie20, Gin86b, Gin90b, Gin99a, Gin01d, Gin04a, Gin05b, Gin11a, Gub88, Hil77, How86, How94, Kre11, Lüt00, Min00, Nor82, dP82, Ram10, Rap11b, Rya71, Sal06, Sch14a, Sch20, Sha07c, Smi97, Tau94, Tau20, Tho73, Tur70a, Tur72, Tur79a, Van99, War08, WW74, Wes78, Che72, GE24, Gre23, Bed80, Ben91b, Bon76, Dob83b, Gen87, Gin96b, Llo90, Ram10, WW82]. instrumenty [Rya71, Hil77]. Integration [Bar05]. Intellectual [Had84, McM74, Mer10, Sha99a, Tur71a]. Inter [SGDO02]. Inter-Union [SGDO02]. Interaction [Moz19, Wes94]. Interdisciplinarity [Azh20, Bar16]. Interdisciplinary [Bru96]. Interest [Ano71c, Bau86, BSL86, Gin75b]. Intermediate [Nor97]. Internal [Het74, Het75, Pow88]. International [Ano70b, Ano79b, Ave95, Jas96b, Sul14, Wal13, Way96, Mun13, Ano78d]. Internationale [Ano77i]. Internationales [Wil75]. Internationalism [Sul14]. Internazionale [Rug95]. Interpolation [Hus12, Van94a]. Interpretation [Ano04b, BGP13, CL22, Gin82a, Gin11b, Gra05b, Gra01, Kru97, Mar03, RS84, Sne98, ZB04, Mil81]. Interpretations [CO13, Dic90, Eas92]. Interpreting [Arn95, Plo77, Kel78]. Interrelationship [You86]. Interrogations [Gol12]. Intersection [Sta13, Gim12]. Interstellar [SB72a, SB72b, Bra90]. Intertexture [Kra99, Hen97]. Intervalic [AMP95, AMP96]. Interveniens [Bon09a]. Interwar [Ban10]. Intrigue [Gle04, How05]. Intro. [Bow06]. Introduction [Bie21, FJ12, Hat08, Her80, JR10, Jas96a, KH12, Lin07, Löh19, Ped73, Rag09, Sal95, Tih07, vB17, Sha87b, Wil04]. Introductions [dSP78, War19, Gra80]. inv [TF16]. Inventar [Eag08a, Eag08b]. Invention [Bol10, Bre18, Hei08, Lin79, Syl07, Lin79]. Investigated [Wil14b]. Investigating [CML23]. Investigation [Ste10b, Bel14a]. Investigations [Hos79b]. Invisible [Kra13, Smi85]. Involving [Had84, SG07]. Iohannes [Not23a]. ionized [RS22]. Iovialis [Van90, Pas15, Van90]. Ireland [Cam11, Hos01a, Rud92, Sch91, Wal13, War98, Cam11, Mol14a, Rud92, Rug94, Rug96, Sto94, Cha15b, Ash00]. Irish [Pal12, Bec18, Ogj11, Pal12]. iron [RS22]. Isaac [Coh80, RA18, Sha87a, Coh80, GC06, Mac10, Nau16, Num72, Ruf10, SD82, Wes77a]. Isagogical [Jon99]. ISBN [Ano15f, Apt14, Ave18, Bak15, Bec15, Bec19, Bel14a, Bel14b, Ben20, Ber14, Boh19, Bon19b, Bra15, Bur15, Cen14, Cen17, CWW17, Cha19, Cha15b, Cro15, Cro23a, Dan19b, Dev14b, Doe19, Dun20, Eva17, Fal23, Fin19, Gan15, Gee18, Ger17, GG14, Gre14, Hal03, Hea15, Hei14a, Hei14b, Hen15b, Joh19, Jus20, Kes15, Kle14, Kra22, Kru17, Lau14, LF15, Lep14, Lip15, Llo15, Löh19, Mal15, Mar14, McC19, Mes15, Mir17, Mit19, Mol14b, ML14, Oes15, Oos15, Pas14, Qua17, Rap15a, Ree20, Rem15, Roc17, Rot17, Sch15, Sha14, Sil23, Smi18a, Sul14, Swe17c, Vet19, War19, Wil15, Wil14b, Wło15, Wło23, Zep19]. Ishaq [Not24b, Sam19]. Isidore [Pal12]. Isis [Ano72b, Hos83b, Hos83b]. Islam [Gin98, Mer96, Ber14, Gin99b, Too77, Ber97b, Ken82]. Islamic [Dal01, Gin99b, Gin99a, Kin79, Kun85, Löh19, Mer96, Rag01, Rut13, Sal06, Ack05, Ans05, BEC^{+}99b, Ber97b, EBB^{+}01, For16, Gin99a, Gol72, Hog88, Kin78, Kin81, Kin87, Kin95b, Kun85, Mer96, Moz19, SS95, Sal87, Sal06, Sal09, SS08, SS91b, Bel14a, Ber14, Kin78, Rag85, SS08]. Islamicate [Kun88, Van24, Kun88]. Island [BH73a, BH73b, GG14, Jor94, Wil00, EB04, GG14, LL87, Lil89, TF77]. Islands [CKE11, Le 08, Le 09]. Islay [Tho74]. Isles [Ash00, McC93]. Isotemporal [Kre11]. issledovanija [Ano15f]. Issues [How02, Iwa03]. Istanbul [Ken82]. istoricheskogo [Rya71]. Istoriia [Wer10]. Istoriko [Ano15f]. Istoriko-astronomiceskie [Ano15f]. Istrus [SM05]. Italia [Cen17, Fin99, Chi16]. Italian [Dup14, Gin11a, Tur07, Bou23, Cen17, Gin11a, TG87, Zui19]. Italiana [Cen14, Hos94]. Italy [AR86, Chi16, Dav17, Fin99, Rap18, Cen17, Min00, Rem11]. Iterative [Swe00]. Itinerarium [Row07]. Iudei [Jus20]. IV [Ano06a, Eva01, Gin90f, Hos83b, Kre07b, Mos09, Pou96, Voe03, DeV94]. IVB [Gin06d]. IVe [Ler11b]. Ives [Mil80, TH79]. Ivory [EM08, Llo89a, Tau94]. Ivy [Loc84, Kno91]. Iwaniszewski [Ave95]. IX [Voe97, Löh19, Gin05a].

J

[Ano91b, Ber01, Bre18, Dun20, Ger17, Gol84, Hos00b, Hoy81, Lan80, Nau98, Oos15, Pet04, She99, Siv10, Smi18a, Swe86a, Too84, TH79]. J. [Dew87, Jar99, Moo94, Pau86, Pow88, Rus02, SCV20, TH79, WDR11]. J.-C [Hos00b, Siv10]. J.-C. [Dew87]. Jaan [Lep14]. Jacob [Hos95a, LH03, RA18]. Jacques [Gra21b]. Jagellonica [Gin87c]. Jaghmīnī [Löh19]. Jahre [Gin80b, Gre00, Kre80a, dC01]. Jahres [Ham06]. Jahrhundert [Ano06a, Jor94, Kre99, Mal90]. Jahrhunderts [Gin05a, Kle14, Voe02, Sha85, Sha94a, Sha99a, Whi72]. Jai [Gin96b, Bla79]. Jail [Bow10]. Jaipur [Sha84]. Jaki [Har92]. Jakob [Sha71b, Sha71b]. Jamaica [Plo93]. James [Gin80a, Hos84b, Smi18a, Túr90b, Wil05b, Ano74d, Gin80a, Hos84b, Kid87, Sim92, UAGGB21, Wil05b, Kid87, Pau88, Túr90b]. Jan [Gin99c, KKSS21, Mon94]. Janaab [Cog08]. Jansky [Hir85]. Janssen [Can09, LH05, Can09]. January [Wis06]. Janus [Sch03]. Japan [Ano03a, Shi11, Har72]. Japanese [Har72, IK80]. Jaws [Gin88d]. Jean [Llo15, Nor85, vH86, Bau87, Cha89, Cha89, Wil03]. Jean-Baptiste [vH86]. Jeanne [Aub16]. Jedrzejewicz [KKSS21]. Jefferson [Waf01]. Jena [Joh19]. Jenseitsvorstellungen [Loc99]. Jérôme [LC10, Hos11a]. Jeremiah [Gin05c, Gra13a, TH79, Gin05c, Sta11]. Jerónimo [Rec22]. Jerusalem [Ano84b, HE82]. Jerzy [Ano04c, Gin04d]. Jesse [Ben09, Ben09]. Jesuit [Has11, Maf05, Aub25, CJ20, CJ22, FS96, Gin96c, Har94, Has11, LN09, Maf05, Nor91a, Rap15b, Shi00, Siv05, SS98a, SF95]. Jesuiten [Har94]. Jesuits [Nee74]. Jesus [Ver12]. Jewel [Tay10a]. Jewish [Ric07, Sta13, Gim12, Ped15, Sta13]. Jh [Ste96]. JHA [Ano92h, Ano93i, Ano94h, Ano95i, Ano96e, Ano97i, Ano98h]. Jñãnarãja [Gan15, Knu14]. Joachim [Kra07a, Hos80a]. Joachimi [MG86]. Joannis [Gin78b]. Joaquín [vH86]. Jodrell [Sul02, Sul02, Sha78]. Jofrancus [GD93]. Johann [Ano15f, Bre13, Gin12b, Hor74, Oes15, RR13, Sch14b, dC01, Ano15f, For80b, JLGL14, Lat14, LGJLF14, Sch14b, Tip88b, Van10]. Johannes [Ano15f, Gin99e, Gin04c, Gin13b, Gin15, Gle04, Hor74, Mil04, Sha89, Sha08, Ste98b, Swe76, Too95, Whi76b, Ano12a, DK96, Eas01, Gin14, Gra22, Gra19b, Ito09, Swe86b, CM10, CJ05, Dij09, Don92, Don00, Gin75a, Gin84a, Hol77, Jar00, Tho71c, Voe97, Voe07, Wil14b, Wil86b, Wil95]. Johannis [Gin89a]. John [Atw79, Bec19, Cas18, Coh79, Coh97, Coh99, Dev14b, Don71b, Fal23, Fei03, Gan15, Gin99d, Gin01b, Gin04c, Hen15b, Hos00c, Hos08c, Kra22, Roc17, Tho84, Wil15, Ben93b, Cas19, CG03, CG09, CG12, CG19, Che72, Che73, Coh79, Dic91a, Don71b, Fal23, For76, Gat24a, Gin99d, Gin09, Hai21, Hos87, Hos04b, Hos08c, Hos12b, Hos15b, KPG03, Kun86a, Not14a, OP92, Orc88, PC01, Sch79, Tay10a, Tur71a, Bur80, Hos04b, Per00, Sha08, Tur71a]. Joined [How88, Ost04]. Joint [Ano09a]. Jones [Roc17, Duk05, Duk11]. Jordan [Bel97]. Jordanus [Pou79, Pou79]. Jørgensen [Wło23]. José [Fal23, FF20, Gra24]. Josefina [Boh19]. Joseph [TH79, Wis06, Ano80d, Swe17c, Tay10b, Wis06, Tay10b]. Jost [Gin80b]. Journal [Hut12, Her71, War93]. Journalism [Gin17]. Journals [Ano09a, For73, Wel05]. Journey [Yeo07]. Journeys [Jon15, Roc17]. Jovis [She74]. JPL [Doe08, Doe08]. Jr [Ost97a]. Jr. [WDR11]. Juan [Cha19, Gin99b]. Jubilee [Gin97a]. Judaïsme [Gol11, Gol12]. Judea [Mol03]. Judgement [HS81]. Judgment [Shr00]. Judgments [War19]. Jugendjahre [Eel83]. Juif [Mol76]. Juillet [Dob86]. Jules [Can09, LH05]. Julian [Bir74, Fat98, McC98a]. julien [Bir74]. July [MSH19a, MSH22a, Wal13, UC23]. June [Jon12, Yeo00]. jungen [Her90]. Jupiter [Cro00, Cro00, GvH11, Hoc92, Ste06a, Van90, Wil95]. Jura [PP82]. Jürgen [Ano15f, DF14, Hei14b]. Jury [Sta07, Cre06]. Justification [McC03]. Juszczakiewicz [Bar19]. Juvenilia [Wes84]. Jyesthadeva [Kel10].

Kūhī [VB01]. Kabylian [SK08]. Kairos [Fat00]. Kaiserhof [Rut12]. Kaiserlichen [Oes15, RR13]. k’al [Vai12]. Kalendar [Gin01b]. Kalendarium [Gin01b, Nor83, CG12, DL15]. Kalenderes [Mea71]. Kalendermachers [Gin06a]. Kalenderrechnung [Pet80]. Kalenderreform [War13]. Kalenderschreiber [Joh19]. Kansas [GDEK24]. Kant [De 13, Fei90, Her90, De 13, Hos72, Jon71, Sch78, Bro83, Her90]. kaocheng [WS19]. Kaos [Hol03]. Kapteyn [Ber01, Ber01, Kra15, Pau81, Pau86]. Karin [Oes15]. Karl [Bru82, Kra12]. Kassel [Kre07b, Mos09, Gin80b, Gin90f, Kre07b, Gin80b]. Katalog [Sch14a]. Katherine [Mes15]. Kavaratti [Mer06]. Keeler [Kid87, Kid87]. keeping [Str23]. Keizer [Nor91a]. Kenelm [Sel06]. Kennedy [Kin10, Nor88, Nor88]. Kennet [TT76]. Kenya [KM91]. Kepler [Ano12a, Ano15f, Bia04, Bon19b, CM10, Coh95, Dij09, Don88a, Don92, Don00, Gin84a, Gin03f, Gin14, Gle04, GB86, Hol77, Hor74, Jar00, Nau98, Ros74a, Sta95, Ste90b, Ste02b, Swe76, Tho71c, Too95, Voe97, Voe07, Wes94, Wil75, Wil86b, Wil95, And17, And21, Ano77a, Bia04, Bia15, Bon05, Bon07, Bon09b, Bon16a, Bow83, CR22, CM10, CJ05, Dan19a, Dav09, Dav12, Dav15, Dij09, Don82a, Don88a, Don88b, Don93, Don96, Don00, Don02, Dra73, Dra84b, Gin84a, Gin14, Gle04, Gol01b, Gra05b, Gra08b, Gra09, Gra11b, Gra22, Gra19b, Gra21b, GB86, Gua09, Hol77, Hol95, How05, Ito09, Jar06, Jar09b, JLGL14, Jar14, Kre19, Kut23, LGJLF14, Mar99, Mes20, Mil08, Num73, Rab05, Rot09, Ruf71, Sha02, She74, Spe17, Sta95, Ste98b]. Kepler [Swe76, Swe00, Tho88, Van97, Voe07, Wes94, Wes01, Whi76b, Wil75, Wil82, Wil86b, Wil95, ZH12, Apt89, Bon09a, Bow83, CJ05, Don02, Gin99e, Gin14, How05, Mar99, Nor94, Sha89, Ste98b, Ste12b, Tho88, Wes93, Whi76b, Don82a, Hol95]. Kepler-Symposium [Wil75]. Keplerian [Apt89, Don73c, Don92, Gin99e, Gin13c, Voe97, VG01, Whi74]. Kepleriana [Gin99e, Mos07]. Keplero [Gua09]. Kerala [Kel10]. Kerlescan [FQ77, TT73a]. Kermario [TT74]. Kerry [PF83]. Kessels [Bre13]. Kew [Mac15]. Key [Nau98, SBF07b, SBF07a]. Keys [UAGGB21]. Khaándakhâdyaka [Pin71a]. Khadim [BSF08]. Khafrī [Sal94b]. Khalīlī [Kin73]. Khalili [Gin99a]. Kiddery [Llo18]. Kiel [Ham09]. Kieler [Ham09]. Kindi [Yaz14]. kinetic [Pop22]. King [Ano17c, Dek12, Kin07, Ste94]. Kingdom [Bel01b, Loc83, Loc92]. Kings [McC98b, PP07, Cau25]. Kirch [Gaa23, Gin06a, Gin06a]. Kircher [Row07, Row07]. Kirkwood [Num73]. Kiskartal [ZL99]. Kit [Gib78]. Kitāb [Kin87]. Kitab [Sal79]. Kitt [Edm84, Edm84]. Klingenstierna [Bre18, Bre18]. Knighthoods [HH13]. Knowing [Bra15, Sul12, Mia12]. Knowledge [Ano06a, Gee18, Gra01, Hei14b, Lip00, McC03, Mil04, NN24, Roc19, Wil14a, Rap15a]. Known [Slu97a, Not24b]. Knowth [Sto94]. Knudsen [Gan15]. Koestler [Dra80, Jar14]. Koestler-Type [Dra80]. Komet [Hen09]. Kometen [Gre00, Mar75]. Kometenforschung [Gre00]. Komposition [Pin92]. końca [Rav72]. Konfessionsbildung [Joh19]. Koninklijke [Hal03]. Konstgenoten [Van02a]. Konstruktion [Ano06a]. Kontroversie [Row07]. Konzepte [Loc99]. Koordinateniiberlieferung [Too77]. Kopal [Phi91]. Kopernik [Gin08, Rav72]. Kopernika [God07]. Kopernikanischen [Har84, Sev00]. Korea [JKKC19, Shi09, Ano03a, Shi11]. Korean [Shi09]. Korrektur [War13]. Korrespondenz [Gin06a, Oes15, RR13]. Kosmographie [Jor94]. Kosmologie [For83, Sha94a, Ham06]. Kosmologische [Eas10, Row07]. Kosmos [Hol03]. Koyré [Ros74a]. KPNO [Edm91]. Krakowskim [Hil77]. Kreamer [Mal15]. Krist’an [Nor02]. Kritik [Sev00, Too77]. Kritischer [Swe88]. Kross [Lep14]. Krupp [Rob86]. Ksiezyc [God13]. Kugelgestalt [Voe02]. K’uk’ulkan [Ald03]. Kulikovsky [Gur04]. Kulturen [Gol95]. Kulturgeschichte [For83, For82]. kul’tury [Rya71]. kulturze [God13, Rav72]. Kunhikunhi [Mer06]. Kunitzsch [Ano21b]. Kunstkammer [War08]. kunstlich [GE24]. Kurtik [Ano15f]. kurz [Swe90]. Kurzbeschreibungen [Mar75]. Kylex [Fal21]. Kyoto [Dic97b].

L [Gat24a, Gen96, Kru17, Rad90, SCV20, Ano15f]. låta [Sha07a]. Laboratories [Sch80]. Laboratory [Yeo88]. Lacaille [Gin04c, Déb13, War93, War93]. Láctea [Omo13]. Lactée [Hos04f]. ladies [Chi21]. Lady [Wel94, Wel94]. laghava [CMP24]. Laid [Whi74]. Lakshadweep [Mer06]. Lalande [Hos11a, Hos11a, LC10, LC10]. Lambert [Hos78a, Hos78b]. Landgraf [Gin90f, Mos09]. Landgrave [Mos09]. Landmarks [Gin01d, Kak95]. Lands [GB11, Ome79]. Landscape [Bel14b, GB11, HGGVE^{+}23]. Landscapes [Bel16a, Bel19, Hos08c]. Lane [Pow88]. Lang [CWW17]. Langley [Edd90]. Language [Wło23]. Languedoc [Fai94]. Lansbergen [Gra13b]. l’Antiquité [Llo80]. Laplace [Bru78a, Eva17, Hei99, Sti07, Eva17, HRVM13, Hei99, Sti07]. Large [Jar99, Ste08a, SMH18, TT71]. laser [TPYR21]. Lassell [Ben81b, Smi83, SB84]. Last [Orc84, Par92, Shr00, Smi07, Lev25]. L’Astrolabe [Ano77i]. l’Astronomia [Ste12b, Hos95a]. L’Astronomie [Llo80, Van92]. Late [Bec15, Bor98, CK07, Hos14c, Hub91, Jar90, Jon99, Mon21, PH96, Péo08, Rem11, Sam13, Ald02, GC06, Nal16, PHP92, Tih21]. Late-Minoan [PH96, PHP92]. lateinische [Eva92]. Lateinischen [Gau17]. Later [Gle07, dP82, Swe04b]. Latin [Ave95, Ber12, Eas97, Jus20, Mal90, Ras13, Rut13, Sal95, Wal13, Ave21, CP23, Dal80, Jus20, Kee91, Lás19, Not24a, Ped15, Pou09, Ves98, Zie23]. Latin-English [Jus20]. Latines [Hat08, Hat08]. Latinorum [Gin87c]. Latins [Cha09]. Latinus [Jus20]. Latitude [Ano01b, Sch01a, Sch02, Sch04a, Vil88]. Latitudes [Rec20, Ste03]. Lauding [Sul99]. Laurentius [Omo17]. Law [Bru78a, Her73, KH73, Tat99, Jak72b]. Laws [Wes94]. Lea [Ost84a]. Leading [Ben09]. Leaf [Loc84, Kno91]. learned [Chi21]. Leavitt [Wil07]. Leben [Dob96, Hol77, Hor74, Sta00a]. Lebenserinnerungen [Hen13]. Lebensläufe [Ano15f, DF14]. Lebeuf [Ave95]. L’Éclisse [Rap11c]. Lectiones [She87]. Lecture [Rug00]. Lectures [Coh79, Eva05, Har94, Hos73, She87, Tur71b, Her72]. Led [App05]. Leeuwarden [Van98b]. Legacy [Bar13b, Ber74, Ber01, Ber22, GHS08, Gin85d, GHHB02, Goe09, Gol87b, Swe13]. Legende [Gin81c]. Legends [Mor97]. Lehre [Wes97]. Leibniz [Gin95]. Leiden [Boh19, Jus20, Mar14, War19, Ban10, Dek08a, KM91]. Leipzig [Ano15f, Kle14, Gin71, RR13, Oes15]. Lemaître [Kra07b, Far05]. Lemaître’s [Bro15]. Length [Bow06, Duk09, Sch92, Swe86c]. Lenina [Rya71]. Lens [Bre18, Meu10, Ost07]. Leo [Fin19, Oos15]. Léon [Déb05, vH86]. Leonardo [RP87]. Leonid [San99]. Leonids [San99]. Leoninus [Ver06]. let [Sha07a, Rya09]. l’Ethiopie [Hos00b]. Letter [Dra86, Gin85g, May11, Moo94, Nor92b, Rus02, Sel06]. Lettera [May11]. Lettere [Rap11c, Tuc00]. Letters [AH83, Brü87, Eva86, Eva17, Gin99c, Hos78a, Hos15b, Mes20, MJT03, Tuc00, War98, Wel94, Mes21, Gre16, Siv05, Wel05]. Lettres [Hat08, Hos78a, Ler06, HRVM13, Hat08]. Leuchtenden [Ash79b]. Léunivers [Cap12]. Level [CR11]. Levenger [Wil15]. lever [BN81]. Levers [Pin81]. Levi [Nor76, Gol75, Gol77, Gol96a, Nor76]. Leviathan [Hos02b]. Levy [Yeo02]. Leyes [Hos00a]. L’horloge [Ano83b]. Li [Cul07]. Libelli [Azz11]. Liber [Hus13, Jus20, Nor99, HH07, Not20, Rut12]. Liberales [Ste96]. Libertas [May11]. Librairies [Gin88b]. Libraries [Con70, Gin87c, Gin77d]. Library [Ash72, Cro15, Gau10, Gin77c, Gin80a, Gin87c, Gin89b, Grö83, Mea81, Spe80, Tau20, Win06, Feu75, LD12, Gin05b, Vil90]. Libration [Pum11, Wło11, Wło25]. libri [Eva00]. Libro [Omo11a, dA74, Gin82c]. Libya [BEBM02]. Lichte [Kle14]. Lick [Pau88, Smi89, Edd71b, Ost02, Ost07, Pau88, Sha71a]. Lid [Loc83]. Liddel [Dob73]. Lids [Loc92]. lieb [Kre86]. Liebniz [Gin95]. Life [Apt99, Bro83, Bro96, Chr98, Cor98, Cro12, Déb05, Dob92, Fei90, Gin82b, Gin05c, Han12, Hei99, Hos88, Hos14b, Hos15b, Kel74, Lon06, Omo20, Ree20, Ros74b, Sch80, Sch01c, Sch12b, SK03, Str99, Túr90b, Ver12, Wel05, Wil00, Wil05b, Str99]. Light [Bat90, BG10, Ber01, Bri01b, Dup03, Eva02, Fab19, For80b, Gin00e, GB09, Hos08f, Hos12a, Jak70, Lat17, Mul82, Naò16, Oss12, Sis85, Swe17c, Tat99, Tho73, ZM20, vH83]. Lights [Far96, Rya09]. Lignères [Fal23, CG19, Fal23]. LIGO [DeV18]. lii [Ano15f]. liii [Ano15f]. Limit [TW25]. Limitations [Moz13b, Moz13c]. Limited [DeV96]. Limits [Azz06, Cor98]. Lindegaard [Gan15]. Lindsay [Bec13]. Linear [Rug85]. Lines [Bre00, Kre11, Moz13a, Rei92, Sha87c, RS22]. l’Infinité [Gra07a]. Linking [PP07]. Linnell [OP92]. l’Inquisition [Fel02]. Lintels [Kel78]. lió [Ano15f, Mal10]. Lion [Loc90, Nau98]. List [Ano77l, Cha15a, Gin81c, Gin81d, How86, How94]. Listed [Gin99d]. Lists [Zso20]. Literary [Eis96]. Literatur [Gau10]. Literature [Dan12b, Har84, Hei00, Her77, Mea86, Sha70, Wło23, Arn95, Mea86]. Literatürü [Hei00]. Literaturund [Har84]. Littéraire [Ler06]. Little [Bar13a, Pin81, SO00]. Liturgical [Kre12]. Liu [Cul02]. Lived [CT12]. Liveright [Vet19]. Lives [DeV89, Dun20, Sch12b, War98]. Living [Ram06, McC13a]. Livre [Eva88, Eva01, JT92, Mol70]. Livres [Eva01, Fel02]. Lixiang [WS19]. Ljubljana [Juz06]. L’Observatoire [CL05, Lan87]. Lockyer [Big10, Big10, Din73, Din73].

L’Oeuvre [Mol76]. Logan [Gin80a, Gin80a]. Logic [Hei16, Mar14, Van89, Sid25]. Lohr [Bur15]. Lomonosov [Kri88, Kri88]. London [Ben09, Ben20, Car86, Gin02b, Ano70h, Cha17, Gin02b, Kun86a, Sim09]. Lonely [Gin77b]. Long [Ber22, Bol10, Edm84, Eva09b, Lan83, LG12, McC83b, Smi81]. Long-Focus [Lan83]. Longer [Fei95]. Longitude [Atw79, Ben93b, Déb87, For71b, For71d, Gin16b, Ker19, Nev96, Por09, Vet12, CT86, For70b, For71c, CS02b, Dic82, Nau01, Sch99]. Longitudes [Ano95b, Car20, Jon91b, Thu95, CT86, Gre23]. Longomontanus [Car20, Has87, Moe75, Rec20]. Look [Fal14, Kno91, Wło15]. Looking [San99, UL95]. Looks [Sch82]. Loomis [SM04]. Loosened [Gen90a]. Lord [Lat93, Far96, Arm72]. Lore [Sch04a, Was77]. Lorena [May11]. Loss [CMO11]. Lost [Atw79, Aub16, Bel01a, Ben93b, Dev14b, Duk08, Eis06, Gin04c, Gle07, Hoo84, Sch05a, Yaz14, Car25, Hut13, Gin04c]. Louis [Ano01j, Gin04c, Mer01, Déb13]. Louvain [Azz06, Dob77, Gin04a, She87, CG18, Eva01]. Louvain-La-Neuve [Eva01]. Louwman [SG16]. Lovanienses [She87]. Lovely [Hos02a]. Low [Azz06, Bak15]. Lowell [DeV02a, Dic95, DeV02a, Dic95, Het78, Str93, Fei90, Het78]. Löwen [Wło25]. Lower [SB05, SBF07b, SBF07a]. LPL [Yeo88]. Ltd [Bel14a]. Lubieniecki [Jas20]. Lucidator [Ves98]. Ludwig [RS18]. Ludwigovich [KI10]. Lügen [Gau10]. Luisa [Ano82d]. lumière [Fel02, Cha85]. Lumières [LC10]. Luminaries [Gol12]. Luminosa} [Cov15]. Luminosity [Wil07]. Luna [Not23a]. Lunar [Car19, CS02a, Coh86a, CS02b, Cow71, Doe94, Dog87, FS96, FSAD99, For70c, Gau14, Gol82, Gol02b, Haw74, Hog88, Hom72, KMD^{+}18, KY95, Llo89b, MEA91, McC93, McC13a, Moz14, Nau01, New77, Ols93, Por09, Pum11, Rug82, Rug83, Sal76, Sal89, Sch92, SS97b, SF93, Ste00b, SMH20, Sto94, Swe99a, TT73b, Tho74, TT75, TTT75, TTG76, TT77, TT78, Tho71a, Van94a, Van71a, Van71b, Van02b, Wes77a, Whi73a, Whi76a, Whi78, Wil87d, Wil03, Wil08a, Wil09, Wil11, Car23, Car24, Gol84, Kin88, SS96, SS97a, TPYR21, Wil87c, Whi73a, Yeo88]. lunationibus [Zep19]. Lundmark [HB71a, Het76b, Tee89]. Lundy [Far93]. Lune [CG12]. Lunette [CL08]. l’Univers [Hos78a]. Lutf [NRKN16]. Luther [Kre89]. lxi [Bur15]. lxviii [Ree20]. lxxiv [Wes97]. lxxviii [Jar90]. Lyman [Ost97a]. Lynn [Nor83]. Lyric [Duf17].

M [Gin03f, KE22, Mer06, Roc17, SCV20, Wes97]. M. [Ano81a, Kri88]. M.ALFA-EB.5.12503 [Fal23]. Mînâ [Loc95]. MA [Dan19b, Dev14b]. maître [Mol76]. Maanen [Het76b, BH73a, BH73b, Gin92a, Het74, HB92]. Mabel [SM04]. Machian [Eva93a]. Machine [Yav99]. Machines [Cap12, fH21]. Machu [DW83]. Mackenzie [Ano74d]. Maclear [Eva86, Eva86]. Macmillan [Ano15f, Kra95]. Macro [Déb96]. Macrocosmica [Rem08, Hos89b]. Macshar [NKM^{+}16]. Madau [PRA87]. Made [Dav17, Dra80, FS96, Shi00, Smi94b, SS98b, SF95, Yeo02]. Madeira [Aug24]. Madeleine [Fal23]. Mädler [Tip88b]. Madrid [BBW97]. Maestlin [Gin03f, Gra07b, Gra14b, Gra12e, Tre04]. Maestlinus [Gin03f]. Maghrib [Sal09]. Maghribī [Sam98]. Maghribi [Mor21]. Magi [Nor08a, GHHB02]. Magic [Gin99a, Gin01g, Par86, Rut13]. Magini [VG01]. Magli [Bel14b]. Magna [AR00]. Magnetic [MG83]. Magnetical [Ben81a]. Magnetism [Bal85]. Magnificent [Sel14]. Magnis [Cha15a]. magnitude [MD21]. Magnitudes [Lil99, Sch13]. Magno [Ped79]. Magnus [Oes15, RR13, Pri91]. Magrib [Hos01b]. Main [She15, Tat99, ZM20]. Maize [Spr93b, Spr93c]. Majanos [BAE93]. Major [Mül91, TT84, Cha21]. Make [Gin03e, Gra19a, Gib82]. Make-it-Yourself [Gib82]. Maker [McK78, Omo17, Spa03, Ben09]. Makers [Dun92, Gin02b, Gin04a, Sch12a, War98, WW74, dC01, Dic21, Tur23, Nor04, WW74]. Makes [Hei12]. Making [Arn11, Car08, Cas18, Eis01, Gin90b, Gin91c, Gin03e, GV03, Hos16a, Joh19, Obr19, Rei18, Rud92, RC16, Rüp15, SS08, Ver13, Woo88, Mar22, Bec19, Gen96, Gin02b, Lip00]. Malaga [HAG94b]. Malayalam [Kel10]. Male [Mil08]. Malet [Ano15f]. Malevolence [Gen92b]. Malinalco [Tre90]. Mallawi [SC13]. Mallorca [AZB02, RBH00]. Malmi [Mer06]. Malta [Hos04e, SHV92, VSH93, Hos04e]. Malvern [Zep19]. Man [Bon09a, Déb05, Gin97c, Hos71b, Rug99, Sch82, Smi94a, Smi94b, Waf04, Wil05b, Yeo02, Bra15, Ber81, Gin00a, Her80, Pau78]. Managing [Mar22]. Manchester [Cha15b, Dic00c]. Manfredi [GB09]. Mania [Mac15]. Maniliu [Tau13]. Manilius [Mer10, Tau13, Mer10]. Manned [Van95a]. Manner [Swe21]. Mansūr [KF70]. Mantua [CG18]. Manual [Cha00, Eva89, Gin03e, Mer06]. Manuel [Cha00]. Manufacture [Sim09, Spa04]. Manuscript [Gin75a, Gin87c, Has11, Kre02, Pou85, Spe80, Sta11, WW71, Sid24b]. Manuscripta [Don92, Don00]. Manuscriptorum [Gin87c]. Manuscripts [Dek20, Don92, For74b, Gin92a, Gin95, Grö83, Har92, Pou96, Sav12, Vil90, CT86, Feu75, Gee18, Gin87c]. Manuscritos [LN09]. manuscrits [CT86, Feu75, Pou96]. Many [Pan09a, Sch00b, Sch12b, WW19]. Map [Pum11, Ver13]. Mapping [Gin80c, Gin01d, Hen99a, Hen99b, Kel99, Whi76a, de 22, Mea02, Van02b, Whi76a]. Maps [Ack18, Blo78, Ste14, Sto94, Van71a, Van71b, Van02b, Whi73a, Rag01, Ash87, Haw74]. Marāgha [DK96]. Maragha [Blå14]. Maraghah [Sal85]. March [Rot01]. Marchia [Sha94a]. Marcus [Kre12]. Mardi [Gin77b]. Margaret [Aub12, Aub12, Bec13]. Maria [Mos11, Ano82d, Wel05, Wel05]. Marie [Fal23]. Marie-Madeleine [Fal23]. Mariner [DeV96, Nor90, Ree11, Nor90]. Marion [Gee18]. Maritime [And00, Gin03d, Ree11]. Marius [Pas15]. Mariusz [Ave95]. Marker [BAE93, EBA94, Hos97c, SS86, Zei85a]. Markers [ESBG96, ESBG97]. Marking [Sho21, SS90]. Markree [Hos82a, MLH84]. Mars [Dic95, Ost89, Sul12, Wil97, Car20, Car21, CR22, Cro15, Don96, Don00, Gin70b, Gol96a, Pet84, Plo93, Rec20, Ste00b, Sul12, TQ17, Wil97, Cro15, Het78, She14]. Marseille [Hus13]. Marseilles [Hus13]. Martianus [Eas00]. Martin [McK78, McK87, Oes15, RR13, Rot17, Wło15, McK78, McK87]. Martis [Don00]. Marum [Ano76c]. Marvellous [Rec20]. Mary [Eis85, Eis03, Hos10a, Ost84a, Nee08]. Maryland [Had84, Bro88b]. Ma’sar [Sal95, Ada21]. Masha’allah [Gol72]. Mashups [Gos16]. Maskelyne [Gen90b, For74b, Gen90b]. Masks [Lad10]. Mass [Nau08, Sal06]. Massimi [Gin02c]. Master [CG19, GD93]. Material [Ano71d, Coo99, Cro23a, Eva99, Het76b, Sch01c]. Materials [Dob83b, Gin06d, Sha96, UL95, WW71, Win06]. Mathematica [Jon99, Too77, Whi70b]. Mathematical [Ack05, Bel14a, Ben75, Dal01, Don71b, Fei84, Gin03a, Grö83, Har78, Kel10, Nau98, Nee71, Nor06, Pin73b, Rem11, Sal79, Sta72, Ste03, Ste13, Swe01, Tay10a, War08, Whi70b, Wil05a, Lin04, Gin91b, Tay10a, Ack05, Gin89c, McC82]. Mathematician [Jon15, Omo17, Rem11, Roc17, Vic85]. Mathematicians [Don88a]. Mathematics [Bur82, Duk10, Duk12, Gin06b, Har94, Jon10, Kak95, Kin78, Kre99, Nor96, Plo16a, Pou79, Rap11b, Rap15b, Ric07, Siv98, Jon10, Mar99]. Mathematicus [Gin03f]. Mathematik [Har94, Kre99]. Mathematiker [Kin83, Sha08]. mathématiques [Wil03]. mathematische [vG05]. Mathematischen [Wil01, Sha85]. Matilda [Car25]. Matteo [Cro23a]. Matter [Fin19, Sch80, Van14b, Fre14, Mes15, McM05]. Matters [Bra90, Bro96, Nor92a, Mes15]. Maturing [Whi70a]. Maupertuis [Waf04]. Mauri [Don79]. Maurolico [VB21]. Maverick [Pet04]. Max [Naz16]. Maximilian [Rut16]. Maximum [Nau08, GC01]. Maxwell [Wil05b, Wil05b]. May [Ave95, Das15, Ma23, Mel89, Swe93]. Maya [Ald11, AB14, Cog08, Kel78, Pin74, Ald03, Ave94, AH94, AMP95, AMP96, ABB03, ASS13, Kno03, McC83b, MS85, Sch84a, Sch84b, Kru17, Har73, vG05]. Mayan [Arn95, Ald11, Bel16b, Har73, Thu94, vG05, Pin74]. Mayans [Ave94, Hos94]. Mayapán [Ald03]. Mayer [Fle73, Gin82b, Her72, Wil11, For70c, For80a, Gin82b, Kle14, Wil11]. Mayers [Kle14]. Mayor [Spr00]. Mayr [Van90]. Mazela [SK08]. McConnell [Bak15]. McDonald [Lan88, Lan88]. MD [Gan15]. Meadows [Big10]. Mean [Byr11, Duk09, Gol03a, CMP24]. Meaning [Coh91, Ler05, Whi00]. Means [Tur70b]. Measure [Dav15, Dun13, Gin03g, Wil07]. Measured [Moz13b, Moz13c, SSM97]. Measurement [And90, Ben91a, Ben99, Bro88a, Hol10a, Lil99, Sha75, Ste12c, Tum74, Wil88, Wil79]. Measurements [MG83, SS95, SS91b, Wün99, Not24a, SS96, SS97a]. Measures [Mul20]. Measuring [Ben91a, Cha83, Déb91, Gin86b, Gin03g, Lil99, Tur72, Vet12, Tho87]. mécanique [Wil03, Wil87b]. Mecca [Rag01, Hos01b]. Mechanic [Bol05]. Mechanical [War08]. Mechanick [Rud92]. Mechanics [Cro07, Don09, Gin88a, Llo15, De 14, Don09, Kre08]. Mechanism [ASE^{+}13, ASCE14, CTE12, Car17a, Edm11, Eva09b, ECT10, Gin14, Mer77, VMV18]. Mechanisms [Gin90f]. Mecklenburg [Eag08a, Eag08b]. Medal [Ano17c]. Media [Kre13]. Mediaeval [Gin06b, Hos06b, Smo09, Tum74]. Medical [Gen87, Gol12]. Medicine [Ano09a]. Medieval [Ano78b, Arn11, Bar13b, Ber92, Ber12, Boh19, Bra20, CK07, Cha12, Dal78, Dal01, Dav17, Dek12, Dek20, Dob83b, Don09, Eas96, Eas10, Gin06b, Gol98, Gol03a, Gra97, Gra00, Gra05c, Har77, Ken87, Kin95b, Lan89, LD12, Mar14, Mar12b, McC91, McC04, McC19, Mer96, Mol93, Mor08, Moz16, Nor86, Nor88, Nor06, Not19, Plo00, Pou75, Pou84, dP82, Ras13, Rut13, SS91a, SS95, Sal89, Sal95, Sal06, Sav12, SS91b, Ste10b, Ste88, Swe73, Tur03, TA02, UC17, UC19, Ves98, Wal99, WP22, CP23, MSH21, MD21, Not21, SS96, SS97a, Cam11, Moz19, War15, Gol11, Gol11, Gol12, O’B88]. Medii [Gin87c]. Mediterranean [Bel01a, Bre11, Bel97, BEG98, Bre11, CLL^{+}00, Hos00a]. medizinischen [Wes97]. Medusa [Mor01b]. Meeting [Ave95, Bre18]. Meetings [Ano71c, Ano71e]. Megalithic [Atk75, BGP13, Bur82, CFMR77, EB04, Ell85, Far93, GB10, GGCF06, Heg78, Hos92b, HAG94b, HP98, HMA^{+}98, HP99, HS99, HVT^{+}01, Le 08, Le 09, Lil89, Por73, PRA87, Ros84, Rug82, Rug83, RH00, SK08, TT73b, Tho74, TTG76, TM78, TT83, TT84, Cow71]. Mégalithismes [Hos00b]. Megaliths [Bel01a]. Meghnad [DeV85b, DeV94]. Mehrheit [Har84]. Meister [Dup08]. Meisterwerk [Oes10]. Melbourne [Hig12b]. Méliténiote [JT92]. mellan [Sha07a]. Melon [Gin01a]. Melon-Shaped [Gin01a]. Melville [Gin77b, Gin77b]. Memoir [Dew87, Ken94, Eva13]. mémoires [Wil03]. Memoirs [Gin96c, Orc88]. Memorable [Gin10c]. Memorial [Gin91d].

Memoriam [Wen99]. Memories

[Bec01, MHS23]. Memory

[Gin91d, Rad90, Bir74]. Memphis [Mas21]. Men [Dra80, Phi91]. Ménage [SM04]. Menec [TT72b]. Menhirs [TT71]. Menorca [Hos85d, HHK90, HN91]. Menschheit [For83]. Mensurations [Pou84]. Mental [SO00, Rav72]. Mentalities [Par86]. mentioned [Jus23]. Menzel [Bog02, DeV02b, Gin02d, Lay02, Lie02, Ost01, Ost02, Pas02]. Mercator [AH83]. Mercurian [Hoy81]. Mercury [Bro10, Bra23, Dic99, Lan84, Lan80, Nev96, Ros82b, SM94, Van76, Whi75, Lan84]. Meridian [Cou18, Waf01, Wło23, Bre00, Sta99, Waf01]. Meridiankreises [Sta99]. Meridians [KD98]. Meriti [Ban13]. Merovingian [PP82]. mérovingiennes [PP82]. Merz [Rap18]. Mesoamerica [Kru17, Spr93a]. Mesoamerican [Arn95, Hen15a, Mal78, Spr93b, Spr93c]. Mesopotamia [Sol09, Jon06b, Kur21, Bri01b]. Mesopotamian [Jon06b, Kur19, Pin73b, Ste01]. Mesopotamiens [Gol95]. Mesopotamii [Sol09]. Mess [Don88b]. Message [Dic00b, Fin10]. Messager [Van94b, Wil95]. Messenger [Wil15, HDG^{+}13, Fin90]. Messier [Gin10b, Bou12, Gin70a, Gin10b, Gin70a]. Mesztechnik [For82]. metaphors [Mes15]. Metaphysics [Liz17]. Meteor [Ano78b, Dal78, Hoy84, San99, Shi11, UC17, Hoy84, Mar89]. Meteoric [UC16]. Meteorite [Mar89]. Meteorites [Ano78b, Dal78, Gin01e, UC17, Woo88]. meteorology [Bel22]. Meteors [Coo99, Dal78, Dal80, UC17, Yeo99, Ano78b]. Method [Bar05, Bri01a, Car21, Kol08, Moz13a, Neu82, Ost01, Ros84, ZH12, Lat00, Smi94c]. Methodological [Iwa03, RS84].

Methodology [GB86, McM74]. Methods [Ano24a, Das15, Gin90a, GC08, Moz13b, Moz13c, Sid04, Van97, WS19, CML23]. Metodo [Bel11]. Metonic [ASCE14]. Metre [Gin03g]. Metrical [Sid04]. Metrology [TTF76]. Meudon [CL08, CL08, Mal23]. Mexican [Ano01c, Ave99, Ave06]. Mexico [Had84, SS90, FA97, Spr92, Spr93a, Spr00, Tre90, Hiv81]. Mexiko [Spr92]. Mialet [Bra15]. Michael [Bec15, Ger17, Gin03f, Rem15, Rut12, Sha14, Ano17c, EK22, Gra07b, Gra14b, Gra12e, Rut12, SS86, Tre04]. Michał [Bar19]. Michelangelo [Shr00]. Michell [Hos15b, Sch79]. Michelson [Gol73a, DeV75, Gol73a, Swe70]. Microfiche [How80]. Microfilms [Win06]. Micrometers [Bro91b]. Micromètre’ [Déb96]. Mid [Gin77c, Kok98, McC89, Cha98, GJ22, RS22]. mid-19th [GJ22]. mid-60s [RS22]. mid-Fifteenth [Cha98]. Mid-Fifteenth-Century [Kok98]. Mid-Quarter [McC89]. Mid-Sixteenth [Gin77c]. Middle [BFAH^{+}10, Eva82, GC06, Loc83, Loc92, Mor08, Obr14, Rag09, Sam13, Ano78b, Dal78, Eas96, Hos97a, MSH21, Moz13b, Moz13c, Pou75, Rag09, Sam13, Ste96, UC17, Bon09a]. Middle-Eastern [MSH21]. Midi [CL05, RS22]. Miethe [Hen13, Hen13]. Miguel [Cha19, Hei14a]. Mikolaj [Gin08]. Mikolaja [God07]. Mikulov [Sch14a, Sch14a]. Mikulove [Sch14a]. mikulovských [Sch14a]. Milestones [Gin75b, Sho21]. Military [Van95b]. Milky [Ber97a, Don73b, Lad10, Smi94a, Whi00, Don73b, Hos75, Hos04f, Hos08g, Jak71, Jak72a, Lat14, Omo13]. Mill [PG72]. Millennia [Osl06]. Miller [Gol73a, Mit19, Swe70]. Mills [Mit19]. Milton [Gos16]. Mind [BS73, Het76a, Het94]. Minds [Dan19b, Gen90a]. mineurs [Pou96]. Miniature [EM08]. Minime [Pal08]. Minoan [PHP92, PH96]. Minor [Gin08, Kno90a, Pou96, Wil02]. Minora [Voe97]. minutes [Bar19]. mir [Kre86]. Mirror [Cha72, Rom06, Tat09, Hei08]. Mirrors [Hys96]. Misc [Azz11]. Miscellanea [Mos07]. Miscellany [Ber94, Voe97, Voe07]. Misconception [BS92]. Mise [Not14a]. Misri [Ack05]. Miss [Wil07]. Missed [Bou12]. Missing [Gin04c, Wag87]. Mission [Dea94, Doe90, Pet20]. Missions [Tat99]. Missouri [Pet04]. Mitchell [Wel05, Wel05]. Mithraic [vdW90]. Mithras [vdW90]. Mithridates [GC06]. Mittelalter [Ste96, Voe02]. Mittelalterliche [Eva92, Gin00b]. Mittelalters [Eas10]. MLC [Kol01a]. Mnajdra [VSH93]. mnemonic [Fal21]. MOAI [Lil89]. Model [And17, Car19, Car20, CG03, Cha02, Hen75, Kra20, Moz14, Nev96, Nor08b, Pin71b, Sal89, WS19, Whi75, Car23, Car24, dB95]. Modèles [Hos04f]. Models [Duk04, Mor16, Mur98, Swe75, Swe99b, Swe17a, Van98a]. Modena [Pap91]. Modern [Ano06a, Axw17, Azh20, Ber01, Bok73, Bon16b, Coo97, Cro23a, Dar91, Don98, Don09, Dra80, Eis96, GHS08, Gin20, Gin83b, Gin96d, Gin98, Gin11c, Gin13b, Goe09, Gre99, Har85, Hir09, Hoc92, HG80, Hos05a, Jar98, JJ10, JR10, Jar99, Jon15, Kra07b, Kre81b, Lei17, Li19, Mar03, Mar12a, Mar14, McM05, Mer09, Nor08b, Omo20, Par92, PP07, Pin88, Rem11, Roc17, Sal97, Sha14, Sha87b, Smi15, Str81, Tho90a, CM23, Far05, Oos15, dH21, Tri97]. moderna [Mal10, Ano15f]. modernity [Gat24b]. modernos [dA17]. Modes [Cro23a]. Modified [McC98a]. Moesgaard [New77]. Moitié [Bla01]. Mollan [Cha15b]. Monacelli [Mal15]. Mond [Fei13, Hei14b, Hub83]. Monde [Ait86, Mol70]. Mondo [Gin02c].

Mondphasen [Swe99a]. Mondtheorie [Swe99a]. Monks [Sav12]. Montagne [CL05]. Monteiro [FF20]. Month [Bow06, Gol03a, GC13, Sch92, Ste08b]. Months [Mea70b]. Montpellier [Not23a]. Montreal [Ano79b]. Monument [Pau88]. Monumental [AH94]. Monuments [BEC^{+}99b, BGP13, GZB14, GGCF06, HN91, Hos03c, Hos07b, Hos07c, Iwa03, WZR23]. Moon [CT12, Doe94, Dog87, Edd82, Fle73, Gin91b, Kre93, Mor97, Rus02, Swe13, Van02b, Vet19, Wes77a, Whi76a, Whi02, Wło10, Ano22a, Blo78, Bon16b, CG16, Cul02, Da 04, Doe94, Dog87, Gap21, God13, Kra09, MS85, Mon94, MSH20, Pum11, RP87, RC18, Rug99, Sch84a, Sch84b, SP12, SB12, Ste08b, Swe99b, TT83, UC19, Van98a, Van02b, Whi02, Wło10, Wło11, Wło25, Zei86, Ste77]. Moondials [Llo89b]. Moone [Gin97c]. Moons [Coh86a, GvH11]. Moore [Cro15]. Moreau [Hoy84]. Morley [Gol73a, Swe70, Gol73a]. Morley-Miller [Gol73a]. Morocco [BEC^{+}99b]. Morrison [Bak15, Swe17c]. Morrison-Low [Bak15]. Morse [Str93]. Mortifications [Hos14b]. Moscow [Ano15f]. Moses [She15]. Most [Swe17c, Woo87]. Mother [How05]. Motion [CT12, CS02b, Eva93a, Gin87b, Gol02b, Hoo84, Jon04, Kre08, Lan84, Llo15, Mac98, Mur95, Nau01, Rap11a, Sal94a, Ser90, Swe86a, TQ17, VvdS19, Ver06, Wes77a, Wil11, Wes77a]. Motions [DE00, Duk02, Duk09, Het74, Het75, Kno90b, Van98a, CMP24, Whi73b]. Motivations [Cul93]. Motives [HAG95c]. motuum [Wil86b]. Moulton [Bru78b, Bru78c]. Mount [Dic06, Ost89, DeV00b, Dic06, Ost96]. Mountain [Ost89, Sha87c, Mar89]. mouvement [Pin81]. Move [Doe90, Mac20]. Moved [Ker19]. Movement [Ram16]. Movements [Gra01, Smi22]. movimiento [Cha19]. Moving [Sta72, Ste13].

MR0685069 [Gol84]. MR1352683 [Smi96]. MR1658137 [Nau98]. MR1658501 [Nau98]. MS [Azz11]. Mss [Ano74d]. Mu’addel [Kin79]. Much [Gap15, Van15, Gat24b, Eva82]. Mueller [Top03]. Muezzin [Sal06]. MUL.APIN [Oss12, Jon91a, Oss12]. Mulakhkhas [Löh19]. Mull [MR93, RMH91, RM92, Rug85]. Mullen [Mal15]. Müller [Wen99, Wen99]. multispectral [GWZ22]. Mumtahan [Vil88]. Mundi [LN09, Ste88]. mundo [Cha19]. Mundus [Pas15, Van90]. Munich [Pal12, Wes97, Gra05a]. Munns [Sul14]. Muñoz [Rec22, Tir22b]. Muralis [Sch25]. Murder [Gle04, LH03]. Murs [CG09, CG12, PC01]. Museum [Ack10, And00, Gin86b, Gin94, Gin03d, Gub88, Jas96b, Llo90, ML07, dP82, Ree11, Sch14a, Sym02, Tur70a, And00, Ano76c, Dav19, Gin04b, Lat20, Gin88c, dP82, Van99]. Music [Coh95, Dan19b, Gin85b, Gua09, Hos02a, Mar97, Swe76, Mar97]. Musiktheoretiker [Swe76]. Musikwissenschaft [Swe76]. Muslim [Ric07, Coo99, Kin83, Kun80, Rag85, SS96, SS97a]. Mutation [Lüt00]. Mutio [Rem11]. muzea [Sch14a]. My [Top03, Dew87]. Mycenae [Mar02]. Mysteries [Kra13, Sha87c, vdW90]. Mysterious [And17, Gin01e]. Mysterium [Ait86, Wil82, Wil82]. Mystery [Ald11, Gra12e, Mor01b, Nor08a]. Mystic [Dav08]. Mysticism [Nor78]. Myth [Ald97, Aub16, Bel22, Mer92, Mil80, Mor01b, Ost92a, PG72, Por73, TH79, TH79]. Mythology [Mor97, Sis85, Zso20, Kru97]. Mythos [Har84, Goe09]. Myths [Ave09, Bow10, Cer99, Ell08, Kru97, Mor97].

N [Bre18, Rad90]. nach [Voe02]. Nachale [Wer10]. Nachrichten [Swe98]. Nachtwolken [Ash79b]. näher [Kre86]. Nahmias [Swe17c]. Nairne [Tur79b].

Najm [Ack05]. Nakovana [FJ11]. nakusatras [PM89]. Named [Kun91]. Names [Gin06b, Kun91, Pin88, Sol09, Wil02, Kun90, Wil02]. Naming [Cas19, Kun90, Van02b, Con24]. Napier [Car86]. Naples [Cen17, Loc95]. Narratio [Gra80, MG86, Tre04, God16, She74]. Narrationes [MG86, Swe86a]. Narrative [Lia19, UC19]. nas [LN09]. Nasīr [Ves73]. NASA [Ram06, Waf91, Han16, Ram06, Smi07, Smi09b, Smi07, Smi09b]. Nascita [Cen17, Chi16]. Nasim [Kes15]. Nasir [Ken94]. Nasiriyya [RPPA15]. Nasser [Gin99a]. Natacha [Fin19]. Nath [Hea15]. Nathaniel [Mon82]. Nation [Hir92, Vet19]. National [And00, Bla98a, Cah19, Dic91b, Edm84, Gat20, Gin03d, Grö83, Gub88, Hod77, Kee91, Lat20, Nee91, Ree11, Smi91, Sul99, Tur70a, Vil90, And00]. Native [Rom06]. Nativitatibus [Jus20]. Nativitatum [Jus20]. Nativities [Jus20]. Natural [Bru78a, Bru09, De 13, Don98, Far04, Gin01f, Her90, Llo94, Mil04, Pin92, Rap15b, Swe01]. Naturales [Pin92]. Naturalistic [Mon94, Oss12]. Nature [Dob79, Gee18, Gin90c, Gin92a, Gra09, Lat00, Lip00, Mar03, Pau78, Cro23b, Sch78, Big10, Gin92b, Lin93]. Naturforschenden [Kel78]. Naturphilosophie [Sha94a]. Naturphilosophische [Pin92]. Naturwissen [Ano06a]. Naturwissen-schaften [Ano06a]. naturwissenchaftlicher [Sev00]. Naturwissenschaft [Sha85, Sha89]. nauce [God13]. Nauchnye [Rya71]. Nautical [How80, Rot01, vdW17, vdW19]. Naval [Dic91a, How88, Huf91, Ost04, Rot01]. Navicula [Sav12, Sav12]. Navigate [Dev14b]. Navigation [And90, dA74]. Navigator [dA74]. navigators [dH21]. Nazca [Rei92, Rei92, Sha87c]. Nazis [Ano97b, Wes97]. nazvaniia [Sol09]. Near [Haw74, McC17, Roc14, Sid10, Bla98b, Ext92, Nor88, PRA87, Roc08, Sch23a]. Near-Eastern [Sid10]. nebo [Sol09, Wer10]. Nebula [Hos82c, Tee89, Van14b]. Nebulae [Ben76a, For80b, Gin70a, Gin92a, HB71b, HB71a, Het74, Het75, Hos76c, Hos79b, Hos90, Hos05e, Hos06c, Hos08g, Hos11b, Hos11c, Hos11d, Hos13, Hos14e, Jon86, Kra24b, Lay02, Mac10, Pau78, SIN85, Kes15, Nas13, Ben76a, DeV96, Gin92a]. Nebular [Bru78a, Num72, Ost01]. Necropolises [Bel97, Bor98, SK08]. Nederlandse [Hal03, Van02a]. Need [MG83]. Neglected [Dra86, Van02a]. Negotiating [Bel20, HS92]. Neighbouring [GB10, HS99]. Neil [Mol14b]. nel [GM03, GM03]. nella [May11]. nell’Astronomia [Swe92b]. Nemore [Pou79]. Nemrud [BG10]. Neolithic [Rug99, Hos03c, Hos07b, Hos07c, Sto94]. Néolithiques [Hos00b]. Neoplatonists [Jon99]. Neptune [Cas19, Cha88, Gap15, HS92, Sad90, Smi83, SB84]. Nernst [Kra95]. Netherlands [Str19, Sch03]. Networks [Mir17, Gre23, Kwa12]. Neuchâtel [GJ22]. Neue [Swe76]. neuerer [Kle14]. Neufeld [Ger17]. Neugebauer [Roc17, Swe93, Jon15, Ken85, Roc17]. Neuve [Eva01]. Neuzeit [Dan11, Har84, War13]. Never [Sha87a]. Nevil [Gen90b]. Newcomb [Smi94c, DOL98, Smi94c]. Newham [Ano82e]. Newly [Moo94, Rus02, Str16a]. Newly-Discovered [Moo94, Rus02]. Newspaper [Nas20]. Newsworthy [Ano15d]. Newton [Ben87b, Dic99, Fei95, Gin95, Mel89, Nau98, Sha87a, SD82, Ste90b, Wes77a, Ben87b, Bro88b, BF13, Coh80, CS02b, Eva93a, Fei95, Gau87, Gin90e, Gli88, Hat04, Hos73, Hos77, Hos86a, KY95, Nau01, Nau16, Ruf10, VWH97, Wes77a, Whi70a, Whi70b, Wil87b, Coh80, CS02b, Nau98, Nau01, Pan15, Smi96, Str10, Wes90]. Newtonian [Hat04, Hos86a, Ben81b, Hos85c, Hos08f, Sha87a]. Newtonianism [Far04]. Next [Yeo00]. NGC [Ash81]. Nguni [Sne98]. Niccolò [Omo11a]. Nicéphore [Dob86]. Nicholas [Gin04c, Nor83, Pou96, Fle92, Pou96, Too81]. Nicholas-Louis [Gin04c]. Nicholaus [GC08]. Nicht [Gau10]. Nick [Rap15a]. Nicolas [Kno90a, Pou96, Cha19, Déb13, Kno90a]. Nicolas-Louis [Déb13]. Nicolaus [Bar19, Dan04b, Gau17, Gin05a, Swe98, God07, Dob96, Ros73, Swe88]. Nicole [Lin07]. Niels [Wło23]. Night [Lad10, Sch07a, Tip88a]. Nights [War80]. Nin [Mul82]. Nine [Hos92a, Mar02]. Nineteenth [Cha15b, Eis85, Gin87a, Hos04d, Kes15, Mer11, Nal16, Rot80, RC11, Tur72, Waf02, Mol14a, Bre13, Bro91b, Chi15, CO13, DOL98, FV17, Gen87, Her73, Hos04d, Mea85, Nas13, RS18, RC11, Smi91, TG87, War98, Sut72]. Nineteenth-Century [Cha15b, Eis85, Hos04d, Nal16, RC11, Waf02, Mol14a, Bre13, Chi15, CO13, DOL98, Gen87, Hos04d, Mea85, RC11, TG87, War98]. ninth [NKM^{+}16, Kin87, Sal94a]. Ninth-Century [Kin87, Sal94a]. NJ [Ree20]. No [How94, Ros86, Ano95a, Ano01c, Ave95, Gol84, Had84, Nak71]. Nobel [Kra17]. Nobis [Wes97]. Nobody [Dan04b]. Nochmals [Hor74]. Noctilucent [Ash79b]. Nocturnal [Gib82, Gib82]. node [Ler80]. Nodes [Tho71a]. Nodus [Ler80]. Noel [KE22]. Noise [McC13b]. Nolthius [Kus11]. Nomenclature [Ash81, Wag88, Van02b]. nommées [Kun91]. Non [HB71b, HB71a, Jon04, Mor16, Moz16, Pal08, Rap11a, Rug04, HRVM13]. Non-Astronomical [Rap11a]. Non-Galactic [HB71b, HB71a]. Non-Ptolemaic [Mor16, Moz16]. Non-Uniform [Jon04]. Non-Western [Rug04]. Noon [All92]. Nordenmark [Bre18]. Nordström [Bre18]. Normal [Rav73, RC92]. Normale [Lip15]. Norman [Big10, Din73]. Norris [DK83b, Doe01, KD90]. Norse [Kru97, Jor94]. North [Ave09, Ben84, Bur82, Had84, Iun02, Vet12, Ano80e, Ave09, BD82, Gin09, Hos00c, MR93, RMH91, RM92, Hos00c]. North-West [Bur82]. Northern [Dav17, Sis85, BEC^{+}99b, Hos97c, HMA^{+}98, Smi18a]. Northridge [Wil87a]. Northwest [EBB^{+}01, HC99]. Northwestern [Lep14]. Norwegen [Jor94]. Notations [ASS13]. Note [Ano24a, Dew82, Eva15, Gau87, Lan80, RS84, Shr00]. Notebook [Don96, Don00]. Notebooks [Wes84]. Notes [Ano70c, Ano70d, Ano71f, Ano71g, Ano71h, Ano72c, Ano72d, Ano72e, Ano73g, Ano73h, Ano73i, Ano74e, Ano74f, Ano74g, Ano75d, Ano75e, Ano75f, Ano76e, Ano76f, Ano76g, Ano77d, Ano77e, Ano77f, Ano78e, Ano78f, Ano78g, Ano79e, Ano79f, Ano79g, Ano80f, Ano80g, Ano80h, Ano81c, Ano81d, Ano81e, Ano82f, Ano82g, Ano82h, Ano83d, Ano83e, Ano84c, Ano86a, Ano86b, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89f, Ano89g, Ano89h, Ano90b, Ano90c, Ano90d, Ano90e, Ano91d, Ano91e, Ano91f, Ano91g, Ano92b, Ano92c, Ano92d, Ano92e, Ano93b, Ano93c, Ano93d, Ano93e, Ano94d, Ano94e, Ano94f, Ano95e, Ano96c, Ano97d, Ano97e, Ano97f, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e]. Notes [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03b, Ano03c, Ano03d, Ano03e, Ano04d, Ano04e, Ano04f, Ano04g, Ano05b, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12b, Ano12c, Ano12d, Ano12e, Ano13b, Ano13c, Ano13d, Ano13e, Ano14c, Ano14d, Dob01, Gin96c, Jar06, Jon86, Kno90a, Rag16b, Sha92, Wło90]. Nothaft [McC19, Zep19]. Nothing [Eva82]. Notices [Ano70e, Ano70f, Ano71i, Ano71j, Ano72f, Ano72g, Ano73j, Ano73k, Ano74h, Ano75g, Ano75h, Ano75i, Ano76h, Ano76i, Ano77g, Ano77h, Ano78h, Ano78i, Ano78j, Ano79h, Ano79i, Ano79j, Ano80i, Ano81f, Ano81g, Ano81h, Ano82i, Ano82j, Ano82k, Ano84d, Ano84e, Ano84f, Ano84g, Ano85a, Ano85b, Ano87e, Ano87f, Ano87g, Ano88e, Ano89i, Ano89j, Ano90f, Ano90g, Ano91h, Ano92f, Ano92g, Ano93f, Ano93g, Ano93h, Ano94g, Ano95f, Ano96d, Ano97g, Ano97h, Ano98g, Ano99f, Ano99g, Ano99h, Ano00f, Ano00g, Ano00h, Ano01i, Ano02g, Ano02h, Ano02i, Ano02j, Ano03f, Ano03g, Ano03h, Ano04h, Ano04i, Ano04j, Ano05c, Ano05d, Ano05e, Ano06b, Ano06c, Ano07f, Ano07g, Ano08e, Ano08f, Ano08g, Ano08h, Ano09f, Ano09g, Ano10e, Ano10f, Ano10g]. Notices [Ano11e, Ano11f, Ano12f, Ano12g, Ano12h, Ano13f, Ano13g, Ano13h, Ano15e, Ano15f, Ano16b, Ano16c, Ano16d]. Notorious [Sch01b]. Notre [Hos04f]. Nouveau [Kus10]. Nouvelle [Gin07b, Kus10]. Nova [Gin82c, Gra05b, Gra11b, Gra14b, Por74, Rec22, Sch23a, Ste12b, Ham06, Ste02b, Ste12b, Dav15, Ste02b]. Novae [Bar13a, Dal80, Hei14a, Hos76c, Hos07a, ZL99, Wil86b]. Novara [Swe73]. Novas [Gra12a, Hei14a]. Novel [Hen08, Ree16, ZH12, Lep14]. Novelist [Don82a]. Novelties [Gra09, Gra11b, Oos15, TB13, Voe94]. Novelty [Wło07]. November [Jas96b]. Novum [Rap11a]. novus [Hos89b]. Nr [For73]. NRAO [Sul99, Wen24]. NT [Fat00]. Nub [Wil08a, Wil09]. Nubia [SB05]. Nuclear [Ost05a]. Nuevo [Gin82c]. Nui [GG14]. Number [AMP95, AMP96, Mul82, Tur79b]. Numbers [Gin06b, Lat17, Moe75, Ste14]. Numerical [Ste23]. Nuncio [She74]. Nuncius [Fin90, Lüt00, Van94b, Hos89b, Hei12, Whi78, Fin10, Fin90, Fin10, GV03, Hei12, Van94b]. Nunes [Mal13, Mal13]. Nuraghes [Ano04b, ZB04]. Nuraghic [HAG93]. Nuremberg [Llo89a, Mos03, Ham11, Kre80a, Llo89a, OP89]. Nürnberg [Ham11, Kre80a]. NV [Llo15]. NW [HGGVE^{+}23].

O [Bee06, Mil08]. O. [EP22, Hei12]. Oases [BS06]. Oasis [SBF07b, SBF07a]. Obervational [Ost96]. Obituary [Ano01j, Atk86, Coh86b, Hos85b, Too84]. Object [Bel20, Ben82b, Low88]. Object-Glass [Low88]. Objective [Meu10]. Objectivity [Ber89]. Objects [Bec19, Gin10b, Hos13, Ste11, Gin03c]. Oblique [Neu82]. Obliquity [Jon02]. Obras [Mal13]. Obscuration [Gau12]. Observable [ABB03]. Observaciones [Ave87]. Observation [Ash80, dB70, Cul07, DN98, Gin11b, Gra13a, Jon12, Ken87, Kre81b, McC90, MS85, Mon94, Mül10, Ost92a, Sal87, SCV20, TT83, Whi02, Wil97, Wil95]. Observational [AH94, Bat15, Eis01, For80a, Gie20, GB09, Jon71, Ost95, Ost97b, Ost98, Shi11, TA02, dJ12, de 22]. Observationes [CR22]. Observations [Ano03a, Ave05, Ave11, Bri05, Chi15, Cro00, Dek08a, Dra76, FS96, Fin08, Gin97c, Gin03c, Gra10, HS81, Har77, Hoc92, HD84, Kol07, Kre81b, KKSS21, Lan80, Liv13, McC83b, Moe75, Moz19, New77, Rug87, SS91a, Sal85, SIN85, Shi00, Sta06, SS98b, SG07, Ste19, Ste77, SF95, SB12, TMG02, Tre90, Wes79, Whi78, Wil11, Wło07, Yaz14, dJ12, vH86, Cas21, KG82, Kre80b, MD21, Not23a, RS22, Sch23b, SP22, TPYR21, Wło23, dB95, SS97a, Dog87, Mor01a]. observatoire [Feu75]. Observatorien [Her76]. Observatories [Ano77l, Bel09a, Bla79, Con70, Dic91b, Fai94, Gin90d, Her73, Her76, How86, How94, Hut12, Kee91, Ken82, Kri94, Mul81, Mül91, Ros84, RC11, Smi07, TTG76, TT84, Cha21, Mal23, Moz21, Hos82a, Bla98a, Bre00, Cow71, Gin90d, Maf05]. Observatórios [Gra12b]. Observatorium [Eel83]. Observatory [Ban13, Bel07, Bla98b, Bog02, Bry72, Cah19, DW83, Dic95, Dic06, Doe19, Dup90, Edd90, Hea99, Hos16a, How88, Jar91b, MH88, Mul20, Nee91, Orc84, Ost04, Rie13, Rot80, Rot01, RS22, Smi89, Smi91, Smi92, Sul99, TT73b, TT75, TTT75, Tur02, SP22, Ano82a, Ano90h, Ban10, Bel07, Bla91a, BH87, Can12, Dic91a, Doe90, DH73, Edd71b, Edm84, Edm98, Eel83, Eis88b, Feu75, For70b, For71b, For71c, For74a, For74b, Gin00c, GDEK24, GJ22, Ham09, Hea94, Hea99, Hig12a, Hod77, How88, Kär24, KM91, Lan87, Lan88, Lie02, Mac15, Mac20, Mac90, MLH84, Orc84, Ost02, Ost04, Ost07, Pau88, Rei82, Rob90, Rot80, Rot90, Sal85, Sch91, Sch18, Sha71a, Ste90a, Tho74, TMG02, TPYR21, War78, ZM20, Zru21, Bel77, Cha90, Dew79]. Observatory [Edm84, Eis88b, Gin00c, Hea94, Hea99, Hig12a, Hig17, Hos92a, Lan88, Pau88, Sch91]. Observed [Apt99, Bou13, Coy07, GC99, JKKC19, Pan12, Ste00b, Wil97, Sch23a, Bec99a]. Observer [Ano01b, Hos05b, Hut12, Sch01a, Ash82]. Observers [Smi12]. Observing [Ben14a, Cro00, Nas13, Pan94b, Pan95, Van15, Wło87, ZJJ^{+}99, Wło22, Hos11b, Kes15]. Obsession [Sch07b]. Obstacles [Cha21]. obstruction [Kut23]. Occhiale [GV03]. Occidens [Gin85d, Bee71]. Occident [Ler11b, Ste05]. Occidental [Hos08b]. Occidentale [Hos00b, Hat88]. Occult [Par86]. Occultation [GC99, NN24, Ste00b, UC19]. Occultations [MS85, MHS17, SB12, WKH18]. Ocean [How88, Ost04]. Octants [Ree11]. October [Sch23a]. Octogenarian [Ber70]. Oddi [Rem11]. Odyssey [Gin03g, Had84]. OEuvres [Pou96, Rag92, Wil03]. off [Eva13, Jar91a]. Office [Rot01, How08, Rot01]. Offset [PC87]. Offusius [GD93]. Ohio [HH82, McC85]. ohne [Mos86]. Oklahoma [Gin77d, Gin77d, Sal97]. Olaf [Nor98]. Olber [Sch70]. Olbers [Har92, CO13, Hos85c, Jak70, Sch70, Tip88a, Tip88b, Har92]. Old [Ber97a, Dob00, Eva09b, Hig12a, Jor94, Kru83, Li19, Mul82, Ste12a, Zso10, CM23, Gra21b, Bel01b]. Oldest [Hos04e, SM05, Sta04]. Ole [Wło23]. Olschki [Fin19, Oos15]. Olsztyn [SW18]. Omar [Kes15]. ombre [Par85]. Omens [Gin88a, Ste99, Ste06a]. Omnia [Hat08, Hus13, Gin08]. Omocentriche [Swe92b]. One [Eis88a, Gle04, Kra22, Llo11, OT13, Goo09]. Online [Ban12]. Ontologies [Mal15]. Onus [Sch15]. Oort [Gin99c, Pau82, Gin99c, Pau82]. Open [Doe19, Fal23, Gen90a, Sil23]. Opera [Gin78b, Hat08, Her72, Hus13, Nor86, Gin08]. Operation [Gra97, Rug95]. Opere [Gin10a, Ser88]. òptica [Ano15f, Mal10]. Optical [O’D06, Sha02, Wło25]. Optician [Eel74]. Optics [CM17, Dij09, Dup14, Lin07, Lüt00, Max12, McC91, Sha02, Dew70, Gin97b, Mor08, Sha02]. Opticus [dC01]. Optik [Dij09, Dup08]. Optiker [Ost05b]. Optimism [Eis01]. optisches [dC01]. Oracular [Ave18]. Orbit [Don73c, Gra19a, Not24a]. Orbital [Moz13b, Moz13c, Moz19]. orbites [Kre81a]. Orbits [FF20, Ruf10, VWH97, Don93]. orbium [Cha19, Eva00, Gin77a]. Orbs [Eas96, Gin04b, Ano83b, Jar82]. Orchiston [Sil23]. ordena [Rya71]. Order [How05, Lat00, LG12, Mer92, Bra23]. Ordering [Cau08]. Oresme [Lin07, Lin07]. Organisation [Hos78a]. organized [Ave95]. Organizing [Ano95g]. Oriel [Spe80]. Oriens [Bee71, Gin85d]. Orient [Ste05]. Oriental [Siv89, Ste94]. Orientalist [Nor80]. Orientation [AZB02, Bel01b, BS06, BSF08, BFAH^{+}10, BGP13, CKE11, GZB14, GGCF06, Haa84, HK82, Hos91b, Kin95b, PP82, RAGGB18, SB05, SBF07b, SBF07a, HGGVE^{+}23, UAGGB21]. Orientations [Ano04b, AR00, Bel01a, Bor98, Che99, EBB^{+}01, GB10, GGCF03, HHK90, HN91, Hos92b, HAG94a, HAG94b, HAG95a, HAG95b, HZ97, HP98, HMA^{+}98, HP99, HS99, HVT^{+}01, Hos02c, Hos03c, Hos07b, Hos07c, Hos08h, Iun02, Le 08, Le 09, Lil89, Mar02, NR14, PHP92, PHP93, PH94, PRA87, RBH00, RH00, SK08, Sau00, SHV92, TBSKO06, TSH92, ZHGM96, ZB04, Zed05]. Origanus [Omo11b]. orígens [Ano15f, Mal10]. Origin [Dic13b, Duk05, Eva87b, Eva87c, Gen92b, God06, Gol02a, Hom72, Ler05, Mul80, PC87, Pin71b, Pin73b, Sch04a, Sch05a, Sch04b, Swe74a, Wil87d, BF13, Kur21, Oss24, Pan15, Wil87c]. Original [Jak72b, Kel10, Swe86c, Van71b, Whi71, Van71a]. Originality [Bar05]. origines [Har85]. Origini [Cen14]. Origins [Ave97, Bec01, Bon07, Bon09b, Cen17, Dog87, Eva93b, Gin90e, Gin01c, Gin01g, Gin02b, Gin12c, Gra07a, Hos85c, Hos10a, Jar91b, Ler06, Mac15, McC11, Pau88, Ped81, Spe13, Wil84, Zuj11, vdW90, Chi16, Smi79, Edd82, Liv13, Pin73a]. Orkney [TT73b]. Orleans [Dew87]. Oronce [Pan10, Rap11b, Rap11b]. Oroscopi [Jon97]. Orr [Num72]. Orthodox [Sev00]. Orthodoxer [Sev00]. OSCURAUIT [UCI16]. Osmanli [Hei00]. Osservatorio [Can12]. Other [Bro83, Lau14, Par92, Ree20, UCI16, Ave05, Ber94, Bow10, Bra90, Cro12, DeV85a, Hei12, Ric07, Sol09, Str99]. Otrazhenii [Wer10]. Ottawa [Jar91b]. Otto [DeV89, Hos80a, Roc17, Jas20, Jon15, KI10, Swe93, KI10]. Ottoman [Hei00, Hei00, Sch20]. Ottone [Cro23a]. Our [Arm73, Bia04, Cro12, Dev14b, Dic13b, Dun20, Hos75, Lad10, Hut13]. Ours [Bro83]. Outer [Cas19, Rea84, TW25]. Outline [DeV05, Doe05, Hol03, Pal84, Ros81, Sam98, Tau09]. Outstanding [Gin12a]. Oval [Don73c, Don96, Don93]. Ovalis [Dav09]. Overseas [Dea94]. Overview [BSF08, Dic91b, Hos02d, Rug07, Wil05a, vG05]. Ovid [McM01, McM01]. Ovidian [McM01]. Owe [Gin85f]. Owen [Jar90, KE23]. Own [Arm73, Gib82, Gin00c, Sha12, Udí21]. Ownership [Dav19]. Oxbow [Kru17]. Oxford [Kru17, McC19, Smi96, Ano94a, Ano96a, Car00, CG16, CG18, Hos96, Spe80, Vic85]. Oxygen [Plo77]. Oxyrhynchus [Gol00].

P [Ave18, Eva87a, Löh19, Mer06, New77, Sul14, JKKC19, OP89]. P. [Jon16, TF16]. P/Crommelin [JKKC19]. P/Halley [OP89]. PA [Ano15f, Bec19, Bre18]. Padua [Cha07, Ves98, Tur75]. Page [GV03]. Pages [Vai12]. Painting [Cam02, Ste00a]. Països [Hos06b]. Pakal [Cog08]. Palenque [Cog08, Ros82a, Cog08]. Palermo [Chi09]. Palgrave [Ano15f]. palimpsest [GH24]. Pallas [Bou12]. Palmyra [Loc95]. Palomar [Yav99, Ost97b, Ost98]. Pamyatniki [Rya71]. Pantelleria [TSH92]. Paper [Kre11, Ano15f, Bec15, Boh19, Bur15, Cen17, Kru17, Llo15, Mit19, Oos15]. paperback [Bra15, Cen14, Hea15, Hei14a, Jar90, Kle14, LF15, Lep14, Mol14b, Mul83, Nau98]. Papers [Bar13b, BH81, Duk05, For70b, For71b, For71c, Gin92a, Gin99c, Had84, Jas96b, Ken85, Mil80, Min00, Ost97a, Rad90, Ste05, War78, Wis06, TH79, Wil87a]. Papirologica [Jon97]. Papyri [Gol00]. Papyrus [Jon97, Jon16]. Paradigm [Rav73]. Paradigmatic [Hei12]. Paradigms [Het94]. Paradosis [Bar18]. Paradox [Hos85c, Sch70, Tip88a, Tip88b]. Paradoxes [Wer13]. Paralipomena [Sha02]. Paralipòmens [Mal10, Ano15f]. Parallax [Atk82, CT93, Sie05, Wil79, Huf91]. Parallaxes [GB09]. Parallel [EZ16, Gol11, Gol12, Jus20, Mil04, War19]. Parameters [Sal85, Wil84, Gin90e, Wil84].

Parapegma [ASE^{+}13]. Parapegmata [Sid10, Sid10]. Paris [Cha90, Gin03a, Gin11c, Lan87, CG18, Déb96, Fab19, Feu75, Hat83, Hen76, Kel83, Lan87, Lau07, Ler00, Mal23, Not24b]. Parisian [Not15b]. parisien [Mol76]. Parisiens [Gin88b]. Park [Had84]. Parmenides [Gre14, Llo15]. Pars [Nor86]. Parsons [Cha15b, Mol14a, Smi02, Smi02]. Parsonstown [Hos02b]. Part [AMP95, Gin86b, Hos11c, Jar90, JLGL14, LGJLF14, Sha02, Ste99, Ste06a, Wes90, Wil86b, Ano76c, Ano05a, AMP96, BH73a, BH73b, Bru78b, Bru78c, ESBG96, ESBG97, Eva87b, Eva87c, For70b, For71b, For71c, Gre04, Hen99a, Hen99b, Hos11d, Ito09, JLS^{+}05, JHLS05, Moz13b, Moz13c, Ost84b, Ost84c, Ost97b, Ost98, SB72b, SBF07b, SBF07a, Smi08, Smi09a, Spr93b, Spr93c]. partes [Pou84]. Partner [Hos04a]. Partnership [Bia04, Cor03]. Parts [Mes15, Fre14]. Partying [RC16]. Paschali [Nor92b]. Passage [CKE11, Dew87, Gen92a]. Passion [Ver12]. Past [Ano76b, Ave94, Doe08, Eng02, Hos94, McC07, Smi20, Whi76a, Wea90]. Patent [Bak15, GMML14]. Path [Ano22a, CB05, Mar14, Moz14, Swe13, WDR11, Gap21]. Paths [Wes09a]. Pathways [Roc16]. Patience [Sta06]. Patras [Gin82d]. Patrick [Cro15, Oos15]. Patrizi [Bon09b]. Patronage [BG03, Bla91b, Blu13, Bru96, Rot90]. Patterns [Ave09, Van89, Whi92]. Paucker [Oes15, RR13]. Paul [Ano21b]. Paulerinus [HH07]. Pauper [Hev94]. Pawnee [Had84]. Paying [Rot17]. Payne [DeV85a, DeV03, DeV85a, DeV03, Tri13]. Payne-Gaposchkin [DeV85a, DeV03, DeV85a, DeV03]. Payne-Scott [Tri13]. pbk [Doe19, Fin19]. PC [Ols93]. Peace [Van95b]. Peak [Edm84, Lie02, Edm84]. Peck [SM04]. Pecked [Ave05, RS84]. Pedersen [Duk11, Nor98]. Pedro [Nor00b, Mal13, Nor00b, Mal13]. Pèlerin [Tur96]. Pèlerin’s [Tur96]. Peloponnes [Eva08]. Pendulum [Déb05]. Penetrating [Ben76b]. Peninsula [GB10, Kre02, RAGGB18, Iwa96]. Penser [Gra07a]. People [Dic00a, Sis87]. Peoples [Sne98]. Perception [Dic90, Lüt00]. Perceptions [Blo78]. Percival [Het78, Str93, DeV02a]. Perfect [Dan12a, Moe75, Yav99]. Perfection [Gin87b]. Perga [Jon99]. Perigee [Whi75]. Perihelion [Dic99, Lan84, Ros82b]. Period [BS92, Heg82, Hei00, Mer09, Nic90, Vai12, Wil07, Don09]. Period-Luminosity [Wil07]. Periods [Gol02b, Hos83b, McC83b, TQ18, Hos83b]. Peripatetic [Hos88]. perit [Ham06]. permanentes [PC01]. Perrin [Wes09a]. Perrine [Hod77]. Perseverence [Bec00]. Persian [Bar18, Gol72]. Person [God07, Dob83a]. Personal [Cah19, Eva13, Gin91b, Hir09, Rie13]. Personalities [For74a]. Personality [Ald02]. Perspective [Bel01a, Cor03, Mer11, Rot01, Ste04]. Perspectives [Ald09, Apt99, Eis96, Gin05d, Gol94, Hir92, Sch90, Tau11, HE82, Eis96, Ano84a]. Persuasion [Bec00, Gin03h, Rot09]. Peru [Ave87, Ave87, Gli88, Sha87c]. pervoy [Rya71]. Pesic [Dan19b]. Peter [CWW17, Dan19b, Mal90, Sil23, Smi18a, GE24, Kul87, Kun87, Ves98]. Petersburg [Kri97, Oes15, RR13]. Petersburgskaya [Kul87]. Petit [Pin80]. Petri [Nor86, Nor86]. Petroglyphs [McC85]. Petrus [Gra11a, Kre07a]. Peucer [Gin77c, Gin77c]. Peurbach [Bar13a, Bar22, DK96, Lan98, Mal13]. Pfaff [Oes15, RR13]. Phaenomena [Eva12, Gin03b, Eva12]. Phanes [Pap91]. Phänomen [Sis87]. Phases [Dra84b, Gin84c, Gol96b, Pal01, Pan12].

Phenomena [Gol97, Sch87, Tih07, HGGVE^{+}23]. Philadelphia [Ano15f, Gin80a, Jar90]. Philipp [McC19, Zep19, Sta00a]. Philippe [Ler11a]. Philolaus [Tau96, Tau96]. Philomenae [Nor86]. Philosophandi [May11]. Philosopher [CJ05, Tes10, Van98b, Wil14b]. Philosophers [Don79, Don98, Kre82, Mil04]. Philosophical [Ano15f, Don71b, Eis96, Jar90, Kel83, Llo94, Mol93, Ree16]. Philosophically [Ber19]. Philosophy [Ben81a, Don02, Gin96d, GM03, GB86, Hos00c, Lat00, Mar14, Mol70, Rap15b, Swe01, Wes94]. phisicos [dA17]. Phoenix [Sch78]. Photoelectric [Lat20]. Photograph [Bec00]. Photographic [Gin12a, Hen99a, Hen99b, SS16, Sta04]. Photographing [Moz21]. Photographique [LH05]. Photographs [Bec13, Ben87a, Kes10, POH96]. Photography [Lan83, Ost94, Zru21, Cro00]. Photomechanical [Pan95]. Photometer [Sta00b]. Photometry [Lil99]. Photosphere [TMG02]. Physical [Ano84b, Bec19, Bur10, Cas18, Dar91, Dog87, Edd82, Gra76, Hos73, Kra12, Mur95, Sha07b, Tho88, Wes84, Wes01]. Physician [Omo17]. Physicist [Ost92b]. Physicists [DeV00b, Sil23, DeV07]. Physics [Bar88, Ano81a, Coo97, DK83a, DK83b, DeV00b, DeV94, Edd71a, Gra76, KD90, Ost05a, Sta13, Tho88, Tur79a, Whi92, Gim12, Mel09, TH79, Coo97, DeV05, Edd71a, Mil80, Sch11, Ste12c]. Physik [For82, Her90, Wil01]. physikalisch [Wes97]. physikalisch-medizinischen [Wes97]. Physiktheologie [Her90]. Physique [Can09, Ste12b]. Piam [Wen99]. Piazzi [Eva87a, Hos88, CMO11, Ser90]. Pic [CL05, RS22]. Picard [Cha89, Cha89]. Picchu [DW83]. Pickering [Bed10, Plo90, Plo93, Sad90]. Pico [Azz06, Van23]. pictorial [WZR23].

Picture [Eva04, Ost94]. Pictures [Hos02a, Smi09b, Van98c]. Picturing [Rem15, Ben14b, McC13c]. Piero [Sha11]. Pierre [Eva17, Hat08, Mal97, HRVM13, Hat08, Hei99, Sti07]. Pierre-Simon [HRVM13, Hei99]. Pies [Jar09a]. Pietro [Rap11c]. Pigott [Hos79a]. Pike [Gen87]. Pillars [Mit19]. Pin [CTE12]. Pin-and-Slot [CTE12]. Pingree [Jon06a]. Pioneer [Ber74, Edd86, Gin82b, Kid87, Lut91, Wło25, Hir09, Bed10]. Pioneering [Bed10, Hir85]. Pioneers [Bec15, Bel77, Gol87b, Hos14c]. Piotr [Gur04]. Piper [Mos86]. Pirámides [Hos12a]. Pisa [Lip15]. Pisma [Gin08]. Pittsburgh [Bec19]. Pivotal [Hos76a]. Pivots [Kok98]. Place [Cha85, De 06, Dew79, O’B88, Ves98]. Placeholders [Bec19]. Places [Jar98]. Plagiarism [Cle71]. plana [Pou79]. Plane [Déb91]. Planet [Cro15, Cro00, Dic99, Gin97c, Lan82, Rea84, Rea92, Sad90, Sch82, Wil97, Wil02, Yeo07, She14, Sul12]. planétaire [Ano83b, Pou83, Nor82]. Planétaires [Yan86]. Planetaria [Kru08]. Planetarians [Mul80]. Planetarium [Ack10, Ano73f, Dun20, Muc75, Tur75, Van99, Dun20]. Planetarum [Byr11, Pan12, Mal13, Ped81, Hus13, Bar13a, Swe73]. Planetary [Ano13a, AH94, ABB03, Ber97b, BSL86, Bri98, CG13, Coo97, CO13, Dek08a, Doe00, Doe05, Duk04, Duk09, Eas00, ECT10, Gol97, GC13, Gra05c, HS81, Hos14e, Hub83, Jon04, KH73, Kno90b, Lau18, LG12, Li19, LD12, Mac93, McC04, Mul80, Mur95, Omo11b, Pin70, Pin71b, Plo16b, Sch80, Smi97, Ste99, Ste01, Ste03, Ste06a, SH81, Swe72, Swe73, Swe75, Swe04a, Van94a, Vil88, Wes90, Whi72, Whi73b, Whi74, WKH18, Yan86, Yeo88, Bra23, CMP24, Pou83, TW25, Gin91b, Coh86a, Gra05c]. Planeten [Hub83]. Planetenlaufuhr [Oes10]. Planetentheorie [Whi72]. Planètes [Nor82, Kre81a]. Planetology [Str93]. Planets [Bau87, CTE12, CH92, Cro12, Dic90, Doe00, Eas01, Gin72, Moz13a, Mul80, Rea84, RC18, SB12, Swe84, Swe89, Swe17a, UC19, Van98a, TW25, Gin91b, Bri98, Eas96, Mor97, Rea81]. Planisferologio [Tur90a, Tur90a]. Planispheric [Gib78]. Planned [Mau13]. Planorum [Gin89b]. Plaskett [Jar99, Smi18a]. Plata [Gra12b, dA17]. Plate [ASCE14, Haz02, Zru21, AKJ^{+}23]. Plateau [BGP13]. Plates [Gin83a, SS16, Sta04]. Platform [Ave05]. Platinum [Cha72]. Plato [Llo11, Bra23, GT00, Kno90b, Llo11]. Platonic [Dra73]. Played [Bre18]. Playwright [Fal14, Wło15]. Pleasing [Het76a]. Pléthon [Cha00, Mer98a, Tih98]. Pliny [McC91, GB95]. Plumbline [CR11]. Plumed [FA97]. Plumian [Moo94]. Plurality [Far04, Fei90, O’B88, Bro83]. Plus [Cha09]. Pluto [Yeo07, Lan82, Rea81, Lan82, Rea81, Rea84, Rea92]. Plymouth [Bre18]. Pocket [Gin91b, McC82, Par90, Wil04]. Pocket-Sized [Wil04]. Poésie [Bla01]. Poet [Dan19a]. Poetic [Dan22, Gee18, Wes93]. Poetics [Wes93]. Poetry [Bla01]. Point [Eva93a, Moz13a, Not14a, PC89, RC18, SCV20]. Pointer [ASCE14, Tur15]. Points [SBF07b, SBF07a]. Poklady [Gin83a]. Poland [Ave95, Bar19, Prz79, Prz79, Rav72]. Polar [Dew82]. Polaris [Far96]. Polarlichts [Sis87]. Pole [TQ17, Wil79]. Polemic [Bal85]. Policies [Tat99]. Policy [Hir92]. Polish [God07, God16, Rav72]. Politecnica [Cen17]. Política [Hei14a, Gli88, Gra12a]. Political [How05, Rüp15, Wes96, CL22]. Politics [Coy07, For74a, Nic90, Rap15a, Sta13, CJ22, Gim12, Wil14a, Waf91]. Polsce [Prz79]. polskiej [Rav72]. Polyhedral [And17]. Polymath [Nor80, Lev25]. Polymathic [Ver17]. Polyphonic [Dan19b, Dan19b]. Polyptych [Kre12].

Pomniejsze [Gin08]. Poniatowski [Dre15]. Ponticus [Gra14a]. Popes [Hos92a]. Popper [Kra12]. populaires [Bir74]. Popular [Edd74, Gin81a, Gre99, Van92]. Popularising [BS73]. Popularization [Dra80]. Popularized [Wil05b]. Popularizing [Nal16, Ram16, Sta15]. Populations [Fea00]. Portable [Arn11, Ges19, Mas21]. portée [Lüt00]. Portents [Cam11]. Portrait [Gin77c, Ome73]. Portraits [Gin81b, Gin81b]. Portrayed [Han12]. Portugal [CG10, Hos97c, HS99, HVT^{+}01, LN09]. Position [Bot76, Gap15]. positional [Wło23]. Positions [Ano13a, CG13, Coh86a, Gin00b, Hub83, Bra23]. Positivisme [Har85]. Possible [Car19, Dib76, Hos97c, TTT75, VSH93, Oss24, UC22]. Post [Gin83b, Hen08, McC04, OT13, Sul02]. Post-Carolingian [McC04]. Post-Copernican [OT13]. Post-Darwinian [Hen08]. Post-Galilean [Gin83b]. Post-War [Sul02]. Postage [LF15, Rap20]. Postclassic [Ald03]. Poszukiwanie [God07]. Potential [RMH91, RM92]. Potsdam [Eel83]. Potsdamer [Eel83]. Poulle [Nor82]. Poverty [PC89]. Power [Ano06a, Ave11, Ben76b, CJ20, DeV98, GH07, McC98b, Rüp15, Tho90a, War08, McC98b]. Pp [Ano15f, Apt14, Ave95, Ave18, Bak15, Bec15, Bec19, Bel14a, Bel14b, Ben20, Boh19, Bon19b, Bra15, Bre18, Bro90, Bur15, Car86, Cen14, Cen17, CWW17, Cha19, Cha15b, Cro15, Cro23a, Dan19b, Dev14b, Doe19, Dun20, Eva17, Fal23, Fin19, Gan15, Gee18, Ger17, GG14, Had84, Hal03, Hea15, Hei14a, Hei14b, Hen15b, Jar90, Joh19, Jus20, Kes15, Kle14, Kra22, Kru17, Lau14, LF15, Lep14, Lip15, Llo15, Löh19, Mal15, Mar14, McC19, Mes15, Mir17, Mit19, Mol14b, ML14, Mul83, Nau98, Oes15, Oos15, Qua17, Rad90, Rap15a, Ree20, Rem15, Rob86, Roc17, Rot17, Sch15, Sha14, Sil23, Smi96, Smi18a, Sul14, Swe17c, Too95, Vet19, War19, Wes97, Wil15, Wło15, Wło23, Zep19]. Prachatic [Nor02]. Practica [Gre10]. Practical [Ben92, Dew00, Gin91b, McC82, Dav08, Dun92, Gin91b]. Practically [Not19]. Prácticas [Iwa96]. Practice [Eva04, Joh10, KD98, McC03, Sha94b, Sta04, Voe01, You86, dJ12, Hol10b]. Practices [Ban13, Ben87b, CK07, MJT03, Pan94b, Pan95]. Practika [Gin03e]. Practique [Tay10a]. Practising [Ano18a]. practitioner [Hei22]. Præsepe [Sch23b]. Prague [Kre05, Tur70a, Mor14, Tur70a]. Praise [Dan11, Don71a]. praká’saka [Pin84]. Pratiques [Hos00b]. Praxis [Gin89b]. Präzisionspendeluhren [Bre13]. Pre [Ano73a, Ave18, BEG98, BEC^{+}99b, Car76, Don73c, ESBG96, ESBG97, EBB^{+}01, Gol72, Gol96b, Hiv81, Kru22, Li19, Pum11, Sal97, Spr92, SS98a, Thi88, WKH18]. Pre-Columbian [Ave18, Car76, Ano73a, Ave18, Hiv81, Kru22, Spr92]. Pre-Hispanic [ESBG96, ESBG97]. Pre-Islamic [BEC^{+}99b, EBB^{+}01, Gol72]. Pre-Jesuit [SS98a]. Pre-Keplerian [Don73c]. Pre-Modern [Li19, Sal97]. Pre-Roman [BEG98]. Pre-Telescopic [Gol96b, Pum11, Thi88, WKH18]. Preacher [Gue12]. Precession [Gol75, Gol94, Har79, Pin72, VWH97]. Precessional [Moe75]. Precision [Gin02b, Rug82, Rug83, SS95, SS91b, GJ22, Cha89]. precursors [GE24]. Predicted [Bro85, SS97b, Ste97]. Prediction [CJ22, Dra79, Far96, Gin70b, Pan94a, SF97, Ste08b]. Predictions [Mor01a, Ste98a, Rut21]. Preface [Ano89c, Ano90a, Ano91a, Ano92a, Ano93a, Ano94b, Ano95c, Ano96b, Ano97a, Blu13, Tho84]. Prehistoria [Iwa96]. Prehistorian [Atk75]. Prehistoric [Ash00, Cow71, GZB14, Heg78, HH82]. Prehistory [Gin01d, HD06, Bel01a].

Prelude [Nee91]. Preludes [McM05].

Premières [Yan86]. Premiers

[Cap12, Wil03]. Premodern [Roc17].

Preparation [Ost95]. Preparatory [Jar06]. Prepared [Gin06d]. Presence [Sha87a]. Present [AKJ^{+}23, Ano76b, Ave94, Eng02, Hos94, McC07, Wea90, Tat99, Ten00, Whi76a]. Presentation [BSL86]. Presented [Gen96, Gin88a, Hos00c, Jas96b]. Preserved [Nor92a]. President [Ano70b]. Presocratic [Tau96]. Presocratics [Gre14]. Press [Ave18, Bec19, Bel14b, Bon19b, Bra15, Cha15b, Dan19b, Dev14b, Doe19, Gan15, Ger17, GG14, Had84, Hen15b, Kes15, Kra22, Lau14, Lep14, Mal15, McC19, Mes15, ML14, Mul83, Nau98, Rad90, Rap15a, Ree20, Rob86, Rot17, Sha14, Smi96, Smi18a, Sul14, Swe17c, Too95, Wil15, Wło15, Gin81a]. Press/Center [Had84]. Pressure [Coh99]. Pretended [DN98]. Previously [Gin92a]. pribory [Rya71]. Prima [MG86, Swe86a, Gra80, Tre04, MG86, God16]. Primal [Hat08]. Primarily [Mer88]. Prime [Waf01]. Primitifs [Cha09]. Prince [Hev94]. Princely [War08]. Princeton [Lau14, Mes15, Ree20, DeV02b]. Principal [Ric07]. principales [Pou84]. Principia [Smi96, Nau98, Den00, Gin95, Nau98, Nau17, Smi96, Swe01, Whi70a, Whi70b, Wil01, Gin95]. Principle [Blu00]. Principles [Ken73, McC93, Nau98, Swe01, Whi70b]. Print [Lip00, Rap11b, Kre02]. Printed [Gin90d, GV03, Pan12]. Printer [Tho71b]. Printing [Gin20, Gin75b, Pan94b, Pan95, Cro23a, Kre13]. Prinzipien [Wil01]. Prior [DeV00b, WR08, Sch23a]. Priority [Bec00, Gin95, Jar90, Ser13]. Prismi [Cen17]. prístroje [Sch14a]. prístroju [Sch14a]. Private [Cha88, Hos04b]. Prize [Gin88d, Kra17]. Probabilities [Fre75]. Probable [Ash80, Loc85]. Problem [DeV75, Nau98, Rug83, Van74a, Wes09a, Wil08a, Wil09, Kut23, Swe86a]. problematic [Wes01]. Problems [Gol82, Gol84, Jas96b, Kin73]. Procedure [Ste13]. Proceedings [Ash86, Ave95, Hos96, How88, Liv13, Mel89, Sal97, Tri10, Wal13, Whi76b, Wil15, Rot01]. Process [De 06]. Processo [She86]. Proconsularis [BEG98, BBMG03]. Procyon [Kol01b]. Prodromus [Dra73]. Production [BG03, Sha12]. Prof [Gin99b]. Professional [Eva04]. Professor [Gin77c]. Professorship [Coo84, Moo94]. Prognostica [Voe97]. Prognostication [Kre13, LG12, RL18]. Program [Ano15f, Bro10, Doe08, Fat98, Fat00, Mer98b, Ger14]. Programme [Cul07, Pau81, Bec22]. Programs [Gin91b, LS05]. Project [BM15, MR93, RMH91, RM92]. Projections [Jar09a]. Projects [Bro10]. Prolegomena [Jon99]. Prominent [RM92]. Promise [O’D06]. pronostication [Hos89b]. Proof [Gin03h, Hei12, Sza18, Nal22]. Proofs [GB09]. Propaganda [Lew07]. Proper [DE00, Duk02, Ser90, VvdS19, Ler80]. Prophecy [Kre13, Mal97, Tho90a]. Prophetic [Kre13]. Proportion [Ano84b, Woo80]. Proposal [For71d, Swe86c]. Proposals [Mac20]. Proposed [SG05]. propre [Ler80]. Propriety [MJT03]. Prosdocimo [Cha07]. Protestantism [Gin01f]. Protomathesis [Pan10]. prototype [CR22]. Proust [Roc17]. Proved [Déb05]. Provenance [GB86, Mor14, SMH18]. Provencal [Sau00]. Provence [Ver87, Ver87]. Providence [Hos77]. Provijest [Mul84]. Provoked [Gin11c]. Prozess [Mos86]. Prusse [Tur96]. Ps. [Bou23]. Ps.-Ptolemy [Bou23]. pseudo [GH24, She15]. pseudo-Hipparchus [GH24]. Pseudo-Ptolemy [She15]. Pseudoscience [Sha14]. PSI [TF16]. Psychology [Dic90]. Ptolemaeum [Gol73b]. Ptolemaic [Bla97, Löh19, Car23, Eva87b, Eva87c, Eva92, EM08, Gin00b, Mor16, Moz16, Sal89, Sal94a, Sal94b, Too77].

Ptolemäus [Eva92, Too77]. Ptolémée [Eva88, Eva01, Ler11b, Nor82, Pin80]. Ptolemy [Eva92, Eva93b, Gin90e, Gol84, Kre93, Llo94, Nor94, Wil84, Bar99, Ber22, Ber19, Bou23, CS21, Car19, Cau25, CT93, DE00, Dob02, Eva93b, Gol73b, Gol82, Gol97, HS81, Jon17, Kno91, Kre93, Lás19, Ler11b, Llo94, Loc84, Mac98, Mer11, Moz14, Mur95, Nev96, OT13, Pin03, SC88, Sch13, SL13, She90, She15, Swe89, Swe92a, Swe04a, Sza18, Thu03, Van94a, Van98a, Whi75, Wil84, Wło90, Llo94, Mer11, Moe80, Pin03, Tih17]. Ptolomäus [Mos09]. Public [Cha88, Cou18, Gol87a, Han16, Hos04b, Nor07, Ste90a, Waf86b]. Publicacions [Hei14a]. Publication [Ave95, Gin06d, Van06]. Publications [Ano16d, Gin87a, Hos84b, Cha75]. Publikationen [For70a]. Publishers [Zep19]. Publishing [Bel14a, Llo15, Vet19, Cro23a]. Pueblos [Zei86]. Pulchro [Not14a]. Pulkovo [Kri97, Kri97]. Pulkovo/St. [Kri97]. Pulpit [Gue12]. Pulsars [Mea70b]. Pulsating [Mea70b]. Punto [Gli88]. Purbachii [Mal13]. Purchase [Bel20, Hos08e]. Purse [Gen90a]. Pursuit [Bon19b]. Putting [Van14b]. Puzzle [Cer96]. Pyramid [Ave05, Wal07]. Pyramidentexten [Loc99]. Pyramids [Bel01b, Haa84, Wal06, NR14]. Pyrénées [BG12, Iun02]. Pythagoras [Was77]. Pythagorean [Tau96, Dib76]. Pythagoreanism [Was77].

Qibla [Rag01]. Qing [CL22]. Quadrans [Not20, Kin02b, Kno97, Pou84, Sch25]. Quadrant [Ben92, Tur02]. Quadrants [Ree11]. quadrivialia [Nor86]. Quadrivium [Ste96]. Quaestiones [Pin92]. Quaker [Dav08]. Quantum [DK83a, DK83b, DeV00b, DeV94, KD90]. Quarter [McC89]. quasi [dA17]. Quattuor [She74]. Queen [Eis85]. Quellen [Jor94, Sha94a, Pin92, Voe02]. Quest [Bon16a, CS02b, Hos71b, Hos79a, Nau01, Ped85, Ruf10, Sch99, Sch07b]. Question [LG12]. questionable [MHS23]. Questione [McM89]. Questions [Jon99, Pin92, Wes84]. qui [Gin87c]. Quincy [Dic91a]. Quinque [Azz11]. Quintants [Ree11]. Qurra [Rag92].

R [Ano15f, Ano81a, Gin90e, Smi18a, Ken78]. R. [Gin90e]. Rūh [NRKN16]. Rêve [Tur07]. Rôle [Har77]. Race [Dun13, Sta07, Vet19, Cre06, Van95a]. Radar [Doe00, Tri13]. radices [GD93, Gin97a]. Radio [Hew74, Hir85, McC13b, Mit86, Mit19, Nee91, Sil23, SL83, Str19, Sul77, Sul99, Sul14, Tri13, Mun13]. Radioactivity [Kra07c]. Ragep [Löh19]. Rahmani [Mer06]. Raihan [Ken78]. Rain [Kno03, Spr93b, Spr93c]. Raised [Gol82, Gol84]. Raison [Tur07]. Raisonné [Gin04a]. Ram [Mol03]. Rambaut [Ben14a]. Ramble [Gin05e]. Ramponi [Sie05]. Rampur [Gin05b]. Ramsden [Ben09, Meu10, Ben09, Chi09, CR11]. Ramus [Swe86a]. Ranganátha [Pin84]. ranging [TPYR21]. Rantzau [Gin07a]. Rapa [GG14]. Raphael [Rem11]. Rapport [Wil95]. Rare [Con70, Gin01a]. RAS [Ron89]. RASC [UL95]. Rather [Gra12d]. Ratio [Car21, Pou09]. Rational [KD90]. Rationales [Kel10]. Raumordnung [Spr92]. Rawh [NRKN16]. Ray [Mul83, Smi85]. Raymond [Hus13, Hus13]. Raza [Gin05b]. rd.springer.com [Sil23]. rd.springer.com/book/10.1007/978 [Sil23]. rd.springer.com/book/10.1007/978-3-319-91843-3 [Sil23]. Re [LL87, MHS17, Pal08, She86, War93]. Re-Appraisal [Pal08]. Re-Edited [She86]. Re-Established [War93]. Re-Evaluation [LL87]. Re-Visited [MHS17]. Reachable [Ave09]. Reaction [Fei86]. Reactions [Fei86]. Read [CMM^{+}15, Dan04b, Hus12, Cro23a, Gra21a]. Reader [Bel09b, Smi96, App02]. Reading [Ano22b, Blu21, Cam02, Den00, Hus12, Lip00, Mar14, Pas03a]. Readings [Tri97, Jas96b]. Ready [Bed80]. Ready-to-Assemble [Bed80]. Reality [Ban13, Rug95, Smi92]. Really [Hos76b]. Realm [Edm84]. Reappraisal [PH96, Ser13, SG05]. Reappraisals [Jar91a]. Reappraising [Tau96]. Reason [Nor78, Tur07]. Reasoning [Don82b]. Reasons [Gol11, Kin87]. Reassessed [Gre14, Ste01]. Reassessing [Bro15]. Reassessment [Dav19, EB04, HC99, Hos79b, Rug82, Rug83, SS86, Zei85a]. Reattribution [GW81]. Received [Ano80b, Ano81b, Ano82b, Ano82c, Ano83a]. Receives [Ano17c]. Reception [Hal03, Jar73, Mos03, Not14b, OT13, Pal12, Siv05, VG01, Wło11]. Rechenmaschine [Nor80]. Rechnen [Swe90]. Reckoning [AMP95, AMP96, Duk12, Qua17, Cul17, Duk12]. Recognition [Ser88]. Recollections [Kra10, DeV85a]. Reconnaissance [Hos85d, Rug94]. Reconsideration [Duk02, Shi00, TT78]. Reconstruction [CB05, Jon76, Mal78, Sta06, VMV18]. Record [ABB03, IK80]. Recorded [OP89, SS91a, Ste98a, SF93, SG09, Yor01, MD21]. Records [Ano03a, Bri05, Edd82, Gau12, Mar87, PG79, Por74, Shi11, SY92, Ste08a, Ste10b, SMH18, SMH20, UCI16, WS16, MSH22b, NN24, Edd80, Pan09a, Rad90]. Recovery [Pin76]. Recreating [Sta04]. Rectilinearity [Mil08]. Recueil [Gra90a]. recurrent [Sch23a]. Red [Cer95, Cer96, De 06, Edd72, Hoc92, Liv13, Sul12]. Red-Shifts [Liv13]. Redefining [Doe90]. rediscovered [Mas21]. Rediscovering [Tau13]. Redshift [Hol10a]. Reducing [Wło23]. reedited [EP22]. reference [UC22]. Referenced [Hol95]. References [Gin81c]. Refined [Whi75]. Refining [Azz11]. Reflecting [Cha72, Jar99, Ree11, Sim92, Sim09, Tur79b]. Reflections [God06, Hal88, Hir85, ML14]. Reflector [Hos03b, Mau13]. Reflectors [Hos08e]. Reform [Ash86, HS16, McC19, Pan09a, Sal89, War13]. Reformation [Gin11b]. Réforme [Kus10]. Reforms [Cul93]. Refraction [Ext92, Lin07, McC17]. Refractor [CL08, Lan83, Kär24]. Refractors [Dew82]. Regard [Bre18]. Regime [Ben14a]. Regiomontani [Gin78b]. Regiomontanus [Swe90, Dob92, Gin78b, Rag05, Sha82, Sha92, Sha98, Sha07b, Sha12, SS98b, Swe99b, Swe17a, Dob92, Sha82]. Regional [Rot03, Sch14a]. Regionálního [Sch14a]. Régionaux [Hos00b]. Regions [GB10]. Regis [Cár88]. Register [Ano71d, Joh19]. Regression [Tho71a]. Rehabilitating [Can09]. Reich [Oes15]. Reihe [Her76]. Reimers [Gau17]. Reinhold [OT13, Kre93]. Reinterpreting [Sch90]. Re’is [Kin79]. rekiho [Nak71]. Rekonstruktionsversuche [Row07]. Related [Gin92c, Hub91, KMD^{+}18, Llo90, Sid10]. Relating [Dal80]. Relation [SW18, Smi79]. Relations [Han16]. Relationship [Chi09, Wil07]. Relative [Eva93a]. Relativitätstheorie [Hen92]. Relativity [Gin82a, Gle07, Kra99, Sch11, Sta07, Cre06, G^{+}99, Eis06, Gle07, Hen92, Hen97, Mel09, Mil81, Nor00a, Sta07]. Relevance [Coh91]. Relevant [Ano92h, Ano93i, Ano94h, Ano95i, Ano96e, Ano97i, Ano98h]. Reliable [Ste08a]. Relief [Pap91]. Religion [Bow10, Cog08, Dav06, Dav08, Gin11c, Lin93, McM05, Osl06, Bro15, Gim12, Sta13, Bow10, Wes94]. Religions [Fek13]. religiosa [Gra12a, Hei14a]. religiöse [Sev00]. Religious [Eis96, Gre09, How05, Kin95b, McM05, Mol93, Ram16, Dav06]. Remaining [HHK90]. Remains [Far93]. Remarks [Aab72, New77]. Remembered [DeV03, LC10, Rag83]. Remembering [Lay02, Tri10, Wen99]. Reminiscences [Ber70, Phi91]. Reminisces [Sul77]. Remnants [Tri03]. Remote [ZJJ^{+}99]. Renaissance [Azz06, Bar13b, Blu00, Gin13b, Hor82, Lat00, Loh75, Oes10, Rap11b, Rem11, Rut13, Wes90, Ban10, Dup14, Gin97d, Gra76, Hor82, Kre99, Kun87, Lüt92, Mar12a, Sha07a, Swe72, Swe90, VB87, dA74, dV91, vH87, Cam02, Cau08, Don98, Dup14, Hal93, Nor04, Par86, SS08, vB17]. Renässansen [Sha07a]. Renato [LF15]. Renovación [Hei14a, Gra12a]. Replacement [Mer92]. Replicability [Hol10b]. Report [Ano94a, Ano96a, Car00, Siv05, UC22]. Reports [Jon76, NRKN16]. Representation [Fle92, Mea02, Hen76]. Representations [Pan94b, Pan95]. Representing [Eva06, fH21]. Reprint [Gin78c]. Reprinted [Hei99, Ken85, Rem08]. Reprints [Mit86]. Reproduction [Pan95]. Republic [Hal03, Siv05, Gre16]. Republik [For70a]. Reputation [Cas16]. Reputed [Ome73]. Rescriptus [GH24]. Research [Ano73f, AR86, Bla98b, Bru96, Cha88, Coo75, Dew00, Gen90a, Mac15, Mea02, Mir17, Pau81, Plo16b, RS18, SB72a, SB72b, TQ17, TW25, Wea90, Cun16, Wło25, Sis85, Wea90]. Researched [Van96]. Researches [FV17, For76, Ano15f]. Residence [Hos12c]. Resolutae [Cha98]. Resolute [DeV89]. Resolution [Hos90]. Resonances [Roc19]. resources [CJ22]. Response [Ano09a]. Responses [Swe17b]. Rest [Sha87a]. Restored [CJ20, Mol14b]. Restoring [Cas16, Mol14b]. Resultados [Ave87]. Resurrection [Bla98b]. Retirement [Ano07h]. rétrogradation [Par85]. Retrogradations [Swe84]. Retrograde [Mac98]. Retrying [Gin05d]. Return [Bro85, Waf86b, Wil93]. Rev [Don71b]. revealed [GWZ22]. Revelation [Mar03]. Revenge [Hos06d]. Reverent [Gin01b].

Revew [Fal23]. Review [Ack05, Ack10, Ait82, Ait86, Ald97, Ald09, Ald11, All71, And90, And00, Ano73e, Ano75c, Ano76b, Ano76c, Ano77a, Ano84b, Ano84a, Ano97b, Ano03a, Ano06a, Ans05, App02, App05, Apt89, Apt99, Apt14, Arm70, Arm72, Arm73, Arm74, Arn95, Ash72, Ash79a, Ash79b, Ash00, Ash82, Ash86, Ash87, Atw79, Aub12, Ave87, Ave92, Ave94, Ave95, Ave97, Ave09, Ave11, Axw17, Azz06, Bak15, Ban13, Bar05, Bar13b, Bat87, Bat90, Bec99a, Bec10, Bec15, Bec19, Bed80, Bed10, Bee71, Bel77, Bel07, Bel09a, Bel14a, Bel01a, Bel06, Bel09b, Bel11, Bel14b, Bel19, Ben76a, Ben82a, Ben84, Ben87a, Ben87b, Ben91a, Ben91b, Ben93a, Ben99, Ben09, Ben13, Ben20, Ber70, BS73, Ber74, Ber81, Ber82, Ber89, Ber94, Ber97a, Ber01, Ber92, Ber97b]. Review [Ber12, Ber14, Ber19, Bia04, Big10, Bla98a, Bla98b, Bla91b, Bla97, Bla01, Bla79, Blu00, Boh19, Bok73, Bol05, Bol10, Bol20, Bon76, Bon09a, Bon11, Bon12, Bon19b, Bon19a, Bow83, Bow10, Bra90, Bra15, Bra05, Bre00, Bre13, Bre11, Bri98, Bri01b, Bri05, Bro91a, Bro83, Bro96, Bro10, Bro90, Brü87, Bru78a, Bru82, Bru96, Bru09, Bry81, Ano81a, Bur82, Bur84, Bur15, Bur20, Bur80, Bur10, Cah19, Cam02, Cam11, Can12, Can09, CL05, CL08, Cap12, Cár88, Car76, Car86, Car97, Car99, Cau08, Cau17, Cen14, Cen17, Cer97, Cer99, Cha00, Cha09, Cha19, Cha85, Cha89, CT12, Cha15b, Cha75, CM10, Chr98, Cog08, Coh79, Coh80, Coh86a, Coh91, Coh95, Coh97, Coh99, CS02b, CJ05, Coo75, Coo97, Coo00, Cor98, Cor03, Cow71, Coy07, Cro15]. Review [Cro00, Cro08, Cro12, Cro23a, Dal01, Dan04b, Dan11, Dan12a, Dan12b, Dan19a, Dan19b, Dar91, Dav06, Dav08, De 13, DeV85b, DeV85a, DeV89, DeV96, DeV98, DeV00a, DeV02a, DeV03, DeV05, DeV07, Dea94, Déb87, Déb91, Déb05, Déb13, Dek12, Den00, Dev14a, Dev14b, Dew70, Dew79, Dew87, Dew00, Dic82, Dic90, Dic95, Dic97a, Dic99, Dic00a, Dic00b, Dic06, Dic09, Dic13b, Dij09, Din73, Dob77, Dob79, Dob83b, Dob83a, Dob86, Dob92, Dob96, Dob99a, Dob00, Doe94, Doe00, Doe01, Doe05, Doe08, Doe19, Dog87, Don71b, Don71a, Don73a, Don73b, Don79, Don82b, Don82a, Don86, Don88a, Don92, Don98, Don00, Don02, Don09, Dra80, Dra84a, Dra90, Dru84, Duk10, Duk11, Duk12, Dun92, Dun13, Dun20, Dup08, Dup14, Eag08a, Eag08b, Eas89, Eas96, Eas10]. Review [Edd71a, Edd74, Edd80, Edd82, Edd86, Edm84, Eel74, Eel83, Eis85, Eis88a, Eis88b, Eis96, Eis03, Ell85, Ell08, Eng02, Eva82, Eva86, Eva87a, Eva88, Eva89, Eva92, Eva93a, Eva93b, Eva00, Eva01, Eva02, Eva05, Eva06, Eva07, Eva08, Eva09a, Eva09b, Eva11, Eva12, Eva13, Eva17, Fan05, Fan09, Fat00, Fei84, Fei86, Fei90, Fei95, Fei03, Fek13, Fel02, Fin90, Fin99, Fin10, Fin19, Fle73, For70a, For73, For81, For82, For83, Gal83, Gan15, Gau10, Gau17, Gen87, Gen90b, Gen92b, Gen96, Ger17, Ges19, Gib78, Gib82, Gie20, Gin70a, Gin72, Gin75b, Gin75a, Gin76, Gin77b, Gin77d, Gin77c, Gin78a, Gin78c, Gin78b, Gin80a, Gin80b, Gin80c, Gin81a, Gin81b, Gin81c, Gin82a, Gin82b, Gin82c, Gin83a, Gin83b, Gin84b, Gin84a, Gin85a]. Review [Gin85c, Gin85b, Gin85d, Gin85e, Gin86a, Gin86b, Gin87b, Gin87a, Gin87c, Gin88a, Gin88b, Gin88c, Gin89a, Gin89b, Gin89c, Gin90c, Gin90d, Gin90a, Gin90b, Gin90e, Gin90f, Gin91a, Gin91b, Gin91d, Gin91c, Gin92b, Gin92a, Gin92c, Gin93, Gin94, Gin95, Gin96b, Gin96c, Gin96a, Gin96d, Gin97a, Gin97b, Gin97c, Gin97d, Gin98, Gin99b, Gin99a, Gin99c, Gin99d, Gin99e, Gin00a, Gin00b, Gin00c, Gin00d, Gin00e, Gin01c, Gin01g, Gin01d, Gin01e, Gin01b, Gin01a, Gin01f, Gin02a, Gin02b, Gin02c, GHHB02, Gin03a, Gin03b, Gin03c, Gin03d, Gin03e, Gin03f, Gin03g, GM03, Gin04b, Gin04a, Gin04c, Gin05b, Gin05a, Gin05c, Gin05d, Gin06a, Gin06b, Gin06c, Gin07a, Gin07b, Gin08, Gin10a, Gin10c, Gin10b, Gin11a, Gin11b]. Review [Gin11c, Gin12b, Gin12c, Gin12a, Gin13b, Gin14, Gin15, Gle04, Gle07, Gli88, God07, God13, God19, Goe09, Gol73a, Gol87a, Gol72, Gol73b, Gol76, Gol87b, Gol95, Gol00, Gol11, Gol12, Gol03b, GG14, Goo09, Gra80, Gra90a, Gra14a, Gra07a, Gra08a, Gra12b, Gra12c, Gra12d, Gra13a, Gra76, GB86, Gra90b, Gra00, Gra01, Gra05c, Gre99, Gre00, Gre02, Gre04, Gre09, Gre14, Gua09, Gub88, Gur06, Had84, Hal88, Hal93, Hal03, Ham06, Ham09, Ham11, HS81, Han12, Han09a, Han09b, Har73, Har94, Har92, Har72, Har75, Har78, Har84, Har85, Has11, Has19, Hat83, Hat88, Hat04, Hat08, Haw74, Hay81, Haz02, Hea94, Hea97, Hea99, Hea15, Heg78, Heg82, Heg85, Hei00, Hei14a, Hei99, Hei08, Hei12, Hei14b, Hen75, Hen76, Hen08, Hen09]. Review [Hen13, Hen15b, Her72, Her76, Her80, Her90, Her94, Het76a, Het78, Het94, Hev94, Hew74, Hig12b, Hig12a, Hig17, Hil77, Hin04, Hir85, Hir92, Hir09, Hiv81, Hol03, Hol10b, Hol77, Hol95, Hor74, Hor82, Hos71b, Hos72, Hos73, Hos75, Hos76a, Hos78a, Hos81, Hos82b, Hos83a, Hos83b, Hos84b, Hos86a, Hos88, Hos89b, Hos92a, Hos94, Hos95a, Hos95b, Hos96, Hos97a, Hos97b, Hos00a, Hos00b, Hos00c, Hos01a, Hos01b, Hos02a, Hos04b, Hos04c, Hos04d, Hos04f, Hos04e, Hos05a, Hos06a, Hos06b, Hos08b, Hos08a, Hos08c, Hos11a, Hos11b, Hos12a, Hos13, How02, How05, How08, How80, How88, Hoy84, Hub83, Hub91, Huf06a, Hug10, Hus13, Hut12, Iwa96, Jar90, Jar91a, Jar00, Jar09a, Jar73, Jas96a, Jas96b, Jhi74, Joh19, Jon91a, JT92, Jon97, Jon99, Jon06b, Jon10, Jor94]. Review [Jus20, Kai07, Kak95, Kaz73, Kel10, Kel78, Kel83, Kel99, Kel74, Ken73, Ken78, Ken82, Ken85, Ken87, Ken94, Kes10, Kes15, Kid87, KH73, Kin78, Kin79, Kin81, Kin83, Kin87, Kin95a, Kle14, Kni70, Kno90a, Kra07a, Kra99, Kra07b, Kra10, Kra11, Kra13, Kra22, Kre80a, Kre81a, Kre82, Kre86, Kre89, Kre93, Kre99, Kre02, Kre05, Kre07b, Kre08, Kre09, Kre13, Kre19, Kre20, Kri88, Kri94, Kri97, KI10, Kru83, Kru97, Kru08, Kru17, Kul80, Kul81, Kul75, Kul87, Kun85, Kun88, Kun90, Kun91, Kus10, Kwa12, Lad10, Lan89, Lan82, Lan84, Lan87, Lan88, LN09, Lás19, Lat93, Lat00, Lat12, Lat19, LF15, LC10, Lep14, Ler06, Ler11b, LG12, Ler17, Lia19, Lil99, Lin79, Lin93, Lin07, Lip15, Lip19, Lip00, Liv13, Liz17, Llo80]. Review [Llo89a, Llo90, Llo94, Llo09, Llo11, Llo15, Loc96, Loc99, Loc00, Loh75, Löh19, Lon06, Lov87, Low09, Lut91, Lüt92, Lüt00, Mac76, Mac93, Maf05, Mal90, Mal97, Mal13, Mal15, Mar75, Mar87, Mar97, Mar99, Mar03, Mar12a, Mar14, Mar12b, Mar89, Max12, May11, McC82, McC83a, McC85, McC91, McC98b, McC99, McC04, McC07, McC11, McC13a, McC19, McC13b, McC13c, McC08, McK78, McK87, McM01, McM74, McM89, McM03, McM05, Mea70b, Mea71, Mea81, Mea85, Mea86, Mea02, Mel89, Mer77, Mer88, Mer92, Mer96, Mer03, Mer06, Mer09, Mer10, Mer11, Mes11, Mes15, Mil80, Mil04, Min00, Mir17, Mit86, Mit13, Mit19, Moe77, Moe80, MG86, Mol14b, Mol70, Mol76, Mol93, ML07, ML14, Mor97, Mor01b, Mor01a, Mor08, Mos03]. Review [Mos07, Mos09, Mos10, Mos86, Mul82, Mul84, Mul80, Mul83, Mul81, Mül91, Nak71, Nau98, Nau01, Nee71, Nee74, Nee08, Neu81, Nor76, Nor78, Nor80, Nor82, Nor83, Nor85, Nor86, Nor88, Nor90, Nor91a, Nor92b, Nor92a, Nor94, Nor96, Nor97, Nor99, Nor00a, Nor00b, Nor02, Nor04, Nor06, Nor07, Nor08b, Nor08a, Not19, O’B88, O’D06, Obr19, Oes10, Oes15, Ogj11, Olo04, Ols93, Ome79, Omo11a, Omo13, Omo19, Oos15, Orc84, Orc88, Orc90, Osl06, Oss12, Ost80, Ost92a, Ost97a, Ost04, Ost05b, Ost05a, Pal12, Pal77, Pal84, Pan09a, Pan09b, Pan15, Par86, Par90, Par92, Pas03a, Pas03b, PP07, Pas14, Pau78, Pau82, Pau88, Pau95, PG72, PG79, Ped73, Ped79, Ped87, Péo08, Per00, Pet80, Phi91, Pin71a, Pin73a, Pin74, Pin80]. Review [Pin81, Pin84, Pin88, Pin90, Pin92, Pin03, Pla91, PdV85, Pou75, Pou79, Pou84, Pou85, Pou96, Pou09, dSP78, dP82, Prz79, Qua17, Rad90, Rag83, Rag85, Rag92, Rag01, Rag09, Ram10, Ram06, Rap11b, Rap13, Rap15a, Rap11c, Ras13, Rav72, Rav73, Rea81, Rea84, Rea92, Ree10, Ree11, Ree20, Rei92, Rem08, Rem11, Rem15, Ren89, Ric07, Rob86, Roc08, Roc17, Roc19, Rom06, Ron89, Ros79, Ros73, Ros74a, Ros81, Ros74b, Ros19, Rot80, Rot01, Rot03, Row07, Rud92, Rug95, Rug99, Rug04, Rug07, RC11, Rut03, Rut12, Rut13, Rya71, Rya09, Sal79, Sal90, Sal95, Sal97, Sal06, Sal09, Sam07, Sam13, San99, Sar78, SS08, Sav12, Saw80, Sch07a, Sch15, Sch82, Sch90, Sch91, Sch99, Sch05b, Sch00b, Sch03, Sch12b, Sch14a, Sch70]. Review [SK03, Sch11, Sch07b, Sel14, Ser88, Sev00, Sha07a, Sha78, Sha82, Sha85, Sha87a, Sha89, Sha94a, Sha94b, Sha99a, Sha99b, Sha08, Sha11, Sha17, Sha71b, Sha02, Sha14, Sha87b, Sha96, Sha70, Sha07c, Sha87c, She74, She79b, She79a, She86, She87, Shi09, Shi11, Shu82, Sid10, Sil23, Sis85, Sis87, Siv77, Siv89, Siv98, Siv05, Siv10, Smi81, Smi85, Smi89, Smi92, Smi93, Smi94a, Smi94b, Smi94c, Smi96, Smi98, Smi02, Smi07, Smi09b, Smi12, Smo09, Sol09, Spr92, Spr93a, Sta72, Sta95, Sta07, Sta13, Sta99, Sta00a, Sta10, Ste99, Ste01, Ste05, Ste06a, Ste12a, Ste13, Ste17, Ste77, Ste90b, Ste98b, Ste00a, Ste02b, Ste06b, Ste12b, Ste12c, Ste88, Ste96, Sti07, Str10, Str99, Str19, Str81, Sul77, Sul99, Sul02, Sul12, Sul14, Sut72, Swe73]. Review [Swe76, Swe86a, Swe88, Swe90, Swe92b, Swe98, Swe99a, Swe01, Swe04b, Swe13, Swe17c, Syl07, Tat99, Tat09, Tau94, Tau96, Tau09, Tau11, Tau13, Tay10b, Ten00, Tes10, Thi88, Tho71b, Tho71c, Tho84, Tho86, Tho87, Tho88, Tho90a, Tho90b, Tih07, Tih17, Tip88a, Too77, Too81, Too95, Too96, Tri97, Tri03, Tri13, Tri10, Tuc00, Tur70b, Tur70a, Tur71b, Tur71a, Tur72, Tur75, Tur79a, Tur84, Tur90a, Túr90b, Tur96, Tur03, Tur07, UL95, UC17, Vai12, Van95a, Van95b, Van99, Van02a, Van85, Van89, Van90, Van92, Van94b, Van96, Van98b, Van98c, Van02b, Ver13, Ver87, Ver09, Ver12, Ver17, Vet12, Vet19, Vic85, Vil90, Voe94, Voe97, Voe01, Voe02, Voe03, Voe07, Waf86a, Waf91, Waf01, Waf02, Waf04, Wag88, Wal77, Wal99, Wal13, War71, War79]. Review [War80, War93, War98, War08, War13, War19, Was72, Was77, Way96, Wea90, WW74, WW82, Wel84, Wel94, Wel05, Wen99, Wer10, Wer13, Wes77a, Wes84, Wes90, Wes94, Wes96, Wes77b, Wes93, Wes97, Wes09b, Whi73a, Whi76a, Whi02, Whi72, Whi73b, Whi76b, Whi92, Whi00, Whi71, Wil15, Wil86a, Wil97, Wil02, Wil07, Wil87a, Wil14b, Wil75, Wil82, Wil84, Wil86b, Wil87b, Wil88, Wil95, Wil00, Wil01, Wil03, Wil04, Wil05b, Wil05a, Wil11, Wis06, Wło10, Wło15, WW19, Wło23, Woo88, Woo87, Yan86, Yav99, Yeo88, Yeo99, Yeo00, Yeo02, Yeo07, Zbi97, Zep19, Zui19, Zuj11, dA74, dC01, dV91, vB10, vG05, vH86, vH87, vB17, vdW78, vdW90, Ave18, Bre18, CWW17, Rot17, Sch25, Smi18a, Gee18]. Reviews [Ano07h]. Revised [Duk11, PF83]. revision [Zie23]. Revisited [BBMG03, BG10, CG12, Gin90e, MSH20, Nau98, Ped83, PC89, Rag01, Rag16b, WS16, Not23a, Fan09]. Revival [McC04]. Révocation [Fel02]. Revolución [Hei14a, Gra12a]. Revolution [Ano06a, App05, Ber97a, Dic13b, Don98, Gin13b, GB09, Jar91a, Lüt92, Mar97, Max12, Mil80, Oos15, Par92, Ros74a, Van14b, Wes09b, TB13, TH79, Tim25, Kus10, App05, Eva82, Gin85c, Gin96d, Hos86a, Jar91a, Kra07a, Lüt92, McM74, Nor78, Pas03b, Ped73, She79b, Tho86]. Revolutionaries [DeV07]. Revolutionibus [Blu13, Eva00, Cha19, Gra08a, Bar19, Mos03, Swe88, vN12, BG03, Gin81d, GW81, Gin84d, Gra08a, Ler00, Mor14, Mos03, SL13, Str16a, Swe74b, Swe81, Wes77b, Gin89c]. Revolutionized [Dan12a]. Revolutions [Coh91, Dan04b, Het76a, Hos86a, Rag05, Wes77b, Coh91, Too81, Wes09b]. Revolver [LH05]. Rheita [Cov17]. Rhetici [MG86]. Rheticu [Swe86a]. Rheticus [Kra07a, Swe86a, Dan11, God16, Hoo84, Kra07a, Swe86a, VB21, Dan11, Gra80, MG86]. Rhetoric [Ban13, Smi94c, Voe94]. Ribera [Ome73]. Riccioli [Gra10, Gra12d, Mar23, Rap11a]. Riccò [Nal22]. Ricerca [Bel11]. Richard [Duk16, PP07, Pau88, Syl07, Ano95h, Gin97e, dSP78, Syl07, dSP78]. richer [Gat24b]. Riches [Not15a]. Riddle [Bee70, Tip88a]. Rider [Bel01a]. Rienk [Hal03]. Righini [Ano82d]. Right [Don71b, GJ22]. Rights [Dew87, Sch99]. Rinascimento [Ano06a]. Ring [TTF76, Fre77, SB84, TT73b, TT77]. Rings [Edd86]. Río [dA17]. Rise [Aub12, Bru78b, Bru78c, Cha02, Dav06, Eis03, Fan05, Gin98, Gin01f, Kol01b, Kol07, Kra07a, Ost84b, Ost84c, Sch87, Sch00a, Wes90]. Risings [Sta72, LS24a]. Ritchey [Hev94, Hos76c]. Ritual [Mer92]. Rituals [Vai12]. Riverview [DH73]. Riviera [Tur07]. Rivieras [Tur07]. Rivo [Not14a]. rivoluzione [Ano06a]. Rivoluzioni [Omo11a]. Road [Don73b]. Robert [Jar90, Mit19, Swe17c, Cro23b, Hal93]. Roberts [Mac10]. Robertson [Sil23]. Robin [Joh19]. Robinson [Ben14a, Hos90]. Rocha [FF20]. Rock [Had84, Mur98, You86, MEA91]. Rocknia [SK08]. Rocky [Doe94]. Roddick [Ave18]. Rodríguez [Boh19]. Rodríguez-Arribas [Boh19]. Roemer [Cha85, vH83, Cha85]. Roeslin [Gra05b, Gra11b]. Roger [Ano01j, Bre18, Eva17, Mer01]. Role [Bre18]. Roles [Fab19]. Rolle [Gol95]. ROM [Bur15, Eva00, Fei13, Hei14b]. Roman [Bar18, Cau08, Eva04, Gra05c, Saw80, AS97, BEG98, Bor98, CLL^{+}00, Jon21, Loc93, Loc95, MHS17, Ram07, RAGGB18, Saw80]. Romanesque [UAGGB21]. Romans [Cer99]. Romantic [Duf17]. Rome [Fan05, Fan05, Han09b, Hei14b, MSH19a, MSH22a]. Rømer [Wło23, Fab19]. Ronchi [Cen14]. Ronchitti [Don79]. Rond [Wil03]. Rong [Cul07]. Root [Dav15]. Roots [Rob86]. Rosen [Coh86b, Kno90a]. Röslin [Sha99b]. Ross [Ost07]. Rosse [Arm72, Cha15b, BH81, Ben14a, DH91, Hos90, Mol14a]. Rossii [Wer10]. Rotates [Déb05]. Rotating [Pal08]. Rotation [BS92, Ste77, Ste12c, SMH18, Ler80, Mor01a]. Rothman [Bon19b]. Rothmann [Kre07b, Sha99b, Sei17]. Roussanova [Oes15]. Route [Bol10, Jon04]. Routledge [Qua17]. Rowan [Hos84a]. Rowley [Che72]. Rows [Rug94, Rug96]. Royal [Apt99, Ash72, Coh99, Dav17, Fei03, Gin81b, Per00, Ron89, Rot80, Tur79a, UL95, Vai12, War78, Ald02, Ano90h, Ash72, For70b, For71b, For71c, For74b, Gin81b, HR92, Mac20, Mul20, Rei18, War78, Coh97, Hea99, NN24]. Rubin [Tho22]. Ruby [Tri13]. Rudolphinae [Tho71c]. Rudolphine [Tho71c, JLGL14, LGJLF14]. Rugged [Ber70, Gin90g]. Ruggero [Tuc00]. Ruggles [Rad90, Dun23, Mag23]. Rule [Arn11, Gin03e]. Rus [Rya09]. Rusi [McC08]. Russell [Doe01, DK83b, Doe01, KD90]. Russia [Bat87, Che72, FV17, Nic90, Rya71]. Russian [FV17, MBH16, McC08, Sho21, Wer10]. Russischen [Kri97]. russkoy [Rya71]. Rutherford [Ano81a]. Ryan [Boh19]. ryght [Hos89b].

S [Ave95, Dav08, Fin19, Gin97e, HB92, Jar90, Jar99, Nor86, Nor88, Oos15, Ost84b, Ost84c, Smi96, Smi18a, Smi22, Jon76]. S.E.

[SS06]. S.J [Gin96c, Row07, Siv05]. S16C [CS14]. S1C [CS14]. sa [Fel02]. Saby [Fal23]. Sachs [Gin91d, Too84, Gin91d]. Sacramento [Lie02]. Sacred [Bel14b, BMA90]. Sacrobosco [Cro23a, CMM^{+}15, Gin88e, Kno97, Ped85, Gin15]. Saff [Ben20]. Saga [Eva09b, Pau88, Tat09, Wło10]. Sagittarius [RC92]. Sagredo [Rap15a, Wil14a]. Saha [DeV85b, DeV94, DeV85b]. Sahara [BEBM02]. Saharan [BEC^{+}99b]. Sahl [VB01]. Sail [Pas03a]. Sailing [Mer06, Lep14]. Saint [Ash03, UAGGB21]. Sainte [Gin07b]. Salamanca [LS24b, Cha98, Hos92b, Kre02]. Sale [Hos15a, Spa04]. Sally [Löh19]. Salt [Kre11]. Salvation [vdW90]. Samarkand [Kul81]. Samarqand [ZM20]. Sammelband [Gin13b]. Sammlung [Gin76]. Samos [Zbi97]. Samplings [Sha07b]. Samsø [CKE11]. Samuel [Bre18, Sch12b]. Sanctuaries [RBH00]. Sanctuary [FJ11]. Sand [Gin86b]. Sand-Glasses [Gin86b]. saniya [Rag85]. Sanskrit [Ans05, Mer88, All71, Plo16a, Sar78, Sha96, Shu82, All71, Sar78, Sha96, Shu82]. Santa [Nau98]. Santalla [Nor99]. Santo [LN09]. Santoni [Lip15]. Sara [Dun20]. Sardinia [HAG93, PRA87]. Sardinian [Ano04b, GZB14, HZ97, ZHGM96, ZB04, Zed05]. Saros [Gin92c, Gin92c]. Satellite [Dra79, Kwa12, Lau14, Ver13, Wło10]. Satellites [Cas19, Gin70b, Ogj11, Wil95]. Saturn [CWW17, DN98, Dob99b, Ste02a, Ste10a, Ste19, Ush15, Van74a, Van74b]. Savante [Hat08]. Savilian [Coo84]. Saving [Gol97, Kre93]. Sawai [Gin96b]. Saxe [Nor85]. Saxon [MSH22b]. Saxony [CG19, Not14a]. sbírkách [Sch14a]. Scale [Dav12, DK83a, Fea00, Gol77]. Scandalous [McC19]. Scandinavian [Sis85]. Scenes [How08]. Scepticism [Bau86]. Schaeberle [Edd71b]. schaften [Ano06a]. Schechner [Dun20]. Scheiner [EZ16]. Scheme [Ste10a]. Schiaparelli [Can12]. Schickard [Nor80]. Schinnagel [Kre12]. Schmeidler [Wes97]. Schmidt [Eel83, Eel74, Eel83, Lep14, Ost94, Ost05b, Eel74, Ost05b]. Scholarly [Doe19, Ger17]. Scholars [Don98]. Scholarship [GM03, Hat04]. Scholastic [Fei86]. Schöner [Ano15f, Van10, Gin84d, Gin13b]. School [Gin04a, Kel10]. Schriften [Dij09, Gin99e, Ren89]. Schrifttums [Kin81]. Schroeter [For80b]. Schulen [Sha85]. Schultz [For71d]. Schumacher [For81, For81, For81]. Schwarzschild [Gin93]. Schwemin [Ano15f]. Science [Ald02, All71, Ano75c, Ano76b, Ano06a, Ano09a, Ash72, Atw79, Bec99a, Ben93b, BS73, Ber89, Bla91b, Blu00, Bow10, Bur82, CL05, Cer97, Cog08, Coh91, Cor98, Coy07, Cro15, Dan12b, Dan17, Dar91, Dav06, Dav08, DeV00a, DeV02a, Déb05, Doo18, Eis85, Eis01, Fal14, Fei84, GHS08, Gin75b, Gin77d, Gin82b, Gin87b, Gin87c, Gin90c, Gin98, Gin00a, Gin01d, Gin01f, GM03, Gin11b, Gin13b, Goe09, Gol03b, Gra07a, Gra12d, GB86, Har84, Har85, Has11, Het94, Hos00c, Hos04b, Hos06b, Jon15, Kin78, Kul81, LN09, Lat19, Lau18, Lin93, Loc96, Lon06, Lov77, Lov87, Lüt92, Mar97, Mar14, Mel89, Mol93, Mos10, Nor88, Not14b, Olo04, Osl06, Oss12, Par86, Par92, Pas14, PG72, Pou75, Rav73, Ree22, Roc17, Rot03, RC11]. Science [Sal97, Sal06, SS08, Sch01c, Sha85, Sha94b, Smi97, Smi02, Spe13, Sta13, Ste00a, Sul02, Sul14, Tho90b, Tur71b, Van95a, Vic85, Voe94, Waf91, Waf04, Wal13, Was77, Wer13, Wes94, Whi92, Wło15, Bro15, Gim12, Hai21, Mun13, Hun18, Rob86, Roc17, Wer13, App02, Big10, Cor98, Dan12b, De 13, Déb05, Din73, Don73b, Gin75b, Gin90c, Gin96d, Gin99a, Gin01f, Gin11b, Gin11c, Gol03b, Gra01, Gre14, Han12, Heg78, Hei99, How02, Kre05, Lov87, Mar03, Min00, ML07, Nee08, Osl06, Pou75, Rag01, Rav73, Rot03, Shi09, Sul99, Van95b, Wel84, Yeo99]. Sciences [All71, Ben75, Bri01b, Dal01, Dar91, Dea94, Gin88a, Gin99b, Gra76, Jar14, Jon06b, Kra17, Kun85, Mul84, Nor08b, Oss18, Plo16a, Roc16, Sar78, Sha96, Shu82, Tur79a, Van95b, Waf04, Bok73, Oes15, Gin92b, Kun85]. Scientific [Ack10, Ano76c, Ano77j, App05, Ben87a, Ben91b, Ben09, Bla98b, Bon76, Bre11, Cha90, Coo00, Cro23a, Cul93, Dob79, Edd86, Eis96, Eva82, Gal83, Gen87, Gen96, Gin87a, Gin90b, Gin91d, Gin91c, Gin99a, Gin04a, Gle04, Gol11, GB86, Het94, Hos82b, Hos86a, Jar91a, Kre09, Low09, Lüt92, Mar97, McM05, Mil04, Mol93, Nor78, Omo17, Oos15, Par86, Pas03b, dP82, Rud92, Rya71, Sch14a, SS16, Sch70, She79b, Smi94c, Smi97, Ste05, Tau94, Tho86, Van95a, Van99, Ver12, Vil90, WW82, Wes09b, Gre23, TB13, Bon76, Pou85, WW74]. scientifica [Ano06a, Gua09]. Scientifique [Hos11a]. Scientist [Eas89, Fei84, Gal83, Gin02d, Smi94c, Aub25, BS73, Loh75, Lut91, Sti07]. Scientists [Gin85a, Gin87a, Het94, Hos04d, Smi12, Hos04d]. Scienza [GM03]. Scienziato [Lat12, Van96]. Scimitars [Hug10]. Scope [Hos83a]. Scot [Rut12]. Scotland [HC99, Sco90, TT84, Gin91c]. Scott [Tri13]. Scottish [Ell85, Gin91c, Rus74a, Rus74b]. Scottus [Eas01]. Scotus [Rut12]. Scrapbook [Gol87b]. Scrape [Cas19]. Scripture [Cou18, Gin07b, Swe86a, Hoo84]. Scrolls [Ste17]. Scrutinizing [Nor04]. Scrutiny [Tih07]. Sea [For71d, Ste17]. SEAC [Ano94c]. Seaman [Gen90b]. Search [Apt14, Ben76a, Cha88, Cro12, Dic99, Dor20, Eva09b, Fon73, Gin01g, GB09, Hos83a, KD90, Mil04, Rot17, Sad90, Sie05, Spe13, God07, Har13, Sel12]. Searching [Bon05, Whi00, Maf05]. Seas [Sch00b]. Seasonal [Kno03]. Seasons [Pan09a]. Secant [VB21]. Secchi [Cen14, Cen14, Udí21]. Second [Car19, Kid90, Kni70, Kra09, Swe17a, Sha92, Ano80a, Kni70]. Seconde [Bla01]. Secrecy [Sha07c]. Secret [Ait86, Ald97, Cam02, Wil82, Mos10]. Secrets [Bar19, Eva09b, Hir09]. Seculo [GM03]. Secunda [MG86, Swe86a]. See [Doe00, Cov17, Hoy81, Lan80, Pet04, She99]. Seeing [Hoc92, Lin07, Sul12]. Seekers [Gol87b]. Seeking [ABB03]. Seeleneinstimmung [Dru84]. Seen [Bla98a, Gra13a, WR08]. Segno [GM03]. Segonds [Ler11a]. Seicento [Van96]. Sein [Brü87, Hor74, Eel83, Rut12, Sha99a]. seine [Gin90f]. seizième [Ler80, Bla01]. Sejong [Ste94]. Sela [Jus20, War19]. selbst [Gin90f]. Selbstsreflexion [Kre82]. Selected [Bro10, Gin83b, Hos76a, Hos97b, Hoy84, Ken85, Ost97a, Rap13, Rob86, Ste05, Bar13b]. Selection [Rug83]. Selections [Rap13, Ree10]. Selene [WW19]. Selenographer [Sta00a]. Selenographia [Pum11, Mül10]. Selenology [Hoy82]. Sellers [Apt14]. Selling [Waf02, Waf02]. selon [Nor82]. Semi [Gin85e]. Semi-Tychonic [Gin85e]. Send [DeV19]. Seneca [Pin92]. Senecas [Pin92]. Sense [Woo88]. Sent [KY95]. September [Liv13]. Sepulchres [Hos92b, HAG94b]. Sépultures [Hos00b]. Serabit [BSF08]. Série [Wil03]. Series [Bel14a, Gin85e, Hos89b, Rob86, Not23b]. Serpent [Car86]. Serpents [FA97]. Service [Edd80, Hos92a, Li19, Mer96, Mor01a, Sho21]. ses [Hen76]. Sesi} [TSH92]. Sesquicentennial [How88, Rot01]. Set [Sal85]. SETI [Cor98]. Setting [Pas03a]. Settings [Rug85, LS24a, Sta72]. Settlement [HB92]. seu [Hos89b]. Seven [CH92, Gin03g]. Seven-Year [Gin03g]. Seventeenth [Apt14, Ben87b, Dan12b, Dun92, Fei86, Min00, WW74, Whi73a, Sel12, Aub16, Bau87, Ben84, Bro88a, Bro91b, Dan12b, Gin97c, GM03, Gin13c, Rus74a, Rus74b, Spe13, Ste90b, Str10, VG01, Wer13, Whi72]. Seventeenth-Century [Min00, Aub16, Ben84, Dan12b, Gin97c, GM03, Rus74a, Rus74b, Spe13, Ste90b, Wer13, Whi72]. Seventh [TQ18]. Seville [Pal12]. Sewers [GDEK24]. Sextants [Ree11, Ree11]. Seydi [Kin79]. Sfere [Swe92b]. Shātir [Rag16b, Sal87]. shade [Par85]. Shadow [Ben87b, Fin08, Pas15, Vet19]. Shadows [Duk12, Hig16]. Shafts [Wal07]. Shaker [Nor92a, Oes15]. Shakespeare [Wło15, Fal14, Ush15, Wło15, CWW17]. Shamans [McC98b]. Shame [Hos02a]. Shams [Sal94b]. Shane [Ost84a]. Shape [Hys96]. Shaped [Ave11, Gin01a]. Shapes [Cam11]. Shaping [Rot17, Har13]. Shapley [Ber70, Gin88d, Het76b, Hos16a, Smi00]. Shatir [Moz22]. Shea [Ano81a]. Sheehan [Cro15]. Sheepshanks [Hos91a, Hos89a]. Shetland [TM78]. Shifts [Liv13]. Shines [Bec99a]. Ships [And00]. Shirazi [Mor16]. Shixian [CML23]. Shkola [Kul87]. Shlomo [Jus20, War19]. Shoemaker [Yeo02]. Shooting [VB01]. Short [CT12, Ler17, Wil04, Pin88]. Short-Lived [CT12]. Shoujing [Ste94]. Should [Bec99a]. Shoulders [Fei95]. Shoushi [LZ99]. Showers [Ano78b, Dal78, Shi11, UC17]. Showman [McK78]. Shrines [Zei85b]. Shrouds [Lad10]. Shukuk [Gol73b]. Shumero [Sol09]. Siblings [Kra11]. sich [Gin90f]. Sicilia [Rap11c]. Sicily [AR00, Chi15]. Sicut [Ler80]. Siddhāntasundara [Gan15, Knu14]. Siddhánta [Pin84]. Sidereal [ABB03, Fin90, Fin10]. Sidereo [She74]. Sidereus [Hos89b, Lüt00, Van94b, Whi78, Fin10, Fin90, Fin10, GV03, Hei12, Van94b]. Siècle [Cha09, Ler80, Ler11b, Lüt00, Gin88b, Nor82, Bla01, Cha89]. Sighting [Gau14, Ste08b, WDR11]. Sighting-Criterion [Gau14]. Sightlines [Rug82, Rug83]. Sigismund [Oes15, RR13]. siglo [Gli88]. Sign [Bar16, Mol03]. signals [BC16]. signes [BN81]. Significance [Ano83b, Gin00a, Gra11b, Het74, Jar82, JLGL14, Rya71, TT71, GB86]. Signis [Rut12]. Signs [Gre99, BN81, War19]. Signshadow [GM03]. Silke [Boh19]. Simms [McC94]. Simon [Eva17, Smi94c, Sti07, DOL98, HRVM13, Pas15, Smi94c, Hei99]. Simone [Wil15]. Simplicius [Eva07, Eva07]. Simplified [Tur70b]. Simultaneous [Het75, Kol08, LS24a]. since [Bon11, Gin11a, Whi92]. Singh [Bla79, Gin96b]. Single [Sul14, Mun13]. Singular [Kni70]. Sir [Bec19, Big10, Din73, Gin99d, Hos08c, Hos13, Cas18, Din73, Sel06]. Sirius [Cer95, Cer96, HW07, Hol10a, Kol01b, Kol07, Sch00a, Sch07a, WR08, Sch07a]. Siromani [Pin84]. Sistemi [Gin02c]. Sister [Sha84]. Sistine [Shr00]. Site [Rom06, SS90, ZJJ^{+}99, Aug24]. Sites [CFMR77, Ell85, HC99, RMH91, TT78, TM78]. Situating [Lat20]. Sitzungberichte [Wes97]. Siwa [SBF07b, SBF07a]. Six [Dic21, Kol07, Mea70b]. Sixteenth [Gin77c, Gin11b, Jar90, Péo08, Ver17, vdW17, CMM^{+}15, DS89, Gin76, Gin83a, Gre00, Grö83, Kre89, Kre11, Ler80, LG12, Nor80, Sal94b, Sha99a, Spe13]. Sixteenth- [Spe13]. Sixteenth-Century [Gin11b, Jar90, Gin76, Gre00, Grö83, Kre89, Kre11, Sal94b, Sha99a]. sixth [MHS23]. Size [Gol96b, Gra13b]. Sized [Wil04]. Sizes [Hei16]. själen [Sha07a]. Skeptical [Sch15, AB13]. Skepticism [LG12]. Sketches [Hos08c]. Sketching [Kes15, Nas13]. Skidi [Had84]. Skies [Gin80c, Rug07, War79, Win25, Hos10b]. Skills [Kwa12]. Sky [Arn95, Ave97, Bel16b, Gin80c, Had84, Hea97, Hos06a, Jar09a, Kra22, McC83a, McC07, Mea86, Smi89, Ste11, Sul14, Ver09, Wil87a, Yeo99, Mun13, Ash82, Gin10b, How88, Ost04, Sch07a]. Skyscapes [Bel16a]. Skywatchers [Gol87b, McC98b, Hiv81]. Skywatching [Ald09]. Slab [SS90]. Slava [Ano15f]. Slavs [Rya09]. Slide [LH05]. Slipher [Liv13]. Slot [CTE12]. Small [Hos05c]. Smithsonian [Doe19, Ger17, Gin87c, Wis06, DeV90, Dic91a, Doe90, Doe19, Dup90, Lat20, Wis06, Gin01e]. Smyth [Eva87a, Eva87a, Hos88, Hos88]. SN1604 [Mes20]. Snatching [Gin88d]. Sneferu [BM15]. Snows [Waf04]. So-Called [Wes84]. sobranii [Rya71]. Sobre [Omo13, Nor00b]. Social [Sch01c]. Société [Ano77i]. Sociétés [Hos00b]. Societies [Sid10]. Society [Ano76b, Ano94c, Ano15f, Ban13, Ben20, Bre18, Gin81a, Gin81b, Heg78, Heg82, Hin04, Jar90, UL95, Vic85, Wło23, Gin81b, HR92, Kel83, Ron89, Tho90b]. Sociologist [Smi18b]. Sociology [Sul02]. Socorro [Had84]. Socrates [Gre14]. Software [Fat00, Ols93]. SOL [UCI16]. solaires [Par85]. Solar [Ald02, Ano95b, Bar88, Bec00, Bre00, BSL86, Bro88a, Bru96, Cas19, Cha02, Coh86a, Cro12, Cul07, DeV05, Duk08, Edd71a, ECT10, FA97, Far93, Gra08b, Gur06, Hen99a, Hen99b, Hos80b, Huf91, Jon91b, Jon12, Jon16, KMD^{+}18, LZ99, Mac15, Moe75, MSH19a, New77, OP92, Pan94a, Pan94b, Pan95, Plo77, SS95, SS96, SS97a, Sak80, Sal85, Sch05b, SS86, Ste97, SH81, SF95, SF97, SS06, Ste08a, Ste10b, SMH18, SMH20, SW18, TQ18, Thu95, Tre90, WS19, Wes79, Whi92, Wło07, Yeo07, Zei85a, GC01, KF70, Ma23, MSH22a, MHZS23, MHS23, Not24a, UC22, Mac15, Bru96, Haw74]. Sold [Smi94a]. Soldiers [Ano15f, Ger14]. Sole [AH83, Rap11c]. Soleil [Dob86, Ler80]. Solid [Gra00]. Solis [CG12]. Solstice [All92, Hos97c, SS90]. Solstices [Thu01]. Solstitial [BAE93, EBA94]. Solution [Swe00, Van74a, Wes09a]. Solving [Cer96, Mor01b, Kin73, Sta95]. Some [Ano04b, Ext92, Jon86, Kak95, KG82, Mea71, ML14, Pin98, Ros82a, Rug87, SG05, TM78, Tur77, Van71a, Van71b, Woo80, ZB04, Not24a, Ano84b]. Somer [Gin01b]. Somerville [Eis85, Nee08, Nee08]. Sometimes [Zso10]. Somnium [Voe97]. Son [Ben20, Hos12b, Lan87, dA17]. Songs [Hos06a]. Sonne [Hub83, Fei13, Hei14b]. Sonnenuhren [Eag08a, Eag08b, Eva08]. Sons [Dew70, Gin97b]. sopra [Gin02c]. Soul [Bon05, Sha07a, Gin14, Ste88]. Soul-Searching [Bon05]. Source [Ano71d, Bar18, Dob83b, Het76b, Kre10, Kun86a, Mer88, Not15a, Ost80, Pou75, SL13, Sha96, She15, Tri97, WW71]. Sourcebook [Kre08, Cer99]. Sources [Ano78b, Arm74, Dal78, Dob83b, Gin07b, Mer88, Sar78, SL83, UC17, UC19, Van71a, Van71b, Wil84, Kno97, Mon22, Oos15, Pou85]. South [Ave05, Hos91a, SCV20, CMP24, Hos89a, Pow95, Rug94, Rug96]. South-West [Rug94, Rug96]. Southern [Bla91a, Edm98, Hea97, Orc90, Sch02, Smi92, War77, War79, Ano97c, Che99, Hos10b, Sne98, Wil96]. Southwest [Rot03, HVT^{+}01, You86, Zei85b]. souvenir [Bir74]. Sovereign [TQ17]. Soviet [Ano15f, Ger14]. Soyria [PP82]. Sozietdt [Wes97]. sozvezdii [Sol09]. Space [Ano15f, Bed10, Ben76b, Bro10, Cro15, Doe08, Eva82, Ger17, Gol03b, Hir92, Huf06a, Kes10, Kru08, Lat20, Lau14, Lov87, Mac93, McC07, McC13c, Rem15, Sis85, Spr92, Tat99, Tat09, Tri10, Tur72, Van95b, Waf91, Ben14b, Ger14, Ano15f, Dic00a, Hos02a, Mac93]. Spacefarers [Neu13, Ger17]. Spacefaring [Dic00a, Hir92]. Spaceflight [Bro10, Neu13, Ger17]. Spain [Ano17c, Hos92b, HVT^{+}01, Iwa96, Kun80, Lei17, RC16, UC16, Mer96]. Spanish [Mos10, Mos10, Rec22, Smi93, UCI16, UC23]. spätere [Gre00]. Speaking [Gra20]. Special [Bre18, Gin82a, Mil81]. Species [Rab05]. Spectacle [Sul02]. Spectra [KD90, Sut72]. Spectrographic [RS22]. Spectrographs [Sta10]. spectroheliograms [Mal23]. Spectroscope [Hos07a]. Spectroscopy [HD06, Mea02, Ost86, Sut72, Hea14, Hen15b, Hen15b, Eis88a]. Spectrum [Hen99a, Hen99b, Mea02, Plo77]. Speculatie [Van98b]. Speculum [Ped79]. Speed [vH83]. Spencer [Bar88]. spera [Pou79]. Spey [Sco90]. Sphaera [Gin15, Cro23a, LN09]. Sphaerica [Ber12]. Sphaerical [Gin03b]. Sphaericorum [Gin89b]. Sphären [Dru84]. Sphere [CMM^{+}15, Sid04, Sta72, Swe99b, Pin81, Gra07a]. Spheres [Bol05, Don73a, Gin85e, Gin03d, Gin04b, Gra06, Hos02a, Nor07, Sal94a, Swe72, Swe92b, Fek24, Bol05, Wes77b]. Spherical [Dru84, Gin03b, GH07, Plo00, Kin73]. Sphericity [Voe02]. Spherics [Ber12, Sid24b, Sid25]. Spheroid [Déb91]. Spin [Van14b]. Spinning [Sch11]. Spiral [Gin92a, Het74, Het75, Hos76c, Hos82c]. Spirals [ASCE14, DH91, Liv13, Mar23]. Spirit [McM05]. Spiritual [GG14]. Spitzer [Ost97a, Smi07, Ost97a]. Spokes [SD03]. Spot [Hoc92, Mac15]. Spotlight [Sul12]. Spots [Pas03b]. Spread [Fin99]. Spring [Da 04]. Springer [Apt14, Bec15, Cro15, Cro23a, Gee18, Hea15, LF15, Löh19, Mir17, Mit19, Mol14b, Roc17, Sil23]. Spuren [Kri97]. Spurious [Wło10]. Sputnik [Bru96]. Spyglass [Sha07c]. Square [AZB02]. St [Oes15, RR13, Wło15]. St. [Bot76, Cro23b, Oes15, Kri97]. Stadt [Wil75, Wil75]. Staffs [Ree11]. Stalinist [Nic90]. Stampa [Gin10a]. Stamping [Dic13a, LF15]. Stamps [LF15, Rap20]. Standard [Kra20]. standardization [Boh23]. Standards [Cah19]. Standing [Sco90, Fei95]. Standstills [MEA91]. Stanisław [Ave95, Jas20]. Star [AB13, Ano01b, Ash81, Bon07, Cer99, CS14, DE00, Dob02, Duk02, Duk03, Eva87b, Eva87c, Eva92, FZ07, Gin70a, Gin81c, Gin00b, GHHB02, Gin06b, Gin12a, Gra07b, Gra10, Gra13b, Hos08g, Hos11b, Hos13, IK80, KPG03, Kri93, Kul80, Kun80, Kun86b, Kun91, Loc83, Loc92, Moz16, Nor08a, Pin88, SC88, Sch01a, Sch13, Smi12, Smi18a, Sol09, Sta06, Ste14, Swe86b, Sym02, SC13, Tho84, Ver13, Wal07, War71, Wri07, Zso94, Bec22, Cas21, Moz22, Pop22, Sch15, She90, Nor08a, Wil79, Cer99, Coo75, GWZ22, Kru97]. Star-Clock [Loc83]. Star-Size [Gra13b]. Stardust [Dic13b]. Starého [Gin83a]. Starlight [Aub12, Chi16, Eis88a, Hen15b, Lil99, Hea14, Cen17, War80]. Starowolski [Ait82]. Starry [DeV03, Gin90f, Gra19b, Gra21b, Wil15, HDG^{+}13]. Stars [And00, Apt99, Ave09, Ave11, Bec19, Ber92, Bon09b, Cas18, Cha15a, Cro12, DK83a, DK83b, DeV00b, De 06, Dek20, DeV94, Eas96, Gau18, Gen90a, Gin04c, Gra13b, Gui85, Had84, Hei16, Hol09, Hos77, Hos79a, Hos08f, Hug10, KD90, Lay02, Lin07, Mal97, Mar12a, Mer92, Mor97, Ogj11, Ost97a, Phi91, RC92, Rug99, RC16, Sch07b, Smi00, Ste16, SG09, Swe92a, Tau13, VB01, Wag87, Whi71, Wil07, Wło90, Zso10, Zso20, CM23, Ber70, Ber81, Bry81, Dic13b, Gin75a, Gin06b, Kre89, Kun91, Mea70b, Tay10b, Wag88]. Started [App05]. Starting [Zso10]. State [Jas96b, Sch91, Smi98]. Statecraft [Ald03]. States [Huf91, Ano97c, Rot17, Smi97, Waf01, Waf02, WW71, Wil96, Waf01]. Stations [Atk76, ABB03]. Statistical [Bel14a, Ber97a, Ell85, GB10, GZB14, Gin88a, HC99, RAGGB18, Ros82a, Ros84]. Statistics [Wil88]. Status [Rug82, AKJ^{+}23]. Stavba [Nor02]. Steady [Smi98]. Steady-State [Smi98]. Steele [Roc17]. Stefano [Ree20, Tuc00]. Steine [Jas96a]. Steiner [Bur15]. Stelae [Ald02]. stele [Lip15]. Stella [Car25]. Stellar [Ber82, DeV75, DK83a, DeV05, Fea00, GB09, Hos86b, HD06, Huf06b, KD90, Mal15, Omo20, Ost86, Pet20, Sch86, Sie05, Smi06, Smi22, Wag88, Wes09a, KHM^{+}12, Ber82]. Stellar-Energy [Wes09a]. Stellarum [Lin07]. stelle [GS13]. Stellibus [She74]. Stephen [Bec19, Bra15, Mia12]. Stephenson [Duk16]. Sterken [Kra22]. Stern [Gin81c]. Sternbilder [Mal90]. Sterne [Gau10, Gin90f, Kri97, Oes15, RR13, Jas96a]. Sternkarten [Eva06]. Sternkatalog [Eva92, Gin00b, Too77]. Sternkoordinaten [Gin00b]. Sternkunde [Siv77]. Sternstunden [Rut12]. Sternwarte [For73, Gin80b]. Sternwarten [Kri94, Her76, Kri94, Mul81]. Sternwarten-Architektur [Her76]. Sterrenkunde [Van02a, Nor91a]. Stew [Cul05]. Stewart [Ano74d, Kin10]. Sticker [Ano78a]. Stieda [Oes15, RR13]. Still [Car17b]. Stillman [Mac94]. Stimulus [Mar99]. Stockholm [Kär24, Tur79a, Tur79a]. Století [Sch14a]. Stone [Bla79, Gin96b, Hos01a, Hos04e, PF83, Rad90, Rug94, Rug96, UCI16, Waf01]. Stonehenge [Ano82e, Atk76, Dib76, Rug99, TTT74, TTT75, Edd74, Rug99]. Stoner [Nau08]. Stones [LL87, Rug96, Sco90, VSH93]. Storer [Bro88b]. Storms [San99]. Story [Cap12, Dic13b, Dra80, Hea15, Heg85, Lau14, Lep14, McC99, ML14, Pas03b, Sul18, Van98c, Wil07, Nat13, Wen24]. Straight [Ruf71]. Strange [Cro12]. Strangeness [Ano84b, Woo80]. Strasbourg [Ano83b, Pou83]. Strategies [Mac90]. Strath [Sco90]. Stratosphere [Van95a]. Strawberries [Cul05]. Street [PC87]. String [Gre00]. Strings [Gen90a]. Strömgren [Reb03]. Structure [Eva93a, Gra97, Hos71b, Huf06b, Mur95, Pow88, Ric07, Siv10, Smi06, Wes93, Wes09b]. Structures [Dan22, Hen15a, NR14, WP22, McC91]. struggle [Kut23]. Struve [DeV89, KI10, Oes15, RR13, KI10, DeV89]. Students [RS18]. Studi [Hos04c]. Studia [Ait82, Jar73, Nee74]. Studien [Ano06a, Jor94, Swe76, Gin15, Sha82]. Studienplan [Har94]. Studies [Ald09, Ave95, Bel14a, Ben99, Ber82, Ber01, Chi15, Dal01, Eas00, Gin91d, Gin99b, Gin11b, Har92, Hos76a, Hos86b, HAG94b, HAG95b, HAG95c, HP98, HMA^{+}98, HS99, HVT^{+}01, Hos02d, Lan83, Lut91, McC07, McM03, Mer96, Mir17, Mor08, Mur98, Nor88, Nor00a, Plo93, Ram10, Rei92, RH00, Sal06, Sis85, Ste90b, Cun16, Bar05, Ben82a, Ben91b, Kun85, Nor06, Ste90b]. Study [Ell85, Eva93a, For74a, Gin03b, Hol10b, Hos97b, Pan09a, Rag85, Roc16, SB84, SG07, Tih07, Waf91, War79, Zep19, Rec22, Van10, Ano15f, Kin78]. Studying [Doe00, Whi00]. Stukeley [Hos85c]. Stuttgart [Bur15]. Suanjing [Siv98]. subject [Mia12, Bra15]. Sublime [McC13c, McC13c]. Sublunary [Sch03]. Substantial [Rab05]. Success [Mue00, Ost86]. successors [GE24]. Sudan [BFAH^{+}10, CLL^{+}00]. Sudhoffs [Bur15]. Sufi [Sch13]. Sui [TQ18]. sulle [Str10]. Sullivan [Sil23]. Sumero [Sol09, Sol09]. Summaries [Ano92h, Ano93i, Ano94h, Ano95i]. Summary [Ano96e, Ano97i, Ano98h, Hos08h, Kin95a, Bec15]. Summits [RM92]. Sun [Bre00, Gin91b, Gra13a, Kre93, LL87, Mor97, PP07, Sch05b, Whi92, UC23, Ave05, Bau87, Byr11, DK83b, Eas01, Gol01a, Gra22, Gra19a, Jon16, Ler80, MSH20, Moz13b, Moz13c, Oss24, Pas03b, PP07, Plo77, Pow88, Rug99, Sha75, SP12, Swe99b, Van98a, Ves98, Zei85b, ZH12]. Sun-Centred [Eas01]. Sundial [EM08, Loc89, Sav12, Mas21]. Sundials [Arn11, CLL^{+}00, Che73, Dun20, Eag08a, Eag08b, Eva08, Gin03e, Llo89a, Llo90, Loc93, Loc95, Sav12, Saw80, Sch01c, SK16, Tau94, Tur70b, Han21, Par85, Llo90, Saw80]. Sundials. [Tur70b]. Sundry [Gin04c]. Sunniya [Rag85]. sunrise [BN81, CML23]. sunset [BN81, CML23]. Sunspot [Bau87, Gin03h, Mue00]. Sunspots [Bau87, Bon12, Pas03b, Top03, Bon12]. sunt [Pou84]. suo [McM89]. Superior [CTE12, Moz13a, Swe17a]. Supernova [Ash80, Ham06, Bot76]. Supernovae [NRKN16, PG79, SG05, Tri03, WS16, PG79]. Supper [She79a]. Supplement [Gin06e, McK87]. Supplements [Ano79a]. Supporting [Gra10]. Supposed [Wło10]. Surface [McC17, RP87]. Surrey [Bel14a]. Survey [Coo99, Duk11, Fre77, Gin96a, Gol76, Rug07, Sam13, Sco90, Tih07, UC16, UC17, Vil90, Mor21, Tur23]. Surveyed [Hos04f, Rut03, Wer10]. Surveying [Kru17, And90]. Surveys [PF83]. survie [PS98]. Survival [McC89, PS98]. Surviving [Ben81b]. Súrya [Pin84]. Suyuti [Rag85]. svetil [Sol09]. Swedish [Tur79a, Sha07a]. Sweeper [Hos05c]. Sweepers [HW81]. Sweeps [Hos05e]. Swenson [Ave18]. Swerdlow [Duk05, KE22]. Sydney [Orc84, Orc84]. Sygkepleriazein [Bia15]. Symbolism [BM15, Gin01f, McC90]. Symbols [RS84, Zei85b]. Symmetry [GH07, HG04]. Symphony [Gin85b]. Symposia [How88, Rob86]. Symposium [Ano77j, Ano78d, Ano84b, Ave95, Ben82a, Ben91b, Cár88, Déb87, Dob79, GHHB02, Hen76, Moe77, Nor78, Rot01, Spr93a, Wil15, Wil75, HE82, Woo80, Ken82, Loh75, Sha82, Wal77, Kra22]. Synchrony [Sal06]. Synodic [Gol03a, Ste10a]. Synopsis [Gin85g, Hos85a, Wag88]. Syntaxis [Too77]. Synthesis [Gin14]. Syria [Ack05]. System [Bru96, Car08, Cas19, CS02a, Gol02a, Gur06, Jar90, Kra17, Ler05, Omo11b, Pin84, Smi06, Ste02a, TQ17, Van06, Kur21, Cro12, Gau87, Yeo07]. Systems [Car99, Cul05, Gin85e, Gra08b, Kel99, Kul75, Mal78, Qua17, Cul17, Qua17, Mul80, Ric07]. Syzygies [Hos95a, Kok98, Kre11, PC01]. Syzygy [CG97, GC19]. Szpiech [Boh19].

T [Sil23, Sch23a]. T. [Hoy81, Lan80, Pet04, She99]. tōn [LS24a].

Tūnisī [Not24b]. Tūsī [dB95, Ves73]. Table [Ano95a, Ano01c, Cul07, Gin16b, Hus12, Mar95, Moz16, SC13, Thu94, Fal23, Moz22, Rem15, Cau25]. Tables [Bar18, Bel14a, BH12, BSL86, CG09, Cha12, CG16, Cha17, CS14, Dob77, Dob83b, Eva88, Eva01, Fal23, Gin91b, Gin06c, Gol82, Gol01b, GC08, GC19, Hog88, Hus12, JLGL14, Jon16, Jon17, Kin87, KMD^{+}18, Kun80, LGJLF14, Mar12b, Nor76, Nor85, Pin70, Pin80, Pou85, Sal76, Sam13, SH81, Swe04b, Van94a, Vil88, VG01, vB10, vdW17, vdW19, CP23, Gol84, Kin73, Mer98a, Not24b, Pou88, Pri91, Ste23, Zie23, Cha07, CG18, GC01, Jus23, Kre07a, Kun86b, Ler06, Nor85, Not15b, SC88, Sam19, Swe04b, Tho71c, CT93, Eva93b, Gin06c, How80]. Tablets [Wei82b]. Tabula [And17, Ler80]. Tabulae [Tho71c, Cha98, Kre07a, PC01, vB10]. Tabularum [Not15b]. Tacchini [Cen14, Rap11c, Cen14]. Tactics [Van85]. Tadhkira [Ans05]. Tahdid [Sal79]. Tähetorn [Hig12a]. Tails [Hen09]. Taking [Eva13]. Talayotic [Hos85d, RBH00]. Talayots [AZB02]. Tale [PP07]. Talent [Smi02]. Tally [VSH93]. Tandem [Bon09b]. Tang [TQ18]. Tangled [Gin05e]. Tanis [EM08]. Tanstetter [Sha99a]. Tappan [Gin00c]. Tarde [Bau87]. Tarihi [Hei00]. Tarnished [Cas16]. Tartu [Hig12a, Hig12a, Hig12a]. Tashkent [TA02]. Tassinari [Cen14]. Taulas [HHK90, Hos91b]. Taurus [Dre15]. Tawantinsuyu [Urt21]. te [Van98b]. Teach [Ban10]. Teaching [Ler00, Rus74a, Rus74b, Voe01, Mea02]. Tears [Gra80]. Tebbutt [Orc88]. Technical [Tur70a, CJ22]. Technicians [Ave18]. Technik [Ano06a, Oes10, Sta99, Sta99]. technique [KF70]. Techniques [Bel14a, Hus12, Mea02]. Technology [Ano09a, Gin87c, Mac93, Pan94b, Pan95, Smi02, Waf91, Pas14, Ben99, O’D06, Tur71b]. Teil [For83]. Teiledition [Kre07b].

Telegraphy [Vet12]. Telescope [Arm72, Bak15, Ben81b, Bol10, Bre18, Cap12, DeV19, Dup03, Dup08, GMML14, Gin12c, Gra10, Hei08, Huf06a, Jar99, Lau07, Lin79, Low88, McC13c, Ree16, RL18, Sim92, Sim09, Slu97b, Spa03, Tat09, Tur77, Waf91, War98, Zui19, Zuj11, Mar22, Hig12b, SD82, Cap12, Bol10, Gin10c, Gin12c, Hig12b, Hos97a, Lin79, Thi88, Yav99]. Telescope-Building [War98]. Telescopes [Ben76b, Bre17, Cha72, Edd86, Gin17, Hev94, Hos15a, Kri94, Kwa12, MBH16, Mol14b, O’D06, Pas14, Rap18, RC16, Sha07c, Slu97a, Spa04, SG16, Tur79b, Tur84, Van15, dV91, Ben93a, Hev94, Van85]. Telescopic [Ben93a, Dob99b, Dra76, Gin03c, Gol96b, Mül10, Pum11, Thi88, WKH18, Dic90]. Telesio [Omo13, Omo13]. Tell [Van98c]. Tempel [Naz16]. Temperature [DK83a]. Temperatures [Pet20]. Temple [AR00, FA97, JLGL14, LGJLF14]. Temples [Bel01a, BS06, BSF08, BFAH^{+}10, EBB^{+}01, Hos04e, PH94, SHV92, SB05, SBF07b, SBF07a, HAG93]. Templo [Spr00]. Templos [Hos12a]. Temporality [Mur98]. Temporary [Mul20]. Temps [Hos04f]. Ten [Mul84]. Tenerife [BAE93, EBA94]. tenmon [Nak71]. Tenochtitlan [Spr00]. Tenth [Arn11, Pou09, For16]. Tenth- [Pou09]. tenth-century [For16]. Tenuous [Bon09b]. Teologia [May11]. Teologo [Lat12]. Teorie [Str10]. Teotihuacan [Ave05, PC87, RS84]. Teratology [Van06]. Tercentenary [Ben93b]. Terminology [Dal80, Kra24b]. Terms [McM89]. Terre [Kus10, Ver09]. Terrestrial [Ben20, Bon19b, Bon19a]. terrestrisque [Ste88]. Tertius [Bon09a]. Tessicini [Oos15]. Test [BG12, Kol08, MS85, Sta07, Cre06]. Testimonia [Pin81]. Testing [BFAH^{+}10, Cah19, Das15, Gli88, Sta07, Tur84]. testo [Omo11a, Swe92b]. Tests [Hys96].

Tetrabiblos [CS21]. Text [Arn11, Gin89c, Gin96c, Hos02a, KMD^{+}18, Ler00, Pal12, Rap11a, Sal94a, SV16, SL13, Siv98, Swe81, Don93, Gol11, Gol11, Cau17, Kel10, Swe88]. Textbook [Eva05, Gin88e, Lat19, Bar22]. Textbooks [Gin88b]. Texte [Eva88, Sha94a, Swe98]. Textes [Dob83b, Wil03]. Texts [Arn95, Ber92, Dal01, Eas92, Gin72, Hub91, Kre13, McC91, Mer88, Pou84, Sid10, SS06, Pin98, Ste13, Gin92c, Hos97b]. textual [Not21]. Teyler [Ano76c]. th [Tri10]. Thabit [Rag92, Rag92]. Thackeray [Fea00]. Thales [Gra14a, Pan94a, SF97]. Theaters [Kru08]. Theatre [Loc89]. Thebes [LB06]. Their [Ano04b, Bel01a, Ber92, Coh91, Dew00, DK96, Eng02, Gen92b, Gin87a, Gin01e, Gin06b, Kru97, Kun88, Mul80, Nor04, Pas03a, Pin88, RM92, Sch05a, Sta10, Tri03, UCI16, VG01, WW74, ZB04, Not15a]. Theistic [How02]. Thema [Har84]. Theme [Van85]. Themes [Mea71]. Thémon [Mol76]. Théodore [JT92]. Theodosios [Sid24b, Sid25]. Theodosius [Ber12, Ras13, Ber12, Ras13]. Theological [Gin11b, Mar99]. Theology [Far04, Gin87b, Gin11b, Mal97, Hay81]. Theon [Eva88, Sha07b, Jon12, Pin80, Eva88, Eva01, Pin80]. Theorem [VWH97]. Theoretical [Aab72, Pin03, Hen92]. theoretical-historical [Hen92]. theoretischen [Swe99a]. Theoreumacie [Not20]. Theoriam [Don00]. Theorica [Byr11, Swe73, Ped81, Gin77a]. Theoricæ [Pan12, Bar13a]. Theoricas [Mal13]. Théorie [Neu81, Nor82]. Theorien [vdW78]. theoriengeschichtlichen [Hen92]. Theories [Ano84b, Ber97b, Bur10, Eva82, Eva93a, Hen09, Het76a, Hos04f, Moe75, Mul80, O’B88, Rut13, Smi98, Sut72, TQ18, Tri97, Yan86]. Theory [BFAH^{+}10, BH73a, BH73b, Bri98, CT12, CS02b, Cul02, Duk08, Eas00, Eis06, For70c, Gin82a, Gle07, Gol97, Gra13b, Gra12e, Hos03a, Jas20, KD90, Kni70, Kra99, Kre81b, McC99, Mil81, Nau01, New77, Ruf71, Sal87, Sch80, Swe72, Swe73, Swe75, Swe76, Swe89, Swe99a, Swe04a, Swe17a, TG87, Whi72, Whi73b, Whi71, Wil87d, Wil03, Wil08b, Wil11, Wło11, vH87, vdW74, CR22, Pop22, TW25, Wil87c, Hen92, Hen97, Wes77a, Ken87, KH73, Mar89]. There [Rya09]. Therkel [Wło23]. Thermodynamics [Gre09]. Theses [Ano70g, Ano70h, Ano71k, Ano72h, Ano72i, Ano73c, Ano77k, Ano80e, Ano82l]. Thesis [Gin85e]. Theutonicus [Not23a]. Thierfelder [Gau17]. Things [Gin03g]. Thinking [GB11, Van89]. Third [Bon09a, Ver06, vN12]. Thirteenth [Gra05a, Fle92, KG82, Sam98]. Thirteenth-Century [Fle92, KG82]. Thirty [Lat23]. Tholos [HAG95b, HVT^{+}01, Mar02]. Thom [Atk86, Fre77, Rad90]. Thomas [Hos08d, Hut12, War98, Edd72, Hen08, Hos70, Hos81, HR92, Hos03a, Don86, Hen08, Loh75]. Thomson [Hoy84]. Thoren [Wes91, Lat93]. Thought [Bra20, Van15, Har13, Rot17, Bru78a, Liz17, Sch70]. Thoughts [Kni70, Whi70a, Don93, Sha92]. Thousand [Sch13]. Threat [Ano09a]. Three [AH83, BBW97, Dic06, Gin85e, Gin87c, Har92, Hog88, Jon21, Lan83, Loc95, Mes15, Qua17, Rug96, Sco90, SS90, Sta95, Thu95, Zso94, Cul17, Fre14, Ano95b, Ben84, Don88a, Har75]. Three-Slab [SS90]. Throwing [Gin96d]. Tibetan [Pet80]. tibetischen [Pet80]. Tides [Coo00, Van85, Coo00]. Tierkreiszeichen [Bur84]. Tierra [Ave87]. Ties [Bro88b]. til [Hol03]. Timaeus [Llo11, Llo11]. Time [Ald11, Ave92, Ave94, AMP95, AMP96, Ave18, BD81, BD82, Bed80, Ben99, Bro90, Cle71, Dav15, Dew79, Dic82, Eva11, Gin86b, Gin88c, Goo09, Han09b, Jhi74, Kel99, Kru08, Kru22, Kul81, Mar12b, McC07, Moe83, Mul82, O’B88, Par90, Roc08, Sak80, Spr92, SF94, Ste12c, Str23, Ter84, Tum74, Waf02, Wło07, Ben14b, BC16, CJ22, GJ22, Rem15, Ald97, Ave92, Ave95, Bru09, Dun20, Eva11, Gin78a, Gin85c, Gra01, HS81, Kre89, Ste12c, Syl07, Tur72, Woo87]. Time-keeping [Str23]. Timekeeping [Ben20, Sal06, Smi81, Waf02]. Timely [Gin88c]. Times [Ano77a, Arm73, Dic09, Gin87a, Hol77, Mor01a, Nor00b, SS95, SSM97, SS97b, Ste97, Ste98a, SF93, SG09, TT83, CL22, CML23, BN81, Nor08b, Pal12, Sha87a]. timings [MSH21]. tinnar [Sha07a]. Tinsley [Eis01]. Titelbilder [Ano06a]. Titius [KH73, Jak72b]. Tobias [Her72, Wil11, For70c, For80a, Kle14, Wil11, Gin82b, Kle14]. Today [Llo11, Moe77]. Todd [SM04]. todos [dA17]. Together [Sal95]. Toke [Gan15]. Toledan [Gin06c, Gin06c]. Toledo [Swe04b, Cha07, Swe04b]. Tomb [GB10]. Tombaugh [Rea92, Rea92]. Tombe [ZHGM96, HAG93]. tombes [PP82]. Tombs [BEG98, BEBM02, HAG95b, HP98, HMA^{+}98, HP99, HS99, HVT^{+}01, Le 08, Le 09, Mar02, PHP92, PHP93, PRA87, RH00, SK08, PP82, Bel01a]. Tome [HRVM13]. Tommy [Eva13]. Too [Zso10]. Took [Ano15f, Ger14]. Toolbox [Gin13b]. Tools [Dan12b, Gin99a, O’D06, RPPA15, Rot17, Har13]. Topical [Rot03]. Topics [Wel84]. Toronto [Smi18a]. torquetum [Wło22]. Torreon [DW83]. Tortzen [Wło23]. Toruń [God24]. Total [MSH19a, MHZS23, SS06, MSH22a, UC23]. Totale [Rap11c]. Tour [Hut12]. Touring [Hut12]. Tower [Kra99, Hen92, Hen97, Rod73]. Towns [RAGGB18]. tra [Bel11]. Trabant [Hei14b]. Tracing [CR11]. Tractatus [Tur03]. Trade [Gin02b, Rud92]. Tradition [AS97, Eva92, Gin81d, Gin00b, Gol98, GC06, KD98, Lan98, Mal90, Mar14, Mer06, Mol93, Mos11, Pan10, Pan12, Rag01, Wło07, Lat00, Sal97].

Traditional [Shi09]. Traditions [Ber92, Hos06a, Moz19, Bir74]. Traducción [Cha19]. Traduction [Eva88]. traduzione [Omo11a]. Tragedy [PP07]. Tragic [Mac76]. Tragicall [Tur75]. Training [dH21]. Trains [Edm11]. Traité [Hus13]. Traités [Cha09, Wil03]. Trajectoire [Hos11a]. Transactions [Ano15f, Jar90, Kel83]. Transcription [Ano15f, Van10]. transfer [NN24]. Transformation [Hod77, Ost05a, Sal97]. Transformations [Jon15, Rag05, Roc17, Wes18]. Transformed [Gin03g, Smi91]. Transgressor [Jar09b]. Transit [Bel20, DOL98, Dic09, Gin05c, Gra13a, Kol08, Sch01b, SM04, Van76, Yeo00, vH86, Yeo00]. Transition [Har92, Kre02]. Transits [Bra05, Dic09, Dun13, Gin13c, Gur06, Gur06]. Transjordan [BGP13]. Translated [Cro15, Lep14, Ste98b, Bre18, Swe17c, Too95]. Translating [Cul05]. Translation [Bon09a, Cha19, Eva05, Fin10, Gan15, Gin96c, Gin03b, Gra01, Has11, Kel10, LGJLF14, Neu81, Pal12, Pan09a, Ped87, Pin03, dSP78, RA18, Ros74a, Sal95, Sha02, Tih07, Too81, War93, Wil82, Zep19, Bou23, Van23, Car99, Don93, Knu14, Swe86a, Zie23, Ans05, Lin07]. Translations [CMM^{+}15, Gra08a, Vai12, Ber12, Ras13]. Translator [Gau87]. Transmission [Ans05, Dal01, DS89, Gee18, Gin03a, GC08, Nor06, Sal97, dB95]. Transmitting [Ano06a]. Transversal [Gol77]. Trapped [Don88a]. Trattati [GM03]. Travails [Kri97, Smi07]. Travel [Ban10]. Traveler [War93]. Traversing [Win25]. Travis [Had84]. TRB [GGCF06]. TRB-West [GGCF06]. Treasure [Haz02]. Treasure-Hunting [Haz02]. Treasures [Gin94, War08]. Treasury [Hos02a, Hos05a]. Treatise [Ack05, Eva89, Gan15, Gin03b, Gra14b, Hoo84, Kin02b, Lin07, Nor92a, Nor02, Oss12, Pin71a, RA18, Sam07, Ste88, Swe86a, Tur96, Yaz14, She15]. Treatises [Gin97c, vdW80, Har75]. Treatment [Aab72, Gol96b, Mac98]. Trebizond [Sha07b]. Tree [Edd86]. Trends [Wea90]. Trent [Ped83]. Trepidation [CG03, Pin72]. Très [Not15a]. Trial [Ano05a, Bla91b, Fan09, How05, JLS^{+}05, JHLS05, Lew07, Wes89, How08]. Triangle [Dib76]. Triangles [GH07]. Triangulorum [Gin89b]. Tribiblos [JT92]. Tribune [Dew87]. Tribute [Naz16]. Triduum [Wło23]. Trigonometrisch [Swe90]. Trigonometry [Duk10, Gin89b, Plo00, Duk10]. Trilogy [Bed80]. Triumph [Déb05]. Trois [SM04]. Trojan [Con24]. Tropics [Thu03]. Trotter [Can09]. Trouble [Sta00b]. Troublesome [Fan05, Gin04c]. Troughton [McC94]. Trudy [Ros74b]. True [CG97, GC19, Rot17, Waf02, Har13, PC01]. Trust [Bel20]. Truszczyny [Bar19]. Truth [Lew07]. Trying [Ser13]. Tübingen [CM10, Mar99]. tubo [Kär24]. Tuckerman [SH81]. Tucson [Gin01e]. Turbines [Smi02]. turketum [Wło22]. Turm [Hen92]. Turn [Ano81a, de 22]. Turner [Gen96, TH79, Gen96]. Turnhout [Eva17, Fal23, Zep19]. Turning [App05]. Tusi [Gin84d, Ken94, Ken94]. Twelfth [Eas89, Not14b]. Twelfth-Century [Not14b]. Twentieth [Cor98, Dar91, Ost80, PdV85, Maa16, Bec10, Kai07, Ker19, Liv16, Pau86, Sho21, Smi97]. Twentieth-Century [Cor98, PdV85, Kai07, Ker19, Liv16, Pau86, Sho21]. Two [Aul16, Ave95, Bel97, Bou23, Dic00b, Gin90h, Gin97c, Gol11, Hen75, Hen15b, Kul80, Kun80, LS05, Lil99, Loc90, Loc92, Loc93, Mar03, McM09, Mes21, Osl06, Ost86, Pou84, Sal97, Sch03, SV16, Smi98, Sym02, TTG76, TT84, Tho22, War15, vdW80, Hea14, Wes89, Gre04, Sch12b]. Two-Faced [Sch03]. Tycho [Don88a, Gle04, GB86, Lat93, Péo08, Sha07a, Wes90, Wil00, Ano05a, Bia04, CR22, Car08, Don88a, Gin82c, Gin89b, GV98a, GV98b, Gin07a, Gra06, Gra13b, Ham06, JLS^{+}05, JHLS05, Kre05, Mes21, MJT03, NKM^{+}16, Péo08, Sch13, Sch23b, Sch25, Ser13, Sha07a, Tho73, Van06, Wes78, Wes79, Bia04, Kre05, Tho71b]. Tychonian [Moe75]. Tychonic [Gin85e, Car21, Car08, GW81, Gin85e]. Tychonis [Gin89b]. Tying [Vai12]. Type [Dra80, Nor97]. Typologies [Cha12].

u [Mul84]. U.R.S.S [Bat87]. U.S [Bro10, How88, Ost04, Rot01]. U.S. [Bro10, Dic91a, How88, Kru08, Ost04, Zei85b]. U.S.A. [Rot03]. UBAI [AKJ^{+}23]. Überlieferung [Too77]. Übersetzung [Bur15, Loh13, Eva92, Gau17, Gra08a, Kre07b]. Übersetzungen [Swe98, Swe98]. UCLA [Wal77]. Ueberwindung [Ham06]. Uhren [Gin80b]. Ulrike [Hei14b]. Ulugh [Kul81, Kri93, She90]. Umberto [Ano80j]. Umfeld [Sha99a]. umyslowej [Rav72]. Unblemished [Mue00]. Uncertainty [Bra23, Wil88]. Uncommon [Bra20]. Uncovering [LB06]. Underhay [Had84]. Underlying [Nau98, Whi70b]. Understand [Hos71b, McC99]. Understanding [Bia04, Coo00, Don73b, Hos71b, Sut72]. Undertaking [DeV89]. Unedited [Arn11]. Unexpected [PP07]. Unification [Ano84b]. Unified [Gra09]. Uniform [Jon04]. Uniformity [Goo09]. Unifying [GH07]. Union [Ano79b, Ano70b, SGDO02, Gin90d, Way96]. Unison [Str93]. Unit [Sha84, SF94]. United [Ano97c, Waf01, Wil96, Huf91, Rot17, Smi97, Waf01, Waf02, WW71]. Unity [Dav09, Blu00]. Univers [Gra07a, Har85, Hen76]. universalis [Hos89b]. Universe [Arm73, BH73a, BH73b, Ber82, Car86, Cor98, Dav12, DeV03, Dob83a, Ell08, Fal14, Gin01d, Gin10c, GG14, Gra07a, Gra08b, Gra10, Gra19b, Gra21b, Gre09, Hen08, Hos71b, Hos77, Hos05a, Hos08f, Kra07c, Kra11, Kra13, Liv13, Llo94, Llo11, McM05, Pas03a, Pau86, Ram06, Rot17, Sch80, Smi85, Smi94b, Smi15, Ste88, Str99, Tat09, Tho87, Tri97, Whi71, dA74, Har13, Wło15, Bec10, Dic99, Gra90b, Her80, Hor82, Hos71b, Huf06a, Kni70, Kra13, Lon06, McC99, Mer92, Pal84, Pau82, Smi98, Str81, Tip88a, Whi71, Wil07, Wil82]. Universitat [Cha19, Hei14a, Sha99a, Kre99]. Universitätsbibliothek [Oes15, RR13]. Universitätssternwarte [Ham09]. Universities [Rus74a, Rus74b, Vic85]. University [Ave95, Ave18, Bec19, Bel09a, Bel14b, Bon19b, Bra15, Cha15b, Dev14b, Gan15, Gin77d, Hen15b, Kes15, Kra22, Lau14, Lep14, LD12, McC19, Mes15, ML14, Nau98, Rad90, Rap15a, Ree20, Rot17, Sal97, Sha14, Smi18a, Swe17c, Too95, Ano70h, Bel09a, Pet04, Tau94]. Uniwersytecie [Hil77]. Unknown [Bro10, Kun86a]. Unlaid [Whi74]. Unlikely [Bia04]. Unloved [Sel14]. Unmaking [Gra00]. Unmanned [Lau14]. Unmasking [Ano14a, Van14a]. Unpublished [AH83, Gin92a, Gin95, Gol00, Gra14b, Har92, Mes20, Tee89, Mes21]. Unraveling [Kra13]. Unravelling [Aub12]. unrecognized [Wło25]. Unreported [Don73c]. Unruly [Jon16]. Unseen [Doe00]. Untersuchungen [Sha94a, Kle14, Pet80]. Untold [Wil07]. Unusual [Déb96, Gin02c, Kok98, Ben93a]. Unveiled [Gin01d]. Updated [Tri03]. Updates [CS14]. Updating [Hen15b, Kre09, Low09]. upon [Kni70, Sal79]. Upper [SB05]. Uppsala [Rag16b]. Urania [JLGL14, LGJLF14]. Uraniborg [Lat93]. Uranographia [KPG03]. Uranometria [Gin12b, Hos89b, Lat14]. Uranus [Bou13, Sch81, Rea84]. Urbanism [RAGGB18]. Urkunden [Swe98]. Ursum [Jar06]. Ursus [GB86, Sha99b, Ano05a, Don88a, JLS^{+}05, JHLS05, Ser13, Gau17, Don88a]. US$25 [Bre18]. US$39.95 [Bec19]. USA [Liv13, Smi85, SS90]. Use [Dav17, Déb96, Kre81b, Mea71, Mer11, RPPA15, Spa04, Mar23, Pri91, Tur77, dB95, Kun88]. Used [Cha83, Das15, McC93, Swe86b]. Useful [LS05]. User [Ano13a, CG13]. User-Friendliness [Ano13a, CG13]. Users [Dew00]. Uses [TT72b]. Usher [CWW17]. Usherwood [POH96]. Using [Cha72, Duk02, Mol14b, Ste02a]. USSR [Ano15f, Ano80a, Ger14]. Utility [Ste90a]. uz [Kur19]. Uzití [Nor02].

V [Ano06a, Bre18, Hos83b, Kri88, Voe03, Ano05a, Gra05b, Hos89a, Hos91a, JLS^{+}05, JHLS05, Jar73, Rya71, Sch14a, Wer10, Bon07, DeV00b]. Vacuum [Eva82]. Valete [Hos14a]. Validity [HS81]. Valleriani [Cro23a]. Valley [Bel97]. Valls [Kra22]. Valls-Gabaud [Kra22]. Value [SS16, Ma23]. Values [Dav06, Gol03a, MS04, MS05, Moz13b, Moz13c, MD21]. Vancouver [Kre01]. Variable [Hos79a, Smi00, Smi12, Zso94, Zso20]. Variables [Hos07a]. Variations [Dre15, Van85]. Variorum [Bel14a]. Vasco [vdW19]. Vat. [Lip15]. Vaterland [Brü87]. Vatican [Ash86, Hos92a, Hos92a]. Vaticano [GS13, Lip15]. vault [LS24b]. ve [Sch14a]. vedecké [Sch14a]. vedeckých [Sch14a]. Vegas [Llo15]. Veil [Cov17]. veka [Rya71, Wer10]. Vekenstyl [Dob77]. Velazquez [vH86]. Velikovsky [Gol87a, Sha14, Gol87a]. Velocities [GC13]. Velocity [Fab19, Smi79]. Venerable [Pal12]. Veneto [AR86, Mos11]. Veneto-Friuli [AR86]. Vengeance [Van95b]. Venice [Hei12, HDG^{+}13, Wil15, Cha07]. Venu [Wło10]. Venus [Dic09, Dun13, Fei13, Gin05c, Gra13a, Gur06, Hei14b, Vai12, Yeo00, vH86, Ald02, Ald03, BS92, Bra20, Bra23, Bra05, Bri01a, BMA90, DOL98, Dic09, Dra84b, Dun13, Gin84c, Gol96b, GC99, Gra13a, Gur06, Hen15a, Mar95, McC83b, MS85, Moz19, NKM^{+}16, Pal01, Pet84, Sch01b, SD03, SM04, Spr93b, Spr93c, Ste99, Vai12, Wei82b, Wło10, Yeo00, dJ12, vH86, AH83, Bra05, Gin97c, Gra13a]. Venustafeln [Ste12a]. Vera [Tho22]. Verbiest [Gin96c, Has11, Nor91a, Siv05]. Verdensbilledets [Hol03]. Verdun [Fle92]. Verfertiger [Bre13]. Verhandlungen [Kel78]. Verification [CT93]. Verlag [Bur15, Hei14b, Joh19, Oes15, Wes97]. Verlagsanstalt [Ano15f, Kle14]. Vermeer [Mil04, Mil04]. Vermij [Hal03]. Vernacular [CMM^{+}15]. Vernet [Gin99b]. Vernuft [Van98b]. Veröffentlichungen [For73]. Veronica [Cov17]. Verrier [Lan84, Sel14, Gap15, Ros82b, Sel14]. Verse [Gan15]. Version [Pou96, Fat00]. Versions [Gol11]. Versus [McC99, Gin95, KD98, Smi98]. Very [Gib82, Ler17, McC83b, Wil04]. Verzeichnis [Gin99e]. Vesmir [Dob83a]. Vetus} [Not20, Kin02b, Pou84]. Vetustissimus [Kin02b]. VI [Kin81, Swe98, Ano15f, Boh19, SL13, Eva00]. VI/1 [Swe98]. VI/2 [Swe98]. Vía [Omo13, Dav09]. vicarious [CR22]. Vicente [vH86]. Vicissitudes [Cha90]. Victor [Car86, Wes91]. Victorian [Dew00, Dic09, Sch05b, Dic09, Pan94b, Pan95, War98]. Vie [CL05, Hat08]. Vienna [Sha85]. Viennese [Sha99a]. Viète [Car23, Car24, Swe75]. View [Ald97, Car86, DeV96, Ell90, Eva93a, Fei95, Gin00a, Gra09, Her80, Hom72, Lad10, Lau18, Sch82, Sch03, Spr93b, Spr93c, AB13, Sch15]. Viewed [Cor03]. Viewing [McC07]. Viewpoint [McC94]. Views [Dic90, Gur06]. Viginti [HH07]. VII [Bro10, Mes20, Mes21, Kru17]. viii [Bec15, Bec19, Bon19b, Dan19b, Hei14b, Jar90]. Vikings [Kru97]. Villalobos [Kru17]. Vimond [CG03]. Vincenzio [GM03]. Vincenzo [Mos11]. Vinci [RP87]. Virginia [Olo04, Olo04]. Virginio [Can12].

Virreinato [Ave87, Gli88]. Virtue [Bec00]. Visa [AH83]. Visibility [FSAD99, Hog88, LS05, Kin88]. Visible [Lüt00]. Vision [Ano84b, Ben87a, Ome79]. Visionaries [Tat09]. Visionary [Bec01]. Visione [Lin07]. Visionis [dJ12]. Visited [MHS17]. Visits [Ash03]. Visual [Lan83, Lüt00, Mea02]. Visualisierungen [Fei13, Hei14b]. Visualization [Wri07]. Visualizing [Hei14b, Rap11a]. Visuelle [Eas10]. Vita [Ser88]. Vitel [Ano15f, Mal10]. Vitesse [Cha85]. Viviani [Ree20]. Voice [Aub16, Smi94c]. Voices [Ger14, Ano15f]. Void [O’B88]. Voie [Hos04f, Wei82a]. Vol [Gin86b, Gre02, Hol77, Kno90a, Lüt00, Tuc00, Ano15f, Cro15, Gin87c, Han09a, Tho71c]. volgens [Van98b]. Volker [Mos07, Mos07, Spr92]. vols [Eva17, Wło23]. Volume [Ano14b, Ano15c, Ano16a, Ano17b, Ano18b, Ano19, Ano20b, Ano21a, Ano23, Ano24b, Gin91d, Nor88, PdV85, Tho90b, Wes90, Ano78c, Ano79d, Ano89d, Ano22c, Rag83]. Volumes [Ano79c, Ano89b, Ano20a, Hos83b]. Volvelles [Llo89b]. vom [Ano06a, Jor94]. vor [Swe90]. Vorgängerinnen [Ham09]. vorkolumbischen [Spr92]. Vorpommern [Eag08a, Eag08b]. Vorstellung [Voe02]. Vortex [Whi73b]. Vortheil [Ham09]. Vortices [Whi73b]. Vorzeit [Eva06]. Voyage [Hos04b]. Voyager [Doe05]. Vulcan [Dic99, Fon73]. Vulgar [Rud92]. Vulparius [Gin77a].

W [Ano81a, Gef75, Kes15, Mit19, Moo94, Ost89, Rus02, Sta06, Van14b, Wil15, Wil08a, Wil09, God13, Hil77, Prz79, Rav72]. W. [Ost89]. Wadi [BEBM02]. Walcher [Zep19]. Walery [KKSS21]. Wales [Cam11, Cam11, Pow95]. Wall [Kre20]. Wallingford [dSP78, Syl07, dSP78, Syl07]. Wallis [Ano95d]. Walter [DL15, HB92, Ost95, Ost96, Ost97b, Ost98, SK03].

Walther [dB70, Kre80b, Kre81b, SS98b]. Wandel [For82]. War [Ald97, Dew87, Hos89a, How05, Kid90, McC94, Sul02, Van95b, Ano80a]. Warburg [Tho90b]. Warfare [Eis88b]. warhafftige [Gin76]. Wars [Kra95, Sha14]. Warsaw [Ave95, Jas96b, Jas96b]. Was [Bou13, CJ20, Eis06, Gin05e, Gle07, GG14, Hos04g, Rab05, Sta13, WR08, Wło22]. Washington [Dic06, Doe19, Ger17, Moe77, Dob79, Dup90, Nee91, Pau95]. Wasn’t [Wło10]. Watch [Eva13]. Watches [Ges19, Gin85c]. Watching [Zei86, Ave97]. Water [Bon09a, Gin86b, Gin96d]. Water-Clocks [Gin86b]. Watermark [Dek08b]. Watkins [Bak15, GMML14]. Wave [Smi18b]. Way [Ber97a, Don73b, Lad10, Smi94a, UAGGB21, Hut13, Don73b, Hos75, Hos04f, Hos08g, Jak71, Jak72a, Lat14, Omo13, Dev14b, Whi00]. Wayne [Sil23]. Ways [Ber70, Gin90g, Hus12]. Wealth [Gin97d]. Weather [Bur20, Sid10]. Website [Ste11]. Wechselspiel [Dup08]. Wechselverhältnisses [Sta99]. Wednesday [She79a]. Weg [Sha89]. Wegbereiter [Dan11]. Weil [Wil75, Wil75]. Weisen [Gin81c]. Wellesley [Mor80]. Welt [Jor94, Spr92, Van90, Mul81]. Weltaneignung [Eas10]. Weltanschauung [For83]. Weltbild [Jor94, Sev00, Siv77, vG05]. Weltbildes [For82]. Welten [Har84]. Welteneinklangs [Dru84]. Welterklärung [Ano06a]. Weltraum [Hen09]. Wende [Har84]. Wendt [Mit19]. Were [Not24b]. Werk [Dob96, Hol77, Hor74, Sta00a]. Werke [Don92, Don00, Gin93, Hol77, Sha71b, Tho71c, Voe97, Wil86b, Gin14, Kin83, Kre19, Voe07]. West [Bur82, Cha09, TT76, Tur71b, Wal13, Gin98, Gin03a, GGCF06, Han09a, Hos02c, Ler11b, Pin90, Pou09, Rug94, Rug96, Tur71b, Gin98]. West-Central [Hos02c]. Western [Bon11, Ell85, Gin01d, Gin06b, God13, Han09a, Har72, Hat88, HMA^{+}98, Hos08h, Mol93, Pin90, Ric07, Rug04, Sal09, Sha87b, Whi02, BS06, Ros19, WS19, Van99]. Westfall [Gin97e]. Westman [Jar90]. Westview [Rob86]. Wetenschap [Van98b]. Wetenschappen [Hal03]. Whatever [Bec99a]. Wheelwright [Túr90b]. Where [Ave95, Cov17]. Whipple [Doe19, ML07, Llo90, Bur10]. Whisper [Ome79]. White [Hol09, Nau08]. Whiteside [Hos08d]. Whitney [Ost89]. Whizz [Llo18]. Whizz-Kiddery [Llo18]. Who [Ano15f, Déb05, Dra80, Smi94a, Smi94b, Tat09, Waf04, Wil07, Wil05b, Yeo02, Ger14]. Whole [Llo11]. Whose [Rug97]. Wide [Ost94]. Wide-Field [Ost94]. Widening [Hos71b]. Wider [RMH91]. Widmung [Ano06a]. wieku [Hil77, Rav72]. Wien [Sha99a]. Wiener [Sha85]. Wierciński [Ave95]. Wiesel [dC01]. Wikiana [Gin76]. Wild [Dra80, Mit19]. Wilding [Rap15a]. Wilhelm [DeV89, Gin90f, Hen09, Kre07b, Mos09, Oes15, RR13, Wil86a, Naz16, DeV89, Nor80]. Wilkins [Tur71a, Don71b, Tur71a, Don71b]. William [Apt14, Aub12, Ben13, Cro15, Hos13, Kra11, Too95, Apt14, Aub12, Bec00, Bec01, Bec13, Ben76b, Ben81b, Ben13, Ben20, Bol20, Bra20, DOL98, Hos79b, Hos84a, Hos04g, Hos05e, HD06, Hos08g, Hos10b, Hos11c, Hos11d, Hos12b, Hos12c, Hos14c, Hos14d, Hos14e, Hos15c, Hos16b, Mau13, MBH16, MH88, Mol14a, Plo93, Pum11, Sad90, Sch80, Sel06, Sel12, Smi83, SB84, Spa04, Ste14, Whi08, Bec15, Bed10, Cha15b]. Williams [Sch12b]. Williamson [Mul83]. Wilson [Dic06, DeV00b, Dic06, Gin13a, MH88, Ost96, Ste90b]. Win [Vet19]. Wind [Lep14]. Window [Ald02]. Windows [Fat00]. Winged [Kre12]. Winthrop [WDR11]. Wire [Vet12]. Wirkung [Dob96]. Wisconsin [Lat20]. Wisdom [War19]. Wissens [Fei13, Hei14b]. Wissenschaft [Kle14]. Wissenschaften [Oes15, RR13]. wissenschaftliche [Kre82, Ost05b]. wissenschaftlichen [Ano06a]. wissenschaftliches [Sha99a]. Wissenschaftsforschung [Kre82]. Wissenschaftsgeschichte [Kre82]. Wissenschaftslegitimierung [Ano06a]. Witch [How05]. Witelo [Sha02]. Within [Bla98a, Chr98, Dre15, GC08, Roc19]. without [Bec00, Far05, Gra80, Kra07b, Nor96]. Witness [AS97, Cha17]. Witnessing [Kol01b]. Wittenberg [Gin77c, OT13]. Wittich [Jar90]. Wolfenbüttel [Gau10, Gau10]. Wolfgang [Ano15f]. Woman [Aub16, Tri13, Wil07]. Women [Bec99a, Mac90, Mul20, Ogj11, Zru21]. Woodcuts [Lüt10]. Woodlawn [Plo93]. Woodruff [Sil23]. Woolf [Olo04, Olo04]. Words [Ano06a, Udí21]. Work [Apt99, Bol20, Bro88b, CM10, Chr98, Coh95, Cou18, FF20, Gal83, Gin82b, Hol10b, Loc00, Mer11, Naz16, Sha71b, Sul02, TT75, Túr90b, Sal87, Sal94b, Dob92]. Working [Ano12a, Kok98, Smi22]. Works [DeV85b, Don71b, Don79, Gin93, Gin08, Hos05a, Kno90a, Lat19, Pou96, Rag92, Ros74b, Sal90, Boh23, Gin93]. Workshops [Gin91c]. World [Ald09, Ano77l, Ano80a, App05, Bel09b, Bon16a, Bre17, Cam02, Car99, Eva11, Fek13, Gin85e, Gol11, Gra22, Has19, Hos78a, Hos04e, How86, How94, Hun18, Jar90, Kid90, Kol92, Kre89, Kru83, Kul75, Ler05, Naò16, Nee08, Ost92b, Pla91, Por09, Rag01, Ric07, Rut13, Sal09, Sch03, Sid10, Spr93b, Spr93c, Ste98b, Swe17c, Van95b, Van06, Wes93, Wil14b, Vet19, Zuj11, Dal01, Gin78c, Gin03g, Hal88, Hen08, Hos06a, Kru83, Mar12b, Mos10, Mül91, vdW90, Gau87]. World-Edifice [Hos78a]. World-maps [Rag01]. Worldly [Gin97d]. Worlds [Ave95, Bro83, Cro12, Don71a, EZ16, Far04, Fei90, Lau14, Nor00a, Par92, Rap11b, Rug00, Str99, G^{+}99, Wes89, Ave95, Gin06b, Han09a, O’B88]. Worldwide [Mor97]. Wren [Ben75, Fei84, Joh10, Fei84]. Wright [Hos70, Hos81, HR92, Hos03a, Jak72a, Kni70, Sch78]. Writing [Den00, HG80, Jar91a, Jon06b, Oss12]. Writings [Gol11, Gol12, Ost97a, Pou85, dSP78, Jus20, Rap13]. Written [Gin03e, Nor92a]. Wrongs [Sch99]. Wszystkie [Gin08]. Wu [Cul93]. Wunderzeichen [Gin76]. Wytze [Van98b].

X [Rea81, Smi85, Wis06, Apt14, Hal03, Hei14b, ML14, Nau98, Sha14, Tho71c, Pou88, Smi85]. X-ray [Smi85, Smi85]. Xaver [Brü87, Mes11]. Xe [Cha09]. XI [Wil86b, Hea15, Mes15, Qua17, Roc17]. xii [Bel14a, Bel14b, GG14, Kra22, LF15, Mar14, Nau98, Rob86, Sul14, Vet19, Gin99a]. XIII [Nor82, Voe03, Ben20, Dun20, Mir17, War19, SL13]. xiv [Ano15f, Ave18, Cro23a, Eva17, Jus20, Lep14, Swe17c, Wil15, Mol76]. XIX [Hol77, Wer10, Zep19]. xlviii [Kle14]. Xochicalco [FA97]. Xultun [ASS13]. XV [Hil77, Wło15]. XVI [Nor82, Hen15b, Mol14b, Oos15, Too95, Wes97, Gin07b]. XVIe [Ler11b, Gin88b]. XVII [GM03, Doe19, Gan15, McC19, Mit19, Rad90, Rot17]. XVIIe [Lüt00, Cha89]. XVIII [Kul87, Rav72, Rya71, Sil23, Gli88]. XVIth [Con70]. XVIth-Century [Con70]. XVth [Con70]. XVth- [Con70]. XX [Don92, Wer10, Smi96, Smi18a]. XX/1 [Don92]. XXI [Ait82]. xxii [CWW17, Cha15b, Cro15]. xxiv [Gee18]. xxv [Fin19, Llo15]. xxvi [Bak15]. xxviii [Mar14]. XXXII [Her76].

Yūnus [Kin88]. Yahyā [KF70]. Yale [Dic97a, Dic97a]. Yard [Por73]. Yarkovsky [Bee06]. Ydea [Cha19]. Year [Duk09, Eva09b, Gin03g, Swe86c, Whi76b, GB95, Oss24]. yearbook [Sho21]. Years [Ban10, Bla91a, DeV90, Dob99b, Edd90, Eis88a, Eva87a, Gin03a, Gin11a, Gle07, Hir85, Hol03, Hos92a, Hos15c, Mac15, Mar12b, McC19, Orc84, Pas14, Roc08, Sch91, Sil23, Ste90a, Whi76b, Aug24, GMW80, Lat23, Mal23, Sho21, ML14, Gin84b, Wis06, Yeo88]. Yemen [NRKN16]. Yerkes [Bla98b, Bla98b]. Yesterday [Kra07b, Moe77, Far05]. Yesteryear [Mol14b]. Yeux [CL08]. yi [Kin79]. yogatārās [PM89]. York [Ano15f, Apt14, Bok73, Bro90, Car86, CWW17, Hea15, LF15, Mal15, Mol14b, Qua17, Rem15, Vet19, Wło15, Hei12]. Yorkshire [Hos15c]. Young [Dob83a, Har79, Ost02]. Yourself [Gib82]. Yugoslavia [Mul84].

z [Nor02, Sch14a]. Zīj [Vil88]. zījat [Kin87]. Zach [Brü87, Brü87, Mes11, Mes11]. Zachodu [God13]. Zacut [Kre02, Gol98, GC99, Kre02]. Zanstra [Ost01]. Zarqālluh [SM94]. Zawiya [RPPA15]. Zdenek [Gin88f]. Zealand [Bat15, CKE11]. Zeichen [Rut12]. Zeilik [SS86]. Zeit [Ste12a]. Zeitalter [Joh19]. Zeiten [Kri97]. Zeitordnung [Spr92]. Zeitrechnung [War13]. Zemnom [Wer10]. Zentralbibliothek [Gin76]. Zeremonialzentren [Har73]. Zero [Déb87]. Zhizn’ [Ros74b]. Zhou [Siv98]. Zierde [Ham09]. Zij [Moz16, Van24]. Zijes [Sam98]. zijn [Nor91a]. Zinner [Wes97, Ano71b, Ano97b, Dob92, Wes97]. Ziółkowski [Ave95]. znacheniya [Rya71]. znanosti [Mul84]. Zodiac [Bur84, FJ11, Gin11c, Mer03, SG07, BN81, Kur21, Gra05a]. Zodiacal [War19]. Zodiacs [WZR23, Gin72]. zodiaque [BN81]. Zöllner [Sta00b]. Zone [Hoc92]. Zones [Goo09]. Zum [For81, Ano06a, Ano15f, DF14, Jor94, Mos07, Voe02]. Zur [Sha85, Too77, Ano06a, Dij09, Dob99a, For70a, Gin15, Har84, Kle14, Pet80, Sha94a, Swe76, Voe03, Whi72, Swe99a]. Zürich [Gin76]. Zvezdnoe [Sol09]. zwischen [Ano15f, Dup08, For81, Sch14b, Sev00, Sha99a, Sta99, Gin03f].

References

Aaboe:1972:RTT

[Aab72] Asger Aaboe. Remarks on the theoretical treatment of eclipses in antiquity. Journal for the History of Astronomy, 3(2):105–118, June 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Adair:2013:SBS

[AB13] Aaron Adair and Bob Berman.

Star of Bethlehem: a skeptical view. Onus Books, Fareham, Hampshire, UK, 2013. ISBN 0-9566948-6-1. 155 pp. LCCN ????

AldanayVillalobos:2014:AM

[AB14] Gerardo Aldana y Villalobos and Edwin Lawrence Barnhart, editors. Archaeoastronomy and the Maya. Oxbow Books, Oxford, UK, 2014. ISBN 1-78297-643-4 (paperback), 1-78297-644-2 (e-book). vii + 165 pp. LCCN F1435.3.C14 A73 2014.

Aveni:2003:SSO

[ABB03] Anthony F. Aveni, Harvey M.

Bricker, and Victoria R. Bricker. Seeking the sidereal: Observable planetary stations and the Ancient Maya record. Journal for the History of Astronomy, 34(2):145–161, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Acerbi:2020:HF

[Ace20]  Fabio Acerbi.  Hipparchus

in French. Journal for the History of Astronomy, 51(3): 376–377, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620930835.

Ackermann:2005:BRI

[Ack05] Silke Ackermann. Book review: Islamic mathematical instruments: Mathematical Instrumentation in Fourteenth-Century Egypt and Syria: The Illustrated Treatise of Najm al-Din al-Misri. Journal for the History of Astronomy, 36(3): 342–343, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ackermann:2010:BRE

[Ack10] Silke Ackermann. Book review: Eastern astrolabes at the Adler: Historic Scientific Instruments of the Adler Planetarium & Astronomy Museum, II: Eastern Astrolabes. Journal for the History of Astronomy, 41(2):276–278, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ackermann:2018:GCM

[Ack18] Silke Ackermann. From globes to celestial maps. Journal for the History of Astronomy, 49(2):264–265, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618757249.

Adamson:2021:AMA

[Ada21] Peter Adamson. Abu Ma’sar’s astrological classic in English. Journal for the History of Astronomy, 52(2):237– 239, May 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211004675.

Applebaum:1983:BMH

[AH83] Wilbur Applebaum and Robert A.

Hatch. Boulliau, Mercator, and Horrocks’s “Venus in Sole Visa”: Three unpublished letters. Journal for the History of Astronomy, 14(3):166–179, October 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1994:MIO

[AH94] Anthony F. Aveni and Lorren D. Hotaling. Monumental inscriptions and the observational basis of Maya planetary astronomy. Journal for the History of Astronomy, 25(19): S21–S54, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aiton:1982:BRE

[Ait82] E. J. Aiton. Book review: Early biographies of Copernicus: Starowolski’s Biographies of Copernicus (Studia Copernicana, XXI). Journal for the History of Astronomy, 13(1): 62–64, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aiton:1986:BRM

[Ait86] E. J. Aiton. Book review: Mysterium cosmographicum in French: Le Secret du Monde. Journal for the History of Astronomy, 17(3):203–205, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Abdumannapovich:2023:PSU

[AKJ^{+}23] Ehgamberdiev Shuhrat Abdumannapovich, Yuldoshev Qudratillo Khabibullaevich, Turaev Sobir Jurakobilovich, Jianhai Zhao, Meiting Yang, Zhenghong Tang, and Yong Yu. Present status of UBAI plate archive. Journal for the History of Astronomy, 54(4):456–468, November 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231207414.

Aldana:1997:BRV

[Ald97] Gerardo V. Aldana. Book review: a view of Andean culture: The Secret of the Incas:

Myth, Astronomy, and the War against Time. Journal for the History of Astronomy, 28(22): S88–S89, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aldana:2002:SSV

[Ald02] Gerardo Aldana. Solar stelae and a Venus window: Science and Royal personality in Late Classic Copán. Journal for the History of Astronomy, 33(27): S29–S50, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aldana:2003:KMV

[Ald03] Gerardo Aldana. K’uk’ulkan at Mayapán: Venus and postclassic Maya statecraft. Journal for the History of Astronomy, 34 (1):33–51, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aldana:2009:ERA

[Ald09] Gerardo Aldana. Essay review: Aveni honoured, Skywatching in the Ancient World: New Perspectives in Cultural Astronomy. Studies in Honor of Anthony F. Aveni. Journal for the History of Astronomy, 40(1):109–113, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aldana:2011:BRM

[Ald11] Gerardo V. Aldana. Book review: a Mayan day of doom: The End of Time: The Maya Mystery of 2012. Journal for the History of Astronomy, 42(1):111–113, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Allchin:1971:BRS

[All71] F. R. Allchin. Book review: Science in Sanskrit: Census of the Exact Sciences in Sanskrit. Journal for the History of Astronomy, 2(3):203–204, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Allen:1992:SDN

[All92] David A. Allen. Solstice determination at noon. Journal for the History of Astronomy, 23 (17):S21–S31, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1995:MNT

[AMP95] Anthony F. Aveni, Steven J. Morandi, and Polly A. Peterson. The Maya number of time: Intervalic time reckoning in the Maya codices, part 1. Journal for the History of Astronomy, 26(20):S1–S28, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1996:MNT

[AMP96] Anthony F. Aveni, Steven J.

Morandi, and Polly A. Peterson. The Maya number of time: Intervalic time reckoning in the Maya Codices, Part II. Journal for the History of Astronomy, 27(21):S1–S32, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Andrewes:1990:BRM

[And90] William Andrewes. Book review: Measurement of angles, the divided circle: A History of Instruments for Astronomy, Navigation and Surveying. Journal for the History of Astronomy, 21(2):224–226, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Andrewes:2000:BRC

[And00] William J. H. Andrewes. Book review: Creating the National Maritime Museum: Of Ships and Stars: Maritime Heritage and the Founding of the National Maritime Museum, Greenwich. Journal for the History of Astronomy, 31(2):173– 175, May 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Andrews:2017:TIK

[And17] Noam Andrews. Tabula III:

Kepler’s mysterious polyhedral model. Journal for the History of Astronomy, 48(3):281–311, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617721545.

Andrews:2021:GKC

[And21] Noam Andrews. Gilding Kepler’s cosmology. Journal for the History of Astronomy, 52(1):3–32, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620987320.

Anonymous:1970:EF

[Ano70a] Anonymous. Editorial foreword. Journal for the History of Astronomy, 1(1):4, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:PHA

[Ano70b] Anonymous. From the President of the History of Astronomy Commission of the International Astronomical Union. Journal for the History of Astronomy, 1(1):3, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:NCa

[Ano70c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 1(1):93–94, February 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:NCb

[Ano70d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 1(2):180, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:NBa

[Ano70e] Anonymous. Notices of books.

Journal for the History of Astronomy, 1(1):89–90, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:NBb

[Ano70f] Anonymous. Notices of books.

Journal for the History of Astronomy, 1(2):164, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:TDH

[Ano70g] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 1(2):167–168, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1970:ULT

[Ano70h] Anonymous. University of London theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 1(1):91–92, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:EF

[Ano71a] Anonymous. Editorial foreword. Journal for the History of Astronomy, 2(1):iv, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:EZ

[Ano71b]  Anonymous.  Ernst Zinner

(1886–1970). Journal for the History of Astronomy, 2(2): 132, June 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:FMI

[Ano71c] Anonymous. Future meetings of interest to historians of astronomy. Journal for the History of Astronomy, 2(1):56–57, February 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:GRS

[Ano71d] Anonymous. Greenwich register of source material on the history of astronomy. Journal for the History of Astronomy, 2(2):132, June 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:MCI

[Ano71e] Anonymous. Meetings of Commission 41 of IAU. Journal for the History of Astronomy, 2(1):55, February 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:NCa

[Ano71f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 2(1):72, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:NCb

[Ano71g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 2(2):146, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:NCc

[Ano71h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 2(3):219, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:NBa

[Ano71i] Anonymous. Notices of books.

Journal for the History of Astronomy, 2(1):50, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:NBb

[Ano71j] Anonymous. Notices of books.

Journal for the History of Astronomy, 2(2):130–131, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1971:TDH

[Ano71k] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 2(1):71, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:EF

[Ano72a] Anonymous. Editorial foreword. Journal for the History of Astronomy, 3(1):iv, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:ICB

[Ano72b] Anonymous. Isis cumulative bibliography. Journal for the History of Astronomy, 3(2): 149–150, June 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:NCa

[Ano72c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 3(1):74, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:NCb

[Ano72d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 3(2):150, June 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:NCc

[Ano72e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 3(3):221, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:NBa

[Ano72f] Anonymous. Notices of books.

Journal for the History of Astronomy, 3(1):72, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:NBb

[Ano72g] Anonymous. Notices of books.

Journal for the History of Astronomy, 3(2):148, June 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:TDHa

[Ano72h] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 3(1):73, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1972:TDHb

[Ano72i] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 3(2):149, June 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:APC

[Ano73a] Anonymous. Archaeoastronomy in pre-Columbian America. Journal for the History of Astronomy, 4(3):214, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:AAA

[Ano73b]  Anonymous.  Astronomy

archives in Australia (1). Journal for the History of Astronomy, 4(1):66–68, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:DT

[Ano73c] Anonymous. Doctoral theses. Journal for the History of Astronomy, 4(2):143, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:EF

[Ano73d] Anonymous. Editorial foreword. Journal for the History of Astronomy, 4(1):iv, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:ERC

[Ano73e] Anonymous. Essay review: C. Doris Hellman (1910–1973). Journal for the History of Astronomy, 4(2):136, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:HRP

[Ano73f] Anonymous. Historical research with planetarium. Journal for the History of Astronomy, 4(3):191, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:NCa

[Ano73g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 4(1):72, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:NCb

[Ano73h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 4(2):144, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:NCc

[Ano73i] Anonymous. Notes on contributors. Journal for the History of Astronomy, 4(3):215, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:NBa

[Ano73j] Anonymous. Notices of books.

Journal for the History of Astronomy, 4(2):143, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1973:NBb

[Ano73k] Anonymous. Notices of books. Journal for the History of Astronomy, 4(3):210, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:CCA

[Ano74a] Anonymous. A computerized checklist of astrolabes. Journal for the History of Astronomy, 5(2):143, June 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:C

[Ano74b]  Anonymous.  Copernicana.

Journal for the History of Astronomy, 5(3):205–207, October 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:EF

[Ano74c] Anonymous. Editorial foreword. Journal for the History of Astronomy, 5(1):iv, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:JSM

[Ano74d] Anonymous. James Stewart Mackenzie (1719–1800) and the Bute Mss. Journal for the History of Astronomy, 5(3): 208–214, October 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:NCa

[Ano74e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 5(1):70, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:NCb

[Ano74f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 5(2):144, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:NCc

[Ano74g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 5(3):214, October 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1974:NB

[Ano74h] Anonymous. Notices of books.

Journal for the History of Astronomy, 5(2):141–142, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:ADH

[Ano75a] Anonymous. American dissertations in history of astronomy. Journal for the History of Astronomy, 6(2):147, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:EF

[Ano75b] Anonymous. Editorial foreword. Journal for the History of Astronomy, 6(1):i, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:ERS

[Ano75c] Anonymous. Essay review: Science in India: A Concise History of Science in India. Journal for the History of Astronomy, 6(2):135–137, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NCa

[Ano75d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 6(1):74, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NCb

[Ano75e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 6(2):148, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NCc

[Ano75f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 6(3):220, October 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NBa

[Ano75g] Anonymous. Notices of books.

Journal for the History of Astronomy, 6(1):71, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NBb

[Ano75h] Anonymous. Notices of books.

Journal for the History of Astronomy, 6(2):142, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1975:NBc

[Ano75i] Anonymous. Notices of books.

Journal for the History of Astronomy, 6(3):218, October 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:AA

[Ano76a] Anonymous. Angus Armitage (1902–1976). Journal for the History of Astronomy, 7(3): 224, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:BRS

[Ano76b] Anonymous. Brief review: Science and society, past, present, and future. Journal for the History of Astronomy, 7(1):70, February 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:BRV

[Ano76c] Anonymous. Brief review: Van Marum’s scientific instruments in Teyler’s Museum, Part II, descriptive catalogue. Journal for the History of Astronomy, 7(1):70, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:CG

[Ano76d] Anonymous. Commission 41 at Grenoble. Journal for the History of Astronomy, 7(3): 216–224, October 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See remarks [New77].

Anonymous:1976:NCa

[Ano76e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 7(1):74, February 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:NCb

[Ano76f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 7(2):151, June 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:NCc

[Ano76g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 7(3):226, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:NBa

[Ano76h] Anonymous. Notices of books.

Journal for the History of Astronomy, 7(1):71–72, February 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:NBb

[Ano76i] Anonymous. Notices of books.

Journal for the History of Astronomy, 7(3):225, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1976:PC

[Ano76j] Anonymous. Per Collinder (1890–1975). Journal for the History of Astronomy, 7(3): 224, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:CER

[Ano77a]  Anonymous.  Corrigendum:

[Essay review: Kepler and his times]. Journal for the History of Astronomy, 8(2):149, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Hol77].

Anonymous:1977:E

[Ano77b] Anonymous. Editorial. Journal for the History of Astronomy, 8 (2):iii, June 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:IC

[Ano77c] Anonymous. Instructions to contributors. Journal for the History of Astronomy, 8(2):iv, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:NCa

[Ano77d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 8(1):75, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:NCb

[Ano77e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 8(2):150, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:NCc

[Ano77f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 8(3):223– 224, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:NBa

[Ano77g] Anonymous. Notices of books. Journal for the History of Astronomy, 8(1):69–70, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:NBb

[Ano77h] Anonymous. Notices of books.

Journal for the History of Astronomy, 8(2):148–149, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:SIA

[Ano77i] Anonymous. Société Internationale de L’Astrolabe. Journal for the History of Astronomy, 8(1):76, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:SSI

[Ano77j] Anonymous. Symposium on scientific instruments. Journal for the History of Astronomy, 8(1):73, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:TDH

[Ano77k] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 8(1):73, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1977:WLA

[Ano77l] Anonymous. World list of astronomical observatories, instruments and clocks 1670– 1850. Journal for the History of Astronomy, 8(1):74, February 1977. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:BS

[Ano78a] Anonymous. Bernhard Sticker (1906–1977). Journal for the History of Astronomy, 9(2): 153, June 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:EMM

[Ano78b] Anonymous. Erratum: [Meteors, meteor showers and meteorites in the Middle Ages: From European medieval sources]. Journal for the History of Astronomy, 9(3):230, October 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Dal78].

Anonymous:1978:IV

[Ano78c] Anonymous. Index to volume 9. Journal for the History of Astronomy, 9(3):228–230, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:ISC

[Ano78d]  Anonymous.  International

Symposium on Christiaan Huygens (1629–1695), Amsterdam, August 1979. Journal for the History of Astronomy, 9(1):75, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NCa

[Ano78e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 9(1):76, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NCb

[Ano78f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 9(2):154, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NCc

[Ano78g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 9(3):227, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NBa

[Ano78h] Anonymous. Notices of books.

Journal for the History of Astronomy, 9(1):72–73, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NBb

[Ano78i] Anonymous. Notices of books.

Journal for the History of Astronomy, 9(2):151–152, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1978:NBc

[Ano78j] Anonymous. Notices of books. Journal for the History of Astronomy, 9(3):226, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:AS

[Ano79a] Anonymous. The archaeoastronomy supplements. Journal for the History of Astronomy, 10(1):iv, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:CMI

[Ano79b] Anonymous. Commission 41 at the Montreal International Astronomical Union Congress. Journal for the History of Astronomy, 10(3):189–192, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:CIV

[Ano79c] Anonymous. Cumulative indices to volumes 1–10. Journal for the History of Astronomy, 10(3):195–206, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:IV

[Ano79d] Anonymous. Index to volume 10. Journal for the History of Astronomy, 10(3):192–194, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NCa

[Ano79e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 10(1): 73–74, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NCb

[Ano79f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 10(2):132, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NCc

[Ano79g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 10(3):192, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NBa

[Ano79h] Anonymous. Notices of books.

Journal for the History of Astronomy, 10(1):70–72, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NBb

[Ano79i] Anonymous. Notices of books.

Journal for the History of Astronomy, 10(2):131, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1979:NBc

[Ano79j] Anonymous. Notices of books. Journal for the History of Astronomy, 10(3):188–189, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:BDH

[Ano80a] Anonymous. Bibliography of dissertations on the history of astronomy in the USSR after the Second World War. Journal for the History of Astronomy, 11(3):221–224, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:BR

[Ano80b] Anonymous. Books received. Journal for the History of Astronomy, 11(2):144, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:IC

[Ano80c] Anonymous. Instructions to contributors. Journal for the History of Astronomy, 11(1):i, February 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:JA

[Ano80d] Anonymous. Joseph Ashbrook (1918–1980). Journal for the History of Astronomy, 11(3): 201, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:NAT

[Ano80e] Anonymous. North American theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 11(2):148–151, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:NCa

[Ano80f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 11(1): 75–76, February 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:NCb

[Ano80g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 11(2):151–152, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:NCc

[Ano80h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 11(3):225–226, October 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:NB

[Ano80i] Anonymous. Notices of books.

Journal for the History of Astronomy, 11(2):142–143, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1980:UD

[Ano80j] Anonymous. Umberto Dall’Olmo (1925–1980). Journal for the History of Astronomy, 11(2): 145, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bunge:1981:BRR

[Ano81a] Anonymous. Book review: M. Bunge and W. R. Shea, Rutherford and Physics at the Turn of the Century. Journal for the History of Astronomy, 12:75, January 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https:

                                        //ui.adsabs.harvard.edu/#
                                        abs/1981JHA....12...75S.

Anonymous:1981:BR

[Ano81b] Anonymous. Books received. Journal for the History of Astronomy, 12(1):75, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NCa

[Ano81c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 12(1): 75–76, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NCb

[Ano81d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 12(2):154, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NCc

[Ano81e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 12(3):229–230, October 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NBa

[Ano81f] Anonymous. Notices of books.

Journal for the History of Astronomy, 12(1):73–75, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NBb

[Ano81g] Anonymous. Notices of books.

Journal for the History of Astronomy, 12(2):153, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1981:NBc

[Ano81h] Anonymous. Notices of books.

Journal for the History of Astronomy, 12(3):214–215, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:AGO

[Ano82a]  Anonymous.  Archives of

Greenwich Observatory. Journal for the History of Astronomy, 13(2):153, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:BRa

[Ano82b] Anonymous. Book received. Journal for the History of Astronomy, 13(2):145, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:BRb

[Ano82c] Anonymous. Books received. Journal for the History of Astronomy, 13(3):226, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:MLR

[Ano82d] Anonymous. Maria Luisa Righini Bonelli (1917–1981). Journal for the History of Astronomy, 13(1):74, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NS

[Ano82e] Anonymous. Newham and Stonehenge. Journal for the History of Astronomy, 13(4): S73–S74, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NCa

[Ano82f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 13(1): 75–76, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NCb

[Ano82g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 13(2):153–154, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NCc

[Ano82h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 13(4):S75, February 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NBa

[Ano82i] Anonymous. Notices of books.

Journal for the History of Astronomy, 13(1):73, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NBb

[Ano82j] Anonymous. Notices of books.

Journal for the History of Astronomy, 13(2):143–145, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:NBc

[Ano82k] Anonymous. Notices of books.

Journal for the History of Astronomy, 13(4):S72, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1982:TDH

[Ano82l] Anonymous. Theses and dissertations on the history of astronomy. Journal for the History of Astronomy, 13(1): 73–74, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1983:BR

[Ano83a] Anonymous. Books received. Journal for the History of Astronomy, 14(1):69, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1983:CSC

[Ano83b]  Anonymous.  Corrigenda:

[“The significance of the Copernican ORBS” and “L’horloge planétaire de la cathédrale de Strasbourg”]. Journal for the History of Astronomy, 14(2): 153, June 1983. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Jar82, Pou83].

Anonymous:1983:EHA

[Ano83c] Anonymous. Ejnar Hertzsprung Archive. Journal for the History of Astronomy, 14(2):101, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1983:NCa

[Ano83d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 14(2):152–153, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1983:NCb

[Ano83e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 14(5):S75–S76, February 1983. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:BREb

[Ano84a] Anonymous. Book review: The 1979 Einstein centennial: Albert Einstein: Historical and Cultural Perspectives. Journal for the History of Astronomy, 15(3):222, October 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:BREa

[Ano84b] Anonymous. Book review: The 1979 Einstein centennial: Some Strangeness in the Proportion: a Centennial Celebration to Celebrate the Achievements of Albert Einstein, to Fulfill a Vision: Jerusalem Einstein Centennial Symposium on Gauge Theories and Unification of Physical Forces. Journal for the History of Astronomy, 15 (3):222, October 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:NC

[Ano84c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 15(1): 75–76, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:NBa

[Ano84d] Anonymous. Notices of books. Journal for the History of Astronomy, 15(1):61–63, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:NBb

[Ano84e] Anonymous. Notices of books. Journal for the History of Astronomy, 15(2):145, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:NBc

[Ano84f] Anonymous. Notices of books. Journal for the History of Astronomy, 15(3):223, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1984:NBd

[Ano84g] Anonymous. Notices of books. Journal for the History of Astronomy, 15(7):S152, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1985:NBa

[Ano85a] Anonymous. Notices of books. Journal for the History of Astronomy, 16(1):73, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1985:NBb

[Ano85b] Anonymous. Notices of books.

Journal for the History of Astronomy, 16(2):152–153, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1986:NCa

[Ano86a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 17(2):153–154, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1986:NCb

[Ano86b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 17(3):213– 214, August 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NCa

[Ano87a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 18(1): 76–77, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NCb

[Ano87b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 18(2):153–154, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NCc

[Ano87c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 18(3):232, August 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NCd

[Ano87d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 18(4):308, November 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NBa

[Ano87e] Anonymous. Notices of books. Journal for the History of Astronomy, 18(1):75, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NBb

[Ano87f] Anonymous. Notices of books. Journal for the History of Astronomy, 18(2):150–152, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1987:NBc

[Ano87g] Anonymous. Notices of books. Journal for the History of Astronomy, 18(3):230–231, August 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1988:NCa

[Ano88a] Anonymous. Notes on contributor. Journal for the History of Astronomy, 19(1):75– 76, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1988:NCb

[Ano88b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 19(2):154, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1988:NCc

[Ano88c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 19(3):216, August 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1988:NCd

[Ano88d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 19(4):291, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1988:NB

[Ano88e] Anonymous. Notices of books.

Journal for the History of Astronomy, 19(2):143–145, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:CBG

[Ano89a] Anonymous. Commission 41 at the Baltimore General Assembly of the IAU. Journal for the History of Astronomy, 20 (1):74–75, February 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:CIV

[Ano89b] Anonymous. Cumulative indices to volumes 1–20. Journal for the History of Astronomy, 20(3):229–250, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:EP

[Ano89c] Anonymous. Editorial preface. Journal for the History of Astronomy, 20(13):i, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:IV

[Ano89d] Anonymous. Index to volume 20. Journal for the History of Astronomy, 20(3):226–228, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:IC

[Ano89e] Anonymous. Instructions to contributors. Journal for the History of Astronomy, 20(1):i, February 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:NCa

[Ano89f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 20(1):76, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:NCb

[Ano89g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 20(2):155–156, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:NCc

[Ano89h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 20(3):225–226, October 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:NBa

[Ano89i] Anonymous. Notices of books.

Journal for the History of Astronomy, 20(1):70–73, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1989:NBb

[Ano89j] Anonymous. Notices of books.

Journal for the History of Astronomy, 20(2):153–154, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:EP

[Ano90a] Anonymous. Editorial preface. Journal for the History of Astronomy, 21(15):i, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NCa

[Ano90b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 21(1):155–156, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NCb

[Ano90c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 21(2):233–234, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NCc

[Ano90d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 21(3):311– 312, August 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NCd

[Ano90e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 21(4):385–386, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NBa

[Ano90f] Anonymous. Notices of books.

Journal for the History of Astronomy, 21(2):231–232, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:NBb

[Ano90g]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 21(4):382–384, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1990:RGO

[Ano90h] Anonymous. Royal Greenwich Observatory archives. Journal for the History of Astronomy, 21(4):352, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:EP

[Ano91a] Anonymous. Editorial preface. Journal for the History of Astronomy, 22(16):i, February 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:EJA

[Ano91b] Anonymous. Eric J. Aiton (1920–1991). Journal for the History of Astronomy, 22(3): 254, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:IC

[Ano91c] Anonymous. Instructions to contributors. Journal for the History of Astronomy, 22(1):i, February 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:NCa

[Ano91d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 22(1):100, February 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:NCb

[Ano91e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 22(2):194, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:NCc

[Ano91f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 22(3):255– 256, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:NCd

[Ano91g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 22(4):331–332, November 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1991:NB

[Ano91h] Anonymous. Notices of books.

Journal for the History of Astronomy, 22(2):193, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:EP

[Ano92a] Anonymous. Editorial preface. Journal for the History of Astronomy, 23(17):i, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NCa

[Ano92b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 23(1):76, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NCb

[Ano92c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 23(2):153–154, May 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NCc

[Ano92d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 23(3):231– 232, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NCd

[Ano92e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 23(4):307–308, November 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NBa

[Ano92f] Anonymous. Notices of books. Journal for the History of Astronomy, 23(2):152–153, May 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:NBb

[Ano92g] Anonymous. Notices of books. Journal for the History of Astronomy, 23(3):230, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1992:SRA

[Ano92h] Anonymous. Summaries of relevant articles in JHA. Journal for the History of Astronomy, 23(17):S73–S74, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:EP

[Ano93a] Anonymous. Editorial preface. Journal for the History of Astronomy, 24(18):i, February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:NCa

[Ano93b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 24(1–2): 155–156, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/155.full.pdf+html.

Anonymous:1993:NCb

[Ano93c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 24(3):234, August 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:NCc

[Ano93d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 24(4):309–310, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:NCd

[Ano93e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 24(18):S75–S76, February 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:NBa

[Ano93f]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 24(1–2): 154, May/February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/24/1-
                                        2/154.full.pdf+html.

Anonymous:1993:NBb

[Ano93g] Anonymous. Notices of books. Journal for the History of Astronomy, 24(3):230–233, August 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:NBc

[Ano93h]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 24(4):307–309, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1993:SRA

[Ano93i] Anonymous. Summaries of relevant articles in JHA. Journal for the History of Astronomy, 24(18):S75, February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:CRO

[Ano94a] Anonymous. Conference report: Oxford 4. Journal for the History of Astronomy, 25 (19):S91, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:EP

[Ano94b] Anonymous. Editorial preface. Journal for the History of Astronomy, 25(19):i, February 1994. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:ESA

[Ano94c] Anonymous. The European Society for Astronomy in Culture (SEAC). Journal for the History of Astronomy, 25(19):S90, February 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:NCa

[Ano94d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 25(1): 75–76, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:NCb

[Ano94e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 25(2):153–154, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:NCc

[Ano94f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 25(3):248, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:NB

[Ano94g] Anonymous. Notices of books.

Journal for the History of Astronomy, 25(3):246–247, August 1994. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1994:SRA

[Ano94h] Anonymous. Summaries of relevant articles in JHA. Journal for the History of Astronomy, 25(19):S90, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:CAH

[Ano95a]  Anonymous.  Corrigenda:

[Aveni and Hotaling in Archaeoastronomy, no. 19, table 1]. Journal for the History of Astronomy, 26(20):S79, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:CTS

[Ano95b]  Anonymous.  Corrigendum:

[Three solar longitudes in the Almagest due to Hipparchus]. Journal for the History of Astronomy, 26(3):274, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Thu95].

Anonymous:1995:EP

[Ano95c] Anonymous. Editorial preface. Journal for the History of Astronomy, 26(20):i, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:HW

[Ano95d]  Anonymous.  Helen Wallis

(1924–1995). Journal for the History of Astronomy, 26(3): 276, August 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:NC

[Ano95e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 26(20):S91, February 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:NB

[Ano95f] Anonymous. Notices of books.

Journal for the History of Astronomy, 26(3):276, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:OCI

[Ano95g] Anonymous. The Organizing Committee of IAU Commission 41. Journal for the History of Astronomy, 26(2):184, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:RA

[Ano95h] Anonymous. Richard Atkinson (1920–1994). Journal for the History of Astronomy, 26 (20):S90, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1995:SRA

[Ano95i] Anonymous. Summaries of relevant articles in JHA. Journal for the History of Astronomy, 26(20):S90, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1996:CRO

[Ano96a] Anonymous. Conference report: Oxford 5. Journal for the History of Astronomy, 27 (21):S80, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1996:EP

[Ano96b] Anonymous. Editorial preface. Journal for the History of Astronomy, 27(21):i, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1996:NC

[Ano96c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 27(21):S90–S91, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1996:NB

[Ano96d] Anonymous. Notices of books. Journal for the History of Astronomy, 27(4):370, November 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1996:SRA

[Ano96e] Anonymous. Summary of a relevant article in JHA. Journal for the History of Astronomy, 27(21):S90, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:EP

[Ano97a] Anonymous. Editorial preface. Journal for the History of Astronomy, 28(22):i, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:EER

[Ano97b] Anonymous. Errata: [Essay review: Zinner, Copernicus, and the Nazis]. Journal for the History of Astronomy, 28(4):369, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Wes97].

Anonymous:1997:EDA

[Ano97c] Anonymous. Errata: [The development of astronomy in the Southern United States, 1840– 1914]. Journal for the History of Astronomy, 28(3):258, August 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Wil96].

Anonymous:1997:NCa

[Ano97d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 28(3):278, August 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:NCb

[Ano97e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 28(4):367–368, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:NCc

[Ano97f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 28(22):S90–S91, February 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:NBa

[Ano97g] Anonymous. Notices of books. Journal for the History of Astronomy, 28(2):182–183, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:NBb

[Ano97h]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 28(4):362–364, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1997:SRA

[Ano97i] Anonymous. Summary of a relevant article in JHA. Journal for the History of Astronomy, 28(22):S90, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:DH

[Ano98a] Anonymous. Derek Howse (1919–1998). Journal for the History of Astronomy, 29(4):

398, November 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NCa

[Ano98b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 29(1): 91–92, February 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NCb

[Ano98c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 29(2):209–210, May 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NCc

[Ano98d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 29(3):307– 308, August 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NCd

[Ano98e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 29(4):398–399, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NCe

[Ano98f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 29(23):–, February 1998. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:NB

[Ano98g]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 29(4):396–397, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1998:SRA

[Ano98h] Anonymous. Summary of a relevant article in JHA. Journal for the History of Astronomy, 29(23):S88, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NCa

[Ano99a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 30(1): 91–92, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NCb

[Ano99b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 30(2):185–186, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NCc

[Ano99c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 30(3):340– 342, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NCd

[Ano99d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 30(4):410–411, November 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NCe

[Ano99e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 30(24):S91–S92, February 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NBa

[Ano99f] Anonymous. Notices of books.

Journal for the History of Astronomy, 30(1):89–91, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NBb

[Ano99g] Anonymous. Notices of books.

Journal for the History of Astronomy, 30(2):184–185, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:1999:NBc

[Ano99h] Anonymous. Notices of books.

Journal for the History of Astronomy, 30(3):339, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NCa

[Ano00a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 31(1): 91–92, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NCb

[Ano00b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 31(2):183–184, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NCc

[Ano00c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 31(3):277– 278, August 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NCd

[Ano00d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 31(4):367–369, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NCe

[Ano00e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 31(25):S91–S92, February 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NBa

[Ano00f] Anonymous. Notices of books.

Journal for the History of Astronomy, 31(1):87–90, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NBb

[Ano00g] Anonymous. Notices of books.

Journal for the History of Astronomy, 31(2):182, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2000:NBc

[Ano00h]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 31(4):362–364, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:AH

[Ano01a] Anonymous. Astronomy and its histories. Science History Publications Ltd., Cambridge, UK, 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). i–ii + 187– 280 pp. Papers from the Special Session to mark the retirement of Owen Gingerich held at the Annual Meeting of the History of Science Society in Vancouver, BC, November 2000, J. Hist. Astronom. 32 (2001), no. 3.

Anonymous:2001:CLO

[Ano01b]  Anonymous.  Corrigendum:

[The latitude of the observer of the Almagest star catalogue]. Journal for the History of Astronomy, 32(2):151, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Sch01a].

Anonymous:2001:EAA

[Ano01c] Anonymous. Errata: [Anthony Aveni, “Astronomy in the Mexican Codex Borgia”, Archaeoastronomy, no. 24 (1999), Table 2]. Journal for the History of Astronomy, 32(26):S90, February 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NCa

[Ano01d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 32(1):92, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NCb

[Ano01e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 32(2):185–186, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NCc

[Ano01f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 32(3):279– 280, August 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NCd

[Ano01g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 32(4):371, November 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NCe

[Ano01h] Anonymous. Notes on contributors. Journal for the History of Astronomy, 32(26):S91–S92, February 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:NB

[Ano01i] Anonymous. Notices of books.

Journal for the History of Astronomy, 32(1):90–91, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2001:ORL

[Ano01j] Anonymous. [Obituary:] Roger Louis Billard. Journal for the History of Astronomy, 32(2): 151, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:IC

[Ano02a] Anonymous. Instructions to contributors. Journal for the History of Astronomy, 33(1):i, February 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NCa

[Ano02b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 33(1): 91–92, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NCb

[Ano02c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 33(2):218, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NCc

[Ano02d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 33(3):310– 312, August 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NCd

[Ano02e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 33(4):402–403, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NCe

[Ano02f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 33(27):S83–S84, February 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NBa

[Ano02g] Anonymous. Notices of books.

Journal for the History of Astronomy, 33(1):88–91, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NBb

[Ano02h] Anonymous. Notices of books.

Journal for the History of Astronomy, 33(2):213, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NBc

[Ano02i] Anonymous. Notices of books.

Journal for the History of Astronomy, 33(3):309–310, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2002:NBd

[Ano02j]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 33(4):401–402, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:ERE

[Ano03a] Anonymous. Essay review:

East Asian observations: East-Asian Archaeoastronomy: Historical Records of Astronomical Observations of China, Japan and Korea. Journal for the History of Astronomy, 34(1): 111–114, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NCa

[Ano03b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 34(1):123–124, February 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NCb

[Ano03c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 34(2):249–250, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NCc

[Ano03d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 34(3):343– 344, August 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NCd

[Ano03e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 34(4):465–466, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NBa

[Ano03f]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 34(1):121–123, February 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NBb

[Ano03g] Anonymous. Notices of books.

Journal for the History of Astronomy, 34(2):248, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2003:NBc

[Ano03h] Anonymous. Notices of books.

Journal for the History of Astronomy, 34(3):341–342, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:EGH

[Ano04a] Anonymous. Erratum: [Galileo’s horoscopes]. Journal for the History of Astronomy, 35(3): 374, August 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Swe04c].

Anonymous:2004:EOS

[Ano04b] Anonymous. Erratum: [On the orientations of Sardinian Nuraghes: Some clues to their interpretation]. Journal for the History of Astronomy, 35(2): 236, May 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [ZB04].

Anonymous:2004:JD

[Ano04c] Anonymous. Jerzy Dobrzycki.

Journal for the History of Astronomy, 35(2):248, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NCa

[Ano04d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 35(1):121–122, February 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NCb

[Ano04e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 35(2):247–248, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NCc

[Ano04f] Anonymous. Notes on contributors. Journal for the History of Astronomy, 35(3):373– 374, August 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NCd

[Ano04g] Anonymous. Notes on contributors. Journal for the History of Astronomy, 35(4):496–497, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NBa

[Ano04h] Anonymous. Notices of books. Journal for the History of Astronomy, 35(2):246, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NBb

[Ano04i] Anonymous. Notices of books.

Journal for the History of Astronomy, 35(3):370, August 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2004:NBc

[Ano04j]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 35(4):494–495, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2005:ETV

[Ano05a] Anonymous. Errata: [Tycho v. Ursus: The build-up to a trial, Part 2]. Journal for the History of Astronomy, 36 (3):358, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [JHLS05].

Anonymous:2005:NC

[Ano05b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 36(1):122–124, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2005:NBa

[Ano05c]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 36(1):121–122, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2005:NBb

[Ano05d] Anonymous. Notices of books. Journal for the History of Astronomy, 36(2):246–247, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2005:NBc

[Ano05e]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 36(4):462–463, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2006:ERI

[Ano06a] Anonymous. Essay review: Imagineering the astronomical revolution: I. Transmitting Knowledge: Words, Images, and Instruments in Early Modern Europe, II. Widmung, Welterklärung und Wissenschaftslegitimierung: Titelbilder und ihre Funktionen in der wissenschaftlichen Revolution, III. The Power of Images in Early Modern Science, IV. Immagini per conoscere: Dal Rinascimento alla rivoluzione scientifica, V. Erkenntnis Erfindung Konstruktion: Studien zur Bildgeschichte von Naturwissen-schaften und Technik vom 16. bis zum 19. Jahrhundert. Journal for the History of Astronomy, 37(4): 471–484, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2006:NBa

[Ano06b] Anonymous. Notices of books.

Journal for the History of Astronomy, 37(2):248–249, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2006:NBb

[Ano06c] Anonymous. Notices of books.

Journal for the History of Astronomy, 37(3):369, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:AHA

[Ano07a] Anonymous. Asger Hartvig Aaboe (1922–2007). Journal for the History of Astronomy, 38(2):261–263, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NCa

[Ano07b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 38(1):138–139, February 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NCb

[Ano07c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 38(2):264–266, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NCc

[Ano07d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 38(3):391– 392, August 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NCd

[Ano07e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 38(4):528–530, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NBa

[Ano07f]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 38(1):136–138, February 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:NBb

[Ano07g] Anonymous. Notices of books. Journal for the History of Astronomy, 38(2):258–260, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2007:RRE

[Ano07h] Anonymous. Retirement of reviews Editor. Journal for the History of Astronomy, 38(4): 502, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NCa

[Ano08a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 39(1):138–140, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NCb

[Ano08b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 39(2):281–282, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NCc

[Ano08c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 39(3):422– 424, August 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NCd

[Ano08d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 39(4):545–546, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NBa

[Ano08e] Anonymous. Notices of books.

Journal for the History of Astronomy, 39(1):137, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NBb

[Ano08f] Anonymous. Notices of books. Journal for the History of Astronomy, 39(2):279–280, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NBc

[Ano08g] Anonymous. Notices of books. Journal for the History of Astronomy, 39(3):421–422, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2008:NBd

[Ano08h] Anonymous. Notices of books. Journal for the History of Astronomy, 39(4):544, November 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:JUT

[Ano09a] Anonymous. Journals under threat: a joint response from history of science, technology and medicine Editors. Journal for the History of Astronomy, 40(1):v–viii, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NCa

[Ano09b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 40(1):134–136, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NCb

[Ano09c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 40(2):244–246, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NCc

[Ano09d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 40(3):372– 374, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NCd

[Ano09e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 40(4):477–478, November 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NBa

[Ano09f] Anonymous. Notices of books.

Journal for the History of Astronomy, 40(1):134, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2009:NBb

[Ano09g] Anonymous. Notices of books.

Journal for the History of Astronomy, 40(2):243, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NCa

[Ano10a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 41(1):153–156, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NCb

[Ano10b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 41(2):281–282, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NCc

[Ano10c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 41(3):423– 424, August 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NCd

[Ano10d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 41(4):528–530, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NBa

[Ano10e]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 41(1):148–152, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NBb

[Ano10f] Anonymous. Notices of books.

Journal for the History of Astronomy, 41(2):279–280, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2010:NBc

[Ano10g] Anonymous. Notices of books.

Journal for the History of Astronomy, 41(4):527, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NCa

[Ano11a] Anonymous. Notes on contributors. Journal for the History of Astronomy, 42(1):123–124, February 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NCb

[Ano11b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 42(2):280–282, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NCc

[Ano11c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 42(3):422– 424, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NCd

[Ano11d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 42(4):544–545, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NBa

[Ano11e] Anonymous. Notices of books. Journal for the History of Astronomy, 42(2):280, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2011:NBb

[Ano11f]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 42(4):540–541, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:JKW

[Ano12a] Anonymous. Johannes Kepler Working Group. Journal for the History of Astronomy, 43(2):231, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NCa

[Ano12b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 43(1):138–140, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NCb

[Ano12c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 43(2):264–266, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NCc

[Ano12d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 43(3):375– 376, August 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NCd

[Ano12e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 43(4):513–514, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NBa

[Ano12f]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 43(1):136–138, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NBb

[Ano12g] Anonymous. Notices of books.

Journal for the History of Astronomy, 43(2):262–263, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2012:NBc

[Ano12h] Anonymous. Notices of books. Journal for the History of Astronomy, 43(4):512, November 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:CCP

[Ano13a]  Anonymous.  Corrigenda:

[Computing planetary positions: User-friendliness and the Alfonsine Corpus]. Journal for the History of Astronomy, 44 (4):479–480, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [CG13].

Anonymous:2013:NCa

[Ano13b] Anonymous. Notes on contributors. Journal for the History of Astronomy, 44(1):122–124, February 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NCb

[Ano13c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 44(2):232–234, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NCc

[Ano13d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 44(3):375– 376, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NCd

[Ano13e] Anonymous. Notes on contributors. Journal for the History of Astronomy, 44(4):497–498, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NBa

[Ano13f] Anonymous. Notices of books.

Journal for the History of Astronomy, 44(1):121, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NBb

[Ano13g] Anonymous. Notices of books.

Journal for the History of Astronomy, 44(2):230–231, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2013:NBc

[Ano13h] Anonymous. Notices of books.

Journal for the History of Astronomy, 44(4):496, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2014:CUG

[Ano14a]  Anonymous.  Corrigendum:

“Unmasking a Galileo forgery”. Journal for the History of Astronomy, 45(4):507, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Van14a].

Anonymous:2014:IV

[Ano14b] Anonymous. Index to volume 45. Journal for the History of Astronomy, 45(4):508–511, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2014:NCa

[Ano14c] Anonymous. Notes on contributors. Journal for the History of Astronomy, 45(1):139–140, February 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2014:NCb

[Ano14d] Anonymous. Notes on contributors. Journal for the History of Astronomy, 45(2):267–268, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:Ca

[Ano15a]  Anonymous.  Corrigendum.

Journal for the History of Astronomy, 46(1):100, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:Cb

[Ano15b]  Anonymous.  Corrigendum.

Journal for the History of Astronomy, 46(4):504, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:IV

[Ano15c] Anonymous. Index to volume 46. Journal for the History of Astronomy, 46(4):505–508, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:N

[Ano15d]  Anonymous.  Newsworthy.

Journal for the History of Astronomy, 46(1):100, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:NB

[Ano15e]  Anonymous.  Notices of

books. Journal for the History of Astronomy, 46(1):109–110, February 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2015:NBV

[Ano15f] Anonymous. Notices of books:

Voices of the Soviet Space Program: Cosmonauts, Soldiers, and Engineers Who Took the USSR into Space. Slava Gerovitch (Palgrave Macmillan, New York, 2014). Pp. xiv + 305. $90. ISBN 978-1-137-48178-8. Lebensläufe und Himmelsbahnen: Festschrift zum 60. Geburtstag von Jürgen Hamel. Edited by Wolfgang R. Dick and Dietmar Fürst (Acta historica astronomiae, lii; Akademische Verlagsanstalt, Leipzig, 2014). Pp. 293. EUR 24. ISBN 978-3-944913-42-1. Paralipòmens a Vitel· lió: Els orígens de l òptica moderna. Johannes Kepler. Edited by Antoni Malet (Institut d Estudis Catalans, Barcelona, 2010). Pp. cxxxv + 265. EUR 28. ISBN 978-84-9965-031-9 (paper). Istorikoastronomiceskie issledovanija = Historical and Astronomical Researches, vol. 36. Edited by Gennady K. Kurtik (Fizmatlit, Moscow, 2012). Pp.

384. ISBN 978-5-94052-225-6. Der Briefwechsel zwischen Carl Friedrich Gauß und Johann Elert Bode. Edited by Friedhelm Schwemin (Acta historica astronomiae, liii; Akademische Verlagsanstalt, Leipzig, 2014). Pp. 143. EUR 14. ISBN 978-3-944913-43-8. Johann Schöner’s Globe of 1515: Transcription and Study. Chet Van Duzer (Transactions of the American Philosophical Society, c/5; Philadelphia, PA, 2005). Pp. vi + 216. $35. ISBN 978-1-60618-005-1. Journal for the History of Astronomy, 46(3):383– 384, August 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2016:IV

[Ano16a] Anonymous. Index to volume 47. Journal for the History of Astronomy, 47 (4):453–458, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616684249.

Anonymous:2016:NBa

[Ano16b] Anonymous. Notices of books.

Journal for the History of Astronomy, 47(3):351, August 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2016:NBb

[Ano16c] Anonymous. Notices of books.

Journal for the History of Astronomy, 47(4):452, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616675043.

Anonymous:2016:NP

[Ano16d] Anonymous. Notices of publications. Journal for the History of Astronomy, 47(1):100– 101, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
100.full.pdf+html.

Anonymous:2017:C

[Ano17a]  Anonymous.  Corrigendum.

Journal for the History of Astronomy, 48(4):497, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617748004.

Anonymous:2017:IV

[Ano17b] Anonymous. Index to volume 48. Journal for the History of Astronomy, 48 (4):498–501, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617748004.

Anonymous:2017:MHR

[Ano17c] Anonymous. Michael Hoskin receives gold medal from the King of Spain. Journal for the History of Astronomy, 48(1):121, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616688789.

Anonymous:2018:GPA

[Ano18a] Anonymous. Galileo as practising astrologer. Journal for the History of Astronomy, 49 (3):388–391, August 2018. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2018:IV

[Ano18b] Anonymous. Index to volume 49. Journal for the History of Astronomy, 49(4): 506–508, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618814909.

Anonymous:2019:IV

[Ano19] Anonymous. Index to volume 50. Journal for the History of Astronomy, 50(4): 501–504, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619889697.

Anonymous:2020:CIV

[Ano20a] Anonymous. Cumulative index to volumes 41–50. Journal for the History of Astronomy, 51(4):500–513, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620972198.

Anonymous:2020:IV

[Ano20b] Anonymous. Index to volume 51. Journal for the History of Astronomy, 51(4): 514–518, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620975318.

Anonymous:2021:IV

[Ano21a] Anonymous. Index to volume 52. Journal for the History of Astronomy, 52(4): 501–504, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211058926.

Anonymous:2021:PK

[Ano21b] Anonymous. Paul Kunitzsch (1930–2020). Journal for the History of Astronomy, 52 (2):235–236, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211000149.

Anonymous:2022:CHP

[Ano22a]  Anonymous.  Corrigendum

to ‘The heliocentric path of the Moon’. Journal for the History of Astronomy, 53(1): 128, February 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211067843.

See [Gap21].

Anonymous:2022:ERG

[Ano22b] Anonymous. Erratum to ‘A Reading Guide for Bruno’s On the Infinite’. Journal for the History of Astronomy, 53(1):129, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211067841.

See [Blu21].

Anonymous:2022:IV

[Ano22c] Anonymous. Index to volume 53. Journal for the History of Astronomy, 53(4): 509–512, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221135958.

Anonymous:2023:IV

[Ano23] Anonymous. Index to volume 54. Journal for the History of Astronomy, 54(4): 487–490, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231216796.

Anonymous:2024:ENH

[Ano24a] Anonymous. Editor’s note:

Hipparchus’s methods of calculation. Journal for the History of Astronomy, 55(2): 156, May 1, 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231226441.

Anonymous:2024:IV

[Ano24b] Anonymous. Index to volume 55. Journal for the History of Astronomy, 55(4):511– 513, November 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241292442.

Ansari:2005:BRT

[Ans05] S. M. R. Ansari. Book review:

The transmission of Islamic astronomy: Arabic Astronomy in Sanskrit: Al-Birjandion Tadhkira II, Chapter 11 and its Sanskrit Translation. Journal for the History of Astronomy, 36(1):112–114, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Applebaum:2002:BRF

[App02] Wilbur Applebaum. Book review: A flawed guide, Reader’s Guide to the History of Science. Journal for the History of Astronomy, 33(3):289–291, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Applebaum:2005:BRC

[App05] Wilbur Applebaum. Book review: Copernicus and the scientific revolution: It Started with Copernicus: How Turning the World inside Out Led to the Scientific Revolution. Journal for the History of Astronomy, 36(2):242–243, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Apt:1989:BRK

[Apt89] Adam J. Apt. Book review: Keplerian cosmology, Kepler’s Geometrical Cosmology. Journal for the History of Astronomy, 20(3):213–215, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Apt:1999:BRF

[Apt99] Adam Jared Apt. Book review: Flamsteed observed: Flamsteed’s Stars: New Perspectives on the Life and Work of the First Astronomer Royal, 1646– 1719. Journal for the History of Astronomy, 30(2):178–180, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Apt:2014:BRW

[Apt14] Adam Jared Apt. Book review:

William Gascoigne: In Search of William Gascoigne: Seventeenth Century Astronomer. David Sellers (Springer, New York, 2012). Pp. x + 222. $100. ISBN 978-1-4614-4096-3. Journal for the History of Astronomy, 45(4):490–492, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1986:ARV

[AR86] A. F. Aveni and G. Romano.

Archaeoastronomical research in Veneto-Friuli, Italy. Journal for the History of Astronomy, 17(10):S23–S31, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:2000:TOM

[AR00] Anthony Aveni and Giuliano Romano. Temple orientations in Magna Graecia and Sicily. Journal for the History of Astronomy, 31(25):S51–S57, February 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Armitage:1970:BRE

[Arm70] A. Armitage. Book review: Edmond Halley: Edmond Halley: Genius in Eclipse. Journal for the History of Astronomy, 1(2): 156–157, August 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Armitage:1972:BRL

[Arm72] A. Armitage. Book review: Lord Rosse’s Telescope: The Astronomy of Birr Castle. Journal for the History of Astronomy, 3(2):146, June 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Armitage:1973:BRH

[Arm73] A. Armitage. Book review: History to our own times: Discovering the Universe: a History of Astronomy. Journal for the History of Astronomy, 4(3):209, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Armitage:1974:BRH

[Arm74] A. Armitage. Book review: an historical anthology: Illustrated Sources in History: Astronomy. Journal for the History of Astronomy, 5(1):65, February 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Arnold:1995:ERI

[Arn95] Philip P. Arnold. Essay review: Interpreting Mesoamerican texts: The Sky in Mayan Literature. Journal for the History of Astronomy, 26(20): S86–S89, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Arnaldi:2011:ARM

[Arn11] Mario Arnaldi. An ancient rule for making portable altitude sundials from an ‘unedited’ medieval text of the tenth century. Journal for the History of Astronomy, 42(2):141–160, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Arnaldi:1997:RCD

[AS97] Mario Arnaldi and Karlheinz Schaldach. A Roman cylinder dial: Witness to a forgotten tradition. Journal for the History of Astronomy, 28(2):107– 117, May 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anastasiou:2014:AMC

[ASCE14] M. Anastasiou, J. H. Seiradakis, C. C. Carman, and K. Efstathiou. The Antikythera mechanism: The construction of the metonic pointer and the back plate spirals. Journal for the History of Astronomy, 45 (4):418–441, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anastasiou:2013:AEP

[ASE^{+}13] Magdalini Anastasiou, John H. Seiradakis, James Evans, Stella Drougou, and Kyriakos Efstathiou. The astronomical events of the Parapegma of the Antikythera mechanism. Journal for the History of Astronomy, 44(2):173–A10, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashbrook:1972:BRA

[Ash72] Joseph Ashbrook. Book review: Astronomy at the Royal Institution: The Royal Institution Library of Science: Astronomy. Journal for the History of Astronomy, 3(1):70–71, February 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashbrook:1979:BRA

[Ash79a] Joseph Ashbrook. Book review: Astronomy after Herschel, Geschichte der Astronomie von Herschel bis Hertzsprung. Journal for the History of Astronomy, 10(1): 67–69, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashbrook:1979:BRN

[Ash79b] Joseph Ashbrook. Book review: Noctilucent clouds, Entwicklungsphasen der Erforschung der Leuchtenden Nachtwolken. Journal for the History of Astronomy, 10(1):69–70, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:1980:PFO

[Ash80] William B. Ashworth, Jr. A probable Flamsteed observation of the Cassiopeia A supernova. Journal for the History of Astronomy, 11(1):1–9, February 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:1981:HDN

[Ash81] William B. Ashworth, Jr. Halley’s discovery of NGC 6231 and the hazards of early star nomenclature. Journal for the History of Astronomy, 12(1): 1–10, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:1982:BRO

[Ash82] William B. Ashworth, Jr. Book review: an Observer’s atlas: Sky Atlas 2000·0. Journal for the History of Astronomy, 13(3):212–213, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:1986:BRG

[Ash86] William B. Ashworth, Jr. Book review: The Gregorian calendar in context: Gregorian Reform of the Calendar: Proceedings of the Vatican Conference to Commemorate its 400th Anniversary, 1582–1982. Journal for the History of Astronomy, 17(2):147–148, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:1987:BRA

[Ash87] William B. Ashworth, Jr. Book review: Astronomical charts: Maps of the Heavens. Journal for the History of Astronomy, 18(4):287–288, November 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashmore:2000:ERA

[Ash00] Patrick Ashmore. Essay review: Archaeoastronomy of the British Isles: Astronomy in Prehistoric Britain and Ireland. Journal for the History of Astronomy, 31(25):S77–S85, February 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ashworth:2003:FSF

[Ash03] William B. Ashworth, Jr. Faujas-De-Saint-Fond visits the Herschels at Datchet. Journal for the History of Astronomy, 34(3):321–324, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:2013:AIM

[ASS13] Anthony F. Aveni, William Saturno, and David Stuart.

Astronomical implications of Maya hieroglyphic notations at xultun. Journal for the History of Astronomy, 44(1):1–16, February 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Atkinson:1975:MAP

[Atk75] R. J. C. Atkinson. Megalithic astronomy — a Prehistorian’s comments. Journal for the History of Astronomy, 6(1):42–52, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Atkinson:1976:SS

[Atk76] R. J. C. Atkinson. The Stonehenge stations. Journal for the History of Astronomy, 7 (2):142–144, June 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Atkinson:1982:EP

[Atk82] R. D’E. Atkinson. The Eros parallax, 1930–31. Journal for the History of Astronomy, 13 (2):77–83, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Atkinson:1986:OAT

[Atk86] R. J. C. Atkinson. Obituary:

Alexander Thom. Journal for the History of Astronomy, 17 (1):73–75, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Atwood:1979:BRL

[Atw79] Seth G. Atwood. Book review: “longitude” Harrison, The Lost Science of John “Longitude” Harrison. Journal for the History of Astronomy, 10 (1):65–66, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aubin:2012:BRM

[Aub12] David Aubin. Book review: Margaret and William Huggins: Unravelling Starlight: William and Margaret Huggins and the Rise of the New Astronomy. Journal for the History of Astronomy, 43(4):510– 511, November 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aubin:2016:JDA

[Aub16] David Aubin. Jeanne Dumée as astronomer and woman in Seventeenth-Century France: The myth and her lost voice. Journal for the History of Astronomy, 47(3):231–255, August 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aubin:2025:JS

[Aub25] David Aubin. Jesuit and scientist. Journal for the History of Astronomy, 56(1):110– 112, February 2025. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241263514.

Augusto:2024:MYA

[Aug24] Pedro Augusto. Madeira: 300 years of an astronomical site. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231220367.

Aulenbacher:2016:TGA

[Aul16] Gerhard Aulenbacher. Two Golden ages of gnomonics. Journal for the History of Astronomy, 47(2):220–222, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1987:BRE

[Ave87]  Anthony F. Aveni.  Book

review: The expedition to Peru: La Geometrización de la Tierra: Observaciones y Resultados de la Expedición Geodésica Hispano–Francesa al Virreinato del Peru (1735– 1744). Journal for the History of Astronomy, 18(1):64– 65, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1992:ERT

[Ave92] A. F. Aveni. Essay review:

Time among the Inca: Time and Calendars in the Inca Empire. Journal for the History of Astronomy, 23(17):S68–S73, February 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1994:BRM

[Ave94] Anthony Aveni. Book review: Mayans past and present: Time and the Highland Maya. Journal for the History of Astronomy, 25(19):S86–S87, February 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1995:ERF

[Ave95] Anthony F. Aveni. Essay review: Frombork 1992: Where worlds and disciplines collide: Time and Astronomy at the Meeting of Two Worlds, Proceedings of the International Symposium held April 27– May 2, 1992, in Frombork, Poland, organized by the Department of Historical Anthropology. Edited by Stanisław Iwaniszewski, Arnold Lebeuf, Andrzej Wierciński and Mariusz S. Ziółkowski (Publication no. 10, Institute of Archaeology, Warsaw University Centre for Latin American Studies (CESLA); Warsaw, 1994). Pp.

527. Journal for the History of Astronomy, 26(20):S74–S79, February 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1997:BRI

[Ave97] Anthony F. Aveni. Book review: The Incas and the sky:

Astronomy and Empire in the Ancient Andes: The Cultural Origins of Inca Sky Watching. Journal for the History of Astronomy, 28(22):S85–S87, February 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:1999:AMC

[Ave99] Anthony Aveni. Astronomy in the Mexican Codex Borgia. Journal for the History of Astronomy, 30(24):S1–S20, February 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See errata [Ave06].

Aveni:2005:OPD

[Ave05] Anthony F. Aveni. Observations on the pecked designs and other figures carved on the south platform of the pyramid of the Sun at Teotihuacan. Journal for the History of Astronomy, 36(1):31–47, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:2006:EAM

[Ave06] Anthony F. Aveni. Errata: [Astronomy in the Mexican Codex Borgia]. Journal for the History of Astronomy, 37(3):372, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Ave99].

Aveni:2009:BRN

[Ave09] Anthony F. Aveni. Book review: North American constellation myths, Reachable Stars: Patterns in the Ethnoastronomy of Eastern North America. Journal for the History of Astronomy, 40(1):124–125, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:2011:BRA

[Ave11] Anthony F. Aveni. Book review: Astronomy and civilization: The Power of Stars: How Celestial Observations Have Shaped Civilization. Journal for the History of Astronomy, 42(3):421–422, August 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Aveni:2018:AOT

[Ave18] Anthony Aveni. Astronomers as oracular technicians in the pre-Columbian Andes: review of Constructions of Time and History in the Pre-Columbian Andes, Edward Swenson & Andrew P. Roddick (University Press of Colorado, Boulder, CO, 2018). Pp. xiv + 288. $72. ISBN 978-1-60732-641-0. Journal for the History of Astronomy, 49(3):393–395, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Avelar:2021:IEH

[Ave21]  Helena Avelar.  Ibn Ezra

from Hebrew to Latin. Journal for the History of Astronomy, 52(2):242–243, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828621995307.

Axworthy:2017:BRE

[Axw17] Angela Axworthy. Book review: Early Modern Cosmological Debates. Journal for the History of Astronomy, 48 (4):489–491, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617724895.

Aramburu-Zabala:2002:AOS

[AZB02] Javier Aramburu-Zabala and Juan Antonio Belmonte. On the astronomical orientation of the square talayots of Mallorca. Journal for the History of Astronomy, 33(27):S67–S74, February 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Azhar:2020:IMC

[Azh20]  Feraz Azhar.  Interdisci-

plinarity and modern cosmology. Journal for the History of Astronomy, 51(4): 485–486, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620945487.

Azzolini:2006:BRA

[Azz06] Monica Azzolini. Book review: Astrology in the Low Countries: The Limits of Influence: Pico, Louvain, and the Crisis of Renaissance Astrology. Journal for the History of Astronomy, 37(3):360–361, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Azzolini:2011:RAA

[Azz11]  Monica Azzolini.  Refining

the Astrologer’s art: Astrological diagrams in Bodleian MS Canon. Misc. 24 and Cardano’s Libelli Quinque (1547). Journal for the History of Astronomy, 42(1):1–25, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:1993:SMT

[BAE93] Juan A. Belmonte, Antonio Aparicio, and César Esteban. A solstitial marker in Tenerife: The “Majanos De Chacona”. Journal for the History of Astronomy, 24(18):S65–S68, February 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [EBA94].

Baker:2015:BRA

[Bak15] Alexi Baker. Book review: The achromatic telescope controversy: Francis Watkins and the Dollond Telescope Patent Controversy. Brian Gee, edited by Anita McConnell and A.D. Morrison-Low (Ashgate, Farnham, 2014). Pp. xxvi + 392. £85. ISBN 978-1-4094-6643-7. Journal for the History of Astronomy, 46(3):376–377, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baldwin:1985:MAC

[Bal85] Martha R. Baldwin. Magnetism and the anti-Copernican polemic. Journal for the History of Astronomy, 16(3):155– 174, October 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baneke:2010:TTL

[Ban10] David Baneke. Teach and travel: Leiden Observatory and the Renaissance of Dutch astronomy in the interwar years. Journal for the History of Astronomy, 41(2):167–198, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baneke:2012:AO

[Ban12] David Baneke. Astronomers of 1908 online. Journal for the History of Astronomy, 43(2): 232, May 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baneke:2013:BRO

[Ban13] David Baneke. Book review: an observatory that failed, Between Rhetoric and Reality: Astronomical Practices at the Observatory of the Amsterdam Society ‘Felix Meritis’, 1786– 1889. Journal for the History of Astronomy, 44(4):489–490, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bartholomew:1988:HSC

[Bar88] C. F. Bartholomew. Herbert Spencer’s contribution to solar physics. Journal for the History of Astronomy, 19(1):1–28, February 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bartha:1994:AIH

[Bar94] Lajos Bartha. Astrophysical instruments in Hungary, 1871– 1911. Journal for the History of Astronomy, 25(2):77– 91, May 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barker:1999:CCP

[Bar99] Peter Barker. Copernicus and the critics of Ptolemy. Journal for the History of Astronomy, 30(4):343–358, November 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barker:2005:BRM

[Bar05] Peter Barker. Book review: Method in Copernicus: Copernicus’s Originality: Towards Integration of Contemporary Copernican Studies. Journal for the History of Astronomy, 36(4):454–456, November 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barker:2013:ABL

[Bar13a]  Peter Barker.  Albert of

Brudzewo’s little commentary on George Peurbach’s ‘Theoricae Novae Planetarum’. Journal for the History of Astronomy, 44(2):125–148, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barker:2013:BRD

[Bar13b] Peter Barker. Book review:

Dobrzycki’s legacy: Selected Papers on Medieval and Renaissance Astronomy. Journal for the History of Astronomy, 44(1):105–107, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barreca:2016:GUS

[Bar16] Francesco Barreca. Galileo under the sign of interdisciplinarity. Journal for the History of Astronomy, 47 (4):424–425, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668749.

Bardi:2018:PPT

[Bar18] Alberto Bardi. The paradosis of the Persian tables, a source on astronomy between the Ilkhanate and the Eastern Roman Empire. Journal for the History of Astronomy, 49(2):239–260, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618762461.

Bardi:2019:CFL

[Bar19]  Alberto Bardi.  Coperni-

cus on film … again. The Secrets of De Revolutionibus Michał Juszczakiewicz & Nicolaus Copernicus Foundation, (Truszczyny, Poland, 2018). 53 minutes. Available from the Nicolaus Copernicus Foundation (copernicus@ekoprussia.home.pl). Journal for the History of Astronomy, 50(1):116–117, February 1, 2019. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL

https://journals.sagepub.
com/doi/full/10.1177/0021828618823981.

Barker:2022:PIT

[Bar22] Peter Barker. Peurbach’s influential textbook. Journal for the History of Astronomy, 53(3):369–370, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221110573.

Batten:1987:BRA

[Bat87]  A. H. Batten.  Book re-

view: Astronomy in Russia: Le Développement de l’Astronomie en U.R.S.S. 1917– 1935. Journal for the History of Astronomy, 18(1):67– 68, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Batten:1990:BRA

[Bat90] Alan H. Batten. Book review: Astronomy in Canada, the cold light of dawn: A History of Canadian Astronomy. Journal for the History of Astronomy, 21(3):307–308, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Batten:2015:OAN

[Bat15]  Alan H. Batten.  Obser-

vational astronomy in New Zealand. Journal for the History of Astronomy, 46(4):496– 497, November 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baumgartner:1986:SFI

[Bau86] Frederic J. Baumgartner. Scepticism and French interest in Copernicanism to 1630. Journal for the History of Astronomy, 17(2):77–88, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baumgartner:1987:SSP

[Bau87] Frederic J. Baumgartner. Sunspots or Sun’s planets: Jean Tarde and the sunspot controversy of the early Seventeenth Century. Journal for the History of Astronomy, 18(1):44–54, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2003:DHA

[BBMG03] Juan Antonio Belmonte, Marie Antonia Perera Betancort, Rita Marrero, and Antonio Tejera Gaspar. The dolmens and hawanat of Africa Proconsularis revisited. Journal for the History of Astronomy, 34 (3):305–320, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bricker:1997:DHT

[BBW97] Harvey M. Bricker, Victoria R. Bricker, and Bettina Wulfing. Determining the historicity of three astronomical almanacs in the Madrid Codex. Journal for the History of Astronomy, 28 (22):S17–S36, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Black:2016:GTS

[BC16] William B. Black and David Clarke. Glasgow time signals. Journal for the History of Astronomy, 47(3):256–293, August 2016. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Bartky:1981:FTB

[BD81] Ian R. Bartky and Steven J.

Dick. The first time balls. Journal for the History of Astronomy, 12(3):155–164, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bartky:1982:FNA

[BD82] Ian R. Bartky and Steven J.

Dick. The first North American time ball. Journal for the History of Astronomy, 13 (1):50–54, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2002:AST

[BEBM02] Juan Antonio Belmonte, César Esteban, Maria Antonia Perera Betancort, and Rita Marrero. Archaeoastronomy in the Sahara: The tombs of the Garamantes at Wadi El Agial, Fezzan, Libya. Journal for the History of Astronomy, 33(27): S1–S19, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:1999:BRF

[Bec99a] Barbara J. Becker. Book review: Five women of science: Whatever Shines Should Be Observed. Journal for the History of Astronomy, 30(3):333– 334, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:1999:PIB

[BEC^{+}99b] J. A. Belmonte, C. Esteban, L. Cuesta, M. A. Perera Betancort, and J. J. Jiménez González. Pre-Islamic burial monuments in Northern and Saharan Morocco. Journal for the History of Astronomy, 30 (24):S21–S34, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2000:PPV

[Bec00] Barbara J. Becker. Priority, persuasion, and the virtue of perseverence: William Huggins’s efforts to photograph the solar corona without an eclipse. Journal for the History of Astronomy, 31(3):223–243, August 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2001:VMW

[Bec01] Barbara J. Becker. Visionary memories: William Huggins and the origins of astrophysics. Journal for the History of Astronomy, 32(1):43–62, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2010:BRC

[Bec10] Barbara Becker. Book review: Cosmology in the early Twentieth century: The Day We Found the Universe. Journal for the History of Astronomy, 41(1):133–135, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2013:HPW

[Bec13]  Barbara J. Becker.  His-

toric photographs by William and Margaret Lindsay Huggins. Journal for the History of Astronomy, 44(4):481–484, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2014:BA

[Bec14] Barbara J. Becker. Biographies of astrophotographers. Journal for the History of Astronomy, 45(3):378–379, August 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2015:BRC

[Bec15]  Barbara J. Becker.  Book

review: A concise summary of the Herschels. William and Caroline Herschel: Pioneers in Late 18th-Century Astronomy. Michael Hoskin (Springer, Dordrecht, 2014). Pp. viii + 122. ISBN 978-94-007-6874-1 (paper, $49.99), 978-94-007-6875-8 (eBook, $39.99). Journal for the History of Astronomy, 46 (2):243–244, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Becker:2018:IAA

[Bec18] Barbara J. Becker. An Irish amateur astronomer. Journal for the History of Astronomy, 49(2):263–264, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618772196.

Becker:2019:BRC

[Bec19] Barbara J. Becker. Book review: From celestial placeholders to physical objects: Making Stars Physical: The Astronomy of Sir John Herschel Stephen Case, (University of Pittsburgh Press, Pittsburgh, PA, 2018). Pp. viii + 319. US$39.95. ISBN 978-0-8229-4530-7. Journal for the History of Astronomy, 50(1):109–111, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618806145.

Becker:2022:GHS

[Bec22] Barbara J. Becker. Gauging the Herschels’ star gauging programme. Journal for the History of Astronomy, 53 (2):231–233, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286221080454.

Bedini:1980:BRR

[Bed80] Silvio A. Bedini. Book review: Ready-to-assemble instruments, A Trilogy of Time Instruments. Journal for the History of Astronomy, 11(1): 72–73, February 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bedini:2010:BRP

[Bed10] Peter D. Bedini. Book review: Pioneering deep space: William H. Pickering: America’s Deep Space Pioneer. Journal for the History of Astronomy, 41(1):146–147, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Beer:1970:HRG

[Bee70] Arthur Beer. Hartner and the riddle of the Golden Horns. Journal for the History of Astronomy, 1(2):139–143, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Beer:1971:BRA

[Bee71] Arthur Beer. Book review: Astronomy and culture: Oriens, Occidens. Journal for the History of Astronomy, 2(3):201– 202, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Beekman:2006:YDH

[Bee06]  George Beekman.  I. O.

Yarkovsky and the discovery of ‘his’ effect. Journal for the History of Astronomy, 37(1):

71–86, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:1998:MAA

[BEG98] Juan A. Belmonte, César Esteban, and José Juan Jiménez González. Mediterranean archaeoastronomy and archaeotopography: Pre-Roman tombs of Africa Proconsularis. Journal for the History of Astronomy, 29(23):S7–S24, February 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bell:1977:BRH

[Bel77] Trudy E. Bell. Book review: Harvard pioneers, The Harvard College Observatory. Journal for the History of Astronomy, 8(3):214–215, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:1997:MAA

[Bel97] Juan Antonio Belmonte. Mediterranean archaeotopography and archaeoastronomy: Two examples of dolmenic necropolises in the Jordan Valley. Journal for the History of Astronomy, 28 (22):S37–S43, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2001:BRR

[Bel01a] Juan Antonio Belmonte. Book review: Rider of the lost megaliths: Tombs, Temples and Their Orientations: a New Perspective on Mediterranean Prehistory. Journal for the History of Astronomy, 32(26): S85–S89, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2001:OOK

[Bel01b] Juan Antonio Belmonte. On the orientation of Old Kingdom Egyptian pyramids. Journal for the History of Astronomy, 32(26):S1–S20, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2006:ERE

[Bel06] Juan Antonio Belmonte. Essay review: Egyptian astronomy: La Astronomía del Antiguo Egipto. Journal for the History of Astronomy, 37(3): 356–359, August 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bell:2007:BRD

[Bel07] Trudy E. Bell. Book review: The Dudley Observatory: Civic Astronomy: Albany’s Dudley Observatory, 1852–2002. Journal for the History of Astronomy, 38(2): 241–244, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bell:2009:BRB

[Bel09a] Trudy E. Bell. Book review: British University observatories, British University Observatories, 1772–1939. Journal for the History of Astronomy, 40(3):364–366, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2009:BRR

[Bel09b] Juan Antonio Belmonte. Book review: a reader in new world astronomy, Foundations of New World Cultural Astronomy: a Reader with Commentary. Journal for the History of Astronomy, 40(3):352–353, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2011:BRA

[Bel11] Juan Antonio Belmonte. Book review: Archaeoastronomy debated: Archeoastronomia: Un Dibattito tra Archaeologi ed Astronomi alla Ricerca di un Metodo Comune. Journal for the History of Astronomy, 42 (1):113–114, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bellhouse:2014:BRS

[Bel14a] David R. Bellhouse. Book review: Statistical techniques for analyzing astronomical tables: Islamic Astronomical Tables: Mathematical Analysis and Historical Investigation. Benno van Dalen (Variorum Collected Studies Series, CS1040; Ashgate Publishing Ltd, Farnham, Surrey, 2013). Pp. xii + 350. £81. ISBN 978-1-4724-2238-5. Journal for the History of Astronomy, 45(4): 500–504, November 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2014:BRA

[Bel14b] Juan Antonio Belmonte. Book review: The archaeotopography of Egypt: Architecture, Astronomy and Sacred Landscape in Ancient Egypt. Giulio Magli (Cambridge University Press, Cambridge, 2013). Pp. xii + 272. £60. ISBN 978-1-10703-208-8. Journal for the History of Astronomy, 45(2): 264–266, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2016:LSA

[Bel16a] Juan Antonio Belmonte. Landscapes, skyscapes, archaeoastronomy. Journal for the History of Astronomy, 47 (4):445–446, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616665949.

Belmonte:2016:MAS

[Bel16b] Juan Antonio Belmonte. Mayan architecture and the sky. Journal for the History of Astronomy, 47(4):439–440, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616665950.

Belmonte:2019:BRC

[Bel19] Juan Antonio Belmonte. Book review: Cultural landscapes. Journal for the History of Astronomy, 50(4):482–483, November 1, 2019. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619871725.

Belteki:2020:TGN

[Bel20]  Daniel Belteki.  Trust in

glass: Negotiating the purchase of the object glass for the Airy transit circle. Journal for the History of Astronomy, 51(4):401–422, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620968631.

Belmonte:2022:MM

[Bel22] Juan Antonio Belmonte. Myth and meteorology. Journal for the History of Astronomy, 53(2):233–234, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221080452.

Bennett:1975:CWA

[Ben75] J. A. Bennett. Christopher Wren: Astronomy, architecture, and the mathematical sciences. Journal for the History of Astronomy, 6(3):149– 184, October 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1976:BRD

[Ben76a] J. A. Bennett. Book review:

Discoverers of nebulae, The Search for the Nebulae. Journal for the History of Astronomy, 7(1):67, February 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1976:PPS

[Ben76b] J. A. Bennett. “On the power of penetrating into space”: The telescopes of William Herschel. Journal for the History of Astronomy, 7(2):75–108, June 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1981:CMP

[Ben81a] J. A. Bennett. Cosmology and the magnetical philosophy, 1640–1680. Journal for the History of Astronomy, 12(3): 165–177, October 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1981:SFW

[Ben81b] J. A. Bennett. A surviving flat from William Lassell’s fourfoot equatorial Newtonian telescope. Journal for the History of Astronomy, 12(3):195–197, October 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1982:BRH

[Ben82a] J. A. Bennett. Book review: The Huygens symposium: Studies on Christiaan Huygens. Journal for the History of Astronomy, 13(2):138– 139, June 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1982:FAO

[Ben82b] J. A. Bennett. The first aplanatic object glass. Journal for the History of Astronomy, 13(3):206–208, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1984:BRS

[Ben84] J. A. Bennett. Book review: Seventeenth-Century English astronomers: Three North Country Astronomers. Journal for the History of Astronomy, 15(1):54–56, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1987:BRH

[Ben87a] J. A. Bennett. Book review: Historic scientific photographs: Beyond Vision. Journal for the History of Astronomy, 18 (1):65, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1987:BRN

[Ben87b] J. A. Bennett. Book review:

Newton’s contemporaries: Under Newton’s Shadow: Astronomical Practices in the Seventeenth Century. Journal for the History of Astronomy, 18(2): 136, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1991:BRM

[Ben91a] J. A. Bennett. Book review: Measuring angles: Dividing the Circle: The Development of Critical Angular Measurement in Astronomy 1500– 1850. Journal for the History of Astronomy, 22(4):319–320, November 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1991:BRS

[Ben91b] J. A. Bennett. Book review:

Symposium on scientific instruments, Studies in the History of Scientific Instruments. Journal for the History of Astronomy, 22(3):243–244, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1992:EQE

[Ben92] J. A. Bennett. The English quadrant in Europe: Instruments and the growth of consensus in practical astronomy. Journal for the History of Astronomy, 23(1):1–14, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1993:BRT

[Ben93a] J. A. Bennett. Book review: Telescopic curiosities: Unusual Telescopes. Journal for the History of Astronomy, 24(4):306, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1993:SLL

[Ben93b] J. A. Bennett. Science lost and longitude found: The tercentenary of John Harrison. Journal for the History of Astronomy, 24(4):281–287, November 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1999:BRE

[Ben99] Jim Bennett. Book review: Essays in the history of horology: Of Time and Measurement: Studies in the History of Horology and Fine Technology. Journal for the History of Astronomy, 30(1):78–79, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:2009:BRJ

[Ben09] Jim Bennett. Book review: Jesse Ramsden, Jesse Ramsden (1735–1800): London’s Leading Scientific Instrument Maker. Journal for the History of Astronomy, 40(2):226– 227, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:2013:BRW

[Ben13] Jim Bennett. Book review:

William Herschel’s cosmology, The Construction of the Heavens: William Herschel’s Cosmology. Journal for the History of Astronomy, 44(2):213– 214, May 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:2014:RRR

[Ben14a] Jim Bennett. Rosse, Robinson and Rambaut: Creating a regime of observing at Birr Castle, 1840–1850. Journal for the History of Astronomy, 45 (4):400–417, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Benson:2014:CPS

[Ben14b] Michael Benson. Cosmigraphics: picturing space through time. Abrams, New York, NY, 2014. ISBN 1-4197-1387-6. 320 pp. LCCN N8217.C62 B46 2014.

Bennett:2018:CG

[Ben18]  Jim Bennett.  Coronelli’s

globes. Journal for the History of Astronomy, 49 (1):123–125, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618755444.

Bennett:2020:BRW

[Ben20] Jim Bennett. Book review: William Bond & Son, Boston: From Celestial to Terrestrial Timekeeping: Clockmaking in the Bond Family. Donald Saff (Antiquarian Horological Society, London, 2019). Pp. xiii + 424. £45. ISBN 978-0-901180-54-4. Journal for the History of Astronomy, 51(1): 125–126, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619900816.

Berendzen:1970:BRS

[Ber70] R. B. Berendzen. Book review: Shapley: Octogenarian Reminiscences: Through Rugged Ways to the Stars. Journal for the History of Astronomy, 1 (1):85–87, February 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1974:ERP

[Ber74] Richard Berendzen. Essay review: Pioneer of American astronomy: The Legacy of George Ellery Hale. Journal for the History of Astronomy, 5(1): 61–63, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1981:BRW

[Ber81] Richard Berendzen. Book review: Why bother with astronomy?: Man and the Stars. Journal for the History of Astronomy, 12(2):149–150, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1982:ERS

[Ber82] Richard Berendzen. Essay review: Stellar astronomy: Stellar Astronomy: Historical Studies, the Expanding Universe: Astronomy’s ‘Great Debate’ 1900–1931. Journal for the History of Astronomy, 13(3):209–211, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1989:BRE

[Ber89] Richard Berendzen. Book review: To err is human, science and objectivity: Episodes in the History of Astronomy. Journal for the History of Astronomy, 20(2):150–152, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berggren:1992:BRA

[Ber92] J. L. Berggren. Book review:

Arabs and the fixed stars: The Arabs and the Stars: Texts and Traditions on the Fixed Stars, and Their Influence in Medieval Europe. Journal for the History of Astronomy, 23 (2):147–149, May 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1994:BRH

[Ber94] Richard Berendzen. Book review: an Historian’s miscellany: The Great Copernicus Chase: And other Adventures in Astronomical History. Journal for the History of Astronomy, 25(1):67–69, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1997:BRC

[Ber97a] Richard Berendzen. Book review: a century-old revolution: The Milky Way Galaxy and Statistical Cosmology 1890– 1924. Journal for the History of Astronomy, 28(4):359–360, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berggren:1997:BRI

[Ber97b] Len Berggren. Book review: Islamic astronomers and the Equant: A History of Arabic Astronomy: Planetary Theories during the Golden Age of Islam. Journal for the History of Astronomy, 28(3):270– 272, August 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:2001:BRL

[Ber01] Richard Berendzen. Book review: Light on Kapteyn: The Legacy of J. C. Kapteyn: Studies on Kapteyn and the Development of Modern Astronomy. Journal for the History of Astronomy, 32(2):183–184, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berggren:2012:BRS

[Ber12] Len Berggren. Book review:

The spherics of Theodosius, Theodosius, Sphaerica: Arabic and Medieval Latin Translations. Journal for the History of Astronomy, 43(2):250– 252, May 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berggren:2014:BRA

[Ber14] J. L. Berggren. Book review: Astronomy in Islam: Islamic Astronomy and Geography, ISBN 978-1-4094-4201-1. Journal for the History of Astronomy, 45(2):258–259, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bernard:2019:BRP

[Ber19]  Alain Bernard.  Book re-

view: a philosophically coherent Ptolemy. Journal for the History of Astronomy, 50 (2):249–250, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619841993.

Berggren:2022:LLP

[Ber22] J. Lennart Berggren. The long legacy of Ptolemy. Journal for the History of Astronomy, 53(2):223–229, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286221081975.

Buchwald:2013:NOC

[BF13] Jed Z. Buchwald and Mordechai Feingold. Newton and the origin of civilization. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-15478-3 (hardcover). ix + 528 pp. LCCN QC16.N7 B93 2013.

Belmonte:2010:OAE

[BFAH^{+}10] Juan Antonio Belmonte, Magdi Fekri, Yasser A. Abdel-Hadi, Mosalam Shaltout, and A. Céasar González García. On the orientation of Ancient Egyptian temples: (5) testing the theory in middle Egypt and Sudan. Journal for the History of Astronomy, 41(1): 65–93, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Barker:2003:PPR

[BG03] Peter Barker and Bernard R. Goldstein. Patronage and the production of De Revolutionibus. Journal for the History of Astronomy, 34(4):345–368, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2010:AHN

[BG10] Juan Antonio Belmonte and A. César González García. Antiochos’s Hierothesion at Nemrud Dag revisited: Adjusting the date in the light of astronomical evidence. Journal for the History of Astronomy, 41(4):469–481, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2012:GAI

[BG12] Juan Antonio Belmonte and A. César González García. The ‘genetic’ analysis of Iberian dolmens: a test of the idea in the Central Pyrénées. Journal for the History of Astronomy, 43(2):227–231, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2013:OMM

[BGP13] Juan Antonio Belmonte, A. César González García, and Andrea Polcaro. On the orientation of megalithic monuments of the Transjordan plateau: New clues for an astronomical interpretation. Journal for the History of Astronomy, 44(4): 429–455, November 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1973:AVMa

[BH73a] R. Berendzen and R. Hart. Adriaan Van Maanen’s influence on the island universe theory: Part 1. Journal for the History of Astronomy, 4(1): 46–56, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1973:AVMb

[BH73b] R. Berendzen and R. Hart. Adriaan Van Maanen’s influence on the island universe theory: Part 2. Journal for the History of Astronomy, 4(2):73– 98, June 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bennett:1981:RPI

[BH81] J. A. Bennett and Michael Hoskin. The Rosse papers and instruments. Journal for the History of Astronomy, 12(3): 216–229, October 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Butler:1987:AAO

[BH87] John Butler and Michael Hoskin. The archives of Armagh Observatory. Journal for the History of Astronomy, 18 (4):295–307, November 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boudet:2012:EAT

[BH12] Jean-Patrice Boudet and Matthieu Husson. The earliest astronomical tables in French (c. 1271). Journal for the History of Astronomy, 43(3):287–298, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bialas:2004:BRT

[Bia04] Volker Bialas. Book review:

Tycho and Kepler: Tycho and Kepler: The Unlikely Partnership That Forever Changed Our Understanding of the Heavens. Journal for the History of Astronomy, 35(1):115–116, February 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bialas:2015:KS

[Bia15] Volker Bialas. Kepler’s Sygkepleriazein. Journal for the History of Astronomy, 46(4):498– 500, November 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bienias:2021:IEC

[Bie21] Barbara Bienias. An introduction to English calendars. Journal for the History of Astronomy, 52(3):367–370, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211010000.

Bigg:2010:BRM

[Big10] C. Bigg. Book review: Meadows on Lockyer: Science and Controversy: a Biography of Sir Norman Lockyer, Founder Editor of Nature. Journal for the History of Astronomy, 41 (4):509–510, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Biraud:1974:SDC

[Bir74] F. Biraud. Le souvenir du calendrier julien dans les traditions populaires. (French) [The memory of the Julian calendar in folk traditions]. Journal for the History of Astronomy, 5(3): 199–200, October 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blanpied:1979:BRJ

[Bla79] William A. Blanpied. Book review: Jai Singh’s Observatories, Stone Observatories in India. Journal for the History of Astronomy, 10(1):56– 57, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blaauw:1991:EYE

[Bla91a] Adriaan Blaauw. The early years of the European Southern Observatory: an effort in Europeanization. Journal for the History of Astronomy, 22 (1):87–99, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blackwell:1991:BRG

[Bla91b] Richard J. Blackwell. Book review: Galileo and the patronage of science, Essays on the Trial of Galileo. Journal for the History of Astronomy, 22 (2):189–190, May 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blair:1997:BRF

[Bla97] Ann Blair. Book review: Focus on Clavius: Between Copernicus and Galileo: Christoph Clavius and the Collapse of Ptolemaic Cosmology. Journal for the History of Astronomy, 28(2):170–172, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blaauw:1998:BRA

[Bla98a] Adriaan Blaauw. Book review:

Aura seen from within: Aura and its US National Observatories. Journal for the History of Astronomy, 29(4):394–396, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blaauw:1998:ERF

[Bla98b] Adriaan Blaauw. Essay review: The first half-century of Yerkes: Yerkes Observatory, 1892–1950: The Birth, near Death, and Resurrection of a Scientific Research Institution.

Journal for the History of Astronomy, 29(1):80–82, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blair:2001:BRH

[Bla01] Ann Blair. Book review: The heavens in French poetry: La Poésie du Ciel en France dans la Seconde Moitié du Seizième Siècle. Journal for the History of Astronomy, 32(2):179– 181, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blaasjo:2014:CAM

[Blå14] Viktor Blåsjö. A critique of the arguments for Maragha influence on Copernicus. Journal for the History of Astronomy, 45(2):183–195, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bloom:1978:BPH

[Blo78] Terrie F. Bloom. Borrowed perceptions: Harriot’s maps of the Moon. Journal for the History of Astronomy, 9(2):117– 122, June 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blum:2000:BRG

[Blu00] Paul Richard Blum. Book review: Giordano Bruno considered: Giordano Bruno and Renaissance Science, Cause, Principle and Unity. Journal for the History of Astronomy, 31(1):72–75, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blumenthal:2013:DPP

[Blu13] Geoffrey Blumenthal. Diplomacy, patronage, and the preface to De revolutionibus. Journal for the History of Astronomy, 44(1):75–92, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Blum:2021:RGB

[Blu21] Paul Richrd Blum. A reading guide for Bruno’s On the Infinite. Journal for the History of Astronomy, 52 (3):366–367, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211022072.

See erratum [Ano22b].

Belmonte:2015:AAS

[BM15] Juan Antonio Belmonte and Giulio Magli. Astronomy, architecture, and symbolism: The global project of Sneferu at Dahshur. Journal for the History of Astronomy, 46(2):173– 205, May 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brosche:1990:SCV

[BMA90] Peter Brosche, Lucrecia Maupomé, and Anthony F. Aveni. The sacred calendar and Venus. Journal for the History of Astronomy, 21(15):S51–S55, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brunet:1981:DLC

[BN81] J.-P. Brunet and R. Nadal. Durées de lever et de coucher des signes du zodiaque. (French) [Times of sunrise and sunset of the zodiac signs]. Journal for the History of Astronomy, 12(3):178–194, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bogdan:2002:DMB

[Bog02] Thomas J. Bogdan. Donald Menzel and the beginnings of the high altitude observatory. Journal for the History of Astronomy, 33(2):157–192, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bohloul:2019:BRM

[Boh19] Hamid Bohloul. Book review: Medieval astrolabes: Astrolabes in Medieval Cultures, Josefina Rodríguez-Arribas, Charles Burnett, Silke Ackermann & Ryan Szpiech, (Brill, Leiden, 2019). Pp. vi + 508. EUR 87. ISBN 978-90-04-38380-7 (paper). Journal for the History of Astronomy, 50(3):373–374, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619866330.

Bohloul:2023:TSH

[Boh23]  Hamid Bohloul.  Toward

a standardization of Hay’a works. Journal for the History of Astronomy, 54(3):363– 366, August 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231191201.

Bok:1973:BRE

[Bok73] Bart J. Bok. Book review:

Education and history: Education in and History of Modern Astronomy, Annals of the New York Academy of Sciences. Journal for the History of Astronomy, 4(3):208– 209, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bolt:2005:BRG

[Bol05] Marvin Bolt. Book review:

Globes and spheres: Spheres: The Art of the Celestial Mechanic. Journal for the History of Astronomy, 36(4):453–454, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bolt:2010:ERI

[Bol10] Marvin Bolt. Essay review:

The invention of the telescope: The Long Route to the Invention of the Telescope. Journal for the History of Astronomy, 41(1):120–122, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bolt:2020:BRW

[Bol20] Marvin Bolt. Book review: William Herschel’s astronomical work. Journal for the History of Astronomy, 51(2): 249–251, May 1, 2020. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619885492.

Bonelli:1976:BRS

[Bon76] Maria Luisa Righini Bonelli. Book review: Scientific instruments: Scientific Instruments. Journal for the History of Astronomy, 7(3):208–210, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2005:SSK

[Bon05]  Patrick J. Boner.  Soul-

searching with Kepler: an analysis of anima in his astrology. Journal for the History of Astronomy, 36(1):7–20, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2007:KVE

[Bon07] Patrick J. Boner. Kepler V. The Epicureans: Causality, coincidence and the origins of the new star of 1604. Journal for the History of Astronomy, 38 (2):207–221, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2009:BRT

[Bon09a] Patrick J. Boner. Book review: Tertius interveniens in translation, Kepler’s Astrology: The Baby, the Bath Water, and the Third Man in the Middle. Journal for the History of Astronomy, 40(1):128–130, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2009:TTP

[Bon09b] Patrick J. Boner. A tenuous tandem: Patrizi and Kepler on the origins of stars. Journal for the History of Astronomy, 40(4):381–391, November 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2011:BRA

[Bon11] Patrick J. Boner. Book review: Astrology since antiquity: A History of Western Astrology. Journal for the History of Astronomy, 42(2):276– 278, May 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2012:BRC

[Bon12] Patrick J. Boner. Book review:

The controversy on sunspots: On Sunspots. Journal for the History of Astronomy, 43(1):

131–132, February 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boner:2016:KQW

[Bon16a]  Patrick J. Boner.  Kepler

and the quest for world harmony. Journal for the History of Astronomy, 47 (4):431–433, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616665947.

Boner:2016:MEM

[Bon16b]  Patrick J. Boner.  The

Moon in early modern imagination. Journal for the History of Astronomy, 47 (4):433–434, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668740.

Boner:2019:BRBb

[Bon19a] Patrick J. Boner. Book review: Between celestial and terrestrial harmony. Journal for the History of Astronomy, 50(2):255–257, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618817137.

Boner:2019:BRBa

[Bon19b] Patrick J. Boner. Book review: Between celestial and terrestrial harmony: The Pursuit of Harmony: Kepler on Cosmos, Confession, and Community Aviva Rothman, (University of Chicago Press, Chicago, IL, 2017). Pp. viii + 355. $55. ISBN 978-0-226-49697-9. Journal for the History of Astronomy, 50(1): 101–103, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618811732.

Borgontoz:1998:OGL

[Bor98] Alfonso López Borgoñtoz. Orientations of graves in the late Roman necropolises of Ampurias. Journal for the History of Astronomy, 29(23):–, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boschiero:2009:GBC

[Bos09] Luciano Boschiero. Giovanni Borelli and the comets of 1664– 65. Journal for the History of Astronomy, 40(1):11– 30, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Botley:1976:PSS

[Bot76] Cicely M. Botley. The position of Supernova 1006 and the St. Gallen Chronicle. Journal for the History of Astronomy, 7(2):139–140, June 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bourtembourg:2012:MMD

[Bou12] René Bourtembourg. Messier’s missed discovery of Pallas in April 1779. Journal for the History of Astronomy, 43(2):209– 214, May 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bourtembourg:2013:WUO

[Bou13]  René Bourtembourg.  Was

Uranus observed by Hipparchus? Journal for the History of Astronomy, 44(4): 377–387, November 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Boudet:2023:TEI

[Bou23] Jean-Patrice Boudet. Two editions of an Italian translation of Ps.-Ptolemy’s Centiloquium. Journal for the History of Astronomy, 54(1): 110–111, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221130756.

Bowden:1983:BRK

[Bow83] Mary Ellen Bowden. Book review: Kepler as astrologer: Kepler astronome astrologue. Journal for the History of Astronomy, 14(3):223–225, October 1983. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Bowen:2006:GLM

[Bow06] Alan C. Bowen. Geminus and the length of the month: The authenticity of Intro. Ast. 8.43– 45. Journal for the History of Astronomy, 37(2):193–202, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bowler:2010:BRM

[Bow10] Peter J. Bowler. Book review:

Myths about science and religion: Galileo Goes to Jail and other Myths about Science and Religion. Journal for the History of Astronomy, 41(2):278– 279, May 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bracher:1990:BRB

[Bra90] Katherine Bracher. Book review: On E. E. Barnard and other matters, Interstellar Matters: Essays on Curiosity and Astronomical Discovery. Journal for the History of Astronomy, 21(3):308–309, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brashear:2005:BRM

[Bra05] Ronald Brashear. Book review: More on Venus transits: The Transits of Venus. Journal for the History of Astronomy, 36(4):459–460, November 2005.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brandt:2015:BRH

[Bra15]  Marisa Brandt.  Book re-

view: Hawking: The man and the corporation. Hawking Incorporated: Stephen Hawking and the Anthropology of the Knowing Subject. Hélene Mialet (University of Chicago Press, Chicago, IL, 2012). Pp.

266. $31.00 (paperback). ISBN 978-0-226-52228-9. Journal for the History of Astronomy, 46 (2):250–252, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brannon:2020:BDV

[Bra20] James Brannon. Bi-daily Venus in the medieval thought of William of Conches: Explaining an uncommon celestial event by circumsolarity. Journal for the History of Astronomy, 51(3):305–323, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620932701.

Brannon:2023:PPO

[Bra23] James Brannon. Plato and planetary order: Uncertainty in the positions of Mercury and Venus. Journal for the History of Astronomy, 54(4): 404–424, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231203198.

Brenni:2000:BRC

[Bre00] Paolo Brenni. Book review:

Cathedral Meridian lines: The Sun in the Church: Cathedrals as Solar Observatories. Journal for the History of Astronomy, 31(2):168–170, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brentjes:2011:BRC

[Bre11]  Sonja Brentjes.  Book re-

view: Cross-cultural exchanges in the Mediterranean: Cross-Cultural Scientific Exchanges in the Eastern Mediterranean, 1560–1660. Journal for the History of Astronomy, 42(3): 411–415, August 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brenni:2013:BRN

[Bre13]  Paolo Brenni.  Book re-

view: a Nineteenth-Century clockmaker, Heinrich Johann Kessels (1781–1849): Ein bedeutender Verfertiger von Chronometern und Präzisionspendeluhren. Journal for the History of Astronomy, 44(2):228–229, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brenni:2017:BWB

[Bre17] Paolo Brenni. Building the world’s best telescopes in Bavaria. Journal for the History of Astronomy, 48 (1):129–131, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617693514.

Brenni:2018:KIA

[Bre18] Paolo Brenni. Klingenstierna and the invention of the achromatic lens: review of The Invention of the Achromatic and Aplanatic Lens: With Special Regard to the Role Played by Samuel Klingenstierna N. V. E. Nordenmark & J. Nordström, edited by Roger C. Ceragioli, translated by Elisabeth Goodwin (Antique Telescope Society, Inc., Plymouth Meeting, PA, 2016). Pp. 142. US$25. Journal for the History of Astronomy, 49(3):392–393, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Britton:1998:ERP

[Bri98] John P. Britton. Essay review: Planetary theory in Babylon: The Babylonian Theory of the Planets. Journal for the History of Astronomy, 29(4):381– 385, November 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bricker:2001:MDV

[Bri01a] Victoria R. Bricker. A method for dating Venus almanacs in the Borgia Codex. Journal for the History of Astronomy, 32 (26):S21–S44, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Britton:2001:BRL

[Bri01b] John P. Britton. Book review:

Light on the East: Astral Sciences in Mesopotamia. Journal for the History of Astronomy, 32(2):169–170, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Britton:2005:BRB

[Bri05] John P. Britton. Book review:

Babylonian eclipse records: Babylonian Eclipse Observations from 750 BC to 1 BC. Journal for the History of Astronomy, 36(2):235–236, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooke:1983:BRO

[Bro83] John Hedley Brooke. Book review: Other worlds than ours: Plurality of Worlds: The Extraterrestrial Life Debate from Democritus to Kant. Journal for the History of Astronomy, 14(1):65–67, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Broughton:1985:FPR

[Bro85] Peter Broughton. The first predicted return of comet Halley. Journal for the History of Astronomy, 16(2):123–133, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooks:1988:EMS

[Bro88a] Randall C. Brooks. Errors in measurement of the solar diameter in the Seventeenth and Eighteenth centuries. Journal for the History of Astronomy, 19(4):239–255, November 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Broughton:1988:ASM

[Bro88b]  Peter Broughton.  Arthur

Storer of Maryland: His astronomical work and his family ties with Newton. Journal for the History of Astronomy, 19(2):77–96, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brotherston:1990:BRE

[Bro90] Gordon Brotherston. Book review: Empires of Time: Calendars, Clocks, and Cultures,A. F. Aveni (Basic Books, New York, 1989). Pp. 371. $24.95. Journal for the History of Astronomy, 21(15):S61–S62, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Broda:1991:ERA

[Bro91a] Johanna Broda. Essay review: Archaeoastronomy in the Americas: New directions in American archaeoastronomy. Journal for the History of Astronomy, 22(16):S84– S88, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooks:1991:DMS

[Bro91b] Randall C. Brooks. The development of micrometers in the Seventeenth, Eighteenth and Nineteenth Centuries. Journal for the History of Astronomy, 22(2):127–173, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooke:1996:ERM

[Bro96] John Hedley Brooke. Essay review: Matters of fact and Faith: Galileo: For Copernicanism and for the Church, Galileo: a Life, Galileo: Decisive Innovator. Journal for the History of Astronomy, 27 (1):68–74, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooks:2010:BRD

[Bro10] Randall C. Brooks. Book review: Documenting the U.S. space program: Exploring the Unknown: Selected Documents in the History of the U.S. Civil Space Program, VII: Human Spaceflight: Projects Mercury, Gemini and Apollo. Journal for the History of Astronomy, 41(1):142–143, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brooke:2015:RLS

[Bro15] John Hedley Brooke. Reassessing Lemaître’s science and religion. Journal for the History of Astronomy, 46(1):102–103, February 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:1978:BRN

[Bru78a] Stephen G. Brush. Book review: The nebular hypothesis in America: Creation by Natural Law: Laplace’s Nebular Hypothesis in American Thought. Journal for the History of Astronomy, 9(1):70–71, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:1978:GAAa

[Bru78b] Stephen G. Brush. A geologist among astronomers: The rise and fall of the Chamberlin– Moulton cosmogony, Part 1. Journal for the History of Astronomy, 9(1):1–41, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:1978:GAAb

[Bru78c] Stephen G. Brush. A geologist among astronomers: The rise and fall of the Chamberlin– Moulton cosmogony, Part 2. Journal for the History of Astronomy, 9(2):77–104, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:1982:BRG

[Bru82] Stephen G. Brush. Book review: G. K. Gilbert: Grove Karl Gilbert. Journal for the History of Astronomy, 13(1): 71–72, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bruck:1987:BRL

[Brü87]  H. A. Brück.  Book re-

view: Letters of Zach: Briefe Franz Xaver von Zach’s in Sein Vaterland. Journal for the History of Astronomy, 18(2):136– 137, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:1996:BRS

[Bru96] Stephen G. Brush. Book review: Solar system astronomy before Sputnik: Solar System Astronomy in America: Communities, Patronage, and Interdisciplinary Research, 1920– 1960. Journal for the History of Astronomy, 27(4):368–370, November 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Brush:2009:BRD

[Bru09] Stephen G. Brush. Book review: Determining the age of the Earth, A Natural History of Time. Journal for the History of Astronomy, 40(3):360– 361, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bryden:1972:BFO

[Bry72] D. J. Bryden. Britain’s first observatory? Journal for the History of Astronomy, 3(3):205, October 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bryden:1981:BRA

[Bry81] D. J. Bryden. Book review: Astronomical clockwork: Geared to the Stars. Journal for the History of Astronomy, 12(3): 207–209, October 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Berendzen:1973:BRP

[BS73] Richard Berendzen and Daniel Seeley. Book review: Popularising history of science: The Mind of the Scientist. Journal for the History of Astronomy, 4 (1):64–65, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Baum:1992:GDC

[BS92] Richard Baum and William Sheehan. G. D. Cassini and the rotation period of Venus: a common misconception. Journal for the History of Astronomy, 23(4):299–301, November 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2006:OAE

[BS06] Juan Antonio Belmonte and Mosalam Shaltout. On the orientation of Ancient Egyptian temples: (2) new experiments at the oases of the Western Desert. Journal for the History of Astronomy, 37(2):173– 192, May 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Belmonte:2008:OAE

[BSF08] Juan Antonio Belmonte, Mosalam Shaltout, and Magdi Fekri. On the orientation of Ancient Egyptian temples: (4) epilogue in Serabit El Khadim and overview. Journal for the History of Astronomy, 39 (2):181–212, May 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bretagnon:1986:PNS

[BSL86] P. Bretagnon, J.-L. L. Simon, and J. Laskar. Presentation of new solar and planetary tables of interest for historical calculations. Journal for the History of Astronomy, 17(1): 39–50, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burnette:1980:BRD

[Bur80] Mark Burnette. Book review: Dee on celestial influences, John Dee on Astronomy. Journal for the History of Astronomy, 11(3):205–207, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burl:1982:BRM

[Bur82] Aubrey Burl. Book review: Megalithic science: Ancient Mathematics and Astronomy in North-West Europe. Journal for the History of Astronomy, 13(4):S71–S72, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burnett:1984:BRZ

[Bur84] Charles S. F. Burnett. Book review: The zodiac in antiquity: Die Eigenschaften der Tierkreiszeichen in der Antike. Journal for the History of Astronomy, 15(1):48–50, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burns:2010:BRT

[Bur10] Elizabeth Burns. Book review: Theories of comets: A History of Physical Theories of Comets: From Aristotle to Whipple. Journal for the History of Astronomy, 41(3):415– 416, August 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burnett:2015:BRG

[Bur15] Charles Burnett. Book review: The Gerland computus. Der Computus Gerlandi: Edition, Übersetzung und Erläuterungen. Alfred Lohr (Sudhoffs Archiv Beihefte, lxi; Franz Steiner Verlag, Stuttgart, 2013). Pp. 493 + CD-ROM. EUR 74. ISBN 978-3-515-10468-5 (paper). Journal for the History of Astronomy, 46(2):239–240, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Burnett:2020:BRA

[Bur20] Charles Burnett. Book review: Astrological weather forecasting. Journal for the History of Astronomy, 51(2): 248–249, May 1, 2020. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620921985.

Byrne:2011:MDS

[Byr11] James Steven Byrne. The mean distances of the Sun and commentaries on the theorica planetarum. Journal for the History of Astronomy, 42(2):205– 221, May 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cahan:2019:BRP

[Cah19]  David Cahan.  Book re-

view: From personal observatory to national standards testing. Journal for the History of Astronomy, 50 (2):258–260, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619834394.

Campion:2002:BRR

[Cam02] Nick Campion. Book review: Reading a painting, The Ambassadors’ Secret: Holbein and the World of the Renaissance. Journal for the History of Astronomy, 33(3):294–295, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Campion:2011:BRA

[Cam11] Nicolas Campion. Book review: Astrology in Medieval Ireland and Wales: Fiery Shapes: Celestial Portents and Astrology in Ireland and Wales 700– 1700. Journal for the History of Astronomy, 42(3):419–421, August 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Canales:2009:BRR

[Can09] Jimena Canales. Book review: Rehabilitating Janssen, Un Globe-Trotter de la Physique Céleste: L’astronome Jules Janssen. Journal for the History of Astronomy, 40(2):233– 235, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Canadelli:2012:BRA

[Can12] Elena Canadelli. Book review: Arcetri Observatory, Giovanni Virginio Schiaparelli e l’Osservatorio di Arcetri. Journal for the History of Astronomy, 43(3):371–372, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Caplan:2012:BRS

[Cap12] James Caplan. Book review:

The story of the telescope, Histoire du Télescope: La Contemplation de Léunivers, des Premiers Instruments aux Actuelles Machines Célestes. Journal for the History of Astronomy, 43(2):256–257, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carlson:1976:BRA

[Car76] John B. Carlson. Book review: Archaeoastronomy in America: Archaeoastronomy in Pre-Columbian America. Journal for the History of Astronomy, 7(3):205–206, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carlson:1986:BRC

[Car86] John B. Carlson. Book review:

The Cosmic Serpent: A Catastrophist View of Earth History, Victor Clube and Bill Napier (Faber and Faber, London, 1982; Universe, New York). Pp. 299. £12.50/$17.95. Journal for the History of Astronomy, 17(10):S70–S72, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cardenas:1988:BRS

[Cár88] Anthony J. Cárdenas. Book review: The symposium on Alfonsine astronomy: De Astronomia Alphonsi Regis. Journal for the History of Astronomy, 19(2):140–141, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carroll:1997:BRH

[Car97] William E. Carroll. Book review: How the heavens go: Galileo, Bellarmine, and the Bible. Journal for the History of Astronomy, 28(3):272– 274, August 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carroll:1999:BRG

[Car99] William E. Carroll. Book review: Galileo’s dialogo: Galileo on the World Systems: a new abridged translation and guide. Journal for the History of Astronomy, 30(1):73–74, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carlson:2000:CRO

[Car00] John B. Carlson. Conference report: Oxford 6. Journal for the History of Astronomy, 31 (25):S89–S90, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carolino:2008:MTC

[Car08] Luís Miguel Carolino. The making of a Tychonic cosmology: Cristoforo Borri and the development of Tycho Brahe’s astronomical system. Journal for the History of Astronomy, 39(3):313–344, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carman:2017:FDA

[Car17a] Christián C. Carman. The final date of the Antikythera mechanism. Journal for the History of Astronomy, 48(3):312–323, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617721553.

Carter:2017:SCC

[Car17b] Christopher Carter. “I still continue convinced”: Expecting the great comet of 1848. Journal for the History of Astronomy, 48(2):180–206, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617702714.

Carman:2019:PDP

[Car19] Christián C. Carman. A possible date for Ptolemy’s development of a model for the second lunar anomaly. Journal for the History of Astronomy, 50(4):398–410, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619871723.

Carman:2020:LML

[Car20] Christián C. Carman. Longomontanus’ model for the longitudes of Mars. Journal for the History of Astronomy, 51(2):162–186, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620923802.

Carman:2021:TMC

[Car21] Christián C. Carman. The Tychonic method for calculating the ratio between the eccentricities of Mars. Journal for the History of Astronomy, 52(2):212–234, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211014311.

Carman:2023:PLM

[Car23] Christián C. Carman. A Ptolemaic lunar model of the 17th century: François Viète and his first lunar model. Journal for the History of Astronomy, 54(4):375–403, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231210546.

Carman:2024:FVH

[Car24] Christián C. Carman. François Viète and his analysis of the Copernican lunar model. Journal for the History of Astronomy, 55(3):314–331, August 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241264614.

Carter:2025:SIC

[Car25] Christopher Carter. Stella Insolita: The comet of 1114, a lost chronicle and the Empress Matilda. Journal for the History of Astronomy, 56(1):63–71, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241294046.

Case:2015:GE

[Cas15] Stephen Case. Galactic encounters. Journal for the History of Astronomy, 46(4):497– 498, November 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Case:2016:RTR

[Cas16] Stephen Case. Restoring a tarnished reputation. Journal for the History of Astronomy, 47(4):443–444, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616668750.

Case:2018:MSP

[Cas18] Stephen Case. Making Stars Physical: The Astronomy of Sir John Herschel. University of Pittsburgh Press, Pittsburgh, PA, USA, 2018. ISBN 0-8229-4530-4. viii + 319 pp. LCCN QB36.H59 C37 2018. URL https:

                                        //www.jstor.org/stable/j.
                                        ctv7n0c4q.

Case:2019:CSJ

[Cas19] Stephen Case. A “confounded scrape”: John Herschel, Neptune, and naming the satellites of the outer solar system. Journal for the History of Astronomy, 50(3):306–325, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619863866.

Case:2021:ADS

[Cas21] Stephen Case. Amici’s double star observations. Journal for the History of Astronomy, 52(4):491–492, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211032417.

Caudano:2008:BRC

[Cau08] Anne-Laurence Caudano. Book review: Carolingian astronomy; Ordering the Heavens: Roman Astronomy and Cosmology in the Carolingian Renaissance. Journal for the History of Astronomy, 39(2):275– 277, May 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Caudano:2017:BRF

[Cau17] Anne-Laurence Caudano. Book review: The First Byzantine Astronomical Text. Journal for the History of Astronomy, 48(4):491–492, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617719386.

Caudano:2025:PTK

[Cau25] Anne-Laurence Caudano. Ptolemy’s Table of Kings. Journal for the History of Astronomy, 56 (1):113–115, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241275233.

Clutton-Brock:2005:CPH

[CB05] Martin Clutton-Brock. Copernicus’s path to his cosmology: an attempted reconstruction. Journal for the History of Astronomy, 36(2):197–216, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cenadelli:2014:BRS

[Cen14] Davide Cenadelli. Book review: The Secchi–Tacchini correspondence: Alle Origini dell Astrofisica Italiana: Il Carteggio Secchi Tacchini 1861 1877. Ileana Chinnici and Antonella Gasperini (Fondazione “Giorgio Ronchi”, cxvi; Edizioni Tassinari, Florence, 2013). Pp.

496. EUR 57 (paperback). ISBN 978-88-88649-38-2. Journal for the History of Astronomy, 45(4):506–507, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cenadelli:2017:EIA

[Cen17]  Davide Cenadelli.  Early

Italian astrophysics on exhibit. Essay review: Starlight: La Nascita dell’Astrofisica in Italia: The Origins of Astrophysics in Italy. Ileana Chinnici (Prismi Editrice Politecnica, Naples, 2016). Pp. 127. EUR 25. ISBN 978-88-569-0522-9 (paper). Journal for the History of Astronomy, 48(3):369–371, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617713525.

Ceragioli:1995:DCR

[Cer95] R. C. Ceragioli. The debate concerning ‘red’ Sirius. Journal for the History of Astronomy, 26(3):187–226, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ceragioli:1996:SPR

[Cer96] R. C. Ceragioli. Solving the puzzle of ‘red’ Sirius. Journal for the History of Astronomy, 27(2):93–128, May 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ceragioli:1997:BRH

[Cer97] Roger Ceragioli. Book review:

History and science in collaboration: Caesar’s Comet: The Comet of 44 B.C. and Caesar’s Funeral Games. Journal for the History of Astronomy, 28(4):348–350, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ceragioli:1999:BRS

[Cer99] Roger C. Ceragioli. Book review: Star myths of antiquity: Star Myths of the Greeks and Romans: a Sourcebook. Journal for the History of Astronomy, 30(3):313–315, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cooke:1977:IDC

[CFMR77] J. A. Cooke, R. W. Few, J. G. Morgan, and C. L. N. Ruggles. Indicated declinations at the Callanish megalithic sites. Journal for the History of Astronomy, 8(2):113–133, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:1997:CAF

[CG97] José Chabás and Bernard R. Goldstein. Computational astronomy: Five centuries of finding true syzygy. Journal for the History of Astronomy, 28(2):93–105, May 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2003:JVA

[CG03] José Chabás and Bernard R. Goldstein. John Vimond and the Alfonsine trepidation model. Journal for the History of Astronomy, 34(2):163–170, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2009:JMT

[CG09] José Chabás and Bernard R. Goldstein. John of Murs’s tables of 1321. Journal for the History of Astronomy, 40(3): 297–320, August 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2010:AAP

[CG10] José Chabás and Bernard R.

Goldstein. Astronomical activity in Portugal in the Fourteenth Century. Journal for the History of Astronomy, 41 (2):199–212, May 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2012:JMR

[CG12] José Chabás and Bernard R.

Goldstein. John of Murs revisited: The Kalendarium Solis Et Lune for 1321. Journal for the History of Astronomy, 43(4):411–437, November 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2013:CPP

[CG13] José Chabás and Bernard R.

Goldstein. Computing planetary positions: User-friendliness and the Alfonsine Corpus. Journal for the History of Astronomy, 44(3):257–276, August 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigenda [Ano13a].

Chabas:2016:MOT

[CG16] José Chabás and Bernard R.

Goldstein. The Moon in the Oxford tables of 1348. Journal for the History of Astronomy, 47(2):159–167, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2018:AOT

[CG18] José Chabás and Bernard R. Goldstein. Adaptations of the Oxford Tables to Paris, Mantua, and Louvain. Journal for the History of Astronomy, 49(1):99–115, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617752698.

Chabas:2019:MDA

[CG19] José Chabás and Bernard R. Goldstein. The master and the disciple: The almanac of John of Lignères and the ephemerides of John of Saxony. Journal for the History of Astronomy, 50(1):82– 96, February 1, 2019. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618820215.

Craig:1992:HSP

[CH92] Edward Craig and Michael Hoskin. Hegel and the seven planets. Journal for the History of Astronomy, 23(3):208– 210, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapin:1972:MBF

[Cha72] Seymour L. Chapin. “In a mirror brightly”: French attempts to build reflecting telescopes using platinum. Journal for the History of Astronomy, 3(2):87– 104, June 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chatterjee:1975:ERA

[Cha75] Bina Chatterjee. Essay review: Astronomy in India: Astronomie Indienne: Publications. Journal for the History of Astronomy, 6(1):65–66, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapman:1983:AAM

[Cha83] Allan Chapman. The accuracy of angular measuring instruments used in astronomy between 1500 and 1850. Journal for the History of Astronomy, 14(2):133–137, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapin:1985:BRR

[Cha85] Seymour L. Chapin. Book review: Roemer’s place in history: Roemer et la Vitesse de la Lumière. Journal for the History of Astronomy, 16(1): 61–63, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapman:1988:PRP

[Cha88] Allan Chapman. Private research and public duty: George Biddell Airy and the search for Neptune. Journal for the History of Astronomy, 19(2):121– 139, May 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapin:1989:BRJ

[Cha89] Seymour L. Chapin. Book review: Jean Picard commemorated, Jean Picard et les Débuts de l’Astronomie de Précision au XVIIe Siècle. Journal for the History of Astronomy, 20 (2):143–145, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chapin:1990:VSI

[Cha90] Seymour L. Chapin. The vicissitudes of a scientific institution: A Decade of Change at the Paris Observatory. Journal for the History of Astronomy, 21(3):235–274, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:1998:ASM

[Cha98]  José Chabás.  Astronomy

in Salamanca in the mid-Fifteenth Century: the Tabulae Resolutae. Journal for the History of Astronomy, 29(2):167– 175, May 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2000:BRM

[Cha00] José Chabás. Book review: a manual of Byzantine astronomy: Georges Gémiste Pléthon: Manuel d’astronomie.

Journal for the History of Astronomy, 31(1):68–69, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Charbonneau:2002:RFF

[Cha02] Paul Charbonneau. The rise and fall of the first solar cycle model. Journal for the History of Astronomy, 33(4):351–372, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2007:TVA

[Cha07] José Chabás. From Toledo to Venice: The Alfonsine Tables of Prosdocimo De’ Beldomandi of Padua (1424). Journal for the History of Astronomy, 38 (3):269–281, August 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2009:BRA

[Cha09] José Chabás. Book review:

Arabic Astrology Enters the West, Les Alchandreana Primitifs: Étude sur les Plus Anciens Traités Astrologiques Latins d’Origine Arabe (Xe Siècle). Journal for the History of Astronomy, 40(3):348–350, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2012:CTM

[Cha12] José Chabás. Characteristics and typologies of medieval astronomical tables. Journal for the History of Astronomy, 43 (3):269–286, August 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2015:LSC

[Cha15a]  José Chabás.  A list of

stars “Correcte Cum 2 Magnis Armillis” in 1362. Journal for the History of Astronomy, 46(2):206–217, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chase:2015:BRW

[Cha15b] Stephen Chase. Book review: William Parsons, 3rd Earl of Rosse: Astronomy and the Castle in Nineteenth-Century Ireland, Edited by Charles Mollan (Manchester University Press, Manchester, 2014). Pp. xxii + 368. £70. ISBN 978-0-7190-9144-5. Journal for the History of Astronomy, 46(3): 381–382, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:2017:EWA

[Cha17] José Chabás. An early witness of Alfonsine astronomy: The London tables for 1336. Journal for the History of Astronomy, 48(3):324–328, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617716556.

Chabas:2019:BRC

[Cha19] José Chabás. Book review: a Castilian translation of Copernicus’ De revolutionibus orbium coelestium: Ydea astronomica de la fabrica del mundo y movimiento de los cuerpos celestiales (Traducción de “De revolutionibus I–III” de Nicolás Copérnico) Juan Cedillo, Miguel Ángel Granada & Félix Gómez Crespo, (Edicions de la Universitat de Barcelona, Barcelona, 2019). Pp. 482. EUR 30. ISBN 978-84-9168-164-9. Journal for the History of Astronomy, 50 (3):378–380, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619864725.

Chassefiere:2021:OEF

[Cha21] Eric Chassefière. Obstacles encountered by four major European astronomical observatories belonging to academies in the 18th century. Journal for the History of Astronomy, 52(4):414–441, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211052194.

Chenakal:1972:AIJ

[Che72] Valentin L. Chenakal. The astronomical instruments of John Rowley in Eighteenth-Century Russia. Journal for the History of Astronomy, 3(2):119– 135, June 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chenakal:1973:JBH

[Che73] Valentin L. Chenakal. John Bradlee and his sundials. Journal for the History of Astronomy, 4(3):159–167, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chevalier:1999:ODS

[Che99] Yves Chevalier. Orientations of 935 dolmens of Southern France. Journal for the History of Astronomy, 30(24):S47–S82, February 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chinnici:2009:RBR

[Chi09] Ileana Chinnici. The relationship between the Ramsden circles at Palermo and Dunsink. Journal for the History of Astronomy, 40(3):321–333, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chinnici:2015:NCC

[Chi15] Ileana Chinnici. Nineteenth-Century comets: Studies and observations in Sicily. Journal for the History of Astronomy, 46(2):130–158, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chinnici:2016:SND

[Chi16] Ileana Chinnici. Starlight: la nascita dell’astrofisica in Italia = the origins of astrophysics in Italy. Arte’m, Napoli, Italy, 2016. ISBN 88-569-0522-1. 127 pp. LCCN QB461 .S73 2016. 25.00 EUR. URL http:

//www.loc.gov/catdir/toc/
casalini15/3117492.pdf.

Chinnici:2021:ADL

[Chi21] Ileana Chinnici. Astronomical dialogues with learned ladies. Journal for the History of Astronomy, 52(2):239– 240, May 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828621998451.

Christianson:1998:BRF

[Chr98] Gale E. Christianson. Book review: The first Barnard biography: The Immortal Fire Within: The Life and Work of Edward Emerson Barnard. Journal for the History of Astronomy, 29(1):84–86, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Connor:2005:BRK

[CJ05] James A. Connor and Petra S.

Jung. Book review: Kepler the philosopher: Johannes Kepler. Journal for the History of Astronomy, 36(3):345–346, August 2005. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:2020:CAH

[CJ20] Christopher Cullen and Catherine Jami. Christmas 1668 and after: How Jesuit astronomy was restored to power in Beijing. Journal for the History of Astronomy, 51 (1):3–50, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620901887.

Cullen:2022:PPB

[CJ22] Christopher Cullen and Catherine Jami. Prediction and politics in Beijing, 1668: a Jesuit astronomer and his technical resources in a time of crisis. Journal for the History of Astronomy, 53(4): 422–474, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221114093.

Chabas:2007:CAP

[CK07] José Chabás and Richard L. Kremer. The circulation of astronomical practices in late medieval Europe. Journal for the History of Astronomy, 38(3): 267–268, August 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Clausen:2011:ODP

[CKE11] Claus Clausen, Per Kjærgaard, and Ole Einicke. The orientation of Danish passage graves on the islands of Samsø and Zealand. Journal for the History of Astronomy, 42(3):339– 351, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Caplan:2005:BRA

[CL05] James Caplan and Françoise Le Guet Tully. Book review: Astronomy at altitude: L’Observatoire du Pic du Midi: Cent Ans de Vie et de Science en Haute Montagne. Journal for the History of Astronomy, 36(2):243–245, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Caplan:2008:BRM

[CL08] James Caplan and Françoise Le Guet Tully. Book review: The Meudon refractor; La Grande Lunette de Meudon: Les Yeux de la Découverte. Journal for the History of Astronomy, 39(1):131–133, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chen:2022:API

[CL22] Ting Chen and Lingfeng Lü.

Astronomical or political: Interpretation of comets in times of crisis in qing China. Journal for the History of Astronomy, 53(1):13–26, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211070275.

Clemence:1971:CET

[Cle71] G. M. Clemence. The concept of ephemeris time: a case of inadvertent plagiarism. Journal for the History of Astronomy, 2(2):73–79, June 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Catamo:2000:FFG

[CLL^{+}00] Mario Catamo, Nicoletta Lanciano, Kurt Locher, Manuel Lombardero, and Manuel Valdés. Fifteen further Greco– Roman sundials from the Mediterranean area and Sudan. Journal for the History of Astronomy, 31(3):203–221, August 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chen-Morris:2010:BRK

[CM10] Raz Chen-Morris. Book review: Kepler’s work in context: Johannes Kepler, from Tübingen to Żagań. Journal for the History of Astronomy, 41(4):516–518, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chen-Morris:2017:HCO

[CM17] Raz Chen-Morris. The historical continuity of optics. Journal for the History of Astronomy, 48(1):122–124, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616681217.

Chen-Morris:2023:NSO

[CM23] Raz Chen-Morris. New stars, old cosmologies in early modern Europe. Journal for the History of Astronomy, 54 (3):360–362, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231193712.

Choi:2023:ICM

[CML23] Go-Eun Choi, Byeong-Hee Mihn, and Ki-Won Lee. Investigating calendrical methods of calculating sunrise and sunset times in the Shixian calendar. Journal for the History of Astronomy, 54 (3):251–272, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231171855.

Crowther:2015:BER

[CMM^{+}15] Kathleen Crowther, Ashley Nicole McCray, Leila Mc-Neill, Amy Rodgers, and Blair Stein. The book everybody read: Vernacular translations of Sacrobosco’s sphere in the Sixteenth Century. Journal for the History of Astronomy, 46 (1):4–28, February 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cunningham:2011:GPC

[CMO11] Clifford J. Cunningham, Brian G. Marsden, and Wayne Orchiston. Giuseppe Piazzi: The controversial discovery and loss of Ceres in 1801. Journal for the History of Astronomy, 42(3): 283–306, August 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cidambi:2024:AHC

[CMP24] Sahana Cidambi, Clemency Montelle, and Kim Plofker. Astronomical handbooks in 16thcentury South Asia: Analysis of mean planetary motions in the 1520 Graha-laghava of Gan.eśa Daivajña. Journal for the History of Astronomy, 55(4):405–453, November 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286241241812.

Cunningham:2013:OPE

[CO13] Clifford J. Cunningham and Wayne Orchiston. Olbers’s planetary explosion hypothesis: Genesis and early Nineteenth-Century interpretations. Journal for the History of Astronomy, 44(2):187–205, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Coggins:2008:BRA

[Cog08]  Clemency Coggins.  Book

review: Astronumerology at Palenque: The Apotheosis of Janaab’ Pakal: Science, History, and Religion at Classic Maya Palenque. Journal for the History of Astronomy, 39 (3):406–408, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1979:BRJ

[Coh79] I. Bernard Cohen. Book review: John Flamsteed’s lectures, The Gresham Lectures of John Flamsteed. Journal for the History of Astronomy, 10 (1):61–63, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1980:BRN

[Coh80] I. Bernard Cohen. Book review:

Newton and Flamsteed, the correspondence of Isaac Newton, The Correspondence of Isaac Newton. Journal for the History of Astronomy, 11(3): 213–214, October 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1986:BRE

[Coh86a] I. Bernard Cohen. Book review: Ephemerides for 1650– 1805: Planetary, Lunar, and Solar Positions, New and Full Moons, A.D. 1650–1805. Journal for the History of Astronomy, 17(3):207–209, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1986:OER

[Coh86b] I. Bernard Cohen. Obituary: Edward Rosen (1906– 1985). Journal for the History of Astronomy, 17(2):150– 152, May 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1991:BRE

[Coh91] I. Bernard Cohen. Book review: Essays on revolutions, Revolutions in Science: Their Meaning and Relevance. Journal for the History of Astronomy, 22(3):243, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1995:ERD

[Coh95] H. Floris Cohen. Essay review: Divinity at work: The Music of the Heavens: Kepler’s Harmonic Astronomy. Journal for the History of Astronomy, 26 (2):165–167, May 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1997:BRF

[Coh97] I. Bernard Cohen. Book review: Flamsteed correspondence: The Correspondence of John Flamsteed, First Astronomer Royal. Journal for the History of Astronomy, 28 (2):172–174, May 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:1999:BRF

[Coh99] I. Bernard Cohen. Book review:

Flamsteed under pressure: The Correspondence of John Flamsteed, First Astronomer Royal, II: 1682–1703. Journal for the History of Astronomy, 30(3): 325–327, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Conti:1970:NCR

[Con70] Giovanna Grassi Conti. A new catalogue of rare XVthand XVIth-Century astronomical books contained in the libraries of European observatories. Journal for the History of Astronomy, 1(2):165–166, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Connors:2024:DNT

[Con24] Martin Connors. The discovery and naming of Trojan asteroids. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231221166.

Cook:1975:BRR

[Coo75] Allan F. Cook. Book review: Research under difficulties: Find a Falling Star. Journal for the History of Astronomy, 6(2):140–141, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cook:1984:EEH

[Coo84] A. H. Cook. The election of Edmond Halley to the Savilian Professorship of Geometry. Journal for the History of Astronomy, 15(1):34–36, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cook:1997:BRM

[Coo97] Alan Cook. Book review: Modern planetary physics: A History of Modern Planetary Physics. Journal for the History of Astronomy, 28(4):357– 359, November 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cook:1999:SMM

[Coo99] David Cook. A survey of Muslim material on comets and meteors. Journal for the History of Astronomy, 30(2):131–160, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cook:2000:BRU

[Coo00] Alan Cook. Book review: Understanding the tides: Tides: a Scientific History. Journal for the History of Astronomy, 31 (1):69–70, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

CorballySJ:1998:BRS

[Cor98] Christopher J. Corbally, S.J.

Book review: ‘SETI’ and the limits of science: The Biological Universe: The Twentieth-Century Extratesrrestrial Life Debate and the Limits of Science. Journal for the History of Astronomy, 29(1):88– 89, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Corbin:2003:ERC

[Cor03] Brenda Corbin. Essay review: Caroline Herschel’s perspective: The Herschel Partnership: As Viewed by Caroline, Caroline Herschel’s Autobiographies. Journal for the History of Astronomy, 34(3): 328–331, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Courtney:2018:HMA

[Cou18] Stephen Courtney. The historical meridian: Antiquity and scripture in the public work of George Biddell Airy. Journal for the History of Astronomy, 49(2):135–157, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618773173.

Covington:2015:HLA

[Cov15] Michael A. Covington. Hodierna’s Luminosa 8 = ψ^{1,2,3} Aquarii. Journal for the History of Astronomy, 46(3):365–366, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Covington:2017:WIA

[Cov17] Michael A. Covington. Where, if anywhere, did Rheita see Veronica’s veil? Journal for the History of Astronomy, 48(2):233–237, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617708625.

Cowan:1971:BRP

[Cow71] Thaddeus M. Cowan. Book review: Prehistoric astronomers: Megalithic Lunar Observatories. Journal for the History of Astronomy, 2(3):202–203, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Coyne:2007:BRM

[Coy07] George V. Coyne. Book review: More on the Galileo affair: Galileo Observed: Science and the Politics of Belief. Journal for the History of Astronomy, 38(2):251–253, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chlench-Priber:2023:HML

[CP23] Kathrin Chlench-Priber. A handbook of medieval Latin astronomical tables. Journal for the History of Astronomy, 54(2):241–242, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231169116.

Chinnici:2011:TRP

[CR11] Ileana Chinnici and Donatella Randazzo. Tracing Ramsden’s ‘plumbline level’. Journal for the History of Astronomy, 42 (2):161–176, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carman:2022:TBA

[CR22] Christián C. Carman and Gonzalo L. Recio. Tycho Brahe’s Appendix ad Observationes anni 1593 and the date of Brahe’s theory of Mars, the prototype for Kepler’s vicarious hypothesis. Journal for the History of Astronomy, 53(3):239–265, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221099993.

Crelinsten:2006:EJR

[Cre06] Jeffrey Crelinsten. Einstein’s jury: the race to test relativity. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12310-1 (hardcover), 1-4008-4967-5 (e-book). xxix + 397 pp. LCCN QC173.585 .C74

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0654/2005032681-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0734/2005032681-
b.html;   http://www.loc.
gov/catdir/toc/ecip063/2005032681.
html.

Crowe:2000:BRE

[Cro00] Michael J. Crowe. Book review: Early observations of Jupiter: Galileo’s Planet: Observing Jupiter before Photography. Journal for the History of Astronomy, 31(4):355–356, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Crowe:2007:MAE

[Cro07] Michael J. Crowe. Mechanics from Aristotle to Einstein. Green Lion Press, Santa Fe, NM, USA, 2007. ISBN 1-888009-32-2 (paperback). xxii + 331 pp. LCCN QA802 .C76 2007. URL http:

//catdir.loc.gov/catdir/
toc/fy0803/2007925664.html.

Crowe:2008:BRH

[Cro08] Michael J. Crowe. Book review: The Herschel family: The Herschels of Hanover. Journal for the History of Astronomy, 39 (3):405–406, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Crowe:2012:BRP

[Cro12] Michael J. Crowe. Book review:

Planets of other stars: Strange New Worlds: The Search for Alien Planets and Life beyond Our Solar System. Journal for the History of Astronomy, 43(1):125–126, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Crossley:2015:BRF

[Cro15] Robert Crossley. Book review:

Flammarion on Mars in English: Camille Flammarion’s the Planet Mars (Astrophysics and Space Science Library, vol. cdix). Translated by Patrick Moore, Edited by William Sheehan (Springer, Heidelberg, 2015). Pp. xxii + 528. EUR 107. ISBN 978-3-319-09640-7. Journal for the History of Astronomy, 46(3):373–375, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Crowther:2023:BRP

[Cro23a] Kathleen Crowther. Book review: Printing the book everybody read: Publishing Sacrobosco’s De sphaera in Early Modern Europe: Modes of Material and Scientific Exchange. Edited by Matteo Valleriani and Andrea Ottone (Cham, Springer, 2022). Pp. xiv + 492. $60. ISBN 978-3-030-86599-3. Journal for the History of Astronomy, 54 (3):371–372, August 1, 2023.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231175992.

Crozier:2023:SAG

[Cro23b] William Crozier. St. Albert the Great and Robert Grosseteste on the nature and causes of comets. Journal for the History of Astronomy, 54(2):220–240, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231170596.

Clarke:2002:CGB

[CS02a] A. J. M. Clarke and J. M. Steele. A computer generated Babylonian system a lunar Ephemeris. Journal for the History of Astronomy, 33(3): 279, August 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cohen:2002:BRN

[CS02b] I. B. Cohen and G. E. Smith. Book review: Newton and the lunar motion, Newton’s Forgotten Lunar Theory: His Contribution to the Quest for Longitude. Journal for the History of Astronomy, 33(2):212, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cockcroft:2014:AED

[CS14] Robert Cockcroft and Sarah L. Symons. Ancient Egyptian diagonal star tables: a new fragment, and updates for S16C and S1C. Journal for the History of Astronomy, 45(2):197– 208, May 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Caballero-Sanchez:2021:CAC

[CS21] Raúl Caballero-Sánchez. On the chronology of the anonymous commentary to Ptolemy’s Tetrabiblos: Analysis of the astronomical evidence. Journal for the History of Astronomy, 52(4):442–461, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211052186.

Chapront-Touze:1986:MDD

[CT86] Michelle Chapront-Touzé. Les manuscrits Delambre du Bureau des Longitudes. (French) [The Delambre manuscripts of the Bureau of Longitude]. Journal for the History of Astronomy, 17(3):183–186, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Chabas:1993:VPP

[CT93] José Chabás and Anne Tihon. Verification of parallax in Ptolemy’s Handy Tables. Journal for the History of Astronomy, 24(1–2): 123–141, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/24/1-
                                        2/123.full.pdf+html.

Chapront-Touze:2012:BRH

[CT12] Michelle Chapront-Touzé. Book review: The Hill–Brown theory: The Hill–Brown Theory of the Moon’s Motion: Its Coming-to-Be and Short-Lived Ascendancy (1877–1984). Journal for the History of Astronomy, 43(1):134–136, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Carman:2012:PSD

[CTE12] Christián C. Carman, Alan Thorndike, and James Evans. On the pin-and-slot device of the Antikythera mechanism, with a new application to the superior planets. Journal for the History of Astronomy, 43(1):93–116, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:1993:MSC

[Cul93] Christopher Cullen. Motivations for scientific change in Ancient China: Emperor Wu and the grand inception astronomical reforms of 104 B.C. Journal for the History of Astronomy, 24(3):185–203, August 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:2002:FCC

[Cul02] Christopher Cullen. The first complete Chinese theory of the Moon: The innovations of Liu Hong c. A.D. 200. Journal for the History of Astronomy, 33 (1):21–39, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:2005:TAC

[Cul05] Christopher Cullen. Translating Ancient Chinese astronomical systems with Excel: How not to stew the strawberries? Journal for the History of Astronomy, 36(3):336–338, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:2007:HRO

[Cul07]  Christopher Cullen.  Huo

Rong’s observation programme of A.D. 102 and the Han Li solar table. Journal for the History of Astronomy, 38(1): 75–98, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Cullen:2017:FCR

[Cul17] Christopher Cullen. The foundations of celestial reckoning: three ancient Chinese astronomical systems. Scientific writings from the ancient and medieval world. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2017.

ISBN 1-138-10117-6 (hardcover), 1-315-65715-5 (e-book). xi + 434 pp. LCCN QB17 .C85 2017.

Cunningham:2016:DFA

[Cun16] Clifford J. Cunningham. Discovery of the first asteroid, Ceres: historical studies in asteroid research. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-21777-1 (e-book), 3-319-21776-3. LCCN QB378.C4. URL

http://link.springer.com/
10.1007/978-3-319-21777-2.

Cetera-Wlodarczyk:2017:DBE

[CWW17] Anna Cetera-Włodarczyk and Jarosław Włodarczyk. Digging (up) the bard. Essay review: Shakespeare and Saturn: Accounting for Appearances. Peter D. Usher (Peter Lang, New York, 2015). Pp. xxii + 244. $87. ISBN 978-1-4331-2860-8. Journal for the History of Astronomy, 48(3):372–375, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617715825.

DaSilva:2004:SFM

[Da 04] C. Marciano Da Silva. The spring full Moon. Journal for the History of Astronomy, 35 (4):475–478, November 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deAlbuquerque:1974:ERR

[dA74] Luis de Albuquerque. Essay review: Renaissance navigation: A Navigator’s Universe, Libro de Cosmographia. Journal for the History of Astronomy, 5(1): 54–56, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deAsua:2017:PMQ

[dA17] Miguel de Asúa. Los phisicos modernos quasi todos son copernicanos: Copernicanism and its discontents in colonial Río De La Plata. Journal for the History of Astronomy, 48(2):160–179, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617701210.

DallOlmo:1978:MMS

[Dal78] Umberto Dall’Olmo. Meteors, meteor showers and meteorites in the Middle Ages: From European medieval sources. Journal for the History of Astronomy, 9(2):123–134, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See erratum [Ano78b].

DallOlmo:1980:LTR

[Dal80] Umberto Dall’Olmo. Latin terminology relating to aurorae, comets, meteors and novae. Journal for the History of Astronomy, 11(1):10–27, February 1980. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Dallal:2001:ERA

[Dal01] Ahmad Dallal. Essay review: Arabic studies assembled, Arabic Mathematical Sciences: Instruments, Texts, Transmission, Astronomy and Astrology in the Medieval Islamic World. Journal for the History of Astronomy, 32(4):362–365, November 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2004:AGB

[Dan04a] Dennis Danielson. Achilles Gasser and the birth of Copernicanism. Journal for the History of Astronomy, 35(4):457– 474, November 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2004:BRC

[Dan04b] Dennis Danielson. Book review: Conducting the census: The Book Nobody Read: Chasing the Revolutions of Nicolaus Copernicus. Journal for the History of Astronomy, 35(3): 360–362, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2011:BRP

[Dan11] Dennis Danielson. Book review: In praise of Rheticus: Rheticus, Wegbereiter der Neuzeit (1514–1574). Journal for the History of Astronomy, 42(4):533–534, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2012:BRC

[Dan12a] Dennis Danielson. Book review: Copernicus in fact and fiction, A More Perfect Heaven: How Copernicus Revolutionized the Cosmos. Journal for the History of Astronomy, 43(3): 363–364, August 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2012:BRF

[Dan12b]  Dennis Danielson.  Book

review: Fictional tools in Seventeenth-Century science, fictions of the cosmos: Science and Literature in the Seventeenth Century. Journal for the History of Astronomy, 43 (3):364–366, August 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Danielson:2017:IAS

[Dan17] Dennis Danielson. Imagination, art and science. Journal for the History of Astronomy, 48(2):249–251, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617705016.

Danielson:2019:BRK

[Dan19a] Dennis Danielson. Book review: Kepler as poet. Journal for the History of Astronomy, 50(4):497–498, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619882216.

Danielson:2019:BRP

[Dan19b] Janet Danielson. Book review: Polyphonic minds: Music of the hemispheres: Polyphonic Minds: Music of the Hemispheres Peter Pesic, (The MIT Press, Cambridge MA, 2017). Pp. viii + 330. $38. ISBN 978-0-262-03691-7. Journal for the History of Astronomy, 50(1):103–105, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618817136.

Danielson:2022:PSC

[Dan22] Dennis Danielson. Poetic structures of the Cosmos. Journal for the History of Astronomy, 53(3):376–378, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221101128.

Darius:1991:BRM

[Dar91] Jon Darius. Book review: Modern physical science illustrated:

Album of Science: The Physical Sciences in the Twentieth Century. Journal for the History of Astronomy, 22(4): 326–327, November 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dash:2015:FTM

[Das15] Glen Dash. On the field testing of methods the Egyptians may have used to find cardinal directions. Journal for the History of Astronomy, 46(3):351– 364, August 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:1985:IAE

[Dav85] Virginia Lee Davis. Identifying Ancient Egyptian constellations. Journal for the History of Astronomy, 16(9):S102– S104, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2006:BRR

[Dav06] Edward B. Davis. Book review:

Religion and the rise of science: Religious Values & the Rise of Science in Europe. Journal for the History of Astronomy, 37(4):489–491, November 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2008:BRE

[Dav08] Edward B. Davis. Book review: Eddington the Quaker, Practical Mystic: Religion, Science, and A. S. Eddington. Journal for the History of Astronomy, 39(3):417–419, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2009:KOC

[Dav09] A. E. L. Davis. Kepler’s ‘Via Ovalis Composita’: Unity from diversity. Journal for the History of Astronomy, 40(1): 55–69, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2012:KAH

[Dav12] A. E. L. Davis. Kepler’s archetypes: How they increased the scale of the universe. Journal for the History of Astronomy, 43(4):491–496, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2015:GRA

[Dav15] A. E. L. Davis. The geometrical root of the area-measure of time (from Kepler’s Astronomia nova). Journal for the History of Astronomy, 46(3):297– 324, August 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2017:REM

[Dav17] John Davis. A royal English medieval astrolabe made for use in northern Italy. Journal for the History of Astronomy, 48(1):3–32, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616681214.

Davis:2019:CAB

[Dav19] John Davis. The “Chaucerian”

astrolabe in the British Museum: a reassessment of its dating and ownership. Journal for the History of Astronomy, 50(2):121–154, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619845585.

Beaver:1970:BWI

[dB70] Donald deB. Beaver. Bernard Walther: Innovator in astronomical observation. Journal for the History of Astronomy, 1 (1):39–43, February 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

diBono:1995:CAF

[dB95] Mario di Bono. Copernicus, Amico, Fracastoro and t.ūsī’s device: observations on the use and transmission of a model. Journal for the History of Astronomy, 26(2):133–154, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deClercq:2001:BRI

[dC01] Peter de Clercq. Book review:

Instrument makers of Augsburg: Augustanus Opticus. Johann Wiesel (1583–1662) und 200 Jahre optisches Handwerk in Augsburg. Journal for the History of Astronomy, 32 (2):178–179, May 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2006:CPR

[De 06] David H. De Vorkin. The changing place of red giant stars in the evolutionary process. Journal for the History of Astronomy, 37(4):429–469, November 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeBianchi:2013:BRK

[De 13] Silvia De Bianchi. Book review: Kant and astronomy, Immanuel Kant, Natural Science. Journal for the History of Astronomy, 44(4):487–488, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeGroot:2014:AEE

[De 14] Jean De Groot. Aristotle’s empiricism: experience and mechanics in the fourth century BC. Parmenides Publishing, Las Vegas, NY, USA, 2014. ISBN 1-930972-83-0 (paperback), 1-930972-84-9 (ebook). xxv + 442 pp. LCCN Q124.95 .D43 2014.

deRezendeVergara:2022:OAM

[de 22] Moema de Rezende Vergara.

Observational astronomy and the mapping of Brazil at the turn of the 20th century. Journal for the History of Astronomy, 53(2):197–208, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221095143.

Dambis:2000:DPS

[DE00] A. K. Dambis and Yu. N. Efremov. Dating Ptolemy’s star catalogue through proper motions: The Hipparchan epoch. Journal for the History of Astronomy, 31(2):115–134, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dear:1994:BRF

[Dea94] Peter Dear. Book review:

French expansion and the exact sciences: Civilizing Mission: Exact Sciences and French Overseas Expansion, 1830– 1940. Journal for the History of Astronomy, 25(4):331–332, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:1987:BRG

[Déb87] Suzanne Débarbat. Book review: The Greenwich symposium: The Longitude Zero Symposium 1984. Journal for the History of Astronomy, 18 (4):288–290, November 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:1991:BRM

[Déb91] Suzanne Débarbat. Book review: Measuring the Earth, From Plane to Spheroid. Journal for the History of Astronomy, 22(2):188–189, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:1995:CHG

[Déb95] Suzanne Débarbat. Commission 41 at the Hague General Assembly of the IAU. Journal for the History of Astronomy, 26(1):90–91, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:1996:UUA

[Déb96] Suzanne Débarbat. An unusual use of an astronomical instrument: The Dreyfus Affair and the Paris ‘Macro–Micromètre’. Journal for the History of Astronomy, 27(1):45–52, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:2005:BRF

[Déb05] Suzanne Débarbat. Book review: Foucault and his pendulum: The Life and Science of Léon Foucault: The Man Who Proved the Earth Rotates, Pendulum: Léon Foucault and the Triumph of Science. Journal for the History of Astronomy, 36(3):354–355, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Debarbat:2013:BRL

[Déb13] Suzanne Débarbat. Book review: Lacaille, Nicolas-Louis De La Caille, Astronomer and Geodesist. Journal for the History of Astronomy, 44(3):367– 368, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dekker:2008:CPO

[Dek08a] Elly Dekker. Carolingian planetary observations: The case of the Leiden planetary configuration. Journal for the History of Astronomy, 39(1): 77–90, February 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dekker:2008:WEA

[Dek08b] Elly Dekker. A ‘watermark’ of Eudoxan astronomy. Journal for the History of Astronomy, 39(2):213–228, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dekker:2012:BRK

[Dek12] Elly Dekker. Book review:

King on astrolabes, Astrolabes from Medieval Europe. Journal for the History of Astronomy, 43(2):249–250, May 2012.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dekker:2020:SCM

[Dek20]  Elly Dekker.  Stars and

constellations in medieval manuscripts. Journal for the History of Astronomy, 51(4): 492–493, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620954563.

DASBC:2023:CBC

[Del23] Delaware Astronomical Society Book Club. The coolest book cover ever. Journal for the History of Astronomy, 54(2):246–248, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231154233.

Densmore:2000:ERW

[Den00] Dana Densmore. Essay review: Writing, and reading, Principia. Journal for the History of Astronomy, 31(4):345–348, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1975:MPS

[DeV75] David H. DeVorkin. Michelson and the problem of stellar diameters. Journal for the History of Astronomy, 6(1):1–18, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1985:BRP

[DeV85a] David H. DeVorkin. Book review: Payne-Gaposchkin: Cecilia Payne-Gaposchkin: an Autobiography and other Recollections. Journal for the History of Astronomy, 16(1): 64–66, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1985:BRC

[DeV85b] David H. DeVorkin. Book review: The collected works of Saha: Collected Works of Meghnad Saha. Journal for the History of Astronomy, 16 (1):71–72, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1989:BRW

[DeV89] David DeVorkin. Book review:

Wilhelm and Otto Struve, Resolute and Undertaking Characters: The Lives of Wilhelm and Otto Struve. Journal for the History of Astronomy, 20 (2):145–146, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1990:DDC

[DeV90] David H. DeVorkin. Defending a dream: Charles Greeley Abbot’s years at the Smithsonian. Journal for the History of Astronomy, 21(1):121–136, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Devorkin:1994:QPS

[DeV94] David H. DeVorkin. Quantum physics and the stars (IV): Meghnad Saha’s fate. Journal for the History of Astronomy, 25(3):155–188, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1996:BRL

[DeV96] David H. DeVorkin. Book review: A Limited View of Hubble: Edwin Hubble: Mariner of the Nebulae. Journal for the History of Astronomy, 27 (2):184–185, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1998:ERA

[DeV98] David DeVorkin. Essay review: The American astronomical community: American Astronomy: Community, Careers, and Power, 1859– 1940. Journal for the History of Astronomy, 29(4):389–392, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2000:BRA

[DeV00a] David DeVorkin. Book review: Astronomical science down the ages: The Cambridge Concise History of Astronomy. Journal for the History of Astronomy, 31(2):161–163, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2000:QPS

[DeV00b] David H. DeVorkin. Quantum physics and the stars (V): Physicists at Mount Wilson prior to 1922. Journal for the History of Astronomy, 31(4): 301–321, November 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2002:BRM

[DeV02a] David DeVorkin. Book review:

More on Lowell, Percival Lowell: The Culture and Science of a Boston Brahmin. Journal for the History of Astronomy, 33 (3):307–308, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2002:MP

[DeV02b] David DeVorkin. Menzel at Princeton. Journal for the History of Astronomy, 33(2):119– 131, May 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2003:BRP

[DeV03] David H. DeVorkin. Book review: Payne-Gaposchkin remembered: The Starry Universe: The Cecilia Payne-Gaposchkin Centenary. Journal for the History of Astronomy, 34(2):244–245, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2005:BRA

[DeV05] David DeVorkin. Book review: Astrophysics in outline: A Concise History of Solar and Stellar Physics. Journal for the History of Astronomy, 36 (3):349–351, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2007:BRC

[DeV07] David DeVorkin. Book review: Changing the face of the Cosmos: Revolutionaries of the Cosmos: The Astro-Physicists. Journal for the History of Astronomy, 38(1): 134–136, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Devorkin:2014:BRE

[Dev14a] David Devorkin. Book review: an encyclopedia: Encyclopedia of the History of Astronomy and Astrophysics. Journal for the History of Astronomy, 45(1):130–131, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Devoy:2014:BRH

[Dev14b] Louise Devoy. Book review: How to navigate: The Lost Art of Finding Our Way. John Edward Huth (Belknap Press of Harvard University Press, Cambridge, MA, 2013). Pp.

528. $35. ISBN 978-0-674-07282-4. Journal for the History of Astronomy, 45(4):488– 490, November 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:2018:DL

[DeV18] David H. DeVorkin. From Dallas to LIGO. Journal for the History of Astronomy, 49(4):497–498, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618808873.

DeVorkin:2019:WSY

[DeV19] David H. DeVorkin. When to send your telescope aloft. Journal for the History of Astronomy, 50(3):265–305, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619864723.

DeVorkin:2020:EB

[DeV20] David H. DeVorkin. An encyclopaedic biography. Journal for the History of Astronomy, 51(3):374–376, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620930836.

Dewhirst:1970:BRA

[Dew70] David W. Dewhirst. Book review: Alvan Clark and sons: Alvan Clark and Sons, Artists in Optics. Journal for the History of Astronomy, 1(2):158– 159, August 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dewhirst:1979:ERT

[Dew79] D. W. Dewhirst. Essay review: Time and place, Francis Place and the Early History of the Greenwich Observatory. Journal for the History of Astronomy, 10(2):122–126, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dewhirst:1982:NPR

[Dew82] D. W. Dewhirst. A note on polar refractors. Journal for the History of Astronomy, 13 (2):119–120, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dewhirst:1987:BRJ

[Dew87] D. W. Dewhirst. Book review: J.-C. Houzeau and civil rights: My Passage at the New Orleans Tribune: a Memoir of the Civil War Era. Journal for the History of Astronomy, 18(1):66, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dewhirst:2000:BRP

[Dew00] David W. Dewhirst. Book review: Practical astronomers: Astronomical Instruments and Their Users, the Victorian Amateur Astronomer: Independent Astronomical Research in Britain 1820–1920. Journal for the History of Astronomy, 31(4):357–359, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2014:LHF

[DF14] Wolfgang Rudolf Dick and Dietmar Fürst. Lebensläufe und Himmelsbahnen: Festschrift zum 60. Geburtstag von Jürgen Hamel, volume 52 of Acta historica astronomiae. AVA, Akademische Verlagsanstalt, Leipzig, Germany, 2014. ISBN 3-944913-42-6 (paperback). ISSN 1422-8521. 293 pp. LCCN QB28 .L43 2014.

Drake:1973:AAA

[DH73] Lawrence A. Drake and Jani Haenke. Astronomy archives in Australia (2): 1. archives of Riverview Observatory. Journal for the History of Astronomy, 4(3):211–213, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dewhirst:1991:RS

[DH91] David W. Dewhirst and Michael Hoskin. The Rosse spirals. Journal for the History of Astronomy, 22(4):257–266, November 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deHilster:2021:TEM

[dH21] Nicolàs de Hilster. Training early modern navigators. Journal for the History of Astronomy, 52(2):244–245, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211004674.

Dibble:1976:PPT

[Dib76] William E. Dibble. A possible Pythagorean triangle at Stonehenge. Journal for the History of Astronomy, 7(2):141– 142, June 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1982:BRG

[Dic82] Steven J. Dick. Book review: Greenwich time: Greenwich Time and the Discovery of the Longitude. Journal for the History of Astronomy, 13 (1):69–70, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1990:BRP

[Dic90] Steven J. Dick. Book review: Psychology and history, planets and perception: Telescopic Views and Interpretations, 1609–1909. Journal for the History of Astronomy, 21 (2):226–228, May 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1991:JQA

[Dic91a] Steven J. Dick. John Quincy Adams, the Smithsonian bequest and the founding of the U.S. Naval Observatory. Journal for the History of Astronomy, 22(1):31–44, February 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1991:NOO

[Dic91b] Steven J. Dick. National observatories: an overview. Journal for the History of Astronomy, 22(1):1–4, February 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1995:BRL

[Dic95] Steven J. Dick. Book review:

Lowell’s observatory commemorated: The Explorers of Mars Hill: a Centennial History of Lowell Observatory, 1894– 1994. Journal for the History of Astronomy, 26(2):181–182, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1997:BRA

[Dic97a] Steven J. Dick. Book review:

Astronomy and astronomers at Yale: Astronomy at Yale 1701– 1968. Journal for the History of Astronomy, 28(2):174–175, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1997:CKG

[Dic97b] Steven J. Dick. Commission 41 at the Kyoto General Assembly of the IAU. Journal for the History of Astronomy, 28(4):365–367, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1999:BRP

[Dic99] Steven J. Dick. Book review: The perihelion of Mercury: In Search of Planet Vulcan: The Ghost in Newton’s Clockwork Universe. Journal for the History of Astronomy, 30(1): 85–87, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2000:BRA

[Dic00a] Steven J. Dick. Book review: Adventures of a spacefaring people: Space and the American Imagination. Journal for the History of Astronomy, 31(2):176–177, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2000:BRT

[Dic00b] Steven J. Dick. Book review: Two centuries of astrometry: The Message of the Angles: Astrometry from 1798 to 1998. Journal for the History of Astronomy, 31(4):354–355, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2000:CMG

[Dic00c] Steven J. Dick. Commission 41 at the Manchester General Assembly of the IAU. Journal for the History of Astronomy, 31(4):365–367, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2006:BRT

[Dic06] Steven J. Dick. Book review: Three generations at Mount Wilson: The Mount Wilson Observatory: Breaking the Code of Cosmic Evolution (Centennial History of the Carnegie Institution of Washington, I). Journal for the History of Astronomy, 37(2):243– 245, May 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2009:BRT

[Dic09] Steven J. Dick. Book review: Transits of Venus in Victorian times, The Transit of Venus Enterprise in Victorian Britain. Journal for the History of Astronomy, 40(2):232– 233, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dicati:2013:STA

[Dic13a]  Renato Dicati.  Stamping

through astronomy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 88-470-2828-0, 88-470-2829-9 (e-book). xi + 373 pp. LCCN QB28 .D53

2013. URL http://www.loc.
gov/catdir/enhancements/
fy1410/2013940073-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1410/2013940073-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1410/2013940073-t.html.

Dick:2013:BRC

[Dic13b] Steven J. Dick. Book review: Cosmic evolution, The Stardust Revolution: The New Story of Our Origin in the Stars. Journal for the History of Astronomy, 44(2):219–220, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:2021:SHC

[Dic21] Wolfgang R. Dick. Six hundred calendar makers. Journal for the History of Astronomy, 52(2):240–242, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828621996481.

Dijksterhuis:2009:BRK

[Dij09] Fokko Jan Dijksterhuis. Book review: Kepler on optics, Johannes Kepler, Schriften zur Optik, 1604–1611. Journal for the History of Astronomy, 40 (3):370–371, August 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dingle:1973:ERS

[Din73] Herbert Dingle. Essay review:

Sir Norman Lockyer: Science and Controversy: a Biography of Sir Norman Lockyer. Journal for the History of Astronomy, 4(2):131–133, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deJong:2012:BOV

[dJ12] Teije de Jong. Babylonian observations of Venus: Arcus visionis, atmospheric extinction and observational practice. Journal for the History of Astronomy, 43(4):391–409, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1983:QPSa

[DK83a] David H. DeVorkin and Ralph Kenat. Quantum physics and the stars (I): The establishment of a stellar temperature scale. Journal for the History of Astronomy, 14(2):102–132, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

DeVorkin:1983:QPSb

[DK83b] David H. DeVorkin and Ralph Kenat. Quantum physics and the stars (II): Henry Norris Russell and the abundances of the elements in the atmospheres of the Sun and stars. Journal for the History of Astronomy, 14(3):180–222, October 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1996:PMA

[DK96] Jerzy Dobrzycki and Richard L. Kremer. Peurbach and Marāgha astronomy? The Ephemerides of Johannes Angelus and their implications. Journal for the History of Astronomy, 27(3):187–237, August 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Davis:2015:EEA

[DL15] John Davis and Michael Lowne. An early English astrolabe at Gonville & Caius College, Cambridge, and Walter of Elveden’s Kalendarium. Journal for the History of Astronomy, 46(3):257–290, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Deiss:1998:POS

[DN98] Bruno M. Deiss and Volker Nebel. On a pretended observation of Saturn by Galileo. Journal for the History of Astronomy, 29(3):215–220, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1973:ACC

[Dob73] Jerzy Dobrzycki. The Aberdeen copy of Copernicus’s Commentariolus (by Duncan Liddel). Journal for the History of Astronomy, 4(2):124– 127, June 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1977:BRB

[Dob77] Jerzy Dobrzycki. Book review: Baers’s astronomical tables, Les tables astronomiques de Louvain de 1528 par Henri Baers ou Vekenstyl. Journal for the History of Astronomy, 8(3):216, October 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1979:BRW

[Dob79] Jerzy Dobrzycki. Book review: The Washington Copernicus symposium, The Nature of Scientific Discovery. Journal for the History of Astronomy, 10(2):129, June 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1983:BRY

[Dob83a] Jerzy Dobrzycki. Book review:

a Young Person’s guide to history of astronomy: Vesmir (The Universe). Journal for the History of Astronomy, 14(2): 148, June 1983. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1983:BRS

[Dob83b] Jerzy Dobrzycki. Book review: Source materials for medieval astronomy: Les Sources Astronomiques: Textes, Tables, Instruments. Journal for the History of Astronomy, 14 (2):147–148, June 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1986:BRB

[Dob86] Jerzy Dobrzycki. Book review: Byzantine astronomy: Nicéphore Grégoras: Calcul de l’eclipse de Soleil du 16 Juillet 1330. Journal for the History of Astronomy, 17(3):199–200, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1992:BRZ

[Dob92] Jerzy Dobrzycki. Book review: Zinner on Regiomontanus: Regiomontanus: His Life and Work. Journal for the History of Astronomy, 23(4):305–307, November 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1996:BRC

[Dob96] Jerzy Dobrzycki. Book review: Copernicus in context: Nicolaus Copernicus: Leben, Werk und Wirkung. Journal for the History of Astronomy, 27(1): 89–91, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1999:BRH

[Dob99a] Jerzy Dobrzycki. Book review: a history of astronomy in German: Geschichte der Astronomie: Von den Anfängen bis zur Gegenwart. Journal for the History of Astronomy, 30 (3):331–332, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1999:SAA

[Dob99b] Jerzy Dobrzycki. Saturn, Aristotelian astronomy, and Cracow astronomers: an episode from the early years of telescopic astronomy. Journal for the History of Astronomy, 30 (2):121–129, May 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:2000:BRC

[Dob00] Jerzy Dobrzycki. Book review: Constellations old and new: The Cambridge Guide to the Constellations. Journal for the History of Astronomy, 31 (3):270–272, August 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:2001:NCE

[Dob01] Jerzy Dobrzycki. Notes on Copernicus’s early heliocentrism. Journal for the History of Astronomy, 32(3):223– 225, August 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). Astronomy and its histories (Vancouver, BC, 2000).

Dobler:2002:DPS

[Dob02] Hermann R. Dobler. The dating of Ptolemy’s star catalogue. Journal for the History of Astronomy, 33(3):265–277, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:1990:RMS

[Doe90] Ron Doel. Redefining a mission: The Smithsonian Astrophysical Observatory on the move. Journal for the History of Astronomy, 21(1):137–153, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:1994:BRE

[Doe94] Ronald E. Doel. Book review: Exploring the Moon: To a Rocky Moon: a Geologist’s History of Lunar Exploration. Journal for the History of Astronomy, 25(2):149– 151, May 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:2000:BRS

[Doe00] Ronald E. Doel. Book review: Studying the planets by radar: To See the Unseen: a History of Planetary Radar Astronomy. Journal for the History of Astronomy, 31(1):79–81, February 2000. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:2001:BRD

[Doe01] Ronald E. Doel. Book review:

DeVorkin on Russell, Henry Norris Russell: Dean of American Astronomers. Journal for the History of Astronomy, 32 (4):366–369, November 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:2005:BRP

[Doe05] Ronald E. Doel. Book review:

Planetary astronomy in outline: Babylon to Voyager and Beyond: a History of Planetary Astronomy. Journal for the History of Astronomy, 36 (3):347–349, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:2008:BRJ

[Doe08] Ronald E. Doel. Book review: JPL’s recent past: Into the Black: JPL and the American Space Program, 1976– 2004. Journal for the History of Astronomy, 39(1):133–135, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Doel:2019:BRB

[Doe19] Ronald E. Doel. Book review: Building the Harvard-Smithsonian Center for Astrophysics: Fred Whipple’s Empire: The Smithsonian Astrophysical Observatory, 1955– 1973, David H. DeVorkin, (Smithsonian Institution Scholarly Press, Washington, DC, 2018). Pp. xvii + 401. Open access. ISBN 978-1-944466-18-3 (pbk). Journal for the History of Astronomy, 50(3): 375–376, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619862970.

Doggett:1987:BRA

[Dog87] LeRoy E. Doggett. Book review: The acceleration of the Moon: The Moon’s Acceleration and its Physical Origins, II: As Deduced from General Lunar Observations. Journal for the History of Astronomy, 18(1):68–70, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dick:1998:SNW

[DOL98] Steven J. Dick, Wayne Orchiston, and Tom Love. Simon Newcomb, William harkness and the Nineteenth-Century American transit of Venus expeditions. Journal for the History of Astronomy, 29(3):221– 255, August 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1971:BRP

[Don71a] W. H. Donahue. Book review: In praise of Bruno: The Infinite Worlds of Giordano Bruno. Journal for the History of Astronomy, 2(3):204–205, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1971:BRJ

[Don71b] W. H. Donahue. Book review:

John Wilkins: The Mathematical and Philosophical Works of the Right Rev. John Wilkins. Journal for the History of Astronomy, 2(1):48–49, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1973:BRC

[Don73a] W. H. Donahue. Book review:

The comet amidst the spheres: The Comet of 1577. Journal for the History of Astronomy, 4 (1):61–62, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1973:BRU

[Don73b] W. H. Donahue. Book review:

Understanding the Milky Way: The Milky Way: an Elusive Road for Science. Journal for the History of Astronomy, 4(3): 200–201, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1973:HUP

[Don73c] W. H. Donahue. A hitherto unreported pre-Keplerian oval orbit. Journal for the History of Astronomy, 4(3):192–194, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1979:ERW

[Don79] W. H. Donahue. Essay review: Works ascribed to Galileo, Galileo against the Philosophers in his “Dialogue of Cecco di Ronchitti” (1605) and “Considerations of Alimberto Mauri” (1606). Journal for the History of Astronomy, 10 (1):44–47, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1982:BRN

[Don82a] William Donahue. Book review: a Novelist’s Kepler: Kepler. Journal for the History of Astronomy, 13(2):135–136, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1982:BRG

[Don82b] William Donahue. Book review: The Galileo of the “dialogue”: Galileo and the Art of Reasoning. Journal for the History of Astronomy, 13(1): 68–69, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1986:BRB

[Don86] W. H. Donahue. Book review: Biography of Harriot: Thomas Harriot: a Biography.

Journal for the History of Astronomy, 17(1):71–73, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1988:BRK

[Don88a] W. H. Donahue. Book review: Kepler, Tycho and Ursus: Three Imperial Mathematicians: Kepler Trapped between Tycho Brahe and Ursus. Journal for the History of Astronomy, 19(1):64–66, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1988:KFF

[Don88b] W. H. Donahue. Kepler’s fabricated figures: Covering up the mess in the New Astronomy. Journal for the History of Astronomy, 19(4):217–237, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1992:ERE

[Don92] W. H. Donahue. Essay review: On editing Keplerian manuscripts: Johannes Kepler Gesammelte Werke, XX/ 1: Manuscripta astronomica (I). Journal for the History of Astronomy, 23(2):141–145, May 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1993:KFT

[Don93] W. H. Donahue. Kepler’s first thoughts on oval orbits: text, translation, and commentary. Journal for the History of Astronomy, 24(1–2):71– 100, May/February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/71.full.pdf+html.

Donahue:1996:KAO

[Don96] William H. Donahue. Kepler’s approach to the oval of 1602, from the Mars notebook. Journal for the History of Astronomy, 27(4):281–295, November 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:1998:BRH

[Don98] William Donahue. Book review: The Hall Festschrift: Renaissance and Revolution: Humanists, Scholars, Craftsmen and Natural Philosophers in Early Modern Europe. Journal for the History of Astronomy, 29(1):89–91, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:2000:BRK

[Don00] William H. Donahue. Book review: Kepler’s notebook on Mars: Johannes Kepler Gesammelte Werke 20,2 (Manuscripta Astronomica (II): Commentaria in Theoriam Martis). Journal for the History of Astronomy, 31(1):71– 72, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:2002:BRP

[Don02] William H. Donahue. Book review: Philosophy in Kepler’s astronomy, Kepler’s Philosophy and the New Astronomy. Journal for the History of Astronomy, 33(3):296–297, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Donahue:2009:BRM

[Don09] William Donahue. Book review: Mechanics in early modern cosmology: Mechanics and Cosmology in the Medieval and Early Modern Period. Journal for the History of Astronomy, 40(4):473–474, November 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dooley:2018:ABS

[Doo18] Brendan Dooley. Astrology as a bookish science. Journal for the History of Astronomy, 49(4):500–501, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618807009.

Doran:2020:SDE

[Dor20] Connemara Doran. The search for dark energy. Journal for the History of Astronomy, 51(4):483–485, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620965806.

Price:1982:ERI

[dP82] Derek deS. Price. Essay review: Instruments in the British museum: A Catalogue of European Scientific Instruments in the Department of Medieval and Later Antiquities of the British Museum. Journal for the History of Astronomy, 13(2):129– 132, June 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1973:GPC

[Dra73] Stillman Drake. Galileo’s “Platonic” cosmogony and Kepler’s Prodromus. Journal for the History of Astronomy, 4(3): 174–191, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1976:GFT

[Dra76] Stillman Drake. Galileo’s first telescopic observations. Journal for the History of Astronomy, 7(3):153–168, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1979:GSP

[Dra79] Stillman Drake. Galileo and satellite prediction. Journal for the History of Astronomy, 10(2):75–95, June 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1980:BRK

[Dra80] Stillman Drake. Book review:

a Koestler-type popularization, This Wild Abyss: The Story of the Men Who Made Modern Astronomy. Journal for the History of Astronomy, 11(2): 141, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1984:BRG

[Dra84a] Stillman Drake. Book review:

Galileian biographies: Galileo Galilei, Galileo Galilei. Journal for the History of Astronomy, 15(3):215–216, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1984:GKP

[Dra84b] Stillman Drake. Galileo, Kepler, and phases of Venus. Journal for the History of Astronomy, 15(3):198–208, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1986:NGL

[Dra86] Stillman Drake. A neglected Galilean letter. Journal for the History of Astronomy, 17 (2):99–108, May 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drake:1990:BRG

[Dra90] Stillman Drake. Book review: Galileian documents, The Galileo Affair: a Documentary History. Journal for the History of Astronomy, 21 (4):374–376, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drennan:2015:VWF

[Dre15] Anthony S. Drennan. Variations within the Fortin– Flamsteed atlas of 1776: Taurus Poniatowski. Journal for the History of Astronomy, 46 (1):29–47, February 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Drubba:1984:BRS

[Dru84] Helmut Drubba. Book review: Spherical harmonies: Die Harmonie der Sphären: Die Geschichte der Idee des Welteneinklangs und der Seeleneinstimmung. Journal for the History of Astronomy, 15(3):215, October 1984. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dobrzycki:1989:TCC

[DS89] Jerzy Dobrzycki and Lech Szczucki. On the transmission of Copernicus’s Commentariolus in the Sixteenth Century. Journal for the History of Astronomy, 20(1):25–28, February 1989. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Price:1978:BRR

[dSP78] Derek de Solla Price. Book review: Richard of Wallingford, Richard of Wallingford: an Edition of his Writings with Introductions, English Translation and Commentary. Journal for the History of Astronomy, 9(3):219–220, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duffy:2017:AGR

[Duf17] Cian Duffy. Astronomy and the “greater romantic lyric”. Journal for the History of Astronomy, 48(1):125–126, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616687599.

Duke:2002:DAS

[Duk02] Dennis W. Duke. Dating the Almagest star catalogue using proper motions: a reconsideration. Journal for the History of Astronomy, 33(1): 45–55, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2003:DAB

[Duk03] Dennis W. Duke. The depth of association between the ancient star catalogues. Journal for the History of Astronomy, 34(2):227–230, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2004:CAA

[Duk04] Dennis Duke. Computer animations of Ancient Greek and Arabic planetary models. Journal for the History of Astronomy, 35(2):225–228, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2005:COE

[Duk05] Dennis W. Duke. Comment on the origin of the Equant papers by Evans, Swerdlow, and Jones. Journal for the History of Astronomy, 36(1):1–6, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2006:AFG

[Duk06] Dennis W. Duke. Analysis of the Farnese globe. Journal for the History of Astronomy, 37(1):87–100, February 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2008:FLE

[Duk08]  Dennis Duke.  Four lost

episodes in ancient solar theory. Journal for the History of Astronomy, 39(3):283–296, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2009:YLH

[Duk09] Dennis Duke. A year length hidden in ancient planetary mean motions. Journal for the History of Astronomy, 40 (2):213–216, May 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2010:BRE

[Duk10] Dennis Duke. Book review:

Early trigonometry: The Mathematics of the Heavens and the Earth: The Early History of Trigonometry. Journal for the History of Astronomy, 41(1): 123–124, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2011:BRP

[Duk11] Dennis Duke. Book review:

Pedersen’s survey of the Almagest revised: A Survey of the Almagest with Annotation and New Commentary by Alexander Jones. Journal for the History of Astronomy, 42(3):416– 418, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2012:BRE

[Duk12] Dennis Duke. Book review:

Eclipse reckoning in antiquity: Chasing Shadows: Mathematics, Astronomy, and the Early History of Eclipse Reckoning. Journal for the History of Astronomy, 43(4):500–502, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Duke:2016:FFR

[Duk16] Dennis W. Duke. Festschrift for F. Richard Stephenson. Journal for the History of Astronomy, 47(2):225–226, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dunn:1992:BRA

[Dun92] Richard Dunn. Book review: Almanacs in America and England: Practical Astronomy during the Seventeenth Century: Almanac-Makers in America and England. Journal for the History of Astronomy, 23(3): 226–227, August 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dunn:2013:BRT

[Dun13] Richard Dunn. Book review: Transits of Venus: Chasing Venus: The Race to Measure the Heavens. Journal for the History of Astronomy, 44(1): 114–115, February 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dunn:2020:BRM

[Dun20] Richard Dunn. Book review: More than a catalogue of sundials. Time of Our Lives: Sundials of the Adler Planetarium. Sara J. Schechner (Adler Planetarium, Chicago, 2019). Pp. xiii + 474. $40. ISBN 978-0-578-49710-5. Journal for the History of Astronomy, 51(1):123–124, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619900815.

Dunn:2023:FCR

[Dun23] Richard Dunn. A festschrift for Clive Ruggles. Journal for the History of Astronomy, 54(1):108–109, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221137803.

Dunn:2024:CH

[Dun24] Richard Dunn. Cosmography and its histories. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231214487.

Dupree:1990:SAO

[Dup90] A. Hunter Dupree. Smithsonian astrophysical observatory: From Washington to Cambridge. Journal for the History of Astronomy, 21(1): 107–110, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dupre:2003:GTC

[Dup03] Sven Dupré. Galileo’s telescope and celestial light. Journal for the History of Astronomy, 34(4):369–399, November 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dupre:2008:BRT

[Dup08] Sven Dupré. Book review: The telescope in history; Der Meister und die Fernrohre: Das Wechselspiel zwischen Astronomie und Optik in der Geschichte. Journal for the History of Astronomy, 39(3): 410–411, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dupre:2014:BRA

[Dup14] Sven Dupré. Book review: Astrology in the Renaissance: Influences: Art, Optics, and Astrology in the Italian Renaissance. Journal for the History of Astronomy, 45(1):132–134, February 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

deVaucouleurs:1991:BRR

[dV91] G. de Vaucouleurs. Book review: The Renaissance of French astronomy, Des hommes, des télescopes, des étoiles. Journal for the History of Astronomy, 22(3):248– 250, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Dearborn:1983:TMP

[DW83] D. S. P. Dearborn and R. E.

White. The “Torreon” of Machu Picchu as an observatory. Journal for the History of Astronomy, 14(5):S37–S49, February 1983. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eagleton:2008:BRGa

[Eag08a] Catherine Eagleton. Book review: German sundials; Inventar der historischen Sonnenuhren in Mecklenburg– Vorpommern. Journal for the History of Astronomy, 39(2): 278, May 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eagleton:2008:BRGb

[Eag08b] Catherine Eagleton. Book review: German sundials; Inventar der historischen Sonnenuhren in Mecklenburg– Vorpommern. Journal for the History of Astronomy, 39(3): 411–412, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:1989:BRA

[Eas89] Bruce Eastwood. Book review:

Adelard Analysed, Adelard of Bath: an English Scientist and Arabist of the Early Twelfth Century. Journal for the History of Astronomy, 20(2):138– 139, June 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:1992:HHT

[Eas92] Bruce Stansfield Eastwood. Heraclides and heliocentrism: Texts, diagrams, and interpretations. Journal for the History of Astronomy, 23(4):233–260, November 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:1996:BRC

[Eas96] Bruce Eastwood. Book review: The Cosmos in the Middle Ages and after: Planets, Stars, and Orbs: The Medieval Cosmos, 1200–1687. Journal for the History of Astronomy, 27(1): 83–85, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:1997:ACL

[Eas97] Bruce Stansfield Eastwood. Astronomy in Christian Latin Europe c. 500–c. 1150. Journal for the History of Astronomy, 28 (3):235–258, August 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:2000:AIP

[Eas00] Bruce S. Eastwood. Astronomical images and planetary theory in Carolingian studies of Martianus Capella. Journal for the History of Astronomy, 31 (1):1–28, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:2001:JSE

[Eas01] Bruce Stansfield Eastwood. Johannes Scottus Eriugena, Suncentred planets, and Carolingian astronomy. Journal for the History of Astronomy, 32(4): 281–324, November 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eastwood:2010:BRM

[Eas10] Bruce S. Eastwood. Book review: Medieval astronomical diagrams: Visuelle Weltaneignung: Astronomische und Kosmologische Diagramme in Handschriften des Mittelalters. Journal for the History of Astronomy, 41(4):510–512, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edwards:2004:MAE

[EB04] Edmundo R. Edwards and Juan Antonio Belmonte. Megalithic astronomy of Easter Island: a reassessment. Journal for the History of Astronomy, 35(4):421–433, November 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Esteban:1994:SMT

[EBA94] César Esteban, Juan A. Belmonte, and Antonio Aparicio. A solstitial marker in Tenerife: Addendum. Journal for the History of Astronomy, 25(19): S84–S86, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [BAE93].

Esteban:2001:OPI

[EBB^{+}01] César Esteban, Juan A. Belmonte, María A. Perera Betancort, Rita Marrero, and José J. Jiménez González. Orientations of pre-Islamic temples of Northwest Africa. Journal for the History of Astronomy, 32 (26):S65–S84, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2010:SAP

[ECT10] James Evans, Christián C. Carman, and Alan S. Thorndike. Solar anomaly and planetary displays in the Antikythera mechanism. Journal for the History of Astronomy, 41(1): 1–39, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1971:BRG

[Edd71a] John A. Eddy. Book review: The Golden age of solar physics: Early Solar Physics. Journal for the History of Astronomy, 2(2):126–128, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1971:SFE

[Edd71b] John A. Eddy. The Schaeberle 40-ft eclipse camera of Lick Observatory. Journal for the History of Astronomy, 2(1):1–22, February 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1972:TEI

[Edd72] John A. Eddy. Thomas A. Edison and infra-red astronomy. Journal for the History of Astronomy, 3(3):165–187, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1974:BRP

[Edd74] John A. Eddy. Book review:

Popular astro-archaeology: Beyond Stonehenge. Journal for the History of Astronomy, 5 (1):66, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1980:BRH

[Edd80] John A. Eddy. Book review:

History in the service of astronomy, Applications of Early Astronomical Records. Journal for the History of Astronomy, 11(3):202–204, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1982:BRA

[Edd82] John A. Eddy. Book review:

Ancient records under attack: The Moon’s Acceleration and its Physical Origins. Journal for the History of Astronomy, 13(2):133–135, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1986:BRP

[Edd86] John A. Eddy. Book review: Pioneer of dendrochronology: Tree Rings and Telescopes: The Scientific Career of A. E. Douglass. Journal for the History of Astronomy, 17(1): 69–71, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eddy:1990:FAO

[Edd90] John A. Eddy. Founding the astrophysical observatory: The Langley years. Journal for the History of Astronomy, 21(1): 111–120, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmondson:1984:BRO

[Edm84] Frank K. Edmondson. Book review: Observatory at Kitt Peak: Realm of the Long Eyes: a Brief History of the Kitt Peak National Observatory. Journal for the History of Astronomy, 15(2):139–141, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmondson:1991:AKE

[Edm91] Frank K. Edmondson. Aura and KPNO: The evolution of an idea, 1952–58. Journal for the History of Astronomy, 22 (1):68–86, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmondson:1998:FFE

[Edm98] Frank K. Edmondson. The Ford Foundation and the European Southern Observatory. Journal for the History of Astronomy, 29(4):309–326, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmunds:2011:IAA

[Edm11] M. G. Edmunds. An initial assessment of the accuracy of the gear trains in the Antikythera mechanism. Journal for the History of Astronomy, 42(3): 307–320, August 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eelsalu:1974:BRS

[Eel74] Heino Eelsalu. Book review:

Schmidt the optician: Bernhard Schmidt. Journal for the History of Astronomy, 5(1):65, February 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eelsalu:1983:BRS

[Eel83]  H. Eelsalu.  Book review:

Schmidt and the Potsdam Observatory: Die Jugendjahre von Bernhard Schmidt und sein Briefwechsel mit dem Potsdamer Observatorium. Journal for the History of Astronomy, 14(1):67, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisberg:1985:BRQ

[Eis85] Joann Eisberg. Book review: “The Queen of Nineteenth-Century Science”: Mary Somerville and the Cultivation of Science, 1815–1840. Journal for the History of Astronomy, 16(2):144–146, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisberg:1988:BRA

[Eis88a]  Joann Eisberg.  Book re-

view: Aspects of astrophysics: The Analysis of Starlight: One Hundred and Fifty Years of Astronomical Spectroscopy. Journal for the History of Astronomy, 19(1):71–72, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisberg:1988:BRW

[Eis88b] Joann Eisberg. Book review: Warfare at the Dudley Observatory: Elites in Conflict: The Antebellum Clash over the Dudley Observatory. Journal for the History of Astronomy, 19(4):288–290, November 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisberg:1996:BRP

[Eis96]  Joann Eisberg.  Book re-

view: Perspectives on cosmology: Cosmology: Historical, Literary, Philosophical, Religious, and Scientific Perspectives, Encyclopedia of Cosmology: Historical, Philosophical, and Scientific Foundations of Modern Cosmology. Journal for the History of Astronomy, 27(1):86–87, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisberg:2001:MSO

[Eis01] Joann Eisberg. Making a science of observational cosmology: the cautious optimism of Beatrice Tinsley [1941–1981]. Journal for the History of Astronomy, 32(3):263–278, August 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). Astronomy and its histories (Vancouver, BC, 2000).

Eisberg:2003:BRA

[Eis03] Joann Eisberg. Book review:

Agnes Clerke in context: Agnes Mary Clerke and the Rise of Astrophysics. Journal for the History of Astronomy, 34(2): 243–244, May 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Eisenstaedt:2006:CHR

[Eis06] Jean Eisenstaedt. The Curious History of Relativity: How Einstein’s Theory of Gravity Was Lost and Found Again, volume 15 of CNRS editions. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11865-5. ix + 363 pp. LCCN QC173.6 EIS; 06.E06047.

Evans:2022:MH

[EK22] James Evans and Richard L. Kremer. Michael Hoskin (1930–2021). Journal for the History of Astronomy, 53(1): 120–123, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221077318.

Ellegaard:1985:ERM

[Ell85] Alvar Ellegård. Essay review: Megalithic astronomy: a new archaeological and statistical study of 300 western Scottish sites. Journal for the History of Astronomy, 16(9):S154– S157, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Elliott:1990:HHA

[Ell90] Clark A. Elliott. The history of Harvard astronomy: A View from the Archives. Journal for the History of Astronomy, 21 (1):3–8, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ellis:2008:BRC

[Ell08] George Ellis. Book review:

Cosmology down the ages, Conceptions of Cosmos, from Myths to the Accelerating Universe: a History of Cosmology. Journal for the History of Astronomy, 39(4):537–538, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2008:MIS

[EM08] James Evans and Marcel Marée. A miniature ivory sundial with equinox indicator from Ptolemaic Tanis, Egypt. Journal for the History of Astronomy, 39(1):1–17, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

English:2002:BRE

[Eng02] Tom English. Book review:

Eclipses illustrated, Glorious Eclipses: Their Past, Present, and Future. Journal for the History of Astronomy, 33(3): 302–303, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Escolano-Poveda:2022:AAF

[EP22] Marina Escolano-Poveda. Astrologica athribitana: Four demotic-hieratic horoscopes from Athribis (O. Athribis 17-36-5/1741 and ANAsh.Mus.D.O.633 reedited). Journal for the History of Astronomy, 53 (1):49–87, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211068573.

Esteban:1996:PHE

[ESBG96] César Esteban, Rosa Schlueter, Juan A. Belmonte, and Oswaldo González. Pre-Hispanic equinoctial markers in Gran Canaria, Part I. Journal for the History of Astronomy, 27(21): S73–S79, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Esteban:1997:PHE

[ESBG97] César Esteban, Rosa Schlueter, Juan A. Belmonte, and Oswaldo González. Pre-Hispanic equinoctial markers in Gran Canaria, Part II. Journal for the History of Astronomy, 28 (22):S51–S56, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1982:BRS

[Eva82] G. R. Evans. Book review: Space and vacuum: Much Ado about Nothing: Theories of Space and Vacuum from the Middle Ages to the Scientific Revolution. Journal for the History of Astronomy, 13(3): 217–218, October 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1986:BRM

[Eva86] David S. Evans. Book review:

Maclear and Herschel at the Cape: Maclear and Herschel: Letters and Diaries at the Cape of Good Hope 1834–1838. Journal for the History of Astronomy, 17(2):144–146, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1987:BRC

[Eva87a] David S. Evans. Book review: C. P. Smyth at the Cape: Charles Piazzi Smyth, Astronomer Artist: His Cape Years 1835–1845. Journal for the History of Astronomy, 18 (2):138–140, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1987:OPSa

[Eva87b] James Evans. On the origin of the Ptolemaic star catalogue: Part 1. Journal for the History of Astronomy, 18(3):155– 172, August 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1987:OPSb

[Eva87c] James Evans. On the origin of the Ptolemaic star catalogue: Part 2. Journal for the History of Astronomy, 18(4):233–278, November 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1988:BRT

[Eva88] James Evans. Book review: Theon of Alexandria’s big commentary: Le Grand Commentaire de Théon d’Alexandrie aux Tables Faciles de Ptolémée: Livre I. Histoire du Texte, Edition Critique, Traduction. Journal for the History of Astronomy, 19(1):56–58, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1989:BRB

[Eva89]  James Evans.  Book re-

view: a Byzantine treatise, An Eleventh-Century Manual of Arabo-Byzantine Astronomy. Journal for the History of Astronomy, 20(3):217–219, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1992:ERP

[Eva92]  James Evans.  Essay re-

view: The Ptolemaic star catalogue: Claudius Ptolemäus: Der Sternkatalog des Almagest. Die arabisch–Mittelalterliche Tradition, II: Die lateinische Übersetzung Gerhards von Cremona, the History of Ptolemy’s Star Catalogue. Journal for the History of Astronomy, 23 (1):64–68, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1993:BRD

[Eva93a] James Evans. Book review:

Dynamics from Aristotle to Newton: Absolute or Relative Motion? A Study from a Machian Point of View of the Discovery and the Structure of Dynamical Theories, I: The Discovery of Dynamics. Journal for the History of Astronomy, 24(3):229–230, August 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:1993:ERP

[Eva93b] James Evans. Essay review:

Ptolemy indicted again: The Origins of Ptolemy’s Astronomical Tables. Journal for the History of Astronomy, 24(1–2): 145–147, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/145.full.pdf+html.

Evans:1999:MCG

[Eva99] James Evans. The material culture of Greek astronomy. Journal for the History of Astronomy, 30(3):237–307, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2000:BRC

[Eva00]  James Evans.  Book re-

view: Copernicus on CD-ROM: De Revolutionibus orbium coelestium libri VI. Journal for the History of Astronomy, 31(1):75–76, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2001:ERG

[Eva01] James Evans. Essay review: Greek astronomy from Alexandria, Byzantium and Louvain-La-Neuve: Le “Grande Commentaire” de Théon d’Alexandrie aux Tables Faciles de Ptolémée: Livres II et III: Le “Grand Commentaire” de Théon d’Alexandrie aux Tables Faciles de Ptolémée: Livre IV: Études d’Astronomie Byzantine. Journal for the History of Astronomy, 32(1): 77–80, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2002:BRN

[Eva02] James Evans. Book review: New light on Babylon, Ancient Astronomy and Celestial Divination. Journal for the History of Astronomy, 33(1): 75–76, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2004:AAP

[Eva04] James Evans. The Astrologer’s apparatus: a picture of professional practice in Greco– Roman Egypt. Journal for the History of Astronomy, 35 (1):1–44, February 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2005:BRG

[Eva05] James Evans. Book review:

a Greek textbook of astronomy: Cleomedes’ Lectures on Astronomy: a Translation of the Heavens. Journal for the History of Astronomy, 36(3): 340–342, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2006:BRR

[Eva06]  James Evans.  Book re-

view: Representing the heavens: Himmelsgloben und Sternkarten: Astronomie und Astrologie in Vorzeit und Altertum. Journal for the History of Astronomy, 37(2):239–240, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2007:BRS

[Eva07] James Evans. Book review:

Simplicius on Aristotle’s astronomy, Simplicius: On Aristotle’s “On the Heavens” 2.1-9 and 2.10-14. Journal for the History of Astronomy, 38(4): 503–505, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2008:BRA

[Eva08]  James Evans.  Book re-

view: Ancient Greek sundials; Die antiken Sonnenuhren Griechenlands: Festland und Peloponnes. Journal for the History of Astronomy, 39(1): 125, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2009:BRE

[Eva09a] James Evans. Book review: Explaining the astrolabe, The Astrolabe. Journal for the History of Astronomy, 40(2):221– 222, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2009:BRS

[Eva09b] James Evans. Book review: The saga of the Antikythera mechanism, Decoding the Heavens: a 2000-Year-Old Computer and the Century-Long Search to Discover its Secrets. Journal for the History of Astronomy, 40(3):362–364, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2011:BRT

[Eva11] James Evans. Book review: Time in the classical world: Time in Antiquity. Journal for the History of Astronomy, 42(4):530–531, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2012:BRP

[Eva12] James Evans. Book review: The phaenomena of Aratus, Aratus. Phaenomena. Journal for the History of Astronomy, 43(3):366, August 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2013:BRT

[Eva13] James Evans. Book review:

Tommy Gold, Taking the Back off the Watch: a Personal Memoir. Journal for the History of Astronomy, 44(3):370– 372, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2015:NE

[Eva15] James Evans. A note from the Editor. Journal for the History of Astronomy, 46(1):3, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Evans:2017:LHL

[Eva17] James Evans. Laplace in his letters. Essay review: Correspondance de Pierre Simon Laplace (1749–1827), 2 vols. Edited by Roger Hahn (Brepols, Turnhout, 2013). Pp. xiv + 1416. EUR 120. ISBN 978-2-503-54129-7. Journal for the History of Astronomy, 48(3):365–367, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617709373.

Exton:1992:FAS

[Ext92] Harold Exton. A fresh analysis of some recent data on atmospheric refraction near the horizon with implications in archaeoastronomy. Journal for the History of Astronomy, 23 (17):S57–S58, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Engvold:2016:PWC

[EZ16] Oddbjørn Engvold and Jack B. Zirker. The parallel worlds of Christoph Scheiner and Galileo Galilei. Journal for the History of Astronomy, 47(3):332– 345, August 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Farhi:1997:SET

[FA97] Brian Farhi and Anthony F. Aveni. Solar eclipses and the Temple of the Plumed Serpents, Xochicalco, Mexico. Journal for the History of Astronomy, 28(22):S83–S85, February 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fabricius:2019:PDV

[Fab19] Claus Fabricius. Paris 1676: The discovery of the velocity of light and the roles of Rømer and Cassini. Journal for the History of Astronomy, 50(4):456–475, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619869445.

Faidit:1994:OL

[Fai94] Jean-Michel Faidit. The observatories of Languedoc. Journal for the History of Astronomy, 25(3):199–206, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Falk:2014:SSN

[Fal14] Dan Falk. The Science of Shakespeare: a New Look at the Playwright’s Universe. Thomas Dunne Books/St. Martin’s Press, New York, NY, USA, 2014. ISBN 1-250-00877-8 (hardcover), 1-250-00878-6 (ebook). xv + 364 pp. LCCN PR3047 .F35 2014.

Falk:2021:CKF

[Fal21] Seb Falk. ‘El Capri Kylex’:

a Franciscan astronomical mnemonic. Journal for the History of Astronomy, 52(3): 267–288, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211026210.

Falk:2023:BRJ

[Fal23] Seb Falk. Book revew: John of Lignères as a table compiler: The Tables of 1322 by John of Lignères: An Edition with Commentary. Chabás José and Saby Marie-Madeleine (Brepols, Turnhout, 2022). Pp. 159. 65. ISBN 978-2-503-59609-9. Open access at doi

10.1484/M.ALFA-EB.5.12503.

Journal for the History of Astronomy, 54(4):481–483, November 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231208663.

Fantoli:2005:BRG

[Fan05] Annibale Fantoli. Book review: Galileo and Rome: Galileo in Rome. The Rise and Fall of a Troublesome Genius. Journal for the History of Astronomy, 36(1):109–111, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fantoli:2009:BRM

[Fan09] Annibale Fantoli. Book review: More on the trial of Galileo, Galileo’s Inquisition Trial Revisited. Journal for the History of Astronomy, 40(3):358– 360, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Farrah:1993:MAL

[Far93] Robert W. E. Farrah. The megalithic astronomy of Lundy: Evidence for the remains of a solar calendar. Journal for the History of Astronomy, 24(18): S69–S72, February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fara:1996:LDL

[Far96] Patricia Fara. Lord Derwentwater’s lights: Prediction and the Aurora Polaris. Journal for the History of Astronomy, 27 (3):239–258, August 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fara:2004:HBN

[Far04] Patricia Fara. Heavenly bodies:

Newtonianism, natural theology and the plurality of worlds debate in the Eighteenth Century. Journal for the History of Astronomy, 35(2):143–160, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Farrell:2005:DYL

[Far05] John Farrell. The day without yesterday: Lemaître, Einstein, and the birth of modern cosmology. Thunder’s Mouth Press, New York, NY, USA, 2005. ISBN 1-56025-660-5, 1-56025-902-7 (paperback). viii + 262 pp. LCCN QB991.E94 F37

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0831/2006272995-d.html.

Fatoohi:1998:CPC

[Fat98] Louay J. Fatoohi. A computer program for the conversion of Babylonian into Julian dates. Journal for the History of Astronomy, 29(4):378–379, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fatoohi:2000:BRS

[Fat00] Louay Fatoohi. Book review: Software for calendar conversion: Kairos, General Calendar Conversion Program, for Windows 9x or NT; current version 2.3. Journal for the History of Astronomy, 31(2):180– 181, May 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feast:2000:SPD

[Fea00] Michael Feast. Stellar populations and the distance scale: The Baade–Thackeray correspondence. Journal for the History of Astronomy, 31(1): 29–36, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feingold:1984:BRW

[Fei84] Mordechai Feingold. Book review: Wren as scientist: The Mathematical Science of Christopher Wren. Journal for the History of Astronomy, 15 (1):56–58, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feingold:1986:BRR

[Fei86] Mordechai Feingold. Book review: Reactions to Copernicus: In Defense of the Earth’s Centrality and Immobility: Scholastic Reaction to Copernicanism in the Seventeenth Century. Journal for the History of Astronomy, 17 (3):205–207, August 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feingold:1990:BRP

[Fei90] Mordechai Feingold. Book review: The plurality of worlds, the extraterrestrial life debate 1750–1900: The Idea of a Plurality of Worlds from Kant to Lowell. Journal for the History of Astronomy, 21(2):228– 229, May 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feingold:1995:BRN

[Fei95] Mordechai Feingold. Book review: Newton and Halley: Standing on the Shoulders of Giants: a Longer View of Newton and Halley. Journal for the History of Astronomy, 26 (1):88–89, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feingold:2003:BRF

[Fei03] Mordechai Feingold. Book review: The Flamsteed correspondence completed: The Correspondence of John Flamsteed, the First Astronomer Royal, III: 1703–1719. Journal for the History of Astronomy, 34(2):241–243, May 2003.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feist:2013:SMV

[Fei13] Ulrike Feist. Sonne, Mond und Venus: Visualisierungen astronomischen Wissens im frühneuzeitlichen Rom, volume X of Actus et Imago. Akademie-Verlag, Berlin, Germany, 2013. ISBN 3-05-006365-3. viii + 259 pp. LCCN QB33.I8 F457 2013.

Feke:2013:BRA

[Fek13] Jacqueline Feke. Book review: Astrology and cosmology, Astrology and Cosmology in the World’s Religions. Journal for the History of Astronomy, 44 (4):492–493, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feke:2024:ACS

[Fek24] Jacqueline Feke. Aristotle on the celestial spheres. Journal for the History of Astronomy, 55(4):508–510, November 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286241262066.

Feldhay:2002:ERC

[Fel02] Rivka Feldhay. Essay review: Copernicus, Galileo and the inquisition, La Condamnation des Livres Coperniciens et sa Révocation: Á la lumière de Documents inédits des Congregation de l’Index et de l’Inquisition. Journal for the History of Astronomy, 33(3): 280–284, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feuillebois:1975:MBO

[Feu75] Geneviéve Feuillebois. Les manuscrits de la bibliothèque de l’observatoire de Paris. (French) [the manuscripts of the library of the Observatory of Paris]. Journal for the History of Astronomy, 6(1):72–74, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Figueiredo:2020:JMR

[FF20] Fernando B. Figueiredo and João Fernandes. José Monteiro da Rocha (1734–1819) and his 1782 work on the determination of comet orbits. Journal for the History of Astronomy, 51(4):461–481, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620947844.

Huang:2021:MRC

[fH21] Hsiang fu Huang. Machines for representing the cosmos. Journal for the History of Astronomy, 52(3):371–372, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211020937.

Finocchiaro:1990:BRS

[Fin90] Maurice A. Finocchiaro. Book review: Sidereus Nuncius, Sidereus Nuncius, or the Sidereal Messenger. Journal for the History of Astronomy, 21 (4):376–377, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Finocchiaro:1999:BRS

[Fin99] Maurice A. Finocchiaro. Book review: The spread of Copernicanism in Italy: La Diffusione del Copernicanesimo in Italia 1543–1610. Journal for the History of Astronomy, 30 (2):182–183, May 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fink:2008:ISC

[Fin08] Rita Fink. Inca shadow casting observations in Cuzco. Journal for the History of Astronomy, 39(3):357–361, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Finocchiaro:2010:BRS

[Fin10] Maurice A. Finocchiaro. Book review: Sidereus Nuncius in translation: Galileo’s Sidereus Nuncius or a Sidereal Message. Journal for the History of Astronomy, 41(4):525–526, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Finocchiaro:2019:BRB

[Fin19] Maurice A. Finocchiaro. Book review: 1616 is before 1633: Copernicus Banned: The Entangled Matter of the Anti-Copernican Decree of 1616, Natacha Fabbri & Federica Favino, (Leo S. Olschki Editore, Florence, 2018). Pp. xxv + 252. EUR 32. ISBN 978-88-222-6584-5 (pbk). Journal for the History of Astronomy, 50(3):376–378, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619862971.

Forenbaher:2011:NZF

[FJ11] Staso Forenbaher and Alexander Jones. The Nakovana zodiac: Fragments of an astrologer’s board from an Illyrian–Hellenistic cave sanctuary. Journal for the History of Astronomy, 42(4):425–438, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fay:2012:I

[FJ12] Isla Fay and Nicholas Jardine.

Introduction. Journal for the History of Astronomy, 43(1):1– 2, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fleckenstein:1973:BRE

[Fle73] J. O. Fleckenstein. Book review: Euler, Mayer and the Moon: The Euler-Mayer Correspondence. Journal for the History of Astronomy, 4(3): 205–206, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fleck:1992:CBE

[Fle92] Robert C. Fleck, Jr. The comet of Bethlehem: an early Thirteenth-Century representation by Nicholas of Verdun. Journal for the History of Astronomy, 23(2):137–140, May 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fontenrose:1973:SV

[Fon73] Robert Fontenrose. In search of Vulcan. Journal for the History of Astronomy, 4(3): 145–158, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1970:BRH

[For70a] Eric G. Forbes. Book review: History of Astronomy in East Germany: “Forschungen und Publikationen zur Geschichte der Astronomie in der Deutschen Demokratischen Republik”. Journal for the History of Astronomy, 1(2):159, August 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1970:IBL

[For70b] Eric G. Forbes. Index of the Board of Longitude papers at the Royal Greenwich Observatory, Part I. Journal for the History of Astronomy, 1(2): 169–179, August 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1970:TMC

[For70c] Eric G. Forbes. Tobias Mayer’s contributions to the development of lunar theory. Journal for the History of Astronomy, 1 (2):144–154, August 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1971:GDC

[For71a] Eric G. Forbes. Gauss and the discovery of Ceres. Journal for the History of Astronomy, 2(3): 195–199, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1971:IBLa

[For71b] Eric G. Forbes. Index of the board of longitude papers at the Royal Greenwich Observatory, Part II. Journal for the History of Astronomy, 2(1): 58–70, February 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1971:IBLb

[For71c] Eric G. Forbes. Index of the Board of Longitude papers at the Royal Greenwich Observatory, Part III. Journal for the History of Astronomy, 2(2):58– 70, June 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1971:SPF

[For71d] Eric G. Forbes. Schultz’s proposal for finding longitude at sea. Journal for the History of Astronomy, 2(1):35–41, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1973:BRA

[For73] Eric G. Forbes. Book review: Astronomical journals in Germany: Veröffentlichungen der Archenhold-Sternwarte Nr.

5. Journal for the History of Astronomy, 4(1):63–64, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1974:FFG

[For74a] Eric G. Forbes. The foundation of the first Göttingen Observatory: a study in politics and personalities. Journal for the History of Astronomy, 5(1): 22–29, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1974:MMR

[For74b] Eric G. Forbes. The Maskelyne manuscripts at the Royal Greenwich Observatory. Journal for the History of Astronomy, 5(1):67–69, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1976:EAR

[For76] Eric G. Forbes. Early astronomical researches of John Flamsteed. Journal for the History of Astronomy, 7(2):124– 138, June 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1980:TMC

[For80a] Eric G. Forbes. Tobias Mayer’s contributions to observational astronomy. Journal for the History of Astronomy, 11(1): 28–49, February 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbush:1980:JSL

[For80b] William B. Forbush III. Johann Schroeter on the light of the nebulae. Journal for the History of Astronomy, 11(2):111– 113, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1981:BRH

[For81] Eric G. Forbes. Book review: Humboldt–Schumacher correspondence: Briefwechsel zwischen Alexander von Humboldt und Heinrich Christian Schumacher. Zum 200. Geburtstag von H. C. Schumacher. Journal for the History of Astronomy, 12(1):67–68, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1982:BRI

[For82] Eric G. Forbes. Book review: an illustrated general history: Wandel des Weltbildes: Astronomie, Physik und Mesztechnik in der Kulturgeschichte. Journal for the History of Astronomy, 13(3): 214–215, October 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forbes:1983:BRH

[For83]  Eric G. Forbes.  Book

review: The history of geodesy: Erdgestalt, Kosmologie und Weltanschauung: Die Geschichte der Geodäsie als Teil der Kulturgeschichte der Menschheit. Journal for the History of Astronomy, 14(1): 68, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Forcada:2016:ATC

[For16] Miquel Forcada. Astronomy in tenth-century Islamic encyclopaedism. Journal for the History of Astronomy, 47(3): 346–347, August 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Freer:1977:KA

[FQ77] Robert Freer and Jean-Luc Quinio. The Kerlescan alignments. Journal for the History of Astronomy, 8(1):52–54, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Freeman:1975:CPC

[Fre75] P. R. Freeman. Carnac probabilities corrected. Journal for the History of Astronomy, 6 (3):219, October 1975. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Freeman:1977:TSA

[Fre77] P. R. Freeman. Thom’s survey of the Avebury Ring. Journal for the History of Astronomy, 8(2):134–136, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Freese:2014:CCT

[Fre14] Katherine Freese. The cosmic cocktail: three parts dark matter. Science essentials. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15335-3 (hardcover). xi + 250 + 8 pp. LCCN QB791.3 .F75 2014.

Fatoohi:1996:ALE

[FS96] L. J. Fatoohi and F. R.

Stephenson. Accuracy of lunar eclipse observations made by Jesuit astronomers in China. Journal for the History of Astronomy, 27(1):61–67, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Fatoohi:1999:BFV

[FSAD99] Louay J. Fatoohi, F. Richard Stephenson, and Shetha S. Al-Dargazelli. The Babylonian first visibility of the lunar crescent: Data and criterion. Journal for the History of Astronomy, 30(1):51–72, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Feklova:2017:BRC

[FV17] Tatiana Yu Feklova and Natalia M. Vekshina. From the borders of Russia to the capital of China: Russian astronomical researches in China in the Nineteenth Century. Journal for the History of Astronomy, 48(4):417–431, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617725931.

Farkas:2007:NSH

[FZ07] Gábor F. Farkas and Endre Zsoldos. The new star of 1572 and Hungary. Journal for the History of Astronomy, 38(4): 477–486, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goenner:1999:EWG

[G^{+}99] Hubert Goenner et al., editors. The expanding worlds of general relativity, volume 7 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4060-6 (hardcover), 3-7643-4060-6. xv + 512 pp. LCCN QC173.6 .E97 1999.

Gaab:2023:BGK

[Gaa23] Hans Gaab. A biography of Gottfried Kirch. Journal for the History of Astronomy, 54(2):245–246, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231154232.

Galluzzi:1983:BRG

[Gal83] Paolo Galluzzi. Book review:

Galileo the scientist: Galileo at Work: His Scientific Biography. Journal for the History of Astronomy, 14(3):225–227, October 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gansten:2015:BRB

[Gan15] Martin Gansten. Book review: A “beautiful treatise” in verse: The Siddhāntasundara of Jñãnarãja: An English Translation with Commentary. Toke Lindegaard Knudsen (John Hopkins University Press, Baltimore, MD, 2014). Pp. xvii + 351. $89.95. ISBN 978-1-4214-1442-3. Journal for the History of Astronomy, 46 (3):367–368, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gapaillard:2015:HMD

[Gap15] Jacques Gapaillard. By how much did Le Verrier err on the position of Neptune? Journal for the History of Astronomy, 46(1):48–65, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gapaillard:2021:HPM

[Gap21] Jacques Gapaillard. The heliocentric path of the Moon. Journal for the History of Astronomy, 52(4):462–490, November 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211049618.

See correction [Ano22a].

Gattei:2020:CNE

[Gat20]  Stefano Gattei.  Complet-

ing the national edition of Galileo. Journal for the History of Astronomy, 51(3):368– 370, August 1, 2020. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620932668.

Gattei:2024:JH

[Gat24a] Stefano Gattei. John L. Heilbron, 1934–2023. Journal for the History of Astronomy, 55(2):279–284, May 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241246543.

Gattei:2024:MRI

[Gat24b] Stefano Gattei. A much richer idea of modernity. Journal for the History of Astronomy, 55(3):396–400, August 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241232617.

Gaustad:1987:NSW

[Gau87] John E. Gaustad. Newton’s System of the world: a note on the identity of the translator. Journal for the History of Astronomy, 18(2):125–129, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gaulke:2010:BRA

[Gau10] Karsten Gaulke. Book review: Astronomy in the Wolfenbüttel library: Die Sterne Lügen Nicht: Astronomische und Astrologische Literatur in der Herzog August Bibliothek Wolfenbüttel. Journal for the History of Astronomy, 41(2):271–273, May 2010.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gautschy:2012:DOB

[Gau12]  Rita Gautschy.  Directions

of obscuration in Babylonian eclipse records. Journal for the History of Astronomy, 43(4): 479–489, November 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gautschy:2014:BSC

[Gau14] Rita Gautschy. On the Babylonian sighting-criterion for the lunar crescent and its implications for Egyptian lunar data. Journal for the History of Astronomy, 45(1):79–90, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gaulke:2017:BRA

[Gau17] Karsten Gaulke. Book review: Astronomischer Grund: Fundamentum Astronomicum 1588 des Nicolaus Reimers Ursus. Übersetzung aus dem Lateinischen ins Frühneuhochdeutsche (1594) durch Caspar Thierfelder. Journal for the History of Astronomy, 48 (1):128–129, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616660050.

Gautschy:2018:CSB

[Gau18]  Rita Gautschy.  Culminat-

ing stars in Babylonian astronomy. Journal for the History of Astronomy, 49(4): 503–504, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618781367.

Grant:1986:ERK

[GB86] Edward Grant and Volker Bialas. Essay review: Kepler, hypothesis and scientific methodology: The Birth of History and Philosophy of Science: Kepler’s “A Defence of Tycho against Ursus” with Essays on its Provenance and Significance. Journal for the History of Astronomy, 17(2):137– 143, May 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1995:PHY

[GB95] Bernard R. Goldstein and Alan C. Bowen. Pliny and Hipparchus’s 600-year cycle. Journal for the History of Astronomy, 26(2):155–158, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gualandi:2009:SSP

[GB09] Andrea Gualandi and Fabrizio Bònoli. The search for stellar parallaxes and the discovery of the aberration of light:

The observational proofs of the Earth’s revolution, Eustachio Manfredi, and the ‘Bologna Case’. Journal for the History of Astronomy, 40(2):155– 172, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Garcia:2010:SAM

[GB10] A. César González Garcia and Juan Antonio Belmonte. Statistical analysis of megalithic tomb orientations in the Iberian peninsula and neighbouring regions. Journal for the History of Astronomy, 41(2):225–238, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Garcia:2011:THA

[GB11] A. César González García and Juan Antonio Belmonte. Thinking Hattusha: Astronomy and landscape in the Hittite lands. Journal for the History of Astronomy, 42(4): 461–494, November 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1999:OVO

[GC99] Bernard R. Goldstein and José Chabás. An occultation of Venus observed by Abraham Zacut in 1476. Journal for the History of Astronomy, 30(3): 187–200, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2001:MSE

[GC01] Bernard R. Goldstein and José Chabás. The maximum solar equation in the Alfonsine Tables. Journal for the History of Astronomy, 32(4):345–348, November 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2006:IIA

[GC06] Bernard R. Goldstein and Joséa Chabás. Isaac Ibn Al-Hadib and Flavius Mithridates: The diffusion of an Iberian astronomical tradition in the Late Middle Ages. Journal for the History of Astronomy, 37 (2):147–172, May 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2008:TCM

[GC08] Bernard R. Goldstein and José Chabás. Transmission of computational methods within the Alfonsine corpus: The case of the tables of Nicholaus De Heybech. Journal for the History of Astronomy, 39(3):345–355, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2013:PVA

[GC13] Bernard R. Goldstein and José Chabás. Planetary velocities and the astrological month. Journal for the History of Astronomy, 44(4):465–478, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2019:IAA

[GC19] Bernard R. Goldstein and José Chabás. Ibn al-H .adib’s tables for finding true syzygy. Journal for the History of Astronomy, 50(4):428–446, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619872666.

Gingerich:1993:MRJ

[GD93] Owen Gingerich and Jerzy Dobrzycki. The master of the 1550 radices: Jofrancus Offusius. Journal for the History of Astronomy, 24(4):235–253, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2024:AAP

[GDEK24] Owen Gingerich, David De-Vorkin, James Evans, and Richard L. Kremer. Ad astra per aspera: From the sewers of Kansas to Harvard College Observatory. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241227058.

Gislen:2024:AEI

[GE24] Lars Gislén and James Evans.

The artful early instruments of Peter Apian: Ein kunstlich Instrument of 1524, its precursors and its successors. Journal for the History of Astronomy, 55(2):207–278, May 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241241070.

Gee:2018:AFP

[Gee18]  Emma Gee.  Astronomical

facts or poetic form? Review of Astronomical Knowledge Transmission through Illustrated Aratea Manuscripts. Marion Dolan (Springer Nature, Cham, 2017). Pp. xxiv + 464. EUR 155. ISBN 978-3-319-56783-9. Journal for the History of Astronomy, 49(3):396– 397, August 2018. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gefwert:1975:FWA

[Gef75] Christoffer Gefwert. F. W. A.

Argelander in Finland (1823– 1837). Journal for the History of Astronomy, 6(3):209– 211, October 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1987:BRN

[Gen87] Sara Schechner Genuth. Book review: Nineteenth-Century apparatus: Pike’s Illustrated Catalogue of Scientific & Medical Instruments. Journal for the History of Astronomy, 18 (2):140–141, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1990:BSO

[Gen90a] Sara Schechner Genuth. Blazing stars, open minds, and loosened purse strings: Astronomical research and its early Cambridge audience. Journal for the History of Astronomy, 21 (1):9–20, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1990:BRB

[Gen90b] Sara Schechner Genuth. Book review: Biography of Maskelyne, Nevil Maskelyne: The Seaman’s Astronomer. Journal for the History of Astronomy, 21(4):380–381, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1992:AIP

[Gen92a] Sara Schechner Genuth. Astronomical imagery in a passage of Homer. Journal for the History of Astronomy, 23(4):293–298, November 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1992:BRC

[Gen92b] Sara Schechner Genuth. Book review: Comets and their malevolence: The Origin of Comets. Journal for the History of Astronomy, 23(3):215– 217, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Genuth:1996:BRG

[Gen96] Sara Schechner Genuth. Book review: The Gerard Turner Festschrift: Making Instruments Count: Essays on Historical Scientific Instruments Presented to Gerard L ’Estrange Turner. Journal for the History of Astronomy, 27 (2):176–177, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gerovitch:2014:VSS

[Ger14] Slava Gerovitch. Voices of the Soviet space program: cosmonauts, soldiers, and engineers who took the USSR into space. Palgrave Macmillan, New York, NY, 2014. ISBN 1-137-48178-1 (hardcover). xiv + 305 pp. LCCN TL539 .V65 2014.

Gerovitch:2017:CIH

[Ger17] Slava Gerovitch. Cultural images of humans in space. Essay review: Spacefarers: Images of Astronauts and Cosmonauts in the Heroic Era of Spaceflight. Edited by Michael J. Neufeld (Smithsonian Institution Scholarly Press, Washington, DC, 2013). Pp. 256. $29.95. ISBN 978-1-935623-19-9. Journal for the History of Astronomy, 48(3):362–364, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617708701.

Gessner:2019:BRE

[Ges19] Samuel Gessner. Book review: Early portable watches. Journal for the History of Astronomy, 50(4):491–493, November 1, 2019. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619880409.

Gonzalez-Garcia:2014:BRE

[GG14] A. César González-García. Book review: Easter Island: When the Universe Was an Island: Exploring the Cultural and Spiritual Cosmos of Ancient Rapa Nui. Edmundo Edwards and Alexandra Edwards (Hangaroa Press, Easter Island, 2013). Pp. xii + 536. $140. ISBN 978-956-353-131-2. Journal for the History of Astronomy, 45(4):487–488, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gonzalez-Garcia:2003:ODH

[GGCF03] A. César González-García and Lourdes Costa-Ferrer. Orientations of the Dutch Hunebedden. Journal for the History of Astronomy, 34(2):219–226, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gonzalez-Garcia:2006:OTW

[GGCF06] A. César González-García and Lourdes Costa-Ferrer. Orientation of TRB-West megalithic monuments. Journal for the History of Astronomy, 37(4): 417–427, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2007:CCS

[GH07] Bernard R. Goldstein and Giora Hon. Celestial charts and spherical triangles: The unifying power of symmetry. Journal for the History of Astronomy, 38(1):1–14, February 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grasshoff:2024:AAD

[GH24] Gerd Grasshoff and Susanne M.

Hoffmann. An astronomical analysis of the data in the pseudo-Hipparchus palimpsest in the Codex Climaci Rescriptus. Journal for the History of Astronomy, 55(3): 332–349, August 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241256345.

Gingerich:2002:RSS

[GHHB02] Owen Gingerich, Michael Hoskin, David W. Hughes, and J. Neville Birdsall. Review symposium: The star of Bethlehem, The Star of Bethlehem: The Legacy of the Magi. Journal for the History of Astronomy, 33(4):386–394, November 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Galison:2008:ECH

[GHS08] Peter Galison, Gerald James Holton, and S. S. (Silvan S.) Schweber, editors. Einstein for the 21st Century: His Legacy in Science, Art, and Modern Culture. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13520-7 (hardcover). xviii + 363 + 16 pp. LCCN QC16.E5 E446 2008.

Gibbs:1978:BRA

[Gib78] Sharon Gibbs. Book review: an astrolabe kit and booklet, The Planispheric Astrolabe. Journal for the History of Astronomy, 9(3):220–221, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gibbs:1982:BRY

[Gib82]  Sharon Gibbs.  Book re-

view: Your very own nocturnal: Make-it-Yourself Nocturnal. Journal for the History of Astronomy, 13(2):142–143, June 1982. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Giesemann:2020:BRH

[Gie20] Rebecca Giesemann. Book review: Hevelius’ observational instruments. Journal for the History of Astronomy, 51(2):247–248, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620923786.

Gimbel:2012:EJS

[Gim12] Steven Gimbel. Einstein’s Jewish science: physics at the intersection of politics and religion. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0554-7 (hardcover), 1-4214-0575-X (e-book). viii + 245 pp. LCCN QC173.55 .G55 2012.

Gingerich:1970:BRM

[Gin70a] Owen Gingerich. Book review:

Messier’s catalogue: Messier’s Nebulae and Star Clusters. Journal for the History of Astronomy, 1(1):82–83, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1970:SMP

[Gin70b] Owen Gingerich. The satellites of Mars: Prediction and discovery. Journal for the History of Astronomy, 1(2):109–115, August 1970. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1971:AAC

[Gin71] Owen Gingerich. Apianus’s Astronomicum Caesareum and its Leipzig facsimile. Journal for the History of Astronomy, 2(3): 168–177, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1972:BRA

[Gin72] Owen Gingerich. Book review: Astronomy in Egypt: Egyptian Astronomical Texts, Decans, Planets, Constellations and Zodiacs. Journal for the History of Astronomy, 3(3): 217, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1975:BRH

[Gin75a] Owen Gingerich. Book review: Hevelius’s catalogue in manuscript: Johannes Hevelius and His Catalog of Stars. Journal for the History of Astronomy, 6(1):69, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1975:BRB

[Gin75b] Owen Gingerich. Book review: Of bibliographical interest: The Awakening Interest in Science during the First Century of Printing, 1450–1550, Milestones of Science. Journal for the History of Astronomy, 6(1):70, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1976:ERS

[Gin76]  Owen Gingerich.  Essay

review: Sixteenth-Century broadsides, Erschröckliche und warhafftige Wunderzeichen 1543–1586. Faksimiledruck von Einblattdrucken aus der Sammlung Wikiana in der Zentralbibliothek Zürich. Journal for the History of Astronomy, 7 (2):145–150, June 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1977:TOH

[Gin77a] Owen Gingerich. The 1582 “Theorica Orbium” of Hieronymus Vulparius. Journal for the History of Astronomy, 8(1): 38–43, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1977:BRA

[Gin77b] Owen Gingerich. Book review:

Astronomy and Melville’s “Mardi”: That Lonely Game, Melville. Journal for the History of Astronomy, 8(1):68, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1977:BRP

[Gin77c] Owen Gingerich. Book review: Peucer’s Library: Caspar Peucer’s Library: Portrait of a Wittenberg Professor of the Mid-Sixteenth Century. Journal for the History of Astronomy, 8(1):66–67, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1977:BRO

[Gin77d] Owen Gingerich. Book review: Oklahoma history of science catalogue, The Catalogue of the History of Science Collections of the University of Oklahoma Libraries. Journal for the History of Astronomy, 8(3):215, October 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1978:BRC

[Gin78a] Owen Gingerich. Book review: Calendars, Time and the Calendars. Journal for the History of Astronomy, 9(3): 221–222, October 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1978:BRRb

[Gin78b] Owen Gingerich. Book review: Regiomontanus facsimiles, Joannis Regiomontani Opera Collectanea. Journal for the History of Astronomy, 9 (2):146–147, June 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1978:BRRa

[Gin78c] Owen Gingerich. Book review: Reprint of Gunther’s Astrolabes: Astrolabes of the World. Journal for the History of Astronomy, 9(1):69–70, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1980:BRJa

[Gin80a] Owen Gingerich. Book review: James Logan’s library, The Library of James Logan of Philadelphia, 1674–1751. Journal for the History of Astronomy, 11(2):140–141, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1980:BRJb

[Gin80b] Owen Gingerich. Book review: Jost Bürgi at Kassel, Die erste Sternwarte Europas mit ihren Instrumenten und Uhren. 400 Jahre Jost Bürgi in Kassel. Journal for the History of Astronomy, 11(3):212–213, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1980:BRM

[Gin80c] Owen Gingerich. Book review: Mapping the skies, The Sky Explored. Journal for the History of Astronomy, 11(1): 73–75, February 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1981:BRA

[Gin81a] Owen Gingerich. Book review: Almanacs and society: English Almanacs 1500–1800: Astrology & the Popular Press. Journal for the History of Astronomy, 12(1):65–66, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1981:BRR

[Gin81b] Owen Gingerich. Book review: Royal Society portraits: The Royal Society Catalogue of Portraits. Journal for the History of Astronomy, 12(2):152, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1981:BRS

[Gin81c] Owen Gingerich. Book review: The star of Bethlehem: The Star of Bethlehem: an Astronomer’s Confirmation, Der Stern der Weisen: Geschichte oder Legende?, the Star of Bethlehem: a List of References. Journal for the History of Astronomy, 12(3):212–213, October 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1981:ETE

[Gin81d] Owen Gingerich. An early tradition of an extended errata list for Copernicus’s “De Revolutionibus”. Journal for the History of Astronomy, 12(1): 47–52, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1982:BRE

[Gin82a] Owen Gingerich. Book review:

Einstein’s special theory analysed: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911). Journal for the History of Astronomy, 13(2):135, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1982:BRM

[Gin82b] Owen Gingerich. Book review:

Mayer’s life and work: Tobias Mayer (1723–62): Pioneer of Enlightened Science in Germany. Journal for the History of Astronomy, 13(1):72, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1982:BRN

[Gin82c] Owen Gingerich. Book review:

The nova of 1572 and Tycho’s hand: Libro del Nuevo Cometa. Journal for the History of Astronomy, 13(3):224–225, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1982:ICP

[Gin82d] Owen Gingerich. IAU Commission 41 at Patras. Journal for the History of Astronomy, 13(3):227–228, October 1982.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1983:BRPa

[Gin83a] Owen Gingerich. Book review: Plates from the Sixteenth Century: Poklady Starého Hvezdarstvi. Journal for the History of Astronomy, 14(1):67, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1983:BRPb

[Gin83b] Owen Gingerich. Book review: Post-Galilean bibliography: The History of Modern Astronomy and Astrophysics: a Selected, Annotated Bibliography. Journal for the History of Astronomy, 14(2):146–147, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1984:BRK

[Gin84a] Owen Gingerich. Book review: Kepler in Graz: Johannes Kepler und Graz. Journal for the History of Astronomy, 15 (1):52, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1984:BRE

[Gin84b] Owen Gingerich. Book review: The 1979 Einstein centennial: Einstein, the First Hundred Years. Journal for the History of Astronomy, 15 (3):222, October 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1984:PV

[Gin84c] Owen Gingerich. Phases of Venus in 1610. Journal for the History of Astronomy, 15(3): 209–210, October 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1984:TCS

[Gin84d] Owen Gingerich. A Tusi couple from Schöner’s De Revolutionibus? Journal for the History of Astronomy, 15(2):128– 133, June 1984. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:BRA

[Gin85a] Owen Gingerich. Book review:

Astronomical biographies: The Biographical Dictionary of Scientists: Astronomers. Journal for the History of Astronomy, 16(1):59, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:BRHa

[Gin85b] Owen Gingerich. Book review:

Herschel’s Music: The Symphony, 1720–1840. Journal for the History of Astronomy, 16 (1):68, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:BRC

[Gin85c] Owen Gingerich. Book review: Clocks and watches: Revolution in Time. Journal for the History of Astronomy, 16(2): 149, June 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:BRHb

[Gin85d] Owen Gingerich. Book review: Hartner’s legacy: Oriens– Occidens II. Journal for the History of Astronomy, 16(2): 138–140, June 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:BRT

[Gin85e] Owen Gingerich. Book review: Three of a thesis series: Tychonic and Semi-Tychonic World Systems, the Dissolution of the Celestial Spheres, 1595–1650. Journal for the History of Astronomy, 16(1): 54–55, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:DCO

[Gin85f] Owen Gingerich. Did Copernicus owe a debt to Aristarchus? Journal for the History of Astronomy, 16(1):37–42, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1985:HLG

[Gin85g] Owen Gingerich. Halley’s letter to Gregory concerning the “Synopsis”. Journal for the History of Astronomy, 16(3): 223–224, October 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1986:BRH

[Gin86a] Owen Gingerich. Book review: Halley’s comet in art: Fire and Ice: a History of Comets in Art. Journal for the History of Astronomy, 17 (1):62–63, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1986:BRW

[Gin86b] Owen Gingerich. Book review:

Water-clocks and sand-glasses: The Time Museum, Vol. I: Time Measuring Instruments. Part 3: Water-clocks, Sandglasses, Fire-clocks. Journal for the History of Astronomy, 17 (3):211–212, August 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1987:BRS

[Gin87a] Owen Gingerich. Book review: Scientists and their publications: A Historical Catalogue of Scientists and Scientific Books from the Earliest Times to the Close of the Nineteenth Century. Journal for the History of Astronomy, 18 (1):70, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1987:BRP

[Gin87b] Owen Gingerich. Book review: The perfection of circular motion: Circles of God: Theology and Science from the Greeks to Copernicus. Journal for the History of Astronomy, 18 (2):130–132, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1987:BRT

[Gin87c] Owen Gingerich. Book review: Three manuscript catalogues: Catalogus Codicum Manuscriptorum Medii Aevi Latinorum qui in Bibliotheca Jagellonica Cracovie Asservantur, vol. III, Manuscripts of the Dibner Collection in the Dibner Library of the History of Science and Technology of the Smithsonian Institution Libraries, 1726–1799, Catalogo della Corrispondenza degli Astronomi di Brera, vol. I. Journal for the History of Astronomy, 18(3):229–230, August 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:BRA

[Gin88a] Owen Gingerich. Book review: Ageing Aaboe apotheosized: From Ancient Omens to Statistical Mechanics: Essays on the Exact Sciences Presented to Asger Aaboe. Journal for the History of Astronomy, 19(2):142–143, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:BRE

[Gin88b] Owen Gingerich. Book review: Early astronomical textbooks in France: Imprimeurs et Librairies Parisiens du XVIe Siècle: Fascicule Cavellat. Journal for the History of Astronomy, 19(4):274–275, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:BRT

[Gin88c] Owen Gingerich. Book review:

Timely astrolabes: The Time Museum. Journal for the History of Astronomy, 19(4):275, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:HSC

[Gin88d] Owen Gingerich. How Shapley came to Harvard or, snatching the prize from the jaws of debate. Journal for the History of Astronomy, 19(3):201–207, August 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:ST

[Gin88e] Owen Gingerich. Sacrobosco as a textbook. Journal for the History of Astronomy, 19(4): 269–273, November 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1988:ZH

[Gin88f]  Owen Gingerich.  Zdenek

Horský (1929–1988). Journal for the History of Astronomy, 19(3):215, August 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1989:BRD

[Gin89a] Owen Gingerich. Book review: The De’ Dondi astrarium, Johannis de Dondis Astrarium. Journal for the History of Astronomy, 20(1):69–70, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1989:BRT

[Gin89b] Owen Gingerich. Book review: Trigonometry from Tycho’s library, Tychonis Brahe Triangulorum Planorum et Sphaericorum Praxis Arithmetica. Journal for the History of Astronomy, 20(2):141– 142, June 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1989:ERT

[Gin89c] Owen Gingerich. Essay review: Text and context in analysing Copernicus, Mathematical Astronomy in Copernicus’s De Revolutionibus. Journal for the History of Astronomy, 20(2):128–131, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRH

[Gin90a] Owen Gingerich. Book review: Horoscopes and Horoscopic Methods, Horoscopes and History. Journal for the History of Astronomy, 21(4): 368–369, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRM

[Gin90b] Owen Gingerich. Book review:

The Making of Scientific Instruments, Early Scientific Instruments: Europe 1400–1800. Journal for the History of Astronomy, 21(2):219–221, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRC

[Gin90c] Owen Gingerich. Book review:

Christianity and science, God & Nature: Historical Essays on the Encounter between Christianity and Science. Journal for the History of Astronomy, 21(4):373–374, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRE

[Gin90d] Owen Gingerich. Book review: Early printed books in European observatories, Union Catalogue of Printed Books of 15th, 16th and 17th Centuries in European Astronomical Observatories. Journal for the History of Astronomy, 21(3): 301–302, August 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRR

[Gin90e] Owen Gingerich. Book review: R. R. Newton revisited, The Origins of Ptolemy’s Astronomical Parameters. Journal for the History of Astronomy, 21(4):364–365, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:BRS

[Gin90f] Owen Gingerich. Book review: Starry mechanisms: The Kassel globes, Astronomen — Sterne — Gerate: Landgraf Wilhelm IV. und seine sich selbst bewegenden Globen. Journal for the History of Astronomy, 21(2):222, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:TRW

[Gin90g] Owen Gingerich. Through rugged ways to the galaxies. Journal for the History of Astronomy, 21(1):77–88, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1990:TAA

[Gin90h] Owen Gingerich. Two astronomical anniversaries. Journal for the History of Astronomy, 21(1):1–2, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1991:BRA

[Gin91a] Owen Gingerich. Book review:

Apianus facsimile, Instrument Buch. Journal for the History of Astronomy, 22(3):244– 245, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1991:BRC

[Gin91b] Owen Gingerich. Book review: Calculating astronomers, Practical Astronomy with Your Calculator, Astronomy with Your Personal Computer, Astronomical Formulae for Calculators, Astronomical Tables of the Sun, Moon, and Planets, Planetary Programs and Tables from -4000 to +2800, Mathematical Astronomy with a Pocket Calculator, Practical Ephemeris Calculations. Journal for the History of Astronomy, 22(3):241–243, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1991:BRSb

[Gin91c] Owen Gingerich. Book review:

Scottish workshops, Brass and Glass: Scientific Instrument Making Workshops in Scotland. Journal for the History of Astronomy, 22(2):192, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1991:BRSa

[Gin91d] Owen Gingerich. Book review: The Sachs memorial volume, A Scientific Humanist: Studies in Memory of Abraham Sachs. Journal for the History of Astronomy, 22(2):186– 187, May 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1992:BRH

[Gin92a] Owen Gingerich. Book review: Hubble, Van Maanen, and spiral nebulae: The Edwin Hubble Papers: Previously Unpublished Manuscripts on the Extragalactic Nature of Spiral Nebulae. Journal for the History of Astronomy, 23(3):223– 224, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1992:BRD

[Gin92b] Owen Gingerich. Book review: The Drake Festschrift: Nature, Experiment, and the Sciences. Journal for the History of Astronomy, 23(2):149– 150, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1992:BRS

[Gin92c] Owen Gingerich. Book review:

The Saros Cycle in Babylonia: Saros Cycle Dates and Related Babylonian Astronomical Texts. Journal for the History of Astronomy, 23(3):229, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1993:BRS

[Gin93] Owen Gingerich. Book review: Schwarzschild’s collected works: Gesammelte Werke / Collected Works. Journal for the History of Astronomy, 24 (3):227–228, August 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1994:BRT

[Gin94] Owen Gingerich. Book review: Treasures of the Germanisches Museum: Focus Behaim Globus. Journal for the History of Astronomy, 25(4): 327–328, November 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1995:BRL

[Gin95] Owen Gingerich. Book review:

Liebniz’s debt to the Principia: Equivalence and Priority: Newton versus Leibniz, including Leibniz’s Unpublished Manuscripts on the Principia. Journal for the History of Astronomy, 26(1):87–88, February 1995. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1996:BRS

[Gin96a] Owen Gingerich. Book review: a survey of early astronomy: Early Astronomy. Journal for the History of Astronomy, 27 (3):277, August 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1996:BRI

[Gin96b] Owen Gingerich. Book review: Instruments of Stone: Sawai Jai Singh and his Astronomy. Journal for the History of Astronomy, 27(1):87–89, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1996:BRM

[Gin96c] Owen Gingerich. Book review: Memoirs of a Jesuit in China: The Astronomia Europaea of Ferdinand Verbiest, S.J. (Dillingen, 1687): Text, Translation, Notes and Commentaries. Journal for the History of Astronomy, 27(2):183, May 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1996:BRT

[Gin96d] Owen Gingerich. Book review: Throwing out the baby with the bath water: Revolution and Continuity: Essays in the History and Philosophy of Early Modern Science. Journal for the History of Astronomy, 27 (4):365–368, November 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1997:BRF

[Gin97a] Owen Gingerich. Book review:

Frankfurt Jubilee Celebrations: Ad Radices. Journal for the History of Astronomy, 28(2): 180, May 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1997:BRM

[Gin97b] Owen Gingerich. Book review: More on Alvan Clark and his sons: Alvan Clark & Sons, Artists in Optics. Journal for the History of Astronomy, 28(2):179, May 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1997:BRT

[Gin97c]  Owen Gingerich.  Book

review: Two Seventeenth-Century treatises: The Man in the Moone, Observations concerning the Planet Venus. Journal for the History of Astronomy, 28(1):86–87, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1997:BRW

[Gin97d] Owen Gingerich. Book review: Wealth in the Renaissance: Worldly Goods. Journal for the History of Astronomy, 28(3):275, August 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1997:RW

[Gin97e] Owen Gingerich. Richard S. Westfall (1924–1996). Journal for the History of Astronomy, 28(2):184–185, May 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1998:BRR

[Gin98] Owen Gingerich. Book review: The rise of science in the West: The Rise of Early Modern Science: Islam, China, and the West. Journal for the History of Astronomy, 29(4):393–394, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1999:BRAb

[Gin99a] Owen Gingerich. Book review: Art in Islamic scientific instruments: Science, Tools and Magic: The Nasser D. Khalili Collection of Islamic Art, XII. Journal for the History of Astronomy, 30(3):336–337, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1999:BRAa

[Gin99b] Owen Gingerich. Book review: Aspects of the exact sciences in Islam: From Baghdad to Barcelona: Studies in the Islamic Exact Sciences in Honour of Prof. Juan Vernet. Journal for the History of Astronomy, 30(1):74–75, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1999:BRC

[Gin99c] Owen Gingerich. Book review:

Calendar of the Oort archive: The Letters and Papers of Jan Hendrik Oort. Journal for the History of Astronomy, 30(3): 335, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1999:BRJ

[Gin99d] Owen Gingerich. Book review: John Herschel’s correspondence listed: a Calendar of the Correspondence of Sir John Herschel. Journal for the History of Astronomy, 30(2):183, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1999:BRK

[Gin99e] Owen Gingerich. Book review:

Keplerian bibliography: Bibliographia Kepleriana: Verzeichnis der Gedruckten Schriften von und über Johannes Kepler. Journal for the History of Astronomy, 30(2):177–178, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2000:BRA

[Gin00a] Owen Gingerich. Book review:

an Amateur’s view of Galileo:

Galileo: His Science and His Significance for the Future of Man. Journal for the History of Astronomy, 31(4):356–357, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2000:BRC

[Gin00b] Owen Gingerich. Book review: Concordance of Ptolemaic star positions: Der Sternkatalog des Almagest: Die arabisch-Mittelalterliche Tradition. III Gesamtkordanz der Sternkoordinaten. Journal for the History of Astronomy, 31(4):350, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2000:BRD

[Gin00c] Owen Gingerich. Book review: Detroit Observatory: A Creation of His Own: Tappan’s Detroit Observatory. Journal for the History of Astronomy, 31(2):181, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2000:BRI

[Gin00d] Owen Gingerich. Book review: Illuminating Eddington: The Eddington Enigma. Journal for the History of Astronomy, 31 (1):78–79, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2000:BRL

[Gin00e] Owen Gingerich. Book review:

Light on Galileo: The Cambridge Companion to Galileo. Journal for the History of Astronomy, 31(2):175–176, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRRb

[Gin01a] Owen Gingerich. Book review:

a rare form of astrolabe construction: The Melon-Shaped Astrolabe in Arabic Astronomy. Journal for the History of Astronomy, 32(2):173, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRRa

[Gin01b] Owen Gingerich. Book review: a Reverent Clerke’s kalendar: The Kalendarium of John Somer. Journal for the History of Astronomy, 32(1): 84–85, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRC

[Gin01c] Owen Gingerich. Book review:

Calculating the birth of Christ: Anno Domini: The Origins of the Christian Era. Journal for the History of Astronomy, 32 (1):81–82, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRI

[Gin01d] Owen Gingerich. Book review: Illustrating the history of astronomy: The Mapping of the Heavens, Landmarks in Western Science: From Prehistory to the Atomic Age, the Universe Unveiled: Instruments and Images through History. Journal for the History of Astronomy, 32(1):86–87, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRM

[Gin01e] Owen Gingerich. Book review: Mysterious meteorites: The Tucson Meteorites, Their History from Frontier Arizona to the Smithsonian. Journal for the History of Astronomy, 32 (1):87–88, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRS

[Gin01f] Owen Gingerich. Book review: Science and biblical symbolism, The Bible, Protestantism, and the Rise of Natural Science. Journal for the History of Astronomy, 32(4):369, November 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2001:BRF

[Gin01g] Owen Gingerich. Book review: The formation of the elements: The Magic Furnace:

The Search for the Origins of Atoms. Journal for the History of Astronomy, 32(2):181– 182, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2002:BRE

[Gin02a] Owen Gingerich. Book review: Early astronomy explained, Episodes from the Early History of Astronomy. Journal for the History of Astronomy, 33(3):301, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2002:BRI

[Gin02b] Owen Gingerich. Book review:

Instrument makers in Elizabethan London, Elizabethan Instrument Makers: The Origins of the London Trade in Precision Instrument Making. Journal for the History of Astronomy, 33(1):81–82, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2002:BRU

[Gin02c] Owen Gingerich. Book review:

Unusual Galileiana, against the donning of the gown; Enigma, Dialogo sopra i Due Massimi Sistemi del Mondo. Journal for the History of Astronomy, 33 (1):77–78, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2002:DHM

[Gin02d] Owen Gingerich. Donald H. Menzel: Scientist, educator, builder. 11 April 1901–14 December 1976. Journal for the History of Astronomy, 33(2): 93–94, May 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRC

[Gin03a] Owen Gingerich. Book review: Contacts between East and West: From China to Paris: 2000 Years Transmission of Mathematical Ideas. Journal for the History of Astronomy, 34(2):247–248, May 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRE

[Gin03b] Owen Gingerich. Book review: Euclid on spherical astronomy: Euclid’s Phaenomena: a Translation and Study of a Hellenistic Treatise in Sphaerical Astronomy. Journal for the History of Astronomy, 34(2):240, May 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRF

[Gin03c] Owen Gingerich. Book review: Fontana’s telescopic observations: New Observations of Heavenly and Earthly Objects. Journal for the History of Astronomy, 34(1):115, February 2003. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRG

[Gin03d] Owen Gingerich. Book review: Greenwich globes and armillary spheres: Globes at Greenwich: a Catalogue of the Globes and Armillary Spheres in the National Maritime Museum, Greenwich. Journal for the History of Astronomy, 34 (3):339–340, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRH

[Gin03e] Owen Gingerich. Book review:

How to make sundials and astrolabes: Hartmann’s Practika: a Manual for Making Sundials and Astrolabes with Compass and Rule, Written from 1518 to 1528 by Georg Hartmann. Journal for the History of Astronomy, 34(1):117, February 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRMa

[Gin03f] Owen Gingerich. Book review: Maestlin commemorated: Zwischen Copernicus und Kepler: M. Michael Maestlinus Mathematicus Goeppingensis 1550–1631. Journal for the History of Astronomy, 34(2):246–247, May 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:BRMb

[Gin03g] Owen Gingerich. Book review: Measuring the metre: The Measure of All Things: The Seven-Year Odyssey and Hidden Error That Transformed the World. Journal for the History of Astronomy, 34(4): 464–465, November 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2003:GSC

[Gin03h] Owen Gingerich. The Galileo sunspot controversy: Proof and persuasion. Journal for the History of Astronomy, 34(1): 77–78, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2004:BRL

[Gin04a] Owen Gingerich. Book review: The Louvain Instrument Makers: a Catalogue Raisonné of Scientific Instruments from the Louvain School, 1530 to 1600. Journal for the History of Astronomy, 35(1):113–115, February 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2004:BRG

[Gin04b] Owen Gingerich. Book review: Globes of the Florence Museum: Catalogue of Orbs, Spheres and Globes. Journal for the History of Astronomy, 35(4):489–490, November 2004.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2004:BRT

[Gin04c] Owen Gingerich. Book review:

Troublesome stars: Lost Stars: Lost, Missing, and Troublesome Stars from the Catalogues of Johannes Bayer, Nicholas-Louis de Lacaille, John Flamsteed, and Sundry others. Journal for the History of Astronomy, 35(2):244–245, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2004:JD

[Gin04d] Owen Gingerich. Jerzy Dobrzycki (1927–2004). Journal for the History of Astronomy, 35 (3):371–372, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2005:BRIb

[Gin05a] Owen Gingerich. Book review: Images of Copernicus: Die Copernicus-Biographien des 16. bis 18. Jahrhunderts (Nicolaus Copernicus Gesamtausgabe), IX. Journal for the History of Astronomy, 36(3): 343–345, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2005:BRIa

[Gin05b] Owen Gingerich. Book review: Indian astronomical instruments: Astronomical Instruments in the Rampur Raza Library. Journal for the History of Astronomy, 36(1):120– 121, February 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2005:BRJ

[Gin05c] Owen Gingerich. Book review: Jeremiah Horrocks: The Transit of Venus: The Brief, Brilliant Life of Jeremiah Horrocks, Father of British Astronomy. Journal for the History of Astronomy, 36(2):233–234, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2005:ERF

[Gin05d] Owen Gingerich. Essay review: Fresh perspectives on the Galileo affair: Retrying Galileo, 1633–1992, the Church and Galileo. Journal for the History of Astronomy, 36(4): 443–447, November 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2005:WHC

[Gin05e] Owen Gingerich. Was Horrocks a curate? A tangled bibliographical ramble. Journal for the History of Astronomy, 36 (2):231–232, May 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2006:BRC

[Gin06a] Owen Gingerich. Book review:

Correspondence of Gottfried Kirch: Die Korrespondenz des Astronomen und Kalendermachers Gottfried Kirch. Journal for the History of Astronomy, 37(4):485–486, November 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2006:BRM

[Gin06b] Owen Gingerich. Book review:

Mediaeval star names and their implications: Stars and Numbers: Astronomy and Mathematics in the Medieval Arab and Western Worlds. Journal for the History of Astronomy, 37(3):363, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2006:BRT

[Gin06c] Owen Gingerich. Book review: The Toledan tables: The Toledan Tables. Journal for the History of Astronomy, 37(1): 116–117, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2006:PMP

[Gin06d] Owen Gingerich. Publication of materials prepared for GHA IVB. Journal for the History of Astronomy, 37(1):i, February 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2006:SCC

[Gin06e]  Owen Gingerich.  Supple-

ment to the Copernican census. Journal for the History of Astronomy, 37(2):232, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2007:BRA

[Gin07a] Owen Gingerich. Book review: an astrologer-friend of Tycho Brahe, Heinrich Rantzau und die Astrologie. Journal for the History of Astronomy, 38(4): 510–511, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2007:BRS

[Gin07b] Owen Gingerich. Book review: Sources for astronomy/scripture: Le Conflit entre l’Astronomie Nouvelle et l’Écriture Sainte aux XVI. Journal for the History of Astronomy, 38(2):253–254, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2008:BRC

[Gin08] Owen Gingerich. Book review: Copernicus’s Opera Omnia completed, Mikolaj Kopernik Dziela Wszystkie, III: Pisma Pomniejsze [Minor Works]. Journal for the History of Astronomy, 39(3):416–417, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2009:JDN

[Gin09] Owen Gingerich. John David North (1934–2008). Journal for the History of Astronomy, 40 (3):335–337, August 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2010:BRB

[Gin10a] Owen Gingerich. Book review: a bibliography of astrology: Astrologia Opere a Stampa (1472–1900). Journal for the History of Astronomy, 41(1):143–144, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2010:BRM

[Gin10b] Owen Gingerich. Book review:

Messier’s objects in colour: Atlas of the Messier Objects: Highlights of the Deep Sky. Journal for the History of Astronomy, 41(3):419–420, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2010:BRC

[Gin10c] Owen Gingerich. Book review:

The catalogue of a memorable exhibition: Galileo: Images of the Universe from Antiquity to the Telescope. Journal for the History of Astronomy, 41(1): 145–146, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2011:BRI

[Gin11a] Owen Gingerich. Book review: Italian astronomy since Galileo: Astrum 2009. Astronomy and Instruments: Italian Heritage Four Hundred Years after Galileo. Journal for the History of Astronomy, 42(1): 116–117, February 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2011:BRS

[Gin11b] Owen Gingerich. Book review: Science and theology in 16thcentury Germany: Science and Theology in the Reformation: Studies in Theological Interpretation and Astronomical Observation in Sixteenth-Century Germany. Journal for the History of Astronomy, 42(2):267, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2011:BRZ

[Gin11c] Owen Gingerich. Book review: The zodiac of Dendera: The Zodiac of Paris: How an Improbable Controversy over an Ancient Egyptian Artifact Provoked a Modern Debate between Religion and Science. Journal for the History of Astronomy, 42(3):418–419, August 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2012:BRO

[Gin12a] Owen Gingerich. Book review:

an outstanding star atlas, The Cambridge Photographic Star Atlas. Journal for the History of Astronomy, 43(3):359, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2012:BRB

[Gin12b] Owen Gingerich. Book review: Bayer’s Uranometria in facsimile, Uranometria von Johann Bayer 1603. Journal for the History of Astronomy, 43 (2):257–259, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2012:BRE

[Gin12c] Owen Gingerich. Book review:

The early history of the telescope, The Origins of the Telescope. Journal for the History of Astronomy, 43(2):252–253, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingeri:2013:CAW

[Gin13a] Owen Gingeri. Curtis Alan Wilson (1921–2012). Journal for the History of Astronomy, 44(1):93–94, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2013:BRS

[Gin13b] Owen Gingerich. Book review: The Schöner sammelband, A Renaissance Globemaker’s Toolbox: Johannes Schöner and the Revolution of Modern Science, 1475–1550. Journal for the History of Astronomy, 44(3):364–365, August 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2013:TSC

[Gin13c] Owen Gingerich. Transits in the Seventeenth Century and the credentialling of Keplerian astronomy. Journal for the History of Astronomy, 44(3):303– 312, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2014:BRK

[Gin14] Owen Gingerich. Book review: Kepler on astrology: Kepler’s Cosmological Synthesis: Astrology, Mechanism and the Soul, Kepler’s Astrology, Johannes Kepler Gesammelte Werke. Journal for the History of Astronomy, 45(1):137–138, February 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2015:BRS

[Gin15] Owen Gingerich. Book review: Studien zur ‘Sphaera’ des Johannes de Sacrobosco. Journal for the History of Astronomy, 46(1):101–102, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2016:AA

[Gin16a] Owen Gingerich. The abridged Almagest. Journal for the History of Astronomy, 47 (4):448–449, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668752.

Gingerich:2016:LCT

[Gin16b] Owen Gingerich. Longitude for the coffee table. Journal for the History of Astronomy, 47 (2):224–225, May 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2017:TJG

[Gin17] Owen Gingerich. Telescopes and journalism in Galileo’s era. Journal for the History of Astronomy, 48(1):124– 125, February 2017. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616687421.

Gindhart:2020:EMC

[Gin20]  Marion Gindhart.  Early

modern comets and printing. Journal for the History of Astronomy, 51(4): 490–492, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620968632.

Gressot:2022:DRT

[GJ22] Julien Gressot and Romain Jeanneret. Determining the right time, or the establishment of a culture of astronomical precision at Neuchâtel Observatory in the mid-19th century. Journal for the History of Astronomy, 53(1):27– 48, February 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211068572.

Gingerich:1972:AAA

[GKS72] Owen Gingerich, David King, and George Saliba. The ‘Abd Al-A’Imma astrolabe forgeries. Journal for the History of Astronomy, 3(3):188–198, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gleiser:2004:BRK

[Gle04] Marcelo Gleiser. Book review: Kepler in the Dock: Heavenly Intrigue: Johannes Kepler, Tycho Brahe, and the Murder behind One of History’s Greatest Scientific Discoveries. Journal for the History of Astronomy, 35(4):487–489, November 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gleiser:2007:BRR

[Gle07] Marcelo Gleiser. Book review: Relativity’s later years, The Curious History of Relativity: How Einstein’s Theory Was Lost and Found Again. Journal for the History of Astronomy, 38(4):522–524, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Glick:1988:BRT

[Gli88] Thomas F. Glick. Book review: Testing Newton: Los Caballeros del Punto Fijo: Ciencia, política y aventura en la expedición geodésica hispanofrancesa al virreinato del Perú en el siglo XVIII. Journal for the History of Astronomy, 19(1):67–69, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gluch:2022:AC

[Glu22] Sibylle Gluch. Astronomers in the chair. Journal for the History of Astronomy, 53(4): 505–507, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221114462.

Gingerich:2003:BRS

[GM03] Owen Gingerich and Alessandro Massarotti. Book review: a Seventeenth-Century Florentine: Nel Segno di Galileo: Erudizione, Filosofia e Scienza a Firenze nel Seculo XVII. I Trattati Accademici di Vincenzio Capponi (In the Signshadow of Galileo: Scholarship, Philosophy and Science in Florence in the XVII Century). Journal for the History of Astronomy, 34(3):334, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gee:2014:FWD

[GMML14] Brian Gee, Anita McConnell, and A. D. Morrison-Low, editors. Francis Watkins and the Dollond Telescope Patent Controversy. Science, technology, and culture. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-6643-4 (hardcover), 1-4094-6644-2 (e-book), 1-4094-6645-0 (epub). ISSN 1700-1945. xxvi + 392 pp. LCCN QB88 .G44 2014.

Goldsmith:1980:EFH

[GMW80] Maurice Goldsmith, Alan L. (Alan Lindsay) Mackay, and James Woudhuysen, editors. Einstein, the first hundred years. Pergamon, New York, NY, USA, 1980. ISBN 0-08-025019-X. xiii + 200 pp. LCCN QC16.E5 E52 1980.

Goddu:2006:ROC

[God06] André Goddu. Reflections on the origin of Copernicus’s cosmology. Journal for the History of Astronomy, 37(1):37– 53, February 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goddu:2007:BRC

[God07]  André Goddu.  Book re-

view: Copernicus in person: Poszukiwanie grobu Mikolaja Kopernika. (Polish) [The search for the grave of Nicolaus Copernicus]. Journal for the History of Astronomy, 38 (2):255–256, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goddu:2013:BRM

[God13] André Goddu. Book review:

The Moon in western culture, Ksiezyc w nauce i kulturze Zachodu. Journal for the History of Astronomy, 44(3):365– 367, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goddu:2015:DC

[God15]  André Goddu.  Debating

Copernicus. Journal for the History of Astronomy, 46(4): 500–501, November 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goddu:2016:RNP

[God16] André Goddu. Rheticus’ Narratio prima in Polish. Journal for the History of Astronomy, 47(2):216–217, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goddu:2019:BRC

[God19]  André Goddu.  Book re-

view: Copernicus from 30,000 feet. Journal for the History of Astronomy, 50(4): 487–489, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619877781.

Goddu:2024:CT

[God24] André Goddu. Copernicus and Toruń. Journal for the History of Astronomy, 55(1): ??, February 1, 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241230602.

Goenner:2009:BRM

[Goe09] Hubert Goenner. Book review: Mythos Einstein, Einstein for the 21st Century: His Legacy in Science, Art, and Modern Culture. Journal for the History of Astronomy, 40(2):235– 237, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1972:BRP

[Gol72] Bernard R. Goldstein. Book review: Pre-Islamic Persian astrology: The Astrological History of Masha’allah. Journal for the History of Astronomy, 3(3):218–219, October 1972. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldberg:1973:ERM

[Gol73a]  Stanley Goldberg.  Essay

review: The Michelson– Morley experiment: The Etherial Aether: a History of the Michelson–Morley-Miller Aether-Drift Experiments, 1880–1930. Journal for the History of Astronomy, 4(3): 196–199, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1973:BRA

[Gol73b] Bernard R. Goldstein. Book review: Alhazen on Ptolemy: Ibn al-Haytham, al-Shukuk ‘ala Batlamyus (Dubitationes in Ptolemaeum). Journal for the History of Astronomy, 4(2): 138–139, June 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1975:LBG

[Gol75] Bernard R. Goldstein. Levi Ben Gerson’s analysis of precession. Journal for the History of Astronomy, 6(1):31–41, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1976:ERA

[Gol76] Bernard R. Goldstein. Essay review: The Almagest analysed, A Survey of the Almagest. Journal for the History of Astronomy, 7(1):54–56, February 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1977:LBG

[Gol77] Bernard R. Goldstein. Levi Ben Gerson: On instrumental errors and the transversal scale. Journal for the History of Astronomy, 8(2):102– 112, June 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1982:PRP

[Gol82] S. J. Goldstein, Jr. Problems raised by Ptolemy’s lunar tables. Journal for the History of Astronomy, 13(3):195–205, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrections [Gol84].

Goldstein:1984:CPR

[Gol84] S. J. Goldstein, Jr. Corrections to: “Problems raised by Ptolemy’s lunar tables” [J. Hist. Astronom. 13 (1982), no. 3, 195–205; MR0685069 (84i:01021)]. Journal for the History of Astronomy, 15(2): 134–135, June 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Gol82].

Goldsmith:1987:BRV

[Gol87a] Donald W. Goldsmith. Book review: Velikovsky and after: Beyond Velikovsky: The History of a Public Controversy.

Journal for the History of Astronomy, 18(1):72–74, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1987:BRA

[Gol87b] Samuel J. Goldstein, Jr. Book review: The Ashbrook legacy: The Astronomical Scrapbook: Skywatchers, Pioneers, and Seekers in Astronomy. Journal for the History of Astronomy, 18(2):147–149, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1994:HPC

[Gol94] Bernard R. Goldstein. Historical perspectives on Copernicus’s account of precession. Journal for the History of Astronomy, 25(3):189–197, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1995:BRA

[Gol95] Bernard R. Goldstein. Book review: Astronomy and astrology in Babylon: Die Rolle der Astronomie in den Kulturen Mesopotamiens. Journal for the History of Astronomy, 26(2):170–171, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1996:LBG

[Gol96a] Bernard R. Goldstein. Levi Ben Gerson and the brightness of Mars. Journal for the History of Astronomy, 27(4):297–300, November 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1996:PTT

[Gol96b] Bernard R. Goldstein. The pre-telescopic treatment of the phases and apparent size of Venus. Journal for the History of Astronomy, 27(1):1–12, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1997:SPB

[Gol97] Bernard R. Goldstein. Saving the phenomena: The background to Ptolemy’s planetary theory. Journal for the History of Astronomy, 28(1):1–12, February 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:1998:AZM

[Gol98] Bernard R. Goldstein. Abraham Zacut and the medieval Hebrew astronomical tradition. Journal for the History of Astronomy, 29(2):177–186, May 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2000:BRU

[Gol00] Bernard R. Goldstein. Book review: Unpublished Greek papyri: Astronomical Papyri from Oxyrhynchus. Journal for the History of Astronomy, 31 (1):82–84, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2001:BSC

[Gol01a] Bernard R. Goldstein. Before the Sun in the Church. Journal for the History of Astronomy, 32(1):73–77, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2001:KHA

[Gol01b] Bernard R. Goldstein. Kepler and Hebrew astronomical tables. Journal for the History of Astronomy, 32(2):130–136, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2002:COH

[Gol02a] Bernard R. Goldstein. Copernicus and the origin of his heliocentric system. Journal for the History of Astronomy, 33 (3):219–235, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2002:BDP

[Gol02b] Bernard R. Goldstein. On the Babylonian discovery of the periods of lunar motion. Journal for the History of Astronomy, 33(1):1–13, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2003:AMV

[Gol03a] Bernard R. Goldstein. Ancient and medieval values for the mean synodic month. Journal for the History of Astronomy, 34(1):65–74, February 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Golub:2003:BRS

[Gol03b]  Leon Golub.  Book re-

view: Space science in depth: The Century of Space Science. Journal for the History of Astronomy, 34(4):461–462, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2011:BRS

[Gol11] Bernard R. Goldstein. Book review: The scientific writings of Abraham Ibn Esra: Abraham Ibn Ezra: The Book of Reasons, a Parallel Hebrew– English Critical Edition of the Two Versions of the Text, Abraham Ibn Ezra’s Astrological Writings, I, Études sur le Judaïsme Médiéval, Abraham Ibn Ezra: The Book of the World, a Parallel Hebrew– English Critical Edition of the Two Versions of the Text, Abraham Ibn Ezra’s Astrological Writings, II, Études sur le Judaïsme Médiéval. Journal for the History of Astronomy, 42(1):120–121, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Goldstein:2012:BRA

[Gol12] Bernard R. Goldstein. Book review: Abraham Ibn Ezra on elections, Abraham Ibn Ezra on Elections, Interrogations, and Medical Astrology: a Parallel Hebrew–English Critical Edition of the Book of Elections, the Book of Interrogations and the Book of the Luminaries, Abraham Ibn Ezra’s Astrological Writings, Études sur le Judaïsme Médiéval. Journal for the History of Astronomy, 43(2):253–255, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gooday:2009:BRG

[Goo09] Graeme Gooday. Book review: Global time zones, One Time Fits All: The Campaigns for Global Uniformity. Journal for the History of Astronomy, 40(1):130–132, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gossin:2016:MM

[Gos16]  Pamela Gossin.  Milton’s

mashups. Journal for the History of Astronomy, 47(2):214– 216, May 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:1976:ERP

[Gra76] Edward Grant. Essay review:

Physical sciences before the Renaissance: Early Physics and Astronomy. Journal for the History of Astronomy, 7(3): 201–204, October 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grafton:1980:ERC

[Gra80] A. T. Grafton. Essay review: Copernicus without tears, Introductions à l’astronomie de Copernic. Le Commentariolus de Copernic. La Narratio prima de Rheticus. Journal for the History of Astronomy, 11 (1):63–68, February 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grafton:1990:BRG

[Gra90a] Anthony Grafton. Book review: Gallic calendars, Recueil des Inscriptions Gauloises. Journal for the History of Astronomy, 21(4):366–367, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:1990:ERC

[Gra90b] Edward Grant. Essay review: Chaucer illuminated, Chaucer’s Universe. Journal for the History of Astronomy, 21(2):215–217, May 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:1997:MCS

[Gra97] Edward Grant. The medieval Cosmos: Its structure and operation. Journal for the History of Astronomy, 28(2):147– 167, May 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:2000:BRD

[Gra00] Edward Grant. Book review:

The dissolution of the Cosmos: The Unmaking of the Medieval Christian Cosmos, 1500–1760: From Solid Heavens to Boundless Aether. Journal for the History of Astronomy, 31(4): 350–352, November 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:2001:BRT

[Gra01] Edward Grant. Book review:

Translation and interpretation: Science in Translation: Movements of Knowledge through Cultures and Time. Journal for the History of Astronomy, 32 (2):174–176, May 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graff:2005:TFM

[Gra05a] Eric Graff. The thirteenth figure in the Munich Computus Zodiac. Journal for the History of Astronomy, 36(3):321– 334, August 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2005:KVR

[Gra05b] Miguel A. Granada. Kepler v. Roeslin on the interpretation of Kepler’s nova: (1) 1604– 1606. Journal for the History of Astronomy, 36(3):299–319, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grant:2005:BRM

[Gra05c] Edward Grant. Book review: Medieval planetary diagrams: Planetary Diagrams for Roman Astronomy in Medieval Europe, ca. 800–1500. Journal for the History of Astronomy, 36 (2):237–238, May 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2006:DTE

[Gra06] Miguel A. Granada. Did Tycho eliminate the celestial spheres before 1586? Journal for the History of Astronomy, 37 (2):125–145, May 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2007:BRO

[Gra07a] Miguel A. Granada. Book review: Origins of the infinite universe, Dieu, l’Univers et la Sphère Infinie: Penser l’Infinité Cosmique à l’Aube de la Science Classique. Journal for the History of Astronomy, 38(4):517–519, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2007:MMN

[Gra07b] Miguel A. Granada. Michael Maestlin and the new star of 1572. Journal for the History of Astronomy, 38(1):99– 124, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2008:ERE

[Gra08a]  Miguel A. Granada.  Es-

say review: Early translations of De Revolutionibus: De Revolutionibus: Die erste deutsche Übersetzung in der Grazer Handschrift. Journal for the History of Astronomy, 39(2):265–271, May 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2008:KBI

[Gra08b] Miguel A. Granada. Kepler and Bruno on the infinity of the universe and of solar systems. Journal for the History of Astronomy, 39(4):469–495, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2009:NHB

[Gra09] Miguel A. Granada. Novelties in the heavens between 1572 and 1604 and Kepler’s unified view of nature. Journal for the History of Astronomy, 40(4): 393–402, November 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graney:2010:TAC

[Gra10] Christopher M. Graney. The telescope against Copernicus: Star observations by Riccioli supporting a geocentric universe. Journal for the History of Astronomy, 41(4):453–467, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grafton:2011:PAD

[Gra11a] Anthony Grafton. Petrus Apianus draws up a calendar. Journal for the History of Astronomy, 42(1):55–72, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2011:ANR

[Gra11b] Miguel A. Granada. After the nova of 1604: Roeslin and Kepler’s discussion on the significance of the celestial novelties (1607–1613). Journal for the History of Astronomy, 42(3): 353–390, August 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2012:NCE

[Gra12a] Miguel Angel Granada, editor. Novas y cometas entre 1572 y 1618: revolución cosmológica y renovación política y religiosa, Biblioteca universitària. Universitat de Barcelona, Publicacions i Edicions, Barcelona, Spain, 2012. ISBN 84-475-3588-6. LCCN QB981 .N787 2012.

Granato:2012:BRA

[Gra12b] Marcus Granato. Book review: Astronomy in Argentina, Los Cielos del Sur: Los Observatórios Astronómicos de Córdoba y de La Plata, 1870– 1920. Journal for the History of Astronomy, 43(2):260–262, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grandin:2012:BRA

[Gra12c] Karl Grandin. Book review:

Astronomy in Finland: History of Astronomy in Finland 1828– 1918. Journal for the History of Astronomy, 43(1):128–129, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graney:2012:SRT

[Gra12d] Christopher M. Graney. Science rather than God: Riccioli’s review of the case for and against the Copernican hypothesis. Journal for the History of Astronomy, 43(2):215– 225, May 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grasshoff:2012:MMM

[Gra12e]  Gerd Grasshoff.  Michael

Maestlin’s mystery: Theory building with diagrams. Journal for the History of Astronomy, 43(1):57–73, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graney:2013:BRH

[Gra13a] Christopher M. Graney. Book review: Horrocks on the transit of Venus: Venus Seen on the Sun: The First Observation of a Transit of Venus by Jeremiah Horrocks. Journal for the History of Astronomy, 44(1):119–121, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graney:2013:SAG

[Gra13b] Christopher M. Graney. Stars as the armies of God: Lansbergen’s incorporation of Tycho Brahe’s star-size argument into the Copernican theory. Journal for the History of Astronomy, 44(2):165–172, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graham:2014:BRG

[Gra14a] Daniel W. Graham. Book review: Greek cosmology: Heaven and Earth in Ancient Greek Cosmology: From Thales to Heraclides Ponticus. Journal for the History of Astronomy, 45(1):131–132, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Granada:2014:MMH

[Gra14b] Miguel A. Granada. Michael Maestlin and his unpublished treatise on the nova of 1604. Journal for the History of Astronomy, 45(1):91–122, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Graney:2019:HME

[Gra19a] Christopher M. Graney. How to make the Earth orbit the Sun in 1614. Journal for the History of Astronomy, 50 (1):16–30, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618818643.

Graney:2019:SUJ

[Gra19b] Christopher M. Graney. The starry universe of Johannes Kepler. Journal for the History of Astronomy, 50 (2):155–173, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619847535.

Granada:2020:SC

[Gra20] Miguel Ángel Granada. Speaking of comets. Journal for the History of Astronomy, 51(4): 488–490, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620958852.

Granada:2021:BER

[Gra21a] Miguel Ángel Granada. The book everybody read. Journal for the History of Astronomy, 52(3):352–365, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211011273.

Graney:2021:SUJ

[Gra21b] Christopher M. Graney. The starry universe of Jacques Cassini: Century-old echoes of Kepler. Journal for the History of Astronomy, 52(2): 147–167, May 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211007954.

Granada:2022:JKS

[Gra22]  Miguel Á. Granada.  Jo-

hannes Kepler. the Sun as the heart of the world. Journal for the History of Astronomy, 53(2):133–140, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221082044.

Granada:2024:JC

[Gra24] Miguel Á. Granada. José Chabás, 1948–2024. Journal for the History of Astronomy, 55(4):504–506, November 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241273953.

Green:1999:BRC

[Gre99] Daniel W. E. Green. Book review: Comets as signs and causes: Comets, Popular Culture, and the Birth of Modern Cosmology. Journal for the History of Astronomy, 30 (3):321–323, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Green:2000:BRS

[Gre00] Daniel W. E. Green. Book review: a string of Sixteenth-Century comets: Die Kometen der Jahre 1531 bis 1539 und ihre Bedeutung für die spätere Entwicklung der Kometenforschung. Journal for the History of Astronomy, 31(3): 267–268, August 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Green:2002:BRA

[Gre02] Daniel W. E. Green. Book review: Another cometography, Cometography: a Catalog of Comets. Vol. I: Ancient–1799. Journal for the History of Astronomy, 33(1):78–81, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Green:2004:BRA

[Gre04] Daniel W. E. Green. Book review: Another cometography: Part Two: Cometography: a Catalog of Comets. Journal for the History of Astronomy, 35 (2):238–240, May 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gregory:2009:BRT

[Gre09] Frederick Gregory. Book review: Thermodynamics and the creation of the universe, Entropic Creation: The Religious Contexts of Thermodynamics and Cosmology. Journal for the History of Astronomy, 40(2):238–239, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Green:2010:FCA

[Gre10] Jonathan Green. The first Copernican astrologer: Andreas Aurifaber’s practica for 1541. Journal for the History of Astronomy, 41(2):157–165, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gregory:2014:BRP

[Gre14] Andrew Gregory. Book review: The Presocratics reassessed: Science Before Socrates: Parmenides, Anaxagoras, and the New Astronomy, ISBN 978-0-19-995978-5. Journal for the History of Astronomy, 45(2): 260–261, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Grell:2016:HGR

[Gre16] Chantal Grell. Hevelius in Gdańsk and the Republic of Letters. Journal for the History of Astronomy, 47(2):226– 228, May 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gressot:2023:ANS

[Gre23] Julien Gressot. Actors, networks and scientific instruments at the Bureau des longitudes. Journal for the History of Astronomy, 54(3):369– 370, August 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231176152.

Grossing:1983:FSC

[Grö83] Helmuth Grössing. Fifteenthand Sixteenth-Century astronomical and mathematical manuscripts in the Austrian national library. Journal for the History of Astronomy, 14 (2):149–151, June 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Guidetti:2013:ASB

[GS13] Fabio Guidetti and Anna Santoni, editors. Antiche stelle a Bisanzio: il codice Vaticano greco 1087, volume 17 of Seminari e convegni. Edizioni della Normale, Pisa, Italia, 2013. ISBN 88-7642-485-7. LCCN QB23 .A58 2013.

Gregory:2000:PAE

[GT00] Andrew Gregory and Virginia Trimble. Plato and Aristotle on eclipses. Journal for the History of Astronomy, 31(3):245– 264, August 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gualandi:2009:BRK

[Gua09] Andrea Gualandi. Book review: Kepler’s astronomy and music, Keplero: Una biografia scientifica. Journal for the History of Astronomy, 40(2):219, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gubbins:1988:BRG

[Gub88] David Gubbins. Book review: Geomagnetic instruments: A Brief History of Geomagnetism and a Catalog of the Collections of the National Museum of American History. Journal for the History of Astronomy, 19(2):141–142, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Guerrini:2012:EPP

[Gue12] Luigi Guerrini. Echoes from the pulpit: a preacher against Galileo’s astronomy (1610– 1615). Journal for the History of Astronomy, 43(4):377–389, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Guicciardini:1985:GS

[Gui85] Niccolo Guicciardini. Gravitation and the stars. Journal for the History of Astronomy, 16(3):221–223, October 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gurshtein:2004:PGK

[Gur04] Alexander Gurshtein. Piotr Grigor’evich Kulikovsky (1910– 2003). Journal for the History of Astronomy, 35(1):120–121, February 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gurshtein:2006:BRV

[Gur06] Alexander A. Gurshtein. Book review: Venus transits: Transits of Venus: New Views of the Solar System and Galaxy. Journal for the History of Astronomy, 37(3):364–365, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1998:TBCa

[GV98a] Owen Gingerich and James R.

Voelkel. Tycho Brahe’s Copernican campaign. Journal for the History of Astronomy, 29(1):1–34, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [GV98b].

Gingerich:1998:TBCb

[GV98b] Owen Gingerich and James R.

Voelkel. Tycho Brahe’s Copernican campaign: Addendum. Journal for the History of Astronomy, 29(4):397, November 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [GV98a].

Gingerich:2003:OPP

[GV03] Owen Gingerich and Albert Van Helden. From Occhiale to printed page: The making of Galileo’s Sidereus Nuncius. Journal for the History of Astronomy, 34(3):251–267, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:2011:HGC

[GvH11] Owen Gingerich and Albert van Helden. How Galileo constructed the moons of Jupiter. Journal for the History of Astronomy, 42(2):259–264, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gingerich:1981:RTA

[GW81] Owen Gingerich and Robert S. Westman. A reattribution of the Tychonic annotations in copies of Copernicus’s “De Revolutionibus”. Journal for the History of Astronomy, 12 (1):53–54, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gysembergh:2022:NEH

[GWZ22] Victor Gysembergh, Peter J. Williams, and Emanuel Zingg.

New evidence for Hipparchus’ Star Catalogue revealed by multispectral imaging. Journal for the History of Astronomy, 53(4):383–393, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221128289.

Garcia:2014:OPS

[GZB14] A. César González García, Mauro Zedda, and Juan Antonio Belmonte. On the orientation of prehistoric Sardinian monuments: a comparative statistical approach. Journal for the History of Astronomy, 45(4):467–481, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Haack:1984:AOE

[Haa84] Steven C. Haack. The astronomical orientation of the Egyptian pyramids. Journal for the History of Astronomy, 15(7):S119–S125, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hadingham:1984:BRC

[Had84] Evan Hadingham. Book review: Crystals in the Sky: An Intellectual Odyssey Involving Chumash Astronomy, Cosmology and Rock Art, Travis Hudson and Ernest Underhay (Ballena Press Anthropological Papers no. 10; Ballena Press, Socorro, New Mexico, 1978). Pp. 163. $8.95. When Stars Came Down to Earth: Cosmology of the Skidi Pawnee Indians of North America. Von Del Chamberlain (Ballena Press Anthropological Papers no. 26, BaHena Press/ Center for Archaeoastronomy, Los Altos, California, and College Park, Maryland, 1983). Pp. 270. $17.95. Journal for the History of Astronomy, 15 (7):S149–S151, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1993:TGT

[HAG93] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. The Tombe Di Giganti and Temples of Nuraghic Sardinia. Journal for the History of Astronomy, 24(18):S1–S26, February 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1994:OCD

[HAG94a] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. Orientations of Corsican dolmens. Journal for the History of Astronomy, 25(4):313–316, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1994:SIA

[HAG94b] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. Studies in Iberian archaeoastronomy: (1) orientations of the megalithic sepulchres of Almeria, Granada and Malaga. Journal for the History of Astronomy, 25(19):S55–S82, February 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1995:FOC

[HAG95a] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. Further orientations of Corsican dolmens. Journal for the History of Astronomy, 26(3): 247–252, August 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1995:SIAa

[HAG95b] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. Studies in Iberian archaeoastronomy: (2) orientations of the Tholos tombs of Almeria. Journal for the History of Astronomy, 26(20):S29–S40, February 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1995:SIAb

[HAG95c] Michael Hoskin, Elizabeth Allan, and Renate Gralewski. Studies in Iberian archaeoastronomy: (3) customs and motives in Andalucia. Journal for the History of Astronomy, 26 (20):S41–S48, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Haine:2021:JHC

[Hai21] John Haine. John Harrison’s clockmaking science. Journal for the History of Astronomy, 52(3):370–371, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211026211.

Hall:1988:BRC

[Hal88] A. Rupert Hall. Book review: Copernican reflections: The Genesis of the Copernican World. Journal for the History of Astronomy, 19(4):276–277, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hallyn:1993:BRF

[Hal93] Fernand Hallyn. Book review: Fludd in context: Robert Fludd and the End of the Renaissance. Journal for the History of Astronomy, 24(3):226– 227, August 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hallyn:2003:BRC

[Hal03] Fernand Hallyn. Book review: Copernicus’s reception in Holland: The Calvinist Copernicans: The Reception of the New Astronomy in the Dutch Republic, 1575–1750. Rienk Vermij (Koninklijke Nederlandse Akademie van Wetenschappen, Amsterdam, 2002). Pp. x +

433. EUR49. ISBN 90-698-3404. Journal for the History of Astronomy, 34(4):453–454, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hamel:2006:BRT

[Ham06] Juergen Hamel. Book review: Tycho’s Nova: Ergo perit coelum…: Die Supernova des Jahres 1572 und die Ueberwindung der aristotelischen Kosmologie. Journal for the History of Astronomy, 37(4):488– 489, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hamel:2009:BRK

[Ham09]  Jürgen Hamel.  Book re-

view: Kiel Observatory, “… eine ausnehmende Zierde und Vortheil”: Geschichte der Kieler Universitätssternwarte und ihrer Vorgängerinnen, 1770–1950. Journal for the History of Astronomy, 40(2): 230–232, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hamel:2011:BRN

[Ham11] Jürgen Hamel. Book review:

Nuremberg astronomy: Astronomie in Nürnberg. Journal for the History of Astronomy, 42(3):415–416, August 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hannah:2009:BRA

[Han09a] Robert Hannah. Book review: Astrology in the West, The Dawn of Astrology: a Cultural History of Western Astrology, vol. I: The Ancient and Classical Worlds. Journal for the History of Astronomy, 40(3):350– 351, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hannah:2009:BRC

[Han09b] Robert Hannah. Book review: Calendars in Ancient Rome, Caesar’s Calendar: Ancient Time and the Beginnings of History. Journal for the History of Astronomy, 40(2):224– 226, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Han:2012:BRH

[Han12] Andrew Han. Book review: Hoyle portrayed, Fred Hoyle: a Life in Science. Journal for the History of Astronomy, 43 (2):259–260, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Han:2016:PRN

[Han16] Andrew Han. Public relations at NASA. Journal for the History of Astronomy, 47(1): 110–112, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
110.full.pdf+html.

Hannah:2021:MAG

[Han21] Robert Hannah. More ancient Greek sundials. Journal for the History of Astronomy, 52(3):374–376, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211011272.

Hartner:1972:ERA

[Har72] Willy Hartner. Essay review:

Astronomy in Japan: A History of Japanese Astronomy: Chinese Background and Western Impact. Journal for the History of Astronomy, 3(2): 139–145, June 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Haramundanis:1973:BRM

[Har73] K. Haramundanis. Book review: Mayan astroarchaeology: Die Zeremonialzentren der Maya. Journal for the History of Astronomy, 4(3): 201–202, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hartner:1975:ERC

[Har75] Willy Hartner. Essay review:

Copernican biography and bibliography: Three Copernican Treatises. Journal for the History of Astronomy, 6(1):62–64, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hartner:1977:ROA

[Har77] Willy Hartner. The Rôle of observations in ancient and medieval astronomy. Journal for the History of Astronomy, 8 (1):1–11, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hartner:1978:ERM

[Har78]  Willy Hartner.  Essay re-

view: Mathematical astronomy in antiquity, A History of Ancient Mathematical Astronomy. Journal for the History of Astronomy, 9(3):201–212, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hartner:1979:YAB

[Har79] Willy Hartner. The Young Avestan and Babylonian calendars and the antecedents of precession. Journal for the History of Astronomy, 10(1):1–22, February 1979. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Harwit:1984:BRL

[Har84]  Martin Harwit.  Book re-

view: The literature of cosmology: Der Mythos der Neuzeit: Das Thema der Mehrheit der Welten in der Literaturund Geistesgeschichte von der Kopernikanischen Wende bis zur Science Fiction. Journal for the History of Astronomy, 15(2):141–142, June 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Harwit:1985:BRE

[Har85] Martin Harwit. Book review:

Early-modern cosmology: La Science de l’Univers à l’Âge du Positivisme: Étude sur les origines de la cosmologie contemporaine. Journal for the History of Astronomy, 16(2): 140–142, June 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Harrison:1992:BRJ

[Har92] Edward R. Harrison. Book review: Jaki on cosmology: Olbers Studies: With Three Unpublished Manuscripts by Olbers, Cosmos in Transition: Studies in the History of Cosmology. Journal for the History of Astronomy, 23(3):222– 223, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Harris:1994:BRJ

[Har94] Steven J. Harris. Book review: a Jesuit’s mathematics lectures: Mathematik im Studienplan der Jesuiten. Journal for the History of Astronomy, 25(4):328–330, November 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Harwit:2013:STU

[Har13] Martin Harwit. In search of the true universe: the tools, shaping, and cost of cosmological thought. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-04406-5 (hardcover), 1-107-62025-2 (paperback). xvii + 393 pp. LCCN QB32 .H27 2013.

Harwit:2014:CA

[Har14] Martin Harwit. Classification in astronomy. Journal for the History of Astronomy, 45(3): 381–383, August 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hashimoto:1987:LAD

[Has87] Keizo Hashimoto. Longomontanus’s Astronomia Danica in China. Journal for the History of Astronomy, 18(2):95– 110, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hashimoto:2011:BRJ

[Has11] Keizo Hashimoto. Book review: a Jesuit astronomer in China: Ferdinand Verbiest and Jesuit Science in 17th-Century China: an Annotated Edition and Translation of the Constantinople Manuscript (1676). Journal for the History of Astronomy, 42(4):537–538, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hasse:2019:BRC

[Has19] Dag Nikolaus Hasse. Book review: Copernicus’s Fifteenth-Century world. Journal for the History of Astronomy, 50(4):485–487, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619871726.

Hatch:1983:BRH

[Hat83] Robert A. Hatch. Book review: Huygens in Paris: Huygens et la France. Journal for the History of Astronomy, 14 (1):61–62, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hatch:1988:BRW

[Hat88] Robert A. Hatch. Book review:

Western astronomy in brief: Histoire de l’Astronomie Occidentale. Journal for the History of Astronomy, 19(1):61, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hatch:2004:BRN

[Hat04] Robert Alan Hatch. Book review: The Newton industry: The Foundations of Newtonian Scholarship. Journal for the History of Astronomy, 35(3): 367–368, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hatch:2008:ERC

[Hat08] Robert Alan Hatch. Essay review: The correspondence of Pierre Gassendi: Lettres Latines, Opera Omnia, and the “Primal Archive”, Pierre Gassendi (1592–1655): Introduction à la Vie Savante, Pierre Gassendi (1592–1655): Lettres Latines. Journal for the History of Astronomy, 39(4): 518–528, November 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hawkins:1974:BRE

[Haw74] Gerald S. Hawkins. Book review: Eclipses in antiquity: Solar and Lunar Eclipses of the Ancient Near East from 3000 B.C. to 0 with Maps. Journal for the History of Astronomy, 5(2):136, June 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hayli:1981:ERC

[Hay81] A. Hayli. Essay review: Cosmology in context: Cosmology, History, and Theology. Journal for the History of Astronomy, 12(2):143–148, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hazen:2002:BRE

[Haz02] Martha L. Hazen. Book review: Exploring plate archives, Treasure-Hunting in Astronomical Plate Archives. Journal for the History of Astronomy, 33(3):303, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hart:1971:HLC

[HB71a] R. Hart and R. Berendzen.

Hubble, lundmark and the classification of non-galactic nebulae. Journal for the History of Astronomy, 2(3):200, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hart:1971:HCN

[HB71b] R. Hart and R. Berendzen.

Hubble’s classification of nongalactic nebulae, 1922–1926. Journal for the History of Astronomy, 2(2):109–119, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1992:WAI

[HB92] Norriss S. Hetherington and Ronald S. Brashear. Walter S. Adams and the imposed settlement between Edwin Hubble and Adriaan Van Maanen. Journal for the History of Astronomy, 23(1):53–56, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Higginbottom:1999:RSN

[HC99] Gail Higginbottom and Roger Clay. Reassessment of sites in Northwest Scotland: a new statistical approach. Journal for the History of Astronomy, 30 (24):S41–S46, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hughes:1984:EHO

[HD84] David W. Hughes and Andrew Drummond. Edmond Halley’s observations of Halley’s Comet. Journal for the History of Astronomy, 15(3):189–197, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2006:WHP

[HD06] Michael Hoskin and David W. Dewhirst. William Herschel and the prehistory of stellar spectroscopy. Journal for the History of Astronomy, 37(4): 393–403, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hessler:2013:SMV

[HDG^{+}13] John W. Hessler, Daniel De Simone, Owen Gingerich, Peter K. Machamer, David Marshall Miller, Paul Needham, and Eileen Adair Reeves, editors. The starry messenger, Venice 1610: “from doubt to astonishment”. Levenger Press in association with the Library of Congress, Delray Beach, FL, USA, 2013. ISBN 1-929154-49-6. LCCN QB41 .G173 2013.

Holton:1982:AEH

[HE82] Gerald James Holton and Yehuda Elkana, editors. Albert Einstein, historical and cultural perspectives: the centennial symposium in Jerusalem. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Hearnshaw:1994:BRA

[Hea94] John Hearnshaw. Book review: The Anglo-Australian Observatory: The Creation of the Anglo-Australian Observatory. Journal for the History of Astronomy, 25(2):147–149, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hearnshaw:1997:BRA

[Hea97] John Hearnshaw. Book review:

Astronomy down under: Explorers of the Southern Sky: a History of Australian Astronomy. Journal for the History of Astronomy, 28(2):177–179, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hearnshaw:1999:BRO

[Hea99] John Hearnshaw. Book review:

The Observatory at the Cape: Royal Observatory, Cape of Good Hope, 1820–1831: The Founding of a Colonial Observatory. Journal for the History of Astronomy, 30(3):332– 333, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hearnshaw:2014:AST

[Hea14] J. B. Hearnshaw. The analysis of starlight: two centuries of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, second edition, 2014. ISBN 1-107-03174-5 (hardcover). xvi + 367 pp. LCCN QB465 .H43 2014.

Hearnshaw:2015:BRS

[Hea15] John Hearnshaw. Book review: The Story of Helium and the Birth of Astrophysics, Biman B. Nath (Springer, New York, 2013). Pp. xi + 274. $29.25 (paperback). ISBN 978-1-4614-5362-8. Journal for the History of Astronomy, 46(3):379– 381, August 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heggie:1978:BRM

[Heg78] Douglas C. Heggie. Book review: Megalithic society: Science and Society in Prehistoric Britain. Journal for the History of Astronomy, 9(1):61–64, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heggie:1982:BRA

[Heg82] Douglas C. Heggie. Book review: Astronomy and Society in Britain during the Period 4000–1500 B.C. Journal for the History of Astronomy, 13(4):

S68–S71, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heggie:1985:BRE

[Heg85] D. C. Heggie. Book review: Edinburgh astronomy: The Story of Astronomy in Edinburgh. Journal for the History of Astronomy, 16(1):63–64, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilbron:1999:BRD

[Hei99] J. L. Heilbron. Book review:

The DSB on Laplace reprinted: Pierre-Simon Laplace, 1749– 1827: a Life in Exact Science. Journal for the History of Astronomy, 30(1):83, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heidarzadeh:2000:BRO

[Hei00] Tofigh Heidarzadeh. Book review: Ottoman astronomy: Osmanli Astronomi Literatürü Tarihi (History of Astronomy Literature during the Ottoman Period). Journal for the History of Astronomy, 31(1):84– 86, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilbron:2008:BRI

[Hei08] J. L. Heilbron. Book review: The invention of the telescope, Galileo’s Glassworks:

The Telescope and the Mirror. Journal for the History of Astronomy, 39(4):530–531, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilbron:2012:ERS

[Hei12] J. L. Heilbron. Essay review: Sidereus Nuncius debated, Galileo’s O., Galileo’s Sidereus Nuncius: a comparison of the proof copy (New York) with other paradigmatic copies, Galileo makes a book: The first edition of the Sidereus Nuncius, Venice 1610. Journal for the History of Astronomy, 43(2):244–246, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heidarzadeh:2014:BRN

[Hei14a] Tofigh Heidarzadeh. Book review: Novae and comets 1572–1618: Novas y Cometas entre 1572 y 1618: Revolución Cosmológica y Renovación Política y Religiosa. Edited by Miguel Á. Granada (Universitat de Barcelona Publicacions i Edicions, Barcelona, 2012). Pp. 363. EUR 32 (paperback). ISBN 978-84-475-3588-0. Journal for the History of Astronomy, 45(4):482–483, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilbron:2014:BRV

[Hei14b] J. L. Heilbron. Book review: Visualizing astronomical knowledge in Rome. Sonne, Mond und Venus: Visualisierungen astronomischen Wissens in frühzeitlichen Rom. Ulrike Feist (Actus et imago, ed. by Horst Bredekamp and Jürgen Trabant, x; Akademie Verlag, Berlin, 2013). Pp. viii + 259. EUR 75. ISBN 978-3-05-006365-2. Journal for the History of Astronomy, 45(4): 485–487, November 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilbron:2016:LSS

[Hei16] John Heilbron. Logic and the sizes of stars. Journal for the History of Astronomy, 47 (2):222–224, May 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Heilen:2022:APA

[Hei22]  Stephan Heilen.  An as-

trological practitioner analyzed. Journal for the History of Astronomy, 53(4): 498–503, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221115498.

Henderson:1975:BRT

[Hen75] Janice Henderson. Book review: Two model astrolabes: The Astrolabe, the Astrolabe. Journal for the History of Astronomy, 6(2):138–140, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Henderson:1976:BRP

[Hen76]  Janice Henderson.  Book

review: The Paris Copernicus symposium: Avant, Avec, Après Copernic. La représentation de l’Univers et ses conséquences épistémologiques. Journal for the History of Astronomy, 7(3):213–215, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hentschel:1992:ETE

[Hen92] Klaus Hentschel. Der Einstein-Turm: Erwin F. Freundlich und die Relativitätstheorie; Ansätze zu einer “dichten Beschreibung” von institutionellen, biographischen und theoriengeschichtlichen Aspekten. (German) [The Einstein Tower: Erwin F. friendly and the Theory of Relativity; Dense approaches to a description of institutional, biographical and theoretical-historical aspects]. Spektrum Akademischer Verlag, Heidelberg, Germany, 1992. ISBN 3-86025-025-6. 192 pp. LCCN QB36.F73 H46 1992. See also English translation [Hen97].

Hentschel:1997:ETI

[Hen97] Klaus Hentschel. The Einstein Tower: an intertexture of dynamic construction, Relativity Theory, and astronomy. Writing science. Stanford University Press, Stanford, CA, USA, 1997. ISBN 0-8047-2824-0. xiv + 226 pp. LCCN QB462.65.

URL http://www.gbv.de/
dms/goettingen/220404755.

pdf. English translation by Ann M. Hentschel of the German original [Hen92].

Hentschel:1999:PMSa

[Hen99a] Klaus Hentschel. Photographic mapping of the solar spectrum 1864–1900, Part I. Journal for the History of Astronomy, 30 (2):93–119, May 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hentschel:1999:PMSb

[Hen99b] Klaus Hentschel. Photographic mapping of the solar spectrum 1864–1900, Part II. Journal for the History of Astronomy, 30 (3):201–224, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Henry:2008:BRT

[Hen08] Holly Henry. Book review:

Thomas Hardy and astronomy; Thomas Hardy’s Novel Universe: Astronomy, Cosmology, and Gender in the Post-Darwinian World. Journal for the History of Astronomy, 39 (3):419–421, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hentschel:2009:BRT

[Hen09] Klaus Hentschel. Book review: Theories of cometary tails, Der Komet in der Entladungsröhre: Eugen Goldstein, Wilhelm Foerster und die Elektrizität im Weltraum. Journal for the History of Astronomy, 40(3):367– 368, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hentschel:2013:BRA

[Hen13] Klaus Hentschel. Book review: Adolf Miethe’s autobiography, Adolf Miethe (1862– 1927): Lebenserinnerungen. Journal for the History of Astronomy, 44(2):223–225, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Henderson:2015:CSV

[Hen15a]  John S. Henderson.  Cal-

endar structures for Venus in Mesoamerican divinatory books: Common approaches to commensuration and correction. Journal for the History of Astronomy, 46(4):387–412, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hentschel:2015:BRU

[Hen15b] Klaus Hentschel. Book review: Updating a handbook on astronomical spectroscopy. The Analysis of Starlight: Two Centuries of Astronomical Spectroscopy, 2nd edition.

John B. Hearnshaw (Cambridge University Press, Cambridge, 2014). Pp. xvi + 367. $120. ISBN 978-1-107-03174-6. Journal for the History of Astronomy, 46(2):242–243, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1971:BGH

[Her71] D. B. Herrmann. B. A. Gould and his Astronomical Journal. Journal for the History of Astronomy, 2(2):98–108, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1972:BRL

[Her72] D. B. Herrmann. Book review:

Lectures by Tobias Mayer: Opera Inedita. Journal for the History of Astronomy, 3(3): 219–220, October 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1973:ELE

[Her73] Dieter B. Herrmann. An exponential law for the establishment of observatories in the Nineteenth Century. Journal for the History of Astronomy, 4 (1):57–58, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1976:BRO

[Her76]  D. B. Herrmann.  Book

review: Observatories as architecture: Sternwarten-Architektur und Geschichte der Astronomischen Observatorien (Europäische Hochschulschriften, Reihe XXXII Astronomie Bd. 1). Journal for the History of Astronomy, 7(3):207–208, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1977:DAL

[Her77]  D. B. Herrmann.  Durch-

musterung of astronomical literature begun. Journal for the History of Astronomy, 8(1):74, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1980:BRI

[Her80] D. B. Herrmann. Book review: an illustrated introduction, Man’s View of the Universe. Journal for the History of Astronomy, 11(3):219–220, October 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1990:BRG

[Her90] D. B. Herrmann. Book review: The Genesis of Kant’s General Natural History, Physik und Physiktheologie des jungen Kant. Journal for the History of Astronomy, 21(3):302– 303, August 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herrmann:1994:BRA

[Her94] Dieter B. Herrmann. Book review: Astronomy in Bavaria: Astronomie in Bayern 1914– 1945. Journal for the History of Astronomy, 25(4):336–337, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Herbst:2020:AA

[Her20] Klaus-Dieter Herbst. Asteroids around 1800. Journal for the History of Astronomy, 51(4):494–495, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620945414.

Herbst:2022:AEB

[Her22] Klaus-Dieter Herbst. Astronomy and enlightenment in Berlin circa 1800. Journal for the History of Astronomy, 53(4):504–505, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221123500.

Hetherington:1974:AVM

[Het74] Norriss S. Hetherington. Adriaan Van Maanen on the significance of internal motions in spiral nebulae. Journal for the History of Astronomy, 5(1): 52–53, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1975:SDI

[Het75] Norriss S. Hetherington. The simultaneous ‘discovery’ of internal motions in spiral nebulae. Journal for the History of Astronomy, 6(2):115– 125, June 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1976:BRR

[Het76a] Norriss S. Hetherington. Book review: Recent Copernican revolutions, The Heritage of Copernicus, Theories “More Pleasing to the Mind”. Journal for the History of Astronomy, 7 (1):63–65, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1976:NSM

[Het76b] Norriss S. Hetherington. New source material on Shapley, Van maanen and lundmark. Journal for the History of Astronomy, 7(1):73–74, February 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1978:BRP

[Het78] Norriss S. Hetherington. Book review: Percival Lowell, Lowell and Mars. Journal for the History of Astronomy, 9(3): 224–225, October 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hetherington:1994:BRS

[Het94] Norriss S. Hetherington. Book review: Science and scientists: Paradigms and Barriers: How Habits of Mind Govern Scientific Beliefs. Journal for the History of Astronomy, 25(4): 325–327, November 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hevly:1994:BRB

[Hev94] Bruce Hevly. Book review:

Building the big telescopes: Pauper & Prince: Ritchey, Hale, & Big American Telescopes. Journal for the History of Astronomy, 25(4):333–334, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hewish:1974:BRH

[Hew74]  A. Hewish.  Book review:

History of radio astronomy: The Evolution of Radio Astronomy. Journal for the History of Astronomy, 5(1):64, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1980:WHM

[HG80] Michael Hoskin and Owen Gingerich. On writing the history of modern astronomy. Journal for the History of Astronomy, 11(2):145–146, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hon:2004:SCG

[HG04] Giora Hon and Bernard R. Goldstein. Symmetry in Copernicus and Galileo. Journal for the History of Astronomy, 35 (3):273–292, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Higginbottom:2023:LOC

[HGGVE^{+}23] Gail Higginbottom, A. César González-García, Benito Vilas-Estévez, Víctor López-López, and Felipe Criado-Boado. Landscape, orientation and celestial phenomena on the ‘Coast of Death’ of NW Iberia. Journal for the History of Astronomy, 54(1): 76–106, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221151163.

Hively:1982:GAP

[HH82] Ray Hively and Robert Horn. Geometry and astronomy in prehistoric Ohio. Journal for the History of Astronomy, 13(4):S1–S20, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hively:1984:HGA

[HH84] Ray Hively and Robert Horn. Hopewellian geometry and astronomy at High Bank. Journal for the History of Astronomy, 15(7):S85–S100, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hadravova:2007:APF

[HH07] Alena Hadravová and Petr Hadrava. Astronomy in Paulerinus’s Fifteenth-Century encyclopaedia Liber Viginti Arcium. Journal for the History of Astronomy, 38(3):305–324, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hanham:2013:HKF

[HH13] Andrew Hanham and Michael Hoskin. The Herschel knighthoods: Facts and fiction. Journal for the History of Astronomy, 44(2):149–164, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1990:OTM

[HHK90] Michael Hoskin, Peter Hochsieder, and Doris Knösel. The orientations of the taulas of Menorca (2): The remaining taulas. Journal for the History of Astronomy, 21(15):S37–S48, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addenda [Hos91b].

Higgitt:2012:BRT

[Hig12a] Rebekah Higgitt. Book review:

Tartu Observatory: Tartu Tähetorn, Tartu Old Observatory. Journal for the History of Astronomy, 43(1):133–134, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Higgitt:2012:BRG

[Hig12b] Rebekah Higgitt. Book review: The Great Melbourne Telescope, The Great Melbourne Telescope. Journal for the History of Astronomy, 43(3):373– 374, August 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Higgitt:2014:CH

[Hig14] Rebekah Higgitt. Caroline Herschel. Journal for the History of Astronomy, 45(3):380– 381, August 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Higgitt:2016:AS

[Hig16]  Rebekah Higgitt.  Astron-

omy in the shadows. Journal for the History of Astronomy, 47(4):425–426, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616668754.

Higgitt:2017:BRC

[Hig17] Rebekah Higgitt. Book review: A Children’s Book on the Greenwich Observatory. Journal for the History of Astronomy, 48(4):481–482, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617724894.

Hilfstein:1977:BRI

[Hil77]  Erna Hilfstein.  Book re-

view: Instruments in Copernicus’s Cracow: Instrumenty astronomiczne na Uniwersytecie Krakowskim w XV wieku. Journal for the History of Astronomy, 8(1):64–66, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hingley:2004:BRC

[Hin04] Peter D. Hingley. Book review:

a century of the AAS: The American Astronomical Society’s First Century. Journal for the History of Astronomy, 35(1):112–113, February 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hirsh:1985:BRP

[Hir85] Richard F. Hirsh. Book review: Pioneering radio astronomy: The Early Years of Radio Astronomy: Reflections Fifty Years after Jansky’s Discovery. Journal for the History of Astronomy, 16(1):66–68, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hirsh:1992:BRA

[Hir92] Richard F. Hirsh. Book review: Americans in space: A Spacefaring Nation: Perspectives on American Space History and Policy. Journal for the History of Astronomy, 23 (1):73–74, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hirsh:2009:BRP

[Hir09] Richard Hirsh. Book review: a Pioneer’s autobiography, Secrets of the Hoary Deep: a Personal History of Modern Astronomy. Journal for the History of Astronomy, 40(1): 132–133, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hively:1981:ERP

[Hiv81] Ray M. Hively. Essay review: Pre-Columbian astronomy: Skywatchers of Ancient Mexico. Journal for the History of Astronomy, 12(3):204–207, October 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hawkins:1982:OCB

[HK82] Gerald S. Hawkins and David A. King. On the orientation of the KA c BA. Journal for the History of Astronomy, 13(2):102– 109, June 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1998:SIAb

[HMA^{+}98] Michael Hoskin, Oriol Mercadal i Fernàndez, Sara Aliaga i Rodrigo, Luix Mari Zaldua, Antón A. Rodríguez Casal, Nuno Miguel Soares, M. Socorro López Plaza, Joao Carlos De Senna-Martinez, Francisco Henriques, Juan Antonio Belmonte, Miguel Lago, Joao Albergaria, Manuel Calado, and Artur Martins. Studies in Iberian archaeoastronomy: (5) orientations of megalithic tombs of Northern and Western Iberia. Journal for the History of Astronomy, 29(23): S39–S87, February 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1991:OBM

[HN91] Michael Hoskin and Juan José Morales Núñez. The orientations of the burial monuments of Menorca. Journal for the History of Astronomy, 22 (16):S15–S42, February 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hockey:1992:SRO

[Hoc92] Thomas A. Hockey. Seeing red: Observations of colour in Jupiter’s equatorial zone on the eve of the modern discovery of the great red spot. Journal for the History of Astronomy, 23(2):93–105, May 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hodge:1977:CDP

[Hod77] John E. Hodge. Charles Dillon Perrine and the transformation of the Argentine National Observatory. Journal for the History of Astronomy, 8(1):12–25, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hogendijk:1988:TIL

[Hog88] Jan P. Hogendijk. Three Islamic lunar crescent visibility tables. Journal for the History of Astronomy, 19(1): 29–44, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holton:1977:ERK

[Hol77] Gerald Holton. Essay review: Kepler and his times: Johannes Kepler Gesammelte Werke, Vol. XIX: Documente zu Leben und Werk. Journal for the History of Astronomy, 8(1): 56–59, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano77a].

Holton:1995:BRC

[Hol95]  Gerald Holton.  Book re-

view: Caspar’s Kepler referenced: Kepler. Journal for the History of Astronomy, 26 (2):174–175, May 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holmberg:2003:BRC

[Hol03] Gustav Holmberg. Book review: Cosmology in outline: Fra Kaos til Kosmos: Verdensbilledets Historie Gennem 3000 År (From Chaos to Cosmos: The History of Cosmology during 3000 Years). Journal for the History of Astronomy, 34 (3):337–338, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holberg:2009:DEW

[Hol09] J. B. Holberg. The discovery of the existence of white dwarf stars: 1862 to 1930. Journal for the History of Astronomy, 40(2):137–154, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holberg:2010:SBM

[Hol10a] J. B. Holberg. Sirius B and the measurement of the gravitational redshift. Journal for the History of Astronomy, 41 (1):41–64, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holmberg:2010:BRH

[Hol10b] Gustav Holmberg. Book review: Hands-on history: Astronomers at Work: a Study of the Replicability of 19th Century Astronomical Practice. Journal for the History of Astronomy, 41(1):132–133, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Home:1972:OLC

[Hom72]  Roderick W. Home.  The

origin of the lunar craters: an Eighteenth-Century view. Journal for the History of Astronomy, 3(1):1–10, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hooykaas:1984:RLT

[Hoo84] R. Hooykaas. Rheticus’s lost treatise on Holy Scripture and the motion of the Earth. Journal for the History of Astronomy, 15(2):77–80, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hooykaas:1987:ABC

[Hoo87]  R. Hooykaas.  The Aris-

totelian background to Copernicus’s cosmology. Journal for the History of Astronomy, 18 (2):111–116, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Horsky:1974:BRN

[Hor74] Zdenek Horský. Book review: Nochmals Über Kepler– Diesmal Illustriert: Johann Kepler. Sein Leben in Bildern und eigenen Berichten, Johannes Kepler. Dokumente zu Leben und Werk. Journal for the History of Astronomy, 5 (2):133–134, June 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Horsky:1982:BRR

[Hor82] Zdenek Horský. Book review:

Renaissance cosmological diagrams: The Cosmographical Glass: Renaissance Diagrams of the Universe. Journal for the History of Astronomy, 13(3): 220–222, October 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1970:CTW

[Hos70] Michael Hoskin. The cosmology of Thomas Wright of Durham. Journal for the History of Astronomy, 1(1):44–52, February 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1971:BA

[Hos71a] M. A. Hoskin. Bibliography of astronomy 1881–1898. Journal for the History of Astronomy, 2 (1):53–54, February 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1971:BRU

[Hos71b] M. A. Hoskin. Book review:

Understanding the universe: Widening Horizons: Man’s Quest to Understand the Structure of the Universe. Journal for the History of Astronomy, 2(2):124–125, June 1971.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1972:BRK

[Hos72] M. A. Hoskin. Book review: Kant’s Cosmogony. Journal for the History of Astronomy, 3(1): 67–68, February 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1973:BRA

[Hos73] M. A. Hoskin. Book review: Astronomy after Newton: Astronomical Lectures, the Elements of Physical and Geometrical Astronomy. Journal for the History of Astronomy, 4(3): 204–205, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1975:BRE

[Hos75]  M. A. Hoskin.  Book re-

view: Explaining the Milky Way: The Discovery of Our Galaxy. Journal for the History of Astronomy, 6(2):141– 142, June 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1976:BRD

[Hos76a]  M. A. Hoskin.  Book re-

view: Dictionary or anthology?, Dictionary of the History of Ideas: Studies of Selected Pivotal Ideas. Journal for the History of Astronomy, 7 (1):62–63, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1976:GDW

[Hos76b] M. A. Hoskin. The ‘great debate’: What really happened. Journal for the History of Astronomy, 7(3):169–182, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1976:RCD

[Hos76c] M. A. Hoskin. Ritchey, Curtis and the discovery of novae in spiral nebulae. Journal for the History of Astronomy, 7(1): 47–53, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1977:NPU

[Hos77] M. A. Hoskin. Newton, Providence and the universe of stars. Journal for the History of Astronomy, 8(2):77–101, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1978:ERL

[Hos78a] Michael Hoskin. Essay review: Lambert’s cosmology, Cosmological Letters on the Arrangement of the World-Edifice, Lettres cosmologiques sur l’Organisation de l’Univers. Journal for the History of Astronomy, 9(2):134–139, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1978:LH

[Hos78b] Michael Hoskin. Lambert and Herschel. Journal for the History of Astronomy, 9(2):140– 142, June 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1979:GPQ

[Hos79a] Michael Hoskin. Goodricke, Pigott and the quest for variable stars. Journal for the History of Astronomy, 10(1): 23–41, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1979:WHE

[Hos79b] Michael Hoskin. William Herschel’s early investigations of nebulae: a reassessment. Journal for the History of Astronomy, 10(3):165–176, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1980:JOF

[Hos80a]  M. A. Hoskin.  Joachim

Otto Fleckenstein (1914–1980). Journal for the History of Astronomy, 11(2):144–145, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1980:HDS

[Hos80b] Michael Hoskin. Herschel’s determination of the solar apex.

Journal for the History of Astronomy, 11(3):153–163, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1981:BRT

[Hos81] Michael Hoskin. Book review:

Thomas Wright as architect: Arbours & Grottos. Journal for the History of Astronomy, 12 (1):66–67, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1982:ADM

[Hos82a] Michael Hoskin. Archives of Dunsink and Markree Observatories. Journal for the History of Astronomy, 13(2):146– 152, June 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1982:BRB

[Hos82b] Michael Hoskin. Book review:

Biographies in brief: Concise Dictionary of Scientific Biography. Journal for the History of Astronomy, 13(3):223–224, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1982:FDS

[Hos82c] Michael Hoskin. The first drawing of a spiral nebula. Journal for the History of Astronomy, 13(2):97–101, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1983:BRA

[Hos83a] Michael Hoskin. Book review: Applied history: Cosmic Discovery: The Search, Scope and Heritage of Astronomy. Journal for the History of Astronomy, 14(1):64–65, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1983:BRI

[Hos83b] Michael Hoskin. Book review: The Isis bibliography by periods: Isis Cumulative Bibliography, volumes IV and V: Civilizations and Periods. Journal for the History of Astronomy, 14(1):63–64, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1984:ACW

[Hos84a] Michael Hoskin. Astronomical correspondence of William Rowan Hamilton. Journal for the History of Astronomy, 15 (1):69–73, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1984:BRJ

[Hos84b] Michael Hoskin. Book review: James Ferguson’s publications: A Bibliography of James Ferguson. Journal for the History of Astronomy, 15(3):218–219, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1985:FEH

[Hos85a] Michael Hoskin. The first edition of Halley’s “Synopsis”. Journal for the History of Astronomy, 16(2):133, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1985:OGF

[Hos85b] Michael Hoskin. Obituary: E.

G. Forbes (1933–1984). Journal for the History of Astronomy, 16(1):74, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1985:SCN

[Hos85c] Michael Hoskin. Stukeley’s cosmology and the Newtonian origins of Olbers’s paradox. Journal for the History of Astronomy, 16(2):77–112, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1985:TCM

[Hos85d] Michael Hoskin. The Talayotic culture of Menorca: a first reconnaissance. Journal for the History of Astronomy, 16 (9):S133–S151, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1986:BRN

[Hos86a] Michael Hoskin. Book review:

Newton and scientific revolutions: The Newtonian Revolution. Journal for the History of Astronomy, 17(1):65– 66, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1986:SAH

[Hos86b] Michael Hoskin. Stellar astronomy: Historical studies. Journal for the History of Astronomy, 17(1):38, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1987:JHC

[Hos87] Michael Hoskin. John Herschel’s cosmology. Journal for the History of Astronomy, 18 (1):1–34, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1988:BRC

[Hos88] Michael Hoskin. Book review: Charles Piazzi Smyth: The Peripatetic Astronomer: The Life of Charles Piazzi Smyth. Journal for the History of Astronomy, 19(4):286–287, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1989:AWS

[Hos89a] Michael Hoskin. Astronomers at war: South v. Sheepshanks. Journal for the History of Astronomy, 20(3):175–212, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addenda [Hos91a].

Hoskin:1989:BRN

[Hos89b] Michael Hoskin. Book review:

a new facsimile series, Uranometria, Harmonia macrocosmica seu Atlas universalis et novus, Atlas celeste, Sidereus nuncius, a pronostication euerlasting of ryght goode effecte. Journal for the History of Astronomy, 20(3):223–224, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1990:RRR

[Hos90] Michael Hoskin. Rosse, Robinson, and the resolution of the nebulae. Journal for the History of Astronomy, 21(4):331– 344, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1991:MSV

[Hos91a] Michael Hoskin. More on “South v. Sheepshanks”. Journal for the History of Astronomy, 22(2):174–179, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Hos89a].

Hoskin:1991:OTA

[Hos91b] Michael Hoskin. The orientation of taulas: Addenda. Journal for the History of Astronomy, 22(16):S89–S90, February 1991. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic). See [HHK90].

Hoskin:1992:BRV

[Hos92a] Michael Hoskin. Book review: Vatican centenary: In the Service of Nine Popes: 100 Years of the Vatican Observatory. Journal for the History of Astronomy, 23(1):74, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1992:OMS

[Hos92b]  Michael Hoskin.  Orienta-

tions of megalithic sepulchres in Salamanca, Spain. Journal for the History of Astronomy, 23(1):57–60, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1994:BRM

[Hos94] Michael Hoskin. Book review: Mayans past and present: Archeoastronomia Italiana. Journal for the History of Astronomy, 25(19):S87– S89, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1995:BRF

[Hos95a] Michael Hoskin. Book review: a Fourteenth-Century cycle of syzygies: L’Astronomia de Jacob ben David Bonjorn. Journal for the History of Astronomy, 26(2):173, May 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1995:BRH

[Hos95b] Michael Hoskin. Book review: a history of astronomy for all: The Fontana History of Astronomy and Cosmology. Journal for the History of Astronomy, 26(1):83–84, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1996:BRP

[Hos96] Michael Hoskin. Book review: Proceedings of Oxford 3: Astronomies and Cultures, Archaeoastronomy in the 1990s. Journal for the History of Astronomy, 27(21):S85–S87, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1997:BRA

[Hos97a] Michael Hoskin. Book review:

Astronomy in antiquity and the Middle Ages: Astronomy before the Telescope. Journal for the History of Astronomy, 28 (2):169–170, May 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1997:BRH

[Hos97b] Michael Hoskin. Book review: Hubble with commentary: Hubble’s Cosmology: a Guided Study of Selected Texts. Journal for the History of Astronomy, 28(1):91, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1997:PSM

[Hos97c] Michael Hoskin. A possible solstice marker in Northern Portugal. Journal for the History of Astronomy, 28(1): 79–82, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2000:BRA

[Hos00a] Michael Hoskin. Book review: Astronomy in Mediterranean cultures: Las Leyes del Cielo: Astronomía y Civilizaciones Antiguas. Journal for the History of Astronomy, 31 (25):S87–S88, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2000:BRD

[Hos00b] Michael Hoskin. Book review: Dolmens in France: Les Dolmens: Sociétés Néolithiques et Pratiques Funéraires. Les Sépultures Collectives d’Europe Occidentale, La France des Dolmens et des Sépultures Collectives (4500-2000 avant J.- C.): Bilans Documentaires Régionaux, Mégalithismes de l’Atlantique à l’Ethiopie. Journal for the History of Astronomy, 31(25):S86–S87, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2000:BRN

[Hos00c] Michael Hoskin. Book review: The North Festschrift: Between Demonstration and Imagination: Essays in the History of Science and Philosophy Presented to John D. North. Journal for the History of Astronomy, 31(2):164– 165, May 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2001:BRC

[Hos01a] Michael Hoskin. Book review: Circles codified: The Stone Circles of Britain, Ireland and Brittany. Journal for the History of Astronomy, 32 (26):S89–S90, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2001:BRF

[Hos01b] Michael Hoskin. Book review:

Facing Mecca: La Alquibla en al-Andalus y al-Magrib al– Aqs.à. Journal for the History of Astronomy, 32(2):173–174, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2002:BRL

[Hos02a] Michael Hoskin. Book review:

Lovely pictures, shame about the text, Celestial Treasury: From the Music of the Spheres to the Conquest of Space. Journal for the History of Astronomy, 33(1):76–77, February 2002. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2002:LPA

[Hos02b] Michael Hoskin. The Leviathan of Parsonstown: Ambitions and achievements. Journal for the History of Astronomy, 33 (1):57–70, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2002:ODW

[Hos02c] Michael Hoskin. Orientations of dolmens of West-Central France. Journal for the History of Astronomy, 33(27):S51–S61, February 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2002:SIA

[Hos02d] Michael Hoskin. Studies in Iberian archaeoastronomy: (9) an overview. Journal for the History of Astronomy, 33(27): S75–S82, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2003:DTW

[Hos03a] Michael Hoskin. Discovery of Thomas Wright’s a new theory of the Earth. Journal for the History of Astronomy, 34 (1):94, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2003:HRF

[Hos03b] Michael Hoskin. Herschel’s 40ft reflector: Funding and functions. Journal for the History of Astronomy, 34(1):1–32, February 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2003:ONM

[Hos03c] Michael Hoskin. Orientations of Neolithic monuments of Brittany: (1) context. Journal for the History of Astronomy, 34(4):401–420, November 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:AHF

[Hos04a] Michael Hoskin. Alexander Herschel: The forgotten partner. Journal for the History of Astronomy, 35(4):387–420, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:BRJ

[Hos04b] Michael Hoskin. Book review: John Herschel’s Cape Expedition in context: John Herschel’s Cape Voyage: Private Science, Public Imagination and the Ambitions of Empire. Journal for the History of Astronomy, 35(3):369, August 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:BRM

[Hos04c] Michael Hoskin. Book review: More on Galileo and the Church: Galileo: For Copernicanism and for the Church, Studi Galileiani. Journal for the History of Astronomy, 35 (3):365–367, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:BRN

[Hos04d] Michael Hoskin. Book review: Nineteenth-Century British scientists: The Dictionary of Nineteenth-Century British Scientists. Journal for the History of Astronomy, 35 (4):490–492, November 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:BRTb

[Hos04e] Michael Hoskin. Book review: The temples of Malta: Malta before History: The World’s Oldest Freestanding Stone Architecture. Journal for the History of Astronomy, 35(4):493, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:BRTa

[Hos04f] Michael Hoskin. Book review: Theories of the Milky Way surveyed: La Voie Lactée: Histoire des Conceptions et des Modèles de Notre Galaxie des Temps Anciens aux Années 1930. Journal for the History of Astronomy, 35(2):245–246, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2004:WWH

[Hos04g] Michael Hoskin. Was William Herschel a deserter? Journal for the History of Astronomy, 35(3):356–358, August 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2005:BRE

[Hos05a] Michael Hoskin. Book review: Extracts ancient and modern: Archives of the Universe: a Treasury of Astronomy’s Historic Works of Discovery. Journal for the History of Astronomy, 36(4):458, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2005:CHO

[Hos05b] Michael Hoskin. Caroline Herschel as observer. Journal for the History of Astronomy, 36 (4):373–406, November 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2005:CHS

[Hos05c] Michael Hoskin. Caroline Herschel’s ‘small’ sweeper. Journal for the History of Astronomy, 36(1):28–30, February 2005. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2005:HAE

[Hos05d] Michael Hoskin. The Herschel archive in electronic formats. Journal for the History of Astronomy, 36(2):248, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2005:WHS

[Hos05e] Michael Hoskin. William Herschel’s sweeps for nebulae. Journal for the History of Astronomy, 36(2):230, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2006:BRC

[Hos06a] Michael Hoskin. Book review: Cosmovisions: Songs from the Sky: Indigenous Astronomical and Cosmological Traditions of the World. Journal for the History of Astronomy, 37(1):115, February 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2006:BRS

[Hos06b] Michael Hoskin. Book review: Science in mediaeval Catalonia: La Ciència en la Història des Països Catalans. Journal for the History of Astronomy, 37(1):119–120, February 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2006:CHC

[Hos06c] Michael Hoskin. Caroline Herschel’s catalogue of nebulae. Journal for the History of Astronomy, 37(3):251–255, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2006:CHR

[Hos06d] Michael Hoskin. Caroline Herschel’s revenge. Journal for the History of Astronomy, 37(1): 109–110, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2007:NVB

[Hos07a] Michael Hoskin. Novae and variables before the spectroscope. Journal for the History of Astronomy, 38(3):365– 379, August 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2007:ONMa

[Hos07b]  Michael Hoskin.  Orienta-

tions of Neolithic monuments of Brittany: (2) the early dolmens. Journal for the History of Astronomy, 38(4):487–492, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2007:ONMb

[Hos07c] Michael Hoskin. Orientations of Neolithic monuments of Brittany: (3) the allées couvertes. Journal for the History of Astronomy, 38(4):493–501, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:BRD

[Hos08a] Michael Hoskin. Book review: a DSB of astronomers; Biographical Encyclopedia of Astronomers. Journal for the History of Astronomy, 39(2):272– 274, May 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:BRC

[Hos08b] Michael Hoskin. Book review: Catalan dolmens in context, Els Dòlmens Catalans en el Context de l’Europa Occidental. Journal for the History of Astronomy, 39(2):274–275, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:BRJ

[Hos08c] Michael Hoskin. Book review: John Herschel’s Cape sketches; Cape Landscapes: Sir John Herschel’s Sketches 1834– 1838. Journal for the History of Astronomy, 39(1):129–130, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:DTW

[Hos08d] Michael Hoskin. Derek Thomas Whiteside (1932–2008). Journal for the History of Astronomy, 39(3):402–404, August 2008. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:GIP

[Hos08e] Michael Hoskin. George III’s purchase of Herschel reflectors. Journal for the History of Astronomy, 39(1):121–124, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:GLN

[Hos08f] Michael Hoskin. Gravity and light in the Newtonian universe of stars. Journal for the History of Astronomy, 39(2):251– 264, May 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:NSC

[Hos08g] Michael Hoskin. Nebulae, star clusters and the Milky Way: From Galileo to William Herschel. Journal for the History of Astronomy, 39(3):363– 396, August 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2008:ODW

[Hos08h] Michael Hoskin. Orientations of dolmens of western Europe: Summary and conclusions. Journal for the History of Astronomy, 39(4):507–514, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2010:MHF

[Hos10a] Michael Hoskin. Mary Herschel’s fortune: Origins and impact. Journal for the History of Astronomy, 41(2):213–223, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2010:WHS

[Hos10b] Michael Hoskin. William Herschel and the Southern Skies. Journal for the History of Astronomy, 41(4):503, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2011:BRE

[Hos11a] Michael Hoskin. Book review: Essays on Lalande: Jérôme Lalande (1732–1807): Une Trajectoire Scientifique. Journal for the History of Astronomy, 42(2):278–279, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2011:BRN

[Hos11b] Michael Hoskin. Book review: Nebulae from Herschel to Dreyer: Observing and Cataloguing Nebulae and Star Clusters: From Herschel to Dreyer’s New General Catalogue. Journal for the History of Astronomy, 42(1):122, February 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2011:WHNa

[Hos11c] Michael Hoskin. William Herschel and the nebulae, part 1: 1774–1784. Journal for the History of Astronomy, 42(2):177– 192, May 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2011:WHNb

[Hos11d] Michael Hoskin. William Herschel and the nebulae, Part 2: 1785–1818. Journal for the History of Astronomy, 42(3):321– 338, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2012:BRL

[Hos12a] Michael Hoskin. Book review:

Light on Egypt: Pirámides, Templos y Estrellas: Astronomía y Arqueología en el Egipto Antiguo. Journal for the History of Astronomy, 43 (4):499–500, November 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2012:WHA

[Hos12b] Michael Hoskin. William Herschel’s agenda for his son John. Journal for the History of Astronomy, 43(4):439–454, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2012:WHR

[Hos12c] Michael Hoskin. William Herschel’s residence in Bath, 1799– 1801. Journal for the History of Astronomy, 43(3):351–358, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2013:BRH

[Hos13] Michael Hoskin. Book review: Herschel’s nebulae: The Complete Guide to the Herschel Objects: Sir William Herschel’s Star Clusters, Nebulae and Galaxies. Journal for the History of Astronomy, 44(1): 110–112, February 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2014:AAV

[Hos14a] Michael Hoskin. Avete atque valete. Journal for the History of Astronomy, 45(4):397–399, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2014:CHL

[Hos14b] Michael Hoskin. Caroline Herschel’s life of “mortifications and disappointments”. Journal for the History of Astronomy, 45(4):442–466, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2014:WCH

[Hos14c] Michael Hoskin. William and Caroline Herschel: Pioneers in Late 18th-Century Astronomy. Springerbriefs in Astronomy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-007-6874-5 (paper), 94-007-6875-3 (e-book), 94-007-6876-1 (paper). vii + 114 pp. LCCN

QB35.  URL http://www.
myilibrary.com?id=551470.

Hoskin:2014:WHG

[Hos14d] Michael Hoskin. William Herschel and God. Journal for the History of Astronomy, 45 (2):247–252, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2014:WHP

[Hos14e] Michael Hoskin. William Herschel and the planetary nebulae. Journal for the History of Astronomy, 45(2):209–225, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2015:HTS

[Hos15a] Michael Hoskin. Herschel telescopes for sale. Journal for the History of Astronomy, 46 (3):291–296, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2015:LLJ

[Hos15b] Michael Hoskin. A life and letters of John Michell. Journal for the History of Astronomy, 46(1):105–106, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2015:WHY

[Hos15c] Michael Hoskin. William Herschel’s Yorkshire years. Journal for the History of Astronomy, 46(2):159–172, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2016:HSM

[Hos16a] Michael Hoskin. Harlow Shapley: The making of an observatory director. Journal for the History of Astronomy, 47 (3):317–331, August 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:2016:WHA

[Hos16b] Michael Hoskin. William Herschel’s application for a coat of arms. Journal for the History of Astronomy, 47(2):210– 213, May 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howse:1980:BRN

[How80] Derek Howse. Book review: Nautical astronomy on microfiche, A History of Nautical Astronomical Tables. Journal for the History of Astronomy, 11(3):210–212, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howse:1986:GLO

[How86] Derek Howse. The Greenwich list of observatories: a world list of astronomical observatories, instruments and clocks, 1670–1850. Journal for the History of Astronomy, 17(4):i–89, November 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howse:1988:BRU

[How88] Derek Howse. Book review:

U.S. Naval Observatory celebrations: Sky with Ocean Joined: Proceedings of the Sesquicentennial Symposia of the U.S. Naval Observatory December 5 and 8, 1980. Journal for the History of Astronomy, 19(1):70–71, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howse:1994:GLO

[How94] Derek Howse. The Greenwich list of observatories: a world list of astronomical observatories, instruments and clocks, 1670–1850 amendment list no.

1. Journal for the History of Astronomy, 25(3):207–218, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howell:2002:BRF

[How02] Kenneth J. Howell. Book review: Faith issues in historical context, Science in Theistic Contexts: Cognitive Dimensions. Journal for the History of Astronomy, 33(4):394–397, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howell:2005:BRK

[How05] Kenneth J. Howell. Book review: Kepler and his God: Kepler’s Witch: an Astronomer’s Discovery of Cosmic Order amid Religious War, Political Intrigue, and the Heresy Trial of His Mother. Journal for the History of Astronomy, 36(3): 346–347, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Howell:2008:BRG

[How08] Kenneth J. Howell. Book review: Galileo and the Holy Office: Behind the Scenes at Galileo’s Trial. Journal for the History of Astronomy, 39(1): 128–129, February 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoyt:1981:JJS

[Hoy81] William Graves Hoyt. T. J. J. See and Mercurian craters. Journal for the History of Astronomy, 12(2):139–142, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoyt:1982:GKG

[Hoy82] William Graves Hoyt. G. K. Gilbert’s contribution to selenology. Journal for the History of Astronomy, 13(3):155– 167, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoyt:1984:BRM

[Hoy84] William Graves Hoyt. Book review: Meteor Crater, Arizona: Heroic Efforts at Meteor Crater, Arizona: Selected Correspondence between Daniel Moreau Barringer and Elihu Thomson. Journal for the History of Astronomy, 15(3): 219–220, October 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1998:SIAa

[HP98] Michael Hoskin and Toni Palomo i Pérez. Studies in Iberian archaeoastronomy: (4) The orientations of megalithic tombs of Eastern Catalunya. Journal for the History of Astronomy, 29(1):63–79, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [HP99].

Hoskin:1999:OMT

[HP99] Michael Hoskin and Toni Palomo i Pérez. The orientations of megalithic tombs of Eastern Catalunya: Addendum. Journal for the History of Astronomy, 30(24):S89–S90, February 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [HP98].

Hoskin:1992:TWR

[HR92] Michael Hoskin and George D. Rochester. Thomas Wright and the Royal Society. Journal for the History of Astronomy, 23 (3):167–172, August 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hahn:2013:CPS

[HRVM13] Roger Hahn, Patricia Radelet de Grave, Jan Vandersmissen, and Thierry Mozdziej, editors. Correspondance de Pierre-Simon Laplace (1749– 1827). Tome I, Années 1769– 1802. Tome II, Années 1803– 1827 et lettres non datées, volume 90 of De Diversis Artibus. Brepols, Turnhout, Bélgium, 2013. ISBN 2-503-54846-6 (vol. 1), 2-503-54847-4 (vol. 2), 2-503-54129-1 (set). xiv + 713 (vol. 1), 716–1416 (vol. 2) pp. LCCN Q143.L36 A3 2013.

Hamilton:1981:ERJ

[HS81] N. T. Hamilton and N. M. Swerdlow. Essay review: Judgement on Ptolemy: Ancient Planetary Observations and the Validity of Ephemeris Time. Journal for the History of Astronomy, 12(1):59– 63, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hubbell:1992:NAN

[HS92] John G. Hubbell and Robert W. Smith. Neptune in America:

Negotiating a discovery. Journal for the History of Astronomy, 23(4):261–291, November 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1999:SIA

[HS99] Michael Hoskin and Carme Sauch i Aparicio. Studies in Iberian archaeoastronomy:

(6) orientations of megalithic tombs of Badajoz and neighbouring Portugal. Journal for the History of Astronomy, 30 (24):S35–S40, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holford-Strevens:2016:CR

[HS16] Leofranc Holford-Strevens. Calendar reform in the 1170s. Journal for the History of Astronomy, 47(4):440–441, November 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL

http://journals.sagepub.
com/doi/full/10.1177/0021828616668759.

Holford-Strevens:2018:FCF

[HS18] Leofranc Holford-Strevens. Fixing the calendar in the fifteenth century. Journal for the History of Astronomy, 49(4):502–503, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618807010.

Huber:1983:BRP

[Hub83] Peter J. Huber. Book review: Planetary positions: Ephemeriden von Sonne, Mond und hellen Planeten von -1000 bis -601. Journal for the History of Astronomy, 14(3):228– 229, October 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Huber:1991:BRL

[Hub91] Peter J. Huber. Book review: Late Babylonian astronomical texts, Astronomical Diaries and Related Texts from Babylonia. Journal for the History of Astronomy, 22(2):184– 186, May 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Huffman:1991:USN

[Huf91] Wendell W. Huffman. The United States Naval Astronomical Expedition (1849–52) for the Solar Parallax. Journal for the History of Astronomy, 22 (3):208–220, August 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hufbauer:2006:BRH

[Huf06a] Karl Hufbauer. Book review: The Hubble space telescope: The Hubble Space Telescope: Imaging the Universe. Journal for the History of Astronomy, 37(2):240–241, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hufbauer:2006:SSE

[Huf06b] Karl Hufbauer. Stellar structure and evolution, 1924–1939. Journal for the History of Astronomy, 37(2):203–227, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hughes:2010:BRC

[Hug10]  David W. Hughes.  Book

review: Comets in history: The Greatest Comets in History: Broom Stars and Celestial Scimitars. Journal for the History of Astronomy, 41(4): 518–519, November 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hunger:2018:SCW

[Hun18] Hermann Hunger. “Science”

in the cuneiform world. Journal for the History of Astronomy, 49(1):129–130, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618755446.

Husson:2012:WRT

[Hus12] Matthieu Husson. Ways to read a table: Reading and interpolation techniques in canons of early Fourteenth-Century double-argument tables. Journal for the History of Astronomy, 43(3):299–319, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Husson:2013:BRR

[Hus13] Matthieu Husson. Book review: Raymond of Marseilles: Raymond de Marseille, Opera omnia, i: Traité de l’astrolabe. Liber cursuum planetarum. Journal for the History of Astronomy, 44(1): 107–108, February 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hutchins:2012:BRT

[Hut12] Roger Hutchins. Book review: Touring observatories in 1777: An Observer of Observatories: The Journal of Thomas Bugge’s Tour of Germany, Holland and England in 1777. Journal for the History of Astronomy, 43(1):126–128, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Huth:2013:LAF

[Hut13] John Edward Huth. The lost art of finding our way. The Belknap Press of Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-07282-0 (hardcover). 528 pp. LCCN VK15 .H87 2013.

Hoskin:2001:SIA

[HVT^{+}01] Michael Hoskin, José Manuel Quinta Ventura, Luis Tirapicos, J. Carlos Vera Rodríguez, Martí Mas Cornellá, R. Cruz-Auñón Briones, L. García Sanjuá, J. A. Belmonte, Fernando Pimenta, J. Carlos Vera Rodríguez, and José C. Martín De La Cruz. Studies in Iberian archaeoastronomy: (8) orientations of megalithic and Tholos tombs of Portugal and Southwest Spain. Journal for the History of Astronomy, 32(26): S45–S64, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1981:CHC

[HW81] Michael Hoskin and Brian Warner. Caroline Herschel’s comet sweepers. Journal for the History of Astronomy, 12 (1):27–34, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Holberg:2007:DCS

[HW07] J. B. Holberg and F. Wesemael. The discovery of the companion of Sirius and its aftermath. Journal for the History of Astronomy, 38(2):161– 174, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hysom:1996:TSM

[Hys96] E. J. Hysom. Tests of the shape of mirrors by Herschel. Journal for the History of Astronomy, 27(4):349–352, November 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Hoskin:1997:OSD

[HZ97] Michael Hoskin and Mauro Zedda. Orientations of Sardinian dolmens. Journal for the History of Astronomy, 28(22): S1–S16, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Imaeda:1980:JRG

[IK80] K. Imaeda and T. Kiang. The Japanese record of the gueststar of 1408. Journal for the History of Astronomy, 11(2): 77–80, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Imhausen:2022:IHA

[Imh22]  Annette Imhausen.  Indi-

cating hours in ancient cultures. Journal for the History of Astronomy, 53(1): 125–127, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211043958.

Itokazu:2009:EHP

[Ito09] Anastasia Guidi Itokazu. On the equivalence of hypotheses in Part 1 of Johannes Kepler’s New Astronomy. Journal for the History of Astronomy, 40 (2):173–190, May 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Iund:2002:ODN

[Iun02] Richard Iund. Orientations of dolmens north of the Eastern Pyrénées. Journal for the History of Astronomy, 33(27): S21–S28, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Iwaniszewski:1996:ERA

[Iwa96] Stanisław Iwaniszewski. Essay review: Archaeoastronomy in Spain: Arqueoastronomía Hispánica. Prácticas Astrónomicas en la Prehistoria de la Península Ibérica y los Archipiélagos Balear y Canario. Journal for the History of Astronomy, 27(21):S81–S85, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Iwaniszewski:2003:AAA

[Iwa03] Stanisław Iwaniszewski. Archaeoastronomical analysis of Assyrian and Babylonian monuments: Methodological issues. Journal for the History of Astronomy, 34(1):79–93, February 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaki:1970:NLO

[Jak70] Stanley L. Jaki. New light on Olbers’s dependence on Chéseaux. Journal for the History of Astronomy, 1(1):53–55, February 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaki:1971:MWB

[Jak71] Stanley L. Jaki. The Milky Way before Galileo. Journal for the History of Astronomy, 2(3): 161–167, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaki:1972:MWG

[Jak72a] Stanley L. Jaki. The Milky Way from Galileo to Wright. Journal for the History of Astronomy, 3(3):199–204, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaki:1972:OFT

[Jak72b] Stanley L. Jaki. The original formulation of the Titius– Bode Law. Journal for the History of Astronomy, 3(2):136– 138, June 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jarrell:1973:BRR

[Jar73] Richard A. Jarrell. Book review: Reception of Copernicus: Studia Copernicana, v: Colloquia Copernicana, I. Journal for the History of Astronomy, 4(3):203–204, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jarrell:1975:AAC

[Jar75] Richard A. Jarrell. Astronomical archives in Canada. Journal for the History of Astronomy, 6(2):143–147, June 1975.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jarrell:1977:AAC

[Jar77] Richard A. Jarrell. Astronomical archives in Canada (2). Journal for the History of Astronomy, 8(1):71–72, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:1982:SCO

[Jar82] Nicholas Jardine. The significance of the Copernican ORBS. Journal for the History of Astronomy, 13(3):168–194, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano83b].

Jardine:1990:ERH

[Jar90] N. Jardine. Essay review: How to appropriate a world system: The Wittich Connection: Conflict and Priority in Late Sixteenth-Century Cosmology, Owen Gingerich and Robert S. Westman (Transactions of the American Philosophical Society, lxxviii, Part 7; Philadelphia, 1988). Pp. viii + 148. $15 (paperback). Journal for the History of Astronomy, 21(4): 353–358, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:1991:ERW

[Jar91a] Nicholas Jardine. Essay review: Writing off the scientific revolution: Reappraisals of the Scientific Revolution. Journal for the History of Astronomy, 22(4):311–318, November 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jarrell:1991:ODO

[Jar91b] Richard A. Jarrell. The origins of the Dominion Observatory, Ottawa. Journal for the History of Astronomy, 22 (1):45–53, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:1998:PAE

[Jar98] Nicholas Jardine. The places of astronomy in early-modern culture. Journal for the History of Astronomy, 29(1):49– 62, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jarrell:1999:JPM

[Jar99] Richard A. Jarrell. J. S. Plaskett and the modern large reflecting telescope. Journal for the History of Astronomy, 30 (4):359–390, November 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2000:BRC

[Jar00] Nick Jardine. Book review: Celestial architect: Johannes Kepler and the New Astronomy. Journal for the History of Astronomy, 31(3):265–266, August 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2006:KCH

[Jar06] Nicholas Jardine. Kepler as castigator and historian: His preparatory notes for Contra Ursum. Journal for the History of Astronomy, 37(3):257– 297, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2009:ERA

[Jar09a] Nick Jardine. Essay review:

Astral projections and pies in the sky, Imagination des Himmels. Journal for the History of Astronomy, 40(1):113–120, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2009:KTA

[Jar09b] Nick Jardine. Kepler as transgressor and amalgamator of disciplines. Journal for the History of Astronomy, 40(4): 375–380, November 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2014:KKE

[Jar14] Nicholas Jardine. Kepler = Koestler: On empathy and genre in the history of the sciences. Journal for the History of Astronomy, 45(3):271– 286, August 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaschek:1996:BRI

[Jas96a] Carlos Jaschek. Book review: an introduction to archaeoastronomy: Sterne und Steine. Journal for the History of Astronomy, 27(21):S89, February 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jaschek:1996:BRP

[Jas96b] Carlos Jaschek. Book review: Papers of the Warsaw conference: Readings in Archaeoastronomy: Papers Presented at the International Conference, “Current Problems and Future of Archaeoastronomy”, Held at the State Archaeological Museum in Warsaw in November 1990. Journal for the History of Astronomy, 27(21):S88–S89, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jasinski:2020:OGC

[Jas20] Maciej Jasiński. Otto von Guericke’s cometary theory in Stanisław Lubieniecki’s correspondence. Journal for the History of Astronomy, 51 (2):131–151, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619891273.

Jhirad:1974:BRC

[Jhi74] David Jhirad. Book review:

The concept of time: What is Time? Journal for the History of Astronomy, 5(2):138– 139, June 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2005:TVUb

[JHLS05] Nicholas Jardine, Katherine Harloe, Dieter Launert, and Alain Segonds. Tycho v. Ursus: The build-up to a trial, Part 2. Journal for the History of Astronomy, 36(2):125– 165, May 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See errata [Ano05a].

Jardine:2010:CEE

[JJ10] Boris Jardine and Nicholas Jardine. Critical editing of early-modern astronomical diagrams. Journal for the History of Astronomy, 41(3):393– 414, August 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jong:2019:CPC

[JKKC19] Yong-Ki Jong, Jik-Su Kim, Chol-Jun Kim, and Chol-Min Choe. Comet 27 P/ Crommelin observed in Ancient Korea. Journal for the History of Astronomy, 50(4): 447–455, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619877965.

Jardine:2014:JBH

[JLGL14] Nicholas Jardine, Elisabeth Leedham-Green, and Christopher Lewis. Johann Baptist Hebenstreit’s idyll on the Temple of Urania, the frontispiece image of Kepler’s Rudolphine Tables, Part 1: Context and significance. Journal for the History of Astronomy, 45(1): 1–19, February 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2005:TVUa

[JLS^{+}05] Nicholas Jardine, Dieter Launert, Alain Segonds, Adam Mosley, and Karin Tybjerg. Tycho v. Ursus: The build-up to a trial, Part 1. Journal for the History of Astronomy, 36(1):81–106, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Johnston:2010:WHG

[Joh10]  Stephen Johnston.  Wren,

Hooke and graphical practice. Journal for the History of Astronomy, 41(3):381–392, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Johnston:2019:BRC

[Joh19] Sky Michael Johnston. Book review: Calendar making in confessional Germany: Die deutschen Kalenderschreiber im Zeitalter der Konfessionsbildung 1531–1630: Register und Analyse Robin B. Barnes, (Verlag HKD, Jena, 2018). Pp. 70. EUR 28. ISBN 978-3-941563-19-3. Journal for the History of Astronomy, 50(1): 100–101, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618823987.

Jones:1971:OBK

[Jon71] Kenneth Glyn Jones. The observational basis for Kant’s cosmogony: a critical analysis. Journal for the History of Astronomy, 2(1):29–34, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1976:AAC

[Jon76] Kenneth Glyn Jones. S Andromedae, 1885: an analysis of contemporary reports and a reconstruction. Journal for the History of Astronomy, 7(1): 27–40, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1986:SNH

[Jon86] Kenneth Glyn Jones. Some notes on Hodierna’s nebulae. Journal for the History of Astronomy, 17(3):187–188, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1991:BRB

[Jon91a] Alexander Jones. Book review: a Babylonian astronomical handbook: MUL.APIN: an Astronomical Compendium in Cuneiform. Journal for the History of Astronomy, 22(4): 327–329, November 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1991:HCS

[Jon91b] Alexander Jones. Hipparchus’s computations of solar longitudes. Journal for the History of Astronomy, 22(2):101– 125, May 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1997:BRM

[Jon97] Alexander Jones. Book review: More papyrus horoscopes: Oroscopi Greci: Documentazione Papirologica. Journal for the History of Astronomy, 28(1):89–90, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1999:BRI

[Jon99] Alexander Jones. Book review: On Isagogical Questions: Prolegomena Mathematica: From Apollonius of Perga to the Late Neoplatonists. Journal for the History of Astronomy, 30(3): 315–316, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2002:EHO

[Jon02]  Alexander Jones.  Eratos-

thenes, Hipparchus, and the obliquity of the ecliptic. Journal for the History of Astronomy, 33(1):15–19, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2004:RAD

[Jon04] Alexander Jones. A route to the ancient discovery of non-uniform planetary motion. Journal for the History of Astronomy, 35(4):375–386, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2006:DP

[Jon06a] Alexander Jones. David Pingree (1933–2005). Journal for the History of Astronomy, 37 (2):229–231, May 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2006:ERA

[Jon06b] Alexander Jones. Essay review: Astral sciences in Mesopotamia: The Heavenly Writing: Divination, Horoscopy, and Astronomy in Mesopotamian Culture. Journal for the History of Astronomy, 37(3):353–356, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2010:BRI

[Jon10] Alexander Jones. Book review: Indian mathematics and astronomy: Mathematics in India. Journal for the History of Astronomy, 41(3):416–417, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2012:TAO

[Jon12] Alexander Jones. Theon of Alexandria’s observation of the solar eclipse of A.D. 364 June

16. Journal for the History of Astronomy, 43(1):117–118, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:2015:MJO

[Jon15] Alexander Jones. A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-25863-X (hardcover), 3-319-25865-6 (e-book). xi + 342 pp. LCCN Q125 .M38

                                        2016. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1608/2015959787-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1608/2015959787-
                                        t.html.

Jones:2016:USS

[Jon16] Alexander Jones. Unruly Sun: Solar tables and calculations in the papyrus P. Fouad 267 A.

Journal for the History of Astronomy, 47(1):76–99, February 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http:/

/jha.sagepub.com/content/
47/1/76.full.pdf+html.

Jones:2017:PHT

[Jon17] Alexander Jones. Ptolemy’s handy tables. Journal for the History of Astronomy, 48(2):238–241, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617706254.

Jones:2021:TGR

[Jon21] Alexander Jones. Three Gallo– Roman bronze disks with astral inscriptions. Journal for the History of Astronomy, 52(4):381–396, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211052655.

Jordan:1994:BRO

[Jor94] Detlev Jordan. Book review:

Old Norse cosmography: Altnordische Kosmographie. Studien und Quellen zu Weltbild und Welt–Beschreibung in Norwegen und Island vom 12. bis zum 14. Jahrhundert. Journal for the History of Astronomy, 25(2):144–145, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jardine:2010:IFF

[JR10] Nicholas Jardine and Renée Raphael. Introduction: Forms and functions of early modern celestial imagery. Journal for the History of Astronomy, 41 (3):283–286, August 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Jones:1992:BRB

[JT92] Alexander Jones and G. J. Toomer. Book review: Byzantine astronomy of the Fourteenth Century: Théodore Méliténiote: Tribiblos Astronomique, Livre I. Journal for the History of Astronomy, 23(3):218–222, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Juste:2020:BRA

[Jus20] David Juste. Book review: Abraham Ibn Ezra in Latin: Abraham Ibn Ezra Latinus on Nativities: A Parallel Latin-English Critical Edition of Liber Nativitatum and Liber Abraham Iudei de Nativitatibus. Abraham Ibn Ezra’s Astrological Writings. Shlomo Sela (Brill, Leiden, 2019). Pp. xiv + 564. EUR 226. ISBN 978-90-04-39234-2. Journal for the History of Astronomy, 51 (1):126–128, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619900818.

Juste:2023:ATM

[Jus23] David Juste. The Alfonsine Tables mentioned in 1304. Journal for the History of Astronomy, 54(2):213–219, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231170402.

Juznic:2006:CL

[Juz06] Stanislav Juznic. Copernicus in Ljubljana. Journal for the History of Astronomy, 37(2):231– 232, May 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kaiser:2007:BRT

[Kai07]  David Kaiser.  Book re-

view: Twentieth-Century astrophysics, The Cosmic Century: A History of Astrophysics and Cosmology. Journal for the History of Astronomy, 38 (3):389–390, August 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kak:1995:BRI

[Kak95] Subhash C. Kak. Book review: Indian astronomy: Indian Mathematics and Astronomy: Some Landmarks. Journal for the History of Astronomy, 26(2):172–173, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Karnfelt:2024:ETD

[Kär24] Johan Kärnfelt. ‘Excellentissimo tubo Dollondiana’: The Stockholm Observatory’s 10foot Dollond achromatic refractor. Journal for the History of Astronomy, 55(1): ??, February 1, 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231225455.

Kazarian:1973:BRA

[Kaz73] K. A. Kazarian. Book review: Armenian chronology: History of Chronology. Journal for the History of Astronomy, 4 (2):137–138, June 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kenat:1990:QPS

[KD90] Ralph Kenat and David H. De-Vorkin. Quantum physics and the stars (III): Henry Norris Russell and the search for a rational theory of stellar spectra. Journal for the History of Astronomy, 21(2):157–186, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1998:AMT

[KD98] Richard L. Kremer and Jerzy Dobrzycki. Alfonsine meridians: Tradition versus experience in astronomical practice c. 1500. Journal for the History of Astronomy, 29(2):187– 199, May 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2022:NMS

[KE22] Richard L. Kremer and James Evans. Noel M. Swerdlow, 1941–2021. Journal for the History of Astronomy, 53(3): 364–368, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221116433.

Kremer:2023:OG

[KE23] Richard L. Kremer and James Evans. Owen Gingerich, 1930– 2023. Journal for the History of Astronomy, 54(3):353– 359, August 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231195409.

Keenan:1991:ENO

[Kee91] Philip C. Keenan. The earliest national observatories in Latin America. Journal for the History of Astronomy, 22 (1):21–30, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kelly:1974:BRB

[Kel74] John T. Kelly. Book review:

Benjamin Banneker: The Life of Benjamin Banneker. Journal for the History of Astronomy, 5(2):140–141, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kelley:1978:BRI

[Kel78] David H. Kelley. Book review: Interpreting Maya Lintels, Verhandlungen der Naturforschenden Gesellschaft Basel. Journal for the History of Astronomy, 9(2):145–146, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kelley:1983:BRP

[Kel83] David H. Kelley. Book review: The Paris Codex: Decoding an Astronomical Ephemeris, Transactions of the American Philosophical Society. Journal for the History of Astronomy, 14(5):S70–S72, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kelley:1999:BRC

[Kel99] David H. Kelley. Book review: Calendrical systems explored: Calendrical Calculations, Mapping Time: The Calendar and its History. Journal for the History of Astronomy, 30(4): 407–409, November 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Keller:2010:BRA

[Kel10]  Agathe Keller.  Book re-

view: Astronomy of the Kerala school: Rationales in Mathematical Astronomy of Jyes.thadeva: a Critical Translation of the Original Malayalam Text. Journal for the History of Astronomy, 41(2):273– 274, May 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1973:ERA

[Ken73] E. S. Kennedy. Essay review: Alpetragius’s astronomy: Al-Bit.ruji: On the Principles of Astronomy. Journal for the History of Astronomy, 4 (2):134–136, June 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1978:BRB

[Ken78] E. S. Kennedy. Book review:

Biography of Al-Biruni: Abur-Raihan al-Biruni, 973–1048. Journal for the History of Astronomy, 9(1):65–67, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1982:BRI

[Ken82] E. S. Kennedy. Book review: The Istanbul Symposium: On the Observatories in Islam. Journal for the History of Astronomy, 13(1): 64–66, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1985:BRN

[Ken85] E. S. Kennedy. Book review: Neugebauer papers reprinted: Astronomy and History: Selected Essays. Journal for the History of Astronomy, 16(1): 55–56, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1987:BRE

[Ken87] E. S. Kennedy. Book review: Essays on early astronomy: Theory and Observation in Ancient and Medieval Astronomy. Journal for the History of Astronomy, 18(2):133– 134, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1994:ERT

[Ken94] E. S. Kennedy. Essay review: T .usi’s cosmology: Nas.ir al-Din al-T .usi’s Memoir on Astronomy. Journal for the History of Astronomy, 25(4):321–324, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kershaw:2019:TCL

[Ker19] Michael Kershaw. Twentieth-Century longitude: When Greenwich moved. Journal for the History of Astronomy, 50(2):221–248, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619848180.

Kessler:2010:BRP

[Kes10] Elizabeth A. Kessler. Book review: Photographs from space: Apollo through the Eyes of the Astronauts. Journal for the History of Astronomy, 41(4): 513–514, November 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kessler:2015:BRS

[Kes15] Elizabeth A. Kessler. Book review: Sketching the nebulae. Observing by Hand: Sketching the Nebulae in the Nineteenth Century. Omar W. Nasim (University of Chicago Press, Chicago, 2013). Pp. 304. $45. ISBN 978-0-226-08440-4. Journal for the History of Astronomy, 46(2):237–239, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kennedy:1970:SET

[KF70] E. S. Kennedy and Nazim Faris. The solar eclipse technique of Yah.yā B. Abī Mans.ūr. Journal for the History of Astronomy, 1(1):20–38, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1982:SAO

[KG82] David A. King and Owen Gingerich. Some astronomical observations from Thirteenth-Century Egypt. Journal for the History of Astronomy, 13 (2):121–128, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King-Hele:1973:BRB

[KH73] D. G. King-Hele. Book review: Bode’s law: The Titius–Bode Law of Planetary Distances: Its History and Theory. Journal for the History of Astronomy, 4(2):141–142, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2012:I

[KH12] Richard L. Kremer and Matthieu Husson. Introduction. Journal for the History of Astronomy, 43(3):267, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kreamer:2012:ACS

[KHM^{+}12] Christine Mullen Kreamer, Erin Haney, Katharine Monsted, Karel Nel, Randall Bird, et al. African cosmos: stellar arts. National Museum of African Art, Smithsonian Institution, Washington, DC, USA, 2012. ISBN 1-58093-343-2 (hardcover). 368 pp. LCCN N7380.5 .K74 2012.

Krisciunas:2010:BRO

[KI10] Kevin Krisciunas and Yulia Ivanova. Book review:

Otto Struve: Otto Ludwigovich Struve, 1897–1963. Journal for the History of Astronomy, 41(1):138–140, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kidwell:1987:BRJ

[Kid87] Peggy Kidwell. Book review:

James E. Keeler: James E. Keeler: Pioneer American Astrophysicist and the Early Development of American Astrophysics. Journal for the History of Astronomy, 18(2):141– 143, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kidwell:1990:HAS

[Kid90] Peggy Kidwell. Harvard astronomers in the second world war. Journal for the History of Astronomy, 21(1):105–106, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1973:AKA

[Kin73] David A. King. Al-Khalīlī’s auxiliary tables for solving problems of spherical astronomy. Journal for the History of Astronomy, 4(2):99–110, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1978:ERI

[Kin78] David A. King. Essay review:

Islamic mathematics and astronomy, Islamic Science: an Illustrated Study. Journal for the History of Astronomy, 9(3): 212–219, October 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1979:BRI

[Kin79] David A. King. Book review: An Islamic Astronomical Instrument, the Da’ire-yi Mu’addel of Seydi ‘Ali Re’is. Journal for the History of Astronomy, 10(1):51–53, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1981:ERE

[Kin81] David A. King. Essay review: Early Islamic astronomy: Geschichte des arabischen Schrifttums, Band VI: Astronomie bis ca. 430H. Journal for the History of Astronomy, 12(1):55–59, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1983:BRB

[Kin83] David A. King. Book review: Bio-bibliographies of Muslim astronomers: Die Mathematiker und Astronomen der Araber und ihre Werke. Journal for the History of Astronomy, 14(1):62–63, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1987:ERN

[Kin87] David A. King. Essay review:

Ninth-Century Islamic astronomy: The Book of the Reasons behind Astronomical Tables (Kitāb fī^{c} ilal al-zījat). Journal for the History of Astronomy, 18(4):284–286, November 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1988:IYL

[Kin88] David A. King. Ibn Yūnus on lunar crescent visibility. Journal for the History of Astronomy, 19(3):155–168, August 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1995:BRE

[Kin95a] David A. King. Book review: an Eleventh-Century summary of the Almagest: Ibn al-Haytham’s. Journal for the History of Astronomy, 26(1): 84–85, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:1995:OMI

[Kin95b] David A. King. The orientation of medieval Islamic religious architecture and cities. Journal for the History of Astronomy, 26(3):253–274, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:2002:FB

[Kin02a] David A. King. Frans Bruin (1922–2001). Journal for the History of Astronomy, 33(2): 214–216, May 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:2002:VAT

[Kin02b] David A. King. A Vetustissimus Arabic treatise on the Quadrans Vetus. Journal for the History of Astronomy, 33 (3):237–255, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:2007:HCK

[Kin07] David A. King. Henry C. King (1915–2005). Journal for the History of Astronomy, 38(4): 526–527, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

King:2010:ESK

[Kin10]  David A. King.  Edward

Stewart Kennedy (1912–2009). Journal for the History of Astronomy, 41(1):117–119, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krolikowska:2021:JWJ

[KKSS21] Małgorzata Królikowska, Małgorzata Królikowska, Andrzej M. Sołtan, and Andrzej M. Sołtan. Jan Walery Jedrzejewicz (1835=- 1887) and his observations of comets. Journal for the History of Astronomy, 52(2): 168–188, May 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211012166.

Kleinert:2014:BRT

[Kle14] Andreas Kleinert. Book review: Tobias Mayer as cartographer and astronomer: Tobias Mayers Beiträge zur Wissenschaft des 18. Jahrhunderts im Lichte neuerer Untersuchungen. Edited by Erhard Anthes and Armin Hüttermann (Acta historica astronomiae, xlviii; Akademische Verlagsanstalt, Leipzig, 2013). Pp. 331. EUR 29 (paperback). ISBN 978-3-944913-39-1. Journal for the History of Astronomy, 45(4):492–494, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Katgert-Merkelijn:1991:KEL

[KM91] J. K. Katgert-Merkelijn. The Kenya expeditions of Leiden Observatory. Journal for the History of Astronomy, 22(4): 267–296, November 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kolachana:2018:CDT

[KMD^{+}18] Aditya Kolachana, Clemency Montelle, Jambugahapitiye Dhammaloka, Keshav Melnad, K. Mahesh, Pravesh Vyas, K. Ramasubramanian, Ms Sriram, and Venketeswara Pai. The Candrārkī of Dinakara: a text related to solar and lunar tables. Journal for the History of Astronomy, 49(3):306– 344, August 2018. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knight:1970:BRW

[Kni70] David M. Knight. Book review: Wright’s second thoughts: Second or Singular Thoughts upon the Theory of the Universe. Journal for the History of Astronomy, 1(2):157–158, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knorr:1990:ERN

[Kno90a] Wilbur R. Knorr. Essay review: Notes on Rosen’s Copernicus, Nicolas Copernicus: Minor Works (Nicolas Copernicus: Complete Works, Vol. III). Journal for the History of Astronomy, 21(2):203–211, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knorr:1990:PEP

[Kno90b] Wilbur R. Knorr. Plato and Eudoxus on the planetary motions. Journal for the History of Astronomy, 21(4):313–329, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knorr:1991:ALP

[Kno91] Wilbur R. Knorr. Another look at Ptolemy’s Ivy Leaf. Journal for the History of Astronomy, 22(2):180–183, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knorr:1997:SQD

[Kno97] Wilbur R. Knorr. Sacrobosco’s Quadrans: date and sources. Journal for the History of Astronomy, 28(3):187–222, August 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knowlton:2003:SIM

[Kno03] Timothy Knowlton. Seasonal implications of Maya eclipse and rain iconography in the Dresden Codex. Journal for the History of Astronomy, 34 (3):291–303, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Knudsen:2014:SJA

[Knu14] Toke Lindegaard Knudsen. The Siddhāntasundara of Jñãnarãja: an English translation with commentary. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1442-2 (hardcover), 1-4214-1443-0 (electronic). xvii + 351 pp. LCCN QB18 J6313

Kokott:1981:C

[Kok81] Wolfgang Kokott. The comet of 1533. Journal for the History of Astronomy, 12(2):95– 112, June 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kokott:1998:SPU

[Kok98]  Wolfgang Kokott.  Syzy-

gies as pivots: an unusual mid-Fifteenth-Century working Ephemeris. Journal for the History of Astronomy, 29 (2):129–135, May 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kollerstrom:1992:HWE

[Kol92] N. Kollerstrom. The hollow world of Edmond Halley. Journal for the History of Astronomy, 23(3):185–192, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kolev:2001:BHM

[Kol01a] Rumen Kolev. Babylonian horoscope MLC 1870. Journal for the History of Astronomy, 32(2):154, May 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kolev:2001:WHR

[Kol01b] Rumen Kolev. Witnessing the heliacal rise of Sirius and Procyon. Journal for the History of Astronomy, 32(2):152–153, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kolev:2007:SOH

[Kol07] Rumen Kolev. Six observations of the heliacal rise of Sirius. Journal for the History of Astronomy, 38(2):227–228, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kolbe:2008:TST

[Kol08]  Geoffrey Kolbe.  A test

of the “Simultaneous Transit Method”. Journal for the History of Astronomy, 39(4):515– 517, November 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kilburn:2003:FSA

[KPG03] Kevin J. Kilburn, Jay M. Pasachoff, and Owen Gingerich. The forgotten star atlas: John Bevis’s Uranographia Britannica. Journal for the History of Astronomy, 34(2):125–144, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:1995:CBW

[Kra95] Helge Kragh. Cosmology between the wars: The Nernst– Macmillan alternative. Journal for the History of Astronomy, 26(2):93–115, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:1999:BRE

[Kra99] Helge Kragh. Book review: Einstein and the Einstein tower: The Einstein Tower: an Intertexture of Dynamic Construction, Relativity Theory, and Astronomy. Journal for the History of Astronomy, 30 (1):87–88, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kraai:2007:BRR

[Kra07a]  Jesse Kraai.  Book re-

view: Rheticus as Copernican, the first Copernican: Georg Joachim Rheticus and the Rise of the Copernican Revolution. Journal for the History of Astronomy, 38(4):516–517, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2007:BRL

[Kra07b] Helge Kragh. Book review: Lemaître in context: The Day without Yesterday: Lemaître, Einstein, and the Birth of Modern Cosmology. Journal for the History of Astronomy, 38(1): 129–130, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2007:CRA

[Kra07c] Helge Kragh. Cosmic radioactivity and the age of the universe, 1900–1930. Journal for the History of Astronomy, 38 (4):393–412, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2009:SME

[Kra09] Helge Kragh. The second Moon of the Earth. Journal for the History of Astronomy, 40(1): 1–10, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2010:BRR

[Kra10] Helge Kragh. Book review:

Recollections of big bang cosmology: Finding the Big Bang. Journal for the History of Astronomy, 41(1):137–138, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2011:BRH

[Kra11] Helge Kragh. Book review:

The Herschel siblings: Discoverers of the Universe: William and Caroline Herschel. Journal for the History of Astronomy, 42(4):538–540, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2012:KPP

[Kra12] Helge Kragh. Karl Popper on physical cosmologies. Journal for the History of Astronomy, 43(3):347–350, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2013:BRI

[Kra13] Helge Kragh. Book review: The invisible universe, Heart of Darkness: Unraveling the Mysteries of the Invisible Universe. Journal for the History of Astronomy, 44(4):490–491, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2015:KD

[Kra15] Helge Kragh. Kapteyn in detail. Journal for the History of Astronomy, 46(4):503–504, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kragh:2017:NPS

[Kra17]  Helge Kragh.  The Nobel

Prize system and the astronomical sciences. Journal for the History of Astronomy, 48(3):257–280, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617721574.

Kragh:2020:BSC

[Kra20] Helge Kragh. Building the standard cosmological model. Journal for the History of Astronomy, 51(4):482–483, November 1, 2020. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620954561.

Kragh:2022:BRC

[Kra22] Helge Kragh. Book review:

Celebrating the centenary of the IAU: Under One Sky: The IAU Centenary Symposium, Christiaan Sterken, John Hearnshaw & David Valls-Gabaud (Cambridge University Press, Cambridge, 2019). Pp. xii + 542. $120.84. ISBN 978-1-108-47159-6. Journal for the History of Astronomy, 53(2):235–236, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221081971.

Kragh:2024:CIH

[Kra24a] Helge Kragh. A comprehensive institutional history. Journal for the History of Astronomy, 55(4):507–508, November 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241259476.

Kragh:2024:NGH

[Kra24b] Helge Kragh. Nebulae or galaxies? The history of a change in astronomical terminology. Journal for the History of Astronomy, 55(3): 287–313, August 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241238732.

Kremer:1980:BRA

[Kre80a] Richard Kremer. Book review: Astronomy in Nuremberg, 600 Jahre Astronomie in Nürnberg. Journal for the History of Astronomy, 11(3):208–210, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1980:BWA

[Kre80b] Richard L. Kremer. Bernard Walther’s astronomical observations. Journal for the History of Astronomy, 11(3):174–191, October 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1981:BRH

[Kre81a] Richard L. Kremer. Book review: Hegel’s dissertation: Les orbites des planètes. Journal for the History of Astronomy, 12(1):69–71, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1981:UBW

[Kre81b] Richard L. Kremer. The use of Bernard Walther’s astronomical observations: Theory and observation in early modern astronomy. Journal for the History of Astronomy, 12(2):124– 132, June 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1982:BRG

[Kre82] Richard Kremer. Book review: Geophysics and the philosophers: Disziplingeschichte als wissenschaftliche Selbstsreflexion der historischen Wissenschaftsforschung. Eine Darstellung unter Heranziehung von Fallstudien der Wissenschaftsgeschichte der Geophysik. Journal for the History of Astronomy, 13(3):215–217, October 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1986:BRB

[Kre86] Richard L. Kremer. Book review: Bessel family correspondence: Ich habe Euch lieb aber der Himmel ist mir näher. Eine Autobiographie in Briefen. Journal for the History of Astronomy, 17(3):210– 211, August 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1989:BRS

[Kre89] Richard Kremer. Book review: Sixteenth-Century astrology, astrologi hallucinati: Stars and the End of the World in Luther’s Time. Journal for the History of Astronomy, 20 (2):140–141, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:1993:BRS

[Kre93] Richard L. Kremer. Book review: Saving Ptolemy: On the Distances between Sun, Moon & Earth: According to Ptolemy, Copernicus & Reinhold. Journal for the History of Astronomy, 24(1–2): 148–150, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/24/1-
                                        2/148.full.pdf+html.

Kremer:1999:BRM

[Kre99] Richard L. Kremer. Book review: Mathematics in Renaissance Ingolstadt: Mathematik und Astronomie an der Universität Ingolstadt im 15. und 16. Jahrhundert. Journal for the History of Astronomy, 30 (2):173–175, May 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2001:VCA

[Kre01] Richard L. Kremer. The Vancouver Conference on “Astronomy and its Histories”. Journal for the History of Astronomy, 32(3):187, August 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2002:BRA

[Kre02] Richard L. Kremer. Book review: Abraham Zacut of Salamanca, Astronomy in the Iberian Peninsula: Abraham Zacut and the Transition from Manuscript to Print. Journal for the History of Astronomy, 33(3):291–292, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2005:BRT

[Kre05] Richard L. Kremer. Book review: Tycho assessed: Tycho Brahe and Prague: Crossroads of European Science. Journal for the History of Astronomy, 36(4):456–458, November 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2007:AAT

[Kre07a] Richard L. Kremer. “Abbreviating” the Alfonsine Tables in Cracow: The Tabulae Aureae of Petrus Gaszowiec (1448). Journal for the History of Astronomy, 38(3):283–304, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2007:ERA

[Kre07b] Richard L. Kremer. Essay review: Astronomy in Kassel: Die astronomische Forschungen in Kassel unter Wilhelm IV, mit einer Teiledition der deutschen Übersetzung des Hauptwerkes von Copernicus um 1586, Christoph Rothmann’s Handbuch der Astronomie von 1589. Journal for the History of Astronomy, 38(1):125–129, February 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2008:BRS

[Kre08] Richard L. Kremer. Book review: a sourcebook on motion and its causes; Mechanics from Aristotle to Einstein. Journal for the History of Astronomy, 39(1):127, February 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2009:ERU

[Kre09] Richard L. Kremer. Essay review: Updating the DSB, New Dictionary of Scientific Biography. Journal for the History of Astronomy, 40(4):467–471, November 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2010:CAA

[Kre10] Richard L. Kremer. Calculating with Andreas Aurifaber: a new source for Copernican astronomy in 1540. Journal for the History of Astronomy, 41(4):483–502, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2011:EPI

[Kre11]  Richard L. Kremer.  Ex-

perimenting with paper instruments in Fifteenth- and Sixteenth-Century astronomy: Computing syzygies with isotemporal lines and salt dishes. Journal for the History of Astronomy, 42(2):223–258, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2012:MSW

[Kre12] Richard L. Kremer. Marcus Schinnagel’s winged polyptych of 1489: Astronomical computation in a liturgical format. Journal for the History of Astronomy, 43(3):321–345, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2013:BRP

[Kre13] Richard L. Kremer. Book review: Prophetic texts: Printing and Prophecy: Prognostication and Media Change, 1450– 1550. Journal for the History of Astronomy, 44(1):118–119, February 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kremer:2018:A

[Kre18] Richard L. Kremer. 29,500 astronomers. Journal for the History of Astronomy, 49(4): 504–505, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618791879.

Kremer:2019:BRC

[Kre19] Richard L. Kremer. Book review: Completing Kepler’s Gesammelte Werke. Journal for the History of Astronomy, 50(4):480–481, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/1939790919871727.

Kremer:2020:BRF

[Kre20] Richard L. Kremer. Book review: Four centuries of wall calendars. Journal for the History of Astronomy, 51 (2):245–246, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620924155.

Krisciunas:1988:BRL

[Kri88] Kevin Krisciunas. Book review: Lomonosov’s contribution to astronomy: M. V. Lomonosov: Astronomer and Astrophysicist. Journal for the History of Astronomy, 19(4): 282–284, November 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krisciunas:1993:MCA

[Kri93] Kevin Krisciunas. A more complete analysis of the errors in Ulugh Beg’s star catalogue. Journal for the History of Astronomy, 24(4):269–280, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krisciunas:1994:BRO

[Kri94] Kevin Krisciunas. Book review: Observatories and telescopes: Sternwarten in Bildern: Architektur und Geschichte der Sternwarten von den Anfängen bis ca. 1950. Journal for the History of Astronomy, 25(1): 63–65, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krisciunas:1997:BRT

[Kri97]  Kevin Krisciunas.  Book

review: The travails of Pulkovo: Pulkovo/St. Petersburg: Spuren der Sterne und der Zeiten; Geschichte der Russischen Hauptsternwarte. Journal for the History of Astronomy, 28(2):176–177, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krupp:1983:BRA

[Kru83]  E. C. Krupp.  Book re-

view: Archaeoastronomy in the old world, Archaeoastronomy in the New World. Journal for the History of Astronomy, 14(5):S67–S70, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krupp:1997:BRM

[Kru97] E. C. Krupp. Book review:

Myths and their interpretation: Star Myths of the Vikings: a New Concept of Norse Mythology. Journal for the History of Astronomy, 28(4):353–354, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krupp:2008:BRP

[Kru08]  Ed C. Krupp.  Book re-

view: Planetaria in the U.S.; Theaters of Time and Space: American Planetaria, 1930– 1970. Journal for the History of Astronomy, 39(1):135–136, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Krupp:2017:SAM

[Kru17]  Ec Krupp.  Surveying ar-

chaeoastronomy in Mesoamerica. Essay review: Archaeoastronomy and the Maya. Gerardo Aldana y Villalobos and Edwin L Barnhart (Oxbow Books, Oxford, 2014). Pp. vii + 165. $50. ISBN 978-1-78297-643-1 (paper). Journal for the History of Astronomy, 48(3):367–369, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617713524.

Krupp:2022:TPC

[Kru22] E. C. Krupp. Time in pre-Columbian America. Journal for the History of Astronomy, 53(1):124–125, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211048989.

Kulikovsky:1975:BRW

[Kul75] P. G. Kulikovsky. Book review: World systems in Armenia: The Geocentric and Heliocentric Systems in Armenia. Journal for the History of Astronomy, 6(1):67, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kulikovskij:1980:BRT

[Kul80] P. G. Kulikovskij. Book review: Two great astronomers, The Star Atlas. Journal for the History of Astronomy, 11(3): 207–208, October 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kulikovskij:1981:BRS

[Kul81] P. G. Kulikovskij. Book review:

Science at Samarkand: From the History of Science in the Time of Ulugh Beg. Journal for the History of Astronomy, 12(3):211–212, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kulikovsky:1987:BRP

[Kul87] P. G. Kulikovsky. Book review:

Peter the Great’s astronomers: Petersburgskaya Astronomicheskaya Shkola XVIII B. Journal for the History of Astronomy, 18(1):71–72, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1980:TST

[Kun80] Paul Kunitzsch. Two star tables from Muslim Spain. Journal for the History of Astronomy, 11(3):192–201, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1985:BRA

[Kun85] Paul Kunitzsch. Book review: Arabic-Islamic exact sciences: Studies in the Islamic Exact Sciences. Journal for the History of Astronomy, 16 (1):49–51, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1986:JLH

[Kun86a] Paul Kunitzsch. John of London and his unknown Arabic source. Journal for the History of Astronomy, 17(1): 51–57, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1986:SCC

[Kun86b] Paul Kunitzsch. The star catalogue commonly appended to the Alfonsine Tables. Journal for the History of Astronomy, 17(2):89–98, May 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1987:PAA

[Kun87] Paul Kunitzsch. Peter Apian and ‘Azophi’: Arabic constellations in Renaissance astronomy. Journal for the History of Astronomy, 18(2):117–124, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1988:BRI

[Kun88] Paul Kunitzsch. Book review:

‘Islamicate’ globes: Islamicate Celestial Globes: Their History, Construction, and Use. Journal for the History of Astronomy, 19(4):278–280, November 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1990:BRN

[Kun90] Paul Kunitzsch. Book review:

Naming the heavens, Dictionary of Astronomical Names. Journal for the History of Astronomy, 21(3):296–297, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kunitzsch:1991:BRS

[Kun91] Paul Kunitzsch. Book review:

Star names: Catalogue des étoiles nommées /Catalogue of Named Stars. Journal for the History of Astronomy, 22(4): 320–322, November 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kurtik:2019:UEH

[Kur19] Gennady E. Kurtik. ^{mul} uz_{3}, muldGula, and the early history of Mesopotamian constellations. Journal for the History of Astronomy, 50(3): 339–359, August 1, 2019.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619853676.

Kurtik:2021:OZC

[Kur21]  Gennady E. Kurtik.  On

the origin of the 12 zodiac constellation system in ancient Mesopotamia. Journal for the History of Astronomy, 52(1):53–66, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620980544.

Kusukawa:2010:BRC

[Kus10] Sachiko Kusukawa. Book review: Copernicanism in Germany: Nouveau Ciel, Nouvelle Terre: La Révolution Copernicienne dans l’Allemagne de la Réforme (1530–1630). Journal for the History of Astronomy, 41(4):519–521, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kusukawa:2011:ANA

[Kus11] Sachiko Kusukawa. Andreas Nolthius’s Almanach for 1575. Journal for the History of Astronomy, 42(1):91–110, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kutrovatz:2023:KSP

[Kut23] Gábor Kutrovátz. Kepler’s struggle with the problem of force obstruction. Journal for the History of Astronomy, 54(3):316–332, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231184672.

Kwan:2012:BRT

[Kwa12] Alistair Kwan. Book review:

On telescopes: From Earth-Bound to Satellite: Telescopes, Skills and Networks. Journal for the History of Astronomy, 43(4):508–510, November 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Kollerstrom:1995:FLD

[KY95] Nick Kollerstrom and Bernard D.

Yallop. Flamsteed’s lunar data, 1692–95, sent to Newton. Journal for the History of Astronomy, 26(3):237–246, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lada:2010:BRI

[Lad10] Charles J. Lada. Book review:

Imaging the heavens: Shrouds of the Night: Masks of the Milky Way and Our Awesome New View of Galaxies. Journal for the History of Astronomy, 41(2):269–271, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lankford:1980:NJJ

[Lan80] John Lankford. A note on T. J. J. See’s observations of craters on Mercury. Journal for the History of Astronomy, 11 (2):129–132, June 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lang:1982:BRD

[Lan82] Kenneth R. Lang. Book review: The discovery of Pluto: Out of the Darkness: The Planet Pluto. Journal for the History of Astronomy, 13(2):136– 138, June 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lankford:1983:PLF

[Lan83] John Lankford. Photography and the long-focus visual refractor: Three American case studies, 1885–1914. Journal for the History of Astronomy, 14(2):77–91, June 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lang:1984:BRM

[Lan84] Kenneth R. Lang. Book review: Mercury’s anomalous motion: Mercury’s Perihelion from Le Verrier to Einstein. Journal for the History of Astronomy, 15(3):216–218, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lankford:1987:BRI

[Lan87] John Lankford. Book review:

Illustrating the Paris Observatory: L’Observatoire de Paris: Son Histoire, 1667–1963. Journal for the History of Astronomy, 18(2):135, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lankford:1988:BRM

[Lan88]  John Lankford.  Book re-

view: The McDonald Observatory: Big and Bright: a History of the McDonald Observatory. Journal for the History of Astronomy, 19(1):72–74, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Landes:1989:ERC

[Lan89] David S. Landes. Essay review:

Chinese astronomical clocks, heavenly clockwork: The Great Astronomical Clocks of Medieval China. Journal for the History of Astronomy, 20(1): 63–65, February 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Langermann:1998:PHT

[Lan98] Y. Tzvi Langermann. Peurbach in the Hebrew tradition. Journal for the History of Astronomy, 29(2):137–150, May 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Laszlo:2019:BRP

[Lás19] Levente László. Book review: Ptolemy from Greek to Latin. Journal for the History of Astronomy, 50 (2):250–252, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619840510.

Lattis:1993:BRT

[Lat93] James M. Lattis. Book review: Thoren on Tycho Brahe: The Lord of Uraniborg: a Biography of Tycho Brahe. Journal for the History of Astronomy, 24 (3):225–226, August 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lattis:2000:BRC

[Lat00] James M. Lattis. Book review: Commenting on the commentators: Method and Order in Renaissance Philosophy of Nature: The Aristotle Commentary Tradition. Journal for the History of Astronomy, 31 (2):163–164, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lattis:2012:BRC

[Lat12] James Lattis. Book review: a contemporary of Galileo: Ilario Altobelli: Scienziato, Teologo, Corrispondente di Galileo Galilei. Journal for the History of Astronomy, 43(4):506–508, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Latusseck:2014:MWJ

[Lat14] Arndt Latusseck. The Milky Way in Johann Bayer’s Uranometria, 1603. Journal for the History of Astronomy, 45 (2):161–181, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Latusseck:2017:MLF

[Lat17] Arndt Latußeck. More light on the Flamsteed numbers. Journal for the History of Astronomy, 48(3):346–353, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617721576.

Lattis:2019:BRT

[Lat19] James M. Lattis. Book review: a textbook about how science works. Journal for the History of Astronomy, 50(4): 489–491, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619880408.

Lattis:2020:CEE

[Lat20] Anthony Lattis. Constructing the electric eye: Situating the Smithsonian National Air and Space Museum’s Wisconsin collection of photoelectric detectors in historical context. Journal for the History of Astronomy, 51(4):423–460, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620942409.

Lattis:2023:TYH

[Lat23] James M. Lattis. Thirty years of the HST. Journal for the History of Astronomy, 54(1): 107–108, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221142543.

Launay:2007:GPE

[Lau07] Françoise Launay. The great Paris exhibition telescope of 1900. Journal for the History of Astronomy, 38(4):459–475, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Launius:2014:ASD

[Lau14] Roger D. Launius. Astronomy by satellite: Dreams of Other Worlds: The Amazing Story of Unmanned Space Exploration. By Chris Impey and Holly Henry (Princeton University Press, Princeton, 2013). Pp. 450. $35. ISBN 978-0-691-14753-6. Journal for the History of Astronomy, 45(3):377– 378, August 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Launius:2018:IVP

[Lau18] Roger D. Launius. An insider’s view of planetary science. Journal for the History of Astronomy, 49(4): 498–499, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618806142.

Layzer:2002:ASN

[Lay02] David Layzer. Atoms, stars, and nebulae: Remembering Donald H. Menzel. Journal for the History of Astronomy, 33 (2):133–138, May 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lull:2006:FAT

[LB06] José Lull and Juan Antonio Belmonte. A firmament above Thebes: Uncovering the constellations of Ancient Egyptians. Journal for the History of Astronomy, 37(4):373–392, November 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

LeGuetTully:2010:BRL

[LC10] Françoise Le Guet Tully and James Caplan. Book review: Lalande remembered: Un Astronome des Lumières: Jérôme Lalande. Journal for the History of Astronomy, 41(1):144– 145, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lippitsch:2012:MPD

[LD12] Max E. Lippitsch and Sonja Draxler. A medieval planetary diagram in Graz University Library. Journal for the History of Astronomy, 43(2):141–149, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

LeConte:2008:OCI

[Le 08] David Le Conte. Orientations of Channel Islands megalithic tombs. Journal for the History of Astronomy, 39(4):497–506, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [Le 09].

LeConte:2009:OCI

[Le 09] David Le Conte. Orientations of Channel Islands megalithic tombs: Addendum. Journal for the History of Astronomy, 40 (2):217–218, May 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Le 08].

Leitao:2017:AEM

[Lei17] Henrique Leitão. Astronomy in early modern Spain. Journal for the History of Astronomy, 48(2):248–249, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617706234.

Leppik:2014:BRS

[Lep14] Lea Leppik. Book review:

a story of Bernhard Schmidt: Sailing Against the Wind: A Novel. Jaan Kross. Translated from Estonian by Eric Dickens (Northwestern University Press, Evanston, IL, 2012). Pp. xiv + 347. $24.95 (paperback). ISBN 978-0-8101-2652-7. Journal for the History of Astronomy, 45(2):263–264, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lequeux:2018:BC

[Leq18] James Lequeux. A biography of Cassini i. Journal for the History of Astronomy, 49(4):496–497, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618808874.

Lerner:1980:SNT

[Ler80]  Michel-Pierre Lerner.  “Si-

cut Nodus in Tabula”: de la rotation propre du soleil au seizième siècle. (French) [“As a node on the board”: on the proper rotation of the Sun in the Sixteenth Century]. Journal for the History of Astronomy, 11(2):114–129, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2000:CPT

[Ler00] Michel-Pierre Lerner. Copernicus in Paris in 1612: a teaching text edition of De Revolutionibus. Journal for the History of Astronomy, 31(1): 55–67, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2005:OMW

[Ler05] Michel-Pierre Lerner. The origin and meaning of “world system”. Journal for the History of Astronomy, 36(4):407–441, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2006:BRO

[Ler06] Michel-Pierre Lerner. Book review: The origins of the Alphonsine Tables: Histoire Littéraire de la France (Académie des Inscriptions et Belles Lettres). Journal for the History of Astronomy, 37(4): 491–493, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2011:APS

[Ler11a] Michel-Pierre Lerner. Alain-Philippe Segonds (1942–2011). Journal for the History of Astronomy, 42(4):542–543, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2011:BRP

[Ler11b] Michel-Pierre Lerner. Book review: Ptolemy’s geography in the West: La Géographie de Ptolémée en Occident (IVe– XVIe siècle). Journal for the History of Astronomy, 42(2): 273–275, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2017:BRV

[Ler17] Michel-Pierre Lerner. Book review: A Very Short Biography. Journal for the History of Astronomy, 48 (4):487–489, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616689320.

Lerner:2025:CCE

[Ler25] Michel-Pierre Lerner. A Copernican classic in English. Journal for the History of Astronomy, 56(1):107–109, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241230370.

Levitt:2017:AA

[Lev17a] Theresa Levitt. Arago the astronomer. Journal for the History of Astronomy, 48(2): 247–248, May 2017. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617706829.

Levy:2017:ECA

[Lév17b]  Tony Lévy.  Early calen-

dars and astronomy. Journal for the History of Astronomy, 48(2):252–253, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617705928.

Levitt:2025:LP

[Lev25] Theresa Levitt. The last polymath. Journal for the History of Astronomy, 56(1):112– 113, February 2025. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286241269052.

Lewis:2007:TPI

[Lew07] John Lewis. Truth and propaganda in images of the trial of Galileo. Journal for the History of Astronomy, 38(1): 15–29, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Laurel:2015:BRA

[LF15] Maryland Laurel and Jerry L. Fields. Book review: Astronomy on postage stamps. Stamping Through Astronomy. Renato Dicati (Springer, New York, 2013). Pp. xii + 373. $40 (paperback). ISBN 978-88-470-2828-9. Journal for the History of Astronomy, 46(2):235– 236, May 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lerner:2012:ERP

[LG12] Michel-Pierre Lerner and Miguel Granada. Essay review: Planetary order in the long Sixteenth Century, the Copernican question: Prognostication, skepticism, and celestial order. Journal for the History of Astronomy, 43(2):233–243, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Leedham-Green:2014:JBH

[LGJLF14] Elisabeth Leedham-Green, Nicholas Jardine, Christopher Lewis, and Isla Fay. Johann Baptist Hebenstreit’s idyll on the Temple of Urania, the frontispiece image of Kepler’s Rudolphine Tables, Part 2: Annotated translation. Journal for the History of Astronomy, 45(1): 21–34, February 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Latusseck:2003:MJH

[LH03] Arndt Latusseck and Michael Hoskin. The murder of Jacob Herschel. Journal for the History of Astronomy, 34(2):233– 234, May 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Launay:2005:JJR

[LH05] Françoise Launay and Peter D.

Hingley. Jules Janssen’s ‘revolver photographique’ and its British derivative, ‘the Janssen Slide’. Journal for the History of Astronomy, 36(1):57– 79, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Li:2019:NAS

[Li19] Liang Li. New astronomy in service of old astrology: Close planetary conjunctions in pre-modern China. Journal for the History of Astronomy, 50(4):411–427, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619866328.

Liang:2019:BRN

[Lia19] Li Liang. Book review: A narrative history of astronomy in Early Imperial China. Journal for the History of Astronomy, 50(1):97–98, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618808875.

Liebowitz:2002:DMC

[Lie02] Ruth Prelowski Liebowitz. Donald Menzel and the creation of the Sacramento Peak Observatory. Journal for the History of Astronomy, 33(2): 193–211, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Liller:1989:MAE

[Lil89] Wm Liller. The megalithic astronomy of Easter Island: Orientations of AHU and MOAI. Journal for the History of Astronomy, 20(13):S21–S48, February 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Liller:1999:BRM

[Lil99]  William Liller.  Book re-

view: Measuring magnitudes: The Measurement of Starlight: Two Centuries of Astronomical Photometry. Journal for the History of Astronomy, 30 (1):81–82, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lindberg:1979:BRI

[Lin79] David C. Lindberg. Book review: The invention of the telescope, the Invention of the Telescope. Journal for the History of Astronomy, 10(1): 57–58, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lindberg:1993:BRR

[Lin93] David C. Lindberg. Book review: Religion and science: The Book of Nature. Journal for the History of Astronomy, 24(1–2):150–151, May/February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/150.full.pdf+html.

Linton:2004:EEH

[Lin04] C. M. (Christopher M.) Linton. From Eudoxus to Einstein: a history of mathematical astronomy. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-82750-7 (hardcover), 0-511-21109-0, 0-511-21286-0, 0-511-20751-4 (ebook), 0-511-21467-7 (e-book). xii + 516 pp. LCCN QB47 .L56 2004; QB47 LIN. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://catdir.loc.gov/catdir/
enhancements/fy0732/2003067569-
b.html;   http://www.loc.
gov/catdir/description/cam041/
2003067569.html;   http:/
/www.loc.gov/catdir/toc/
cam041/2003067569.html.

Lindberg:2007:BRO

[Lin07] David Lindberg. Book review: Oresme on optics, Nicole Oresme’s De Visione Stellarum (On Seeing the Stars): a Critical Edition of Oresme’s Treatise on Optics and Atmospheric Refraction, with an Introduction, Commentary, and English Translation. Journal for the History of Astronomy, 38 (4):509–510, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lipton:2000:BRR

[Lip00] Joshua Lipton. Book review:

Reading and believing: The Nature of the Book: Print and Knowledge in the Making. Journal for the History of Astronomy, 31(1):76–78, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lippincott:2015:BRA

[Lip15] Kristen Lippincott. Book review: The Aratean corpus in Vat. Grec. 1087. Antiche stele a Bisanzio. Il codice Vaticano Greco 1087. Edited by Fabio Guidetti and Anna Santoni (Edizioni della Normale), Pisa, 2013. Pp. 208 + 31 illustrations. EUR 20. ISBN 978-88-7642-485-4. Journal for the History of Astronomy, 46 (2):249–250, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lippincott:2019:BRI

[Lip19]  Kristen Lippincott.  Book

review: Imaging the Cosmos. Journal for the History of Astronomy, 50(4): 483–485, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/1939790919871737.

Livio:2013:BRS

[Liv13] Mario Livio. Book review:

Slipher’s observations of redshifts of spirals, Origins of the Expanding Universe, 1912– 1932: Proceedings of a Conference Held at Flagstaff, Arizona, USA, 13–15 September 2012. Journal for the History of Astronomy, 44(4):493–494, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Livio:2016:ICT

[Liv16] Mario Livio. Imagined conversations with Twentieth-Century cosmologists. Journal for the History of Astronomy, 47(1):109–110, February 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
109.full.pdf+html.

Lizzini:2017:BRA

[Liz17] O. L. Lizzini. Book review: Astronomy and Metaphysics in al-Farabi’s Thought. Journal for the History of Astronomy, 48(4):483–485, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617730692.

Lee:1987:EIS

[LL87] Georgia Lee and William Liller. Easter Island’s “sun stones”: a re-evaluation. Journal for the History of Astronomy, 18(11): S1–S11, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:1980:BRC

[Llo80] G. E. R. Lloyd. Book review: Colloquium on ancient astronomy, L’Astronomie dans l’Antiquité Classique. Journal for the History of Astronomy, 11(3):204–205, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:1989:BRN

[Llo89a] Steven A. Lloyd. Book review:

Nuremberg sundials, The Ivory Sundials of Nuremberg, 1500– 1700. Journal for the History of Astronomy, 20(3):221–222, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:1989:LVM

[Llo89b]  Steven A. Lloyd.  Lunar

volvelles and moondials in Baroque Germany. Journal for the History of Astronomy, 20 (2):121–127, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:1990:BRS

[Llo90] Steven A. Lloyd. Book review:

Sundials in the Whipple museum, Sundials and Related Instruments. Journal for the History of Astronomy, 21(2):221, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:1994:BRP

[Llo94] Geoffrey Lloyd. Book review: Ptolemy’s astronomy in context: Ptolemy’s Universe: The Natural Philosophical and Ethical Foundations of Ptolemy’s Astronomy. Journal for the History of Astronomy, 25(1): 62–63, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:2009:BRC

[Llo09] Geoffrey Lloyd. Book review: Cosmogony in antiquity, Ancient Greek Cosmogony. Journal for the History of Astronomy, 40(2):220–221, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:2011:BRP

[Llo11] Geoffrey Lloyd. Book review: Plato’s Timaeus: One Book, the Whole Universe: Plato’s Timaeus Today. Journal for the History of Astronomy, 42 (1):117–118, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:2015:BRA

[Llo15] G. E. R. Lloyd. Book review: Aristotle on circular motion: Aristotle’s Empiricism: Experience and Mechanics in the 4th Century BC. Jean De Groot (Parmenides Publishing, Las Vegas, NV, 2014). Pp. xxv + 442. $127 (paper). ISBN 978-1-930972-83-4. Journal for the History of Astronomy, 46(3): 372–373, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lloyd:2018:AWK

[Llo18]  Geoffrey Lloyd.  Ancient

whizz-kiddery. Journal for the History of Astronomy, 49 (1):131–132, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617738387.

Lanuza-Navarro:2009:BRJ

[LN09] Tayra Lanuza-Navarro. Book review: Jesuit science in Portugal, Sphaera Mundi: a Ciência na Aula da Esfera. Manuscritos científicos do Colégio de Santo Antão nas colecções da BNP. Journal for the History of Astronomy, 40 (3):368–369, August 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1983:FCL

[Loc83] Kurt Locher. A further coffinlid with a diagonal star-clock from the Egyptian Middle Kingdom. Journal for the History of Astronomy, 14(2):141– 144, June 1983. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1984:PIL

[Loc84] Kurt Locher. Ptolemy’s ivy leaf. Journal for the History of Astronomy, 15(1):32– 34, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1985:PIA

[Loc85] Kurt Locher. Probable identification of the Ancient Egyptian circumpolar constellations. Journal for the History of Astronomy, 16(9):S152– S153, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1989:FHC

[Loc89] Kurt Locher. A further Hellenistic conical sundial from the Theatre of Dionysos in Athens. Journal for the History of Astronomy, 20(1):60–62, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1990:AEC

[Loc90] Kurt Locher. The Ancient Egyptian constellation group “the lion between the two crocodiles” and the bird. Journal for the History of Astronomy, 21(15):S49–S51, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1992:TFC

[Loc92] Kurt Locher. Two further coffin lids with diagonal star clocks from the Egyptian Middle Kingdom. Journal for the History of Astronomy, 23(3): 201–207, August 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1993:TGR

[Loc93] Kurt Locher. Two Greco– Roman sundials from Alexandria and Dion. Journal for the History of Astronomy, 24(4): 300–302, November 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1995:TFG

[Loc95] Kurt Locher. Three further Greco–Roman conical sundials from Palmyra, Naples, and Abû Mînâ. Journal for the History of Astronomy, 26(2):159– 163, May 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1996:BRA

[Loc96] Kurt Locher. Book review:

Ancient Egyptian astronomy: Ancient Egyptian Science, II: Calendars, Clocks, and Astronomy. Journal for the History of Astronomy, 27(2):173–174, May 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:1999:BRE

[Loc99] Kurt Locher. Book review:

The ecliptic in Ancient Egypt: Astronomische Konzepte und Jenseitsvorstellungen in den Pyramidentexten. Journal for the History of Astronomy, 30 (1):75–76, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Locher:2000:BRE

[Loc00] Kurt Locher. Book review: The Egyptian calendar in context: The Egyptian Calendar: a Work for Eternity. Journal for the History of Astronomy, 31(4):349, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lohne:1975:ERH

[Loh75] J. A. Lohne. Essay review: The Harriot Symposium: Thomas Harriot, Renaissance Scientist. Journal for the History of Astronomy, 6(2):131–134, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lohr:2013:CGE

[Loh13] Alfred Lohr. Der Computus Gerlandi: Edition, Übersetzung und Erläuterungen, volume 61 of Sudhoffs Archiv. Beihefte, ; Wissenschaftsgeschichte. Franz Steiner, Stuttgart, Germany, 2013. ISBN 3-515-10468-2 (paperback). ISSN 0341-0773. 493 pp. LCCN B765.G343 C66 2013.

Lohr:2019:BRJ

[Löh19]  Nadine Löhr.  Book re-

view: Jaghmīnī’s Mulakhkhas.: an Islamic Introduction to Ptolemaic Astronomy, Sally P. Ragep, (Springer, Cham, 2016). Pp. ix + 347. EUR 91. ISBN 978-3-319-31992-6. Journal for the History of Astronomy, 50(1): 105–107, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618817150.

Longair:2006:ERH

[Lon06] Malcolm Longair. Essay review: Hoyle assessed: Conflict in the Cosmos: Fred Hoyle’s Life in Science, Fred Hoyle’s Universe. Journal for the History of Astronomy, 37(1):111– 114, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Longair:2015:CIA

[Lon15] Malcolm Longair. Creating infrared astronomy. Journal for the History of Astronomy, 46(1):103–105, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lovell:1977:EDS

[Lov77] Sir Bernard Lovell. The effects of defence science on the advance of astronomy. Journal for the History of Astronomy, 8(3): 151–173, October 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lovell:1987:BRB

[Lov87] Bernard Lovell. Book review: British space science: History of British Space Science. Journal for the History of Astronomy, 18(2):144–146, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lowne:1988:OGG

[Low88] C. M. Lowne. The object-glass of the Greenwich “great equatorial telescope”. Journal for the History of Astronomy, 19 (3):169–182, August 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lowenstein:2009:ERU

[Low09] Noah Lowenstein. Essay review: Updating the DSB, Complete Dictionary of Scientific Biography. Journal for the History of Astronomy, 40(4): 471–472, November 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Lange:2005:TPE

[LS05] R. Lange and N. M. Swerdlow. Two programs, for Ephemeris and visibility calculations, useful for historical applications. Journal for the History of Astronomy, 36(3):335–336, August 2005. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Landi:2024:DTG

[LS24a] Enrico Landi and Francesca Schironi. Dia tōn grammōn: Hipparchus on simultaneous risings and settings. Journal for the History of Astronomy, 55(2):157–178, May 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241234682.

Leon:2024:BHA

[LS24b] José Guillermo Sánchez León and Pablo Recio Sánchez. The bizarre history of the astrological vault “El Cielo de Salamanca”. Journal for the History of Astronomy, 55(1): ??, February 1, 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231215795.

Luthy:1991:BRG

[Lut91] Christoph Luthy. Book review:

Galilean studies: Galileo: Pioneer Scientist. Journal for the History of Astronomy, 22(4): 324–325, November 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Luthy:1992:BRR

[Lüt92] Christoph Lüthy. Book review:

Renaissance art and the scientific revolution: The Heritage of Giotto’s Geometry: Art and Science on the Eve of the Scientific Revolution. Journal for the History of Astronomy, 23 (3):227–229, August 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Luthy:2000:BRO

[Lüt00] Christoph Lüthy. Book review: Optics and visual perception: La Mutation du Visible. Essai sur la portée épistémologique des Instruments d’optique au XVIIe siècle. Vol. I: Du Sidereus Nuncius de Galilée à la Dioptrique cartésienne. Journal for the History of Astronomy, 31(3):273–274, August 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Luthy:2010:CCC

[Lüt10] Christoph Lüthy. Centre, circle, circumference: Giordano Bruno’s astronomical woodcuts. Journal for the History of Astronomy, 41(3):311–327, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Li:1999:ASE

[LZ99] Y. Li and C. Z. Zhang. Accuracy of solar eclipse calculations, 1644–1785, with the Chinese Shoushi calendar. Journal for the History of Astronomy, 30(2):161–167, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ma:2023:SED

[Ma23] Lihua Ma. The solar eclipse of A.D. 1221 May 23 and the value of ΔT. Journal for the History of Astronomy, 54(4):425–435, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231200486.

Maas:2016:DAT

[Maa16] Ad Maas. Dutch astronomy in the twentieth century. Journal for the History of Astronomy, 47(3):347–349, August 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

MacLachlan:1976:BRT

[Mac76]  James MacLachlan.  Book

review: The tragic Galileo: Galileo. Journal for the History of Astronomy, 7(3):210– 212, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mack:1990:SCW

[Mac90] Pamela E. Mack. Strategies and compromises: Women in astronomy at Harvard College Observatory, 1870–1920. Journal for the History of Astronomy, 21(1):65–76, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mack:1993:BRP

[Mac93] Pamela E. Mack. Book review: Planetary astronomy in the space age: Space Technology and Planetary Astronomy. Journal for the History of Astronomy, 24(4):306–307, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

MacLachlan:1994:SD

[Mac94] James MacLachlan. Stillman Drake (1910–1993). Journal for the History of Astronomy, 25 (1):73–75, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

MacMinn:1998:APT

[Mac98] Donn MacMinn. An analysis of Ptolemy’s treatment of retrograde motion. Journal for the History of Astronomy, 29 (3):257–270, August 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

MacDonald:2010:IRB

[Mac10] Lee T. MacDonald. Isaac Roberts, E. E. Barnard and the nebulae. Journal for the History of Astronomy, 41(2):239– 259, May 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Macdonald:2015:SSM

[Mac15] Lee T. Macdonald. “Solar spot mania”: The origins and early years of solar research at Kew Observatory, 1852–1860. Journal for the History of Astronomy, 46(4):469–490, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Macdonald:2020:PMR

[Mac20] Lee T. Macdonald. Proposals to move the Royal Observatory, Greenwich, 1836–1944. Journal for the History of Astronomy, 51(3):272–304, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620936625.

Maffeo:2005:BRJ

[Maf05] S. J. Sabino Maffeo. Book review: Jesuit astronomers: Searching the Heavens and the History of Jesuit Observatories. Journal for the History of Astronomy, 36(4):460–461, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Magli:2023:FCR

[Mag23] Giulio Magli. A Festschrift for Clive Ruggles. Journal for the History of Astronomy, 54(1):111–112, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221142546.

Malmstrom:1978:RCM

[Mal78] Vincent H. Malmström. A reconstruction of the chronology of Mesoamerican calendrical systems. Journal for the History of Astronomy, 9 (2):105–116, June 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malewicz:1990:BRC

[Mal90] Małgorzata H. Malewicz. Book review: The Constellations in the Arabic-Latin Tradition, Peter Apian und Azophi: Arabische Sternbilder in Ingolstadt im frühen 16. Jahrhundert. Journal for the History of Astronomy, 21(3):300–301, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malewicz:1997:BRA

[Mal97] Małgorzata Malewicz. Book review: Astrology and theology: History, Prophecy, and the Stars: The Christian Astrology of Pierre d’Ailly, 1350– 1420. Journal for the History of Astronomy, 28(4):354–355, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malet:2010:PVL

[Mal10] Antonio Malet. Paralipòmens a Vitel· lió: els orígens de l’òptica moderna, volume 9 of Clàssics de la ciència. Institut d’Estudis Catalans, Barcelona, Spain, 2010. ISBN 84-9965-031-7. cxxxv + 265 pp. LCCN ????

Malpangotto:2013:BRP

[Mal13] Michela Malpangotto. Book review: Pedro Nunes on Peurbach: Pedro Nunes — Obras, in Theoricas planetarum Georgii Purbachii annotationes. Journal for the History of Astronomy, 44(1): 112–114, February 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malville:2015:BRA

[Mal15] J. McKim Malville. Book review: Alternate ontologies and cultural astronomies: African Cosmos: Stellar Arts. Edited by Christine Mullen Kreamer (The Monacelli Press, New York, 2012). Pp. 367. $60. ISBN 978-1-58093-343-8. Journal for the History of Astronomy, 46(3):370–371, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malherbe:2023:YSP

[Mal23] Jean-Marie Malherbe. 130 years of spectroheliograms at Paris–Meudon observatories (1893–2023). Journal for the History of Astronomy, 54 (3):274–315, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231184193.

Marsden:1975:BRB

[Mar75] Brian G. Marsden. Book review: Bright Comets: Helle Kometen -86 bis +1950: Ephemeriden und Kurzbeschreibungen. Journal for the History of Astronomy, 6(1):68–69, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Marsden:1987:BRR

[Mar87]  Brian G. Marsden.  Book

review: Records of Halley’s comet: Halley’s Comet in History. Journal for the History of Astronomy, 18(4):292–294, November 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Marvin:1989:ERC

[Mar89]  Ursula B. Marvin.  Essay

review: Controversies over craters, meteorite craters, Coon Mountain controversies: Meteor Crater and the Development of Impact Theory. Journal for the History of Astronomy, 20(2):132–135, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martin:1995:VDC

[Mar95] Frederick Martin. Venus and the Dresden Codex eclipse table. Journal for the History of Astronomy, 26(20):S57–S73, February 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martens:1997:BRM

[Mar97] Rhonda Martens. Book review:

Music and the scientific revolution: Music and Science in the Age of Galileo. Journal for the History of Astronomy, 28(4):355–357, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martens:1999:BRE

[Mar99] Rhonda Martens. Book review: Educating Kepler: Kepler’s Tübingen: Stimulus to a Theological Mathematics. Journal for the History of Astronomy, 30(3):324–325, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Maravelia:2002:ONT

[Mar02] Amanda-Alice Maravelia. The orientations of the nine Tholos tombs at Mycenae. Journal for the History of Astronomy, 33 (27):S63–S66, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martens:2003:BRN

[Mar03] Rhonda Martens. Book review: Nature and revelation: God’s Two Books: Copernican Cosmology and Biblical Interpretation in Early Modern Science. Journal for the History of Astronomy, 34(1):115–117, February 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martens:2012:BRN

[Mar12a] Rhonda Martens. Book review: New stars in the Renaissance, Change and Continuity in Early Modern Cosmology. Journal for the History of Astronomy, 43(2):255–256, May 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martzloff:2012:BRC

[Mar12b] Jean-Claude Martzloff. Book review: Calendars and chronological tables, Calendars and Years II: Astronomy and Time in the Ancient and Medieval World. Journal for the History of Astronomy, 43(3):360– 361, August 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Martin:2014:BRA

[Mar14] Craig Martin. Book review: Aristotle’s logic and Copernicus’s astronomy: Copernicus and the Aristotelian Tradition: Education, Reading, and Philosophy in Copernicus’s Path to Heliocentricism. André Goddu (Medieval and Early Modern Science, xii; Brill, Leiden, 2010). Pp. xxviii + 545. EUR 129. ISBN 978-90-04-18107-6. Journal for the History of Astronomy, 45(4): 497–500, November 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Margolis:2022:MIT

[Mar22] Emily A. Margolis. Managing innovation in telescope making. Journal for the History of Astronomy, 53(4): 497–498, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221111100.

Marcacci:2023:GBR

[Mar23] Flavia Marcacci. G. B. Riccioli’s geo-heliocentric use of epicepicycles, ellipses and spirals. Journal for the History of Astronomy, 54(2):171– 192, May 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231165331.

Maslikov:2021:GPS

[Mas21] Sergei J. Maslikov. The Greek portable sundial from Memphis rediscovered. Journal for the History of Astronomy, 52(3):311–324, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211033068.

Maurer:2013:WHP

[Mau13] Andreas Maurer. William Herschel’s planned 30-ft reflector. Journal for the History of Astronomy, 44(2):211–212, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Max:2012:BRH

[Max12] Claire E. Max. Book review: a history of adaptive optics, The Adaptive Optics Revolution: a History. Journal for the History of Astronomy, 43(3):370– 371, August 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mayer:2011:BRG

[May11]  Thomas F. Mayer.  Book

review: On Galileo’s letter to the Grand Duchess: Libertas Philosophandi: Teologia e Filosofa nella Lettera alla Granduchessa Cristina di Lorena di Galileo Galilei. Journal for the History of Astronomy, 42(4):534–536, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Maurer:2016:WHR

[MBH16] Andreas Maurer, Vasily Belokurov, and Michael Hoskin. William Herschel’s Russian telescopes. Journal for the History of Astronomy, 47(1): 52–60, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/47/1/
                                        52.full.pdf+html.

McCluskey:1977:AHI

[McC77] Stephen C. McCluskey. The astronomy of the Hopi Indians.

Journal for the History of Astronomy, 8(3):174–195, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1982:BRC

[McC82] Stephen McCluskey. Book review: Calculators for astronomers: Mathematical Astronomy with a Pocket Calculator, Practical Astronomy with Your Calculator. Journal for the History of Astronomy, 13(3):218–220, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1983:BRC

[McC83a] Stephen C. McCluskey. Book review: At the crossroads of the Earth and the sky: An Andean Cosmology. Journal for the History of Astronomy, 14(5):S72–S73, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1983:MOV

[McC83b] Stephen C. McCluskey. Maya observations of very long periods of Venus. Journal for the History of Astronomy, 14 (2):92–101, June 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1985:BRP

[McC85] Stephen C. McCluskey. Book review: Petroglyphs of Ohio.

Journal for the History of Astronomy, 16(9):S158, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1989:MQD

[McC89] Stephen C. McCluskey. The mid-quarter days and the historical survival of British folk astronomy. Journal for the History of Astronomy, 20(13): S1–S19, February 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1990:CSF

[McC90] Stephen C. McCluskey. Calendars and symbolism: Functions of observation in Hopi astronomy. Journal for the History of Astronomy, 21(15):S1–S16, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1991:BRE

[McC91] Stephen C. McCluskey. Book review: Essays in medieval astronomy and optics: Astronomy and Optics from Pliny to Descartes: Texts, Diagrams and Conceptual Structures. Journal for the History of Astronomy, 22(4):322–324, November 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCarthy:1993:EPF

[McC93] Daniel McCarthy. Easter principles and a Fifth-Century lunar cycle used in the British Isles. Journal for the History of Astronomy, 24(3):204–224, August 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McConnell:1994:AWV

[McC94] Anita McConnell. Astronomers at war: The viewpoint of Troughton & Simms. Journal for the History of Astronomy, 25(3):219–235, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCarthy:1998:JMJ

[McC98a] Dennis D. McCarthy. The Julian and modified Julian dates. Journal for the History of Astronomy, 29(4):327–330, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1998:BRC

[McC98b] Stephen McCluskey. Book review: Cosmovisions and power: Skywatchers, Shamans & Kings: Astronomy and the Archaeology of Power. Journal for the History of Astronomy, 29(23):S89–S90, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:1999:BRT

[McC99] Stephen McCluskey. Book review: Theory versus instrumentation: Astronomy Through the Ages: The Story of the Human Attempt to Understand the Universe. Journal for the History of Astronomy, 30(2):180–181, May 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2003:CCC

[McC03] Stephen C. McCluskey. Changing contexts and criteria for the justification of computistical knowledge and practice. Journal for the History of Astronomy, 34(2):201–217, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2004:ERE

[McC04] Stephen McCluskey. Essay review: Early medieval planetary astronomy: The Revival of Planetary Astronomy in Carolingian and Post-Carolingian Europe. Journal for the History of Astronomy, 35(1):109–111, February 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2007:ERA

[McC07] Stephen McCluskey. Essay review: Archaeoastronomy at the crossroads: Current Studies in Archaeoastronomy: Conversations across Time and Space, Viewing the Sky through Past and Present Cultures. Journal for the History of Astronomy, 38(2):229–236, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCutcheon:2008:BRE

[McC08] Robert A. McCutcheon. Book review: Early Russian astronomy; Astronomiya drevnei Rusi. Journal for the History of Astronomy, 39(3):412–414, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2011:BRC

[McC11] Stephen C. McCluskey. Book review: The Christian era: The Easter computus and the origins of the Christian era. Journal for the History of Astronomy, 42(2):265–266, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2013:ERL

[McC13a] Stephen McCluskey. Essay review: Lunar calendars, Living the Lunar Calendar. Journal for the History of Astronomy, 44(4):485–487, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McConnell:2013:BRR

[McC13b] Craig Sean McConnell. Book review: Radio astronomy in infancy: Cosmic Noise: a History of Early Radio Astronomy. Journal for the History of Astronomy, 44(1):109–110, February 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCray:2013:BRS

[McC13c] W. Patrick McCray. Book review: Sublime images from Hubble, Picturing the Cosmos: Hubble Space Telescope Images and the Astronomical Sublime. Journal for the History of Astronomy, 44(3):372–374, August 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McCluskey:2017:ARN

[McC17] Stephen C. McCluskey. Archaeoastronomy and refraction near the Earth’s surface. Journal for the History of Astronomy, 48(3):329–345, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617713398.

McCluskey:2019:BRY

[McC19] Stephen McCluskey. Book review: 1000 years of calendar reform: Scandalous Error: Calendar Reform and Calendrical Astronomy in Medieval Europe C. Philipp E. Nothaft, (Oxford University Press, Oxford, 2018). Pp. xvii + 357. £75. ISBN 978-0-19-879955-9. Journal for the History of Astronomy, 50(1):98–100, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618810567.

McKnight:1978:BRB

[McK78] John L. McKnight. Book review: Benjamin Martin, Benjamin Martin: Author, Instrument-Maker and “Country Showman”. Journal for the History of Astronomy, 9 (2):147–149, June 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McKnight:1987:BRM

[McK87] John L. McKnight. Book review: More on Benjamin Martin: Benjamin Martin: Supplement. Journal for the History of Astronomy, 18(4):294, November 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMullin:1974:ERG

[McM74] Ernan McMullin. Essay review: Galileo’s methodology: Galileo’s Intellectual Revolution. Journal for the History of Astronomy, 5(1):56–58, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMullin:1989:ERC

[McM89] Ernan McMullin. Essay review:

Coming to terms with Copernicanism, Il suo Diario e la Questione Galileiana. Journal for the History of Astronomy, 20 (1):66–69, February 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMahon:2001:BRO

[McM01] John M. McMahon. Book review: Ovidian astronomy: Ovid, Aratus and Augustus: Astronomy in Ovid’s Fasti. Journal for the History of Astronomy, 32(2):176–178, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMullin:2003:BRG

[McM03] Ernan McMullin. Book review: Galileian studies: Galileo in Context. Journal for the History of Astronomy, 34(4): 456–458, November 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMullin:2005:ERR

[McM05] Ernan McMullin. Essay review: Religion and modern cosmology: Matter and Spirit in the Universe: Scientific and Religious Preludes to Modern Cosmology. Journal for the History of Astronomy, 36(4):448–452, November 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McMullin:2009:GAT

[McM09] Ernan McMullin. The Galileo affair: Two decisions. Journal for the History of Astronomy, 40(2):191–212, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2021:AAC

[MD21] S. Mohammad Mozaffari and Jeremy J. Drake. Algol anomaly or careful observations of its brightness? the values recorded for the magnitude of Algol in the medieval astronomical corpus. Journal for the History of Astronomy, 52(1):77–103, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828621990915.

Meadows:1970:A

[Mea70a] A. J. Meadows. Abstracts.

Journal for the History of Astronomy, 1(2):160–163, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:1970:BRP

[Mea70b] A. J. Meadows. Book review: Pulsars: Six Months after Discovery: Pulsating Stars. Journal for the History of Astronomy, 1(1):88–89, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:1971:BRC

[Mea71] A. J. Meadows. Book review: Chaucer’s use of astronomy: Kalenderes Enlumyned Ben They: Some Astronomical Themes in Chaucer. Journal for the History of Astronomy, 2(1):45, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:1981:BRA

[Mea81] A. J. Meadows. Book review: Astrology in the British library: Astrology. Journal for the History of Astronomy, 12(3):213–214, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:1985:BRN

[Mea85]  Jack Meadows.  Book re-

view: Nineteenth-Century astronomy: The History of Astronomy from Herschel to Hertzsprung. Journal for the History of Astronomy, 16(1): 69–71, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:1986:BRA

[Mea86] A. J. Meadows. Book review: Astrology in English literature: The Forgotten Sky: a Guide to Astrology in English Literature. Journal for the History of Astronomy, 17(3):202–203, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malville:1991:LSC

[MEA91] J. McKim Malville, Frank W. Eddy, and Carol Ambruster. Lunar standstills at Chimney Rock. Journal for the History of Astronomy, 22(16):S43–S50, February 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meadows:2002:BRE

[Mea02] Jack Meadows. Book review:

Early spectroscopy, Mapping the Spectrum: Techniques of Visual Representation in Research and Teaching. Journal for the History of Astronomy, 33(3):305–307, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meli:1989:BRC

[Mel89] Domenico Bertoloni Meli. Book review: The Cracow Newton Conference, Newton and the New Direction in Science: Proceedings of the Cracow Conference 25 to 28 May 1987. Journal for the History of Astronomy, 20(3):222–223, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Melia:2009:CEC

[Mel09] Fulvio Melia. Cracking the Einstein code: Relativity and the birth of black hole physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-51951-1. xi + 137 pp. LCCN QC173.6 .M434 2009. With an afterword by Roy Kerr.

Mercier:1977:BRA

[Mer77] Raymond Mercier. Book review: The Antikythera mechanism, Gears from the Greeks. Journal for the History of Astronomy, 8(2):143–145, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:1988:ERS

[Mer88] R. P. Mercier. Essay review: Sources for Indian astronomy: Indian Astronomy: a Source Book (Based Primarily on Sanskrit Texts). Journal for the History of Astronomy, 19(1): 49–51, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:1992:BRM

[Mer92] Raymond Mercier. Book review: Myth, ritual and the stars: The Myth of Replacement: Stars, Gods and Order in the Universe. Journal for the History of Astronomy, 23(4): 303–305, November 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:1996:BRS

[Mer96] Raymond Mercier. Book review: Studies in Islamic astronomy: Astronomy in the Service of Islam, Islamic Astronomy and Medieval Spain. Journal for the History of Astronomy, 27(3):275–276, August 1996. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:1998:ATG

[Mer98a] Raymond Mercier. The astronomical tables of George Gemistus Plethon. Journal for the History of Astronomy, 29 (2):117–127, May 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:1998:GCC

[Mer98b] Raymond Mercier. A general calendar conversion program. Journal for the History of Astronomy, 29(4):379–380, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2001:RLB

[Mer01] Raymond Mercier. Roger Louis Billard (1922–2000). Journal for the History of Astronomy, 32(4):369–370, November 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2003:BRI

[Mer03] Raymond Mercier. Book review: Indian astronomy and chronology: Hindu Zodiac and Ancient Astronomy. Journal for the History of Astronomy, 34(2):240–241, May 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2006:BRS

[Mer06] Raymond Mercier. Book review: a sailing manual in the Arabic tradition: The Rahmani of M. P. Kunhikunhi Malmi of Kavaratti: a Sailing Manual of Lakshadweep. Journal for the History of Astronomy, 37 (2):241–243, May 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2009:BRA

[Mer09] Raymond Mercier. Book review: Astrology down the ages, A History of Horoscopic Astrology: From the Babylonian Period to the Modern Age. Journal for the History of Astronomy, 40(1):125–126, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2010:BRM

[Mer10] Raymond Mercier. Book review: Manilius in context: Manilius and His Intellectual Background. Journal for the History of Astronomy, 41(3): 420–422, August 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mercier:2011:BRP

[Mer11] Raymond Mercier. Book review: Ptolemy in perspective: Ptolemy in Perspective: Use and Criticism of His Work from Antiquity to the Nineteenth Century. Journal for the History of Astronomy, 42(3):405– 408, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meschiari:2011:BRZ

[Mes11] Alberto Meschiari. Book review: Zach’s collected correspondence: The Collected Correspondence of Baron Franz Xaver von Zach. Journal for the History of Astronomy, 42 (2):267–269, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Messeri:2015:BRD

[Mes15] Lisa Messeri. Book review: On dark matters and metaphors: The Cosmic Cocktail: Three Parts Dark Matter. Katherine Freese (Princeton University Press, Princeton, 2014). Pp. xi + 250. $29.95. ISBN 978-0-691-15335-3. Journal for the History of Astronomy, 46 (2):247–249, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Meskens:2020:AAU

[Mes20]  Ad Meskens.  Astronomy

and the Archduke: Unpublished letters on SN1604 by Brengger, Coignet, and Kepler in the archives of Albert VII of Austria. Journal for the History of Astronomy, 51(3):324–367, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620943990.

Meskens:2021:AAT

[Mes21] Ad Meskens. Astrology and the Archduke: Two unpublished letters by Tycho Brahe on the horoscope of Albert VII of Austria. Journal for the History of Astronomy, 52 (3):249–266, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211029054.

Meurs:2010:OLD

[Meu10] E. J. A. Meurs. The objective lens of the Dunsink Ramsden Circle. Journal for the History of Astronomy, 41(4):504–506, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Malin:1983:NAM

[MG83] S. R. C. Malin and D. Gubbins. The need for archival magnetic measurements. Journal for the History of Astronomy, 14(1): 70–75, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Moesgaard:1986:ERR

[MG86] K. P. Moesgaard and Owen Gingerich. Essay review: Rheticus: Narrationes Prima et Secunda: Georgii Joachimi Rhetici Narratio Prima. Journal for the History of Astronomy, 17(2):130–133, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McNally:1988:WWO

[MH88] Derek McNally and Michael Hoskin. William E. Wilson’s observatory at Daramona House. Journal for the History of Astronomy, 19(2):146– 153, May 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2017:AGR

[MHS17] Leslie V. Morrison, Catherine Y. Hohenkerk, and F. Richard Stephenson. The Almagest Greek and Roman occultations re-visited. Journal for the History of Astronomy, 48 (4):405–416, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617739992.

Murata:2023:CAQ

[MHS23] Koji Murata, Hisashi Hayakawa, and Mitsuru Sôma. A critical assessment of questionable solar eclipse memories in the Byzantine Empire from the fourth to sixth centuries CE. Journal for the History of Astronomy, 54(2):193– 212, May 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231167185.

Morrison:2023:TSE

[MHZS23] Leslie V. Morrison, Catherine Y. Hohenkerk, Marek Zawilski, and F. Richard Stephenson. Total solar eclipse of AD 1133 and ΔT. Journal for the History of Astronomy, 54(4):469–480, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231209030.

Mialet:2012:HIS

[Mia12] Hélène Mialet. Hawking incorporated: Stephen Hawking and the anthropology of the knowing subject. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-52226-1 (hardcover), 0-226-52228-8 (paperback). ix + 266 pp. LCCN QC16.H33 M53 2012.

Miller:1980:BRA

[Mil80] Arthur I. Miller. Book review: Anti-Einstein, The Einstein Myth and the Ives Papers: a Counter-Revolution in Physics. Journal for the History of Astronomy, 11(3):217– 218, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Miller:1981:AES

[Mil81] Arthur I. Miller. Albert Einstein’s Special Theory of Relativity: emergence (1905) and early interpretation, 1905– 1911. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-04680-6, 0-201-04679-2 (paperback). xxviii + 466 pp. LCCN QC173.52 .M54.

Mills:2004:BRV

[Mil04] Allan Mills. Book review: Vermeer and his scientific contemporaries: Giants of Delft: Johannes Vermeer and the Natural Philosophers. The Parallel Search for Knowledge during the Age of Discovery. Journal for the History of Astronomy, 35(2):242–244, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Miller:2008:MFR

[Mil08] David Marshall Miller. O male factum: Rectilinearity and Kepler’s discovery of the ellipse. Journal for the History of Astronomy, 39(1):43–63, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Miniati:2000:BRB

[Min00]  Mara Miniati.  Book re-

view: Bedini’s collected papers: Science and Instruments in Seventeenth-Century Italy. Journal for the History of Astronomy, 31(1):81–82, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mirwald:2017:AND

[Mir17]  Benjamin Mirwald.  As-

tronomers’ networks and the discovery of Ceres. Essay review: Discovery of the First Asteroid, Ceres: Historical Studies in Asteroid Research. Clifford Cunningham (Springer, Heidelberg, 2016). Pp. xiii + 333. $35. ISBN 978-3-319-21776-5. Journal for the History of Astronomy, 48(3):375–377, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617717584.

Mitton:1973:CAA

[Mit73] Simon Mitton. Conference on ancient astronomy. Journal for the History of Astronomy, 4(1): 68–71, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mitton:1986:BRR

[Mit86] Simon Mitton. Book review: Radio astronomy reprints: Classics in Radio Astronomy. Journal for the History of Astronomy, 17(3):212, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mitton:2013:BRC

[Mit13] Simon A. Mitton. Book review: Cultural evolution: Cosmos & Culture: Cultural Evolution in a Cosmic Context. Journal for the History of Astronomy, 44(1):104–105, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mitton:2019:BRR

[Mit19] Simon Mitton. Book review:

Radio astronomy in Australia: Four Pillars of Radio Astronomy: Mills, Christiansen, Wild, Bracewell. Robert H Frater, W Miller Goss and Harry W Wendt (Springer, Cham, 2017). Pp. xvii + 199. EUR 32. ISBN 978-3-319-65598-7 (paper). Journal for the History of Astronomy, 50(1):111–112, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618823985.

Mosley:2003:ECE

[MJT03] Adam Mosley, Nicholas Jardine, and Karin Tybjerg. Epistolary culture, editorial practices, and the propriety of Tycho’s astronomical letters. Journal for the History of Astronomy, 34(4):421–451, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison-Low:2007:BRF

[ML07] A. D. Morrison-Low. Book review: Festschrift for a museum: The Whipple Museum of the History of Science. Journal for the History of Astronomy, 38 (2):245–247, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison-Low:2014:BRA

[ML14] A. D. Morrison-Low. Book review: Astronomy in Glasgow: Reflections on the Astronomy of Glasgow: A Story of Some Five Hundred Years. David Clarke (Edinburgh University Press, Edinburgh, 2013). Pp. x + 373. £90. ISBN 978-0-7486-7889-1. Journal for the History of Astronomy, 45(4):495–496, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

McKenna-Lawlor:1984:CMO

[MLH84] Susan McKenna-Lawlor and Michael Hoskin. Correspondence of Markree Observatory. Journal for the History of Astronomy, 15(1):64–68, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Moesgaard:1975:TOP

[Moe75] Kristian Peder Moesgaard. Tychonian observations, perfect numbers, and the date of creation: Longomontanus’s solar and precessional theories. Journal for the History of Astronomy, 6(2):84–99, June 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Moesgaard:1977:BRC

[Moe77] Kristian Peder Moesgaard.

Book review: Copernicus Washington Symposium: Copernicus, Yesterday and Today. Journal for the History of Astronomy, 8(1):61–62, February 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Moesgaard:1980:BRP

[Moe80] Kristian Peder Moesgaard.

Book review: Ptolemy’s Failings, the Crime of Claudius Ptolemy. Journal for the History of Astronomy, 11(2):133– 135, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Moesgaard:1983:AET

[Moe83] Kristian Peder Moesgaard. Ancient ephemeris time in Babylonian astronomy. Journal for the History of Astronomy, 14 (1):47–60, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Molland:1970:BRP

[Mol70] A. G. Molland. Book review:

Philosophy of Earth and heavens: Le Livre du Ciel et du Monde. Journal for the History of Astronomy, 1(1):79–80, February 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Molland:1976:BRF

[Mol76] A. G. Molland. Book review: Fourteenth-Century astronomy, L’Oeuvre astronomique de Thémon Juif, maître parisien du XIV. Journal for the History of Astronomy, 7(1):68–69, February 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Molland:1993:BRM

[Mol93] A. G. Molland. Book review: Medieval science in context: The Beginnings of Western Science: The European Scientific Tradition in Philosophical, Religious, and Institutional Context, 600 B.C. to A.D. 1450. Journal for the History of Astronomy, 24(4): 304–305, November 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Molnar:2003:EAR

[Mol03] Michael R. Molnar. The evidence for Aries the Ram as the astrological sign of Judea. Journal for the History of Astronomy, 34(3):325–327, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mollan:2014:WPE

[Mol14a] R. C. (R. Charles) Mollan, editor. William Parsons, 3rd Earl of Rosse: astronomy and the castle in nineteenth-century Ireland, volume 4 of Royal Dublin society — Science and Irish culture series. Manchester University Press, Manchester, UK, 2014. ISBN 0-7190-9144-6 (hardcover). xxii + 368 pp. LCCN QB36.R77 W55 2014.

Mollan:2014:BRT

[Mol14b] R. Charles Mollan. Book review: Telescopes restored: Classical Telescopes: A Guide to Collecting, Restoring, and Using Telescopes of Yesteryear. Neil English (Springer, New York, 2013). Pp. xvi + 239 (paperback). $39.95. ISBN 978-1-4614-4423-7. Journal for the History of Astronomy, 45(4): 494–495, November 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Montgomery:1982:NBC

[Mon82] James W. Montgomery, Jr.

Nathaniel Bowditch’s classical “black hole” calculation of 1808. Journal for the History of Astronomy, 13(1):54– 55, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Montgomery:1994:FND

[Mon94] Scott L. Montgomery. The first naturalistic drawings of the Moon: Jan Van Eyck and the art of observation. Journal for the History of Astronomy, 25(4):317–320, November 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Monroe:2021:LBA

[Mon21] M. Willis Monroe. Late Babylonian astronomy and astrology. Journal for the History of Astronomy, 52(4): 492–494, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211042962.

Monroe:2022:AAD

[Mon22] M. Willis Monroe. Astronomical and astrological diagrams from cuneiform sources. Journal for the History of Astronomy, 53(3):338–361, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221110919.

Moore:1994:NDL

[Moo94]  Sarah Moore.  A newly-

discovered letter of J. F. W. Herschel concerning the Plumian Professorship. Journal for the History of Astronomy, 25(2):142–143, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morgan:1980:HAW

[Mor80] Julie Morgan. The Huggins Archives at Wellesley College. Journal for the History of Astronomy, 11(2):147, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:1997:BRC

[Mor97] Philip Morrison. Book review:

Celestial mythology worldwide: Beyond the Blue Horizon: Myths and Legends of the Sun, Moon, Stars, and Planets. Journal for the History of Astronomy, 28(1):87–89, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2001:ERH

[Mor01a] Leslie Morrison. Essay review: History in the service of astronomy: Observations and Predictions of Eclipse Times by Early Astronomers, Historical Eclipses and Earth’s Rotation. Journal for the History of Astronomy, 32(2):160– 162, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2001:BRA

[Mor01b] Philip Morrison. Book review: Astronomy in myth: Medusa: Solving the Mystery of the Gorgon. Journal for the History of Astronomy, 32 (1):89–90, February 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2008:BRE

[Mor08] Robert Morrison. Book review: Epicycles and Eccentrics in the Middle Ages, Studies in Medieval Astronomy and Optics. Journal for the History of Astronomy, 39(4):535–537, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morawiec:2014:PPC

[Mor14]  Adam Morawiec.  On the

provenance of a Prague copy of Copernicus’s De Revolutionibus. Journal for the History of Astronomy, 45(2):243–246, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2016:SNP

[Mor16] Robert G. Morrison. Shirazi’s non-Ptolemaic models. Journal for the History of Astronomy, 47(2):218–219, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2021:DSI

[Mor21] Robert G. Morrison. A definitive survey of Iberian and Maghribi astronomy. Journal for the History of Astronomy, 52(4):499–500, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211043317.

Moss:1986:BRG

[Mos86] Jean Dietz Moss. Book review: Galileo the iconoclast: Galileo Galilei — Prozess ohne Ende: Eine Biographie Piper. Journal for the History of Astronomy, 17(1):67–69, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mosley:2003:ERH

[Mos03]  Adam Mosley.  Essay re-

view: The history and reception of Copernicus’s De Revolutionibus: An Annotated Census of Copernicus’ De Revolutionibus (Nuremberg, 1543 and Basel, 1566). Journal for the History of Astronomy, 34(1): 97–111, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mosley:2007:BRF

[Mos07]  Adam Mosley.  Book re-

view: Festschrift for Volker Bialas, Miscellanea Kepleriana: Festschrift für Volker Bialas zum 65. Geburtstag. Journal for the History of Astronomy, 38(3):380–381, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mosley:2009:BRL

[Mos09] Adam Mosley. Book review:

Landgrave Wilhelm IV, Der Ptolomäus von Kassel: Landgraf Wilhelm IV. Von Hessen– Kassel und die Astronomie. Journal for the History of Astronomy, 40(3):357–358, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mosley:2010:BRS

[Mos10] Adam Mosley. Book review: Spanish cosmography: Secret Science: Spanish Cosmography and the New World. Journal for the History of Astronomy, 41(4):507–508, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mosley:2011:VMC

[Mos11] Adam Mosley. Vincenzo Maria Coronelli’s Atlante Veneto and the diagrammatic tradition of cosmography. Journal for the History of Astronomy, 42(1): 27–53, February 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2013:BFP

[Moz13a] S. Mohammad Mozaffari. Biruni’s four-point method for determining the eccentricity and the direction of the apsidal lines of the superior planets. Journal for the History of Astronomy, 44(2):207–211, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2013:LMAa

[Moz13b] S. Mohammad Mozaffari. Limitations of methods: The accuracy of the values measured for the Earth’s/Sun’s orbital elements in the Middle East, A.D. 800–1500, Part 1. Journal for the History of Astronomy, 44 (3):313–336, August 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2013:LMAb

[Moz13c] S. Mohammad Mozaffari. Limitations of methods: The accuracy of the values measured for the Earth’s/Sun’s orbital elements in the Middle East, A.D. 800–1500, Part 2. Journal for the History of Astronomy, 44(4):389–411, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2014:BEP

[Moz14] S. Mohammad Mozaffari.

Biruni’s examination of the path of the centre of the epicycle in Ptolemy’s lunar model. Journal for the History of Astronomy, 45(1):123–127, February 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2016:MBS

[Moz16] S. Mohammad Mozaffari. A medieval bright star table: The non-Ptolemaic star table in the Ilkhani Zij. Journal for the History of Astronomy, 47(3):294– 316, August 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mozaffari:2019:OEV

[Moz19] S. Mohammad Mozaffari. The orbital elements of Venus in Medieval Islamic astronomy:

Interaction between traditions and the accuracy of observations. Journal for the History of Astronomy, 50(1):46– 81, February 1, 2019. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618808877.

Mozaffari:2021:PIO

[Moz21] S. Mohammad Mozaffari. Photographing Indian observatories. Journal for the History of Astronomy, 52(4): 494–495, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211032416.

Mozaffari:2022:AIA

[Moz22] S. Mohammad Mozaffari. An analysis of Ibn al-Shat.ir’s star table. Journal for the History of Astronomy, 53(2):163– 196, May 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221090918.

Martlew:1993:NMP

[MR93] Roger D. Martlew and Clive L. N. Ruggles. The North Mull Project (4): Excavations at Ardnacross 1989–91. Journal for the History of Astronomy, 24(18):S55–S64, February 1993. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Meeus:1985:NTM

[MS85] Jean Meeus and Virginia G.

Smith. A new test for Maya astronomical observation: Occultations of Venus by the Moon. Journal for the History of Astronomy, 16(9):S97–S101, February 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Morrison:2004:HVE

[MS04] L. V. Morrison and F. R.

Stephenson. Historical values of the Earth’s clock error Δ and the calculation of eclipses. Journal for the History of Astronomy, 35(3):327–336, August 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See erratum [MS05].

Morrison:2005:EHV

[MS05] L. V. Morrison and F. R.

Stephenson. Erratum: Historical values of the Earth’s clock error. Journal for the History of Astronomy, 36(3):339, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [MS04].

Morrison:2019:RTS

[MSH19a] L. V. Morrison, F. R. Stephenson, and C. Y. Hohenkerk. Rome and the total solar eclipse of BC 188 July 17. Journal for the History of Astronomy, 50(3):366–372, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619863243.

Morrison:2019:EH

[MSH19b] Leslie V. Morrison, F. Richard Stephenson, and Catherine Y. Hohenkerk. On the eclipse of Hipparchus. Journal for the History of Astronomy, 50 (1):3–15, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618817175.

Morrison:2020:FAS

[MSH20] Leslie V. Morrison, F. Richard Stephenson, and Catherine Y. Hohenkerk. Fotheringham’s 1920 accelerations of the Sun and Moon revisited. Journal for the History of Astronomy, 51(2):209–224, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620918438.

Morrison:2021:AMC

[MSH21] Leslie V. Morrison, F. Richard Stephenson, and Catherine Y. Hohenkerk. Accuracy of medieval Chinese and Middle-Eastern timings of eclipses. Journal for the History of Astronomy, 52(4):397–413, November 1, 2021. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211048263.

Morrison:2022:RTS

[MSH22a] L. V. Morrison, F. R. Stephenson, and C. Y. Hohenkerk. Rome and the total solar eclipse of BC188 July 17: Apology. Journal for the History of Astronomy, 53(3):362– 363, August 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221107706.

Morrison:2022:AER

[MSH22b] Leslie V. Morrison, F. Richard Stephenson, and Catherine Y. Hohenkerk. Accuracy of eclipse records in the Anglo-Saxon Chronicle. Journal for the History of Astronomy, 53(2):209–216, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221097111.

Mucke:1975:PAC

[Muc75] Hermann Mucke. The planetarium as an analogue computer. Journal for the History of Astronomy, 6(1):53–57, February 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

MuellerSJ:2000:USG

[Mue00] Paul R. Mueller, S.J. An unblemished success: Galileo’s sunspot argument in the Dialogue. Journal for the History of Astronomy, 31(4):279–299, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mulholland:1980:BRP

[Mul80] J. Derral Mulholland. Book review: Planets and their inhabitants, planets and planetarians: A History of Theories of the Origin of Planetary Systems. Journal for the History of Astronomy, 11(1):68–69, February 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Muller:1981:BRO

[Mul81]  Peter Muller.  Book re-

view: Observatories illustrated: Sternwarten der Welt. Journal for the History of Astronomy, 12(1):72–73, February 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mulaomerovic:1982:BRB

[Mul82] Jasminko Mulaomerović. Book review: Building in Time: Number from the Light — The Old Croatian Church of the Holy Cross at Nin. Journal for the History of Astronomy, 13(3):213–214, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mulholland:1983:BRA

[Mul83] J. Derral Mulholland. Book review: Archaeoastronomy in the Americas. Edited by Ray A. Williamson (Ballena Press, Los Altos, California, 1981). Pp.

405. $19.95 (paperback). Journal for the History of Astronomy, 14(5):S73–S75, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mulaomerovic:1984:BRT

[Mul84] Jasminko Mulaomerović. Book review: Ten centuries of astronomy in Croatia (Yugoslavia): Provijest egzaktnih znanosti u Hrvata (The History of Exact Sciences of the Croats). Journal for the History of Astronomy, 15(1):51–52, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Muller:1991:BRM

[Mül91] Peter Müller. Book review:

Major observatories, Astronomical Centers of the World. Journal for the History of Astronomy, 22(2):187–188, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Muller:2010:HCT

[Mül10]  Kathrin Müller.  How to

craft telescopic observation in a book: Hevelius’s Selenographia (1647) and its images. Journal for the History of Astronomy, 41(3):355–379, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Mullen:2020:TMW

[Mul20]  Kane Mullen.  Temporary

measures: Women computers at the Royal Observatory, Greenwich, 1890–1895. Journal for the History of Astronomy, 51(1):88–121, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620901358.

Munns:2013:SSH

[Mun13] David P. D. Munns. A single sky: how an international community forged the science of radio astronomy. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01833-0 (hardcover). xi + 247 pp. LCCN QB475.A25 M86 2013.

Murschel:1995:SFP

[Mur95] Andrea Murschel. The structure and function of Ptolemy’s physical hypotheses of planetary motion. Journal for the History of Astronomy, 26(1): 33–61, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Murray:1998:MTA

[Mur98] William Breen Murray. Models of temporality in archaeoastronomy and rock art studies.

Journal for the History of Astronomy, 29(23):S1–S6, February 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nakayama:1971:BRC

[Nak71]  Shigeru Nakayama.  Book

review: Chinese astronomy: Chugoku no tenmon rekiho. Journal for the History of Astronomy, 2(1):42–43, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nall:2016:PAL

[Nal16] Joshua Nall. Popularizing astronomy in Late Nineteenth-Century America. Journal for the History of Astronomy, 47(4):442–443, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616670025.

Nali:2022:ARC

[Nal22] Pier Franco Nali. Annibale Riccò and the catoptric proof of the Earth’s curvature. Journal for the History of Astronomy, 53(1):88–119, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211049617.

Naohmias:2016:LWA

[Naò16] Joseph Ibn Naòhmias. The Light of the World: Astronomy in al-Andalus, volume 1 of Berkeley series in postclassical Islamic scholarship. University of California Press, Berkeley, CA, USA, 2016. ISBN 0-520-28799-1 (hardcover). xiv + 429 pp. LCCN QB23 .J6713 2016. Edited, translated, and with a commentary by Robert G. Morrison.

Nasim:2013:OHS

[Nas13] Omar W. Nasim. Observing by hand: sketching the nebulae in the Nineteenth Century. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-08437-X (hardcover), 0-226-08440-X (ebook). xii + 304 pp. LCCN QB32 .N37 2013.

Nasim:2020:NA

[Nas20] Omar W. Nasim. The newspaper astronomers. Journal for the History of Astronomy, 51(3):371–372, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620921992.

Nath:2013:SHB

[Nat13] Biman Nath. The story of helium and the birth of astrophysics. Astronomers’ universe. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-5362-3 (paperback), 1-4614-5363-1 (e-book). ISSN 1614-659X. xi + 274 pp. LCCN QD181.H4 N38 2013.

Nauenberg:1998:ERM

[Nau98]  Michael Nauenberg.  Es-

say review: The mathematical principles underlying the Principia revisited: The Key to Newton’s Dynamics: The Kepler Problem and the Principia. J. Bruce Brackenridge (University of California Press, Berkeley, 1995). Pp. x +299. $19.95 (paperback) [MR1658137 (99j:01006)]. Newton’s Principia: The Central Argument. Dana Densmore (Green Lion Press, Santa Fe, 1995). Pp. xii + 425. $42; [MR1658501 (99h:01021)]. Journal for the History of Astronomy, 29(3):286–300, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nauenberg:2001:ERN

[Nau01] Michael Nauenberg. Essay review: Newton and the lunar motion: Newton’s Forgotten Lunar Theory: His Contribution to the Quest for Longitude. Journal for the History of Astronomy, 32(2):162– 168, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nauenberg:2008:ECS

[Nau08] Michael Nauenberg. Edmund C. Stoner and the discovery of the maximum mass of white dwarfs. Journal for the History of Astronomy, 39(3):297– 312, August 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nauenberg:2016:INE

[Nau16] Michael Nauenberg. Isaac Newton and Émilie du Châtelet. Journal for the History of Astronomy, 47(3):349–350, August 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nauenberg:2017:FP

[Nau17]  Michael Nauenberg.  The

French Principia. Journal for the History of Astronomy, 48(1):126–128, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616668758.

Naze:2016:TAW

[Naz16] Yaël Nazé. Tribute to an astronomer: The work of Max Ernst on Wilhelm Tempel. Journal for the History of Astronomy, 47(2):115–135, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Needham:1971:BRC

[Nee71] Joseph Needham. Book review: Chinese astronomy: Cosmos and Computation in Early Chinese Mathematical Astronomy. Journal for the History of Astronomy, 2(1):43–45, February 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Needham:1974:BRJ

[Nee74] Joseph Needham. Book review: The Jesuits in China: Studia Copernicana: Colloquia Copernicana. Journal for the History of Astronomy, 5(3): 204–205, October 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Needell:1991:CIW

[Nee91] Allan A. Needell. The Carnegie institution of Washington and radio astronomy: Prelude to an American national observatory. Journal for the History of Astronomy, 22(1):55–67, February 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neeley:2008:BRS

[Nee08] Kathryn A. Neeley. Book review: Somerville in context; Mary Somerville and the World of Science. Journal for the History of Astronomy, 39(1): 130–131, February 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neugebauer:1981:BRT

[Neu81] O. Neugebauer. Book review:

Translation of Cleomedes’s handbook: Cléomède: Théorie élémentaire. Journal for the History of Astronomy, 12(1): 64–65, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neugebauer:1982:GAM

[Neu82]  O. Neugebauer.  A Greek

arithmetical method for finding oblique ascensions. Journal for the History of Astronomy, 13 (1):19–22, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neufeld:2013:SIA

[Neu13] Michael J. Neufeld, editor. Spacefarers: images of astronauts and cosmonauts in the heroic era of spaceflight, A Smithsonian contribution to knowledge. Smithsonian Institution Scholarly Press, Washington, DC, USA, 2013. ISBN 1-935623-19-2 (hardcover). LCCN P96.A792 S63 2013.

Nevalainen:1996:AEL

[Nev96] Jukka Nevalainen. The accuracy of the ecliptic longitude in Ptolemy’s Mercury model. Journal for the History of Astronomy, 27(2):147–160, May 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Newton:1977:RHS

[New77]  R. R. Newton.  Remarks

on “Hipparchus’s Solar Theory Derived from Lunar Eclipse Observations” by K. P. Moesgaard. Journal for the History of Astronomy, 8(3):200– 203, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Ano76d].

Nicolaidis:1990:APR

[Nic90] Efthymios Nicolaïdis. Astronomy and politics in Russia in the early Stalinist period (1928–1932). Journal for the History of Astronomy, 21(4): 345–351, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neuhauser:2016:TBA

[NKM^{+}16] Ralph Neuhäuser, Paul Kunitzsch, Markus Mugrauer, Daniela Luge, and Rob van Gent. Tycho Brahe, Abu Macshar, and the comet beyond Venus (ninth century A.D.). Journal for the History of Astronomy, 47(2):136–158, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neuhauser:2024:ORR

[NN24] Ralph Neuhäuser and Dagmar L. Neuhäuser. Occultation records in the Royal Frankish Annals for A.D. 807: Knowledge transfer from Arabia to Frankia? Journal for the History of Astronomy, 55(3):364–395, August 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241238731.

North:1976:BRL

[Nor76] J. D. North. Book review: Levi’s astronomical tables: The Astronomical Tables of Levi Ben Gerson. Journal for the History of Astronomy, 7(3): 212–213, October 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1978:BRC

[Nor78] J. D. North. Book review: The Capri Symposium: Reason, Experiment, and Mysticism in the Scientific Revolution. Journal for the History of Astronomy, 9 (1):67–68, February 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1980:BRS

[Nor80] J. D. North. Book review: a Sixteenth Century polymath, Wilhelm Schickard, 1592–1635: Astronom, Geograph, Orientalist, Erfinder der Rechenmaschine. Journal for the History of Astronomy, 11(2):138– 140, June 1980. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1982:ERP

[Nor82]  J. D. North.  Essay re-

view: Poulle on Equatoria: Les Instruments de la Théorie des Planètes selon Ptolémée: Équatoires et Horlogerie Planétaire du XIII e au XVI e siècle. Journal for the History of Astronomy, 13(1): 56–60, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1983:BRF

[Nor83] J. D. North. Book review: a Fourteenth-Century calendar: The Kalendarium of Nicholas of Lynn. Journal for the History of Astronomy, 14(3): 227–228, October 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1985:BRA

[Nor85] J. D. North. Book review: The Alphonsine Tables: Les Tables Alphonsines avec les Canons de Jean de Saxe. Journal for the History of Astronomy, 16(1): 51–52, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1986:BRA

[Nor86] J. D. North. Book review:

Astronomy in medieval Denmark: Petri Philomenae de Dacia et Petri de S. Audomaro: Opera quadrivialia. Pars I: Opera Petri Philomenae; pars II: Opera Petri de S. Audomaro. Journal for the History of Astronomy, 17(3):198– 199, August 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1988:BRK

[Nor88] John North. Book review: The Kennedy Festschrift: From Deferent to Equant: a Volume of Studies in the History of Science in the Ancient and Medieval near East in Honor of E. S. Kennedy. Journal for the History of Astronomy, 19 (1):55–56, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1990:BRM

[Nor90] J. D. North. Book review: The Mariner’s astrolabe, The Mariner’s Astrolabe. Journal for the History of Astronomy, 21(3):298–300, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1991:BRJ

[Nor91a] John North. Book review: a Jesuit in China, Astronoom van de Keizer: Ferdinand Verbiest en zijn Europese Sterrenkunde. Journal for the History of Astronomy, 22(3):245–246, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1991:CBA

[Nor91b] John D. North. Commission 41 at the Buenos Aires General Assembly of the IAU. Journal for the History of Astronomy, 22(4):329–330, November 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1992:BRC

[Nor92a] John North. Book review:

a Christian Arab’s chronology: Abu Shaker’s “Chronography”: a Treatise of the 13th Century on Chronological, Calendrical, and Astronomical Matters, Written by a Christian Arab, Preserved in Ethiopic. Journal for the History of Astronomy, 23(1):69–70, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1992:BRA

[Nor92b] John North. Book review: Arguments over Easter: Cummian’s Letter: De Controversia Paschali. Journal for the History of Astronomy, 23(1): 70–71, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1994:ERG

[Nor94] J. D. North. Essay review: a Gingerich anthology: The Eye of Heaven: Ptolemy, Copernicus, Kepler. Journal for the History of Astronomy, 25(3): 243–244, August 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1996:BRA

[Nor96] John North. Book review: Astrology without mathematics:

Ancient Astrology. Journal for the History of Astronomy, 27 (1):81–82, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1997:BRE

[Nor97] J. D. North. Book review: an encyclopedia of intermediate type: History of Astronomy: an Encyclopedia. Journal for the History of Astronomy, 28(4):346–348, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1998:OP

[Nor98] J. D. North. Olaf Pedersen (1920–1997). Journal for the History of Astronomy, 29 (2):211–214, May 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1999:BRB

[Nor99] J. D. North. Book review: “The Book of Aristotle”: The Liber Aristotilis of Hugo of Santalla. Journal for the History of Astronomy, 30(3):319– 321, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2000:BRF

[Nor00a] J. D. North. Book review: Further Einstein studies: The Expanding Worlds of General Relativity. Journal for the History of Astronomy, 31(4):359–360, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2000:BRP

[Nor00b] J. D. North. Book review: Pedro Alfonso and his times: Estudios sobre Pedro Alfonso de Hiesca. Journal for the History of Astronomy, 31(3):269– 270, August 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2002:BRT

[Nor02] John North. Book review:

a treatise on the astrolabe, Krist’an z Prachatic: Stavba a Uzití Astrolábu. Journal for the History of Astronomy, 33 (3):293–294, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2004:BRS

[Nor04] John North. Book review:

Scrutinizing astrolabes: Renaissance Astrolabes and Their Makers. Journal for the History of Astronomy, 35(4):485– 487, November 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2006:BRM

[Nor06]  John North.  Book re-

view: Mathematical astronomy across cultures: Studies on the Transmission of Medieval Mathematical Astronomy. Journal for the History of Astronomy, 37(1):117–119, February 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2007:BRA

[Nor07] John North. Book review: Astrology in history, Horoscopes and Public Spheres: Essays on the History of Astrology. Journal for the History of Astronomy, 38(3):383–386, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2008:BRM

[Nor08a] John North. Book review: More on the Star of Bethlehem: The Star of the Magi: The Mystery That Heralded the Coming of Christ. Journal for the History of Astronomy, 39 (3):408–409, August 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:2008:BRA

[Nor08b] John D. North. Book review: Astronomy as a model?, Astronomy as a Model for the Sciences in Early Modern Times. Journal for the History of Astronomy, 39(4):532–534, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nothaft:2014:JPR

[Not14a] C. Philipp E. Nothaft. John of Pulchro Rivo and John of Saxony: a mise au point. Journal for the History of Astronomy, 45(2):227–242, May 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nothaft:2014:RAA

[Not14b] C. Philipp E. Nothaft. The reception and application of Arabic science in Twelfth-Century computistics: New evidence from Bavaria. Journal for the History of Astronomy, 45(1): 35–60, February 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nothaft:2015:ADT

[Not15a] C. Philipp E. Nothaft. The astronomical data in the Très Riches Heures and their Fourteenth-Century source. Journal for the History of Astronomy, 46(2):113–129, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nothaft:2015:CAA

[Not15b] C. Philipp E. Nothaft. Critical analysis of the Alfonsine Tables in the Fourteenth Century: The Parisian Expositio Tabularum Alfonsii of 1347. Journal for the History of Astronomy, 46(1):76–99, February 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Nothaft:2019:BRM

[Not19] C. Philipp E. Nothaft. Book review: Medieval astrology practically applied. Journal for the History of Astronomy, 50(2):252–253, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619834390.

Nothaft:2020:LTE

[Not20] C. Philipp E. Nothaft. The Liber Theoreumacie (1214) and the early history of the Quadrans Vetus. Journal for the History of Astronomy, 51(1):51–74, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619885504.

Nothaft:2021:EHM

[Not21] C. Philipp E. Nothaft. Ephemerides in high medieval Europe: The textual evidence. Journal for the History of Astronomy, 52(1):33–52, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620981432.

Nothaft:2023:AOB

[Not23a] C. Philipp E. Nothaft. Astronomical observations in Bologna, Montpellier, and Genoa in the early 14th century: Iohannes de Luna Theutonicus revisited. Journal for the History of Astronomy, 54(3):333–352, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231188734.

Nothaft:2023:NSA

[Not23b] C. Philipp E. Nothaft. A new series on Alfonsine astronomy. Journal for the History of Astronomy, 54(1): 113–118, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221150437.

Nothaft:2024:SEL

[Not24a] C. Philipp E. Nothaft. On some early Latin European measurements of the eccentricity of the solar orbit (1308–1314). Journal for the History of Astronomy, 55(2):143–155, May 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241228260.

Nothaft:2024:WTI

[Not24b] C. Philipp E. Nothaft. Were the tables of Ibn Ish.aq al-Tūnisī known in Paris c. 1300? Journal for the History of Astronomy, 55(4):490– 503, November 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286241269044.

Nell:2014:OGP

[NR14] Erin Nell and Clive Ruggles. The orientations of the Giza Pyramids and associated structures. Journal for the History of Astronomy, 45(3):304– 360, August 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Neuhauser:2016:ARA

[NRKN16] Ralph Neuhäuser, Wafiq Rada, Paul Kunitzsch, and Dagmar L. Neuhäuser. Arabic reports about supernovae 1604 and 1572 in Rawh. al-Rūh. by ^{c}Īsā b. Lut.f Allāh from Yemen. Journal for the History of Astronomy, 47 (4):359–374, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616669894.

Numbers:1972:NHI

[Num72] Ronald L. Numbers. The nebular hypothesis of Isaac Orr. Journal for the History of Astronomy, 3(1):49–51, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Numbers:1973:AKD

[Num73] Ronald L. Numbers. The American Kepler: Daniel Kirkwood and his analogy. Journal for the History of Astronomy, 4 (1):13–21, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

OBoyle:1988:BRD

[O’B88] Cornelius O’Boyle. Book review: Duhem Englished: Medieval Cosmology: Theories of Infinity, Place, Time, Void and the Plurality of Worlds. Journal for the History of Astronomy, 19(1):63–64, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Obrist:2014:CMA

[Obr14] Barbara Obrist. Constellations in the middle ages. Journal for the History of Astronomy, 45 (2):253–256, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Obrist:2019:BRM

[Obr19] B. Obrist. Book review: Making Christian the images of ancient constellations in Carolingian Europe. Journal for the History of Astronomy, 50(4):495–497, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619880412.

ODell:2006:BRT

[O’D06] C. R. O’Dell. Book review:

Tools for optical astronomers:

Giant Telescopes: Astronomical Ambition and the Promise of Technology. Journal for the History of Astronomy, 37 (2):245–247, May 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Oestmann:2010:BRE

[Oes10]  Günther Oestmann.  Book

review: Elector August I’s clock: Die Planetenlaufuhr: Ein Meisterwerk der Astronomie und Technik der Renaissance geschaffen von Eberhard Baldewein, 1563– 1568. Journal for the History of Astronomy, 41(1):129–130, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Oestmann:2015:BRG

[Oes15] Günther Oestmann. Book review: German astronomers and the St. Petersburg Academy of Sciences. Formeln und Sterne, Korrespondenz deutscher Gelehrter mit der Kaiserlichen Akademie der Wissenschaften zu St. Petersburg: Briefe von Johann Wilhelm Andreas Pfaff, Johann Sigismund Gottfried Huth, Wilhelm Struve, Martin Bartels, Magnus Georg Paucker aus der Autographensammlung von Wilhelm Stieda in der Universitätsbibliothek Leipzig. Karin Reich and Elena Roussanova (Aachen: Shaker Verlag, Aachen, 2013). Pp. 435. EUR 62,50. ISBN 978-3-8440-2450-0. Journal for the History of Astronomy, 46(2):236–237, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ogjlvie:2011:BRW

[Ogj11] Marilyn Bailey Ogjlvie. Book review: Women in British astronomy: Women in early British and Irish astronomy: Stars and satellites. Journal for the History of Astronomy, 42 (2):271–273, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Olowin:2004:BRV

[Olo04] Ronald P. Olowin. Book review: Virginia Woolf and astronomy: Virginia Woolf and the Discourse of Science: The Aesthetics of Astronomy. Journal for the History of Astronomy, 35(2):240–242, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Olson:1993:BRC

[Ols93]  Donald W. Olson.  Book

review: Computing lunar eclipses: Canon of Lunar Eclipses 1500 B.C.–A.D. 3000, Computer Software for IBM-PC and Compatibles. Journal for the History of Astronomy, 24(4):303–304, November 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omer:1973:RPG

[Ome73] Guy C. Omer, Jr. A reputed portrait of Galileo by Ribera. Journal for the History of Astronomy, 4(3):194–195, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omer:1979:BRA

[Ome79] Guy C. Omer, Jr. Book review: Astronomers in distant lands, The Whisper and the Vision. Journal for the History of Astronomy, 10(1):66–67, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omodeo:2011:BRF

[Omo11a] Pietro Omodeo. Book review: The foundations of Copernicus’s Book I: Niccolò Copernico e la Fondazione del Cosmo Eliocentrico, con testo, traduzione e commentario del Libro I de Le Rivoluzioni Celesti. Journal for the History of Astronomy, 42(4):527–529, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omodeo:2011:DOP

[Omo11b] Pietro Daniel Omodeo. David Origanus’s planetary system (1599 and 1609). Journal for the History of Astronomy, 42 (4):439–454, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omodeo:2013:BRT

[Omo13] Pietro Daniel Omodeo. Book review: Telesio on comets and the Milky Way, Bernardino Telesio: Sobre los Cometas y la Vía Láctea (edición Bilingüe). Journal for the History of Astronomy, 44(2):216–217, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omodeo:2017:SCB

[Omo17] Pietro Daniel Omodeo. The scientific culture of the Baltic mathematician, physician, and calendar-maker Laurentius Eichstadt (1596–1660). Journal for the History of Astronomy, 48(2):135–159, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617703847.

Omodeo:2019:BRA

[Omo19] Pietro Daniel Omodeo. Book review: an apology for Giordano Bruno. Journal for the History of Astronomy, 50 (2):253–255, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619834392.

Omodeo:2020:ISA

[Omo20] Pietro Daniel Omodeo. Ideas on stellar agency and life in Early Modern Europe. Journal for the History of Astronomy, 51(4):486–488, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620967448.

Oosterhoff:2015:BRN

[Oos15] Richard J. Oosterhoff. Book review: New sources for history of early modern astronomy. Celestial Novelties on the Eve of the Scientific Revolution, 1540–1630. Edited by Dario Tessicini and Patrick J. Boner (Biblioteca di Galilaeana, iii: Leo S. Olschki Editore, Florence, 2013). Pp. xvi + 282. EUR 32. ISBN 978-88-222-6254-7 (paper). Journal for the History of Astronomy, 46 (2):245–247, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Olson:1989:CPH

[OP89] Roberta J. M. Olson and Jay M. Pasachoff. Is comet P/Halley of A.D. 684 recorded in the Nuremberg Chronicle? Journal for the History of Astronomy, 20(3):171–174, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Olson:1992:SEC

[OP92] Roberta J. M. Olson and Jay M. Pasachoff. The 1816 solar eclipse and Comet 1811 I in John Linnell’s astronomical album. Journal for the History of Astronomy, 23(2):121– 133, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Orchiston:1984:BRL

[Orc84] Wayne Orchiston. Book review: Last years of Sydney Observatory: Sydney Observatory 1958 to 1981. Journal for the History of Astronomy, 15 (2):142–143, June 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Orchiston:1988:BRB

[Orc88] Wayne Orchiston. Book review: Biography of John Tebbutt: Astronomical Memoirs. Journal for the History of Astronomy, 19(1):69–70, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Orchiston:1990:BRS

[Orc90] Wayne Orchiston. Book review: Southern hemisphere astronomy, under Capricorn: A History of Southern Hemisphere Astronomy. Journal for the History of Astronomy, 21 (3):304–306, August 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osler:2006:BRS

[Osl06] Margaret J. Osler. Book review: Science and religion over two millennia: Science and Religion, 400 B.C.–A.D.

1550: From Aristotle to Copernicus. Journal for the History of Astronomy, 37(4):486–487, November 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ossendrijver:2012:BRN

[Oss12] Mathieu Ossendrijver. Book review: New light on MUL.APIN: Writing Science before the Greeks: a Naturalistic Analysis of the Babylonian Astronomical Treatise MUL.APIN. Journal for the History of Astronomy, 43(4):497–498, November 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ossendrijver:2018:AAS

[Oss18] Mathieu Ossendrijver. Ancient astral sciences. Journal for the History of Astronomy, 49(1):125–129, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618755447.

Ossendrijver:2024:BBA

[Oss24] Mathieu Ossendrijver. BM 47886+47914, a Babylonian astral compendium with possible implications for the origin of the “year of the Sun”. Journal for the History of Astronomy, 55(3):350–363, August 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241233769.

Osterbrock:1980:BRA

[Ost80] Donald E. Osterbrock. Book review: Astronomy in the Twentieth Century, a Source Book in Astronomy and Astrophysics, 1900–1975. Journal for the History of Astronomy, 11(3):215–217, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1984:MLS

[Ost84a] Donald E. Osterbrock. Mary Lea Shane (1897–1983). Journal for the History of Astronomy, 15(1):74, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1984:RFEa

[Ost84b] Donald E. Osterbrock. The rise and fall of Edward S. Holden: Part 1. Journal for the History of Astronomy, 15(2):81– 127, June 1984. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1984:RFEb

[Ost84c] Donald E. Osterbrock. The rise and fall of Edward S. Holden: Part 2. Journal for the History of Astronomy, 15(3):151–176, October 1984. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1986:FST

[Ost86] Donald E. Osterbrock. Failure and success: Two early experiments with concave gratings in stellar spectroscopy. Journal for the History of Astronomy, 17(2):119–129, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1989:CHM

[Ost89] Donald E. Osterbrock. To climb the highest mountain: W. W. Campbell’s 1909 Mars Expedition to Mount Whitney. Journal for the History of Astronomy, 20(2):77–97, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1992:BRH

[Ost92a] Donald E. Osterbrock. Book review: The history of cometary astronomy: Comets: a Chronological History of Observation, Myth and Folklore. Journal for the History of Astronomy, 23(3):217–218, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1992:APD

[Ost92b] Donald E. Osterbrock. “The Appointment of a Physicist as Director of the Astronomical Center of the World”. Journal for the History of Astronomy, 23(3):155–165, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1994:GPW

[Ost94] Donald E. Osterbrock. Getting the picture: Wide-field astronomical photography from Barnard to the achromatic Schmidt, 1888–1992. Journal for the History of Astronomy, 25(1):1–14, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1995:WBO

[Ost95] Donald E. Osterbrock. Walter Baade, observational astrophysicist, (1): The preparation 1893–1931. Journal for the History of Astronomy, 26(1):1–32, February 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1996:WBO

[Ost96] Donald E. Osterbrock. Walter Baade, obervational astrophysicist, (2): Mount Wilson 1931–1947. Journal for the History of Astronomy, 27(4): 301–348, November 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1997:BRP

[Ost97a] Donald E. Osterbrock. Book review: Papers by Spitzer: Dreams, Stars, and Electrons: Selected Writings of Lyman Spitzer, Jr. Journal for the History of Astronomy, 28(3):276– 277, August 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1997:WBO

[Ost97b] Donald E. Osterbrock. Walter Baade, observational astrophysicist, (3): Palomar and Göttingen 1948–1960 (Part A). Journal for the History of Astronomy, 28(4):283–316, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:1998:WBO

[Ost98] Donald E. Osterbrock. Walter Baade, observational astrophysicist, (3): Palomar and Göttingen 1948–1960 (Part B). Journal for the History of Astronomy, 29(4):345–377, November 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2001:HZD

[Ost01] Donald E. Osterbrock. Herman Zanstra, Donald H. Menzel, and the Zanstra method of nebular astrophysics. Journal for the History of Astronomy, 32(2):93–108, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2002:YDM

[Ost02] Donald E. Osterbrock. Young Don Menzel’s amazing adventures at Lick Observatory. Journal for the History of Astronomy, 33(2):95–118, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2004:BRC

[Ost04] Donald E. Osterbrock. Book review: Chronicling the U.S. Naval Observatory: Sky and Ocean Joined: The U.S. Naval Observatory, 1830–2000. Journal for the History of Astronomy, 35(2):237–238, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2005:BRD

[Ost05a] Donald E. Osterbrock. Book review: a decade of nuclear physics: From Nuclear Transformation to Nuclear Fission, 1932–1939. Journal for the History of Astronomy, 36(2): 245–246, May 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2005:BRB

[Ost05b] Donald E. Osterbrock. Book review: Bernhard Schmidt: Den Himmel fest im Blick: Eine wissenschaftliche Biografie über dem Astro-Optiker Bernhard Schmidt. Journal for the History of Astronomy, 36(1):118–120, February 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Osterbrock:2007:FRH

[Ost07] Donald E. Osterbrock. Frank Ross, his Ross lens design, and the Lick Observatory 20-inch astrograph. Journal for the History of Astronomy, 38(1): 31–73, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Omodeo:2013:PCR

[OT13] Pietro Daniel Omodeo and Irina Tupikova. The post-Copernican reception of Ptolemy: Erasmus Reinhold’s commented edition of the Almagest, Book One (Wittenberg, 1549). Journal for the History of Astronomy, 44(3): 235–256, August 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Palter:1977:BRH

[Pal77]  Robert Palter.  Book re-

view: Hanson on Early Astronomy, Constellations and Conjectures. Journal for the History of Astronomy, 8(2):140– 142, June 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Palter:1984:BRC

[Pal84] Robert Palter. Book review: Cosmology in outline: Frame of the Universe. Journal for the History of Astronomy, 15 (2):143–144, June 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Palmieri:2001:GDP

[Pal01] Paolo Palmieri. Galileo and the discovery of the phases of Venus. Journal for the History of Astronomy, 32(2):109– 129, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Palmieri:2008:GDN

[Pal08] Paolo Palmieri. Galileus Deceptus, non minime decepit: a re-appraisal of a counterargument in dialogo to the extrusion effect of a rotating Earth. Journal for the History of Astronomy, 39(4):425–452, November 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Palmer:2012:BRI

[Pal12]  James Palmer.  Book re-

view: Irish computistics: The Munich Computus: Text and Translation. Irish Computistics between Isidore of Seville and the Venerable Bede and its Reception in Carolingian Times. Journal for the History of Astronomy, 43(1):129–131, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Panchenko:1994:TPS

[Pan94a] Dmitri Panchenko. Thales’s prediction of a solar eclipse. Journal for the History of Astronomy, 25(4):275–288, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pang:1994:VOP

[Pan94b] Alex Soojung-Kim Pang. Victorian observing practices, printing technology, and representations of the solar corona, (1): The 1860s and 1870s. Journal for the History of Astronomy, 25(4):249–274, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pang:1995:VOP

[Pan95] Alex Soojung-Kim Pang. Victorian observing practices, printing technology, and representations of the solar corona, (2): The age of photomechanical reproduction. Journal for the History of Astronomy, 26 (1):63–75, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pankenier:2009:BRR

[Pan09a] David W. Pankenier. Book review: The 1280 reform of Chinese astronomy: Granting the Seasons: The Chinese Astronomical Reform of 1280, with a Study of its Many Dimensions and an Annotated Translation of its Records. Journal for the History of Astronomy, 40(4):476–477, November 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pantin:2009:BRB

[Pan09b] Isabelle Pantin. Book review: Bruno’s astronomy, I Dintorni dell’Infinito: Giordano Bruno e l’Astronomia del Cinquecento. Journal for the History of Astronomy, 40(3): 353–355, August 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pantin:2010:ADO

[Pan10] Isabelle Pantin. The astronomical diagrams in oronce Finé’s protomathesis (1532): Founding a French tradition? Journal for the History of Astronomy, 41(3):287–310, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pantin:2012:FPT

[Pan12] Isabelle Pantin. The first phases of the Theoricæ planetarum printed tradition (1474– 1535): The evolution of a genre observed through its images. Journal for the History of Astronomy, 43(1):3–26, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Panza:2015:BRN

[Pan15] Marco Panza. Book review:

Newton and the origin of civilization. Journal for the History of Astronomy, 46(2):252– 254, May 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Papathanassiou:1991:AEP

[Pap91] Maria Papathanassiou. On the astronomical explanation of Phanes’s Relief at Modena. Journal for the History of Astronomy, 22(16):S1–S13, February 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Parisot:1985:ROC

[Par85] Jean-Paul Parisot. La rétrogradation de l’ombre dans les cadrans solaires analemmatiques. (French) [Downshift of shade in Analemmatic sundials]. Journal for the History of Astronomy, 16(1):43–48, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Park:1986:BRS

[Par86] Katharine Park. Book review: Science and magic: Occult and Scientific Mentalities in the Renaissance. Journal for the History of Astronomy, 17 (3):200–202, August 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Park:1990:BRP

[Par90] David Park. Book review: A Pocket Encyclopedia of Time, Time in History. Journal for the History of Astronomy, 21 (3):297–298, August 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Parrinder:1992:BRO

[Par92] Patrick Parrinder. Book review: Other worlds: The Last Frontier: Imagining Other Worlds from the Copernican Revolution to Modern Science Fiction. Journal for the History of Astronomy, 23(2):150– 151, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2002:ME

[Pas02] Jay M. Pasachoff. Menzel and eclipses. Journal for the History of Astronomy, 33(2):139– 156, May 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2003:BRB

[Pas03a] Jay M. Pasachoff. Book review:

Bedside reading: Setting Sail for the Universe: Astronomers and Their Discoveries. Journal for the History of Astronomy, 34(2):238–239, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2003:BRS

[Pas03b] Jay M. Pasachoff. Book review:

The spots on the Sun: The Enigma of Sunspots: a Story of Discovery and Scientific Revolution. Journal for the History of Astronomy, 34(4):458–459, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2014:BRH

[Pas14] Jay M. Pasachoff. Book review: a history of telescopes: 400 Years of Astronomical Telescopes: A Review of History, Science and Technology, ISBN 978-90-481-2232-5. Journal for the History of Astronomy, 45 (2):256–258, May 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2015:SMM

[Pas15]  Jay M. Pasachoff.  Simon

Marius’s Mundus Iovialis: 400th anniversary in Galileo’s shadow. Journal for the History of Astronomy, 46(2):218– 234, May 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1978:BRN

[Pau78] E. Robert Paul. Book review: The nature of the nebulae, Man Discovers the Galaxies. Journal for the History of Astronomy, 9(3):222–224, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1981:DRP

[Pau81] E. Robert Paul. The death of a research programme: Kapteyn and the Dutch astronomical community. Journal for the History of Astronomy, 12(2): 77–94, June 1981. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1982:BRF

[Pau82] E. Robert Paul. Book review:

Festschrift for Oort: Oort and the Universe. Journal for the History of Astronomy, 13(2): 141–142, June 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1986:JCK

[Pau86] E. Robert Paul. J. C. Kapteyn and the early Twentieth-Century universe. Journal for the History of Astronomy, 17 (3):155–182, August 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1988:BRO

[Pau88] Robert Paul. Book review:

The origins of Lick Observatory: James Lick’s Monument: The Saga of Captain Richard Floyd and the Building of the Lick Observatory. Journal for the History of Astronomy, 19 (3):212–213, August 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Paul:1995:BRC

[Pau95] E. Robert Paul. Book review: The Carnegie institution and astronomy: The Earth, the Heavens and the Carnegie Institution of Washington. Journal for the History of Astronomy, 26(2):183–184, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Peterson:1987:AOO

[PC87] Cynthia W. Peterson and Bella C. Chiu. On the astronomical origin of the offset street grid at Teotihuacan. Journal for the History of Astronomy, 18(11):S13–S18, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Purrington:1989:PPR

[PC89] Robert D. Purrington and Colby Allan Child, Jr. Poverty point revisited: Further consideration of astronomical alignments. Journal for the History of Astronomy, 20(13):S49–S60, February 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Porres:2001:JMT

[PC01] Beatriz Porres and José Chabás. John of Murs’s Tabulae permanentes for finding true syzygies. Journal for the History of Astronomy, 32(1):63–72, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Porter:1985:ERV

[PdV85] Roy Porter and G. de Vaucouleurs. Essay review: Volume 4A of The General History: Astrophysics and Twentieth-Century Astronomy to 1950. Journal for the History of Astronomy, 16(3):224–231, October 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1973:BRI

[Ped73] Olaf Pedersen. Book review:

an introduction to Copernicus: The Background to Copernicus. The Copernican Revolution. Journal for the History of Astronomy, 4(2):139– 140, June 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1979:BRA

[Ped79] Olaf Pedersen. Book review:

Albertus (?) on astronomy, Alberto Magno: Speculum Astronomiae. Journal for the History of Astronomy, 10(2):126– 128, June 1979. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1981:OTP

[Ped81] Olaf Pedersen. The origins of the “Theorica Planetarum”. Journal for the History of Astronomy, 12(2):113–123, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1983:GCT

[Ped83] Olaf Pedersen. Galileo and the Council of Trent: The Galileo Affair revisited. Journal for the History of Astronomy, 14 (1):1–29, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1985:QS

[Ped85] Olaf Pedersen. In quest of Sacrobosco. Journal for the History of Astronomy, 16(3): 175–221, October 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:1987:ERA

[Ped87] Olaf Pedersen. Essay review: The Almagest in translation. Journal for the History of Astronomy, 18(1):59–63, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pedersen:2015:LAD

[Ped15] Fritz S. Pedersen. Latin authors discuss the Jewish calendar. Journal for the History of Astronomy, 46(4):491–493, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Peoux:2008:BRT

[Péo08] Gérald Péoux. Book review: Tycho and his correspondence, Bearing the Heavens: Tycho Brahe and the Astronomical Community of the Late Sixteenth Century. Journal for the History of Astronomy, 39(4): 528–529, November 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Perkins:2000:BRF

[Per00]  Adam Perkins.  Book re-

view: a Flamsteed biography:

John Flamsteed: The First Astronomer Royal at Greenwich. Journal for the History of Astronomy, 31(3):274–276, August 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Petri:1980:BRT

[Pet80]  Winfried Petri.  Book re-

view: Tibetan calendars, Untersuchungen zur Geschichte der tibetischen Kalenderrechnung. Journal for the History of Astronomy, 11(1):69– 72, February 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Peters:1984:AVM

[Pet84] William T. Peters. The appearances of Venus and Mars in 1610. Journal for the History of Astronomy, 15(3):211–214, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Peterson:2004:EAM

[Pet04] Charles J. Peterson. The education of an astronomical maverick: T. J. J. See and the University of Missouri. Journal for the History of Astronomy, 35 (3):293–304, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Petrovay:2020:DST

[Pet20] Kristóf Petrovay. The determination of stellar temperatures from Baron B. Harkányi to the Gaia Mission. Journal for the History of Astronomy, 51(2):152–161, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620918961.

Patrick:1983:RSC

[PF83] J. Patrick and P. R. Freeman. Revised surveys of Cork–Kerry stone circles. Journal for the History of Astronomy, 14(5): S50–S56, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Payne-Gaposchkin:1972:ERM

[PG72] Cecilia Payne-Gaposchkin. Essay review: Myth and science: Hamlet’s Mill. Journal for the History of Astronomy, 3(3): 206–211, October 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Payne-Gaposchkin:1979:BRE

[PG79] Cecilia Payne-Gaposchkin. Book review: Early Records of Supernovae, the Historical Supernovae. Journal for the History of Astronomy, 10(1):47– 50, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Papathanassiou:1994:OGT

[PH94] Maria Papathanassiou and Michael Hoskin. Orientations of the Greek temples on Corfu.

Journal for the History of Astronomy, 25(2):111–114, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Papathanassiou:1996:LMC

[PH96] Maria Papathanassiou and Michael Hoskin. The late-Minoan cemetery at Armenoi: a reappraisal. Journal for the History of Astronomy, 27(1): 53–59, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Philip:1991:BRK

[Phi91] A. G. Davis Philip. Book review: Kopal’s autobiography, Of Stars and Men: Reminiscences of an Astronomer. Journal for the History of Astronomy, 22(2):191–192, May 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Papathanassiou:1992:OTL

[PHP92] Maria Papathanassiou, Michael Hoskin, and Helen Papadopoulou. Orientations of tombs in the Late-Minoan cemetery at Armenoi, Crete. Journal for the History of Astronomy, 23(17): S43–S55, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addenda [PHP93].

Papathanassiou:1993:OTA

[PHP93] Maria Papathanassiou, Michael Hoskin, and Helen Papadopoulou.

Orientations of tombs at Armenoi, Crete: Addenda. Journal for the History of Astronomy, 24(18):S54, February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [PHP92].

Pingree:1970:CIP

[Pin70] David Pingree. On the classification of Indian planetary tables. Journal for the History of Astronomy, 1(2):95–108, August 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1971:BRA

[Pin71a]  David Pingree.  Book re-

view: an astronomical treatise of Brahmagupta: The Khaán.d.akhâdyaka (An Astronomical Treatise) of Brahmagupta with the Commentary of Bhat.t.otpala. Journal for the History of Astronomy, 2 (2):121–122, June 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1971:GOI

[Pin71b] David Pingree. On the Greek origin of the Indian planetary model employing a double epicycle. Journal for the History of Astronomy, 2(2):80–85, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1972:PTI

[Pin72] David Pingree. Precession and trepidation in Indian astronomy before A.D. 1200. Journal for the History of Astronomy, 3 (1):27–35, February 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1973:BRA

[Pin73a] David Pingree. Book review:

Astrology in antiquity: Origins of Astrology. Journal for the History of Astronomy, 4(1):59, February 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1973:MOE

[Pin73b] David Pingree. The Mesopotamian origin of early Indian mathematical astronomy. Journal for the History of Astronomy, 4 (1):1–12, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1974:BRM

[Pin74]  David Pingree.  Book re-

view: Mayan Astronomy: a Commentary on the Dresden Codex: a Maya Hieroglyphic Book. Journal for the History of Astronomy, 5(2):137– 138, June 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1976:REG

[Pin76] David Pingree. The recovery of early Greek astronomy from India. Journal for the History of Astronomy, 7(2):109– 123, June 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1980:BRE

[Pin80] David Pingree. Book review: Edition of a commentary by Theon, Le “Petit Commentaire” de Théon d’Alexandrie aux Tables Faciles de Ptolémée. Journal for the History of Astronomy, 11(2):137–138, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1981:BRL

[Pin81]  David Pingree.  Book re-

view: The “little astronomy” of Autolycos: La sphère en mouvement. Levers et couchers héliaques. Testimonia. Journal for the History of Astronomy, 12(2):148–149, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1984:BRH

[Pin84] David Pingree. Book review: Hindu astronomy: The Súrya Siddhánta, or An Ancient System of Hindu Astronomy. With the Exposition of Ranganátha, the Gúd.hártha-praká’saka, the Súrya Siddhánta, or An Ancient System of Hindu Astronomy. Followed by the Siddhánta Siromani. Journal for the History of Astronomy, 15(1): 47–48, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1988:BRS

[Pin88] David Pingree. Book review:

Star names and their etymologies: Short Guide to Modern Star Names and Their Derivatives. Journal for the History of Astronomy, 19(1):59, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1990:BRA

[Pin90]  David Pingree.  Book re-

view: Astrology in the West, A History of Western Astrology. Journal for the History of Astronomy, 21(2):229–230, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1992:BRS

[Pin92]  David Pingree.  Book re-

view: Seneca’s natural questions: Senecas Naturales Quaestiones: Komposition, Naturphilosophische Aussagen und Ihre Quellen. Journal for the History of Astronomy, 23 (2):146–147, May 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1996:BCI

[Pin96] David Pingree. Bīja-corrections in Indian astronomy. Journal for the History of Astronomy, 27(2):161–172, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:1998:SFC

[Pin98] David Pingree. Some Fourteenth-Century Byzantine astronomical texts. Journal for the History of Astronomy, 29(2):103– 108, May 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pingree:2003:ERP

[Pin03] David Pingree. Essay review: Ptolemy’s geographical guide explored: Ptolemy’s Geography. An Annotated Translation of the Theoretical Chapters. Journal for the History of Astronomy, 34(2):235–237, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Platt:1991:ERA

[Pla91] Tristan Platt. Essay review: The anthropology of astronomy: World archaeoastronomy. Journal for the History of Astronomy, 22(16):S76–S83, February 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Plotkin:1977:HDD

[Plo77]  Howard Plotkin.  Henry

Draper, the discovery of oxygen in the Sun, and the dilemma of interpreting the solar spectrum. Journal for the History of Astronomy, 8(1):44–51, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Plotkin:1990:ECP

[Plo90]  Howard Plotkin.  Edward

Charles Pickering. Journal for the History of Astronomy, 21 (1):47–58, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Plotkin:1993:WHP

[Plo93] Howard Plotkin. William H.

Pickering in Jamaica: The founding of Woodlawn and studies of Mars. Journal for the History of Astronomy, 24(1–2): 101–122, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/101.full.pdf+html.

Plofker:2000:AST

[Plo00] Kim Plofker. The astrolabe and spherical trigonometry in medieval India. Journal for the History of Astronomy, 31(1): 37–54, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Plofker:2016:MCS

[Plo16a]  Kim Plofker.  Mathemat-

ics in the classical Sanskrit sciences. Journal for the History of Astronomy, 47 (4):434–436, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616668744.

Plotkin:2016:PRG

[Plo16b] Howard Plotkin. Planetary research in Germany. Journal for the History of Astronomy, 47(1):106–107, February 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/47/1/
                                        106.full.pdf+html.

Pingree:1989:IYI

[PM89] David Pingree and Patrick Morrissey. On the identification of the yogatārās of the Indian naku.satras. Journal for the History of Astronomy, 20 (2):99–119, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:1996:ECP

[POH96] Jay M. Pasachoff, Roberta J. M. Olson, and Martha L. Hazen. The earliest comet photographs: Usherwood, Bond, and Donati 1858. Journal for the History of Astronomy, 27 (2):129–145, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Popp:2022:EAK

[Pop22] Bruce D. Popp. Early application of kinetic theory of gases to star clusters. Journal for the History of Astronomy, 53(2):141–162, May 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221078338.

Porteous:1973:MYM

[Por73] Hugh L. Porteous. Megalithic yard or megalithic myth? Journal for the History of Astronomy, 4(1):22–24, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Porter:1974:NDE

[Por74] N. A. Porter. The nova of A.D. 1006 in European and Arab records. Journal for the History of Astronomy, 5(2):99– 104, June 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Portuondo:2009:LEL

[Por09] María M. Portuondo. Lunar eclipses, longitude and the new world. Journal for the History of Astronomy, 40(3):249– 276, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1975:ERS

[Pou75] Emmanuel Poulle. Essay review: Science in the Middle Ages: A Source Book in Medieval Science. Journal for the History of Astronomy, 6(3):

212–217, October 1975. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1979:ERJ

[Pou79] Emmanuel Poulle. Essay review: Jordanus on the astrolabe: Jordanus de Nemore and the Mathematics of Astrolabes: De plana spera. Journal for the History of Astronomy, 10(3): 185–188, October 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1983:LPC

[Pou83] Emmanuel Poulle. L’horloge planétaire de la cathédrale de Strasbourg. (French) [The planetary clock of the Cathedral of Strasbourg]. Journal for the History of Astronomy, 14(1):33–46, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano83b].

Poulle:1984:ERT

[Pou84] Emmanuel Poulle. Essay review: Two medieval texts: Medieval mensurations: “Quadrans vetus” and “Geometrie Due sunt partes principales”. Journal for the History of Astronomy, 15(2):135–138, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1985:ERA

[Pou85] Emmanuel Poulle. Essay review: Astronomy in Cracow: Scientific Writings and Astronomical Tables in Cracow: A Census of Manuscript Sources. Journal for the History of Astronomy, 16(2):134–138, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1988:ATA

[Pou88] Emmanuel Poulle. The Alfonsine tables and Alfonso X of Castille. Journal for the History of Astronomy, 19(2):97– 113, May 1988. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1996:ERC

[Pou96]  Emmanuel Poulle.  Es-

say review: Copernicus’s minor works in facsimile: Nicholas Copernicus, Complete Works, IV: The Manuscripts of Nicholas Copernicus’ Minor Works Facsimiles, Nicolas Copernic, OEuvres complètes, II (Version française): Facsimilés des manuscrits des écrits mineurs. Journal for the History of Astronomy, 27(1): 75–78, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:2009:BRA

[Pou09] Emmanuel Poulle. Book review: The astrolabe in the Latin West, Aspects of the Astrolabe: ‘Architectonica Ratio’ in Tenth- and Eleventh-Century Europe. Journal for the History of Astronomy, 40 (2):240–242, May 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Powell:1988:JHL

[Pow88] Corey Stevenson Powell. J. Homer Lane and the internal structure of the Sun. Journal for the History of Astronomy, 19(3):183–199, August 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Powell:1995:AIB

[Pow95] Martin J. Powell. Astronomical indications at a bell-barrow in South Wales. Journal for the History of Astronomy, 26(20): S49–S56, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Parisot:1982:OAT

[PP82] J. P. Parisot and P. Petrequin. Orientation astronomique des tombes mérovingiennes du cimétiére de Soyria (Jura). (French) [Astronomical orientation of Merovingian tombs in the cemetery of Soyria (Jura)]. Journal for the History of Astronomy, 13(4):S41–S48, February 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pasachoff:2007:BRL

[PP07] Jay M. Pasachoff and Naomi Pasachoff. Book review: Linking Earth and Sun, The Sun Kings: The Unexpected Tragedy of Richard Carrington and the Tale of How Modern Astronomy Began. Journal for the History of Astronomy, 38 (4):524–525, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Proverbio:1987:AOF

[PRA87] Edoardo Proverbio, Giuliano Romano, and Anthony Aveni. Astronomical orientations of five megalithic tombs at Madau, near Fonni in Sardinia. Journal for the History of Astronomy, 18(11):S55–S65, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Price:1991:UAT

[Pri91] B. B. Price. The use of astronomical tables by Albertus Magnus. Journal for the History of Astronomy, 22(3):221– 240, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Przybylski:1979:BRA

[Prz79] A. Przybylski. Book review:

Astronomy in Poland, Historia Astronomii w Polsce (History of Astronomy in Poland). Journal for the History of Astronomy, 10(1):58–61, February 1979. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Poulle:1998:SAA

[PS98] Emmanuel Poulle and Denis Savoie. La survie de l’astronomie alphonsine. (French) [The survival of Alphonsine astronomy]. Journal for the History of Astronomy, 29(2):201– 207, May 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Pumfrey:2011:SWG

[Pum11] Stephen Pumfrey. The selenographia of William Gilbert: His pre-telescopic map of the Moon and his discovery of lunar libration. Journal for the History of Astronomy, 42 (2):193–203, May 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Quan:2017:CAS

[Qua17] Tang Quan. Chinese astronomical systems. Essay review: The Foundations of Celestial Reckoning: Three Ancient Chinese Astronomical Systems. Christopher Cullen (Routledge, New York, 2017). Pp. xi + 434. £84. ISBN 978-1-138-10117-3. Journal for the History of Astronomy, 48(3):371–372, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617713526.

Rodriguez-Arribas:2018:TCA

[RA18] Josefina Rodríguez-Arribas.

A treatise on the construction of astrolabes by Jacob ben Abi Abraham Isaac al-Corsuno (Barcelona, 1378): Edition, translation and commentary. Journal for the History of Astronomy, 49 (1):27–82, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618755773.

Rabin:2005:WKS

[Rab05] Sheila J. Rabin. Was Kepler’s species immateriata substantial? Journal for the History of Astronomy, 36(1): 49–56, February 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Radoslavova:1990:BRR

[Rad90] Tsvetanka Radoslavova. Book review: Records in Stone: Papers in Memory of Alexander Thom. Edited by C. L. N. Ruggles (Cambridge University Press, Cambridge, 1988). Pp. xvii + 519. £50/$100. Journal for the History of Astronomy, 21(15):S62–S66, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:1983:BRB

[Rag83] F. Jamil Ragep. Book review:

Biruni remembered: Al-Biruni Commemorative Volume. Journal for the History of Astronomy, 14(2):144–146, June 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:1985:BRM

[Rag85] F. Jamil Ragep. Book review: a Muslim cosmology: Islamic Cosmology: a Study of as-Suyuti’s al-Hay’a as-saniya fi l’hay’a as-Sunniya. Journal for the History of Astronomy, 16(2):146–149, June 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:1992:ERT

[Rag92] F. Jamil Ragep. Essay review: Thabit’s astronomical works: Thabit Ibn Qurra: Oeuvres d’Astronomie. Journal for the History of Astronomy, 23 (1):61–63, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:2001:BRQ

[Rag01] F. Jamil Ragep. Book review: The Qibla Revisited: Worldmaps for Finding the Direction and Distance to Mecca: Innovation and Tradition in Islamic Science. Journal for the History of Astronomy, 32(2):171– 172, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:2005:IRE

[Rag05] F. Jamil Ragep. ^{c}Alīqushjī and Regiomontanus: Eccentric transformations and Copernican revolutions. Journal for the History of Astronomy, 36(4): 359–371, November 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:2009:BRA

[Rag09] F. Jamil Ragep. Book review: Astronomy in the Middle East, A Brief Introduction to Astronomy in the Middle East. Journal for the History of Astronomy, 40(3):346–347, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ragep:2016:EC

[Rag16a] F. Jamil Ragep. Egyptian curiosities. Journal for the History of Astronomy, 47 (4):428–429, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616665946.

Ragep:2016:IAS

[Rag16b] F. Jamil Ragep. Ibn al-Shāt.ir and Copernicus: The Uppsala notes revisited. Journal for the History of Astronomy, 47(4):395–415, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616678508.

Rodriguez-Anton:2018:ARU

[RAGGB18] Andrea Rodríguez-Antón, Antonio César González-García, and Juan Antonio Belmonte. Astronomy in Roman urbanism: a statistical analysis of the orientation of Roman towns in the Iberian Peninsula. Journal for the History of Astronomy, 49(3):363–387, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ramberg:2006:BRN

[Ram06] Peter J. Ramberg. Book review: NASA and astrobiology: The Living Universe: NASA and the Development of Astrobiology. Journal for the History of Astronomy, 37(3):365– 366, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ramsey:2007:CGR

[Ram07] John T. Ramsey. A catalogue of Greco–Roman comets from 500 B.C. to A.D. 400. Journal for the History of Astronomy, 38(2):175–197, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ramasubramanian:2010:BRE

[Ram10] K. Ramasubramanian. Book review: Early Indian astronomical instruments: The Archaic and the Exotic: Studies in the History of Indian Astronomical Instruments. Journal for the History of Astronomy, 41(1): 126–127, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ramberg:2016:PAG

[Ram16] Peter J. Ramberg. Popularizing astronomy in the German free religious movement, 1851–1852. Journal for the History of Astronomy, 47(1): 30–51, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
30.full.pdf+html.

Raphael:2011:NAI

[Rap11a]  Renée Raphael.  A non-

astronomical image in an astronomical text: Visualizing motion in Riccioli’s Almagestum Novum. Journal for the History of Astronomy, 42(1): 73–90, February 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Raphael:2011:BRO

[Rap11b] Renée J. Raphael. Book review: Oronce Fine considered: The Worlds of Oronce Fine: Mathematics, Instruments, and Print in Renaissance France. Journal for the History of Astronomy, 42(3):408–409, August 2011. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Raposo:2011:BRE

[Rap11c] Pedro M. P. Raposo. Book review: an 1870 eclipse expedition: L’Éclisse Totale di Sole del 1870 in Sicilia: Lettere di Pietro Tacchini a Gaetano Cacciatore. Journal for the History of Astronomy, 42(2):275– 276, May 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Raphael:2013:BRG

[Rap13] Renée Raphael. Book review: Galileo selections, Galileo, Selected Writings. Journal for the History of Astronomy, 44 (2):225–226, May 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Raphael:2015:BRG

[Rap15a] Renee Raphael. Book review: The global Galileo: Galileo’s Idol: Gianfrancesco Sagredo & the Politics of Knowledge. Nick Wilding (University of Chicago Press, Chicago, IL, 2014). Pp. 200. $35. ISBN 978-0-226-16697-1. Journal for the History of Astronomy, 46(3): 378–379, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Raphael:2015:CCJ

[Rap15b] Renée J. Raphael. Copernicanism in the classroom: Jesuit natural philosophy and mathematics after 1633. Journal for the History of Astronomy, 46(4):419–440, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Raposo:2018:MTI

[Rap18] Pedro M. P. Raposo. Merz telescopes in Italy. Journal for the History of Astronomy, 49(2):265–266, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618773232.

Raposo:2020:BPS

[Rap20] Pedro M. P. Raposo. A book with postage stamps. Journal for the History of Astronomy, 51(3):378–379, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620926758.

Rashed:2013:BRT

[Ras13] Roshdi Rashed. Book review: Theodosius’s on habitations, Theodosius, De habitationibus: Arabic and Medieval Latin Translations. Journal for the History of Astronomy, 44 (2):226–227, May 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ravetz:1972:BRC

[Rav72] J. R. Ravetz. Book review: Copernicanism in Poland: Kopernik i heliocentryzm w polskiej kulturze umyslowej do końca XVIII wieku. (Polish) [Copernicus and heliocentrism in the Polish mental culture to the end of the Eighteenth Century]. Journal for the History of Astronomy, 3(3):220, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ravetz:1973:BRP

[Rav73] J. R. Ravetz. Book review: a paradigm case of normal science: Science at the Crossroads. Journal for the History of Astronomy, 4(3):206– 208, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rossello:2000:OTS

[RBH00] Jaume García Rosselló, Joan Fornés Bisquerra, and Michael Hoskin. Orientations of the Talayotic sanctuaries of Mallorca. Journal for the History of Astronomy, 31(25):S58–S64, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Roughton:1992:BNS

[RC92] N. A. Roughton and G. L. Canzoneri. Babylonian normal stars in Sagittarius. Journal for the History of Astronomy, 23 (3):193–200, August 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruis-Castell:2011:BRN

[RC11] Pedro Ruis-Castell. Book review: Nineteenth-Century observatories in context: The Heavens on Earth: Observatories and Astronomy in Nineteenth-Century Science and Culture. Journal for the History of Astronomy, 42(4): 531–532, November 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruiz-Castell:2016:MTP

[RC16] Pedro Ruiz-Castell. Making telescopes and partying with the stars: Amateur astronomy in Spain during Franco’s dictatorship. Journal for the History of Astronomy, 47(2):194– 209, May 2016. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Recio:2018:EPP

[RC18] Gonzalo L. Recio and Christián C.

Carman. On the equant point for the planets and the Moon. Journal for the History of Astronomy, 49(4): 401–424, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618809222.

Reaves:1981:BRD

[Rea81] Gibson Reaves. Book review: The discovery of Pluto: Planets X and Pluto. Journal for the History of Astronomy, 12 (1):68–69, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reaves:1984:BRD

[Rea84] Gibson Reaves. Book review: Discovering planets: Uranus and the Outer Planets, the Planet Pluto. Journal for the History of Astronomy, 15(1): 60–61, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reaves:1992:BRB

[Rea92] Gibson Reaves. Book review: Biography of Clyde Tombaugh: Clyde Tombaugh: Discoverer of Planet Pluto. Journal for the History of Astronomy, 23(1): 71–72, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rebsdorf:2003:BSG

[Reb03] Simon Olling Rebsdorf. Bengt Strömgren: Growing up with astronomy, 1908–1932. Journal for the History of Astronomy, 34(2):171–199, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Recio:2020:MCL

[Rec20] Gonzalo L. Recio. A marvellous connection: Longomontanus’ battle with the latitudes of Mars. Journal for the History of Astronomy, 51(4):383–400, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620954562.

Recio:2021:FEA

[Rec21] Gonzalo Recio. A Festschrift on early astronomy. Journal for the History of Astronomy, 52(3):376–378, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211026212.

Recio:2022:SSN

[Rec22] Gonzalo Luis Recio. A Spanish study of the 1572 nova: Jerónimo Muñoz and his Book on the New Comet. Journal for the History of Astronomy, 53(1):3–12, February 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221077337.

Reeves:2010:BRG

[Ree10] Eileen Reeves. Book review:

Galileian selections: The Essential Galileo. Journal for the History of Astronomy, 41(1):

127–129, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reeves:2011:BRG

[Ree11] Nicky Reeves. Book review: The Greenwich sextants in context: Sextants at Greenwich: a Catalogue of the Mariner’s Quadrants, Mariner’s Astrolabes, Cross-Staffs, Backstaffs, Octants, Sextants, Quintants, Reflecting Circles, and Artificial Horizons in the National Maritime Museum, Greenwich. Journal for the History of Astronomy, 42(2):269–271, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reeves:2016:GTP

[Ree16] Eileen A. Reeves. Galileo’s telescope in a philosophical novel. Journal for the History of Astronomy, 47 (4):430–431, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668753.

Reeves:2020:BRG

[Ree20] Eileen Reeves. Book review: Galileo’s early biographers. On the Life of Galileo: Viviani’s Historical Account and Other Early Biographies. Stefano Gattei (Princeton University Press, Princeton, NJ, 2019). Pp. lxviii + 348. $49.95. ISBN 978-0-691-17489-1. Journal for the History of Astronomy, 51(1):122–123, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619900817.

Reeves:2022:DS

[Ree22] Eileen Reeves. Drawing science. Journal for the History of Astronomy, 53(3):378– 379, August 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221101326.

Reid:1982:COA

[Rei82] John S. Reid. The Castlehill Observatory, Aberdeen. Journal for the History of Astronomy, 13(2):84–96, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reinhard:1992:ERS

[Rei92] Johan Reinhard. Essay review:

Studies of Nazca: The Lines of Nazca. Journal for the History of Astronomy, 23(17):S62–S67, February 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reid:2018:DGF

[Rei18] John S. Reid. David Gill FRS (1843–1914): The making of a Royal Astronomer. Journal for the History of Astronomy, 49(1):3–26, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617751290.

Remmert:2008:BRH

[Rem08] Volker R. Remmert. Book review: Harmonia macrocosmica reprinted, Andreas Cellarius, Harmonia Macrocosmica of 1660: The Finest Atlas of the Heavens. Journal for the History of Astronomy, 39(4):534–535, November 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Remmert:2011:BRE

[Rem11] Volker R. Remmert. Book review: an early-modern mathematician: Between Raphael and Galileo: Mutio Oddi and the Mathematical Culture of Late Renaissance Italy. Journal for the History of Astronomy, 42(4):536–537, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Remmert:2015:BRC

[Rem15] Volker Remmert. Book review: For the cultivated coffee table. Cosmigraphics: Picturing Space through Time. Michael Benson (Abrams, New York, 2014). Pp. 320. $50. ISBN 978-1-4197-1387-3. Journal for the History of Astronomy, 46 (2):244–245, May 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Renn:1989:BRG

[Ren89] Jürgen Renn. Book review:

Galileo in German, Galileo Galilei: Schriften, Briefe, Dokumente. Journal for the History of Astronomy, 20(2): 142–143, June 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruiz:2000:SIA

[RH00] Adriano Gómez Ruiz and Michael Hoskin. Studies in Iberian archaeoastronomy:

(7) orientations of megalithic tombs of Huelva. Journal for the History of Astronomy, 31 (25):S41–S50, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ribeiro:2024:BCC

[Rib24] Luís Campos Ribeiro. From biblical chronology to criticism of astrology. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241229155.

Richards:2007:BRC

[Ric07] E. G. Richards. Book review:

Calendars of the western world: The Structure and Mathematics of the Principal Calendars of the Western World: Muslim, Gregorian, Jewish, and other Systems. Journal for the History of Astronomy, 38(2):247– 248, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rieznik:2013:COH

[Rie13] Marina Rieznik. The Córdoba observatory and the history of the ‘personal equation’ (1871– 1886). Journal for the History of Astronomy, 44(3):277– 301, August 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ribeiro:2018:ANE

[RL18] Luís Campos Ribeiro and Henrique Leitão. Astrology with new eyes: The telescope in astrological prognostication. Journal for the History of Astronomy, 49(3):345–362, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1992:NMP

[RM92] Clive L. N. Ruggles and Roger D. Martlew. The North Mull Project (3): Prominent hill summits and their astronomical potential. Journal for the History of Astronomy, 23 (17):S1–S13, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1991:NMP

[RMH91] Clive L. N. Ruggles, Roger D.

Martlew, and Peter D. Hinge. The North Mull Project (2): The wider astronomical potential of the sites. Journal for the History of Astronomy, 22(16): S51–S75, February 1991. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Robinson:1986:BRA

[Rob86]  Jack H. Robinson.  Book

review: Archaeoastronomy and the Roots of Science, Edited by E. C. Krupp (The AAAS Selected Symposia Series 71; Westview Press, Boulder, 1984). Pp. xii + 336. $40. Journal for the History of Astronomy, 17(10):S73–S75, February 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Robinson:1990:EHC

[Rob90] Leif J. Robinson. Enterprise at Harvard College Observatory. Journal for the History of Astronomy, 21(1):89–103, February 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rochberg:2008:BRC

[Roc08] Francesca Rochberg. Book review: Calendrics in the ancient near East, Calendars and Years: Astronomy and Time in the Ancient near East. Journal for the History of Astronomy, 39(4):541–543, November 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rochberg:2014:AAN

[Roc14]  Francesca Rochberg.  As-

tronomy in the Ancient Near East. Journal for the History of Astronomy, 45(1):127–129, February 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rochberg:2016:PSA

[Roc16] Francesca Rochberg. Pathways into the study of ancient sciences. Journal for the History of Astronomy, 47(1): 102–105, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
102.full.pdf+html.

Rochberg:2017:NPS

[Roc17] Francesca Rochberg. Neugebauer and premodern science. Essay review: A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science. Edited by Alexander Jones, Christine Proust, and John M. Steele (Springer, Berlin, 2016). Pp. xi + 342. EUR 83. ISBN 978-3-319-25863-8. Journal for the History of Astronomy, 48(3):354–357, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617717126.

Rochberg:2019:BRR

[Roc19] Francesca Rochberg. Book review: Resonances within ancient astronomical knowledge. Journal for the History of Astronomy, 50(4): 476–480, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619871738.

Rodger:1973:GIB

[Rod73] Ian Rodger. Er Grah identical with Bregon’s Tower? Journal for the History of Astronomy, 4(2):130, June 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Romain:2006:BRN

[Rom06] William F. Romain. Book review: a native American site: Cahokia: Mirror of the Cosmos. Journal for the History of Astronomy, 37(1):120–121, February 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ronan:1989:BRH

[Ron89] Colin Ronan. Book review:

The history of the RAS, History of the Royal Astronomical Society. Journal for the History of Astronomy, 20(2):146– 148, June 1989. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosen:1973:BRH

[Ros73] Edward Rosen. Book review: Hoyle on Copernicus: Nicolaus Copernicus. Journal for the History of Astronomy, 4(3): 202–203, October 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosen:1974:ERK

[Ros74a] Edward Rosen. Essay review: Koyré in translation: The Astronomical Revolution: Copernicus–Kepler– Borelli. Journal for the History of Astronomy, 5(3):201–203, October 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosenfeld:1974:BRB

[Ros74b] B. Rosenfeld. Book review: Biruni: Zhizn’ i Trudy Beruni (Life and Works of al-Biruni). Journal for the History of Astronomy, 5(2):135–136, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rose:1979:BRF

[Ros79] P. L. Rose. Book review: Fictional Copernicus, Doctor Copernicus. Journal for the History of Astronomy, 10(2): 128, June 1979. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosen:1981:BRG

[Ros81] Edward Rosen. Book review: Galileo in outline: Galileo.

Journal for the History of Astronomy, 12(2):150–152, June 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosenfeldt:1982:SAS

[Ros82a] G. Rosenfeldt. A statistical analysis of some Palenque dates. Journal for the History of Astronomy, 13(4):S49–S67, February 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Roseveare:1982:MPV

[Ros82b] N. T. Roseveare. Mercury’s perihelion, from Le Verrier to Einstein. Oxford science publications. Clarendon Press, New York, NY, USA, 1982. ISBN 0-19-858174-2. viii + 208 pp. LCCN QB611 .R67 1982. US$20.00.

Rosenfeldt:1984:SME

[Ros84] G. Rosenfeldt. A statistical method of evaluating megalithic observatories. Journal for the History of Astronomy, 15(7):S111–S118, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosen:1986:NEC

[Ros86] Edward Rosen. No edition of Copernicus in 1640 or 1646. Journal for the History of Astronomy, 17(1):58–59, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rosson:2019:BRW

[Ros19] Lois Rosson. Book review: When the Western frontier met the final frontier. Journal for the History of Astronomy, 50(4):499–500, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619884473.

Rothenberg:1980:BRA

[Rot80] Marc Rothenberg. Book review: Astronomers at the Cape Observatory, Astronomers at the Royal Observatory, Cape of Good Hope: a History with Emphasis on the Nineteenth Century. Journal for the History of Astronomy, 11(3):214–215, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rothenberg:1990:PHC

[Rot90] Marc Rothenberg. Patronage of Harvard College Observatory, 1839–1851. Journal for the History of Astronomy, 21 (1):37–46, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rothenberg:2001:BRN

[Rot01] Marc Rothenberg. Book review: The nautical almanac Office in perspective: Proceedings: Nautical Almanac Office Sesquicentennial Symposium: U.S. Naval Observatory, March 3–4, 1999. Journal for the History of Astronomy, 32 (2):182–183, May 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rothenberg:2003:BRR

[Rot03] Marc Rothenberg. Book review: Regional science in the U.S.A.: Science in the American Southwest: a Topical History. Journal for the History of Astronomy, 34(4):462–463, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rothman:2009:FPK

[Rot09] Aviva Rothman. Forms of persuasion: Kepler, Galileo, and the dissemination of Copernicanism. Journal for the History of Astronomy, 40(4):403–419, November 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rothenberg:2017:PAU

[Rot17]  Marc Rothenberg.  Paying

for astrophysics in the United States. Essay review: In Search of the True Universe: The Tools, Shaping, and Cost of Cosmological Thought. Martin Harwit (Cambridge University Press, Cambridge, 2013). Pp. xvii + 393. £35. ISBN 978-1-107-04406-7. Journal for the History of Astronomy, 48(3):364–365, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617708700.

Rothenberg:2018:AAA

[Rot18] Marc Rothenberg. Amateurs in American astronomy. Journal for the History of Astronomy, 49(1):119–120, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617736156.

Rowland:2007:BRK

[Row07] Ingrid D. Rowland. Book review: Kircher in context, Die Große Kosmologische Kontroversie: Rekonstruktionsversuche anhand des Itinerarium Exstaticum von Athanasius Kircher, S.J. (1602–1680). Journal for the History of Astronomy, 38(3):388–389, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reaves:1987:LVD

[RP87] Gibson Reaves and Carlo Pedretti. Leonardo da Vinci’s drawings of the surface features of the Moon. Journal for the History of Astronomy, 18(1): 55–58, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rius-Pinies:2015:AAD

[RPPA15] Mònica Rius-Piniés and Roser Puig-Aguilar. Al-Asfi’s description of the Zawiya Nasiriyya:

The use of buildings as astronomical tools. Journal for the History of Astronomy, 46(3): 325–342, August 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Reich:2013:FSK

[RR13] Karin Reich and Elena Roussanova. Formeln und Sterne: Korrespondenz deutscher Gelehrter mit der Kaiserlichen Akademie der Wissenschaften zu St. Petersburg: Briefe von Johann Wilhelm Andreas Pfaff, Johann Sigismund Gottfried Huth, Wilhelm Struve, Martin Bartels, Magnus Georg Paucker aus der Autographensammlung von Wilhelm Stieda in der Universitätsbibliothek Leipzig, volume Bd. 13 of Relationes. Shaker, Aachen, 2013. ISBN 3-8440-2450-6. 435 pp. LCCN Q141 .R45 2013.

Ruggles:1984:IPC

[RS84] C. L. N. Ruggles and N. J.

Saunders. The interpretation of the pecked cross symbols at Teotihuacan: a methodological note. Journal for the History of Astronomy, 15(7):S101– S110, February 1984. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Remchin:2018:RAC

[RS18] Julia Remchin and Andreas Schrimpf. Research on asteroids of Christian Ludwig Gerling and his students in the Nineteenth Century. Journal for the History of Astronomy, 49(1):83–98, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617746805.

Rozelot:2022:SOI

[RS22] Jean-Pierre Rozelot and Jagdev Singh. Spectrographic observations of the ionized iron coronal emission lines at Pic du Midi Observatory (F) in the mid-60s. Journal for the History of Astronomy, 53 (3):300–327, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221101604.

Rudd:1992:BRI

[Rud92] M. Eugene Rudd. Book review: Instrument making in Ireland: Vulgar & Mechanick: The Scientific Instrument Trade in Ireland, 1650–1921. Journal for the History of Astronomy, 23 (1):74, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruffner:1971:CSC

[Ruf71] J. A. Ruffner. The curved and the straight: Cometary theory from Kepler to Hevelius. Journal for the History of Astronomy, 2(3):178–194, October 1971. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruffner:2010:INH

[Ruf10] J. A. Ruffner. Isaac Newton’s historia cometarum and the quest for elliptical orbits. Journal for the History of Astronomy, 41(4):425–451, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1982:RHP

[Rug82] C. L. N. Ruggles. A reassessment of the high precision megalithic lunar sightlines, 1: Backsights, indicators and the archaeological status of the sightlines. Journal for the History of Astronomy, 13(4): S21–S40, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1983:RHP

[Rug83] C. L. N. Ruggles. A reassessment of the high precision megalithic lunar sightlines, 2: Foresights and the problem of selection. Journal for the History of Astronomy, 14(5): S1–S36, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1985:LSA

[Rug85] C. L. N. Ruggles. The linear settings of Argyll and Mull. Journal for the History of Astronomy, 16(9):S105– S132, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1987:BCS

[Rug87] C. L. N. Ruggles. The Borana calendar: Some observations. Journal for the History of Astronomy, 18(11):S35–S53, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1994:SRS

[Rug94] Clive L. N. Ruggles. The stone rows of South-West Ireland: a first reconnaissance. Journal for the History of Astronomy, 25(19):S1–S20, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1995:ERC

[Rug95] Clive Ruggles. Essay review: Co-operation: The ideal and the reality: Colloquio Internazionale: Archeologia e Astronomia. Journal for the History of Astronomy, 26(20): S80–S86, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1996:SRT

[Rug96] Clive L. N. Ruggles. Stone rows of three or more stones in South-West Ireland. Journal for the History of Astronomy, 27(21):S55–S71, February 1996. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1997:WE

[Rug97] Clive L. N. Ruggles. Whose equinox? Journal for the History of Astronomy, 28(22): S44–S50, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:1999:ERS

[Rug99] Clive Ruggles. Essay review:

Sun, Moon, stars, and Stonehenge: Stonehenge: Neolithic Man and the Cosmos. Journal for the History of Astronomy, 30(24):S83–S88, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:2000:ILA

[Rug00] Clive Ruggles. Inaugural lecture: Ancient astronomies — ancient worlds. Journal for the History of Astronomy, 31(25): S65–S76, February 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:2004:ERC

[Rug04] Clive Ruggles. Essay review:

Cross-cultural astronomy: Astronomy across Cultures: The History of Non-Western Astronomy. Journal for the History of Astronomy, 35(2):229– 236, May 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruggles:2007:BRA

[Rug07]  Clive Ruggles.  Book re-

view: Archaeoastronomy in overview: Exploring Ancient Skies: an Encyclopedic Survey of Archaeoastronomy. Journal for the History of Astronomy, 38(1):133–134, February 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rupke:2015:PPC

[Rüp15] Jörg Rüpke. Political power and calendar-making in antiquity. Journal for the History of Astronomy, 46(4):494–495, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Russell:1974:CTSa

[Rus74a] John L. Russell. Cosmological teaching in the Seventeenth-Century Scottish universities. I. Journal for the History of Astronomy, 5(2):122–132, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Russell:1974:CTSb

[Rus74b] John L. Russell. Cosmological teaching in the Seventeenth-Century Scottish universities. II. Journal for the History of Astronomy, 5(3):145–154, October 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ruskin:2002:NDL

[Rus02] Steven W. Ruskin. A newlydiscovered letter of J. F. W. Herschel concerning the “Great Moon Hoax”. Journal for the History of Astronomy, 33(1): 71–74, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rutkin:2003:BRA

[Rut03] H. Darrel Rutkin. Book review:

Astrology surveyed: Astrology: a History. Journal for the History of Astronomy, 34(3):335– 336, August 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rutkin:2012:BRM

[Rut12] Darrel Rutkin. Book review:

Michael Scot’s Liber de Signis: Sternstunden am Kaiserhof: Michael Scotus und sein Buch von den Bildern und Zeichen des Himmels. Journal for the History of Astronomy, 43(1):119–120, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rutkin:2013:BRA

[Rut13] H. Darrel Rutkin. Book review: Astrology and magic, Astrology and Magic from the Medieval Latin and Islamic World to Renaissance Europe: Theories and Approaches. Journal for the History of Astronomy, 44(2):220–222, May 2013.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Rutkin:2016:MHA

[Rut16] H. Darrel Rutkin. Maximilian I and his astrologers. Journal for the History of Astronomy, 47(4):436–439, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668747.

Rutkin:2021:APA

[Rut21] H. Darrel Rutkin. Astronomical predictions of the Antichrist. Journal for the History of Astronomy, 52(1): 108–110, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620985096.

Ryan:1971:BRH

[Rya71] W. F. Ryan. Book review: Historical scientific instruments in Russia: Nauchnye pribory. Pribory i instrumenty istoricheskogo znacheniya (Scientific Instruments. Instruments of Historic Significance), Pamyatniki russkoy kul’tury pervoy chetverti XVIII veka v sobranii Gosudarstvennogo ordena Lenina Ermitazha. Journal for the History of Astronomy, 2(2):128–129, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ryan:2009:BRC

[Rya09] W. F. Ryan. Book review: Cosmology among the Slavs, ‘Let There Be Lights in the Firmament of the Heaven’: Cosmological Depictions in Early Rus. Journal for the History of Astronomy, 40(2):222–224, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sadler:1990:WPS

[Sad90] Philip M. Sadler. William Pickering’s search for a planet beyond Neptune. Journal for the History of Astronomy, 21(1): 59–64, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sakurai:1980:SAT

[Sak80] Kunitomo Sakurai. The solar activity in the time of Galileo. Journal for the History of Astronomy, 11(3):164–173, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1976:DAL

[Sal76] George A. Saliba. The doubleargument lunar tables of Cyriacus. Journal for the History of Astronomy, 7(1):41–46, February 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1979:BRB

[Sal79]  George Saliba.  Book re-

view: Biruni’s mathematical geography, A Commentary upon Biruni’s Kitab Tah.did al-’Amakin. Journal for the History of Astronomy, 10(1): 50–51, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1985:SOM

[Sal85] George Saliba. Solar observations at the Maraghah Observatory before 1275: a new set of parameters. Journal for the History of Astronomy, 16 (2):113–122, June 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1987:TOI

[Sal87] George Saliba. Theory and observation in Islamic astronomy: The work of ibn al-Shāt.ir of Damascus. Journal for the History of Astronomy, 18(1): 35–43, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1989:MAR

[Sal89] George Saliba. A medieval Arabic reform of the Ptolemaic lunar model. Journal for the History of Astronomy, 20(3): 157–164, October 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1990:ERA

[Sal90] George Saliba. Essay review:

Arabic astronomy in Byzantium, The Astronomical Works of Gregory Chioniades. Journal for the History of Astronomy, 21(2):211–215, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1994:EAC

[Sal94a] George Saliba. Early Arabic critique of Ptolemaic cosmology: a Ninth-Century text on the motion of the celestial spheres. Journal for the History of Astronomy, 25(2):115– 141, May 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1994:SCA

[Sal94b] George Saliba. A Sixteenth-Century Arabic critique of Ptolemaic astronomy: the work of Shams al-Dīn al-Khafrī. Journal for the History of Astronomy, 25(1):15–38, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1995:ERI

[Sal95]  George Saliba.  Essay re-

view: an introduction to astrology: Abu Ma’sar, The Abbreviation of the Introduction to Astrology: Together with the Medieval Latin Translation of Adelard of Bath. Journal for the History of Astronomy, 26 (2):168–170, May 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:1997:BRP

[Sal97]  George Saliba.  Book re-

view: Pre-modern astronomy and culture: Tradition, Transmission, Transformation: Proceedings of Two Conferences on Pre-Modern Science Held at the University of Oklahoma. Journal for the History of Astronomy, 28(4):360–362, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:2006:ERI

[Sal06] George Saliba. Essay review: Islamic science at its best: In Synchrony with the Heavens: Studies in Astronomical Timekeeping and Instrumentation in Medieval Islamic Civilization, I: The Call of the Muezzin, In Synchrony with the Heavens: Studies in Astronomical Timekeeping and Instrumentation in Medieval Islamic Civilization, II: Instruments of Mass Calculation. Journal for the History of Astronomy, 37(2):233– 238, May 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saliba:2009:BRA

[Sal09] George Saliba. Book review: Astronomy in the western Islamic world, Astronomy and Astrology in Al-Andalus and the Maghrib. Journal for the History of Astronomy, 40(1): 122–124, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:1998:OHM

[Sam98] Julio Samsó. An outline of the history of Maghribī Zijes from the end of the Thirteenth Century. Journal for the History of Astronomy, 29(2):93– 102, May 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:2007:BRA

[Sam07] Julio Samsó. Book review: Al-Farghani’s treatise on the astrolabe, Al-Farghani on the Astrolabe. Journal for the History of Astronomy, 38(4):507–509, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:2013:ERA

[Sam13] Julio Samsó. Essay review: Astronomical tables in the Middle Ages: A Survey of European Astronomical Tables in the Late Middle Ages. Journal for the History of Astronomy, 44(1):95–98, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:2019:IIA

[Sam19] Julio Samsó. Ibn Ish.aq and the Alfonsine Tables. Journal for the History of Astronomy, 50(3):360–365, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619829242.

Sanderson:1999:BRL

[San99] Richard Sanderson. Book review: Looking at the Leonids: The Heavens on Fire: The Great Leonid Meteor Storms. Journal for the History of Astronomy, 30(3):329–330, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sarma:1978:BRS

[Sar78] S. R. Sarma. Book review: Sources in Sanskrit: Census of the Exact Sciences in Sanskrit. Journal for the History of Astronomy, 9(1):64, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sauzade:2000:OPD

[Sau00]  Gérard Sauzade.  Orienta-

tions of the Provençal dolmens. Journal for the History of Astronomy, 31(25):S1–S10, February 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Savoie:2012:BRN

[Sav12] Denis Savoie. Book review: The ‘Navicula’ sundial: Monks, Manuscripts and Sundials: The Navicula in Medieval England. Journal for the History of Astronomy, 43(1):120–122, February 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sawyer:1980:BRC

[Saw80] Frederick W. Sawyer III. Book review: Catalogue of Greek and Roman sundials, Greek and Roman Sundials. Journal for the History of Astronomy, 11 (2):135–137, June 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Seeley:1972:DRIa

[SB72a] D. Seeley and R. Berendzen.

The development of research in interstellar absorption, c. 1900– 1930. Journal for the History of Astronomy, 3(1):52–64, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Seeley:1972:DRIb

[SB72b] D. Seeley and R. Berendzen.

The development of research in interstellar absorption, c. 1900– 1930: Part 2. Journal for the History of Astronomy, 3(2):75– 86, June 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1984:WLR

[SB84] Robert W. Smith and Richard Baum. William Lassell and the Ring of Neptune: a case study in instrumental failure. Journal for the History of Astronomy, 15(1):1–17, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shaltout:2005:OAE

[SB05] Mosalam Shaltout and Juan Antonio Belmonte. On the orientation of Ancient Egyptian temples: (1) Upper Egypt and Lower Nubia. Journal for the History of Astronomy, 36(3): 273–298, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2012:ECO

[SB12] F. Richard Stephenson and Joshua T. Baylis. Early Chinese observations of occultations of planets by the Moon. Journal for the History of Astronomy, 43(4):455–477, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shaltout:2007:OAEb

[SBF07a] Mosalam Shaltout, Juan Antonio Belmonte, and Magdi Fekri. On the orientation of Ancient Egyptian temples: (3) key points at Lower Egypt and Siwa Oasis, Part II. Journal for the History of Astronomy, 38(4):413–442, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shaltout:2007:OAEa

[SBF07b] Mosalam Shaltout, Juan Antonio Belmonte, and Magdi Fekri. On the orientation of Ancient Egyptian temples: (3) key points in Lower Egypt and Siwa Oasis, Part I. Journal for the History of Astronomy, 38 (2):141–160, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:1988:HEP

[SC88] Julio Samsó and Francisco Castelló. An hypothesis on the epoch of Ptolemy’s star catalogue according to the authors of the Alfonsine Tables. Journal for the History of Astronomy, 19(2):115–120, May 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Symons:2013:AED

[SC13] Sarah Symons and Robert Cockcroft. An Ancient Egyptian diagonal star table in Mallawi, Egypt. Journal for the History of Astronomy, 44 (4):457–463, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schlegel:1970:BRP

[Sch70] Richard Schlegel. Book review:

The paradox of Olbers’s paradox: The Paradox of Olbers’ Paradox: a Case History of Scientific Thought. Journal for the History of Astronomy, 1(1): 83–84, February 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1978:PNF

[Sch78] Simon Schaffer. The Phoenix of Nature: Fire and evolutionary cosmology in Wright and Kant. Journal for the History of Astronomy, 9(3):180–200, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1979:JMB

[Sch79] Simon Schaffer. John Michell and black holes. Journal for the History of Astronomy, 10 (1):42–43, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1980:GLU

[Sch80] Simon Schaffer. “The great laboratories of the universe”: William Herschel on matter theory and planetary life. Journal for the History of Astronomy, 11(2):81–111, June 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1981:UEH

[Sch81] Simon Schaffer. Uranus and the establishment of Herschel’s astronomy. Journal for the History of Astronomy, 12(1): 11–26, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1982:BRA

[Sch82] Simon Schaffer. Book review:

Astronomy in history: The View from Planet Earth: Man Looks at the Cosmos. Journal for the History of Astronomy, 13(2):139–141, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schove:1984:AMC

[Sch84a] D. J. Schove. Addendum:

[Maya correlations, Moon ages and astronomical cycles]. Journal for the History of Astronomy, 15(1):29, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Sch84b].

Schove:1984:MCM

[Sch84b] D. J. Schove. Maya correlations, Moon ages and astronomical cycles. Journal for the History of Astronomy, 15(1): 18–29, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [Sch84a].

Schaefer:1986:AEE

[Sch86] Bradley E. Schaefer. Atmospheric extinction effects on stellar alignments. Journal for the History of Astronomy, 17 (10):S32–S42, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:1987:HRP

[Sch87] Bradley E. Schaefer. Heliacal rise phenomena. Journal for the History of Astronomy, 18 (11):S19–S33, February 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1990:ERG

[Sch90] Simon Schaffer. Essay review: Galilean perspectives, Reinterpreting Galileo. Journal for the History of Astronomy, 21(4): 359–361, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1991:BRA

[Sch91] Simon Schaffer. Book review: Armagh Observatory, Church, State and Astronomy in Ireland: 200 Years of Armagh Observatory. Journal for the History of Astronomy, 22(3):247– 248, August 1991. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:1992:LLM

[Sch92] Bradley E. Schaefer. The length of the lunar month. Journal for the History of Astronomy, 23(17):S32–S42, February 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:1999:BRH

[Sch99] Simon Schaffer. Book review: Harrison’s rights and wrongs:

The Quest for Longitude. Journal for the History of Astronomy, 30(1):76–78, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2000:HRS

[Sch00a] Bradley E. Schaefer. The helical rise of Sirius and Ancient Egyptian chronology. Journal for the History of Astronomy, 31(2):149–155, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2000:BRM

[Sch00b] Sara Schechner. Book review:

The many facets of Halley: Edmond Halley: Charting the Heavens and Seas. Journal for the History of Astronomy, 31 (2):172–173, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2001:LOA

[Sch01a] Bradley E. Schaefer. The latitude of the observer of the Almagest star catalogue. Journal for the History of Astronomy, 32(1):1–42, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano01b].

Schaefer:2001:TVN

[Sch01b] Bradley E. Schaefer. The transit of Venus and the notorious black drop effect. Journal for the History of Astronomy, 32(4):325–336, November 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2001:MCA

[Sch01c] Sara Schechner. The material culture of astronomy in daily life: Sundials, science, and social change. Journal for the History of Astronomy, 32(3): 189–222, August 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2002:LEF

[Sch02] Bradley E. Schaefer. The latitude and epoch for the formation of the southern Greek constellations. Journal for the History of Astronomy, 33(4): 313–350, November 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2003:BRC

[Sch03] Sara Schechner. Book review: Comets: Sublunary or celestial?: The Age of Two-Faced Janus: The Comets of 1577 and 1618 and the Decline of the Aristotelian World View in the Netherlands. Journal for the History of Astronomy, 34(1): 118–119, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2004:LEO

[Sch04a] Bradley E. Schaefer. The latitude and epoch for the origin of the astronomical lore of Eudoxus. Journal for the History of Astronomy, 35(2):161– 223, May 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaldach:2004:AAO

[Sch04b]  Karlheinz Schaldach.  The

Arachne of the Amphiareion and the origin of gnomonics in Greece. Journal for the History of Astronomy, 35(4):435–445, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2005:ECF

[Sch05a] Bradley E. Schaefer. The epoch of the constellations on the Farnese atlas and their origin in Hipparchus’s lost catalogue. Journal for the History of Astronomy, 36(2):167–196, May 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:2005:BRI

[Sch05b] Simon Schaffer. Book review: Imperial eclipse expeditions: Empire and the Sun: Victorian Solar Eclipse Expeditions. Journal for the History of Astronomy, 36(1):114–116, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2007:BRS

[Sch07a] Bradley E. Schaefer. Book review: Sirius, Sirius: Brightest Diamond in the Night Sky. Journal for the History of Astronomy, 38(3):386–387, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schweber:2007:BRE

[Sch07b] Silvan Sam Schweber. Book review: Eddington, Chandra, et al.; Empire of the Stars: Obsession, Friendship, and Betrayal in the Quest for Black Holes. Journal for the History of Astronomy, 38(4):519–522, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schutz:2011:BRS

[Sch11] Bernard F. Schutz. Book review: Spinning black holes: Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics. Journal for the History of Astronomy, 42(1):118–120, February 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaffer:2012:BCA

[Sch12a] Simon Schaffer. The Bombay Case: Astronomers, instrument makers and the East India Company. Journal for the History of Astronomy, 43 (2):151–180, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2012:BRA

[Sch12b] Sara J. Schechner. Book review: an American of many lives: Two Brides for Apollo: The Life of Samuel Williams, 1743–1817. Journal for the History of Astronomy, 43(1): 122–123, February 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schaefer:2013:TSM

[Sch13] Bradley E. Schaefer. The thousand star magnitudes in the catalogues of Ptolemy, Al sufi, and Tycho are all corrected for atmospheric extinction. Journal for the History of Astronomy, 44(1):47–A91, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2014:BRH

[Sch14a] Sara J. Schechner. Book review: Historical instruments in the Mikulov collections: Historické vedecké prístroje v mikulovských sbírkách: Katalog vedeckých prístroju z 16. az 19. Století ve sbírkách Regionálního muzea v Mikulove [Historical Scientific Instruments in the Mikulov Collections: Catalogue of Scientific Instruments from the 16th to 19th Centuries in the Collection of the Regional Museum in Mikulov]. Journal for the History of Astronomy, 45(1): 134–135, February 2014. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schwemin:2014:BZC

[Sch14b] Friedhelm Schwemin, editor. Der Briefwechsel zwischen Carl Friedrich Gauß und Johann Elert Bode. (German) [The correspondence between Carl Friedrich Gauss and Johann Elert Bode], volume 53 of Acta Historica Astronomiae. AVA, Leipzig, Germany, 2014. ISBN 3-944913-43-4 (paperback). 143 pp. LCCN ????

Schaefer:2015:BRS

[Sch15] Bradley E. Schaefer. Book review: The star of Bethlehem. The Star of Bethlehem: A Skeptical View. Aaron Adair (Onus Books, Fareham, Hampshire, 2013). Pp. 155. $14. ISBN 978-0-9566948-6-7. Journal for the History of Astronomy, 46(2):240–242, May 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2018:GHH

[Sch18] Sara J. Schechner. “Girlhours” at the Harvard Observatory. Journal for the History of Astronomy, 49 (1):117–119, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617749121.

Schmidl:2020:AIO

[Sch20] Petra G. Schmidl. Astronomical instruments in the Ottoman Empire. Journal for the History of Astronomy, 51(4):497–499, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620943749.

Schaefer:2023:RNC

[Sch23a] Bradley E. Schaefer. The recurrent nova T CrB had prior eruptions observed near December 1787 and October 1217 AD. Journal for the History of Astronomy, 54(4): 436–455, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231200492.

Schou:2023:TBO

[Sch23b] Ole Schou. Tycho Brahe’s observations of Præsepe Cancri. Journal for the History of Astronomy, 54(1):34– 75, February 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221126362.

Schou:2025:TBQ

[Sch25] Ole Schou. Tycho Brahe’s Quadrans Muralis — a detailed review. Journal for the History of Astronomy, 56 (1):87–106, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241264261.

Scott:1990:AST

[Sco90] Douglas Scott. Astronomical survey of three groups of standing stones in Strath Spey, Scotland. Journal for the History of Astronomy, 21(15):S56–S58, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Scott:2016:DEE

[Sco16]  Douglas Scott.  Dark en-

ergy for everyone. Journal for the History of Astronomy, 47(4):427–428, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616668745.

Sanhueza-Cerda:2020:FPO

[SCV20] Carlos Sanhueza-Cerda and Lorena B. Valderrama. Finding a point of observation in the global south: The C. L. Gerling and J. M. Gilliss correspondence (1847–1856). Journal for the History of Astronomy, 51(2):187–208, May 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620919536.

Smith:1982:INT

[SD82] F. Graham Smith and J. Dudley. The Isaac Newton Telescope. Journal for the History of Astronomy, 13(1):1–18, February 1982. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sheehan:2003:SVI

[SD03] William Sheehan and Thomas Dobbins. The spokes of Venus: an illusion explained. Journal for the History of Astronomy, 34(1):53–63, February 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Seidengart:2017:RCE

[Sei17] Jean Seidengart. Rothmann on the comet of 1585, in English. Journal for the History of Astronomy, 48(2):251–252, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617705015.

Sellers:2006:LWG

[Sel06] David Sellers. A letter from William Gascoigne to Sir Kenelm Digby. Journal for the History of Astronomy, 37(4): 405–416, November 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sellers:2012:SWG

[Sel12]  David Sellers.  In search

of William Gascoigne: seventeenth century astronomer, volume 390 of Astrophysics and space science library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-4096-3. ISSN 0067-0057. x + 222 pp. LCCN QB36.G373 S45 2012.

Sellers:2014:BRU

[Sel14] David Sellers. Book review: The unloved Le Verrier: Le Verrier: Magnificent and Detestable Astronomer. Journal for the History of Astronomy, 45(1):135–137, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Serio:1988:BRR

[Ser88] Giorgia Fodera‘ Serio. Book review: Recognition for Hodierna: La Vita e le Opere di Giovan Battista Hodierna. Journal for the History of Astronomy, 19 (1):66–67, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Serio:1990:GPD

[Ser90] Giorgia Foderà Serio. Giuseppe Piazzi and the discovery of the proper motion of 61 Cygni. Journal for the History of Astronomy, 21(3):275–282, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Serrano:2013:TUR

[Ser13] Juan D. Serrano. Trying Ursus:

a reappraisal of the Tycho– Ursus priority dispute. Journal for the History of Astronomy, 44(1):17–46, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sevcenko:2000:BRC

[Sev00]  Ihor Sevcenko.  Book re-

view: Copernicus in Orthodox Greece: Die religiöse Kritik am kopernikanischen Weltbild in Griechenland zwischen 1794 und 1821. Aspekte griechisch– Orthodoxer Apologetik angesichts naturwissenchaftlicher Fortschritte. Journal for the History of Astronomy, 31(2): 165–168, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1993:LET

[SF93] F. Richard Stephenson and Louay J. Fatoohi. Lunar eclipse times recorded in Babylonian history. Journal for the History of Astronomy, 24(4):255–267, November 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1994:BUT

[SF94] F. Richard Stephenson and Louay J. Fatoohi. The Babylonian unit of time. Journal for the History of Astronomy, 25(2):99–110, May 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1995:ASE

[SF95] F. R. Stephenson and L. J. Fatoohi. Accuracy of solar eclipse observations made by Jesuit astronomers in China. Journal for the History of Astronomy, 26(3):227–236, August 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1997:TPS

[SF97] F. Richard Stephenson and Louay J. Fatoohi. Thales’s prediction of a solar eclipse. Journal for the History of Astronomy, 28(4):279–282, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2005:RSP

[SG05] F. Richard Stephenson and David A. Green. A reappraisal of some proposed historical supernovae. Journal for the History of Astronomy, 36(2):217– 229, May 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2007:SBO

[SG07] J. M. Steele and J. M. K. Gray. A study of Babylonian observations involving the zodiac. Journal for the History of Astronomy, 38(4):443–458, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2009:CGS

[SG09] F. Richard Stephenson and David A. Green. A catalogue of “guest stars” recorded in East Asian history from earliest times to A.D. 1600. Journal for the History of Astronomy, 40 (1):31–54, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Strano:2016:LCH

[SG16] Giorgio Strano and Museo Galileo. The Louwman Collection of historic telescopes. Journal for the History of Astronomy, 47(2):219–220, May 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2002:IUC

[SGDO02] F. Richard Stephenson, Alexander Gurshtein, Steven J. Dick, and Wayne Orchiston. The Inter-Union Commission for History of Astronomy. Journal for the History of Astronomy, 33(2):216–217, May 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1981:ATS

[SH81] F. Richard Stephenson and Michael A. Houlden. The accuracy of Tuckerman’s solar and planetary tables. Journal for the History of Astronomy, 12 (2):133–138, June 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sharratt:1970:BRA

[Sha70] Bernard Sharratt. Book review: Astronomy in English literature: The High Firmament. Journal for the History of Astronomy, 1(1):80–81, February 1970. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shane:1971:ALO

[Sha71a] Mary Lea Shane. The archives of Lick Observatory. Journal for the History of Astronomy, 2 (1):51–53, February 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shapiro:1971:BRJ

[Sha71b] Alan E. Shapiro. Book review: Jakob Bernoulli’s early work: Die Werke von Jakob Bernoulli. Band 1. Journal for the History of Astronomy, 2(3):207, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shapiro:1975:AMS

[Sha75] Alan E. Shapiro. Archimedes’s measurement of the Sun’s apparent diameter. Journal for the History of Astronomy, 6(2): 75–83, June 1975. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shakeshaft:1978:BRJ

[Sha78] J. R. Shakeshaft. Book review: Jodrell Bank and Cambridge. Journal for the History of Astronomy, 9(2):142– 145, June 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1982:BRR

[Sha82] Michael Shank. Book review:

The Regiomontanus Symposium: Regiomontanus–Studien. Journal for the History of Astronomy, 13(1):66–67, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sharma:1984:GAJ

[Sha84] Virendra Nath Sharma. The Great Astrolabe of Jaipur and its sister unit. Journal for the History of Astronomy, 15 (7):S126–S128, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1985:BRH

[Sha85] Michael H. Shank. Book review: Humanist science in Vienna: Humanistische Naturwissenschaft: Zur Geschichte der Wiener mathematischen Schulen des 15. und 16. Jahrhunderts. Journal for the History of Astronomy, 16(1): 52–54, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1987:ERN

[Sha87a] Michael H. Shank. Essay review: Newtonian biographies: Never at Rest: a Biography of Isaac Newton, in the Presence of the Creator: Isaac Newton and His Times. Journal for the History of Astronomy, 18 (3):224–228, August 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sharma:1987:BRE

[Sha87b] Virendra N. Sharma. Book review: European astronomers in India: Introduction of Modern Western Astronomy in India during 18–19 Centuries. Journal for the History of Astronomy, 18(2):137–138, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shawcross:1987:BRL

[Sha87c] William E. Shawcross. Book review: Lines to the mountain gods: Nazca and the Mysteries of Peru. Journal for the History of Astronomy, 18(11): S70–S71, February 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1989:BRC

[Sha89] Michael H. Shank. Book review: Commemorating Gmunden, Der Weg der Naturwissenschaft von Johannes von Gmunden zu Johannes Kepler. Journal for the History of Astronomy, 20(3):215–217, October 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1992:NAB

[Sha92] Michael H. Shank. The “Notes on al-Bit.rūjī” attributed to Regiomontanus: second thoughts. Journal for the History of Astronomy, 23(1):15–30, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1994:ERC

[Sha94a] Michael H. Shank. Essay review: The Cosmology of Francis of Marchia: Die Kosmologie des Franciscus de Marchia: Texte, Quellen und Untersuchungen zur Naturphilosophie des 14. Jahrhunderts. Journal for the History of Astronomy, 25(1):59–61, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1994:ERG

[Sha94b] Michael H. Shank. Essay review: Galileo’s day in court: Galileo, Courtier: The Practice of Science in the Culture of Absolutism. Journal for the History of Astronomy, 25(3):236– 243, August 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sharma:1996:BRS

[Sha96] Virendra N. Sharma. Book review: Source materials in Sanskrit: Census of the Exact Sciences in Sanskrit. Journal for the History of Astronomy, 27 (1):82–83, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1998:RHA

[Sha98] Michael H. Shank. Regiomontanus and homocentric astronomy. Journal for the History of Astronomy, 29(2):157–166, May 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1999:BRS

[Sha99a]  Michael H. Shank.  Book

review: a Sixteenth-Century Viennese intellectual: Humanismus zwischen Hof und Universität: Georg Tanstetter (Collimitius) und sein wissenschaftliches Umfeld im Wien des frühen 16. Jahrhunderts. Journal for the History of Astronomy, 30(1):83– 85, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:1999:ERC

[Sha99b] Michael H. Shank. Essay review: Cosmology in 1588: El Debate Cosmológico en 1588: Bruno, Brahe, Rothmann, Ursus, Röslin. Journal for the History of Astronomy, 30(3):

309–311, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shapiro:2002:BRK

[Sha02] Alan E. Shapiro. Book review: Kepler’s optics in translation, Optics: Paralipomena to Witelo & Optical Part of Astronomy. Journal for the History of Astronomy, 33(3):299– 301, August 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shackelford:2007:BRT

[Sha07a] Jole Shackelford. Book review: Tycho in context, Att låta själen flyga mellan himlens tinnar: Tycho Brahe och Renässansen. (Swedish) [To let the soul fly from the battlements of heaven: Tycho Brahe and the Renaissance]. Journal for the History of Astronomy, 38(4):511–513, November 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:2007:RPA

[Sha07b] Michael H. Shank. Regiomontanus as a physical astronomer: Samplings from The Defence of Theon against George of Trebizond. Journal for the History of Astronomy, 38(3):325–349, August 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sharratt:2007:BRG

[Sha07c] Michael Sharratt. Book review: Galileo and his Spyglass, Galileo’s Instruments of Credit: Telescopes, Images, Secrecy. Journal for the History of Astronomy, 38(3):382– 383, August 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:2008:BRJ

[Sha08]  Michael H. Shank.  Book

review: John of Gmunden Commemorated; Johannes von Gmunden (ca. 1384–1442): Astronom und Mathematiker. Journal for the History of Astronomy, 39(1):126–127, February 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:2011:ERP

[Sha11] Michael H. Shank. Essay review: Piero’s flagellation of Christ elucidated? Journal for the History of Astronomy, 42 (3):391–403, August 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:2012:GDR

[Sha12] Michael H. Shank. The geometrical diagrams in Regiomontanus’s edition of his own disputationes (c. 1475): Background, production, and diffusion. Journal for the History of Astronomy, 43(1):27– 55, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shapiro:2014:BRP

[Sha14] Adam R. Shapiro. Book review: The pseudoscience wars: Immanuel Velikovsky and the Birth of the Modern Fringe. Michael D. Gordin (University of Chicago Press, Chicago, 2012). Pp. x + 291. $29. ISBN 978-0-226-30442-7. Journal for the History of Astronomy, 45 (4):504–506, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shank:2017:ERF

[Sha17] Michael H. Shank. Essay review: The French Copernicus edition. Journal for the History of Astronomy, 48 (4):460–480, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617726399.

Shank:2022:CCG

[Sha22] Michael H. Shank. Completing the Copernicus Gesamtausgabe. Journal for the History of Astronomy, 53(2): 217–222, May 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221079824.

Shea:1974:ERK

[She74] William R. Shea. Essay review: Kepler and Galileo: Dissertatio cum Nuncio Sidereo and Narratio de Quattuor Jovis Stellibus. Journal for the History of Astronomy, 5(1):58–60, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shea:1979:ERI

[She79a] William Shea. Essay review: Improving on Copernicus, The Ash Wednesday Supper. Journal for the History of Astronomy, 10(2):117–122, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shea:1979:BRH

[She79b] William R. Shea. Book review: Hermes and Copernicus, Hermeticism and the Scientific Revolution. Journal for the History of Astronomy, 10(1): 54–56, February 1979. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shea:1986:BRG

[She86] William R. Shea. Book review: Galileo documents reedited: I Documenti del Processo di Galileo Galilei. Journal for the History of Astronomy, 17(1):66–67, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shea:1987:BRB

[She87] William Shea. Book review:

Bellarmine on the heavens: The Louvain Lectures (Lectiones Lovanienses) of Bellarmine and the Autograph Copy of His 1616 Declaration to Galileo. Journal for the History of Astronomy, 18(1):74, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shevchenko:1990:AES

[She90] M. Shevchenko. An analysis of errors in the star catalogues of Ptolemy and Ulugh Beg. Journal for the History of Astronomy, 21(2):187–201, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sherrill:1999:CCA

[She99] Thomas J. Sherrill. A career of controversy: The anomaly of T. J. J. See. Journal for the History of Astronomy, 30 (1):25–50, February 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sheehan:2014:CFP

[She14] William Sheehan. Camille Flammarion’s The planet Mars. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-09640-0. xxiii + 528 pp. LCCN QB641 .F63 2015.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy1413/
                                        2014950070-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy1413/2014950070-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1413/2014950070-
                                        t.html.

Sheynin:2015:CPP

[She15] Hayim Y. Sheynin. Claudius Ptolemy? Pseudo-Ptolemy? The main source of Moses Almosnino’s Treatise on Astrolabe. Journal for the History of Astronomy, 46(3):343–350, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shi:2000:EOM

[Shi00] Yunli Shi. Eclipse observations made by Jesuit astronomers in China: a reconsideration. Journal for the History of Astronomy, 31(2):135–147, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shi:2009:BRA

[Shi09] Yunli Shi. Book review: Astronomy in Korea, Astronomy: Traditional Korean Science. Journal for the History of Astronomy, 40(3):355–357, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shi:2011:BRF

[Shi11] Yunli Shi. Book review: Far-Eastern observational records:

Archaeoastronomy in East Asia: Historical Observational Records of Comets and Meteor Showers from China, Japan and Korea. Journal for the History of Astronomy, 42(3):404– 405, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shor:2021:TCM

[Sho21] Viktor Abramovich Shor.

Twentieth-century milestones in the history of the Russian ephemeris service: Marking 100 years of the Calculation Institute and astronomical yearbook. Journal for the History of Astronomy, 52(3):325–351, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211035537.

Shrimplin:2000:MCN

[Shr00] Valerie Shrimplin. Michelangelo and Copernicus: a note on the Sistine Last Judgment. Journal for the History of Astronomy, 31(2):156–160, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Shukla:1982:BRE

[Shu82] K. S. Shukla. Book review:

The exact sciences in Sanskrit: Census of the Exact Sciences in Sanskrit. Journal for the History of Astronomy, 13(3): 225–226, October 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Serio:1992:OTM

[SHV92] Giorgia Foderà Serio, Michael Hoskin, and Frank Ventura. The orientations of the temples of Malta. Journal for the History of Astronomy, 23(2):107– 119, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sidoli:2004:HAM

[Sid04] Nathan Sidoli. Hipparchus and the ancient metrical methods on the sphere. Journal for the History of Astronomy, 35(1): 71–84, February 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sidoli:2010:BRP

[Sid10]  Nathan Sidoli.  Book re-

view: Parapegmata: Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Near-Eastern Societies. Journal for the History of Astronomy, 41(1):124–126, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sidoli:2024:HAA

[Sid24a] Nathan Sidoli. Hipparchos and the ancient analemma. Journal for the History of Astronomy, 55(2):179–206, May 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231226269.

Sidoli:2024:MDT

[Sid24b] Nathan Sidoli. The manuscript diagrams of Theodosios’ Spherics. Journal for the History of Astronomy, 55(4):454– 489, November 2024. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241248639.

Sidoli:2025:LDS

[Sid25] Nathan Sidoli. The “logic”

of diagrams in the Spherics of Theodosios. Journal for the History of Astronomy, 56(1):3–25, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241248638.

Siebert:2005:ESS

[Sie05] Harald Siebert. The early search for stellar parallax: Galileo, Castelli, and Ramponi. Journal for the History of Astronomy, 36(3):251–271, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Silva:2023:BRP

[Sil23] Barbara Kirsi Silva. Book review: Physicists becoming astronomers: Golden Years of Australian Radio Astronomy: An Illustrated History. Orchiston Wayne, Robertson Peter, and Sullivan Woodruff T. III (Springer, Cham, 2021). Pp. xviii + 268. $55. ISBN 978-3-319-91841-9. Open access:

rd.springer.com/book/10.1007/978-3-319-91843-3.

Journal for the History of Astronomy, 54(4):483–484, November 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231208152.

Simpson:1992:JGR

[Sim92] A. D. C. Simpson. James Gregory and the reflecting telescope. Journal for the History of Astronomy, 23(2):77– 92, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Simpson:2009:BCM

[Sim09] A. D. C. Simpson. The beginnings of commercial manufacture of the reflecting telescope in London. Journal for the History of Astronomy, 40(4): 421–466, November 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Serio:1985:GBH

[SIN85] G. Foderà Serio, L. Indorato, and P. Nastasi. G. B. Hodierna’s observations of nebulae and his cosmology. Journal for the History of Astronomy, 16(1):1–36, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Siscoe:1985:BRS

[Sis85] George L. Siscoe. Book review:

Scandinavian studies of the Aurora: The Northern Light: From Mythology to Space Research. Journal for the History of Astronomy, 16(2):150– 151, June 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Siscoe:1987:BRH

[Sis87] George L. Siscoe. Book review:

History and the geosciences: Das Phänomen des Polarlichts, Historical Events and People in Geosciences. Journal for the History of Astronomy, 18 (2):146–147, May 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sivin:1977:BRC

[Siv77] N. Sivin. Book review: Chinese astronomy: Sternkunde und Weltbild im alten China. Gesammelte Aufsätze. Journal for the History of Astronomy, 8 (1):59–60, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sivin:1989:BRN

[Siv89] Nathan Sivin. Book review:

The New Delhi Colloquium, History of Oriental Astronomy.

Journal for the History of Astronomy, 20(2):136–138, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sivin:1998:BRC

[Siv98] Nathan Sivin. Book review: a Chinese text analysed: Astronomy and Mathematics in Ancient China: The Zhou bi Suanjing. Journal for the History of Astronomy, 29(3):302– 304, August 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sivin:2005:BRJ

[Siv05] Nathan Sivin. Book review: a Jesuit Astronomer’s report from China: Ferdinand Verbiest, S.J. (1623–1688) and the Chinese Heaven: The Composition of the Astronomical Corpus, its Diffusion and Reception in the European Republic of Letters. Journal for the History of Astronomy, 36(2):240– 241, May 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sivin:2010:BRC

[Siv10] Nathan Sivin. Book review: Computing Ephemerides in China: Le Calendrier Chinois: Structure et Calculs (104 av. J.-C.–1644). Journal for the History of Astronomy, 41(4): 521–523, November 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Straizys:1997:CAB

[SK97] Vytautas Straizys and Libertas Klimka. The cosmology of the Ancient Balts. Journal for the History of Astronomy, 28(22): S57–S81, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schmidt-Kaler:2003:ERB

[SK03] Theodor Schmidt-Kaler. Essay review: Baade and his influence: Walter Baade: a Life in Astrophysics. Journal for the History of Astronomy, 34 (3):331–333, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Santucci:2008:OMT

[SK08] Jean-François Santucci and El Hadi Khoumeri. Orientations of megalithic tombs in Algeria (1): Djebel Mazela and Rocknia necropolises, and the Kabylian allées couvertes. Journal for the History of Astronomy, 39(1):65–76, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Symons:2016:CAE

[SK16] Sarah Symons and Himanshi Khurana. A catalogue of ancient Egyptian sundials. Journal for the History of Astronomy, 47(4):375–385, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616675962.

Smith:1983:DER

[SL83] F. Graham Smith and Bernard Lovell. On the discovery of extragalactic radio sources. Journal for the History of Astronomy, 14(3):155–165, October 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Segonds:2013:GTP

[SL13] Alain-Philippe Segonds and Concetta Luna. The Greek text of Ptolemy, Almagest XIII, as source of Book VI of Copernicus’s De Revolutionibus. Journal for the History of Astronomy, 44(4):413–427, November 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sluiter:1997:FKT

[Slu97a] Engel Sluiter. The first known telescopes carried to America, Asia and the Arctic, 1614– 39. Journal for the History of Astronomy, 28(2):141–145, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sluiter:1997:TBG

[Slu97b] Engel Sluiter. The telescope before Galileo. Journal for the History of Astronomy, 28(3): 223–234, August 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Samso:1994:IAZ

[SM94] Julio Samsó and Honorino Mielgo. Ibn al-Zarqālluh on Mercury. Journal for the History of Astronomy, 25(4):289– 296, November 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sheehan:2004:MTD

[SM04] William Sheehan and Anthony Misch. Ménage à trois: David Peck Todd, Mabel Loomis Todd, Austin Dickinson, and the 1882 transit of Venus. Journal for the History of Astronomy, 35(2):123–134, May 2004. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Saslaw:2005:DHI

[SM05] William C. Saslaw and Paul Murdin. The double heads of istrus: The oldest eclipse on a coin? Journal for the History of Astronomy, 36(1): 21–27, February 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2018:PEC

[SMH18] F. Richard Stephenson, Leslie V.

Morrison, and Catherine Y. Hohenkerk. The provenance of early Chinese records of large solar eclipses and the determination of the Earth’s rotation. Journal for the History of Astronomy, 49(4): 425–471, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618789850.

Stephenson:2020:AAS

[SMH20] F. Richard Stephenson, Leslie V. Morrison, and Catherine Y. Hohenkerk. Astronomical assessment of the solar and lunar eclipse records in Bede’s ecclesiastical history and its continuation. Journal for the History of Astronomy, 51 (1):75–87, February 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619899188.

Smith:1979:OVD

[Smi79] R. W. Smith. The origins of the velocity–distance relation. Journal for the History of Astronomy, 10(3):133–165, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1981:BRT

[Smi81] Humphry M. Smith. Book review: Timekeeping in Canada: The Beginning of the Long Dash: a History of Timekeeping in Canada. Journal for the History of Astronomy, 12 (1):71–72, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1983:WLD

[Smi83] Robert W. Smith. William Lassell and the discovery of Neptune. Journal for the History of Astronomy, 14(1):30– 32, February 1983. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1985:BRX

[Smi85] Robert W. Smith. Book review: X-ray astronomy in the USA: Glimpsing an Invisible Universe: The Emergence of X-ray Astronomy. Journal for the History of Astronomy, 16 (2):142–144, June 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1989:BRH

[Smi89] Robert W. Smith. Book review: The History of Lick Observatory, Eye on the Sky: Lick Observatory’s First Century. Journal for the History of Astronomy, 20(2):148–150, June 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1991:NOT

[Smi91] Robert W. Smith. A national observatory transformed: Greenwich in the Nineteenth Century. Journal for the History of Astronomy, 22(1):5–20, February 1991. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1992:BRE

[Smi92] Robert W. Smith. Book review: ESO’s early history: The European Southern Observatory from Concept to Reality. Journal for the History of Astronomy, 23(4):302–303, November 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1993:BRG

[Smi93] A. Mark Smith. Book review: a Galileian anthology in Spanish: Galileo: Antología. Journal for the History of Astronomy, 24 (1–2):153–154, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http:/

/jha.sagepub.com/content/
24/1-2/153.full.pdf+html.

Smith:1994:BRB

[Smi94a] Elske V. P. Smith. Book review: Bart Bok: The Man Who Sold the Milky Way: a Biography of Bart Bok. Journal for the History of Astronomy, 25(4):334–336, November 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1994:BRF

[Smi94b] Robert W. Smith. Book review: Friedmann as hero: Alexander A. Friedmann: The Man Who Made the Universe Expand. Journal for the History of Astronomy, 25(2):145–147, May 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1994:BRR

[Smi94c] Robert W. Smith. Book review: The rhetoric of Simon Newcomb: A Scientist’s Voice in American Culture: Simon Newcomb and the Rhetoric of Scientific Method. Journal for the History of Astronomy, 25 (1):71–72, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1996:ERC

[Smi96] George E. Smith. Essay review: Chandrasekhar’s Principia: Newton’s Principia for the Common Reader. S. Chandrasekhar (Clarendon Press, Oxford, 1995). Pp. xx + 593. $135.00; [MR1352683 (96h:01010)]. Journal for the History of Astronomy, 27(4): 353–362, November 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1997:EDS

[Smi97] Robert W. Smith. Engines of discovery: Scientific instruments and the history of astronomy and planetary science in the United States in the Twentieth Century. Journal for the History of Astronomy, 28 (1):49–77, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:1998:BRS

[Smi98] Robert W. Smith. Book review: Steady-state versus big bang: Cosmology and Controversy: The Historical Development of Two Theories of the Universe. Journal for the History of Astronomy, 29(1): 86–87, February 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2000:BSV

[Smi00] Horace A. Smith. Bailey, Shapley, and variable stars in globular clusters. Journal for the History of Astronomy, 31(3): 185–201, August 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2002:BRT

[Smi02] Robert W. Smith. Book review: Talent in the Parsons family, From Galaxies to Turbines: Science, Technology and the Parsons Family. Journal for the History of Astronomy, 33(3):304–305, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2004:BC

[Smi04] George E. Smith. I. Bernard Cohen (1914–2003). Journal for the History of Astronomy, 35(1):117–119, February 2004.

CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2006:BBG

[Smi06] Robert W. Smith. Beyond the big galaxy: The structure of the stellar system 1900–1952. Journal for the History of Astronomy, 37(3):307–342, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2007:BRT

[Smi07]  Howard Smith.  Book re-

view: Travails at NASA: The Last of the Great Observatories: Spitzer and the Era of Faster, Better, Cheaper at NASA. Journal for the History of Astronomy, 38(2):248– 250, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2008:BGD

[Smi08] Robert W. Smith. Beyond the galaxy: The development of extragalactic astronomy 1885– 1965, Part 1. Journal for the History of Astronomy, 39(1): 91–119, February 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2009:BGD

[Smi09a] Robert W. Smith. Beyond the galaxy: The development of extragalactic astronomy 1885– 1965, Part 2. Journal for the History of Astronomy, 40(1):

71–107, February 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2009:BRN

[Smi09b] Robert W. Smith. Book review: NASA’s history in pictures, NASA: The Complete Illustrated History. Journal for the History of Astronomy, 40 (2):237–238, May 2009. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2012:BRC

[Smi12]  Horace Smith.  Book re-

view: Citizen-scientists and the AAVSO, Advancing Variable Star Astronomy: The Centennial History of the American Association of Variable Star Observers. Journal for the History of Astronomy, 43(3):367– 368, August 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2015:DMU

[Smi15] Robert W. Smith. Discovering the modern universe. Journal for the History of Astronomy, 46(4):502–503, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2018:EAC

[Smi18a] Robert W. Smith. Early astrophysics in Canada: review of Northern Star: J. S. Plaskett R. Peter Broughton with a foreword by & James E Hesser (University of Toronto Press, Toronto, ON, Canada, 2018). Pp. xx + 539. CAD 67.50. ISBN 978-1-4426-3017-8. Journal for the History of Astronomy, 49(3):395–396, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Smith:2018:SGW

[Smi18b] Robert W. Smith. A sociologist in the gravitational wave community. Journal for the History of Astronomy, 49(1):116–117, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618755445.

Smith:2020:APF

[Smi20] Robert W. Smith. Astronomy’s past and future. Journal for the History of Astronomy, 51(3):379–380, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620940154.

Smith:2022:SMW

[Smi22]  Robert W. Smith.  Stel-

lar movements and working hypotheses: A. S. Eddington’s early astronomical career. Journal for the History of Astronomy, 53(4): 394–414, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221121913.

Smoller:2009:BRC

[Smo09] Laura A. Smoller. Book review: a classic of mediaeval astrology, The Book of Astronomy. Journal for the History of Astronomy, 40(1):126–128, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Snedegar:1998:FFC

[Sne98] Keith Snedegar. First fruits celebrations among the Nguni peoples of Southern Africa: an ethnoastronomical interpretation. Journal for the History of Astronomy, 29(23):S31–S38, February 1998. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sheehan:2000:HLE

[SO00] William Sheehan and Donald E. Osterbrock. Hale’s “Little Elf”: The mental breakdowns of George Ellery Hale. Journal for the History of Astronomy, 31(2):93–114, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Soltysiak:2009:BRS

[Sol09] Arkadiusz Soltysiak. Book review: Sumero–Akkadian star names, Zvezdnoe nebo drevnei Mesopotamii: Shumero–Akkadskie nazvaniia sozvezdii i drugikh svetil [The Star Heaven of Ancient Mesopotamia: Sumero– Akkadian Names of Constellations and other Heavenly Bodies]. Journal for the History of Astronomy, 40(1):120–122, February 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Silva:2012:CSM

[SP12] Fabio Silva and Fernando Pimenta. The crossover of the Sun and the Moon. Journal for the History of Astronomy, 43(2):191–208, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sisulak:2022:IOA

[SP22] Stanislav Sisulák and Ladislav Pastorek. Instrumentation and observations at the astronomical observatory in Hurbanovo in 1871–1918. Journal for the History of Astronomy, 53(4):475–496, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221127966.

Spaight:2003:AHT

[Spa03] John Tracy Spaight. Alexander Herschel as telescope maker. Journal for the History of Astronomy, 34(1):95–96, February 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Spaight:2004:GAM

[Spa04] John Tracy Spaight. “For the good of astronomy”: The manufacture, sale, and distant use of William Herschel’s telescopes. Journal for the History of Astronomy, 35(1): 45–69, February 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Speller:1980:AML

[Spe80] John L. Speller. Astronomical manuscript from the library of oriel College, Oxford. Journal for the History of Astronomy, 11(2):113, June 1980. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Spelda:2013:SAA

[Spe13] Daniel Spelda. The search for antediluvian astronomy: Sixteenth- and Seventeenth-Century astronomers’ conceptions of the origins of the science. Journal for the History of Astronomy, 44(3):337–362, August 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Spelda:2017:KEH

[Spe17]  Daniel Spelda.  Kepler in

the early historiography of astronomy (1615–1800). Journal for the History of Astronomy, 48(4):381–404, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617740948.

Sprajc:1992:ERS

[Spr92]  Ivan Sprajc.  Essay re-

view: Space and time in pre-Columbian Mexico: Die geordnete Welt indianischer Völker: Ein Beispiel von Raumordnung und Zeitordnung im vorkolumbischen Mexiko. Journal for the History of Astronomy, 23(17):S59–S61, February 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sprajc:1993:ERM

[Spr93a] Ivan Sprajc. Essay review:

The 1984 Mexico City symposium: Arqueoastronomía y Etnoastronomía en Mesoamérica. Journal for the History of Astronomy, 24(18):S73–S74, February 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sprajc:1993:VRMa

[Spr93b]  Ivan Sprajc.  The Venus–

Rain–Maize complex in the Mesoamerican world view: Part I. Journal for the History of Astronomy, 24(1–2):17– 70, May/February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/17.full.pdf+html.

Sprajc:1993:VRMb

[Spr93c]  Ivan Sprajc.  The Venus–

Rain–Maize complex in the Mesoamerican world view: Part II. Journal for the History of Astronomy, 24(18):S27–S53, February 1993. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sprajc:2000:AAT

[Spr00] Ivan Sprajc. Astronomical alignments at the Templo Mayor of Tenochtitlan, Mexico. Journal for the History of Astronomy, 31(25):S10–S40, February 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sofaer:1986:AMZ

[SS86] Anna Sofaer and Rolf M. Sinclair. An appraisal of Michael Zeilik’s “A Reassessment of the Fajada Butte Solar Marker”. Journal for the History of Astronomy, 17(10): S59–S69, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Zei85a].

Sofaer:1990:CSM

[SS90] Anna Sofaer and Rolf M. Sinclair. Changes in solstice marking at the three-slab site, New Mexico, USA. Journal for the History of Astronomy, 21(15): S59–S60, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Said:1991:AEO

[SS91a] S. S. Said and F. R. Stephenson. Accuracy of eclipse observations recorded in medieval Arabic chronicles. Journal for the History of Astronomy, 22 (4):297–310, November 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1991:PMI

[SS91b] F. R. Stephenson and S. S.

Said. Precision of medieval Islamic eclipse measurements. Journal for the History of Astronomy, 22(3):195–207, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Said:1995:PMI

[SS95] S. S. Said and F. R. Stephenson. Precision of medieval Islamic measurements of solar altitudes and equinox times. Journal for the History of Astronomy, 26(2):117–132, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Said:1996:SLE

[SS96] S. S. Said and F. R. Stephenson. Solar and lunar eclipse measurements by medieval Muslim astronomers. I. Background. Journal for the History of Astronomy, 27(3):259– 273, August 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Said:1997:SLE

[SS97a] S. S. Said and F. R. Stephenson. Solar and lunar eclipse measurements by medieval Muslim astronomers. II. Observations. Journal for the History of Astronomy, 28(1): 29–48, February 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1997:LET

[SS97b] J. M. Steele and F. R. Stephenson. Lunar eclipse times predicted by the Babylonians. Journal for the History of Astronomy, 28(2):119–131, May 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1998:AEA

[SS98a] J. M. Steele and F. R. Stephenson. Astronomical evidence for the accuracy of clocks in pre-Jesuit China. Journal for the History of Astronomy, 29(1): 35–48, February 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1998:EOM

[SS98b] John M. Steele and F. Richard Stephenson. Eclipse observations made by Regiomontanus and Walther. Journal for the History of Astronomy, 29(4): 331–344, November 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2006:ADB

[SS06] F. R. Stephenson and J. M.

Steele. Astronomical dating of Babylonian texts describing the total solar eclipse of S.E. 175. Journal for the History of Astronomy, 37(1):55–69, February 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Savage-Smith:2008:BRI

[SS08] Emilie Savage-Smith. Book review: Islamic influence on Copernicus, Islamic Science and the Making of the European Renaissance. Journal for the History of Astronomy, 39(4): 538–541, November 2008. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schechner:2016:SHV

[SS16] Sara J. Schechner and David Sliski. The scientific and historical value of annotations on astronomical photographic plates. Journal for the History of Astronomy, 47(1):3– 29, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
3.full.pdf+html.

Steele:1997:AET

[SSM97] J. M. Steele, F. R. Stephenson, and L. V. Morrison. The accuracy of eclipse times measured by the Babylonians. Journal for the History of Astronomy, 28(4):337–345, November 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stahlman:1972:BRG

[Sta72] W. D. Stahlman. Book review: Greek mathematical astronomy: The Books of Autolykos: On a Moving Sphere and on Risings and Settings. Journal for the History of Astronomy, 3(2):147, June 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Standish:1995:BRK

[Sta95] E. M. Standish. Book review: Kepler’s equation down the centuries: Solving Kepler’s Equation: Over Three Centuries. Journal for the History of Astronomy, 26(2):175– 177, May 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:1999:BRG

[Sta99] Klaus Staubermann. Book review: The Genesis of the Meridian circle: Die Entwicklung des Meridiankreises 1700–1850: Genesis eines astronomischen Hauptinstrumentes unter Berücksichtigung des Wechselverhältnisses zwischen Astronomie, Astro-Technik und Technik. Journal for the History of Astronomy, 30(3): 330, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2000:BRA

[Sta00a] Klaus Staubermann. Book review: an amateur selenographer: Philipp Fauth: Leben und Werk. Journal for the History of Astronomy, 31(1):86, February 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2000:TIZ

[Sta00b]  Klaus Staubermann.  The

trouble with the instrument: Zöllner’s photometer. Journal for the History of Astronomy, 31(4):323–338, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2004:OCA

[Sta04]  Klaus Staubermann.  The

oldest collection of astronomical plates: Recreating photographic practice of the 1880s. Journal for the History of Astronomy, 35(4):447–456, November 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2006:EPR

[Sta06] Klaus Staubermann. Exercising patience: On the reconstruction of F. W. Bessel’s early star chart observations. Journal for the History of Astronomy, 37(1):19–36, February 2006. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Stanley:2007:BRT

[Sta07] Matthew Stanley. Book review: Testing relativity: Einstein’s Jury: The Race to Test Relativity. Journal for the History of Astronomy, 38(2):236– 237, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2010:BRA

[Sta10] Klaus Staubermann. Book review: Astronomical spectrographs: Astronomical Spectrographs and Their History. Journal for the History of Astronomy, 41(1):130–132, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stancliffe:2011:EAM

[Sta11] Terry Stancliffe. An early astronomical manuscript of Jeremiah Horrocks. Journal for the History of Astronomy, 42(4):455–460, November 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stanley:2013:BRW

[Sta13] Matthew Stanley. Book review: Was Einstein’s science Jewish?, Einstein’s Jewish Science: Physics at the Intersection of Politics and Religion. Journal for the History of Astronomy, 44(2):214–215, May 2013. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Staubermann:2015:PAG

[Sta15] Klaus Staubermann. Popularizing astronomy in Germany. Journal for the History of Astronomy, 46(4):493–494, November 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1977:BRC

[Ste77] F. Richard Stephenson. Book review: Chronicles and the Earth’s rotation: Ancient Astronomical Observations and the Accelerations of the Earth and Moon. Journal for the History of Astronomy, 8(1):63–64, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stevens:1988:BRM

[Ste88] Wesley M. Stevens. Book review: a medieval cosmology: De mundi celestis terrestrisque constitutione: a Treatise on the Universe and the Soul. Journal for the History of Astronomy, 19(1):62–63, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephens:1990:APU

[Ste90a] Carlene E. Stephens. Astronomy as public utility: The bond years at the Harvard College Observatory. Journal for the History of Astronomy, 21 (1):21–36, February 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1990:BRC

[Ste90b] Bruce Stephenson. Book review: Curtis Wilson’s Seventeenth-Century studies, Astronomy from Kepler to Newton: Historical Studies. Journal for the History of Astronomy, 21(4): 377–378, November 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1994:OAG

[Ste94] F. Richard Stephenson. Oriental astronomy from Guo Shoujing to King Sejong. Journal for the History of Astronomy, 25(2):152–153, May 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stevens:1996:BRQ

[Ste96] Wesley M. Stevens. Book review: The quadrivium in the early Middle Ages: Die Artes Liberales im frühen Mittelalter (5.–9. Jh.). Journal for the History of Astronomy, 27(4): 362–364, November 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1997:SET

[Ste97] J. M. Steele. Solar eclipse times predicted by the Babylonians. Journal for the History of Astronomy, 28(2):133– 139, May 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1998:PET

[Ste98a] John M. Steele. Predictions of eclipse times recorded in Chinese history. Journal for the History of Astronomy, 29(3): 275–285, August 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:1998:BRK

[Ste98b] Bruce Stephenson. Book review: Kepler’s Harmony translated: The Harmony of the World by Johannes Kepler. Journal for the History of Astronomy, 29(3):306–307, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:1999:BRV

[Ste99] John Steele. Book review:

Venus omens in Babylon: Babylonian Planetary Omens, Part 3. Journal for the History of Astronomy, 30(3):312– 313, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2000:BRC

[Ste00a] Bruce Stephenson. Book review: a celestial canvas: Painting the Heavens: Art and Science in the Age of Galileo. Journal for the History of Astronomy, 31(2):170–171, May 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2000:LOM

[Ste00b] F. Richard Stephenson. A lunar occultation of Mars observed by Aristotle. Journal for the History of Astronomy, 31(4):342–344, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2001:ERB

[Ste01] John M. Steele. Essay review: Babylonian astronomy reassessed, Mesopotamian Planetary Astronomy–Astrology. Journal for the History of Astronomy, 32(4):356–362, November 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2002:BSE

[Ste02a] John M. Steele. BM 36948: a Saturn Ephemeris calculated using system a from Babylon. Journal for the History of Astronomy, 33(3):261–264, August 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2002:BRC

[Ste02b] Bruce Stephenson. Book review: Composing the Astronomia Nova, The Composition of Kepler’s Astronomia nova. Journal for the History of Astronomy, 33(3):297–299, August 2002. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2003:PLB

[Ste03] J. M. Steele. Planetary latitudes in Babylonian mathematical astronomy. Journal for the History of Astronomy, 34 (3):269–289, August 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2004:AHA

[Ste04] J. M. Steele. Applied historical astronomy: an historical perspective. Journal for the History of Astronomy, 35(3):337– 355, August 2004. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2005:BRC

[Ste05] John Steele. Book review:

Crossing boundaries: Astronomy in the Orient and Occident: Selected Papers on its Cultural and Scientific History. Journal for the History of Astronomy, 36(3):351–352, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2006:BRJ

[Ste06a] John M. Steele. Book review:

Jupiter omens in Babylon: Babylonian Planetary Omens, Part 4. Journal for the History of Astronomy, 37(3):362– 363, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2006:BRG

[Ste06b] Bruce Stephenson. Book review: Greenwich astrolabes: Astrolabes at Greenwich. Journal for the History of Astronomy, 37(3):367–368, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2008:HRA

[Ste08a] F. Richard Stephenson. How reliable are archaic records of large solar eclipses? Journal for the History of Astronomy, 39(2):229–250, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stern:2008:BMN

[Ste08b] Sacha Stern. The Babylonian Month and the new Moon: Sighting and prediction. Journal for the History of Astronomy, 39(1):19–42, February 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2010:NSB

[Ste10a] J. M. Steele. A new scheme from Babylon for the synodic arc of Saturn. Journal for the History of Astronomy, 41 (2):261–268, May 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2010:IME

[Ste10b] F. Richard Stephenson. Investigation of medieval European records of solar eclipses. Journal for the History of Astronomy, 41(1):95–104, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steinicke:2011:WDS

[Ste11] Wolfgang Steinicke. A Website of deep-sky objects. Journal for the History of Astronomy, 42 (4):520, November 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2012:BRC

[Ste12a] John Steele. Book review: The chronology of old Babylon, Die Venustafeln des Ammi-· aduqa und ihre Bedeutung für die astronomische Datierung der altbabylonischen Zeit. Journal for the History of Astronomy, 43 (2):247–249, May 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2012:BRI

[Ste12b] Bruce Stephenson. Book review: Insights into Astronomia Nova: Kepler: La Physique Celeste, Autour de l’Astronomia Nova (1609). Journal for the History of Astronomy, 43(4): 505–506, November 2012. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stephenson:2012:BRM

[Ste12c] F. Richard Stephenson. Book review: The measurement of time, Time: From Earth Rotation to Atomic Physics. Journal for the History of Astronomy, 43(3):361–363, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2013:BRM

[Ste13] J. M. Steele. Book review: Moving beyond ‘act’, Babylonian Mathematical Astronomy: Procedure Texts. Journal for the History of Astronomy, 44 (3):363–364, August 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steinicke:2014:WHF

[Ste14] Wolfgang Steinicke. William Herschel, Flamsteed numbers and Harris’s star maps. Journal for the History of Astronomy, 45(3):287–303, August 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Steele:2016:SBA

[Ste16] John Steele. Stars in Babylonian astronomy. Journal for the History of Astronomy, 47(4):446–448, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828616670026.

Steele:2017:BRA

[Ste17] John Steele. Book review:

Astronomy in the Dead Sea Scrolls. Journal for the History of Astronomy, 48 (4):486–487, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617730696.

Steele:2019:ECS

[Ste19] John Steele. An early compilation of Saturn observations from Babylon. Journal for the History of Astronomy, 50(1):31–45, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828618820216.

Steele:2023:NTH

[Ste23] John Steele. Numerical tables in the history of astronomy. Journal for the History of Astronomy, 54(2):242– 244, May 1, 2023. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231168349.

Stigler:2007:BRL

[Sti07] Stephen M. Stigler. Book review: Laplace: Pierre Simon Laplace 1749–1827: a Determined Scientist. Journal for the History of Astronomy, 38(1):

131–132, February 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stooke:1994:NLM

[Sto94] Philip J. Stooke. Neolithic lunar maps at Knowth and Baltinglass, Ireland. Journal for the History of Astronomy, 25 (1):39–55, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stroup:1981:BRE

[Str81] Alice Stroup. Book review: Early modern clocks: The Clockwork Universe. Journal for the History of Astronomy, 12(3):209–211, October 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Strauss:1993:FFU

[Str93] David Strauss. “Fireflies flashing in unison”: Percival Lowell, Edward Morse and the birth of planetology. Journal for the History of Astronomy, 24(3): 157–169, August 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Strauss:1999:BRB

[Str99]  David Strauss.  Book re-

view: The biological universe abridged, Life on other Worlds: The 20th-Century Extraterrestrial Life Debate. Journal for the History of Astronomy, 30 (4):409–410, November 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Strano:2010:BRC

[Str10] Giorgio Strano. Book review:

Comets in the Seventeenth Century: Teorie sulle Comete da Galileo a Newton. Journal for the History of Astronomy, 41(4):523–525, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Stratan:2016:NDC

[Str16a] Gheorghe Stratan. On a newly discovered copy of Copernicus’ De Revolutionibus 1543 edition. Journal for the History of Astronomy, 47 (4):386–394, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616676720.

Strick:2016:ACB

[Str16b] James Strick. Astrobiology and cultural beliefs. Journal for the History of Astronomy, 47(1): 107–109, February 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/47/1/
107.full.pdf+html.

Strom:2019:BRB

[Str19] Richard Strom. Book review:

The birth of radio astronomy in The Netherlands. Journal for the History of Astronomy, 50(2):260–262, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619834391.

Strano:2023:TKD

[Str23] Giorgio Strano. Time-keeping devices and astronomy. Journal for the History of Astronomy, 54(3):366–367, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231191419.

Sullivan:1977:BRR

[Sul77] Woodruff T. Sullivan III. Book review: A radio astronomer reminisces, Big Ear. Journal for the History of Astronomy, 8(3):213, October 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sullivan:1999:BRL

[Sul99] Woodruff T. Sullivan III. Book review: Lauding the NRAO: The History of Radio Astronomy and the National Radio Astronomy Observatory: Evolution toward Big Science. Journal for the History of Astronomy, 30(1):79–81, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sullivan:2002:BRS

[Sul02] W. T. Sullivan III. Book review: The sociology of Jodrell Bank, science and spectacle: The Work of Jodrell Bank in Post-War British Culture. Journal for the History of Astronomy, 33(4):399–401, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sullivan:2012:BRM

[Sul12] Woody Sullivan. Book review: Mars in the spotlight, Geographies of Mars: Seeing and Knowing the Red Planet. Journal for the History of Astronomy, 43(3):368–370, August 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sullivan:2014:BRR

[Sul14] Woodruff Sullivan. Book review: Radio astronomy and internationalism: A Single Sky: How an International Community Forged the Science of Radio Astronomy. David P. D. Munns (The MIT Press, Cambridge, 2013). Pp. xii + 247. $34. ISBN 978-0-262-01833-3. Journal for the History of Astronomy, 45(4):483–485, November 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sullivan:2018:HSF

[Sul18] Wt Sullivan. Herschel’s story on film. Journal for the History of Astronomy, 49(2): 261–262, May 2018. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618764458.

Sutton:1972:BRE

[Sut72]  M. A. Sutton.  Book re-

view: Early theories of spectroscopy: Nineteenth Century Spectroscopy. Development of the Understanding of Spectra 1802–1897. Journal for the History of Astronomy, 3(1): 69–70, February 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Schmidl:2016:TEE

[SV16] Petra G. Schmidl and Flora Vafea. Two editions of the earliest text on the astrolabe. Journal for the History of Astronomy, 47 (4):449–451, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616670027.

Szubiakowski:2018:SDO

[SW18] Jacek P. Szubiakowski and Jarosław Włodarczyk. The solar dial in the Olsztyn Castle: Its construction and relation to Copernicus. Journal for the History of Astronomy, 49(2):158–195, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618776057.

Swenson:1970:MMM

[Swe70] Loyd S. Swenson, Jr. The Michelson–Morley–Miller experiments before and after 1905. Journal for the History of Astronomy, 1(1):56–78, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1972:APT

[Swe72] Noel Swerdlow. Aristotelian planetary theory in the Renaissance: Giovanni Battista Amico’s homocentric spheres. Journal for the History of Astronomy, 3(1):36–48, February 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1973:BRP

[Swe73] N. Swerdlow. Book review:

The planetary theory of Campanus: Campanus of Novara and Medieval Planetary Theory: Theorica Planetarum. Journal for the History of Astronomy, 4(1):59–61, February 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1974:OGC

[Swe74a] Noel Swerdlow. The origin of the Gregorian civil calendar. Journal for the History of Astronomy, 5(1):48–49, February 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1974:HRC

[Swe74b] Noel M. Swerdlow. The holograph of De Revolutionibus and the chronology of its composition. Journal for the History of Astronomy, 5(3):186–198, October 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1975:PTF

[Swe75] Noel M. Swerdlow. The planetary theory of François Viète. I. the fundamental planetary models. Journal for the History of Astronomy, 6(3):185– 208, October 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1976:ERK

[Swe76] N. M. Swerdlow. Essay review: Kepler and the theory of music: Der Musiktheoretiker Johannes Kepler (Neue Heidelberger Studien zur Musikwissenschaft, Band 5). Journal for the History of Astronomy, 7(3): 198–201, October 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1981:ETR

[Swe81] Noel M. Swerdlow. On establishing the text of “De Revolutionibus”. Journal for the History of Astronomy, 12(1):

35–46, February 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1984:RP

[Swe84] Noel M. Swerdlow. On the retrogradations of planets. Journal for the History of Astronomy, 15(1):30–32, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1986:ERR

[Swe86a] N. M. Swerdlow. Essay review:

Rheticus: Narrationes Prima et Secunda: G. J. Rheticus’ Treatise on Holy Scripture and the Motion of the Earth: With translation, annotations, commentary and additional chapters on Ramus–Rheticus and the development of the problem before 1650. Journal for the History of Astronomy, 17 (2):133–136, May 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1986:SCU

[Swe86b] N. M. Swerdlow. A star catalogue used by Johannes Bayer. Journal for the History of Astronomy, 17(3):189–197, August 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1986:LYO

[Swe86c] Noel M. Swerdlow. The length of the year in the original proposal for the Gregorian calendar. Journal for the History of Astronomy, 17(2):109–118, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1988:ERE

[Swe88] N. M. Swerdlow. Essay review: On editing Copernicus; Nicolaus Copernicus Gesamtausgabe, II: De Revolutionibus Kritischer Text. Journal for the History of Astronomy, 19 (3):208–211, August 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1989:PTI

[Swe89] N. M. Swerdlow. Ptolemy’s theory of the inferior planets. Journal for the History of Astronomy, 20(1):29–60, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1990:BRR

[Swe90]  N. M. Swerdlow.  Book

review: Renaissance correspondence, Trigonometrisch– Astronomisches Rechnen kurz vor Copernicus: Der Breifwechsel Regiomontanus–Bianchini. Journal for the History of Astronomy, 21(4):362–363, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1992:EPC

[Swe92a] N. M. Swerdlow. The enigma of Ptolemy’s catalogue of stars. Journal for the History of Astronomy, 23(3):173–183, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1992:ERA

[Swe92b] N. M. Swerdlow. Essay review: Amico’s homocentric spheres: Le Sfere Omocentriche di Giovan Battista Amico nell’Astronomia del Cinquecento: Con il testo del. Journal for the History of Astronomy, 23(3):211–215, August 1992. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1993:ONM

[Swe93] N. M. Swerdlow. Otto E.

Neugebauer: (26 May 1899– 19 February 1990). Journal for the History of Astronomy, 24(4):289–299, November 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1998:ERD

[Swe98] N. M. Swerdlow. Essay review: Documenta Copernicana: Nicolaus Copernicus Gesamtausgabe, VI/1: Briefe: Texte und Übersetzungen, Nicolaus Copernicus Gesamtausgabe, VI/2: Urkunden, Akten und Nachrichten: Texte und Übersetzungen. Journal for the History of Astronomy, 29(4): 386–389, November 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1999:ERE

[Swe99a] N. M. Swerdlow. Essay review: Empirical foundations of Babylonian lunar theory: Zur Entstehung der babylonischen Mondtheorie, Beobachtung und theoretischen Berechnung von Mondphasen. Journal for the History of Astronomy, 30(2): 169–172, May 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:1999:RCS

[Swe99b] N. M. Swerdlow. Regiomontanus’s concentric-sphere models for the Sun and Moon. Journal for the History of Astronomy, 30(1):1–23, February 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2000:KIS

[Swe00] N. M. Swerdlow. Kepler’s iterative solution to Kepler’s equation. Journal for the History of Astronomy, 31(4):339–341, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2001:ERN

[Swe01] N. M. Swerdlow. Essay review: The new Principia, the Principia: Mathematical principles of natural philosophy. Journal for the History of Astronomy, 32(4):349–356, November 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2004:EFP

[Swe04a] N. M. Swerdlow. The empirical foundations of Ptolemy’s planetary theory. Journal for the History of Astronomy, 35 (3):249–271, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2004:ERA

[Swe04b] N. M. Swerdlow. Essay review:

Alfonsine Tables of Toledo and later Alfonsine Tables: The Alfonsine Tables of Toledo, Archimedes. Journal for the History of Astronomy, 35(4): 479–484, November 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2004:GH

[Swe04c] N. M. Swerdlow. Galileo’s horoscopes. Journal for the History of Astronomy, 35(2): 135–141, May 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See erratum [Ano04a].

Swerdlow:2013:ERB

[Swe13] N. M. Swerdlow. Essay review: Babylonian astronomy and celestial divination: In the Path of the Moon: Babylonian Celestial Divination and its Legacy. Journal for the History of Astronomy, 44(1): 98–100, February 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Swerdlow:2017:CDH

[Swe17a] N. M. Swerdlow. Copernicus’s derivation of the heliocentric theory from Regiomontanus’s eccentric models of the second inequality of the superior and inferior planets. Journal for the History of Astronomy, 48(1):33–61, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617691203.

Swerdlow:2017:ERC

[Swe17b] N. M. Swerdlow. Early responses to Copernicus: an essential edition. Journal for the History of Astronomy, 48(2):242–246, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617706803.

Swerdlow:2017:MCH

[Swe17c] N. M. Swerdlow. The most complex homocentric astronomy. Essay review: Joseph Ibn Nah.mias, The Light of the World: Astronomy in al-Andalus. Edited, translated, and with a commentary by Robert G. Morrison (University of California Press, Berkeley, CA, 2016). Pp. xiv + 429. $95. ISBN 978-0-520-28799-0. Journal for the History of Astronomy, 48(3):358–361, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617721625.

Swerdlow:2021:AME

[Swe21] N. M. Swerdlow. The Almagest in the manner of Euclid. Journal for the History of Astronomy, 52(1): 104–107, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620977214.

Stephenson:1992:ARC

[SY92] F. Richard Stephenson and Kevin K. C. Yau. Astronomical records in the Ch’un-Ch’iu chronicle. Journal for the History of Astronomy, 23 (1):31–51, February 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Sylla:2007:BRR

[Syl07] Edith Dudley Sylla. Book review: Richard of Wallingford: God’s Clockmaker: Richard of Wallingford and the Invention of Time. Journal for the History of Astronomy, 38(2):238– 241, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Symons:2002:TFD

[Sym02] Sarah Symons. Two fragments of diagonal star clocks in the British museum. Journal for the History of Astronomy, 33 (3):257–260, August 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Symons:2015:CAE

[Sym15] Sarah Symons. Classification of ancient Egyptian astronomical ‘diagrams’. Journal for the History of Astronomy, 46 (1):66–75, February 2015. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Szapiro:2018:ABP

[Sza18]  Aníbal Szapiro.  Assump-

tions behind Ptolemy’s proof of the Earth’s centrality. Journal for the History of Astronomy, 49(2):196–215, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618765197.

Tursunov:2002:MOI

[TA02] O. S. Tursunov and S. H. Azizov. A medieval observational instrument in Tashkent. Journal for the History of Astronomy, 33(1):41–44, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tatarewicz:1999:BRE

[Tat99] Joseph N. Tatarewicz. Book review: European space missions: A New Force at a New Frontier: Europe’s Development in the Space Field in the Light of its Main Actors, Policies, Law and Activities from its Beginnings up to the Present. Journal for the History of Astronomy, 30(3): 337–338, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tatarewicz:2009:BRH

[Tat09] Joseph N. Tatarewicz. Book review: The Hubble space telescope: The Universe in a Mirror: The Saga of the Hubble Space Telescope and the Visionaries Who Built it. Journal for the History of Astronomy, 40(4):474–475, November 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:1994:BRD

[Tau94] Liba Taub. Book review: Diptych sundials: Ivory Diptych Sundials 1570–1750: a Catalogue of the Collection of Historical Scientific Instruments, Harvard University. Journal for the History of Astronomy, 25(1):65–67, February 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:1996:BRR

[Tau96] Liba Taub. Book review: Reappraising Philolaus: Philolaus of Croton: Pythagorean and Presocratic. Journal for the History of Astronomy, 27(1): 79–81, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:2009:BRA

[Tau09] Liba Taub. Book review: Ancient astrology in outline, A Brief History of Ancient Astrology. Journal for the History of Astronomy, 40(2):227– 229, May 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:2011:BRI

[Tau11] Liba Taub. Book review: Insights into de Caelo: New Perspectives on Aristotle’s De caelo. Journal for the History of Astronomy, 42(3):409– 411, August 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:2013:BRM

[Tau13] Liba Taub. Book review: Manilius elucidated, Forgotten Stars: Rediscovering Manilius’ Astronomica. Journal for the History of Astronomy, 44(3): 368–370, August 2013. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taub:2020:IL

[Tau20] Liba Taub. Instruments in the library. Journal for the History of Astronomy, 51(4): 495–497, November 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620945413.

Taylor:2010:PDM

[Tay10a] Katie Taylor. A “practique discipline”? Mathematical arts in John Blagrave’s The Mathematical Jewel (1585). Journal for the History of Astronomy, 41(3):329–353, August 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Taylor:2010:BRJ

[Tay10b] Stephanie L. Taylor. Book review: Joseph Cornell and astronomy: Joseph Cornell and Astronomy: a Case for the Stars. Journal for the History of Astronomy, 41(1):140–141, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tessicini:2013:CNE

[TB13] Dario Tessicini and Patrick Boner, editors. Celestial novelties on the eve of the scientific revolution: 1540–1630, volume 3 of Biblioteca di Galilana. L. S. Olschki, Firenze, Italia, 2013. ISBN 88-222-6254-9. xvi + 282 pp. LCCN QB29 .C45 2013. 32.00 EUR. URL http:

                                        //www.loc.gov/catdir/toc/
                                        casalini12/13017786.pdf.

Thury-Bouvet:2006:OCD

[TBSKO06] Ghjasippina Thury-Bouvet, Jean-François Santucci, El Hadi Khoumeri, and Antoine Ottavi. Orientations of Corsican dolmens. Journal for the History of Astronomy, 37(3):299– 306, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Teerikorpi:1989:LUN

[Tee89] Pekka Teerikorpi. Lundmark’s unpublished 1922 nebula classification. Journal for the History of Astronomy, 20(3): 165–170, October 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tenn:2000:BRC

[Ten00] Joseph S. Tenn. Book review: a century of astronomy: A History of Astronomy from 1890 to the Present. Journal for the History of Astronomy, 31 (2):179–180, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Teres:1984:TCD

[Ter84] Gustav Teres. Time computations and Dionysius Exiguus. Journal for the History of Astronomy, 15(3):177–188, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tessicini:2010:BRG

[Tes10] Dario Tessicini. Book review: Giordano Bruno: Giordano Bruno: Philosopher / Heretic. Journal for the History of Astronomy, 41(3):418– 419, August 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1977:IE

[TF77] A. S. Thom and T. R. Foord.

The Island of Eday. Journal for the History of Astronomy, 8(3): 198–199, October 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tihon:2016:NFP

[TF16] Anne Tihon and Jean-Luc Fournet. A new fragment of P. Fouad inv. 267A: The PSI inv. 2006. Journal for the History of Astronomy, 47 (4):355–358, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616666599.

Turchetta:1987:NCI

[TG87] M. Turchetta and G. Gavazzi.

Nineteenth-Century Italian contributions to galactic theory. Journal for the History of Astronomy, 18(3):196–208, August 1987. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1979:EMI

[TH79] Dean Turner and Richard Hazelett, editors. The Einstein myth and the Ives papers: a counter-revolution in physics: with excerpts from Ives’ correspondence, The Einstein myth by Dean Turner, a condensation of Euclid or Einstein by J. J. (Jeremiah Joseph) Callahan, and papers and comments by others. Devin-Adair Co., Old Greenwich, CT, USA, 1979. ISBN 0-8159-5823-4. vii + 313 pp. LCCN QC173.59.S65 I93 1979.

Thibodeau:1988:BRP

[Thi88] Sharon Gibbs Thibodeau. Book review: Pre-telescopic astronomy: Astronomy before the Telescope. Journal for the History of Astronomy, 19(1): 58–59, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1971:AER

[Tho71a] Victor E. Thoren. Anaxagoras, Eudoxus, and the regression of the lunar nodes. Journal for the History of Astronomy, 2 (1):23–28, February 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1971:BRB

[Tho71b] Victor E. Thoren. Book review: Brahe as astronomer and printer: Tycho Brahe: a Biography. Journal for the History of Astronomy, 2(3): 205–207, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1971:BRR

[Tho71c] Victor E. Thoren. Book review: The Rudolphine Tables: Johannes Kepler Gesammelte Werke, vol. x: Tabulae Rudolphinae. Journal for the History of Astronomy, 2(2):123– 124, June 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1973:NLT

[Tho73] Victor E. Thoren. New light on Tycho’s instruments. Journal for the History of Astronomy, 4 (1):25–45, February 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1974:MLO

[Tho74] A. Thom. A megalithic lunar Observatory in Islay. Journal for the History of Astronomy, 5 (1):50–51, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1984:BRF

[Tho84] Victor E. Thoren. Book review: Flamsteed and his star catalogue: The Preface to John Flamsteed’s Historia Coelestis Britannica or British Catalogue of the Heavens. Journal for the History of Astronomy, 15(1):58–59, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1986:BRC

[Tho86] Victor Thoren. Book review: Copernicus in context: Copernicus and the Scientific Revolution. Journal for the History of Astronomy, 17(2):143– 144, May 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1987:BRD

[Tho87] Victor E. Thoren. Book review: Distances in the heavens: Measuring the Universe: Cosmic Dimensions from Aristarchus to Halley. Journal for the History of Astronomy, 18(2):132– 133, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1988:BRE

[Tho88]  Victor E. Thoren.  Book

review: Elucidating Kepler’s physics: Kepler’s Physical Astronomy. Journal for the History of Astronomy, 19(4):280– 282, November 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1990:BRA

[Tho90a] Victor E. Thoren. Book review: Astrology in England, Prophecy and Power: Astrology in Early Modern England. Journal for the History of Astronomy, 21(4):378–379, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thoren:1990:BRW

[Tho90b] Victor E. Thoren. Book review: The Warburg Astrology Conference Volume, Astrology, Science and Society. Journal for the History of Astronomy, 21(4):370–371, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thompson:2022:TBV

[Tho22] Samantha M. Thompson.

Two biographies of Vera Rubin. Journal for the History of Astronomy, 53(3):370– 372, August 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221107618.

Thurston:1994:MTE

[Thu94] Hugh Thurston. A Mayan table of eclipses. Journal for the History of Astronomy, 25(19): S83–S84, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thurston:1995:TSL

[Thu95] Hugh Thurston. Three solar longitudes in the Almagest due to Hipparchus. Journal for the History of Astronomy, 26(2): 164, May 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano95b].

Thurston:2001:EGS

[Thu01] Hugh Thurston. Early Greek solstices and equinoxes. Journal for the History of Astronomy, 32(2):154–156, May 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thurston:2003:PAB

[Thu03] Hugh Thurston. Ptolemy and the arc between the tropics. Journal for the History of Astronomy, 34(2):231–232, May 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tihon:1998:AGG

[Tih98] Anne Tihon. The astronomy of George Gemistus Plethon. Journal for the History of Astronomy, 29(2):109–116, May 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tihon:2007:BRG

[Tih07] Anne Tihon. Book review: Geminos under scrutiny, Geminos’s Introduction to the Phenomena: a Translation and Study of a Hellenistic Survey of Astronomy. Journal for the History of Astronomy, 38(4):

505–507, November 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tihon:2017:BRB

[Tih17] Anne Tihon. Book review:

A Byzantine Exposition of Ptolemy. Journal for the History of Astronomy, 48 (4):492–494, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617730693.

Tihon:2021:LBA

[Tih21] Anne Tihon. Late Byzantine astronomy. Journal for the History of Astronomy, 52(4): 496–498, November 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211043318.

Timberlake:2020:BF

[Tim20] Todd K. Timberlake. The “big four”. Journal for the History of Astronomy, 51(3): 373–374, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620940172.

Timberlake:2025:DCR

[Tim25] Todd K. Timberlake. Documenting the Copernican Revolution. Journal for the History of Astronomy, 56(1):109– 110, February 2025. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241262281.

Tipler:1988:BRM

[Tip88a] Frank J. Tipler. Book review: More on Olbers’s paradox: Darkness at Night: a Riddle of the Universe. Journal for the History of Astronomy, 19(4):284–286, November 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tipler:1988:OPB

[Tip88b] Frank J. Tipler. Olbers’s Paradox, the beginning of creation, and Johann Mädler. Journal for the History of Astronomy, 19(1):45–48, February 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tirapicos:2022:BGB

[Tir22a] Luís Tirapicos. Bridging the gap between archaeology and archaeoastronomy. Journal for the History of Astronomy, 53(4):507–508, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221122935.

Tirapicos:2022:MNB

[Tir22b] Luís Tirapicos. A Muñoz biography. Journal for the History of Astronomy, 53(2): 230–231, May 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286221092362.

Thom:1978:SMS

[TM78] A. Thom and Robert L. Merritt. Some megalithic sites in Shetland. Journal for the History of Astronomy, 9(1):54–60, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toth:2002:HOP

[TMG02] Lásló Tóth, György Mezõ, and Ottó Gerlei. Haynald Observatory photosphere observations 1880–1919. Journal for the History of Astronomy, 33(3):278, August 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toomer:1977:ERP

[Too77] G. J. Toomer. Essay review:

Ptolemaic astronomy in Islam, Der Almagest: Die Syntaxis Mathematica des Claudius Ptolemäus in arabischlateinischer Überlieferung, Zur Kritik der Koordinateniiberlieferung im Sternkatalog des Almagest. Journal for the History of Astronomy, 8(3):204–210, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toomer:1981:ERC

[Too81] G. J. Toomer. Essay review: Copernicus in translation: Nicholas Copernicus on the Revolutions. Journal for the History of Astronomy, 12(3): 198–204, October 1981. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toomer:1984:OJS

[Too84] G. J. Toomer. Obituary: A. J. Sachs (1914–1983). Journal for the History of Astronomy, 15(2):146–149, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toomer:1995:ERN

[Too95] G. J. Toomer. Essay review: The new astronomy Englished: New Astronomy. Johannes Kepler, translated by William H. Donahue (Cambridge University Press, Cambridge, 1992). Pp. xvi + 665. £85/140. Journal for the History of Astronomy, 26(1):77–82, February 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Toomer:1996:BRH

[Too96] G. J. Toomer. Book review: Hypatia of Alexandria: Hypatia of Alexandria. Journal for the History of Astronomy, 27 (2):174–176, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Topper:2003:CGM

[Top03] David Topper. Colluding with Galileo: On Mueller’s critique of my analysis of Galileo’s sunspots argument. Journal for the History of Astronomy, 34 (1):75–77, February 2003. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tryapitsyn:2021:LLR

[TPYR21] V. N. Tryapitsyn, D. A. Pavlov, E. I. Yagudina, and V. V. Rumyantsev. The 1970–1984 lunar laser ranging observations in the Crimean Astrophysical Observatory. Journal for the History of Astronomy, 52(1):67–76, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828621989110.

Tang:2017:RMM

[TQ17] Quan Tang and Anjing Qu.

Research on the motion of Mars in the sovereign pole system (A.D. 600). Journal for the History of Astronomy, 48(4):432–459, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617734421.

Tang:2018:CCI

[TQ18] Quan Tang and Anjing Qu.

A comparison of Chinese and Indian solar theories in the Sui and Early Tang periods (seventh–eighth centuries C.E.). Journal for the History of Astronomy, 49(4): 472–495, November 1, 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618808876.

Trejo:1990:SOA

[Tre90] J. Galindo Trejo. Solar observations in Ancient Mexico: Malinalco. Journal for the History of Astronomy, 21(15):S17–S36, February 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tredwell:2004:MMF

[Tre04] Katherine A. Tredwell. Michael Maestlin and the fate of the Narratio Prima. Journal for the History of Astronomy, 35 (3):305–325, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Trimble:1997:BRS

[Tri97] Virginia Trimble. Book review: Source readings in cosmology: Modern Theories of the Universe from Herschel to Hubble. Journal for the History of Astronomy, 28(2):180– 181, May 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Trimble:2003:BRH

[Tri03] Virginia Trimble. Book review:

Historical supernovae updated: Historical Supernovae and Their Remnants. Journal for the History of Astronomy, 34 (1):119–121, February 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Trischler:2010:BRS

[Tri10] Helmuth Trischler. Book review: The space age: Remembering the Space Age: Proceedings of the 50 th Anniversary Conference. Journal for the History of Astronomy, 41(1): 135–137, February 2010. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Trimble:2013:BRF

[Tri13] Virginia Trimble. Book review: The first woman radio astronomer, Under the Radar: The First Woman in Radio Astronomy, Ruby Payne-Scott. Journal for the History of Astronomy, 44(2):222–223, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tusa:1992:OSP

[TSH92] Sebastiano Tusa, Giorgia Foderà Serio, and Michael Hoskin. Orientations of the Sesi of Pantelleria. Journal for the History of Astronomy, 23(17):S15–S20, February 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1971:ASL

[TT71] A. Thom and A. S. Thom. The astronomical significance of the large Carnac menhirs. Journal for the History of Astronomy, 2(3):147–160, October 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1972:CA

[TT72a] A. Thom and A. S. Thom. The Carnac alignments. Journal for the History of Astronomy, 3(1): 11–26, February 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1972:UAM

[TT72b] A. Thom and A. S. Thom. The uses of the alignments at Le Menec Carnac. Journal for the History of Astronomy, 3(3): 151–164, October 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1973:KC

[TT73a] A. Thom and A. S. Thom. The Kerlescan Cromlechs. Journal for the History of Astronomy, 4(3):168–173, October 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1973:MLO

[TT73b] A. Thom and A. S. Thom. A megalithic lunar observatory in Orkney: The Ring of Brogar and its cairns. Journal for the History of Astronomy, 4 (2):111–123, June 1973. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1974:KA

[TT74] A. Thom and A. S. Thom. The Kermario alignments. Journal for the History of Astronomy, 5 (1):30–47, February 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1975:FWB

[TT75] A. Thom and A. S. Thom. Further work on the Brogar lunar observatory. Journal for the History of Astronomy, 6 (2):100–114, June 1975. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1976:AIW

[TT76] Alexander Thom and Alexander Strang Thom. Avebury. (II). The West Kennet Avenue. Journal for the History of Astronomy, 7(3):193–197, October 1976. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1977:FLF

[TT77] A. Thom and A. S. Thom.

A fourth lunar foresight for the Brogar Ring. Journal for the History of Astronomy, 8(1): 54–55, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1978:RLS

[TT78] A. Thom and A. S. Thom. A reconsideration of the lunar sites in Britain. Journal for the History of Astronomy, 9(3): 170–179, October 1978. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1983:OMM

[TT83] A. Thom and A. S. Thom. Observation of the Moon in megalithic times. Journal for the History of Astronomy, 14(5): S57–S66, February 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1984:TMM

[TT84] A. Thom and A. S. Thom. The two major megalithic observatories in Scotland. Journal for the History of Astronomy, 15 (7):S129–S148, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1976:ANA

[TTF76] Alexander Thom, Archibald S. Thom, and T. R. Foord. Avebury (I): a new assessment of the geometry and metrology of the ring. Journal for the History of Astronomy, 7(3): 183–192, October 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1976:TML

[TTG76] Alexander Thom, Archibald S.

Thom, and J. M. Gorrie. The two megalithic lunar observatories at Carnac. Journal for the History of Astronomy, 7(1): 11–26, February 1976. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1974:S

[TTT74] Alexander Thom, Archibald Stevenson Thom, and Alexander Strang Thom. Stonehenge. Journal for the History of Astronomy, 5(2):71–90, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Thom:1975:SPL

[TTT75] Alexander Thom, Archibald Stevenson Thom, and Alexander Strang Thom. Stonehenge as a possible lunar observatory. Journal for the History of Astronomy, 6(1):19–30, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tucci:2000:BRL

[Tuc00] Pasquale Tucci. Book review:

Letters to Boscovich: Giovan Stefano Conti: Lettere a Ruggero Giuseppe Boscovich. Vol. II: (1771–1784). Journal for the History of Astronomy, 31 (3):276–277, August 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Tumanian:1974:MTA

[Tum74] Benik E. Tumanian. Measurement of time in Ancient and Mediaeval Armenia. Journal for the History of Astronomy, 5(2):91–98, June 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1970:BRP

[Tur70a] G. L’E. Turner. Book review: Prague instrument collection: Astronomy Gnomonics. A Catalogue of Instruments of the 15th to the 19th Centuries in the Collections of the National Technical Museum, Prague. Journal for the History of Astronomy, 1(1): 81–82, February 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1970:BRC

[Tur70b] G. L’E. Turner. Book review: The construction of sundials: Sundials. A Simplified Approach by Means of the Equatorial Dial. Journal for the History of Astronomy, 1(2): 155–156, August 1970. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1971:BRJ

[Tur71a] A. J. Turner. Book review:

John Wilkins: John Wilkins, 1614–1672: an Intellectual Biography. Journal for the History of Astronomy, 2(1):46–48, February 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1971:BRC

[Tur71b] G. L’E. Turner. Book review: China and the West: Clerks and Craftsmen in China and the West. Lectures and Addresses on the History of Science and Technology. Journal for the History of Astronomy, 2(2):120–121, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1972:BRM

[Tur72] G. L’E. Turner. Book review:

Measuring instruments: Time and Space. Measuring Instruments from the Fifteenth to the Nineteenth Century. Journal for the History of Astronomy, 3 (1):65–66, February 1972. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1975:ERT

[Tur75] A. J. Turner. Essay review:

The tragicall history of Giovanni De’ Dondi: The Planetarium of Giovanni de Dondi, Citizen of Padua. Journal for the History of Astronomy, 6 (2):126–131, June 1975. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1977:SCC

[Tur77] A. J. Turner. Some comments by Caroline Herschel on the use of the 40ft telescope. Journal for the History of Astronomy, 8(3):196–198, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1979:BRH

[Tur79a] G. L’E. Turner. Book review: Historic instruments in Stockholm, The Chamber of Physics. Instruments in the History of Sciences Collections of the Royal Swedish Academy of Sciences, Stockholm. Journal for the History of Astronomy, 10(2):130–131, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1979:NCR

[Tur79b] G. L’E. Turner. The number code on reflecting telescopes by Nairne and Blunt. Journal for the History of Astronomy, 10(3):177–184, October 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1984:BRE

[Tur84] G. L’E. Turner. Book review: Early testing of telescopes: Divini and Campani: a Forgotten Chapter in the History of the Accademia del Cimento. Journal for the History of Astronomy, 15(1):52–54, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1990:BRF

[Tur90a] A. J. Turner. Book review: The Farnese Planisferologio, clockwork Cosmos: Bernardo Facini and the Farnese Planisferologio. Journal for the History of Astronomy, 21(2):222–224, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1990:BRJ

[Túr90b] Gerard L’E. Túrner. Book review: James Ferguson, Wheelwright of the Heavens: The Life and Work of James Ferguson, FRS. Journal for the History of Astronomy, 21(3):303–304, August 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:1996:BRP

[Tur96] A. J. Turner. Book review:

Pèlerin’s treatise on the astrolabe: Pèlerin de Prusse on the Astrolabe. Journal for the History of Astronomy, 27(2):177– 180, May 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:2002:OQE

[Tur02]  A. J. Turner.  The ob-

servatory and the quadrant in Eighteenth-Century Europe. Journal for the History of Astronomy, 33(4):373–385, November 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:2003:BRM

[Tur03]  A. J. Turner.  Book re-

view: Medieval craftsmanship elucidated: Tractatus Astrarii. Journal for the History of Astronomy, 34(4):455–456, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:2007:BRA

[Tur07] Anthony Turner. Book review: Architectural contrasts: Les Riviera de Charles Garnier et Gustave Eiffel: Le Rêve de la Raison /Charles Garnier and Gustave Eiffel on the French and Italian Rivieras: The Dream of Reason. Journal for the History of Astronomy, 38(2):256–258, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:2015:CPA

[Tur15] Anthony Turner. Concerning a pointer on the astrolabe. Journal for the History of Astronomy, 46(4):413–418, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Turner:2023:SAA

[Tur23]  Anthony Turner.  A sur-

vey of Arabic astrolabe makers. Journal for the History of Astronomy, 54(1): 119–121, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221149856.

Tang:2025:REC

[TW25] Quan Tang and Zhenhua Wang. Research on the expansion-contraction difference and limit degree in ancient Chinese planetary theory: The case of outer planets. Journal for the History of Astronomy, 56(1):26–62, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241249776.

Urrutia-Aparicio:2021:EEK

[UAGGB21] Maitane Urrutia-Aparicio, A. César González-García, and Juan Antonio Belmonte. East or Easter? Keys to the orientation of Romanesque churches along the way of Saint James. Journal for the History of Astronomy, 52(3):289–310, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/00218286211026209.

Uso:2016:SMA

[UC16] María José Martínez Usó and Francisco J. Marco Castillo. A survey of meteoric activity over Spain during the eighth– fifteenth centuries. Journal for the History of Astronomy, 47 (2):168–193, May 2016. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Uso:2017:RDS

[UC17] María José Martínez Usó and Francisco J. Marco Castillo. A review of the Dall’Olmo survey of meteors, meteor showers and meteorites in the Middle Ages from medieval European sources. Journal for the History of Astronomy, 48(1):62–120, February 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617693081.

Uso:2019:OPM

[UC19] María José Martínez Usó and Francisco J. Marco Castillo. Occultation of planets by the Moon in European narrative medieval sources. Journal for the History of Astronomy, 50(2):192–220, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619845950.

Uso:2022:PRS

[UC22] María José Martínez Usó and Francisco J. Marco Castillo. A possible reference to the solar corona in a contemporary report of the AD1239 eclipse. Journal for the History of Astronomy, 53(4): 415–421, November 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221125189.

Uso:2023:TES

[UC23] María José Martínez Usó and Francisco J. Marco Castillo. The total eclipse of the sun of July 29, AD1478, in contemporary Spanish documents. Journal for the History of Astronomy, 54(2):153–170, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286231167157.

Uso:2016:OSS

[UCI16] María José Martínez Usó, Francisco J. Marco Castillo, and Loli Ibáñez. OSCURAUIT SOL: Stone engravings and other contemporary Spanish records for the A.D. 1239 and A.D. 1354 eclipses and their astronomical implications. Journal for the History of Astronomy, 47(1):61–75, February 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http:/

                                        /jha.sagepub.com/content/
                                        47/1/61.full.pdf+html.

Udias:2021:SOW

[Udí21] Agustín Udías. In Secchi’s own words. Journal for the History of Astronomy, 52(1): 111–112, February 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620979755.

University:1995:BRM

[UL95] Kansas State University and John Lankford. Book review: Materials for a history of the RASC: Looking up: a History of the Royal Astronomical Society of Canada. Journal for the History of Astronomy, 26(3): 275, August 1995. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Urton:2021:AT

[Urt21] Gary Urton. The astronomy of Tawantinsuyu. Journal for the History of Astronomy, 52(3):373–374, August 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211020938.

Usher:2015:SSA

[Ush15] Peter D. Usher. Shakespeare and Saturn: accounting for appearances, volume 41 of American University Studies XIX: General Literature. Peter Lang, New York, NY, USA, 2015. ISBN 1-4331-2860-8 (hardcover), 1-4539-1451-X (ebook). xix + 243 pp. LCCN PR3047 .U84 2015.

Vail:2012:BRD

[Vai12] Gabrielle Vail. Book review:

The Dresden Codex Venus pages: Tying Headbands or Venus Appearing: New Translations of k’al, the Dresden Codex Venus Pages and Classic Period Royal ‘Binding’ Rituals. Journal for the History of Astronomy, 43(4):503–505, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanDeVyverSJ:1971:AOS

[Van71a] O. Van De Vyver, S.J. Addendum: [Original sources of some early lunar maps]. Journal for the History of Astronomy, 2(2):97, June 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Van71b].

VanDeVyverSJ:1971:OSS

[Van71b] O. Van De Vyver, S.J. Original sources of some early lunar maps. Journal for the History of Astronomy, 2(2):86–97, June 1971. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See addendum [Van71a].

VanHelden:1974:ACS

[Van74a] Albert Van Helden. “Annulo Cingitur”: The solution of the problem of Saturn. Journal for the History of Astronomy, 5(3): 155–174, October 1974. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1974:SHA

[Van74b] Albert Van Helden. Saturn and his anses. Journal for the History of Astronomy, 5(2):105– 121, June 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1976:ITM

[Van76] Albert Van Helden. The importance of the transit of Mercury of 1631. Journal for the History of Astronomy, 7(1):1–10, February 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1977:DCE

[Van77] Albert Van Helden. The development of compound eyepieces, 1640–1670. Journal for the History of Astronomy, 8(1):26–37, February 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1985:BRV

[Van85] Albert Van Helden. Book review: Variations on a Galileian theme: Telescopes, Tides and Tactics. Journal for the History of Astronomy, 16(1):60– 61, February 1985. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1989:BRC

[Van89] Albert Van Helden. Book review: Copernicanism and Logic, Patterns, Thinking, and Cognition. Journal for the History of Astronomy, 20(3): 219–221, October 1989. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1990:BRM

[Van90] Albert Van Helden. Book review: Mayr’s Mundus Iovialis in German, Mundus Iovialis. Die Welt des Jupiter. Journal for the History of Astronomy, 21(4):371–372, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1992:BRP

[Van92] Albert Van Helden. Book review: a popular history of astronomy: Une Histoire de l’Astronomie. Journal for the History of Astronomy, 23(2): 151–152, May 1992. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanBrummelen:1994:LPI

[Van94a] Glen Van Brummelen. Lunar and planetary interpolation tables in Ptolemy’s Almagest. Journal for the History of Astronomy, 25(4):297–311, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1994:ERS

[Van94b] Albert Van Helden. Essay review: Sidereus Nuncius in French: Galileo Galilei: Le Messager des Étoiles, Galileo Galilei: Sidereus Nuncius, Le Messager Céleste. Journal for the History of Astronomy, 25 (1):56–59, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanAllen:1995:BRS

[Van95a] James A. Van Allen. Book review: Science from balloons: Race to the Stratosphere: Manned Scientific Ballooning in America. Journal for the History of Astronomy, 26 (2):177–179, May 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanAllen:1995:BRW

[Van95b] James A. Van Allen. Book review: War and peace: Science with a Vengeance: How the Military Created the US Space Sciences after World War II. Journal for the History of Astronomy, 26(2):179–181, May 1995. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1996:BRC

[Van96] Albert Van Helden. Book review: Cassini researched: Gio: Domenico Cassini: Uno Scienziato del Seicento. Journal for the History of Astronomy, 27 (2):182–183, May 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vanderburgh:1997:EEA

[Van97] William L. Vanderburgh. Empirical equivalence and approximative methods in the New Astronomy: a defence of Kepler against the charge of fraud. Journal for the History of Astronomy, 28(4):317–336, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanBrummelen:1998:CAP

[Van98a] Glen Van Brummelen. Computer animations of Ptolemy’s models of the motions of the Sun, Moon and planets. Journal for the History of Astronomy, 29(3):271–274, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1998:BRD

[Van98b] Albert Van Helden. Book review: a Dutch farmer– philosopher: Speculatie, Wetenschap en Vernuft: Fysica en astronomie volgens Wytze Foppes Dongjuma (1707–1778), Instrumentmaker te Leeuwarden. Journal for the History of Astronomy, 29(3):301–302, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:1998:BRP

[Van98c] Albert Van Helden. Book review: Pictures that tell a story: The Cambridge Illustrated History of Astronomy. Journal for the History of Astronomy, 29 (1):83–84, February 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanGent:1999:BRA

[Van99] Robert H. Van Gent. Book review: Adler astrolabes: Western Astrolabes: Historic Scientific Instruments of the Adler Planetarium & Astronomy Museum. Journal for the History of Astronomy, 30(3):316– 317, August 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanGent:2002:BRN

[Van02a] Robert Van Gent. Book review: Neglected Dutch amateurs, Van ‘Konstgenoten’ en Hemelse Fenomenen: Nederlandse Sterrenkunde in de Achttiende Eeuw. Journal for the History of Astronomy, 33 (1):85–86, February 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanHelden:2002:ERM

[Van02b] Albert Van Helden. Essay review: Moon maps in history, Mapping and Naming the Moon: a History of Lunar Cartography and Nomenclature. Journal for the History of Astronomy, 33(3):285– 289, August 2002. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VandenBroecke:2006:TPT

[Van06] Steven Vanden Broecke. Teratology and the publication of Tycho Brahe’s new world system (1588). Journal for the History of Astronomy, 37(1): 1–17, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanDuzer:2010:JSG

[Van10] Chet A. Van Duzer. Johann Schoner’s globe of 1515: transcription and study, volume 100 (part 5) of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 2010. ISBN 1-60618-005-3 (paperback). vi + 217 pp. LCCN GA271 .V36 2010.

VanHelden:2014:UGF

[Van14a] Albert Van Helden. Unmasking a Galileo forgery. Journal for the History of Astronomy, 45(3):361–376, August 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Ano14a].

Vanderburgh:2014:PNS

[Van14b] William L. Vanderburgh. Putting a new spin on galaxies: Horace W. Babcock, the Andromeda nebula, and the dark matter revolution. Journal for the History of Astronomy, 45(2):141– 159, May 2014. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vanin:2015:MBT

[Van15] Gabriele Vanin. Much better than thought: Observing with Galileo’s telescopes. Journal for the History of Astronomy, 46(4):441–468, November 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Broecke:2023:NET

[Van23] Steven Vanden Broecke. A new edition and translation of Pico’s Disputationes. Journal for the History of Astronomy, 54(3):368–369, August 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231175987.

VanBrummelen:2024:IIZ

[Van24] Glen Van Brummelen. An important Islamicate Zij. Journal for the History of Astronomy, 55(1):??, February 1, 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231226225.

Vargha:1987:ARH

[VB87] Magda Vargha and Előd Both. Astronomy in Renaissance Hungary. Journal for the History of Astronomy, 18(4):279– 283, November 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

VanBrummelen:2001:ASA

[VB01] Glen Van Brummelen and J. L.

Berggren. Abū Sahl al-Kūhī on the distance to the shooting stars. Journal for the History of Astronomy, 32(2):137– 151, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanBrummelen:2010:BRB

[vB10] Glen van Brummelen. Book review: Bianchini’s Tabulae Astronomiae: The Astronomical Tables of Giovanni Bianchini. Journal for the History of Astronomy, 41(4):514–516, November 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vandenBroecke:2017:BRI

[vB17] Steven vanden Broecke. Book review: An Introduction to Astrology in the Renaissance. Journal for the History of Astronomy, 48 (4):494–496, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616670028.

VanBrummelen:2021:MRB

[VB21] Glen Van Brummelen and James Steven Byrne. Maurolico, Rheticus, and the birth of the secant function. Journal for the History of Astronomy, 52(2):189–211, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211013202.

vanderWaerden:1974:EFE

[vdW74] B. L. van der Waerden. The earliest form of the epicycle theory. Journal for the History of Astronomy, 5(3):175– 185, October 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanderWaerden:1978:BRG

[vdW78] B. L. van der Waerden. Book review: Greek astronomy, Theorien und Ansätze in der griechischen Astronomie. Journal for the History of Astronomy, 9(2):149–150, June 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanderWaerden:1980:TTI

[vdW80] B. L. van der Waerden. Two treatises on Indian astronomy. Journal for the History of Astronomy, 11(1):50–62, February 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). With a reply by David Pingree.

vanderWaerden:1990:BRM

[vdW90] B. L. van der Waerden.

Book review: Mithras and the Bull, the Origins of the Mithraic Mysteries: Cosmology and Salvation in the Ancient World. Journal for the History of Astronomy, 21(4):365–366, November 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanderWerf:2017:HCA

[vdW17] Siebren van der Werf. History and critical analysis of fifteenth and sixteenth century nautical tables. Journal for the History of Astronomy, 48(2):207–232, May 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617705244.

vanderWerf:2019:NTV

[vdW19] Siebren van der Werf. Nautical tables for Vasco da Gama, 1497–1500? Journal for the History of Astronomy, 50 (3):326–338, August 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619864472.

Verdet:1987:BRA

[Ver87] J. P. Verdet. Book review:

Astronomy in Provence: Astronomie et Astronomes en Provence, 1680–1730. Journal for the History of Astronomy, 18(2):134–135, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vermij:2006:ALC

[Ver06] Rienk Vermij. Albertus Leoninus (1543–1614) and Copernicus’s “third motion” of the Earth. Journal for the History of Astronomy, 37(1):101–109, February 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vermij:2009:BRG

[Ver09] Rienk Vermij. Book review: Gemma Frisius on Earth and sky, Gemma Frisius, Arpenteur de la Terre et du Ciel. Journal for the History of Astronomy, 40(2):229–230, May 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vermij:2012:BRD

[Ver12] Rienk H. Vermij. Book review: On the date of Easter: Dating the Passion: The Life of Jesus and the Emergence of Scientific Chronology (200– 1600). Journal for the History of Astronomy, 43(4):502–503, November 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Verbunt:2013:BRH

[Ver13] Frank Verbunt. Book review: The Hipparcos satellite, The Making of History’s Greatest Star Map. Journal for the History of Astronomy, 44(2):217– 218, May 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vermij:2017:BRP

[Ver17] Rienk Vermij. Book review:

A Polymathic Astronomer of the Sixteenth Century. Journal for the History of Astronomy, 48(4):482–483, November 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828617730695.

Veselovsky:1973:CNA

[Ves73] I. N. Veselovsky. Copernicus and Nas.īr al-Dīn al-T .ūsī. Journal for the History of Astronomy, 4(2):128–130, June 1973. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vescovini:1998:PSM

[Ves98] Graziella Federici Vescovini.

The place of the Sun in medieval Arabo–Latin astronomy: The Lucidator Dubitabilium Astronomiae (1303–10) of Peter De Padua. Journal for the History of Astronomy, 29 (2):151–155, May 1998. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vetter:2012:BRM

[Vet12] Jeremy Vetter. Book review:

Measuring longitude by telegraphy: Finding North America: Longitude by Wire. Journal for the History of Astronomy, 43(1):123–125, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vetter:2019:BRA

[Vet19] Jeremy Vetter. Book review: American Eclipse: a Nation’s Epic Race to Catch the Shadow of the Moon and Win the Glory of the World, David Baron, (Liveright Publishing Corporation, New York, 2017). Pp. xii + 330. $38. ISBN 978-1-63149-016-3. Journal for the History of Astronomy, 50(1): 114–115, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619826758.

Voelkel:2001:GAM

[VG01] James R. Voelkel and Owen Gingerich. Giovanni Antonio Magini’s “Keplerian” tables of 1614 and their implications for the reception of Keplerian astronomy in the Seventeenth Century. Journal for the History of Astronomy, 32(3):237– 262, August 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). Astronomy and its histories (Vancouver, BC, 2000).

vanGent:2005:BRO

[vG05] Robert H. van Gent. Book review: an overview of Mayan astronomy: Das mathematische Weltbild der Maya. Journal for the History of Astronomy, 36(2):234, May 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanHelden:1983:RSL

[vH83] Albert van Helden. Roemer’s speed of light. Journal for the History of Astronomy, 14 (2):137–141, June 1983. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanHelden:1986:BRO

[vH86] Albert van Helden. Book review: Observations of Venus’s 1769 transit: The 1769 Transit of Venus: The Baja California Observations of Jean-Baptiste Chappe d’Auteroche, Vicente de Doz, and Joaquín Velazquez Cárdenas de León. Journal for the History of Astronomy, 17 (2):146–147, May 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

vanHelden:1987:BRC

[vH87] Albert van Helden. Book review: Comets in the Renaissance: Cometary Theory in Fifteenth-Century Europe. Journal for the History of Astronomy, 18(4):290–291, November 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Vickers:1985:BRS

[Vic85] Brian Vickers. Book review: Science in Oxford and Cambridge: The Mathematician’s Apprenticeship: Science, Universities and Society in England 1560–1640. Journal for the History of Astronomy, 16 (1):56–59, February 1985. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Viladrich:1988:PLT

[Vil88] Mercè Viladrich. The planetary latitude tables in the Mumtah.an Zīj. Journal for the History of Astronomy, 19(4): 257–268, November 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Viladrich:1990:BRC

[Vil90] Mercè Viladrich. Book review: Cairo manuscripts, A Survey of the Scientific Manuscripts in the Egyptian National Library. Journal for the History of Astronomy, 21(2):218–219, May 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voulgaris:2018:CFR

[VMV18] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. Conclusions from the functional reconstruction of the Antikythera mechanism. Journal for the History of Astronomy, 49(2):216–238, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828618762460.

vanNetten:2012:AIT

[vN12] Djoeke van Netten. Astronomia instaurata? The third edition of Copernicus’s De Revolutionibus (Amsterdam, 1617). Journal for the History of Astronomy, 43(1):75–91, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:1994:BRG

[Voe94] James R. Voelkel. Book review: Galileian rhetoric: Novelties in the Heavens: Rhetoric and Science in the Copernican Controversy. Journal for the History of Astronomy, 25(1): 69–70, February 1994. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:1997:ERK

[Voe97] James R. Voelkel. Essay review: Keplerian miscellany: Johannes Kepler: Gesammelte Werke, Band IX, 2: Calendaria et Prognostica, Astronomica Minora, Somnium. Journal for the History of Astronomy, 28(1):83–86, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:2001:BRT

[Voe01] James R. Voelkel. Book review: Teaching ancient astronomy: The History and Practice of Ancient Astronomy. Journal for the History of Astronomy, 32(1):82–84, February 2001. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:2002:BRS

[Voe02]  James R. Voelkel.  Book

review: The sphericity of the Earth, Die Vorstellung von der Kugelgestalt der Erde in europäischen Mittelalter bis zum Ende des 13. Jahrhunderts — Dargestellt nach den Quellen. Journal for the History of Astronomy, 33(4):397– 398, November 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:2003:BRH

[Voe03] James R. Voelkel. Book review: History of astronomy in Germany: Beiträge zur Astronomiegeschichte, I, II and IV (Acta Historica Astronomiae, I, V and XIII). Journal for the History of Astronomy, 34(3):338–339, August 2003. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Voelkel:2007:BRK

[Voe07] James R. Voelkel. Book review: a Kepler miscellany, Johannes Kepler Gesammelte Werke. Journal for the History of Astronomy, 38(4):513–515, November 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Ventura:1993:PTS

[VSH93] Frank Ventura, Giorgia Foderà Serio, and Michael Hoskin. Possible tally stones at Mnajdra, Malta. Journal for the History of Astronomy, 24(3): 171–183, August 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Verbunt:2019:WHD

[VvdS19] Frank Verbunt and Marc van der Sluys. Why Halley did not discover proper motion and why Cassini did. Journal for the History of Astronomy, 50(4):383–397, November 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619877967.

Valluri:1997:NAP

[VWH97] Sree Ram Valluri, Curtis Wilson, and William Harper. Newton’s apsidal precession theorem and eccentric orbits. Journal for the History of Astronomy, 28(1):13–27, February 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:1986:BRB

[Waf86a] Craig B. Waff. Book review: Bibliographies of Halley’s comet: Halley’s Comet: a Bibliography, Halley’s Comet, 1755–1984: a Bibliography. Journal for the History of Astronomy, 17(1):60–62, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:1986:CHF

[Waf86b] Craig B. Waff. Comet Halley’s first expected return: English public apprehensions, 1755–58. Journal for the History of Astronomy, 17(1):1–37, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:1991:BRH

[Waf91] Craig B. Waff. Book review: How the space telescope came to be, The Space Telescope: a Study of NASA, Science, Technology, and Politics. Journal for the History of Astronomy, 22(3):250–252, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:2001:ERP

[Waf01] Craig B. Waff. Essay review: a Prime Meridian for the United States?: The Jefferson Stone: Demarcation of the First Meridian of the United States. Journal for the History of Astronomy, 32(2):157– 159, May 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:2002:BRS

[Waf02] Craig B. Waff. Book review:

Selling time in the United States, Selling the True Time: Nineteenth-Century Timekeeping in America. Journal for the History of Astronomy, 33(4): 398–399, November 2002. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Waff:2004:BRS

[Waf04] Craig B. Waff. Book review:

Science amidst the snows: The Man Who Flattened the Earth: Maupertuis and the Sciences in the Enlightenment. Journal for the History of Astronomy, 35 (3):363–365, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wagman:1987:FMS

[Wag87] Morton Wagman. Flamsteed’s missing stars. Journal for the History of Astronomy, 18(3): 209–223, August 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wagman:1988:BRS

[Wag88] Morton Wagman. Book review:

Stellar identification: Synopsis of the Nomenclature of the Fixed Stars. Journal for the History of Astronomy, 19(1):

59–61, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wagman:1992:HC

[Wag92] Morton Wagman. Hercules, the champion. Journal for the History of Astronomy, 23(2):134– 136, May 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wallace:1977:BRU

[Wal77] William A. Wallace. Book review: The UCLA Copernicus Symposium, The Copernican Achievement. Journal for the History of Astronomy, 8(3): 211–213, October 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wallis:1999:BRE

[Wal99] Faith Wallis. Book review: Early medieval astronomies: Astronomies and Cultures in Early Medieval Europe. Journal for the History of Astronomy, 30(3):317–319, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Walker:2006:AEP

[Wal06] Merle F. Walker. The alignment of the Egyptian pyramids. Journal for the History of Astronomy, 37(3):349–352, August 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wall:2007:SAH

[Wal07] John J. Wall. The star alignment hypothesis for the Great Pyramid shafts. Journal for the History of Astronomy, 38 (2):199–206, May 2007. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wallis:2013:BRC

[Wal13] Faith Wallis. Book review:

Computus: Computus and its Cultural Context in the Latin West, AD 300–1200. Proceedings of the 1st International Conference on the Science of Computus in Ireland and Europe, Galway, 14–16 July, 2006. Journal for the History of Astronomy, 44(1):101–104, February 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1971:BRH

[War71] Deborah J. Warner. Book review: Hevelius’s star atlas: The Star Atlas. Journal for the History of Astronomy, 2(2):125– 126, June 1971. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1977:AAS

[War77] Brian Warner. Astronomical archives in southern Africa. Journal for the History of Astronomy, 8(3):217–222, October 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1978:CGH

[War78] Brian Warner. Cape of Good Hope Royal Observatory papers in the archives of the Royal Greenwich Observatory. Journal for the History of Astronomy, 9(1):74–75, February 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1979:BRS

[War79] Brian Warner. Book review: The study of southern skies, The Astronomy of Southern Africa. Journal for the History of Astronomy, 10(1):63– 65, February 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1980:BRD

[War80] Deborah Jean Warner. Book review: a distinguished amateur, Starlight Nights. Journal for the History of Astronomy, 11(3):220, October 1980. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:1993:BRL

[War93] Brian Warner. Book review: Lacaille re-established: Lacaille: Astronomer, Traveler: With a New Translation of His Journal. Journal for the History of Astronomy, 24(1–2): 152–153, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.

sagepub.com/content/24/1-
2/152.full.pdf+html.

Warner:1998:BRT

[War98] Deborah Jean Warner. Book review: Telescope-building in Nineteenth-Century Ireland: Victorian Telescope Makers: The Lives and Letters of Thomas and Howard Grubb. Journal for the History of Astronomy, 29(3):304–305, August 1998. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warner:2008:BRT

[War08] Deborah Jean Warner. Book review: Treasures of the Dresden Kunstkammer, the Geometry of Power /the Power of Geometry: Mathematical Instruments and Princely Mechanical Devices from around 1600. Journal for the History of Astronomy, 39(3):414–415, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Warntjes:2013:BRG

[War13] Immo Warntjes. Book review: The Gregorian calendar reform: Die Gregorianische Kalenderreform von 1582: Korrektur der christlichen Zeitrechnung in der Frühen Neuzeit. Journal for the History of Astronomy, 44(1):115– 117, February 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wartenberg:2015:TMH

[War15] Ilana Wartenberg. Two Medieval Hebrew astrological classics. Journal for the History of Astronomy, 46(1):106–108, February 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wartenberg:2019:BRI

[War19] Ilana Wartenberg. Book review: Ibn Ezra on astrology: Abraham Ibn Ezra’s Introductions to Astrology: A Parallel Hebrew-English Critical Edition of the Book of the Beginnings of Wisdom and the Book of the Judgments of the Zodiacal Signs Shlomo Sela, (Brill, Leiden, 2017). Pp. xiii + 822. EUR 225. ISBN 978-90-04-34227-9. Journal for the History of Astronomy, 50 (1):108–109, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828618823979.

Wasserstein:1972:ERE

[Was72] A. Wasserstein. Essay review: Early Greek astronomy: Early Greek Astronomy to Aristotle. Journal for the History of Astronomy, 3(3):212–216, October 1972. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wasserstein:1977:BRP

[Was77] A. Wasserstein. Book review: Pythagoras, Lore and Science in Ancient Pythagoreanism. Journal for the History of Astronomy, 8(2):137– 140, June 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wayman:1996:BRC

[Way96] P. A. Wayman. Book review:

Chronicling the IAU: History of the IAU: The Birth and First Half-Century of the International Astronomical Union. Journal for the History of Astronomy, 27(3):278–279, August 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wesemael:2011:JWJ

[WDR11] François Wesemael, Karl Del Duchetto, and René Racine. From J. Winthrop, Jr., to E. E. Barnard: The arduous path to the first sighting of Amalthea. Journal for the History of Astronomy, 42(2):125–139, May 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Weart:1990:BRA

[Wea90] Spencer R. Weart. Book review: Aspects of geophysical research, Past, Present and Future Trends in Geophysical Research. Journal for the History of Astronomy, 21(3):309– 310, August 1990. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Weimer:1982:IVD

[Wei82a]  Th. Weimer.  Une instru-

ment en voie de disparition: l’équatorial coude. (French) [An endangered instrument: the elbow equatorial]. Journal for the History of Astronomy, 13(2):110–118, June 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Weir:1982:VTF

[Wei82b] John D. Weir. The Venus tablets: a fresh approach. Journal for the History of Astronomy, 13(1):23–49, February 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Welther:1984:BRT

[Wel84] Barbara L. Welther. Book review: Topics in brief: Dictionary of the History of Science. Journal for the History of Astronomy, 15(2):138–139, June 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Welther:1994:BRL

[Wel94] Barbara L. Welther. Book review: Lady Herschel at the Cape: Lady Herschel: Letters from the Cape 1834– 1838. Journal for the History of Astronomy, 25(4):330–331, November 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Welther:2005:BRL

[Wel05] Barbara Welther. Book review: a life of Maria Mitchell: Maria Mitchell: a Life in Journals and Letters. Journal for the History of Astronomy, 36 (1):117, February 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wentzel:1999:BRP

[Wen99] Donat G. Wentzel. Book review: In piam memoriam Edith Müller: Remembering Edith Alice Müller. Journal for the History of Astronomy, 30(3): 335–336, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wendt:2024:ISN

[Wen24]  Harry Wendt.  An inside

story of the NRAO. Journal for the History of Astronomy, 55(3):401–402, August 2024. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241239536.

Werrett:2010:BRR

[Wer10] Simon Werrett. Book review:

Russian astronomy surveyed: Nebo v Zemnom Otrazhenii: Istoriia Astronomii v Rossii v XIX — Nachale XX veka. Journal for the History of Astronomy, 41(2):274–276, May 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Werrett:2013:BRP

[Wer13] Simon Werrett. Book review: Paradoxes in Seventeenth-Century science, Baroque Science. Journal for the History of Astronomy, 44(4):495–496, November 2013. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1977:BRN

[Wes77a] Richard S. Westfall. Book review: Newton on the lunar motion: Isaac Newton’s “Theory of the Moon’s Motion”. Journal for the History of Astronomy, 8 (1):67–68, February 1977. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westman:1977:BRR

[Wes77b] Robert S. Westman. Book review: “De Revolutionibus” in English, Copernicus: On the Revolutions of the Heavenly Spheres. Journal for the History of Astronomy, 8(2):145– 147, June 1977. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wesley:1978:ATB

[Wes78] Walter G. Wesley. The accuracy of Tycho Brahe’s instruments. Journal for the History of Astronomy, 9(1):42–53, February 1978. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wesley:1979:TBS

[Wes79] Walter G. Wesley. Tycho Brahe’s solar observations. Journal for the History of Astronomy, 10(2):96–101, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1984:BRG

[Wes84] Richard S. Westfall. Book review: Galileo’s so-called “juvenilia”: Galileo’s Early Notebooks: The Physical Questions. Journal for the History of Astronomy, 15(3):220–221, October 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1989:TGB

[Wes89] Richard S. Westfall. The trial of Galileo: Bellarmino, Galileo, and the clash of two worlds. Journal for the History of Astronomy, 20(1):1–23, February 1989. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1990:BRV

[Wes90] Richard S. Westfall. Book review: Volume 2A of the GHA, The General History of Astronomy, II: Planetary Astronomy from the Renaissance to the Rise of Astrophysics, Part A: Tycho Brahe to Newton. Journal for the History of Astronomy, 21(4):372–373, November 1990. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1991:VT

[Wes91] Richard S. Westfall. Victor E. Thoren (1935–1991). Journal for the History of Astronomy, 22(3):253–254, August 1991. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westman:1993:ERP

[Wes93] Robert S. Westman. Essay review: a poetics of the Cosmos: The Poetic Structure of the World: Copernicus and Kepler. Journal for the History of Astronomy, 24(1–2): 142–145, May/February 1993. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

                                        tronic). URL http://jha.
                                        sagepub.com/content/24/1-
                                        2/142.full.pdf+html.

Westfall:1994:ERK

[Wes94] Richard S. Westfall. Essay review: Kepler and Kepler’s laws: The Discovery of Kepler’s Laws: The Interaction of Science, Philosophy, and Religion. Journal for the History of Astronomy, 25(3):244–245, August 1994. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westfall:1996:BRG

[Wes96] R. S. Westfall. Book review: Galileo in the cross-fire: Galileo and the Church: Political Inquisition or Critical Dialogue? Journal for the History of Astronomy, 27(2):180–181, May 1996. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westman:1997:ERZ

[Wes97] Robert S. Westman. Essay review: Zinner, Copernicus, and the Nazis: Entstehung und Ausbreitung der copernicanischen Lehre. Ernst Zinner, ed. by. Heribert M. Nobis and Felix Schmeidler (Verlag C. H. Beck, Munich, 1988; first edition, Sitzungberichte der physikalisch-medizinischen Sozietdt zu Erlangen, lxxiv (1943)). Pp. xvi + 636. DM 198. Journal for the History of Astronomy, 28(3):259–270, August 1997. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See errata [Ano97b].

Westman:2001:KEP

[Wes01] Robert S. Westman. Kepler’s early physical–astrological problematic. Journal for the History of Astronomy, 32(3):227– 236, August 2001. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). Astronomy and its histories (Vancouver, BC, 2000).

Wesemael:2009:HPA

[Wes09a] François Wesemael. Harkins, Perrin and the alternative paths to the solution of the stellar-energy problem, 1915– 1923. Journal for the History of Astronomy, 40(3):277–296, August 2009. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Westman:2009:ERC

[Wes09b] Robert S. Westman. Essay review: Cognizing the Copernican revolution, The Cognitive Structure of Scientific Revolutions. Journal for the History of Astronomy, 40(3):339– 345, August 2009. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Westman:2018:CT

[Wes18] Robert S. Westman. Copernican transformations. Journal for the History of Astronomy, 49(1):120–123, February 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828617748379.

Whiteside:1970:BPM

[Whi70a]  D. T. Whiteside.  Before

the Principia: The maturing of Newton’s thoughts on dynamical astronomy, 1664–1684. Journal for the History of Astronomy, 1(1):5–19, February 1970. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1970:MPU

[Whi70b] D. T. Whiteside. The mathematical principles underlying Newton’s Principia Mathematica. Journal for the History of Astronomy, 1(2):116–138, August 1970. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitrow:1971:BRH

[Whi71] G. J. Whitrow. Book review:

Heaven and the stars: An Original Theory or New Hypothesis of the Universe, a Theory of the Universe. Journal for the History of Astronomy, 2(3): 208–209, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1972:BRS

[Whi72] D. T. Whiteside. Book review:

Seventeenth-Century planetary theory: Hypothesen zur Planetentheorie des 17. Jahrhunderts. Journal for the History of Astronomy, 3(1):66–67, February 1972. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitaker:1973:BRE

[Whi73a] Ewen A. Whitaker. Book review: Early lunar maps: “Lunar Maps of the Seventeenth Century”. Journal for the History of Astronomy, 4(2):140– 141, June 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1973:BRC

[Whi73b] D. T. Whiteside. Book review:

Cartesian vortices: The Vortex Theory of Planetary Motions. Journal for the History of Astronomy, 4(1):62–63, February 1973. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1974:KPE

[Whi74] D. T. Whiteside. Keplerian planetary eggs, laid and unlaid, 1600–1605. Journal for the History of Astronomy, 5(1):1–21, February 1974. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1975:RCP

[Whi75] D. T. Whiteside. A refined computation of the perigee angle in Ptolemy’s Mercury model. Journal for the History of Astronomy, 6(1):57, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitaker:1976:BRL

[Whi76a] Ewen A. Whitaker. Book review: Lunar mapping, Mapping of the Moon — Past and Present. Journal for the History of Astronomy, 7(1):65–67, February 1976. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whiteside:1976:ERY

[Whi76b] D. T. Whiteside. Essay review: The year of Kepler: Kepler: Four Hundred Years. Proceedings of Conferences Held in Honour of Johannes Kepler. Journal for the History of Astronomy, 7(1):57–61, February 1976. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitaker:1978:GLO

[Whi78] Ewan A. Whitaker. Galileo’s lunar observations and the dating of the composition of “Sidereus Nuncius”. Journal for the History of Astronomy, 9(3):155–169, October 1978. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitney:1992:BRS

[Whi92] Charles A. Whitney. Book review: Solar physics and patterns of history: Exploring the Sun: Solar Science since Galileo. Journal for the History of Astronomy, 23(3):224– 226, August 1992. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitney:2000:BRS

[Whi00] Charles A. Whitney. Book review: Studying the galaxy: The Alchemy of the Heavens: Searching for Meaning in the Milky Way. Journal for the History of Astronomy, 31(2): 177–178, May 2000. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitaker:2002:BRO

[Whi02] Ewen A. Whitaker. Book review: Observation of the Moon, The Moon and the Western Imagination. Journal for the History of Astronomy, 33(1):82–83, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Whitson:2008:WHE

[Whi08] Bruce N. Whitson. William Herschel’s “Eccho Catch”. Journal for the History of Astronomy, 39(3):397–401, August 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1975:ERK

[Wil75] Curtis Wilson. Essay review: Kepler symposium in Weil Der Stadt: Internationales Kepler-Symposium, Weil der Stadt, 1971. Journal for the History of Astronomy, 6(1):58–62, February 1975. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williams:1979:FAM

[Wil79] M. E. W. Williams. Flamsteed’s alleged measurement of annual parallax for the Pole Star. Journal for the History of Astronomy, 10(2):102–116, June 1979. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1982:ERK

[Wil82]  Curtis Wilson.  Essay re-

view: Kepler’s “mysterium” in translation: Mysterium Cosmographicum — The Secret of the Universe. Journal for the History of Astronomy, 13(1):

60–62, February 1982. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1984:ERS

[Wil84]  Curtis Wilson.  Essay re-

view: The sources of Ptolemy’s parameters: The Origins of Ptolemy’s Astronomical Parameters. Journal for the History of Astronomy, 15(1):37– 47, February 1984. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williams:1986:BRB

[Wil86a] Mari E. W. Williams. Book review: a brief biography of Bessel: Friedrich Wilhelm Bessel. Journal for the History of Astronomy, 17(3):209– 210, August 1986. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1986:BRK

[Wil86b]  Curtis Wilson.  Book re-

view: Kepler’s ephemerides: Johannes Kepler Gesammelte Werke, XI, Part 1: Ephemerides novae motuum coelestium. Journal for the History of Astronomy, 17(1):63–65, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williamson:1987:BRE

[Wil87a] Ray A. Williamson. Book review: Earth and sky: Papers from the Northridge Conference on Archaeoastronomy.

Journal for the History of Astronomy, 18(11):S67–S69, February 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1987:BRN

[Wil87b] Curtis Wilson. Book review: Newton’s “De gravitatione”: De la Gravitation ou Les Fondements de la Mécanique Classique. Journal for the History of Astronomy, 18(4):291– 292, November 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1987:COH

[Wil87c] Curtis Wilson. Corrigendum: “On the origin of Horrocks’s lunar theory”. Journal for the History of Astronomy, 18 (3):172, August 1987. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Wil87d].

Wilson:1987:OHL

[Wil87d] Curtis Wilson. On the origin of Horrocks’s lunar theory. Journal for the History of Astronomy, 18(2):77–94, May 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Wil87c].

Wilson:1988:ERC

[Wil88] Curtis Wilson. Essay review: a conceptual history of statistics: The History of Statistics:

The Measurement of Uncertainty before 1900. Journal for the History of Astronomy, 19 (1):51–54, February 1988. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:1993:CCE

[Wil93] Curtis Wilson. Clairaut’s calculation of the Eighteenth-Century return of Halley’s Comet. Journal for the History of Astronomy, 24(1–2):1– 15, May/February 1993. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (elec-

tronic). URL http://jha.
sagepub.com/content/24/1-
2/1.full.pdf+html.

Wilson:1995:BRK

[Wil95] Curtis Wilson. Book review:

Kepler and Galileo: Johannes Kepler: Discussion avec Le Messager Céleste et Rapport sur l’Observation des Satellites de Jupiter. Journal for the History of Astronomy, 26(1): 86–87, February 1995. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williams:1996:DAS

[Wil96] Thomas R. Williams. The development of astronomy in the Southern United States, 1840–1914. Journal for the History of Astronomy, 27(1): 13–44, February 1996. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See errata [Ano97c].

Williams:1997:BRM

[Wil97] Tom Williams. Book review: Mars observed: The Planet Mars: a History of Observation & Discovery. Journal for the History of Astronomy, 28 (3):274–275, August 1997. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2000:BRL

[Wil00] Curtis Wilson. Book review: Life on Hven: On Tycho’s Island: Tycho Brahe and His Assistants, 1570–1601. Journal for the History of Astronomy, 31(4):352–353, November 2000. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2001:BRP

[Wil01]  Curtis Wilson.  Book re-

view: Principia in German, Die Mathematischen Prinzipien der Physik. Journal for the History of Astronomy, 32(4): 365–366, November 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williams:2002:BRN

[Wil02] Gareth Williams. Book review: Names of asteroids, Dictionary of Minor Planet Names. Journal for the History of Astronomy, 33(1):83–84, February 2002. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2003:BRD

[Wil03]  Curtis Wilson.  Book re-

view: D’Alembert’s lunar theory: Jean Le Rond D’Alembert, Oeuvres Complètes, Série I, Traités et mémoires mathématiques, 1736–1756, Premiers textes de mécanique céleste (1747– 1749). Journal for the History of Astronomy, 34(4):459–461, November 2003. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2004:BRP

[Wil04] Curtis Wilson. Book review: a pocket-sized history: The History of Astronomy: a Very Short Introduction. Journal for the History of Astronomy, 35 (3):359–360, August 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2005:BRM

[Wil05a]  Curtis Wilson.  Book re-

view: Mathematical astronomy in overview: From Eudoxus to Einstein: a History of Mathematical Astronomy. Journal for the History of Astronomy, 36 (2):238–240, May 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2005:BRJ

[Wil05b] David B. Wilson. Book review:

James Clerk Maxwell popularized: The Man Who Changed Everything: The Life of James Clerk Maxwell. Journal for the History of Astronomy, 36(3): 352–354, August 2005. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Williams:2007:BRF

[Wil07] Thomas R. Williams. Book review: Finding the Cepheid period-luminosity relationship: Miss Leavitt’s Stars: The Untold Story of the Woman Who Discovered How to Measure the Universe. Journal for the History of Astronomy, 38(2):244– 245, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2008:NLP

[Wil08a] Curtis Wilson. The nub of the lunar problem: From Euler to G. W. Hill. Journal for the History of Astronomy, 39(4):453–468, November 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See corrigendum [Wil09].

Wilson:2008:HCT

[Wil08b] Malcolm Wilson. Hippocrates of Chios’s theory of comets. Journal for the History of Astronomy, 39(2):141–160, May 2008. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2009:CNL

[Wil09] Curtis Wilson. Corrigendum:

[The nub of the lunar problem: From Euler to G. W. Hill]. Journal for the History of Astronomy, 40(1):108, February 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See [Wil08a].

Wilson:2011:ERT

[Wil11] Curtis Wilson. Essay review:

Tobias Mayer’s lunar theory: Between Theory and Observations: Tobias Mayer’s Explorations of Lunar Motion, 1751– 1755. Journal for the History of Astronomy, 42(4):521–526, November 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilding:2014:GIG

[Wil14a] Nick Wilding. Galileo’s idol:

Gianfrancesco Sagredo and the politics of knowledge. University of Chicago Press, Chicago, IL, USA and London, UK, 2014. ISBN 0-226-16697-X (hardcover). 200 pp. LCCN Q143.S17 W55 2014.

Willmoth:2014:BRH

[Wil14b] Frances Willmoth. Book review: Hevelius investigated: Johannes Hevelius and his World: Astronomer, Cartographer, Philosopher and Correspondent, ISBN 978-83-86062-94-2. Journal for the History of Astronomy, 45(2):261–262, May 2014. CODEN JHSAA2.

ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilding:2015:BRG

[Wil15] Nick Wilding. Book review: A Galileo facsimile: The Starry Messenger, Venice 1610 … with Symposium Proceedings. Galileo Galilei, edited by John W. Hessler and Daniel De Simone (Levenger Press, Delray Beach, FL, 2013). Pp. xiv + 209. $79. ISBN 978-1-929154-49-4. Journal for the History of Astronomy, 46(3):368–370, August 2015. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Winterburn:2006:BLM

[Win06] Emily Winterburn. British library microfilms of Herschel materials. Journal for the History of Astronomy, 37(3):343– 348, August 2006. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Winkler:2022:DHH

[Win22] Andreas Winkler. On the demotic-hieratic horoscopes from Athribis. Journal for the History of Astronomy, 53 (3):328–337, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286221109257.

Winkler:2025:TAE

[Win25] Andreas Winkler. Traversing the ancient Egyptian skies. Journal for the History of Astronomy, 56(1):115– 121, February 2025. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286241277908.

Wise:2006:BRJ

[Wis06] George Wise. Book review:

Joseph Henry at the Smithsonian: The Papers of Joseph Henry, X: January 1858– December 1865, the Smithsonian Years. Journal for the History of Astronomy, 37(1): 121–122, February 2006. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2018:EGC

[WKH18] Jarosław Włodarczyk, Richard L.

Kremer, and Howard C. Hughes. Edward Gresham, Copernican cosmology, and planetary occultations in pretelescopic astronomy. Journal for the History of Astronomy, 49(3):269–305, August 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:1987:OAA

[Wło87] Jarosław Włodarczyk. Observing with the armillary astrolabe. Journal for the History of Astronomy, 18(3):173–195, August 1987. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). With an appendix by Jerzy Dobrzycki.

Wlodarczyk:1990:NCP

[Wło90] Jarosław Włodarczyk. Notes on the compilation of Ptolemy’s catalogue of stars. Journal for the History of Astronomy, 21 (3):283–295, August 1990. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2007:SEO

[Wło07] Jarosław Włodarczyk. Solar eclipse observations in the time of Copernicus: Tradition or novelty? Journal for the History of Astronomy, 38(3):351– 364, August 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2010:BRS

[Wło10] Jarosław Włodarczyk. Book review: The supposed Moon of Venus: The Moon That Wasn’t: The Saga of Venus’ Spurious Satellite. Journal for the History of Astronomy, 41 (4):512–513, November 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2011:LMH

[Wło11] Jarosław Włodarczyk. Libration of the Moon, Hevelius’s theory, and its early reception in England. Journal for the History of Astronomy, 42(4): 495–519, November 2011. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2015:BRS

[Wło15] Jarosław Włodarczyk. Book review: Shakespeare and astronomy, yet again: The Science of Shakespeare: A New Look at the Playwright’s Universe. Dan Falk (St. Martin’s Press, New York, 2014). Pp. xv + 364. $27. ISBN 978-1-250-00877-0. Journal for the History of Astronomy, 46(3):375– 376, August 2015. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2022:TTW

[Wło22] Jarosław Włodarczyk. The torquetum (or turketum): Was it an observing instrument? Journal for the History of Astronomy, 53(3):266– 299, August 1, 2022. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221105653.

Wlodarczyk:2023:BRR

[Wło23] Jarosław Włodarczyk. Book review: Reducing meridian circle observations in positional astronomy: Ole Rømer’s Triduum. Fabricius Claus, Jørgensen Niels Therkel, and Tortzen Chr. Gorm. 3 vols (Society for Danish Language and Literature, Copenhagen, 2023). Pp. 819. 799 DKK. ISBN 978-87-7533-060-7. Journal for the History of Astronomy, 54(4):485–486, November 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231212665.

Wlodarczyk:2025:ELC

[Wło25] Jarosław Włodarczyk. Elias von Löwen (Crätschmair): an unrecognized pioneer of the research on optical libration of the Moon. Journal for the History of Astronomy, 56(1):72–86, February 2025. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286241240072.

Woolf:1980:SSP

[Woo80] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wes-ley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Woolley:1987:BRE

[Woo87] Richard Woolley. Book review: Eddington commemorated: Eddington: The Most Distinguished Astrophysicist of His Time. Journal for the History of Astronomy, 18(2):143– 144, May 1987. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wood:1988:BRM

[Woo88] John A. Wood. Book review: Making sense of meteorites: Cosmic Debris: Meteorites in History. Journal for the History of Astronomy, 19(4):287–288, November 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Weill-Parot:2022:MSA

[WP22] Nicolas Weill-Parot. Medieval structures of astrology. Journal for the History of Astronomy, 53(3):372–376, August 1, 2022. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221105652.

Wesemael:2008:WWC

[WR08] François Wesemael and René Racine. Why was the companion of Sirius not seen prior to 1862? Journal for the History of Astronomy, 39(2):161– 179, May 2008. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wright:2007:VAS

[Wri07] Ernie Wright. Visualization of the Almagest star catalogue. Journal for the History of Astronomy, 38(2):222–226, May 2007. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wang:2016:CRS

[WS16] Guangchao Wang and Xiaochun Sun. Chinese records of supernovae in A.D. 1408 revisited. Journal for the History of Astronomy, 47 (4):416–423, November 2016. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828616680965.

Wang:2019:CIB

[WS19] Guangchao Wang and Xiaochun Sun. A Chinese innovation based on Western methods: The doubleepicycle solar model in the Lixiang kaocheng, 1722. Journal for the History of Astronomy, 50(2):174–191, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619843431.

Wunsch:1999:AHA

[Wün99] J. Wünsch. The accuracy of Hevelius’s astrometric measurements. Journal for the History of Astronomy, 30(4): 391–406, November 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Weiner:1971:SMR

[WW71] Charles Weiner and Joan N. Warnow. Source materials for the recent history of astronomy and astrophysics: a checklist of manuscript collections in the United States. Journal for the History of Astronomy, 2(3): 210–218, October 1971. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Webster:1974:BRI

[WW74] Roderick Webster and Marjorie Webster. Book review: Instruments and instrumentmakers: Scientific Instruments of the Seventeenth and Eighteenth Centuries and Their Makers. Journal for the History of Astronomy, 5(2):140, June 1974. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Webster:1982:BRI

[WW82] Roderick S. Webster and Marjorie Webster. Book review: an instrument Collector’s guide: Antique Scientific Instruments. Journal for the History of Astronomy, 13(3):225, October 1982. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wlodarczyk:2019:BRS

[WW19] Jarosław Włodarczyk and Jarosław Włodarczyk. Book review: Selene’s many faces. Journal for the History of Astronomy, 50(4):493–495, November 1, 2019. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828619880411.

Winkler:2023:ZME

[WZR23] Andreas Winkler and Michael Zellmann-Rohrer. Zodiacs and monuments: an early pictorial “horoscope” from Egypt. Journal for the History of Astronomy, 54(2):125–152, May 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        abs/10.1177/00218286231159361.

Yano:1986:BRE

[Yan86]  Michio Yano.  Book re-

view: Early Chinese planetary astronomy: Les Premières Théories Planétaires Chinoises. Journal for the History of Astronomy, 17(2):148–150, May 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yavetz:1999:BRB

[Yav99] Ido Yavetz. Book review: Big astronomy: The Perfect Machine: Building the Palomar Telescope. Journal for the History of Astronomy, 30(2):175– 177, May 1999. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yazdi:2014:FAK

[Yaz14] Hamid-Reza Giahi Yazdi. The fragment of Al-Kindi’s lost treatise on observations of Halley’s Comet in A.D. 837. Journal for the History of Astronomy, 45(1):61–77, February 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yeomans:1988:BRH

[Yeo88] Donald K. Yeomans. Book review: a history of LPL: Lunar and Planetary Laboratory: Its Founding and Early Years. Journal for the History of Astronomy, 19(3):213–214, August 1988. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yeomans:1999:BRI

[Yeo99] Don Yeomans. Book review:

Imaging comets and meteors: Fire in the Sky: Comets and Meteors, the Decisive Centuries, in British Art and Science. Journal for the History of Astronomy, 30(3):327–328, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yeomans:2000:BRB

[Yeo00] Don Yeomans. Book review:

Briefing for the next Venus transit: June 8, 2004: Venus in Transit. Journal for the History of Astronomy, 31(4):361–362, November 2000. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yeomans:2002:BRC

[Yeo02] Don Yeomans. Book review:

The creator of astrogeology, Shoemaker by Levy: The Man Who Made an Impact. Journal for the History of Astronomy, 33(1):86–87, February 2002. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Yeomans:2007:BRW

[Yeo07] Donald Yeomans. Book review: When is a planet not a planet?: Is Pluto a Planet? A Historical Journey through the Solar System. Journal for the History of Astronomy, 38(2):250– 251, May 2007. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

York:2001:ACC

[Yor01] Tom J. York. An analysis of close conjunctions recorded in Ancient China. Journal for the History of Astronomy, 32(4): 337–344, November 2001. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Young:1986:IRA

[You86] M. Jane Young. The interrelationship of rock art and astronomical practice in the American Southwest. Journal for the History of Astronomy, 17(10): S43–S58, February 1986. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zedda:2004:OSN

[ZB04] Mauro Zedda and Juan Antonio Belmonte. On the orientations of Sardinian Nuraghes:

Some clues to their interpretation. Journal for the History of Astronomy, 35(1):85– 107, February 2004. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See erratum [Ano04b].

Zbikowska:1997:BRA

[Zbi97] Isabela Zbikowska. Book review: Aristarchus and his influence: Aristarch von Samos. Journal for the History of Astronomy, 28(4):351–352, November 1997. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zedda:2005:OAS

[Zed05] Mauro Zedda. Orientations of additional Sardinian dolmens. Journal for the History of Astronomy, 36(1):107–108, February 2005. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zeilik:1985:RFB

[Zei85a] Michael Zeilik. A reassessment of the Fajada Butte solar marker. Journal for the History of Astronomy, 16(9):S69–S85, February 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). See appraisal [SS86].

Zeilik:1985:SSS

[Zei85b] Michael Zeilik. Sun shrines and Sun symbols in the U.S. Southwest. Journal for the History of Astronomy, 16(9):S86–S96, February 1985. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zeilik:1986:EHP

[Zei86] Michael Zeilik. The ethnoastronomy of the historic Pueblos, II: Moon watching. Journal for the History of Astronomy, 17(10):S1–S22, February 1986. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zepeda:2019:BRW

[Zep19] Henry Zepeda. Book review: Walcher of Malvern: De lunationibus and De dracone: Study, Edition, Translation, and Commentary C. Philipp & E. Nothaft, (Brepols Publishers, Turnhout, 2017). Pp. xix + 346. EUR 90. ISBN 978-2-503-56876-8. Journal for the History of Astronomy, 50(1):112–114, February 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828619826752.

Zik:2012:ESK

[ZH12] Yaakov Zik and Giora Hon. The eccentricity of the Sun: Kepler’s novel method of calculation. Journal for the History of Astronomy, 43(2):181– 189, May 2012. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zedda:1996:OST

[ZHGM96] Mauro Zedda, Michael Hoskin, Renate Gralewski, and Giacobbe Manca. Orientations of 230 Sardinian Tombe Di Giganti. Journal for the History of Astronomy, 27(21):S33–S54, February 1996. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zieme:2023:GCL

[Zie23] Stefan Zieme. Gerard of Cremona’s Latin translation of the Almagest and the revision of tables. Journal for the History of Astronomy, 54 (1):3–33, February 1, 2023. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/00218286221140848.

Zhongwei:1999:COS

[ZJJ^{+}99] Hu Zhongwei, Ren Jinyi, Guo Jiuli, Hu Xianliang, and Xu Dengli. A Chinese observing site from remote antiquity. Journal for the History of Astronomy, 30(3):231–235, August 1999. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zsoldos:1999:NK

[ZL99] E. Zsoldos and Zs. Lévai. “Novae” over Kiskartal. Journal for the History of Astronomy, 30 (3):225–230, August 1999. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zotti:2020:NLM

[ZM20] Georg Zotti and S. Mohammad Mozaffari. New light on the main instrument of the Samarqand Observatory. Journal for the History of Astronomy, 51(3):255–271, August 1, 2020. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/0021828620943986.

Zrull:2021:WGW

[Zru21] Lindsay Smith Zrull. Women in glass: Women at the Harvard Observatory during the era of astronomical glass plate photography, 1875–1975. Journal for the History of Astronomy, 52(2):115–146, May 1, 2021. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

                                        journals.sagepub.com/doi/
                                        full/10.1177/00218286211000470.

Zsoldos:1994:TEV

[Zso94] E. Zsoldos. Three early variable star catalogues. Journal for the History of Astronomy, 25(2): 92–98, May 1994. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zsoldos:2010:SCN

[Zso10] Endre Zsoldos. Starting the classification: New and old stars, and sometimes comets, too. Journal for the History of Astronomy, 41(1):105–116, February 2010. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Zsoldos:2020:MAL

[Zso20] Endre Zsoldos. From mythology to astronomy: Lists and catalogues of variable stars. Journal for the History of Astronomy, 51(2):225– 244, May 1, 2020. CO-DEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://

journals.sagepub.com/doi/
full/10.1177/0021828620912872.

Zuidervaart:2019:BRI

[Zui19] Huib J. Zuidervaart. Book review: an Italian inspiration for the Dutch telescope? Journal for the History of Astronomy, 50(2):257–258, May 1, 2019. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/
full/10.1177/0021828619838887.

Zujdervaart:2011:BRO

[Zuj11] Huib J. Zujdervaart. Book review: The origins of the telescope: Galileo’s Telescope: The Instrument That Changed the World. Journal for the History of Astronomy, 42(1):114–116, February 2011. CODEN JH-SAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).